A Bibliography of Publications in Linear Algebra and its Applications: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.40

Title word cross-reference

(+, −) [EHJL97]. (−1, 0) [Sau90]. (0, 1) [CC92b, Chi93, CCR97, Gol97c, LMWZ90, Liu90]. (0, 1, 2) [Hou99]. (0, \frac{1}{2}, 1) [Bor95a]. (3, n − 3) [Gut98]. (A, B) [BZ95, BS99c, II98]. (A, B)^{t} [FPP98]. (A, B)t [CF97a]. (A^{T}XA, B^{T}XB) = (C, D) [CW93]. (J, J′) [XK94]. (J.J′) [HC98c].

(λ, G) [DdPC90]. (p, q) [Li94a]. ({AB \atop C0}) [CZ90]. (u, i) [Scu93]. ∗ [Har94]. −1 − \sqrt{2} [WN95]. 0 [Liv95]. \{0, \frac{1}{2}, 1\} [NP96]. 0, ±1 [CCF97]. 1 [HM98, Lev97, Liv95, VW90]. 2 [Ant98, AH97, Col96, FV97, MP97a, Oua91b, WW91]. 2K [Hil92]. 2 × 2 [Dai96, FI92, LS99, Mia91, Nov97, NO98]. 3 [Cos91]. 36 [Spe95]. 3D [KKY91]. 3 × 3 [Bon97, KRS97]. 4 [GGM96b]. 45, 694 [GM97b]. 4 × 4 [Nov97]. 5 [Hig95b, LS93b]. 5 × 5 [LM99]. 6 [Ste95]. < 9 [Bon97]. [tr(Q^{2p})]^{1/2p} [LS90a]. ^{∗} [BT97a, GMŠ98, Lav99a, MS93b]. ^{T} [Hig99]. A [MCK95]. A, B [Won90]. A_{T,S}^{(2)} [WMP98]. A^{2} [EL97]. A^{k} = J − I [WLH99]. A^{l} + A^{l+k} = J_{n} [WL98]. A^{T}A − B^{T}B = L^{T}DL [Atk93]. A_{1}XB_{1} = C_{1}, A_{2}XB_{2} = C_{2} [Mit90]. A_{2}(R, S) [Hou99]. AB [Wer94, Vis95]. A ∘ B [Vis95]. A ⊗ B [Wan97a]. AX + Y A = C [CW93]. AX − XB = C [SN91]. AX = B [Wu92a, WC96]. AX = XAX [HNRR99]. AXA^{T} + BY B^{T} = C [CW93]. AXB + CY D = F [XWZ98]. AXB − GXD = E [Hua96a]. B [EHL97, Mol99a]. B(H) [GMŠ98, Mol99b]. B^{−1} [EHL97]. B^{−}A^{−} [Wer94]. Z^{d} [Fri97a]. C [Daz94, HZ99, CN99b, CLLT98, LT91b, Mar96, Mer91a]. C[y] [Sch97]. C^{∗} [DS98]. C^{p} [EG90]. C_{0} [BLS97]. C_{p} [nCjH98a]. B(H) [Mol96]. S(X) = AX − XB [Hei97]. CS [PW94]. CS_{m} [Wu97]. cu + u^{∗}c = 2d [Dob99]. D [BC95b, BK93, FGL^{+}98, KMS96, NWZ98a, Hol96, HKL96, Rud98]. E′ [Dan97]. e(p, q) [TW99]. ℓ_{1} [Dax91]. ℓ_{∞} [AG94]. ℓ_{p} [MBI94]. ℓ_{p_{1},p_{2}} [Mer90]. EM [MR94]. f(X) = A [EU93]. f(Y) = B [EU93]. F = P G [Woo96]. F_{0} [Che90a]. f_{0} ≥ 2 [Jia96]. F_{4} [CW95b]. G [LT91a, Nie99b, CZ90, LC96]. G(c) [LT91a]. GF(2) [Cop93, BD91, GM97b]. GF(3) [Ell94]. GL(2, Z) [CT91, Hen97]. GL_{2}F [VW98]. GL_{n}(F) [Lev94, Lev97]. GL_{n}(Z) [Ish95]. GR [Tyr97]. H [Buo90b, Wan97b, EH98, KP95, Kol95, LLH^{+}98, Mol99a, hXyY98]. H(∞) [GTÖ96]. H^{2} [Eng95, FFT94]. H^{∞} [FFT94]. H_{2}|H_{∞} [PS94]. H_{∞} [Sto96, ST93, LWX99]. HR [BFW98]. J [ABGR94a, CR94b, CG97, HKL96]. J_{0}(z) − iJ_{1}(z) [IKF^{+}93]. J_{m}(z) [IKF^{+}93]. j_{pq} [FFK93a]. J_{q} [FKK95]. j_{qq} − J_{q} [FKK97]. K [Hil92, LLS97, BL90, Hol96, LS98, Li91a, Lim91b, MK98, MD92, Sha91a, Sha95, ST95, SEK99, Stu99]. K3 [Ste95]. K [FLL97]. L [BY95a, BCS94, Chr99]. l_{∞}^{(n)} [CS98a]. L^{1} [AG97]. l_{1} [Har96]. L_{2} [De 93]. l_{∞}^{n} [Ost99]. λ [EM91]. LDL^{T} [Sun92a]. LDM^{t} [Haf95]. ≤ 4 [PZ90b]. LU [BH92a, HLY92, JOv95, Par98a, Yip96, HLP97, BH91]. LULU [Str97a]. M [Hua91, LC91a, Väl91, BCT93, CC96, CC92a, CKN96, CN97, Els97, Fie98, Fri92, mGhW92, GhW96, Goe96, Hog98a, Hog98b, JS96, JS99b, KN95, KN99, Li92b, LZ96, Mar95a, MNS93, McD94, MNST95, MM93a, NV95a, Neu98, O’N93, Peñ95, Väl93, WZ99, Yip96, CC94, HY96, IKF^{+}93, Ote90, Rea97]. M_{n} [PŠ98a]. M_{n}(R) [Mar99b]. M_{n}(S) [YZ95]. M_{n}Z [Qin96]. C[z, z^{−1}] [– DL95]. F_{q} [Fle93]. GL_{n}(K) [BN99]. P [SG95]. P_{∗} [Väl96b]. P_{0} [GC95, SG95]. P_{3} [San96a]. R^{2n} [Yiu93]. Z [HHR93]. min [BCG91]. m × n [BI92, Tia98]. N [OPR91, PR90, Gał91, Hav93, Kit93, WM90, Yiu93]. ≠ 2 [YW90]. N ≡ 1 mod 4 [MCK95]. n × n [Chi93, Dom96, Ma93]. O(n^{3}L) [Miz91]. O(n^{ρ}L) [Miz91]. ω [Dan91b, FM93c]. Ω_{3}(y ≺ x) [CS98b]. P [FHP90, MP90, Szu90a, JOv95, MS98, Mor97, Woo96, EU93, GHN97, GHN99, Gam97, HN93b, HNT96, rKMMR95, KPS91, Par96, Pee95, WZCZ93]. p(A) = 0 [Sem92]. P (X, Y) [Per98]. P_{0} [EK93]. P_{c} [CF96]. PG(3, 2) [Jeu95]. PG(M_{n,n}(GF(q))) [Coo97]. Φ [MLS98, Hon91]. ±1 [NR97]. Pol(n, m) [ABM93]. PSL(n, F) [Lev93]. p × q × 2 [TK99]. Q [Ebi95, BL94a, EU93, Lal98, Pay95]. q = 2e [Pay95]. Q_{k}^{∗} [LLS97]. Q^{+} [KR91]. Q_{6}^{t} [LC96]. Q_{0} [MPS97, PSS97]. Q_{6} [LC96]. QR [ABD92, BOC95, Bar94, Bat94, Gla98, Kni95, MS93c, MVV93, Sun95, Tyr97, Wat96b]. r [Mic93]. R^{2n} [WM90]. R^{6} [Mes97a]. R^{n} [AGK^{+}98, CCK92, DK92, JB92, Lav98a, MLS91, MLS95]. R^{p} [Le 94]. rank(I − AZ) = 1 [BW91]. ρ [NO98]. S [KvL93, HZ99, LMS95, Mit99, jWySmJzZ95]. S(B^{+}) [KS94c]. S_{p} [OS94]. sin Θ [Xie97b]. SL(3, 3) [CW95b]. SL(n, Z) [GT93]. SL(r + s, Z) [Cha96b]. SL_{n}(F) [Lev97]. SL_{n}(Z) [Ish95]. SN [HKS96a]. Sp(2n, q) [EM91]. Sp_{n}(Z) [Koh97]. SR [BF98, BFW98, BFW99, Cha98]. SU(n)/SO(n) [Pap90]. SV D [MVV93]. SZ [BFW99]. T [Oua91c, Oua91a, BS95b, HHL92]. T (H^{1/n}T)^{n} = K [BF92a]. T_{2}n [Ma93]. τ [Hol99]. tr(AX^{2}) [Kur97]. tr[(A^{−1} − B^{−1})(A − B)] ≤ 0 [Won90]. △ [BD91]. T → ST S^{−1} + S^{−1}T S [LO93]. u [Scu91]. ULU [Tof97]. URV [MVV93]. V [Rus94]. w [Mos96, Xu99]. ˆ3McL [LS93b]. X + A ∗ X − 2A = I [IEs98]. X + A^{∗}X^{−1}A = Q [ERR93].

X + A^{T}X^{−1}A = I [Eng93, ZX96]. x − axb = w [CN92]. X = p(A) [EU93]. X = Q + NX^{−1}N [FL96]. x^{[3]} = ω(x)^{3}x [MV95, MV95]. ξ [Pat99b]. XX^{tr} = A [Ple95]. Y = q(B) [EU93]. Z [FM92, EK93, Smi90, FJSv98, FN97, NV95b, Nab97, Smi95]. Z^{n} [RGS98]. Z_{4} [CM95]. Z_{l} [Shi99b].

-algebras [Oua91c, Oua91a, ABM93, DS98]. -algèbres [Oua91c, Oua91a, ABM93]. -antinorms [Mer91a]. -approximate [MBI94]. -Arcs [ST95]. -Bezoutians [Rus94]. -Clan [Pay95]. -classes [OS94]. -Competition [rKMMR95]. -Congruence [HHL92]. -Contraction [NO98]. -convexity [FHP90, MLS98]. -core-nilpotent [Har94]. -critical [DdPC90]. -Cubes [HKL96]. -Cyclic [AH97, GHN97, HN93b, HNT96, GHN99, KPS91, Lev97, BS99c]. -D [Gał91]. -diagonally [FM93c]. -diagrams [Mit99]. -Dimensional [GM97b, Rud98]. -eigenvalue [FLL97]. -Elementary [FFK93a, FKK95, FKK97]. -EP [MK98]. -equivalence [Hon91]. -factorizations [BH91]. -functions [Pat99b]. -Generalized [LLS97]. -Groups [Par96]. -Invariant [CF97a, FPP98, LT91a, BZ95, II98, WZ99]. -Inverse [LMS95]. -inverses [CZ90]. -Isometric [WZCZ93]. -iteration [Miz91]. -Lanczos [FLL97]. -linear [CM95]. -lossless [XK94, CR94b, CG97, HC98c]. -Major [MD92]. -majorization [Nie99b]. -matrices [Buo90b, Che90a, EK93, Hol99, Hou99, Hua91, KP95, OPR91, PR90, Smi90, Szu90a, Väl91, Wan97b, BY95a, CF96, CC92a, CN97, Dan97, Els97, Fri92, FN97, mGhW92, GhW96, Goe96, JS96, KN95, LLH^{+}98, MNS93, MM93a, NV95b, NV95a, Nab97, O’N93, Peñ95, Smi95, Yip96, BCS94, EH98, FJSv98, Fie98, JOv95, KN99, MS98, McD94, MPS97, Neu98, Väl93, Väl96b]. -matrix [LC91a, MP90, BCT93, CC96, CKN96, CN97, Ebi95, Kol95, Li92b, LZ96, MNST95, Hog98a, Hog98b, JS99b, hXyY98]. -matroids [BD91]. -Minor [MD92]. -modules [HHR93]. -Multigraphs [Mic93]. -Norm [CLLT98, Har96, WZCZ93, Ant98]. -Norms [Mar96, Mer91a]. -numerical [CN99b, Daz94, Li91a]. -operation [Miz91]. -Operators [Mar95a]. -Optimal [KMS96, NWZ98a]. -Optimization [MCK95]. -pairs [Mor97]. -partite [BS95b]. -Permanent [BL94a, Lal98]. -Planes [Yiu93, WM90]. -potent [SEK99, Stu99]. -Properties [SG95]. -radii [LT91a]. -Rank [Pee95]. -Row [FM97a]. -semigroup [GMŠ98]. -semirings [BT97a]. -Simplices [HKL96]. -Spaces [Scu93, Scu91]. -Spectral [Mar96, LT91b, CR94b]. -Stability [BK93]. -stable [FGL^{+}98]. -stage [CC94]. -Step [HY96]. -symétrique [BC95b]. -symmetric [BC95b, LT91b]. -systems [KvL93]. -tessellation [LC96]. -Toeplitz [MP97a]. -transvections [EM91]. -uniqueness [Xu99]. -unitary [ABGR94a]. -values [Gam97]. -Variate [Hol96]. -Way [LS98].

119 [BPS95]. 121 [MV93]. 126 [SJ95]. 127 [NF95]. 132 [BT95b]. 134 [Chi96b]. 145 [GH95a]. 153 [MLS95]. 177 [KP95]. 179 [Mar95b]. 184 [Gil95, Mat95]. 187 [Spi95]. 199 [Ikr95, Mos95]. 1998 [Uhl98].

2 [Gov94, Kiw95b, McC93, MUV^{+}97].

2-Chordal [McC93]. 2-Norm

[Kiw95b, MUV^{+}97]. 2-Toeplitz [Gov94]. 20-Volume [Ano97]. 201 [Gro95a]. 212 [KR96a, Li96]. 212/213 [KR96a, Li96]. 213 [KR96a, Li96]. 215 [Ano96b]. 237 [Mit97]. 237/238 [Mit97]. 238 [Mit97]. 240 [Bha97a]. 2D [For90b, FV96].

3rd [ORS94].

5 [Rez95]. 5-Designs [Rez95].

6-8 [BS92a].

75 [Tho96].

81b [Bon94]. 87k [Tho96].

90g [MV93]. 90h [BPS95]. 91d [BT95b, Chi96b, NF95]. 91f [SJ95]. 92d [GH95a]. 92e [MLS95]. 93k [KP95, Mar95b]. 94a [Gil95]. 94b [Mat95]. 94m [Spi95]. 95h [Li96]. 95i [Ikr95]. 95k [KR96a]. 96i [Ano96b]. 97a [Mit97]. 97b [Bha97a].

= [Eva95].

A. [Far97, ZG94]. Abaffy [Saa92]. Abelian [BWS95, BT96, DIO97, HWZ90, HWZ92, Mos96, TW99, dRWZ90]. abéliens [Mos96]. ABS [Saa92]. Absolute [BKL97, Bha94a, Bha95, Wat90a]. absolute-value [Wat90a]. absolutely [Osa93]. Abwandlungen [Pus90]. Academic [Bon94]. Acausal [PP94]. Accelerated [Kiw97]. Acceleration [LY93c, Dax90c, Dax90a, HN90, LC91b]. Accumulated [ST96a]. accuracy [D^{+}99]. Accurate [DG93]. Acknowledgements [Ano98a]. Act [AP94a, Hil92]. Action [Mus94, Pay95]. Actions [CS98a, PR94a, Ost99]. Acyclic [DG93, Fie97]. adaptation [CG98]. adaptive [Che98, Mai91, SS91]. Addendum [AB91, DL01]. Addition [BBD^{+}97, FM93b]. Additional [Has96]. Additive [CZ96, OŠ93, Prz95, CP90, DS98, OŠ91]. Additivity [BŠS96, HOŠS96, Sem92]. ADI [Kwa97, Sta93]. Adjoint [ABBN94, RZ97, Kar99, LN91, Rob98, vdMRR99]. adjoints [Del99]. adjugate [Ste98]. adjustment [AS99]. admissibilité [Téc97]. admissibility [Téc97]. Admissible [Lis92]. admit [Fri91a]. Admittance [SSS97]. Advanced [BB92, TZ98]. Affine [Fri97b, GT96b, Kes92, Man92, TY97, Uhr92, YC92, Din93b, ER90, RGS99, Tod91]. affine-scaling [Tod91]. After [BBD^{+}97, BZ98]. Aggregation [Kol98]. ahead [FZ93]. Aitken [Far97]. Alarm [BF92b]. Alegebras [Cla92b]. Algebra [Ano96b, ABCP97, BPS95, BP95, BHR91, BB92, BLS97, Bha97a, BO93, BP92c, BT95b, Cec92, Cec96, CK97, Chi96b, Coe93, Da 94, Dat94, DD97a, DD97b, DIO97, Dia96b, DSW96, Far96, Gil95, GM93, GH95a, Gro95a, Ikr95, KP95, KLN96, KR96a, Li96, Mal93, Mar95b, MLS95, Mat95, Mit97, MP98d, Mos95, Nie93a, NF95, Sei96, Sob96, Spi95, SJ95, SSHS92, Tab93, Tho96, Tho92, Vro93, Wu94, YM94, Yas96, Zha97b, LAA68, BS92a, BT99, BCG91, Cap98b, Cec90, CJY91, Che91, EvdD99, Har99a, Hua99, KT91, KS94c, Laf98b, Lei99b, Mar91, MR98b, MO91, OY94, ORvE99, Var96]. Algebraic [BW97, Bai97b, BC95a, Coh96, Gee93b, GKS94, Had94, HC98b, Jon95b, KEK97, KU93, Kar94, LRT97, LV91a, LY93c, MV93, MM96a, MM96b, RR96a, RW94, Sch95, Sto96, gS97a, US92, Vro93, Wim94b, XL95, ZL98, dGK^{+}92, BH90, BGM98, Bol98, DC93, FP99, Fie90b, FG99, Fli94, FY98, Gur94, HNR90, Ikr99, Jód91, KTK99, KSP98, Kuc91, KM90, Li95b, LL93b, MM98, NT91a, Rum91, Sch91, Ste90, SS98, gS98b, Vin90, Wim91]. algebraic-analytic [HNR90]. Algebras [AL96, AG93, AM95, BVB93, BBB95, BB96, BZ96, Boa97, BT96, Boz95, Chi95, Cor93, CG93b, CG94b, Cos95, DZ95, DG95b, FBO97, FRLM97, FLM94, GM93, GGM96a, GGM96b, Gra96, HKH97, HWZ92, Jam94, Kop96b, KR96b, Kye95, LS93a, Le 98, MP98d, Mrč95, Sob96, Spi94, Tur92, ZD92, Ala99, ABM93, AMM99, Bae90, Bha99, BM95, CD99, CM96, Cos91, DR91a, DS98, FNW91, GJMNV98, HP91, Hir90, HWZ90, Lam90, MMO91, MV95, MS99a, MO95, MVKR99, Osa91, Osa93, Oua91b, Oua91c, Oua91a, Oua96, PZ90a, PZ90b, RS90, TW99, TB97, dRWZ90]. algèbre [MO91, Var96]. Algèbres [CM96, MO97, ABM93, BM95, EK99, MMO91, MV95, MO95, Oua91b, Oua91c, Oua91a]. Algébrique [DC93, Fli94]. AlgLat [DC93]. Algorithm [AH97, BOC95, Bar96, BK97a, Bat94, BLL92, BO93, BCU94, BF92b, DS91, Fan93, FWWJL92, FZ93, FM97b, GN93, Gem96, Gem97a, Gen96a, Gla98, Hon92, HO93, Jia98a, Jia92, KY93, Kat94, LA94, Mal93, MS93c, MR94, MVV93, NP92a, PPY95, RS93, Tur91, Tyr97, WL93, Wat96b, AR99, Ans91, BF98, BRMM91, BKO99, Cao97, Che98, Cop93, Dai99, Din93a, DPPR90, Erx90, FLL97, Fre91a, GHN99, GJMNV98, GW99, Han90b, Hof98b, JM99, Jia99, JTU98, JL98b, LL93b, Mel99, MS99b, O’L91, ORvE99, Ple98, PW90, SS90, SS91, SL97, VB97b, Wan91, Xu91, CG92a]. Algorithms [DdlP96, EG97, Fay94, FHR93, Fle93, GTVV96, GK93, HB98, Jia97b, LV97, Lan96, Mes94, Mes97b, Mey97, Pap92, PYH91, Rja94, Saa92, SK95, Sun96b, BFW98, BFW99, BTT91, CG98, DA90, Egg90, KKS90, KN91, Miz91, OY94, PST98, RV98, Ruh94, Sem91, SP99, TZ91, Van91b, WE91]. All-Pass [Mic97]. Allow [JMS95, Wat96a, EHL97]. Allowable [AL95b]. Almost [AP96a, BT96, FK94, HR93a, HWH95, KSW99, Kir97a, Mer91b, OPR91, Tof97, BKM91, Har91, Ols98, Ri99]. Alston [Ano95a]. alternate [PI90]. Alternating [CKN96]. Alternative [CMLS95, HK97b, Kal96, GPM99b]. Ambiguity [ZX99]. Among [KS94b]. Analog [HM93]. Analogies [FP97a]. Analogue [BR95]. analyse [RSDT97]. analyses [CP98a]. Analysis [BW96, Bt97c, BdV92, CWX96, DN97, Eij91, KPGP97, LS98, Mey97, NP92b, Pan93, PN91, Put98, Rei94, Sla98, SY95, gS96c, gS97b, Uhl92, YT92, ZZ98, AFK90, CP99a, Din93b, GZ94, GM98b, GZ98b, Har99a, HS90b, Kri99, MN91, Not91, RSDT97, See99, Sid98, Sun90a, gS98b, TK99, WP98, Li98a]. Analytic [CG97, Els95, GH92, GKv93, Ala99, Blo99, GV98, HNR90, NO90]. Analytic-Geometric [Els95]. Analytical [MM93c, Nie98b]. Analyticity [TFV94]. Analyzing [McE94]. and/or [Fun96a, Has96]. Ando [FFK93b, Fur95]. Ando-Hiai [Fur95]. angle [Obe91]. Angles [JB92, MBI96, Wim99]. Announcement [Ano95a]. ANOVA [AU99]. Answer [BT95a]. anti [KBDZ98]. anti-majorizing [KBDZ98]. Antidiagonals [LM94b]. Antieigenvalues [Gus94]. antimatroids [Die90]. Antinegative [BP92b]. Antinorms [MD92, Mer90, Mer91a]. antitonicity [BNS90]. Any [IKF^{+}93]. AOR [Jam91b, JZ91, Son97, Wan91, Wan98b]. Aperiodicities [Uhr96]. appearing [FFT90]. Appell [Mat93a]. Appl [Ano96b, BPS95, Bha97a, BT95b, Gil95, GH95a, Gro95a, Ikr95, KP95, KR96a, Li96, Mar95b, MLS95, Mat95, MV93, Mit97, Mos95, NF95, Spi95, SJ95, Tho96]. Appl. [Chi96b]. Application [AD96a, BB92, BM98b, EG97, FFK93b, Gra96, GMJ95, HB90, IKF^{+}93, IAMC96, Jia94, Koš96, Kur97, Lag92, Li97b, Met97, PD97, Rol96, SW95, Wan94, Wat95b, BM90, Cow90, Cra99, EKN90, Gem98, GMŠ98, HR98b, HR98c, HR90, LS90a, Rei91, Ste90, Str98, RG99]. Applications [AGR92, AS97, BM94c, BD95, BdV92, CZ96, CC92a, Chi92a, Chi92b, Cox95, DHL97, DN97, DZ98, Dia96b, – DL94, EF95, FB96, FH93, GP94a, Gem96, GKK96, Gol93, GS93, GT97c, HRR92, HM95b, Hua96c, Huy96, Jen97, KO97a, Kar97, KH95b, Kir97c, KNS98, KW95, LL93a, cL93, LVBL98, Lom99, McD94, Mil95, Neu97a, OSiA96, PPS96, Tab93, US92, ABD92, LAA68, BP90c, BA– DR91, Bha91, BG98a, Che90b, Dah90, EG90, FRS99, FY98, HP90, Hun90, RR90a, Rus94, Tia98, Til98a, Tyr91, Wan97a, WI99, WW97, WMP98, hXyY98, Gro95b, Pre98, Bon94, Kir98]. Applied [BL93b, ŽWB93, Wan97b]. Applying [San91]. Appreciation [GPPS94, PS96b]. Approach [Ant94, CG92a, Coh96, DR93, Els94, FH94b, HFK94, IW93, Jan97a, KO97b, Oar96, Rea97, Rei94, Rud96, Sou94, Tyr96, Wei94, WY96a, AFK90, App99, Bak99, Bro90, Chu91, Del90, Fuh91, Gem98, GL91, HNR90, JS99a, RSDT97, Sch97, Sid99, Vav91]. Approche [RSDT97]. approximants [BHK91]. Approximate [HTP96, KLN96, LS92b, MBI95, MM93c, GM98a, Jia99, MN91, MBI94, Sha98b]. Approximating [SY98, LST98]. Approximation [Abd95, Ant98, Chu98, De 93, Mic97, Nie97b, SO94a, Zie96, vD94, Che99, GZ98b, RR99b, Téc97, XB90]. Approximations [Alt94, Chi96d, Dah90, GTZ97, HL96a, Zie93b, Cap98b, Fuh91, KN98]. Arc [Kir95]. Arcs [ST95, Die90]. Argument [Chi92b, MP98c, Zie93a]. Arguments [Chi92a, Mat93a]. Arise [LM97]. Arising [BL93b, Gem97a, Jon95b, RS97, SS95b, HC98a, KLS90, Sha98b]. Arithmetic [AMPV97, Hor95, Kit92, Kra98, Mat93b, MP96b, Fur94, LV91b, Mat95]. arithmetic-geometric [Fur94]. Arithmetic-Geometric-Harmonic [Mat93b, Mat95]. Arithmetic-Geometric-Harmonic-Mean [AMPV97]. Arithmetic-Geometric-Mean [Kit92]. Arithmetic-Harmonic-Mean [Kra98]. Arithmetic-Mean-Harmonic-Mean [MP96b]. arithmetical [Hon98]. ARMAX [KS96]. Arnoldi [DS91, Jia98a, Jia99, Kni99, ZZ96]. Arnoldi’s [Jia97b]. Arov [FKK97]. Arov-normalized [FKK97]. Arrangement [BBLP95]. Arrangements [CHL95]. Array [HM96, LS98]. Arrays [Rud98, Col96, TK99]. artiniennes [EK99].

Artzi [Ion98]. Artzi-Gohberg [Ion98].

Ascent [KS91]. Aspects

[CG94a, D’A97, FOV97, BGM98, Laf98b, Li94b, Li96, Téc97]. Assessing [LA94]. Assigned [KPRV97]. Assigning [LMZZ98]. Assignment [BP95, BCU94, DS91, DER97, FH94b, HRW92, AG98, ACWY99, BLL94, CH99c, HDTR99, LN98, RZ98]. assignments [DD90]. Associated [BPY92, BI92, BS97, BL95, DS91, Fri92, Goe96, HN93b, HX92, HS97b, KN95, Kno92, Kuz97, NU97, SSS97, WAS97, Fer99, FFK91, Hon98, HM90b, JZ99, KSP98, Le 94, Van91b]. Association [GL95, Hig95b]. Associative [BVB93, Mir97, Hua99, Oua96]. associatives [Oua96]. associée [Le 94]. assumption [Cra99]. asymmetry [GZ98b]. Asymptotic [GKK96, Hua96c, Lam93, McL90, Neu96a, PV94, SWP97, Sil97a, Tre99a, Tre99b, Fre91b, Haf95, NW90, PI90]. Asymptotically [GV92]. asymptotics [Ger98]. Asynchronized [EN93]. Asynchronous [BKM91, Str97b, Su96, EKN90, KBŠ90, Pot98, SBKK98]. Attainable [GV92]. attained [SW98a]. Audits [HL95]. August [BS92a, BHR91, DLdO94]. Author [Ano97, Ano95b]. auto [Del99]. auto-adjoints [Del99]. automata [LV91a]. Automorphism [Coe93, Kop96b]. Automorphisms [Ala99, Jøn95a, KR92, Rei93, RS93, ABGR94a, CX98a, MS99a]. Autoregressive [WR94]. Avoiding [Cao97]. Award [Ano95a, Ano98c]. AX=B [Gro99a]. Axelsson [Str96]. Axis [Bol97, CALM97].

B [YM94, PS96b]. Backward

[gS97a, gS99b]. badly [HHR93]. Baer [BT97a]. Balanced [HF94, HF96, LH97, LP94, LP95b, MM93b, PP94, Wei94, Fuh91, HO98, HPWZ98, Loe90]. Balancing [Ach93, Hod93, CZ91, Sou91]. Ball [Mar94b, Mar94c, Mes97a]. Balls [Fri96, Mar96, Mar97, NW92]. Banach [Ala99, BC91b, Beh94, CJY91, Che91, HR93b, KR96b, Wu94]. Band [FW96, KW95, Lin92, Ng97a, Dan91a, GKW91, RBRF91]. Band-Diagonal [FW96]. Banded [GKL96, GS93, LP97, Haf95]. bands [Mar99b]. Bapat [YM94]. Baric [AL96, BVB93, CG93b, CG94b]. Barnett [GV96, Gro95b]. barrier [Fre91a]. Barycentric [ACE95]. base [LHE94]. Based [Bai97a, BK96b, BK96c, CNW96, CG92a, FC97, FH94b, Jia97b, Jia98a, LV97, MH92, Oar96, Bap99, Che98, CL91, Di 98, Gad90, GRO98a, GM98a, KV98, RR98, Ng97b]. Bases [Bar96, CGK96, CF97a, ENN98, HX92, HPW94, Kar94, KM95, Koš96, Kuz97, MH97, BRS91, HNR90]. Basic [HM92b, MBI95, Shi99b]. Basis [BKS93, Lam93, Sun94, Ple98, TZ91]. Bateman [Roe97]. Bauer [Pry92]. Bayesian [ZD92]. BCH [AP99a]. be [Mol96, van98]. Behavior [BG94, Chi96a, GT96a, GKK96, HK93b, PV94, SV96, PI90, TG97]. Behavioral [Wei94]. Behrens [Gam97]. Bellman [Xu95]. Belong [Hua96b]. Ben [Ion98]. Beneficial [PN92]. Berezin [Eng95]. Bergstrom [FM96b]. Berlekamp [Fle93, Lan96]. Bernstein [Bae90, BM95, Cor93, GM93, GGM96a, GGM96b, Gut98, HP91, MMO91, MO95, Oua91b, Oua91c, TB97, Var96]. Bessel [IKF^{+}93, Pav99]. Best [CW95a, EV93, MUV^{+}97, BP90b, Dan91b, RR99b]. Best-Conditioned [CW95a]. Betti [Par96]. Betti-Mathieu [Par96]. Between [AH94a, DK92, Fle93, GT96b, Her93, HLM96, HC98b, JB92, JT95, Kye95, Lau98, Lis92, MBI96, Thi95, Wim94b, Yu95, AH90, BFW98, Han90a, Hou99, KM91, LT91b, MOPvdD99, Obe91]. beyond [Han98]. Bezout [PV93]. Bezoutian [MH97, Wim90]. Bezoutians [Rus94, CZ95, GS95, Gem97b, GS90, Gou94, Hei95b, LR99, Ros93a, Ros93b, BG98a, Gem98, HR98b]. Bhatia [Li98a, Loe92]. BI [Cao97]. BI-CGSTAB [Cao97]. bialternate [GS99a]. biased [AU99, Chi99]. bibliography [PS96b]. bicirculant [Hla99]. Biclique [BS95b, GWS99]. bidiagonal [JOvdD99]. Bidiagonalization [PV95]. bidiagonals [Par98b]. Bidisc [Pat97c, Blo99]. Biehler [HDB99]. Bifurcation [GPM99a]. Bifurcations [WmL97]. Bijective [Zen93]. Bilinear [BVB93, GM95a, Leo94, Wat97b, HP91, Kov99, TB97, von91]. bimatrix [JJPT90]. bimatroids [Bap94b]. Binary [LS95a]. Binet [HM90a, Zen93]. binomial [RR99a]. binormal [BHK91]. Biography [GPPS94, Gro94, Gro95a, PS96b]. Biorthogonal [FZ93, KT95a, SV98]. Bipartite [Mur95a, GP90, JR99]. Birkhoff [LS93a]. Bitangential [BK96a]. Bivalent [CGB95]. bivariate [Gem98]. Björck [BKO99]. Black [MH96]. Block [Adu98, ABL98, AP96a, BZ95, BLJW95, BZ94, BG94, BM98b, BNP93, Bom96, BCJ96, BJW98, CHS92, Dai96, EGL96, FH96a, FP97a, FM97b, FK93, GTVV96, Gem97b, GK93, Gor92, GTÖ96, HLWT96, HJ92, Her92a, HM92b, IW96, Iwa98, Jia98a, Jia94, Juh96, Kam94, KKY91, KT95a, Kiw95a, KP92b, KNW91, KPS91, Lab92, LZ96, LW98b, MS93c, MM93c, Ng97a, Rja94, RR97, RR92b, Ser98b, Shm94, Sil96, Sil97d, SG96, Tak96, Til98b, Woź97, YP96, Yip96, da 98, BKM91, Bol98, CU99, CH99a, Cop93, DHS94a, ENV91, GV98, GW99, Guo91, HO98, Hig99, Jam91b, JM99, JS91, KSW99, KBŠ90, KP95, LST98, MN91, Mia91, Not91, OY94, RR98, Tia98, VB97b, hXyY98, Yan98, Zab94]. block-balanced [HO98]. Block-Form [Kam94]. Block-Iterative [Kiw95a, BKM91, KBŠ90]. Block-Matrix [HM92b]. Block-Toeplitz [BJW98, Ng97a]. Block-Tridiagonal [RR92b]. Blocks [BD93, GKv98, Gol97c, ME98, dS97, da 97, Cao98b, DPPR90, GS90, GKW91, Sil91b, da 91]. Blowing [Hur90]. BLUE [LPW92, NP92c, Sea97, WY96b]. BLUP [Sea97]. Blurring [Wat96b]. bond [HDTR99, RSDT97, RSDT97]. Book [Ano94b, Bar98]. book-graph [Bar98]. books [PS96b]. Boolean [BJP97, BP90d, BH97, CZ97, Cho93a, GKP93, GKP95, Hua90, Hua93, Hua96b, Jia96, KP92a, Liu99a, TZ96, Väl93, Zha91, Zho99]. bordered [CZ90]. Bordering [PR96]. Borel [Mal96]. Bott [Che90b]. Bottleneck [ABCP97, Cec92, Cec96, CK97, Cec90]. Bottleneck/Fuzzy [Cec96]. Boulder [LS95b]. Bound [Dru94a, GKP95, Hol95, HP92, JL97, Li93a, LZ97a, LT97b, Mae96, MTS96, MV98b, McE93, MUV^{+}97, Neu96b, Sch96b, She96a, Tuc96, Xu97b, BTT91, Chi90, HLP97, HKM98, JZ91, LC91a, Li97d, SV99, TS99, YhG97]. Boundary [BL93b, CN99b, KO97a, MP97b, SS95a, BM98c]. Bounded [CK97, HF96, LR94b, OŠ96a, SS96, CP98b, FL99, HOR98, Lav98b, MS91a, Nie90, SNC98]. bounded-shape [HOR98]. Boundedness [Chi96b]. Bounding [Roh97]. Bounds [AL98, An98, BH94b, Bar94, BEK90, Bha94a, Bha95, Cao98a, CZ97, CL91, CHR97b, Drm96, Els97, Fri92, GPRL94, HN93d, Hon98, Hor95, Hor97, Jan97b, Jen97, Jia94, JZ97, JS98a, KH95a, Kir97c, KNS98, Kol93, Kol95, Kol98, KT96b, Kra94, Kuz97, Li97c, Lin98c, LPW92, Liu96, LS97, Loe92, Lop94, Man92, MST94, MV97, MM93c, Neu95, O’L90, O’N93, OK96, Pal90, Roh95, Sch92, Sha91a, SJ96, Smi90, Son92, Sun92a, Sun95, Szy92, TK98, VSB90, Xie97a, Fri98b, HS99, Kni99, KSP98, Mal98, MP90, Nab99, Phi91, SW98a]. Bowtie [AH97]. Box [MH96, JS90b]. Boyle [AD97]. Bracket [Lag92]. Branch [KR90]. Branch-point [KR90]. Branching [Sch96b]. Brauer [Chi91, ZG94]. breakdown [Cao97]. Bridge [FC97]. Bridges [AS97]. Brief [GPPS94, Lun95, PS96b]. Brigham [BHR91]. Brownian [AS97, FC97]. Broyden [Oga99]. Brualdi [Sha92, Kir96a, Kir97c]. Bruhat [Tyr97]. Brunovsky [CF97a, GV98, HAPSG96]. Brunowsky [FGP97]. BSSOR [KKY91]. bundles [Lom99]. butterfly [BF98, BFW99]. buys [Gay91]. BVPs [SS95b].

C [Far97]. C. [Bha97a]. Cache [Mir97]. Calculation [Jan97b, RR99b]. calculus [Cra99]. Call [Ano96a, Nie98a]. Calyampudi [PS96b]. Camion [BKS93]. Campbell [SSS97]. Campbell-Youla [SSS97]. Can [EL97, Mus94, BCCR99, van98]. cancellation [Lev91, Van90]. Canonical [BPY92, BS95a, DS94a, DN97, DE95, DR91b, FH96a, FRLM97, FL98, HAPSG96, HM95b, HM99, Hua96c, KPGP97, LW96, LMX99, Van97b, Wim99, YT92, BRMM91, DD90, GZ94, HO98, HPWZ98, Hon90, Hon91, HH91, Sha91b, SAZ91, SW91c, WP98, ZJ91]. Carathéodory [CG94a, nCjH98a, FKK95, FKK98]. cardinalities [Zho99]. Career [Lun95]. Carlo [BP99]. Carlson [CF99]. Cartan [KS94c]. Cartesian [Lei99a, Zha99]. Case [AL95a, BJ94, nCZpZ96, D’A97, DSW96, Gee93a, Gee93b, Gee97, MH96, Wor97, BW91, Koš94, Ška90, dOGH99]. Cases [Che92, McE94, nCjH98b, Nie90]. categorial [Par98a]. categorical [SRT90]. Cauchy [BD95, BKO99, FHR93, GMJ90, GMJ95, HR95, HST97, Hei98a, HM90a, HM90b, JL98b, KO97a, LN97, MP98b, PPS96, RS98, Tyr91, Vav97, Zam94, Zen93, vD98]. Cauchy-Binet [Zen93]. Cauchy-like [KO97a]. causality [HDTR99]. Cayley [GM98a]. Cell [WR94]. Cells [BBLP95]. cellular [LV91a]. Centers [BE93, KM91].

Central [AB94b, Bon97, Le 98]. Centralizer [dH97b]. Centralizers [– Dok95]. centrohermitian [Pre98]. Centrosymmetric [CP96]. Certain [AB94a, CCK92, CRS94, CCR97, Coh96, Dru94a, DdO94, EHS93, FH96b, Fd92, Fri92, Har93a, Hil92, HS97b, JPR^{+}96, Lav97, cL93, Lim97, Rou93, SJ98, da 98, de 96a, Cai99b, fC91, El 98, FFT90, For90a, Hwa90a, JS91, Lav98b, MS98, Smi90, Son91a]. certains [El 98]. Cesáro [CH97, SW91d]. CG [San91, vV90]. CG-type [vV90]. CGSTAB [Cao97]. Chain [Scu93, Zie93a, Zie98, BS99b, BG98a, Scu91, Bou98]. Chain-group [Bou98]. Chains [Bai97a, CC96, Chi96a, Har93b, KNS98, Kri95, Nie97a, Sem95, Xue97, Flo90, For90b, Guo99, KPS91, NS99b, Nie99a, Vil98]. Challenges [Ano96a, Uhl98]. Challenging [Jia98b]. Chandler [CR94a]. change [DF98]. change-point [DF98]. Changes [JS99a]. Changing [DMB95, Kru92]. Chaotic [Fur96, FMP96, Str97b]. Character [Mal96]. characterisation [KSP98]. Characteristic [BCT93, DD97a, Fis95, FV97, HWV93, Her92a, Liu96, Sch93, Wan93b, dS97, Bot99, CU99, Fie90a, FS99b, Her91, HPWZ98, LS93b, YW90]. Characteristics [Wan93a, AH90]. Characterization [AP94a, Ata92, CL90, Ebi95, FPP96, FM93a, ID92, Kra95, KL98a, LS97, LMMR95, NY93, Nie98b, PŠ98a, RW97, Wat92b, AZ90, DSG98, GdOH98, HL91, Hua99, KS99, Kra91, LM99, Lim91b, MPS97, NY91, Pap90, Per91, WMP98, dOGH99]. Characterizations [BP90b, CG92b, FI95, Fri90a, HOŠS96, MS93b, NF95, Swe93, CRS99, Fie98, HOR98, Jia99]. characterize [CS98a, Ost99]. Characterizing [Bel99, dS95]. Characters [Cha96a, Chi95, Chi91]. Charles [Uhl92]. Chebychev [CR92]. Chebychev-Vandermonde [CR92].

Chebyshev [EFP98, Li95c]. Checking [Roh95]. Chemical [BL93b]. Chemnitz [Meh98]. Chief [Bru93, Bru95, Bru96, Bru97, Bru98, Bru99a]. Choi [SSHS92]. Cholesky [AM93, Gay91, Pan93, Ste97c, Sun92a]. Chordal [PPY93, PPY95, SM93, BC91a, McC93]. Chromatic [Cao98a, DF95, LMR95, dJ95a]. Circle [Gen96b]. circles [LS99]. Circuit [Duc95, jWySmJzZ95]. Circuits [GW95, ORvE99]. Circulant [CW95a, Hua93, Hua96b, KT95a, Laz95, Rja94, TZ96, Tyr95, Wai97, BCCR99, CJY91, CGV98, Hua90, KS90, SW94, Tis91b, Zha91]. Circulant-Block [Rja94]. Circulants [ABCP97, CKN96, Chi95, GN97, Gad90, LC96]. Circular [CIM98, TY99b]. Circularity [CT94]. Circum [THW96]. Circum-Euclidean [THW96]. Clan [Pay95]. Class [AD93, ABGR94b, AD96a, AL96, Bai97a, BW97, EGL96, Fd92, Gol93, GTÖ96, HZ98, HS96b, JPR^{+}96, Lop94, MS93a, Mic93, NV95a, OK96, OST93, RW94, Ser97, Wan98b, Yan93, AB90, AB91, Bai98, BHL99, CC94, nCjH98a, Din93a, Dru98, Ell94, EG90, Far99, FM92, GTT99, HP99, Hou99, LW98a, LWYS99, Loe90, Mal98, Oga99, Oua96, Tre99b, YP91]. classe [Oua96]. Classes [DZ95, Eva95, GOV96, Lav97, LW97, LP94, LP95b, NV95b, Nag95, Sil96, Sil97d, TY97, VW95, War90, App98, Cap98a, FJ99, FJK98, HO98, Hon98, Hwa90a, Lei99a, Lev93, Lev94, Lev97, Lin98b, OS94, Pvv98, San91]. Classical [BLS97, – DL94, GM95b, HK97b, MPŠV97, RR96a, Hir90]. Classification [AHS98, FPP96, GKv98, HS96a, HS97a, Pan99b, FM92]. Classifying [Wat90b]. Clifford [Sob96]. Cliffordean [Yas96]. Clique [PPY95, vLS95]. Clique-Tree [PPY95]. Closed [Beh94, ERS95, LWYS99, Wer90a]. closed-loop [LWYS99]. Closing [McE93].

Closure [Sun92b]. Clustered [hCjXL96, Xie97a]. Clustering [Tyr96, PPG91]. Clusters [TYZ97b, TZ99]. co [Mer91b]. co-immanantal [Mer91b]. Cochran [BH90, Šem96, WCM99]. Cocycle [LS95a]. Cocyclic [BCJ96, Yan98]. Codes [Chi95, CF93, Gol97b, Kan95, PT95, CM95, DSG98, QOQO98, Shi99b, VSB90, von91, AP99a]. Codimensional [WmL97]. Coefficient [AL95b, ZM92, nCjH98a, RR99a]. Coefficients [CRS94, FN98a, Kit93, Rat97, Rho93, Rho97, RR92a, Sen93, – Dok90, HR98a]. Cofactor [Bes92]. Cogrowth [Hum97]. Coherent [Hob95]. Cohesive [FH96b]. Cohomology [Hap97, RS93]. Colin [vLS95]. Collatz [FN98a, Mar92]. collection [Hou99]. Collections [DR96]. collinear [Mub99]. collocation [San96a]. Color [BC92]. Colorability [BF95]. Colorado [LS95b]. Column [AL95b, BKGK94, BCT93, FM97a, HP99, NP93, SH97b, van96, BH94a, LMZZ98]. Column-Allowable [AL95b]. columns [Van91c]. combination [Soh90]. Combinations [Alt94, DJ95b, AYC91, CW90, D’A98, Hou91, KS99]. Combinatorial [BS94a, BCT93, CL96b, FV96, JM97, MOTV92, Neu97a, CGV98, Her99, KS94c, Sha92]. Combinatorially [GZ98a]. combined [FFT94]. Comments [Bon94, WAS97]. Common [BGR90, CV97, GI99, GV96, GKP92, Ton97, Bha91, ÖA91]. Communication [O’N93]. Commutative [BR92, BP92b, BZ96, DZ95, HAPSG96, Jel92, Kam94, Koš93, Koš97, Wu94, ZM92, AP99a, Bha99, Che91, Gur91, HASG90, II98, LL91a, Pra94]. Commutativity [Eva95, MS99a, So94b, Pet99]. Commutators [Hir90, LR94a, MS97, PR94b, VW90, YC99]. Commute [AEG93, SW91e]. Commuting [Els94, GK97, LM94a, Pry92, Wil93a, BP90d, BR99]. Compact [FK94, God97, IKF^{+}93, dS95, LZ99]. Compactly [BP98]. Companion [BK96b, BK96c, CL96b, DS94c, Lin98c, Pta92, Ptá95, Ros93a, Wim93, BW91, Car91, Sol99, VPV99, VW90]. Comparative [HL95]. comparing [BP90a]. Comparison [Chu98, CP99b, Daz98, EU93, FP97a, JL94, Mil95, Nab96, Ste97a, Wan98b, Woź97, CP98b, VV92]. Comparisons [BKS95, LN97, Pap92, Son91b]. Compatible [DIO97, FV93, Lav98a, Li92b, MLS91, MLS95, Li91c]. Competing [PT94b]. Competition [BMV95, rKMMR95, LMR95]. Complement [Smi92, MS98]. Complementarity [CF96, DCPS96, DRT93, Ebi95, JS98b, KY93, LT97b, SG95, TT95, AM98, Bro91b, Din93a, HSW99, MP90, OPR91, PI91, See99, Xu99, YP91]. Complementary [BT94, Goe96, Rie97, Wim93, dSZ97, Fri91a]. Complemented [HP93]. Complements [BT94, Gem97b, LZ97b, RRR93, Bel99, Bon94, CRS99, JR99, Liu99b, LW99, Liu99c, Par98a, WZZ99, Wim99]. Complemetary [AH94b]. Complete [BS95b, DIO97, Gay91, JR99]. Completely [Bar98, LS93a, LW97, SZ93a, XX98, JJPT90, Str98]. completeness [Str98]. Completing [BLJW95, Dai96]. Completion [BJT93, CHR97a, GKv98, GRS92, Hou96, JT95, JS96, Lau97, Lau98, Sil97c, Gee99, JS99b]. Completions [BJ94, CS92, CD96, Dan92, Ghe92, Hog98a, Hog98b, JTU97, KW95, KR96a, KR94, Kru95, RS92, Sil97b, Woe94, CS91, Dan91a, GHJT99, JTU98, LMZZ98, LJ91, Zab99]. Complex [BRL98, GHN97, HKS96b, IKF^{+}93, IAMC96, KH95b, LR95, Li95c, Nov97, SW95, ST96b, Szu92b, Szu92a, Tho97, BKGK94, BGS99, DR91b, Gab97, GHN99, Gro99b, Gu99b, Hon90, HM99, HWZ90, Mar99a, Pav99, ST91, Sch98a, Tho91, Wer90a]. Complexity [Bsh95, GO94, KK96, vD94, BS91, LV91b, NR99, von91]. complicated [Nie90]. Component [LS98, Buf91, TK99]. Components [Lab92, ZD92, Far99, KR91, McL90]. Componentwise [HH92, MR94, Sla98]. Composite [Mai91, Ott93]. Compounds [Dru92a, Dru94b]. Computation [ACR96, AC97, Bal99, GS98a, Har93a, HWV93, Hod93, IKF^{+}93, Kar97, LZW97, PS97, Var98, van94, BG98a, Di 97, Gem98, LWX99, Ols98]. Computational [LMS91]. Computationally [Gem96]. Computations [BP93, Boz95, Huc98, Nas92, Ng97b, KBŠ90]. compute [BCCR99]. Computer [LY93c, LC91b]. computers [Gay91]. Computing [ACF96, Atk93, BK99, Boy95, CK94, DG93, D^{+}99, Die96, Fri98a, Gri96, HL96a, Hon92, IW92, Mor91b, Pat93, Sto98, SZ93b, Tur91, Uhl94, Wan92, Wat96c, At91, LC91b, MV98a, PC90, Ple98, QOQO98, SL97, Van91b, Vos99]. Concave [CD92]. Concept [Par92, Ser97]. Concepts [Gee96]. Concerning [Mah92, Sin90, Mol99a, Neu98, Sch99b]. concordant [LL91b]. Condensed [EI97]. Condition [CGK96, Ded97, GL97, Hig95a, JL97, MUV^{+}97, Neu97b, WZCZ93, Yip96, BG98b, DR91a, LS99, Mol96, SV99]. Conditional [DF95]. Conditioned [CW95a]. Conditioning [LP97, Not91, MN91]. Conditions [BH92a, BCHS97, CR94b, Dia96c, EF97, Gee93b, Hob95, JM93, Wei97a, van96, BK90, CGP99, ERR93, GRO98a, Li91b, Li95b, Per98, See99, Soh91]. Condorcet [FR95b]. Cone [AP94a, Lyu95, NO95, Nie98b, Tam92, Wat93b, BT97a, BHL99, CP99b, HWLT91, LVBL98, Mat91, PSW98]. cone-preserving [BT97a]. Cones [CL97b, SH97a, AIY90, Her98, Kov99, Lew99, Mer91a, Obe91, Pat99a, SW91b, SW91c, SW91a, ST90, vvS99]. Conference [Ano94c, BHR91, BGH92a, BGH92b, BKMS97, Da 94, DLdO94, HPUV97, LS95b, ORS94, Meh98, OPSS98]. Configurations [Hob95, Sha98a]. Confluent [GMJ90, Vav97]. Congruence [– DL95, HHL92, LW96, Mer94c, Pie99, Wat94, Cha96b]. Coninvolutory [HM93]. Conjecture [AD97, BM94a, CH92, DC92, Dru96, Gre92, KA95, Kir97c, Kov94, LW95, LSS98, She95, Sil97b, Son95, Sou94, Taş95, VP92, vH96, Ano96b, BKdP94, CA98, Har90, Kov99, MV93, Neu98, Pus90, Sch99b, YW99, Wan99]. Conjugacy [Gol93, VW95, Ell94, Lev93, Lev97]. Conjugate [SV96, BGS99, Jam91b, KH90, O’L91, Str91]. Conjugate-Gradients [SV96]. Connected [Gro97b, LM98b, Bel91, NT94, Pal90, Sch99b, Ste99]. Connectedness [JS98b, MP98a, NWZ98b]. Connecting [HN93b]. Connection [BK96c, CG92a, nC98, Lau98]. Connections [Auj97, Fle93, JT95, BFW98]. connectivity [AdlP98]. Conpseudosimilarity [BHH90]. Consecutive [FM97a]. Consecutive-Column [FM97a]. consemisimilarity [BHH90]. consequences [Dan90]. Conservative [Els94]. Considerations [Die96]. consimilarity [Hon90]. Consistency [Gee93b, Gee96]. Consistent [Chu98, Rob95, Sid92, CWL99, GJS91, Pot98]. Consolidation [Far97]. Constant [BF92b, – Dok93, LY93a, RZ97, FD99]. Constant-False-Alarm-Rate [BF92b]. Constants [CS96, DO94, KT96a, Li93b]. constellation [Hou99]. Constrained [AP96b, Bal95, DH98, Hu95, Jan97a, Lew98, Mah96, Nie97b, Shi95, WY96a, CZ91, Fir90, Jam91b, LJ91, Wei98, Wer90b]. Constraints [CD92, Fra95, JM97, Leo94, RZ92, Wim94a, YT92, BTT91, Rya90]. Construct [Mir92]. Constructing [Bon97, Chu99, HRW99, JM12]. Construction [AP99a, DS94c, Gol93, GM97b, HA94, HX92, HKS96a, KPRV97, KT95b, Kra95, SV98, Yas96, Car91, LS93b, PZ90a]. Constructions [Rez95, AHJP91]. Constructive [MPS97]. Contain [– Dok95]. Containing [BK97b, CF99, FL99]. contenant [CF99]. Continuation [ACR96, Din93a]. Continued [Bau92, nCjH98b]. Continuity [DS94b, DH97a, JS90a]. Continuous [AS97, AH97, CR94b, Fau92, Gee94, Gee96, HC98b, IW93, IO96, Meh96, MDD93, Wim94b, XL95, Che91, Chu91, ETS90, KTK99, Str98]. Continuous-Time [Gee94, Gee96, HC98b, IO96, Wim94b, XL95, Che91, ETS90, KTK99]. Contours [MM94]. contractante [Abd95, Che99]. Contraction [CRS94, CP90, He94, NO98, Abd95, VSB90]. Contractions [CHR97a, HR93b, NW92, BAG99, Dru90]. Contragredient [HM95b, ŽWB93, GPS99, Gel99, RG99]. Contribution [Kap95, Szu90a, Zho99]. Contributions [Far96]. Control [Ant94, BK97c, Dat94, FP96, Gee94, Gee97, Gom95, HFK94, Kam94, Meh96, Mus94, PS94, SO94a, ETS90, LWX99, LN98, SNC98, ST93]. Controllability [Fun96a, Gee96, GRS92, RW97, SDT97, Che91, GRO98a, LWYS99, Loe97, OTV93]. Controllable [Fun96b, FPP98]. Controlled [Log98]. Controls [Gee94]. Convection [Ott93]. Convection-Diffusion [Ott93]. converge [DL92, DL01]. Convergence [Ach93, Bat94, BCP93, Cao95, EN93, EF97, HN93a, HNT96, HR98a, HK93a, KS91, KS90, Lav98b, Li95c, Liu91, MR94, Ott93, RR94, SG96, SV96, Son97, Str97b, TS97, TYZ97b, Wan94, Wan97b, WE91, Xin93a, Xin93b, ZZ98, CH97, Dax90b, EKN90, ENV91, Fre91b, mJS91, Jia96, JL98b, KBŠ90, Mar90, NS99b, Nie99a, PI91, Pot98, Sou91, Str91, SBKK98, Wan91]. Convergent [CAHJ98, Cai99a, Fan93, GKP93, Cai99b, HRZ91, Paa91]. Convertible [HKS96b, Sha93]. Convex [Alt94, AV96, Bal95, CW90, CL97b, Don92, Kiw95a, Kiw97, KS99, Lew99, MT94, MP95, Prz95, AYC91, Egg90, Hur90, KPY91, MLS98, Mat91, PI91, Soh90]. Convexity [Auj93, HR94, KH95b, Kop96a, Tam97, Vol99, FHP90, MLS98]. Convolutions [BT97b, Ste90]. Coordinate [Gee94, Sim95]. Coordinate-Free [Gee94]. Coordinates [ACE95, Cox95]. Copositive [ACE95, Had97, LF93]. Copositivity [Bom96]. Coprime [BK97a, CG97, Var98]. corank [GMŠ98, Mol99a]. core [Gay91, Har94, TK99]. Correcting [CF93, Pie96]. correction [CP90]. Correlation [Abt94, Neu96a, YT92, GR90, GPW90, NW90, SB90, tKWS99]. correlation-preserving [tKWS99]. Correlations [BPY92, DS94a, GZ94]. Correspondence [Wim94b]. Corrigendum [DL01]. Corrigendum/addendum [DL01]. coset [Cha96b]. Cosets [– Dok95]. Cosine [BT97b, FBO97]. Cosines [MBI96]. Costs [Gee94]. Cottle [DCPS96]. Counterexample [BF92a, Dru92a, NT94]. Counterexamples [Mir92]. Counting [Kri94]. Coupled [KT96b]. Coupling [BRW94, LN98]. Course [BB92]. Courtship [AR95]. covariables [WLN99]. Covariance [FH94b, LPW92, Ton97, WMD95, Cra99, HS90b, Jin90, SB90, Won90]. Covering [DFF98, BDV98, Fer99]. Covers [BNP96]. Coxeter [BL94b, Bol95, Hap97, Lak99]. Craig [Olk97]. Craig-Sakamoto [Olk97]. Cray [Pap92]. Criss [DRT93]. Criss-Cross [DRT93]. Criteria [ACE95, Dom96, Far95b, mGhW92, GhW96, LF93, Väl96a, FHP90, HL99]. Criterion [Gil97, LLH^{+}98, Szu92b, Szu92a, WmL97, BM90, BP90c, Szu90b]. Critical [HS95, DdPC90, ORvE99]. Cross [DRT93, GKT95]. Cross-Positive [GKT95]. Crossing [dS95]. Crystallography [LM97]. Cubes [HKL96]. Curriculum [Ano98b]. Curvatures [KV90]. Curve [WMD95, WLN99]. Curves [dS95, CN99b, Fie90b, Vin90, Mir98]. Cut [DG95a, LP95a]. Cycle [BJT93, EHS93, Fri97b, SK97, Kat99, Rum98a]. Cycles [DF95, van95b, AGK^{+}98, Tho96, WLL99]. Cyclic [And95a, AC97, AH97, Chi95, – Dok90, GHN97, GT96a, GW96, HN93b, HNT96, KS94b, MO96, YP96, BS99c, GHN99, KPS91, Lev93, Lev94, Lev97, Tam99]. cyclization [Sch97, Sch98a]. Cyclotomic [Wat93a].

D [BB92, FV97, Gał91, Kat99]. DAE [Cla92a]. DAEs [RR96b]. Dagger [HOŠS96]. d’algèbres [Oua96]. damped [PC90]. damper [Nyl99]. dans [ABM93, BM95, Téc96]. d’applications [Téc97]. Data [AM93, AGR92, HFK94, Has96, LS98, Mey97, TK95, YA94, BKGK94, EKN90, GKR90, HN99, HB90, Jin90, Roh90, SNC98, SRT90]. Data-Fitting [AGR92]. David [Nas92]. Davis [CR94a, Xie97b]. Davis-Kahan [Xie97b]. – Doković [Wan99]. December [BMH94]. decoding [AP99a]. Decomposability [DHM95, Mar99b, Dia93, LN91]. Decomposable [Dia96c, FN93, GPRL94, Lei97, Lim97, Sha93, Fon99, For90a, Gro98b, LW98a, Lim91a]. Decomposing [Cam93, Rya91, McL90]. Decomposition [BoG92, BOS93, BR95, Cla92a, Cop92, De 92, DD97b, GM97a, Har93c, HM92a, Hon92, HM93, JM97, Kni95, Lag92, LS98, MS93c, MVV93, MZ91, PV95, Ptá95, Ren96, WY96b, Zha96, Bai98, BN99, Cha96b, Cha98, Del90, D^{+}99, GHM99, Har94, Hig94, HKS96a, IHM99, KN91, Lev98, Mar99a, PW94, Par98a, Rob91, Tis91b, Vil98, WE91]. decomposition-type [Bai98]. Decompositions [BvR^{+}97, BS95b, Boz95, DZ95, DZ98, DHS94b, FH93, GVS96, Peñ97, Tyr97, CR93, Gad90, GWS99, Pie91, Zha99, vdMRR99]. Decoupling [IW96, YC92, van94]. Decreasing [RSR94, – Dok90]. Dedekind [App98]. Deepest [Kiw97]. défaut [Del99]. defect [Gro98b]. defective [Pav99]. Deficiency [PR94b]. Definability [Cor93]. Defined [FP92, Gon97, Kam94, Wu94, Che91, See99, Wu97]. Defining [Gov97]. Definite [AP94a, AI96, BJT93, BM93a, BHP97, DER97, Die94, FP97b, GT97b, HR93a, Hu95, Kra98, Li93a, Li94c, LM98a, MV98b, Nab96, Neu97a, OSiA96, gS96b, Val92, WC97, Wu92a, BP91a, BG98b, Buf91, BT95b, CL98a, CH99b, Eng93, ERR93, Erx90, Fri98b, GHJT99, HNN91, Har99b, IEs98, LR90, LJ91, MV98a, Mal98, Pat99a, SY98, Sun90b, WC96]. Definiteness [KSV95, BT91a, Lei99b]. Definitions [Dra92, XB90]. Definitizable [LMY94]. Deflation [GLW95b]. Deformation [GPM97]. deformations [GPS99]. degeneracy [Bro91b]. Degenerate [Ste93, nCjH98b]. Degenerated [Wor97]. Degree [AP94b, BK94a, Bon97, Mer94a, Mic93, OS96b, Ste95, de 96a, AP99b, BR90, BK90]. Degrees [Wat93a]. Delaunay [DG95a]. Delays [Log98]. Deleting [NP92c]. Delsarte [AS98, CM95]. Delsarte-Goethals [CM95]. Denominator [MM93b]. dense [Van91c]. Density [CZ97, Chi97, NL96]. Depend [HN93d]. Dependence [Zho96]. Dependency [Flo90]. Dependent [Has96, MV98c, RR96a, RR96b, CMH90, CC94]. depending [SW90]. Derivation [GM93, Mec98, SC98, SC98].

Derivations [BS99b, CM93, Jøn95a]. Derivative [AEG93, KN97, Par92, Rat97]. derivatives [BS99b]. Derived [GPRL94, CK91]. Derogatory [MH96]. descent [Bre99]. described [Dat91]. Description [BK96b, BK97a, DC93, Gał96, GP94a, Kan93, RAB93, DC93]. Descriptions [SZ93b]. Descriptor [Fun96b, Kat94, Log98, Rat97, Rei94, RW97, BGBMN99, CMN99, KTK99, LWYS99, XK94]. Design [SZ93b, BGBMN99]. Designs [BS92a, Boy95, KMS96, MCK95, PT95, Yan93, AHJP91, Cho98, NWZ98a, Wie90, YW90, Rez95]. Destabilizing [Log98]. Detecting [Bom96]. detection [DF98]. Determinancy [JOv95]. Determinant [Chi93, Chu97, Dru94a, HWS97, HKL96, JT97, MV97, Mur95a, Mur95b, NR97, OK96, Pat97a, BP99, Gat91, GB99, GHJT99, Mar98, Sal98]. Determinantal [BM94b, BM94c, DC92, Dru96, GMJ95, Olk97, Rob96, Fri90b, MV93, Müh90, Vin90]. Determinantal-Rank [Rob96]. Determinants [BKMS97, Els97, Rou93, Vel95, CGP99, Hon98, Le 94, RR99a, Le 94]. Determination [KU93, RR98]. Determine [Bol95, TG97]. Determined [RR97]. Determining [Ito97, LZW97, Val92, Väl97, Fri98b]. deterministic [Nie90]. Deterministically [HL96a]. d’Etherington [MO91]. Development [Sid92]. Developments [Ton94]. DFT [HR98b]. diagnosing [HN99]. Diagonability [SG99]. Diagonal [AP96a, And96a, At91, BD93, BLJW95, BLD93, BK93, CHS92, Fie94, FW96, FM93d, HR93a, HHNS96, JLOvdD96, KO97a, Kal96, Kha96, Min94, MT94, NS97, Tam98, dS97, da 97, Chu99, Daz97, DHS94a, Har91, JSOvdD94, KN99, Kra91, Mar95b]. diagonale [Daz97]. diagonalisability [OŠ98]. diagonalizability [HL99]. Diagonalizable [Bot98, JL97, Bot99].

Diagonalization [AGMP97, BL91b, Kuc91]. diagonalize [BLV98, Bro91a]. Diagonally [BDB92, Far95b, mGhW92, GhW96, GMJ95, Hua95, LT97a, SW94, FL91, FM93c, GM98b, GMJ90, Nab99, Paa91, hXyY98]. Diagonals [Kye95, Lon95, Lon96]. diagram [Hua91]. Diagrammatic [Mit99]. diagrams [GPM99a, Mit99]. Diameter [Neu96b, She96a]. Dichotomy [GS96a, Ion98]. Dicycle [BNP96]. did [Hof98a]. Difference [Car96, D’A97, FJAK96, Kit93, LPW92, Swe93, WW94, BP98, CP99a, fC91, Dat91]. Different [LT95]. Differentiable [FGP94, FPP98, MDD93, NO95]. Differential [AEG93, Bro94, Coh96, Eva95, Fd92, FJAK96, Gee93b, HJRS97, Jod92, JAG98, Now96, OSiA96, RR96a, ZM92, Cra99, D’A98, Dat91, EG90, Jód91, KM90, Lub98, ZJ91]. Differential-Algebraic [Gee93b, RR96a]. differential-difference [Dat91]. Diffusion [Ott93]. Digraph [RW97, RR98]. Digraphs [BNP96, EHS93, Fis95, She96b, SN98a, jWySmJzZ95, Yan92, Sha98a, WLL99]. Dilatations [Bau92, YC99]. dilation [FFT90]. Dimension [Boz95, DE95, Fri98a, GGM96b, Kan93, Laz95, BC91c, KB99, Loe97, PZ90b, VSB90]. Dimensional [BvR^{+}97, CS96, CR94b, CG97, DIO97, GM97b, Gra96, Hau92, JPR^{+}96, Lyu95, OST93, Rud98, SSHS92, Bha99, Bog91, BDV98, DBRW91, Koš94, Osa91, Str98]. Dimensionally [MO97]. Dimensionally-nilpotent [MO97]. Dimensionnellement [MO97]. Dimensions [MO96, Wil93b, YC92, SJ95]. diophantine [VT99]. Dipolynomial [HPW94]. Dirac [RK96]. Direct [Bsh95, FNW96, HW91, Lev91, RRR93, Sla98, FJ99, Wim99]. Direct-sum [Lev91]. Directed [CP95, FV97, Neu96b, Her94, Tho96]. direction [Che98, Tod91]. Directions [Her92b, JS90b, Tod91]. Discontinuous [RR96b]. Discrete [AS97, BK93, BCHS97, CK97, CBSH94, CR94b, FBO97, Fri96, Gee97, Gol97a, Grö93, Gur95, HC98b, IW92, IW93, Ion96, Kam94, Kat94, KT95a, KP94, Meh96, MHR93, Oar96, Raz97, Spa95a, gS97a, Zam94, Adk91, Cec90, DR99, FI99, GRO98a, GL91, Han90a, HPWZ98, Ion98, LL93b, MM98, Nie90, Ols98, PST98, Ste90, SS98, gS98b, VS94, Wim91, dOGH99, BT97b]. Discrete-Time [BK93, CBSH94, Gee97, HC98b, IW92, IW93, Ion96, Kam94, Kat94, KP94, Oar96, gS97a, GL91, HPWZ98, LL93b, gS98b, VS94, Wim91, dOGH99]. Discretized [FC97]. Disparity [Mos94, Mos95]. Dispersion [BPY92, Jen96, HS90b, NF95]. dispersion-matrix [NF95]. Displacement [BM98b, Boz95, CK91, DZ95, DZ98, Gad90, HH94a, HH94b, KO97b, Ros93b, SK95, Wat92a, Woo93, ZZ96, Hei98b]. Displacement-Structure [KO97b]. disrespectful [Bee90]. Dissipation [Gee94]. Dissipative [OH96, RT94]. Distance [BS92c, Boy95, CF97b, DO94, DK92, FC97, GN93, GK95, GT96b, Hae96, HT93, HLWT96, HS95, HOŠ96, JT95, Lau98, THW96, AFK90, BŠ99a, FG99, HN99, HWLT91, HRW99, JM12, Mom98]. Distance-Based [FC97]. Distance-Regular [GK95]. Distance-Regularity [Hae96]. Distances [HLM96, KS94b]. Distinct [Soh97, HL99, Soh91]. Distributed [GTÖ96, WL94, Coo98]. Distribution [CY95, Eve96, ST96a, ST96b, Tyr94, Tyr96, Ede91, Jin90, KK90, LS99, San96a, Tre99a, Tre99b, Yon99]. Distributions [Boy95]. Distributive [Spa97, Tan98]. Disturbance [van94]. Dittert [CH92]. Divergence [HX92]. Divisibility [Thi95]. divisible [AHJP91]. division [AG98, BHH90, MVKR99]. Divisor [GV96, Bha91]. Divisors [AM95, BGR90]. Do [TG97, Fri91a]. Dobrushin [Rho97]. Domain [CBSH94, HAPSG95, KU93, Put98, BK91b, DR99, LN91, ÖA91]. Domains [Gou97, HN93a, HNT96, App98, BRP90, CLW90, II98, Liu97, PRB91]. Dominance [Sou94]. Dominant [BDB92, Far95b, mGhW92, GhW96, Hua95, JSOvdD94, LT97a, NS97, FL91, FM93c, GM98b, Jar90, Nab99, Paa91, hXyY98]. Dominating [FMS98]. donné [CF99]. d’opérateur [BC95b]. d’opérateurs [CF99, Del99]. d’ordre [Col96, Oua91b]. Double [Cha96b, Jia92, PD97]. Double-Shift [Jia92]. Doubly [BNP96, BGS92, BS94b, EF95, Fie95a, Gib92, Hwa95, HS97b, LS93a, LT97a, Lon96, Min94, nM94, PT95, SMB92, SM93, SHJK97, WC97, BR99, GP91b, NT91b, Son91a]. Downdating [AGR92, Pan93, PV95, WZ99]. Drazin [PRB91, Sid92, Sid99, Wan97a, WW97]. Drazin-inverse [Sid99]. Dual [Bd97, Zie93c, AIY90, LMS91]. Dualistic [OSiA96]. Duality [Fuh94a, LT91b, Mar94a, Rud96, ACWY99, HASG90, Rak99]. Duffin [Che90b]. d’un [BC91b]. d’une [Daz97, Le 94, MO91, SC98, Var96]. Duplicate [Var96]. Duplication [MO91]. Dupliquée [MO91, Var96]. durch [Gab97]. Dursun [Ano96b]. Dynamic [CHX92, II98, BJW98, ST93]. Dynamical [Bro91a, GM97a, HAPSG95, HAPSG96, Kno92, Sch94, Wei94, CHASG98, HASG90].

E. [Saa92]. Early [Far96]. ebene [Gab97]. Eckart [Wei98]. Economic [Mos94, O’N93, Mos95]. Edge [FP96]. Edgeworth [Chi96d]. Editor [Ano96b, Bru93, Bru95, Bru96, Bru97, Bru98, Bru99a]. Editor-in-Chief [Bru93, Bru95, Bru96, Bru97, Bru98, Bru99a]. Effect [NP92c, ENV91]. Effective [Nie98b, Rja94, BS91]. Effects [FP96, Log98, MG99, SP99]. Efficiency [LN97, Fre91a, Jen91]. Efficient [Gem96, GK93, Lam93, Mae96, KPY91, ORvE99]. efficiently [NF95, Tre91]. Eidson [GHN99]. eigen [GM98b]. eigen-functionals [GM98b]. eigenbounds [Im97]. Eigenmatrices [Wat97a]. eigenpairs [Erx90, Kni99]. Eigenpolynomial [Men97]. Eigenpolynomials [YW97]. Eigenproblem [CGB95, Gov94, BFW99, BGE91]. Eigenproblems [Jia97b, Jia98a, Kat94, MP97a]. Eigenprojections [SV95]. Eigenspaces [CRS93, Her99]. Eigenstructure [Mom98, LN98]. Eigenstructures [HW93]. Eigenvalue [BF97, BY95c, Bor95b, BCJ96, Coh96, Ded97, Don92, EH93, Eve96, FJSv98, Fan93, FN95a, Fio99, FWWJL92, Fri92, FN97, GKv98, GLW95a, Guo96, GT97c, HM98, HRS98, IKF^{+}93, IAMC96, Jia94, KS91, Kol98, Koš96, KT96b, Lew98, Li94c, Li97b, Li97c, ME98, Mah96, NU97, Rad96, RSR94, See99, SY95, WG93, WN95, AH95, An98, AB90, AB91, At91, BF98, BFW98, CH99b, CH99c, Dai99, DD90, Ede91, FLL97, Fri98b, HRW99, HS99, IC94, JM12, Jar90, Jia98b, Li91b, Li95b, Li97a, LC91b, MV98a, Nyl99, Ple98, Roe90, Ruh94, San96a, SP99, Sun90a, Ten90, Tre91, TZ99, Vos99, WE91, ZGL99]. Eigenvalue-Constrained [Lew98]. Eigenvalues [AL98, And96a, BP92a, BM96, BM94a, BE93, CHL95, CY95, hCjXL96, Cao98a, Car97, DG93, Drm96, Dru92b, EJ93, Fie95a, FM93d, Gil93, GK95, GM98b, Hae95, HRR92, Hof97, Ito97, Juh96, KN95, Kir97b, Kra94, KST99, Li92a, Li95a, LZ97a, LM98a, Lin98c, LZ97b, Loe92, MOTV92, MV97, MP93, Mor91b, NY93, OŠ96a, Pat93, Per96, Rus95, Ser98b, Sha95, Tak96, Tan98, Til98b, TR92, TFV94, Tyr94, Wan92, WmL97, Wuw97, Xie97a, dGK^{+}92, van95a, van95b, Bak99, BEK90, Di 97, – Dok90, Gil95, HL99, JSOvdD94, Li98b, LZ98b, Liu99b, Mar95b, Mer98b, NY91, Paa91, Pav99, Pin98, PP99, Row98, Sha91a, Sid98, Sil91b, Sti91, Str98, Sun90b, Tar91, Yon99]. Eigenvector [Har97, Men97, Mey97, Ols98, At91, SV99, Str98]. Eigenvectors [AK97, Cec92, DR94, HWV93, KN97, Laz95, Mas94, CL91, Jia99, LSJ91, Mer98a, ORvE99, Tan98]. eight [HP99]. Either [PR94b]. Element [BR93, CHX92, Huy96, MT94, DHS94a, RR99b, San96a]. Elementary [EI97, FFK93a, FKK95, GV96, JOTV93, JOvdD99, Pat97a, Que94, Vin90, jWySmJzZ95, FKK97, WLL99]. Elements [BLD93, DFF98, – Dok95, FRLM97, HHS97, SJ96, Tam98, TV92, CT91, Chu99, Fer99, HL92, Ish95, LW98a, Lim91a, MR98b, RGS98, TV90]. Elimination [GP92, GP94a, Mil95, NP92b, ON96, Pap92, PN91, PN92, Tuc96, Coo98, Par98a, Rya91]. ellipsoidal [SW91b]. Ellipsoids [Hol95]. Elliptic [Fie94, GS96a, YW97, CP99a, Idr99]. Elliptical [Neu96a]. Embedding [For90a, FS99a]. emission [SS90]. EMS [LA94]. End [Gee97]. End-Point [Gee97]. Endomorphism [HAP93, HAP95]. endomorphisms [GM98b]. Entries [AB95, EHJL97, EL97, FM93d, Gib96b, GK93, GOV96, Min94, Rea97, KN99, Mar95b, Sil91a]. Entropy [CRS94, CIR^{+}93, FFT90, Fri97a]. entry [OTV93]. Entrywise [Xue97]. Enumerating [BC92]. Enumeration [Lin93, Yan93]. EP [HK97a, Kol99, MK98]. Equal [TV92, Mom98, TV90]. Equalities [Auj97, Gil93, HRS98, Lin98a, SSS97]. Equality [AP96b, Che92, Dia96a, Dia96c, DH98, GP95, LSS98, NP92c, Fon99, Ong91]. Equality-Constrained [AP96b, DH98]. Equation [And96a, BF92a, BG96, Bha97b, CG92a, CL97b, D’A97, DD97a, Eva95, FL96, FH94b, HTP96, HR92, Hua96a, IW92, Ion96, Jod92, JL94, Kit93, LY93c, Oar96, Pat97c, RW94, RK96, Roe97, Sem92, Sto96, gS97a, XL95, D’A98, Dob99, Eng93, ERR93, FI99, Gro99e, Guo99, HM90a, HNRR99, Ion98, IEs98, JZ99, KTK99, KSP98, Loe97, LL93b, Sch91, SS98, gS98b, Wim91, WL98, WLH99, XWZ98, ZX96]. Equations [Abe97, AB94a, Bai97a, BG94, Bro94, Car96, DHL97, DJ95b, EU93, FWWJL92, Fd92, Fra95, FJAK96, Gee93b, Gus95b, HX92, HP96, HR95, HM96, HL96b, HO93, HL97, JV92, Jon95b, KO97a, KT96a, Koš96, KT96b, LS92a, LS94, LRT97, LR93, MBI95, Now96, RR96a, Roh95, Roh97, Ros93b, RV95, RV96, Sak93, SS95b, Swe93, Tis92, US92, Von93, WF92, Wim94b, Wim94a, YC92, ZM92, ŽWB93, Bai98, BA– DR91, BTT90, BP98, BKM91, BKO99, CW93, fC91, Cop93, Dai90, Dat91, Di 98, EG90, GM98a, Jód91, JL98b, KM90, LR99, Lew99, Lub98, MM99, MBI94, Mit90, ÖA91, VT99]. equidimensional [PZ90a]. Equidistance [Wer96]. Equidistant [IS95]. Equilibrium [Sei96, See99]. Equispaced [TK95]. Equitable [God97]. Equivalence [BR95, CL96a, CRS94, Fle90, HHNS96, HM95b, Mer92, San96b, Sil97c, Wye97, YA94, Gel99, Hon91, KR91, Mat91, Oga99, RG99]. Equivalent [HM92a, Kru97, MR98c, Wei97a, VS94]. Ergodic [Vil98, Cap98a]. Ergodicity [AL95b, Rho93, Rho97, Sen93, HR98a, NS99b]. Erratum [BPS95, Bha97a, BT95b, Chi96b, Gil95, GH95a, Gro95a, Ikr95, KP95, KR96a, Li96, Mar95b, MLS95, Mat95, MV93, Mit97, Mos95, NF95, Spi95, SJ95, Tho96]. Error [AL98, CF93, Fri98b, Kni99, LT97b, Man92, MTS96, MP90, Sun92a, gS97a, JZ91, Li97d, Wei90a]. Error-Correcting [CF93]. Errors [Has96, gS99b]. espace [BC91b, CF99, Téc96]. espaces [BC91b, CF99, El 98]. Essay [Li98a]. essentially [BKdP94]. Establish [HK97b]. estimable [NF95]. Estimate [Fie95a, Kwa97, OS94, Str98, Wat90a, Wei90a].

Estimated [She96c]. Estimates

[KS91, LZ97b, SS96, BP99, Li98c, Tar91]. Estimating [GN93, Ves98, Wat95b, WMD95]. Estimation [Chi97, Cra99, Hu95, PT94b, Rum91, SNC98, Ton97, WL94, Xu97b, ZD92, AU99, Bap99, Chi99, Jen91, LS90a, Sch99a, SRT90, Wer90b]. Estimator [Car97, DSW96, LS92b, BP90b]. Estimators [Lis92, BP90a, Far99]. Etherington [MO91]. Euclid [Gen96a]. Euclidean [Téc96, AR99, BDV98, CG99, DN97, Gab97, Gem96, Lau98, SZ90, Ser98a, Shi99b, THW96, van95c]. euclidien [Téc96]. Euklidischen [Gab97]. Euler [Jod92]. Eulerian [BNP96, Dom96]. Evaluation [MHR93, RR99b]. Even [FK94, Tre93, Wan93a, Tho96, Tre99a]. Eventually [DS92, ZT99]. Every [Str97a, Tof97, Fur94, MMP99]. everywhere [FP92]. Exact [DO94, GM98a, HNT96, Jan97b, KS91, MCK95, Sak93, Shi99b]. exactly [Sau90]. except [GKR90]. Exceptional [Rin98]. Existence [BOS93, DHS94a, Ebi95, Gee94, HHS97, Jon95b, Zha96, Eng93, ERR93, Nyl99]. Expanded [Aya99]. Expansion [Rou93, SS93, PI90]. Expansions [ERS95, HRR92, HS97b, Sun91]. Expected [HJ90, Pal90]. Expected-value [HJ90]. experience [LMS91]. Experimental [BS92a]. Experiments [Ste97a]. Explained [CL98b]. Explicit [CY97, Gro99a, Haf95, Jod92, JS91, KPRV97, Rho93, Sen93, Wol96]. Exponent [BK97b, GKP95, HN93d, Kir97b, Neu96b, She95, She96a, She96b, Wie97, Yan92, Li90a, LMWZ90, Liu99a, Vol99]. Exponential [ACF96, CY97, EGLL95, Xu97a, Ves98]. Exponentially [Ng97b]. Exponentials [FS94, Lu98, AZ90, Lie90]. Exponents [Gra96, SB97, SN98a, SGN99, jWySmJzZ95, Liu90, SH98]. Exposed [Mar94a, ST90]. Expressing [Fie90a]. expression [CX98b]. expressions [Ish95, Mia91]. Extended [DD97b, HJK97, HC98c, Ion96, Mor97, SY95, SG95, AM98, Nie99a]. Extendible [CF97a]. Extending [Sea94, VPV99, Müh90]. Extension [ABL98, BC95b, CC96, CS96, CKL96, CG94a, DS98, Fur95, HR93b, HL97, KY93, LM92, MP98d, SV98, Ser97, WL93, WY98, Wat93a, ZL98, BZ98]. Extensions [BRS91, Buo90a, Chi95, CMLS95, Del99, HW93, MG94, PPS96, Won90, CF99, FFT90, GKW91]. Exterior [Pat97a]. External [Coo97, SZ93b]. Extraction [Ote90]. Extrapolated [FMP96]. Extrapolation [Mes97b, Sad93, Sid92, Xin93a, Din93a]. Extrema [BMTZ98]. Extremal [CGB95, Dra92, GP90, GPW90, GP91b, KBDZ98, KS94a, Ptá96, SMB92, CD99, Daz97, LR90, Ška90, vvS99]. extrémale [Daz97]. Extreme [AI96, Laf98a, Pat99a, Ser98b, SM93, Bor95a, CS98b, Kni99, NP96].

F [Bha97a]. Face [HS97b, SHJK97]. Faces [BS94b, Hwa95, Lew98, Mar94a, Mar94b, Mar94c, Mar96, Prz95, THW96, Zie93c, LL91b]. facial [AIY90]. Factor [Bt97c, BCP93, Sob96, Kri99, NO90, RZ98, tKWS99]. Factoring [Fle93]. Factorization [AB92, Bar94, BK96b, BK96c, BLL92, CG92a, CALM97, CG97, – Dok93, EL92, Est91, FH96a, FM97a, GNP97, GKL96, GS93, Hen97, HC98c, IOv98, JOv95, Kat94, KT95b, KP92b, LM92, LM94a, Lan96, MM94, MM93c, NP92a, Nie93a, Pan93, PD97, PL95, SLK93, Sla98, Ste97c, Sun95, Wan95a, Wan96a, Yip96, ABGR94a, ABD92, BKZ98, Blo99, Buo90b, Guo91, Haf95, Hig99, Hof98b, KSW99, KP95, Lom99, MN91, RR90a, SW98b, XK94, vvS99]. Factorizations [BH92a, Bha94b, Bot96, Cho93a, CR94b, Har95, HLY92, LR95, MP98b, RRR93, Sun92a, Var98, BH91, CP98a, CR93, Eij91, Gay91, HLP97, JOvdD99, LR99, Not91, Van90]. Factorized [SV95, Oga99]. Factors [CHS92, FFK93a, FKK95, HAP93, Men97, Pot93, QS95, Thi95, Fre91b, FKK97, GdH99]. Faddeev [HP96, HWV93]. Faddeev-Leverrier [HWV93]. fails [Wan99]. Failure [MM94, Wai97]. False [BF92b]. Falsity [CA98]. Families [Beh94, Bro94, FGP94, FMP95, IS95, Mes95a, SS96, Fon99, GPM99a, GJMNV98, Oga99, SW90]. Family [FGP97, FC97, Nor97, Gem98, Kuc91, Mub99]. Fan [Li98b]. Fast [AP96b, BOC95, BT97b, BG98a, FHR93, Gem97a, Kni95, MHR93, Ohs95, PST98, SLK93, SK95, BCCR99, BLV98, BKO99, HR98b, HR98c, JL98b, KV98, RV98, Ste90, SL97]. FE [KKY91]. Feasibility [Kiw95a, Kiw97]. Feasible [Li93b]. Fecundity [Kir93]. Feedback [BGBMN99, DR93, HA94, HFK94, HAPSG95, Log98, Rat97, Sil97c, SZ93b, Zab99, van94, CHASG98, CMN99, FP99, II98, LWX99, LWYS99, RZ98, ST93]. Feedback-Controlled [Log98]. Fermat [Qin96]. Fernando [PD97]. Few [KN95]. Fewer [Wil93b]. FFT [CG98, Ng97b]. FFT-Based [Ng97b]. Fibonacci [LLS97]. Fiedler [SW98b, Vav95, YW99]. Field [BM92, Chi95, Hua96a, Wan93b, Wan93a, Bot99, KK90]. Fields [AEG93, BGM97, CZ96, CPR93, CP96, ERS95, Eie93, Fle97, Mes95b, Nie93a, Sta93, Wat93a, BP91b, Cho98, ER90, Ger98, Ote90, PZ90b, PT99, Win99]. Figures [HT93]. Fike [Pry92]. filiform [GJMNV98]. Filtering [SC94, DB99]. filters [DF98]. Finding [ACN92, CFJ^{+}96, Erx90, Meh91]. finis [Mos96]. Finite [ATST95, BGM97, BWS95, BvR^{+}97, BM92, CHX92, CP96, EL92, EI97, FRLM97, Fle97, Gem97a, Gra96, Gu99a, HAP93, HK93b, HLMT98, Kes92, Koš94, Kri95, Lyu95, Mes95b, MM96a, MM96b, Moo95, Neu90, Nie93a, Nie97a, PR94b, SJ96, Spa97, Wan93b, Wan93a, YW90, dJ95a, Adu99, AP99a, Bha99, Bog91, BDV98, CP99a, Cho98, Cig99, Fri97a, Ger98, KK90, Mos96, ORvE99, PT99, RR99b, San96a, ST93, Win99]. Finite-Dimensional [Gra96, Lyu95, Koš94, Bha99, Bog91, BDV98]. finite-horizon [ST93]. Finite-Moment [Gem97a]. Finite-State [MM96a, MM96b, Nie97a]. finitely [Rya91]. Finiteness [LW95, Mub99]. Finland [BS92a]. First [BHR91, BB92, Bha94a, Fri92, MCK95, Now96, HKS96a]. First-Order [MCK95]. Fischer [Pat97a]. Fisher [BM94a, Gam97, KS96]. Fitting [AGR92]. Five [DIO97, SJ95]. Five-Dimensional [DIO97]. Fix [Kye95]. Fixed [GKP92, ID92, JLOvdD96, Wes96, Auj99, LS99, LT91b, Mal98, WZ98, WLL99]. Flags [Dia96a]. Flats [DK92, Coo97]. flexible [MVKR99]. Floating [VS93]. Floating-Point [VS93]. Flocks [ST95]. Florida [ORS94]. flow [BKZ98]. Flows [ADH97, Chu95, Fay94, HM92a, Lag92]. following [dRT91]. fondamentaux [Téc97]. Forbidden [DF95, Die90, Sha98a]. Forests [BC92, Sha91a]. Form [BD93, BM98b, Cla92a, EI97, Fay92, FRLM97, FL98, FGP97, Fie96, FH93, GT97c, HAPSG96, HM95b, Hua96c, JTU97, Kam94, KMS96, Kur97, LZW97, Nor97, Rus95, Sch96a, Sil97b, Wan94, BF98, BH97, CF99, DD90, El 98, FHP90, Fri91a, GV98, Han90a, HPWZ98, Hon90, Hon91, HH91, HR90, JS91, New97, Sha99b, SAZ91, SW91c, SL97, Wer90a, ZT99, GKS95]. formal [Bal99]. Formalism [Wil93b, Dor98]. Formally [FZ93]. Forme [CF99]. formes [El 98]. Forms [AC97, BVB93, BRL98, BMTZ98, CMLS95, DdlP96, DE95, GM95a, GPRL94, HK97b, KPGP97, Kuz97, LW96, Leo94, MP95, NM96, OS96b, Pis97, Rho93, RS92, ST96a, Sen93, gS96c, Wat95a, Wim93, AD90a, AM91, AMM99, BW91, BZ98, BA– DR91, BR99, BRMM91, DR91b, Fri91b, HO98, HP91, Hun90, KKS90, Kar99, Kov99, LMX99, Sha91b, Sto98, TB97, Wat90a, Wat90b, Wim91, ZJ91, HM99]. Formula [Bai97a, FHR93, Kan93, LV97, Mal96, SWP97, Ste93, Wie97, Zen93, Bog91, KV98, Mic91, Tre90, Wer90a, von90]. Formulas [AD96a, CY97, GH95b, Huc98, KT93, She96c, Haf95, RV98]. Formulation [MM96a, MM96b]. Forney [BRW94]. Four [LF93, Wil93b, van95a]. Fourier [AMM97, Mil95, PST98, Ste90]. Fractals [Wen96]. Fraction [BK97a, RAB93, BG98a]. fraction-free [BG98a]. Fractional [ERS95, ETS90]. Fractions [Bau92, KS92, Mos96, nCjH98b]. frames [FP99]. Framework [PS94, Lew99]. Francis [Bat94]. Frechet [HN99]. Fredholm [Had94]. Free [ATZ95, Gee94, HAP93, Jel92, BG98a]. French [CF99, SC98, Abd95, ABM93, BC91b, BC95b, BM95, Col96, CM96, Daz97, DC93, El 98, Fli94, Le 94, MMO91, MV95, MO91, MO95, MO97, Mos96, Oua91b, Oua91c, Oua91a, Oua96, RSDT97, Téc96, Téc97, Var96]. fréquences [Col96]. Frequency [CBSH94, RAB93, YA94, Col96, Ric90a]. Friedrichs [Put98]. Frobenius [AC97, Bap98a, DR99, ERS95, Lyu95, MOPvdD99, Rum97, She96c]. Fu [Mar95b]. Fulkerson [CKL96]. Full [BCT93, FFK93a, FKK95, SL96, KS99]. Full-Rank [FFK93a, FKK95]. Function [BC92, Bap93, CD92, CH92, Don92, Fau92, GH92, Mar94b, MP98c, Ser97, gS97b, BHL99, CN99a, Fre91a, Nar91, Sti91, da 91, van98, IW93]. Functional [EHP90, Mil95, BT99, HM90a]. Functionally [ZM92]. Functionals [Car96, Ota97, GM98b, NF95]. Functions [AK97, ABGR94b, AB95, AB96b, Auj93, AV96, BK96a, BK96b, BK96c, BL95, BM92, DFF98, Dru92b, Fay94, FMP95, Fuh94b, GKv93, Gon97, Gov97, Har93a, HS95, HF98, HF96, IKF^{+}93, JOTV93, Kal96, Kra93, LR95, Lei97, LR93, MM94, Mat93a, MP97c, Mic97, Pat97a, Pat97b, Ptá96, Que94, RRR93, Rol96, Shm94, SV94, Sun94, gS96b, Thi95, TFV94, Wat93b, Zie93a, de 96b, Ala99, AD90b, AD90a, ABGR94a, BGR90, BR90, BR91, BKGK94, Blo99, CL99, Dav99a, Dav99b, EH98, FN98b, FFK91, GKR90, GM90, Hon98, HZ99, KT93, KSW99, Kat99, KM94, Lei99a, LR99, MLS98, Neu90, Pat99a, Pat99b, RR90a, Ric90a, Wu97, WSCP91, Zie98]. Fundamental [PS97, Sch94, Bal99, HKS96a, KT91, Téc97]. Fundamentals [Nas92]. Further [BM96, JS98a, PPS96, PS96a, SV96, SBKK98, BH90, MV93, Wer91, Pus90]. Furuta [FFK91, FFK93b, Fur95]. fusion [Bha99]. Fuzzy [Cec96, AS99].

Gains [HA94]. Galois [Rea97, Wim94b]. Galton [Car97]. Games [FR95a, FR95b, Kap95, JJPT90]. GAOR [Wan94, Wan97b]. Gap [Fei94a, McE93, Nak97, AP99b, Hab94]. Gaps [Raz97]. Gauge [Mar94b]. Gauss [Sch99a, BP90b, Cop92, Hof98b, KKNU97, WY96a]. Gaussian [dR98, Coo98, NP92b, ON96, Pap92, Par98a, PD92, PN91, PN92, SV94, Tuc96]. GCD [BL92, Hau92]. Gelfand [SWP97]. General [AB94a, BM94c, Bol97, BvR^{+}97, nCZpZ96, nC98, – Dok95, Fle97, FZ93, Gee93a, Gee93b, Gee94, Mar92, Mia91, Rao96, San96b, Sch95, Wan94, WY96b, WY96a, BP90b, FS99a, Han90a, Her99, JS99a, NS99b, Rao91, Wan97b]. généralisée [SC98]. generality [PW94]. Generalization [AD93, BB93, Far98, Kim95, KS94b, Næv97, Ste90, AD90b, AD90a, BC98a, EG99, LS90b, Mic91]. Generalizations [Fur96, HDB99, HM92b, MZ91, RZ92, SG95]. Generalized [Adu98, Adu99, ABD92, AM95, BPY92, BP92a, BM96, BRW94, BRP90, Bap94a, Bap98b, BoG92, Bar94, BC93, BM93a, BE93, BGS92, BDV98, BN93, Che95, CN97, CZ97, Chi92a, Chi94, Chi95, Ded97, Di 97, DFF98, DH97a, Dra92, EK93, Ebi95, Els95, EHL94, Fay92, FM93b, FJAK96, Gam97, mGhW92, GhW96, GS95, Gem97a, Gen96a, GL95, Goe96, Gon96, Gon97, Gou93, GT97a, Gro97b, HH94a, Hei95a, Hei98a, HS90b, Hua95, Huy96, Ikr94, IO96, Kar97, KZ93, Kat94, KP92b, LW95, LLS97, Lei97, LT95, LM98a, Mae96, Mah92, MH97, MS91b, MQ95, MNST95, MM98, MP94a, MP94b, NV95a, Pat97b, Pay95, Pie96, PRB91, PB92, Pra94, RR96a, Rao91, Rao96, Ree93, Rie97, Ros93a, Ros93b, SK95, SB97, SH98, SWP97, Shm94, Sku98, Son97, SSS97, Su96, SSHS92]. Generalized [TZ96, Tia94, Wat93b, Wei97a, Wer94, WY96a, Xin93a, Xin93b, Zha97b, ZL98, ZW98, Atk93, BNS90, BMH94, BT99, Bor95a, Che90b, CX98b, CH99a, CH99b, Dai99, Dav99a, Dav99b, DW98, Erx90, FLL97, FP99, Gu99b, Her99, Hun90, Hwa99, Ikr95, Jia96, JN90, KTK99, KP95, Lei99a, LZ98a, Li91a, LC91b, Liu99b, MR98a, Mit91, NP96, Pie91, Ric90a, SC98, SW91c, Tre99b, VB97a, WZZ99, WMP98, Wei99, Wer90b, Wu97, Zha91]. Generalized-Order [Goe96]. Generated [GES96, Spi94, LL91a, Mub99, Rya91]. Generating [Le 98, Ser97, CN99b]. Generation [GT93]. generator [KPS91]. Generators [AE93, Jam91a, Koh97]. Generic [ATZ95, Bt97c]. Generically [Gut93]. Genetic [CM96, Cos95, CMH90, MO97]. genétiques [CM96, MO97]. geodesic [Pap90]. Geometric [AMPV97, ACS97, BC98b, CG94a, CYYL98, D’A97, Els95, FPP96, FP97b, Gee96, Hor95, Kar94, Kit92, Kru95, LRT99, Mat93b, NP92b, PN91, Rav97, BTT90, Fur94, GL91, Mat95, PW90]. Geometrical [BCHS97, Gus99]. Geometries [Kes92, Moo95, YW90]. Geometry [CDS95, De 92, GN97, HN93c, MBI95, OSiA96, Tab93, Yan93, Gro91, Har99a]. German [Gab97, Pus90]. Gersgorin [NT94]. Gershgorin [Sal99, Far98]. Gibbs [Har99b]. Given [DE95, HAP95, KU93, Lon95, Lon96, Sch93, SM93, Tak96, BKGK94, CF99, Gat91, HP91, TB97]. Givens [Ste97b]. GKK [Hol99]. Global [Bt97c, GN93, GV98, LT97b, Man92, MR94]. Globally [GH92]. GMRES [SG96, VV92]. GMRES-like [VV92]. Goals [KY93]. Goethals [CM95]. Gohberg [Har90, Huc98, Ion98, MO96]. Gohberg-Kaashoek [MO96]. Goldbach [Qin96]. Goldberg [Ano96b, Taş95]. Golden [AH94b, And96b, HP93]. Golden-Thompson [AH94b, And96b, HP93]. Gonzaga [Ikr99]. Graded [Le 98, Dru90, TS99]. Gradient [HM92a, BGS99, O’L91, Str91]. gradient-type [BGS99]. Gradients [SV96, Jam91b]. gradings [HPP98]. Gram [Bjö94, Cap98b, GMJ95, HL99, Kiw97, Le 94]. Gram-Cauchy [GMJ95]. grammar [Par98a]. Graph [ATST95, BBD^{+}97, Bro95, CC96, CV93, DS94b, DMB95, FOV97, Fei94b, Her93, Li92b, Li95a, LZ97a, Neu96b, Nyl96, PPY93, PPY95, Pol92, Rei94, RR97, Rob95, Shi96, Ste97b, Wat94, vLS95, van94, AdlP98, AP99b, Bap99, Bar98, Fri91b, Gro91, GKS94, Hab94, Hua91, Li91c, LZ98b, Li98c, Mer98a, Mer98b, Pal90, RSDT97, Stu90, Yon99]. Graph-Theoretic [Rei94, Ste97b]. Graph-Theoretically [RR97]. Graphical [HDTR99]. Graphs [And95a, BS95b, BMV95, CY95, CRS93, DG93, EHJL97, Far95a, FV97, Fri92, GK95, God97, Hae95, Hae96, HLM96, rKMMR95, Kir95, Lau97, LMMR95, LMR95, LS95b, McC93, Mer92, Mer94a, Mer94b, Mer94c, MP93, Moo95, Pee95, Row95, SB97, WN95, dJ95a, van95a, van95b, BC91a, BL91a, Bel91, Bel99, CGV98, CIM98, CRS99, Fio99, GZ98a, HDTR99, Her94, HS99, JR99, Pan99b, RSDT97, Row98, Sch99b, SH98, Spe95, Ste99, Tho96]. Grassmann [Dav99a, Dav99b, KL96]. great [LS99]. Greatest [Don92, GV96, Liu96, Bha91]. Green [YZ95]. Grids [Ott93]. Griffel [BB92]. Gross [CKL96]. Group [CL96a, CH96b, CKN96, Chi95, Chu93, Coe93, CW95b, – Dok95, GM97b, Gou94, KN95, Lim93a, LFMT94, Mah96, MT94, MW95, Nie98c, Par96, PRB91, Wat92a, ZPJZ98, Ell94, Gad90, GR90, GP91a, Hei97, KN99, Lei99b, NW91, Rya90, Wan97a, Bou98]. groupes [Mos96]. Groups [AP94a, BWS95, BKN97, Cam95, – DL94, Fle97, Fri96, Fri97b, Hum97, Idr96, Nie98b, Par96, PR94a, Sim92, Wil93b, AMM99, BL94b, Cig99, FJK98, Gur94, Jam91a, JW99, Lev93, Lev94, Lev97, LZ99, Mos96, Thǎ99]. Growth [Tuc96, WMD95, Art91, WLN99, Zha93]. GSVD [Han90a]. Gudkov [Szu95]. Guralnick [DG95b].

H [BB92, KP92b]. H-Matrices [KP92b]. Hadamard [Mat95, AD96b, And95b, And96b, ACN92, CN99a, Che92, Cow90, CDS95, CFJ^{+}96, Dru94a, Dru96, EH98, FM95, FM96a, HM92b, Hor95, HKL96, Im97, Kra91, LC91a, Li98d, Mat93b, Neu98, Sto99, Vis98, Wat95b]. Hahn [HR93b]. Hahn-Banach [HR93b]. Haifa [BGH92b, BGH92a]. Half [ATST95, Bol97]. Hall [Cho93b]. Hamburger [nCjH98b]. Hamilton [van95b]. Hamiltonian [AM91, AP99b, BF97, Die96, Die98, FMMX99, LMX99, RW94]. Hamiltonian-Symplectic [RW94]. Hamiltonians [HKH97]. Hand [FM97b, HH92, GW99]. Handelman [AD97]. Handicap [Väl97]. Handling [Sea97]. Hankel [Adu99, BBB95, BP93, BLL92, BLV98, Boz95, CZ95, nCZpZ96, CK91, EL92, FH96a, Fuh91, Gem97a, Gem97b, Gom95, Gou93, Gou97, Gu99a, Hei92, HJ92, Hei95a, HB97, HR98b, HR98c, HB98, HW93, KV98, Lab92, LBC95, Mes95b, Mic97, PV93, Sal98, SLK93, SO94a, Shm94, Spa95b, VK98, Vav91]. Hankel-like [BP93, Lab92]. Hankel-Norm [SO94a]. Hans [Bru99c]. Hard [Kha96, Roh95]. Hardness [CGV98]. Hardy [TV99, AB95, Cow90, WY98]. Harmonic [AMPV97, Kra98, Mat93b, MP96b, Mat95]. Hartwig [BM94a]. Hausdorff [Fri98a, Kes99]. Having [CF97a, KN95, Pee95, PR94b, Gro98b, Lim91b]. Hecke [Jam94]. Height [BCT93, Her92a, AH90]. Heinz [FN95b]. Held [BGH92b, LS95b, BS92a, BMH94]. Herbert [Sha92]. Hereditary [AHS98, Est91]. Hermite [Chi92a, Chi92b, Chi94, HDB99, Hol96, HR90, Sto98, Sun94, US92, Wye97]. Hermite-Biehler [HDB99]. Hermitian [Mar95b, Blo99, Adk91, AD90b, Aya99, BNS90, BL93a, BKdP94, BA– DR91, Bha99, BL91b, Cai99b, CWL99, CN93, CH99b, CD96, CG93a, CG99, Dan92, DST98, – Dok93,

– DL95, DR92, Drm96, Dru92b, Fal97, FL96, FM95, FP97c, Fre91b, FH93, FM97b, FK93, FM93d, Ghe92, GMJ90, GWS99, Gro99b, Gut93, HW92, Hon90, HH91, Hu96, HQ98, JZ97, Kar99, KW95, KSV91, Li97a, LSS98, LM98a, LZ97b, Nab96, NY91, NY93, OK96, Pat99a, Pet99, ST91, Ser98b, SH97a, SV95, Sla98, Smi92, So94b, gS96b, Tho95, Til98b, Tre91, TS99, TZ99, VS93, da 98]. Hermitian-Preserving [SH97a]. Hessenberg [KR96a, Xu91, AM91, AH95, BM98b, Chi93, Cla92a, Fay92, Hwa99, KR94, Kuz97, Ols98, gS96c]. Hessenberg-Triangular [gS96c]. Heterogeneous [BMTZ98]. Hiai [Fur95]. Hidden [AB96b]. High [GTVV96, Tho92, WF92, WmL97, CP99a, D^{+}99, KKY91]. High-Codimensional [WmL97]. high-order [KKY91]. Higher [Boz95, DV95, Gee96, Hau92, KT95b, LSS98, OS96b, TB97, Hei98b].

Higher-Dimensional [Hau92].

Higher-Order [DV95, Gee96, TB97].

Highly [de 96a]. Hilbert

[AS94, CWX96, CX98b, DH94, DH97a, Dru90, Lin98a, MG99, Wer96]. historical [Har99a]. History [Lun95, PW94, Wim90]. hitting [Pal90]. Hochschild [Hap97]. Hoffman [BB93, Els93, HRW92]. Hoffman-Wielandt [BB93, Els93, HRW92]. Holens [Wan99]. Holens-– Doković [Wan99]. Holomorphic [FGP97, Thi95]. Homeomorphism [Ral93]. homogeneity [AIY90]. Homogeneous [Von93]. Homomorphisms [PŠ98a, KB99]. Homotopy [dS95]. honor [LS95b]. Hook [CL97a, CL98b, CN99a]. Horizon [Gee97, ST93]. Horizontal [SG95, TT95]. Horn [Uhl92, LRT99]. Horvitz [DSW96]. Householder [Ano95a, Ano98c, BNP93, Dru99b]. Hua [EK99]. Huang [BS98]. Huard [Hof98b]. Hull [LTZ97]. Hulls [MT94]. Humbert [Mat93a]. Hurwitz [BM98a, Gen96a]. Hwang [KA95]. Hybrid [BW97, Bai97b, HFK94, Neu95, Dax90c, Gem98]. Hyperbolic [BOS93, Fri96, Lyu95, Zha96, Atk93, Lev98, SV99]. Hypercube [AP96a]. Hyperdominant [KPRV97]. Hyperellipsoid [WM94]. Hypergeneralized [GT97a]. hypergeometric [Ric90a]. Hypergraphs [CP95]. Hyperinvariant [Wu92b]. Hyperlat [JS90a]. hypernormal [Sin90]. Hyperplanes [BBLP95, Kes92, AGK^{+}98]. Hyperpower [CH96a]. Hyponormal [FI92]. hypotheses [von90]. Hypothesis [Car92, Roe90].

Ideal

[HAPSG95, BK91b, Liu97, Mol96, ÖA91]. Ideals [Cla92b, GMŠ98, HW91, VT99]. Idempotence [Esc93, Liu97]. Idempotency [Gro99b]. Idempotent [BJS97, CZ96, Hua96b, WW94, BP91b, GH91, GH95a, HO91, SG99, Wu90]. Idempotents [LMR94, Rob94, Rob96, Wan95b, ZPJZ98]. Identification [Bt97c, TZ91]. identités [ABM93]. Identities [Bon97, Cos95, DdO94, Eğe95, Fle97, Gol97b, Ilt97, Sak93, Sim95, ABM93, Müh90, RS90]. Identity [BM94c, HN93b, Lan92, Ma93, Mur95a, Mur95b, Zen93]. II [BC95b, CG94b, FL91, GhW96, HPP98, HB98, Kap95, KKY91, Koš97, Lal98, LS95a, MM96b, MP94b, NY93, NP92b, PRB91, Rao91, RV95, Ruh94, Scu93, Sil97d, Wan95a, Wes96]. II. [MH96]. III [Bap91b, LT91b, Li97c, Rao96, Roe90, Spa97, Wan96a]. ILAS [DLdO94, ORS94]. Ill [San96b, Fle90, HN98]. Ill-Posed [San96b, Fle90, HN98]. ILU [Aya99]. image [DA90, Fle90, KN98, Rob98, SC98]. Imaginary [CALM97, WmL97]. Imaginary-Axis [CALM97]. Imitations [AMM97]. immanantal [Mer91b, CL97a, CL98b, CN99a]. Immanants [Coe96, CD97, Duf94, Gre92, Jam94, Pat97b, Pat94, Pat99b]. Implementation [SS91]. Implications [Ach93, Ste97b, Har99a]. Implicit [Gee94, Gee96, Ter98, DB99]. Implicitly [BF97, Jia99]. Improved [AU99, BP93, HLP97]. improvement [Li98b]. improvements [BS91]. Improving [KKNU97]. Impulse [Gee96]. in-core [Gay91]. Incidence [GKS95]. Incident [DHM95]. Include [Has96]. including [BS92a]. Inclusion [Fuj92, Gur95]. Incomplete [KKY91, KP92b, Buo90b, Eij91, EKN90, KP95, Not91]. Incompressible [HX92]. Inconsistent [Fra95, Sid92, HN90, Sha99a]. Indecomposability [de 96a, Liu99a, SGN99]. Indecomposable [CG93b, CG94b, Hol97, Osa91, Osa93]. Indefinite [ABBN94, BAG96, BY95c, BR95, BvR^{+}97, HN94, HS96a, HS97a, Hol97, SSS97, BAG99, Fre91b, PC90, TS99, VK98, vdMRR99]. Independence [BCP93, HK97b, Cra99, JS90b]. Independent [BGM97, Jel92, WM90]. Index [Ano97, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, CZ97, Had94, McD94, Ste93, Sto99, Ano94a, AdlP98, Bel91, Cro98, HP99, MLS90, VSB90, WW91, Ano95b]. Indian [BMH94]. Indices [Fri97b, LW98a, SDT97, mJS91, Jia96]. Induced [AO91, BF95, CLLT98, Gol93, Rei93, AR99]. Inducing [Nie98b]. Inefficiency [WAS97]. inégalités [Le 94]. Inequalities [AMPV97, AK98, AH94b, And94, And96b, ACS97, AB96b, AS98, BL94a, BLS97, Bon94, Bro94, CHL95, CL97a, CL98b, CHR97b, Dru96, Fal97, FM96a, FM96b, FN95b, Fur96, GMJ95, He94, HR94, HRS98, HZ99, HC98b, Kit94, Kiw97, KT96a, Kol98, Lal98, LM97, Lin98a, LC92, LN97, LO93, LB96, Mah92, Mal92, MQ95, MNST95, MT94, MP94a, MP94b, MPŠV97, Nie98c, NM96, Pat94, Pat97b, PPS96, SSS97, Ton94, Vel95, WG93, WX97, WAS97, AB90, AB91, BHL99, Bou99, BT95b, CN99a, EHP90, Gil95, Gus99, HM90b, Hou91, Kom90, Le 94, LZ98a, Liu99b, LW99, MS91b, Nie99b, Pat99b, WZZ99, WXZ99, Zha99]. Inequality [AD96b, AGL98, AH94a, BM94b, BD95, Cha96a, Che92, FFK93b, Fur95, Gee94, HRW92, HP93, Hor95, Ikr94, JZ96, KNS98, Kit92, Kra98, LL93a, Man92, Mat93b, McE94, MT93, MT96, MP95, MP96b, Neu97a, Pat97a, Sch95, Tam97, WY98, Wan98a, WY96a, Xu95, Auj99, Bha97a, CPR90, FP99, Fir90, Fri90b, FFK91, Fur94, Ikr95, MPdS97, MS98, Mat95, Sch91, hSyLtP98, Ška90, SS98, TV99, wWS99, WI99, Wer90b, vD98]. inequality-constrained [Fir90]. Inertia [BAG96, DJ95b, DeA96, LP94, LP95b, da 98, BC98a, Dan90, Dat99, Gem98, HC98a, Loe90, Loe97]. Inertias [Uhl94]. inexact [SP99]. infeasibility [CD99]. Inference [Jen96]. Infima [MG99]. Infinite [CH99c, CR94b, CG97, EF97, Gee97, GKL96, IKF^{+}93, IAMC96, Lei92, OST93, Pta92, Ptá95, SJ96, TV90, TV92, Art91, BL94b, DL92, DL01, FL91, MPdS97, Str98, VPV99]. Infinite-Dimensional [CR94b, CG97, OST93, Str98]. Infinite-Horizon [Gee97]. infinitesimal [KPS91]. Infinitesimals [Idr96]. infinity [GKR90, van91a]. Influence [Tyr94]. Information [KS96, O’N93, AS99, Jin90]. Ingram [GPPS94]. Inheritance [BC91a]. inhomogeneous [HR98a]. Inner [Fuh94b, LMZ97, SH97a, Uhr96, Mit99]. Input [KP94, Ter98, CLR99, Nie90]. Input-Output [KP94, Ter98]. Inputs [RR96b]. INS [PSS97]. Instability [SW95]. Institute [BMH94, BGH92b]. Integer [Bal95, BH91, BH92a, Far95a, BK91b, HR90, JS90b, Sto98]. integer-valued [BK91b]. Integers [BS95a, PL95, NT91a, Ple95, RR99a]. Integral [EG93, Gou97, JV92, LR94a, Mer94a, Mus94, BRP90, JL98b, PRB91, Rya90, Rya91]. Integrals [Now96]. interbattery [NO90]. Interchange [GN97]. Interesting [Wie97]. Interior [Mor91b, Sun92b, DBRW91, KPY91, LMS91, Meh91, TZ91]. interior-point [DBRW91, TZ91]. Interlace [EH93]. Interlacement [Tre93]. Interlacing [Hae95, Smi92, Bap91a, Fio99, HJK97, MS91b, Pin98]. International [BS92a, BHR91, Da 94]. Interpenetration [Hol95]. Interpolants [BK94a, FFT94]. Interpolation [ABGR94b, AB95, BK96a, BSK94, CS97, CK94, nCZpZ96, nC98, Die94, KO97a, LR93, MH96, Mes94, MDD93, Pat97c, Ptá95, Rav97, SV94, Sun94, TK95, Wor97, ABGR94a, AA90, ABKW90, BK90, nCjH98a, CH99a, CL99, Gu99b, JM99]. Interpretation [BS98, Fli94, Fli94].

Interpretations [Gee94, GM95b]. Intersecting [Far92]. intersection [PPG91, Wer90a]. Intersections [NW92]. Interval [Abe97, AM93, AL95a, CZ91, Jan97b, LMMR95, Roh95, Wol96, Bea98, Dei91, hSyLtP98, Soh91]. Interval-constrained [CZ91]. Intervals [Har97, HJK97, Zha93]. Intractable [KLN96]. Introduction [Sha98c]. invalid [Pus90]. Invariance [BP92a, BM96, CV97, DR94, Eva95, vLS95, BM90, BP90c, BL90, Lim91a]. Invariant [ADH97, BZ95, BC91b, BC95a, Cab95, CF97a, CG99, D’A98, DS92, Gee93a, Gee93b, GV95, Gro98b, HAP93, Hor95, Kam94, LM95, Mar94a, Mar94c, Mar97, MS97, QS95, RRR93, RR94, Sil94, SW91b, Thi95, VL97, Wye97, Xie97a, Zam94, dSZ97, da 97, AD90b, AD90a, BRS91, CL90, FPP98, Fur94, GI99, GdH99, II98, LT91a, Lin98b, LST98, Ong91, RR90a, RT94, RZ98, SW91c, SW91a, Sun91, TS99, VS94, WZ99]. Invariants [BZ95, BZ94, CS92, GM95a, HAPSG95, KL96, Rei93, Wan93b, Wan93a, dJ95a, BC91b, BT91b, GPM99b, Gel99, KS94c, LMZZ98, Sha91b, Zab99]. Inverse [BR92, BJP97, BLJW95, CH96b, Dai96, Fay92, FM93a, Fie97, FK93, Gro97b, Guo96, HN98, Hei92, HH94a, HL96b, Huy96, JM93, JS96, Kar97, Kol95, Li97b, LMS95, LMOV96, MNST95, MBI93, NV95b, NV95a, Nab97, NP92a, NL96, NU97, Nyl99, Pie96, PB92, Que94, Rad96, SJ96, Sid92, SSS97, Ten90, TW95, WZ99, Wer94, Wu92a, AH95, Buf91, Che90b, CX98b, Dah90, Dai99, EHL97, Fie98, GP91a, Gro99a, Gu99b, HB90, HRW99, Hei97, Hog98a, IC94, JM12, JS99b, LC91a, Li91b, Li95b, Li97a, Mia91, Per91, SY98, WW97, WMP98, WC96, Sid99]. inverse-positive [Per91]. Inverses [BPY92, BP92a, BM96, BRW94, Bap98b, BM93a, BL93b, CKN96, CN97, DH97a, Dra92, EHL94, FM93b, Hei95a, Jen94, KN95, LM98b, Mah92, MH92, MP94a, MP94b, O’L90, Peñ95, Rao96, Rie97, Sku98, Ste89, Tia94, Tyr95, Wei97a, Xu92, BNS90, BRP90, Bap94a, BMH94, CZ90, DW98, Hun90, JL98a, KN99, LV91b, Mit91, Mit99, Nab99, NW91, Neu98, PRB91, Pra94, Rao91, RBRF91, Tia98, Wan97a, Wei99]. Inversion [Adu98, CH99a, FHR93, Gem97b, GM97a, Gor92, GH95b, Lab92, LS92a, LS94, LBC95, RV98, Sil97a, Usm94, Adu99, GS90, JL98a, KV98, Ols98, PT94a, Tre90, Xu91]. invert [Har99b]. Invertibility [Dat91, Far95b, Gee93a, Gee96, GS95, Gou93, Gou94, KR96b, LO93, SL96, TYZ97a, Chr99]. Invertible [BJ94, CL97b, Dan91a, Gur94, Lim97, Wan96b, GKW91, Lew99]. Inverting [BM98b]. Involutary [Ish95]. Involution [Lan92]. Involutions [Sku98]. Involutory [AVH95]. Involved [Shi95]. Involving [Auj97, BP92a, BM96, LN97, Mal96, Mat93b, MP94a, Pat97b, TFV94, Mat95, Pat94, Pat99b, Vis95]. Iohvidov [AD90b, AD90a, AD93]. Irreducible [Alb96, BGM97, CZ97, FV97, Kir96b, Lon99, Nie98b, OR97, SEK99, TYZ97a, At91, nM94, ZT99]. Irregular [Grö93, Raz97]. isoclinic [WM90]. Isometric [WZCZ93]. Isometries [CCK92, CLLT98, Fri96, LM97]. Isomorphic [Wu92b]. isomorphism [Fri91b]. isomorphisms [GMŠ98]. Isospectral [ADH97, Bro94]. Isotone [Ota97]. Isotropic [Rie97, BA– DR91]. Israel [BGH92b]. issue [BS92a, BMH94, ORS94]. Issues [Ano94d]. Iteration [Bai97a, CH96a, EV93, EHS93, GM95b, HN93b, HK93b, HTP96, Pie96, San96b, Su96, Dan91b, Fle90, LST98, Miz91, Pot98, Rei91]. Iterations [Mil95, SBKK98, Zha93]. Iterative [Bai97b, BCP93, Cao95, Eie93, EN93, EG97, Jia97b, Jia98a, Jin98, Kiw95a, Kiw95b, LLH^{+}98, MDD93, Sem91, Sem95, Sid92, SS95b, Str96, Van91b, WB97, Wan98b, Xin93a, Xin93b, BKM91, BTT91, CWL99, Dax90c, Dax90a, Dax90b, Egg90, GJS91, Guo91, KBŠ90, Liu91, LL93b, PI90, Wan97b, ZGL99, vV90]. IV [Bou98, MMO91].

J [Sha92]. J. [DCPS96, Saa92]. J.-S. [DCPS96]. Jacobi [AB91, AB90, Chu91, CPR93, Dav99a, Dav99b, ENV91, FMP96, Gre92, Paa91, Shm94, Usm94, ZZ98]. Jacobi-like [Chu91, FMP96]. Jacobi-Trudi [Gre92]. Jacobian [Que94, vH96]. Jacobians [Yu95]. JII [San91]. JII-SI [San91]. Job [BK96c]. John [Lun95, LS95b]. Johnson [Bha97a, Uhl92]. Joint [BL91b, Boa97, Car97, CC92b, Els94, Gri96, Pry92, BR99]. JOR [PI91]. JOR-type [PI91]. Jordan [KR96a, AMM99, AG93, ABBN94, BLS97, BRS91, CC96, CF99, DE95, DPPR90, FRLM97, GS99a, GT93, Her93, HH91, HM99, Hua91, Ilt97, Jia94, JS91, JTU97, JTU98, KR94, Kru97, LZW97, ME98, Nor97, Oua91a, PŠ98a, RR97, RR98, RS92, Sil97b, ZT99, ZD92]. Jordan-Block-Size [RR97]. Jump [BH92b, CL98d]. June [BGH92b]. just [Väl96b].

K. [Mit97]. Kaashoek [MO96]. Kac

[Gra96, TT91, Tre99b]. Kaczmarz [HN90]. Kahan [hCjXL96, Xie97b]. Kalman [AB96b, DF98]. Kalman-Yakubovich [AB96b]. Kantorovich [Gus99, LN97, PPS96]. Kantorovich-Wielandt [Gus99]. kappa [HW92, HW92]. kappa-Hermitian [HW92]. kappa-Real [HW92]. Karmarkar [Tur91]. Kerdock [CM95, Kan95]. Kernel [AD93, Fau92, HJ92, KP94, MO96, SC98]. kernels [Hun90, Ric90a]. Khatri [Liu99d]. Khinchin [vD98]. kind [HKS96a]. Kippenhahn [LSS98]. KKT [gS99b]. knapsack [PYH91]. Knot [TK95]. komplexen [Gab97]. König [Bap94b]. Korovkin [Cap98b]. Kosaki [Fur96]. KQZ [FWWJL92]. Krawtchouk [FF96]. Krein [Hob95, MR98a, Bog91]. Kreiss [SW91d].

Kronecker [DE95, DH90, KN98, KNW91, Koš96, Liu99c, LMZZ98, Spa95a, SW91c, Sun96b, Wan97a]. Krylov [CGK96, HS98, HR92, HO93, Kuz97, Le 94, Ruh94, Sid98, Sid99]. Krylov-Subspace [HR92]. Kumar [BH94a]. Kutta [CC94]. Ky [Li98b].

Lagrange [BK90, Gu99b]. Lagrange-type [Gu99b]. Lagrangian [Han90b, RR90a, RRR93]. Laguerre [Chi92a, Chi92b, Chi94]. lambda [KR92]. lambda-Matrices [KR92]. Lambek [Par98a]. Lanczos [FLL97, BF97, BF98, BLL92, Bre96, Cop93, LY93b]. Laplace [Fli94, MP97c]. Laplacian [Gro91, LZ97a, LZ98b, Mer94a, Mer94b, Mer98a, Mer98b, Wat94]. Laplacians [CL97a]. Lara [Ikr99]. Large [DS91, GLW95a, Jia97b, Jia98a, Kir97b, Mor91b, PS97, Bai98, BP99, BG98b, BC91c, Dax91, Har99b, Mai91, Roc98, Sem91, Tis91a]. Larger [Raz97]. Largest [Fan93, FN95a, Hof97, HKL96, Sha95, Soh97, An98, Fri98b]. Latent [GLW95b]. Latice [Cor93]. Latimer [Han98]. Latin [Pit97]. Lattice [LS95a, BDV98, Cho98, Lav98a, Nar91]. Lattices [DG95a, PL95, Spa97, Ste95, Wu92b, Jam91a, Ong91, Tan98]. Laurent [AG98, SS93]. Law [Eğe95, DW98]. Laws [AD93, Tia94, AD90b, AD90a, GJMNV98, Wei99]. Layer [BL93b]. LCM [BL92]. LCP [Gow90]. LCPs [Ste93]. LDL [Hig99]. Lead [ENN98]. Leading [CKN96]. Least [AP96b, CWX96, De 93, DH98, FB96, Fra95, GT96b, Gro97b, HS95, Hu95, Jan97a, Kiw95b, Ng97b, SZ93a, Ton97, WAS97, Wei90b, WY96a, WN95, Woo96, BP90a, BM90, BTT90, BTT91, Che98, DH94, HS99, Jam91b, Oga99, SRT90, WZ99, Wei98, Wer90b, vV90]. Least-Squares [Fra95, Hu95, Jan97a, Ton97, Woo96, BP90a, Jam91b, SRT90, Wer90b, vV90]. Leave [Fun96b]. Lee [Shi99b]. Left [BM91, DIO97, GKP92]. Left-star [BM91]. Left-Symmetric [DIO97]. Lei [Tam98]. Leja [Rei91]. Lemke [Bro90]. Length [WW96, Wan95b, jWySmJzZ95, WLL99]. Leontief [CP95]. Leslie [Kir93]. less [SJ95]. Level [Kri95, AH90]. Levels [Ins97]. Leverrier [Bar96, HWV93, WL93]. Levinger [AK98, Fie95b]. Levinson [FZ93]. Lewin [SN98b]. lexicographic [FY98]. Lexicographical [MLS90]. Li [Kir96a, Kir97c]. Liapunov [Car96, RR90b]. Lie [AE93, AMM99, Boa97, BT96, Cla92b, CW95b, DIO97, GJMNV98, Gra96, HKH97, HPP98, Hir90, HWZ90, HWZ92, Lag92, Mah96, MS99a, MR98b, PZ90a, PZ90b, Tur92, dRWZ90]. Lieb [HZ99]. life [Vav95]. Light [Nov97]. Like [DS91, Hil92, AMM99, ADH97, BP93, Chu95, FMP96, KO97a, Lab92, Nie93b, VV92, de 91, Chu91, Fay94]. Likelihood [BK94b, McL90]. Likelihood-Ratio [BK94b, McL90]. Limit [AGK^{+}98, Kra93, Nie97a, Tod91]. Limiting [Har93b, Sil93a]. Line [Dax90a, Kan95, Li95c, MS93a, RZ92, Bel99, DF98]. Line-Product [RZ92]. Line-Segment [Li95c]. Line-Sets [Kan95]. linéaires [CF99, RSDT97, Téc97]. Linear [AR95, AB96a, Abe97, AB92, AP96b, AGL98, LAA68, Ano96b, BPS95, Bap99, BHR91, BL93a, BL90, BP90d, BP91b, BB92, BS94a, BLL94, BY95a, BY95b, BKL97, BC95a, BLS97, BK97c, Bha97a, BS97, Boa97, BP92c, BT95b, CF96, CGP99, Car91, Car96, CL96a, Chi96b, Chi99, CL98e, Coe96, CD97, CBSH94, Cru91, CG97, Da 94, Dat94, DHL97, DCPS96, DRT93, DZ98, DH97a,

– DL94, – Dok95, Duf94, Ebi95, EGLL95, Far96, Fei94a, Fei94b, FP92, FWWJL92, Fle97, FM97b, Fri97b, FM93c, Fun96a, Fun96b, Gee93b, Gee94, Gee97, Gil95, GK93, Goe96, GH95a, Gro95a, Gur95, GMŠ98, HY96, Has96, HR93b, He94, HAPSG95, HAPSG96, HH92, HR94, HPW94, HHL92, HL96b, HO93, Hua96c, Ikr95, IW96, JV92, Jan97b, Jbi95, Jen96, JPR^{+}96, JS98b, KO97a]. Linear [KEK97, KY93, Kam94, Kan93, KL96, Kit93, Kiw95b, KT96a, Kno92, KP95, KSH94, Koš96, KLN96, KR96a, KH92, Kye95, LY93a, Lam93, LMZ97, Lan96, LS92b, LPS98, Len94, Leo94, Li93b, Li94a, LT95, Li96, Lim91a, Lim93b, Lis92, LN97, Liu97, Loe90, LP94, LP95b, Lop94, LP97, LT97b, Mal93, Mar95b, MOR91a, Mat95, MV93, MBI95, Mit97, Mos95, NP92a, Neu95, Nie93a, Nie93b, Nor97, Now96, PT95, Per98, PT94b, Pra91, PS96a, RR96a, RR96b, Rat97, Rei94, RW97, Roh95, Roh97, RR92a, Rud96, Sea94, Sea97, Sei96, Shi95, Sid92, SH97a, SS95b, So99, Spi95, Ste97a, SJ95, Sun96b, Swe93, SG95, Tab93, Ter98, Tho96, Tho92, TS97, TT95, Van97b, Von93, WY96b, Wol96, Wu92b, Wu94, YT92, YM94, Zal99, Zam94, Zha97b, Zho96, ZM92, Zie93b]. Linear [de 96b, van97a, AM98, Ans91, Ant98, Aya99, Bai98, BP90b, BS92a, BT99, Bea98, BTT90, BKM91, BTT91, BKO99, Bro91a, Bro91b, BGS99, BJW98, CM95, Cap98b, CHASG98, CF99, Che91, CC94, Chr99, Cop93, Cro98, Dai90, Dat91, Dax90c, Dax90b, DH90, Din93a, DBRW91, DSG98, ETS90, Fir90, FJK98, Fre91a, FS99a, GRO98a, GPM99a, GL91, GW99, GJS91, Hab94, HN90, HO98, Har99a, HASG90, HRZ91, Hou91, Hua99, II98, JS99a, JM99, JS90b, JZ91, KN91, KT91, KM91, LS90a, Lav98a, LV91a, LN91, LT91b, LN98, Lub98, LMS91, Mai91, MPdS97, MS99a, Mar91, MLS90, MLS91, MLS95, MP90, MR98b, MBI94, Mit90, Miz91, Mol99a, Mub99, Nie90, Nie98a, Nie99b, NF95, OY94, OPR91, ÖA91, Pan99a, PI91, Pot98, PP99, RSDT97]. linear [Ric90b, Rum91, San91, See99, Sha99a, Shi99b, Sid99, SW90, SS98, SBKK98, TZ91, Téc97, Tis91a, Tod91, Van91c, VT99, WZ99, Wei90a, Wie90, Xu99, YP91, Zha93, de 91, dRT91, vV90, von90, BB92, YM94]. Linear-Algebra [Nie93a].

Linear-Quadratic [Gee94, Gee97]. Linear-Response [FWWJL92]. Linearizations [KL98b]. Linearly [Bal95, BGM97, GO93, Jel92, LJ91, Nie97b, WM90]. Lines [Far92, Jeu95, Ols98]. Lipschitz [Cru91, KT96a, Li93b, Mrč95, PR94a]. Lipschitzian [MPS97, PSS97]. Lisbon [Da 94, DLdO94]. lists [Bro91a]. Little [Che95, Gay91]. Littlewood [App99, Aze95, MS99b]. LMI [GdOH98, dOGH99]. Local [BG94, Fan93, FGP97, FN98a, GKv93, JL98a, KL98a, Mey97, SN98a, BR90, BK91b, FP99, SS90, Xu99]. Localization [RSR94]. localizations [– DL95]. Locally [AGL98, Til98a]. Location [Gem96, Shm94, FJSv98]. loci [CMH90]. Loewner [BB96, CD92, CZ94, CH99a, FV93, FM97a, FV92, KV98, PV93, Vav91]. Log [AH94b, Fur95, Tam97, BP99]. Log-Majorization [Fur95]. logarithm [Die98]. Logarithms [Die96]. Long [AG98, Chi96a]. Long-Run [Chi96a]. Look [FZ93, KM94, TT91]. Look-ahead [FZ93]. Lookahead [Gut93, VB97b]. Loop [And95a, LWYS99]. Lorentz [Nov97, Ree93]. Lorentzian [Dru99c]. Loss [Ste97a]. Lossless [CR94b, CG97, HC98c, SK95, XK94]. Low [LMY94, Tho92, vD94, BK90, Osa91]. Low-Complexity [vD94]. low-dimensional [Osa91]. Lower [Chi93, Fri92, GKL96, HP92, Hor97, JZ97, JS98a, Kol93, Mae96, OK96, SJ96, Sil93b, Son92, Szy92, Xu97b, HS99, Hua91, YhG97]. Löwner [BNS90, Liu99c, RV95, RV96, RV98, BP91a]. Löwner-ordering [BNS90]. Löwner-Vandermonde [RV95, RV96, RV98]. LR [Fay94]. LR-like [Fay94]. LU [Hof98b]. Lyapunov [Bha97b, CHS92, CL97b, DJ95b, HP96, HTP96, KH95a, KT96b, Kra91, LR93, Lew99, Loe97, Ros93b, Vol99]. Lying [KU93]. Lyons [GM97b].

M [Mit97]. Macduffee [Han98]. Machine [Pat97b, BKZ98]. Machines [MM96a, MM96b]. MacMahon [Chu97]. MacWilliams [Gol97b, Sim95]. Magic [Jia97a]. Mahalanobis [Jen97]. Main [Min94]. Major [MD92]. Majorization [AH94b, And95b, AP94b, Bap91b, BHP97, CC92a, Dah98, Dah99a, FP97c, Fur95, GW96, Hwa99, MT94, Mos94, Mos95, Nie98c, Per96, Ton94, WC97, AP99b, Bon94, CS98b, Dah99b, GR90, Nie99b, Vis95]. Majorizations [And94, Vis98]. majorizing [KBDZ98]. manifold [Dav99a, Dav99b, Wei90a]. Manifolds [Fri96, Gov97, Sch96b, D’A98, GS99a, Rak99]. Manteuffel [AH97]. Manufacturing [Wai97]. Many [EL97]. Map [BKL97, HJRS97, LO93]. mapping [Rob98]. Mappings [CIR^{+}93, MP97a, Mrč95, Nor97, Pet99, Pit97, Chr99, OŠ93, Téc97]. Maps [Cam93, FP92, GK97, Kye95, LS93a, LT95, LW97, Lim93b, Ral93, SSHS92, TY97, VP92, vH96, BT97a, ER90, GMŠ98, Hoc94, HPWZ98, Lim91a, Liu97, Loe90, MS99a, Nie99b, OŠ91, Osa91, Osa93]. March [ORS94]. Marcus [Gro95a, BKdP94, DC92, Gro94, Kov94, Kov99, MV93]. Marcus-de [BKdP94, Kov99, MV93]. marginal [Col96, Cra99]. marginales [Col96]. Mark [TT91]. Marked [FPP96]. Markets [O’N93]. Markham [Mar95b, YW99]. Marking [Rob95]. Markov [Bai97a, BP90b, Chi96a, Flo90, For90b, GV92, GT96a, Guo99, Har93b, KNS98, KPS91, Kri95, NS99b, Nie99a, Nie97a, PV94, Sch99a, Sem95, VS99, Vil98, WY96a, Xue97]. Markovian [Hun90]. Marshall [Bon94]. Marvin [Gro95a, Gro94]. Mass [NU97, Mes97a, Nyl99]. mass-damper-spring [Nyl99]. Massive [Gay91]. master [BH94a, Chu97]. Matching [Mur95a, Mur95b]. Matchings [AB96a, An98]. Mates [Gib92, KA95, Hwa90b, Per98]. mathematical [BH94a]. Mathematician [Gol97a]. Mathieu [Par96]. matrice [Daz97, Le 94]. Matrices [AEG93, Adu98, AHS98, Alb96, ABL98, AP94a, ACE95, AH94a, AB94b, AGMP97, AVH95, ATZ95, Ata92, BJ94, BDB92, BJS97, BOC95, BH94b, BL93a, BK97a, Bat94, BP92b, BY95a, BY95b, BLD93, BAG96, BI92, Bes92, BBB95, BB96, BZ96, BH92a, BM93a, BP93, BM98b, Bon97, BNP96, BM97, BSK94, Bot96, Bot97, Bot98, BL92, BL95, Boz95, BM92, BCJ96, BGS92, BH92b, BM93b, BS94b, BH95, BHP97, BO92, BKMS97, Cab95, CZ96, CF96, CHS92, Cec96, CK97, CC92a, CZ94, CZ95, Che95, CN97, CZ97, CL98d, CC92b, Chi94, Chi96d, Chi97, Cho93a, Cho93b, CHR97a, CL98e, CF93, Chu93, Chu98, fC91, CCR97, CD97, CIR^{+}93, CD96, CLLT98, CCF97, CG93a, CHR97b, CG92b, CF97b, CP96, CGB95, Dan97, Dan92, DeA96, DR96, DG93, DE95]. Matrices [DRT93, DHS94b, Die94, Die96, Dom96, DS94c, Dru92a, Dru92b, Dru94a, Dru94b, DdO94, EL92, EGL96, EHS93, Els94, EF95, Els97, EI97, ENN98, Esc93, EHL94, EHJL97, EG93, Eve96, Fal97, FOV97, Far92, Far95b, FF96, FH96a, FGP96, FGP97, FM93b, Fie93, Fie94, Fie95a, Fie95b, FM95, FM96a, FP97a, FM97a, FP97b, Fie97, FP97c, FV92, FHR93, FH96b, FMS98, FMP95, FC97, FP96, FH93, FI92, FI95, Fri92, Fri93, FK94, FN97, Fri98a, FK93, FFK96, GTVV96, mGhW92, GhW96, Gem97a, Gem97b, GM97a, Ghe92, Gib92, Gib96a, Gib96b, Gil93, Gil97, Goe96, GO94, GKL96, Gol97c, GS93, Gor92, Gou93, Gou97, Gov97, GN97, GMJ95, GOV96, GKP92, GKP93, GKT95, Gro95b, Gro97b, GT97b, Gro97a, Guo96, GRS92, GH95b, Had97]. Matrices [Har93a, Har95, Har97, HM98, HK97a, HM95a, Hau92, Hau96, Hau95, HRR92, HT93, HLWT96, Hei94, Hei92, HJ92, HH94a, Hei95a, HST97, HB97, Hei98a, Hei98b, Her92a, HHNS96, HHS97, HW92, HJRS97, HWH95, Hol92, Hol97, HHL92, HZ98, HLM96, Hou96, Hu95, Hu96, Hua93, Hua95, Hua96c, Hua96b, Huc98, Hwa95, HKS96b, HS97b, IKF^{+}93, IAMC96, Ilt97, Ito97, Iwa98, JS93, JS96, JLOvdD96, JTU97, Juh96, KO97a, KO97b, Kal96, KPRV97, KZ93, KT95a, KRS97, KR92, KS94a, KH95b, KN95, Kir95, Kir96a, Kir97a, Kir98, Kit92, Kiw97, KS96, Knü96, KKNU97, KP92b, Kol98, KW95, Koš93, Koš97, Kra98, Kru95, Kru97, Lab92, LS92a, LS94, LBC95, LLM95, LM94a, LR94a, Lag92, LW95, LY93a, Laz95, LPS98, Lei92, LF93, cL93, LM94b, Li95a, LT97a, LTZ97, LSS98]. Matrices [LLH^{+}98, Lim93b, Lin98c, LPW92, LN97, LZ97b, LR94b, Lon95, Lon96, LW98b, LS95b, LB96, LMS95, LMV96, LMOV96, LM98b, Ma93, Mae96, MP97a, Mas94, MV98b, MOTV92, MNS93, MNST95, MOTV97, MK98, Mer94b, Mes95b, Mes95a, MM93a, Min94, MT94, Mit96, Mor91b, MV98c, Mrč95, NV95b, NV95a, Nab96, Nab97, Nag95, Neu97a, NR97, Neu96a, NP93, NT91a, New94, Nie97a, NS97, Nov97, Nyl96, O’N93, OSiA96, Ohs95, Okn97, OH96, ORR94, OK96, PSS97, Peñ95, Peñ97, PR96, Pre98, Pry92, Pta92, PV93, Rad96, Rea97, Rho97, Rja94, RS92, Rol96, Ros93a, Ros93b, RZ92, RS97, RR92b, Rum97, SZ93a, SLK93, SL96, Scu93, Sen93, Ser97, Ser98b, Sha93, SMB92, SM93, SS95a, She95, SJ96, SJ98, Shm94, Sil93a, Sil91a, Sil91b, Sil93b, Sil94]. Matrices [Sil97b, Sku98, Sla98, Smi95, SK97, SHJK97, SH97b, SS96, Sta93, Ste97c, SW91e, Sun92b, Szu92b, Szu92a, Szy92, Tak96, TYZ97a, TZ96, THW96, Thi97, Til98b, Tre93, TV92, TW95, Tyr94, Väl96a, Var98, VW95, Vav97, VS93, Wan92, WZCZ93, WL94, Wan95a, Wan95c, Wan96a, Wat92a, Wat96a, Wat97b, Wei97b, Wen96, Wes96, Wim93, WC97, Woo93, Wuw97, XX98, Xu92, Xu93, Xu97b, Yip96, YW97, Zha97a, ZM92, Zie93b, Zie93c, Zie96, da 98, da 91, de 96a, vD94, van94, AG98, Adk91, AD90b, AH95, AG94, Art91, AB90, AB91, AYC91, AZ90, Bak99, BNS90, BP91a, BT91a, Bap91a, Bar98, BM98a, BHL99, BP99, BT91b, BW91, BL90, BP90d, BP91b, BKdP94, BCCR99, Bha91, BŠ99a, BL91b, Bol98, BR99, Bor95a, BM98c, Bot99, BG98b, Bou99, Bro91a]. matrices [BC91c, BCS94, Buo90b, BCG91, CX98a, Cap98a, Cap98b, CP99a, Car91, Cec90, CJY91, CP98a, Che90a, CL98c, CLW90, Chi91, Chi90, Chu91, Chu99, CK91, CP99b, CR93, Dah90, Dan91a, DL92, DL01, Daz98, DW98, Di 97, DdPC90, DHS94a, DL99, DR91b, – Dok90, – DL95, Dru98, Dru99a, Dru99c, EK93, EFP98, EHP90, EH98, EI98, ETS90, EL99, FJSv98, FJ99, FL91, FRS99, FMMX99, FP99, Fie90b, FM92, Fie98, FGL^{+}98, FD99, Fli94, For90a, Fre91b, Fri91a, FL99, FKK98, FM93c, FS99b, Gad90, GPM97, GPM99b, GPS99, Gat91, GB99, Gel99, GI99, Ger98, GZ98a, Gil95, GR90, GS90, GKW91, GMJ90, GP90, GPW90, GP91b, GHM99, GTT99, Gu99b, Gur91, Gus91, Haf95, HL91, HNN91, HO91, Har91, HS90a, HR98a, Har99b, HWLT91]. matrices [HRW99, HP90, HP99, HR98c, HB98, HS90b, Her94, Her99, HM90a, HC98a, Hla99, Hoc94, Hol99, Hon90, Hon98, HM90b, HM99, HL99, Hou99, HJ90, HKS96a, HQ98, Hua90, Hua91, Hwa90a, Hwa99, IHM99, JM12, mJS91, Jia96, JB90, JS91, JOv95, JL98a, JTU98, KT93, KBDZ98, KH90, Kes99, KR91, KN99, Kir99, Kit91, Kol99, KS99, KP95, KLS90, KSV91, KV98, Kri94, KR96a, KR94, Kuc91, KST99, Laf98a, LM99, Lew99, Li90a, LR90, LT91b, Li97a, LHZ97, Li98e, Lin98b, LMX99, Lin99, LV91b, LMWZ90, Liu90, Liu99a, Liu99b, Liu99c, Lom99, Lon99, Lu98, MS98, Mar99a, MP98b, MOR91a, Mat91, McD94, MR99, Mer90, MS99b, MO97, Mom98, nM94, MPS97, Nab99, NDS99, NS99a, Neu98, NS99b, Nyl98, OPR91, Ols98, Paa91, PR90, Pat99a, Pav99, Per91]. matrices [PT94a, Pet99, Pvv98, Pie91, Pie99, Pin98, Pra94, PT99, PSW98, RT94, RR99a, Ri99, Rob91, Roc98, RS98, RV98, RBRF91, Rum98b, Sal99, SW98a, Scu91, Sem91, Sha98a, Sid98, Sin90, Smi90, So94b, Soh90, Son91b, Son91a, SG99, Sto99, Sto98, SW94, SW98b, SEK99, Stu99, Szu90b, Szu90a, Tam99, TY99b, Tan98, Tei99, Ten90, Tho96, Tho91, Tia98, Tis91b, TV99, TR92, Tre91, Tre99a, Tre99b, TS99, TV90, Tyr91, TZ98, TZ99, Väl91, Väl93, Väl96b, VW90, Vav91, Vis95, WW91, Wan97b, WZ98, wWS99, WXZ99, War90, Wer90a, Won90, WM90, Wu90, WSCP91, hXyY98, Xu91, Yan98, Zha91, Zho99, ZX99, ZJ91, Zie98]. Matricial [FFK96, GP94a, SH97a, FKK98]. Matrix [AL98, AMPV97, AB94a, ABGR94b, AL95a, AB95, And96a, AP94a, And94, Ano96a, AHJP91, ACF96, AB96b, AL95b, ADH97, Auj93, AV96, Aze95, BPY92, BP92a, BS92a, BM96, BK90, BK94a, BK96a, BJP97, Bap98b, BLJW95, BK96b, BK96c, BK97a, BK97b, Beh94, BG94, BG96, BGH92a, BGH92b, Bha94b, BLL92, BC98b, Bot96, BP92c, Bro94, BdV92, BCT93, CC96, Car97, CK94, CL96b, CH96b, CKN96, CN97, CY97, CH92, Chi92a, Chi92b, Chi93, Coe93, CM93, Cop92, Dah99b, Dai96, DHL97, DR91a, DR94, DE95, DZ95, DZ98, DFF98, – Dok93, DSW96, DSG98, Don92, Drm96, DV95, Ebi95, EU93, Eva95, Fan93, FN95a, Far97, Fau92, FL98, FV93, Fie96, FHP90, FV97, FN93, FS94, FM93d, FH94b, Gab97, Gał96, GRO98a, Gee94, GH92, GS96a, GKv93, GKK96, Gol98]. Matrix [GS99a, Gov94, GT93, GV95, GES96, Gre92, GKP95, GKS95, GLW95b, GT97c, Gus95a, Gus95b, HN93b, Hap97, HR93a, Har92, Hav93, Hei95b, HWV93, HS95, HF98, Her92b, Her93, Hol95, Hon92, HM92b, Hor97, dH97b, HS96b, HL96b, Hua96a, HL97, Hum97, HS97b, JT97, JM99, JZ97, JL97, JT95, JZ96, JPR^{+}96, Jøn95a, Jon95b, JL94, KO97a, KK96, KEK97, Kar97, Kar94, KL96, Kha96, Kir93, Kir96b, Kir97b, Kir97c, KS96, Kni95, KB99, KN97, Kol93, Kol95, Kop96b, Koš96, Kra93, Kra94, KL98b, Kye95, LBC95, LM92, LY93b, LMY94, LR95, LS93a, Lan92, Lau98, LW96, LLS97, LR93, Li92a, Li92b, LL93a, cL93, Li94c, LZ96, LZW97, Li97b, LM97, LZ97a, Li98a, LM98a, Liu96, Liu99d, Loe92, Lon95, Lon96, MS93a, MM94, MP97b, MS97, Mat93a]. Matrix [MP97c, MP98c, MNST95, MUV^{+}97, Mes94, Mes97b, Mic93, Mit91, MP96a, MP95, MP96b, NW92, NO98, NO90, Nas92, Neu96a, NL96, Nie97b, NM96, Oar96, OŠ96a, Par92, PD92, Pat97b, Pat97c, PPS96, Per96, Pet96, PS97, PR94a, Pie96, Put98, QS95, RRR93, RZ97, Rei93, RAB93, RR97, Rob94, Rob96, Ros93a, Ruh94, SV98, Sch98a, Sch93, Sch92, Sea97, Šem96, Sha92, SW95, Sil96, Sil97d, SG96, Sim96, Sim92, SV94, SV95, Smi92, Son92, Ste97a, Ste94, Str97a, SSS97, gS96b, Sun96a, gS97b, SSHS92, TK98, Thi95, Tho95, Tho97, Tia94, Tof97, Ton97, Tyr94, Tyr97, Uhl92, Uhl94, Uhl98, Val92, Väl97, Var93, Vis98, Von93, WG93, WX97, Wat97a, Wat93b, Wat94, Wat92b, Wie97, Wim94a, Wu92a, XB90, XL95, Xu97a, YZ95, Zha97b, ZL98, Zie93a]. Matrix [dS97, dSZ97, da 97, dGK^{+}92, van96, vH96, Adu99, AH90, AR99, ABGR94a, AMM99, App99, At91, BP90b, BH90, BPS95, BGR90, BR90, BR91, BKGK94, BT99, BEK90, Bha97a, BC98a, Blo99, BLV98, BH97, BRS91, Buf91, BN99, BT95b, CS91, CU99, CL98a, Cap98b, CZ91, CW93, CZ90, nCjH98a, nCjH98b, CH99a, CL99, CL90, Cig99, Cra99, Dai90, DST98, Daz94, Daz97, Dei91, Del90, Die98, Ede91, Eng93, ERR93, Erx90, EG90, Fie90a, FN98b, Fri91c, FL99, Fri98b, Fur94, GPS99, Gee99, GS90, GKR90, GZ94, GH91, GH95a, Gow90, GM90, Gro99a, Gro99b, Guo99, HB90, HR98b, HR98c, Hig94, Hig99, Hog98a, Hog98b, HH91, HKS96a, HKM98, IHM99, Jar90, Jin90, Joh98, JS99b, KKS90, KSW99, Kom90, KS90, KM94, LL91a, Le 94, LS90b, LR99]. matrix [LS99, Li90b, LC91a, LHE94, Li98b, Li98c, Lie90, Liu97, LMZZ98, LST98, MV98a, MMP99, MS99a, MM99, Mar95b, MP98a, MP90, Mat90, MOPvdD99, Mia91, Mic91, Mit90, Mit97, NR99, NWZ98b, NW90, NF95, NT91b, Osa91, Osa93, Ote90, ÖA91, Pat99a, RR90a, RR98, RG99, San96a, ST91, SZ90, hSyLtP98, SY98, Ška90, SV99, Sol99, Ste98, Sti91, SS98, SL97, Str98, SW91d, Stu90, SW91e, Sun90b, TT91, Thǎ99, TG97, TZ98, Usm94, VPV99, Van90, Van91b, Vel99, Vos99, WI99, Wei99, Wer91, Wim90, WC96, Wu97, WL98, WLH99, hXyY98, XWZ98, Zab99, ZT99, ZX96, ZW98, vD98, vvS99, von90, SGR99]. Matrix-Closed [Beh94]. matrix-decomposition [BN99]. Matrix-Fraction [RAB93]. Matrix-Pencil-Based [Oar96]. Matrix-Valued [CK94, SV94, CL99]. Matrix-Variate [MP97c]. matrix-vector [HR98b, HR98c]. Matroid [DFF98, Fer99, HDTR99]. Matroids [BF95, Duc95, LS95a, BD91]. Max [BP95, BO93, DD97a, DD97b, Bap98a, EvdD99, ORvE99, GM98b]. Max-Plus [DD97a, ORvE99]. maxima [DHS94a]. Maximal [BR93, CW95b, CD96, Fd92, FK94, HL92, HOŠ96, HWZ90, HWZ92, Jel92, LMV96, Mer94a, Scu91, Scu93, TW99, Yiu93, dRWZ90, von91, At91, Bel91, GKW91, RT94, Zha91]. Maximization [CH92, GHJT99]. Maximizing [CD92, FOV97, JLOvdD96, Leo94]. Maximum [BH92b, CZ97, Chi93, Gee99, Hu96, HKL96, Hwa90a, NR97, Rho97, TW95, WW96, WC97, CL98d, Kri94, LS99, Mic91, PSW98]. May [LS95b]. Maybee [Lun95, LS95b]. McMillan [BR90, BK90, BK94a]. MDS [DSG98]. me [Hof98a]. Mean [AMPV97, Car97, FP97b, Hor95, Kit92, Kra98, Mat93b, MP96b, MR98c, Ton97, Cra99, Fur94, Mat95]. meaning [Gus99]. Means [Auj97, Bap93, BM98b, EHS93, Fuj92, Auj99, GM90, HKS96a, SW91d]. Measure [HR94]. Measurement [FH94b, ST93]. Measures [Mos94, Mos95]. Meet [Hau96]. Meeting [BdV92]. Membership [Nar95]. Memory [Mir97, Coo98, Gay91]. Merikoski [Dru92a]. meromorphic [PT94a]. Method [Abe97, AM93, ACF96, BAG96, BF97, BE93, DRT93, EV93, HN93a, HP96, HWV93, HK93a, KS91, KKNU97, Kri95, Kwa97, LS92b, LY93c, Mah96, Sak93, Son97, Sta93, US92, Wan94, ZZ98, BF98, BLV98, Bro90, BGS99, Buo90b, Dan91b, Dax91, DBRW91, ENV91, EvdD99, Fri98b, HN90, HRZ91, Jia99, JZ91, Kni99, LL91b, LC91b, LMS91, MV98a, Mar90, Pan99a, SS91, Str91, Wan97b, de 91, dRT91]. Methods [ACR96, Bai97a, BW97, Bai97b, BS98, Bol95, Bre96, Cao95, CHX92, DS91, Eie93, EN93, FB96, FMP96, GM97a, GM95b, Gro95b, HY96, HRW99, He94, HS98, HR92, JM12, Jbi95, Jin98, Kiw95a, Kiw97, KP92b, LY93b, Lin92, MTS96, MMPS96, MM93c, MR98c, RW94, Rav97, Sad93, Sem95, Sid92, WB97, Wan98b, Xin93a, Xin93b, van94, At91, Bai98, BKM91, Bre99, Cao98b, CWL99, CP90, Dax90a, DB99, GJS91, Guo91, KP95, KPY91, Li97d, Liu91, MN91, Mar91, Meh91, PI90, PI91, Pot98, San91, Sid98, Sid99, VV92, ZGL99, tKWS99, vV90, Str96]. Metric [BAG96, Fei94a, HN94, HLMT98, BAG99]. Metrics [Pet96]. Min [GM98b]. Min-Max [GM98b]. Minimal [Alb96, AC97, BR90, BK96b, Cho98, EG93, FH94a, Ghe92, GN97, HN93d, HW93, HPW94, Jen94, JMS95, Kar94, LR95, McC93, Pee95, RRR93, Woe94, Yu95, ABKW90, BGR90, CF99, CH99a, DH90, FFT90, HPWZ98, So99, VB97a].

Minimality [LFMT94]. Minimally [Yan92]. minimaux [CF99]. Minimax [AS99, GP95, LS92b, Lis92, NY93, LS90a, NY91, Wie90]. Minimax-Linear [Lis92]. Minimierung [Gab97]. Minimization [Har93c, HHNS96, Jbi95, Kno92, KSV95]. Minimizing [BE93, Fan93, FN95a, Hwa95, Gab97]. Minimum [BH92b, BS94b, CMN99, GN93, Kir96b, LZ97b, MBI95, Min94, Nyl96, Pol92, Son91a, SHJK97, WZCZ93, AFK90, MV98a, Vos99]. Minimum-Norm [MBI95]. Minimum-Rank [Nyl96]. Ministrong [jWySmJzZ95]. miniversal [GPS99]. Minkowski [FM96b, Prz95, Ree93, RK96, Ren96]. Minor [JM93, MD92]. Minorant [Wei97b]. Minors [GKS95, JMOV97, MBI93, Bap94a]. Minus [BJS97, MNS93, BP91a]. Miranda [Bd97, Tam97, Tam98]. Miroslav [Vav95]. Mirsky [Nie97b, Wei98]. Miscellany [Mit96]. missing [Jin90]. Mitra [BH94a, Mit97]. Mixed [FMS98, MP96b, OST93, PS94, Sea97, CMN99, JJPT90]. Mixed-Sensitivity [OST93]. Mobius [Wil93b, Ptá96]. Modal [LN98]. Model [BS97, Fuh94a, HL95, LN97, PT94b, PS96a, Sea94, Sea97, Ste97b, Sun96b, Van97b, WY96b, WY96a, WMD95, AU99, BP90b, Fir90, JS99a, NO90, Sch99a]. Modeling [Ant94]. Models [BS92a, BLS97, BS97, FV97, Has96, PP94, Sch96a, Von93, WW94, YM94, Bap99, DF98, Kri99, Nie90, NF95, SRT90, WLN99, von90]. modes [PC90]. modifications [Pus90]. Modified [GJS91, KKNU97, Wan92, BT91a, HY91, LST98, Not91]. Modifying [HC98a]. modular [JW99, Kri94]. Module [GM97b, Jel92, MM96b, Rud96]. Modules [FNW96, Gol93, Gus91, Chi91, DSG98, HHR93, Lin99, Liu97, Rin98]. Modulo [Idr96]. Moivre [Ree93]. Moment [AB94a, AL95a, Bol97, nC98, Gem97a, Hei94, Næv97, nCjH98b, CL99, HL99]. Moments [ST96b, Ede91]. Monge [Die90, Rud98]. Monic [FN98a]. Monoid [MM96a, Rya90]. Monoid-Ring [MM96a]. monoids [Rya91]. Monomial [MHR93]. Monotone [Ata92, AV96, BAG96, Kiw97, Kra93, Man92, Mil95, Pet96, Sem95, TW95]. Monotonic [CHL95, EN93, Lav99a, MS91b, MS93b]. Monotonicity [JN91, KS94b, Lav97, Ans91, FI99, Wer91]. Monte [BP99]. Moody [Gra96]. Moore [BR92, BJP97, DS98, FM93a, GP91a, Gou94, HB90, Jen94, Mia91, MBI93, NL96, Pie96, PB92, Rob94, Sil97a, Tia98, Wan97a, Xu92]. Moore-Penrose [Sil97a]. Morphismes [BM95]. morphisms [BM95]. Morris [Gol97a]. Morrison [LV97]. Mosaic [Hei95a, LBC95]. Moser [Bai97a]. Most [Gib96b, SW90]. Motion [Fra95]. motions [SZ90]. MP [Pap92]. MR [Ano96b, BPS95, Bha97a, Bon94, BT95b, Chi96b, Gil95, GH95a, Ikr95, KP95, KR96a, Li96, Mar95b, MLS95, Mat95, MV93, Mit97, NF95, Spi95, SJ95, Tho96]. MSOR [AH97, Mar90, Wan94, Wan97b]. Much [Mus94]. Muir [Zen93]. Muirhead [Kim95]. multi [DF98]. multi-process [DF98]. Multicomponent [EG97]. Multidimensional [AK97, CF93, MM93b, MDD93, Zam94]. Multigraphs [Mic93]. Multigrid [Kri95, Ott93]. Multihomogeneous [Kal96]. multiindex [Kes99]. multiindexed [HM90a]. Multiinput [BCU94, Fun96b]. Multilevel [BW97, Bai97b, WB97, TZ98]. Multilinear [AG97, Dia96b, Mar91, Wat90a]. Multimatroids [Bou98]. Multiparameter [Bre99, Jen91, Koš94, Mal98, Ple98]. Multiple [Chi92a, Chi94, FM97b, HH92, Sun90a, Tia94, nCjH98a, GW99, HB90, Pre98, Row98, Wei99, Wie90]. multiples [BGR90, Cai99a]. Multiplication [BT97b, GO94, Had94, Kni95, Cow90, HR98b, HR98c, Sto99]. Multiplicative [BL95, Bsh95, CAHJ98, CX98a, Egg90, Hoc94, OR97, DA90, Mol99b]. Multiplicities [Kru95, MO96, PP99]. Multiplicity [Far95a, GK95, KT95b, Lub98, dGK^{+}92]. multiplied [Cai99b]. Multiplier [ACN92, AO91]. Multipliers [CDS95, CFJ^{+}96, Lac95, Liv95, Dru98, OS94]. Multiresolution [LZ99]. Multispherical [HLWT96]. Multisplitting [HY96, HO93, MMPS96, Su96, Wan98b, Cao98b, CL98a, Wan91]. Multisplittings [Nab96]. Multitype [Car97]. Multivariable [Gał96, Gom95, BRMM91]. Multivariate [Jia97a, Mey97, ST96b, Von93, WCM99, Flo90, Gam97, HN99, JZ99, Jin90, WLN99]. Multiwavelet [SV98]. Multiway [LS98]. Murdock [Tre99b]. mutation [CM96].

Narrower [Im97]. natural [PC90]. Navier [HX92]. Near [HM98, Kir93, Woo93]. Near-Periodic [Kir93]. Near-Toeplitz [Woo93]. Nearest [GdH99, BHM98, CH99b]. Nearly [LLM95, LMOV96, Pav99, Sha93, Wil93a, For90a, Li90a]. Necessary [CR94b, ERR93, Soh91, Yip96]. Necessity [HK97b]. Negative [CG93a, EL97, Ghe92, HLMT98, NS97, TV92, Jar90, Pat99a, TV90, wWS99]. Negative-Dominant-Diagonal [NS97]. Nehari [Ion96, Spa95b]. Neighborhood [JT97]. Nekrasov [Li98e]. NEPS [Ste99]. Nested [BH92a, Cao95, JMOV97]. Networks [BL93b, SSS97, CIM98, Lub98]. Neumann [Kap95, MT96]. Neutrality [LMZ97]. Nevanlinna [CK94, nC98, nCjH98a, CL99, Pat97c, Wor97]. Neville [FM97a, GP92, GP94a]. Newman [Ano96b, Gol97a, Taş95]. Newton [LST98, Mah96, Mil95, OY94]. Newton-Fourier [Mil95]. Niche [And95a].

Niederreiter [Fle93, Lan96]. Nilpotent [Boa97, BT96, vH96, App98, BC91b, EL99, Har94, JTU98, MOR91a, MR98b, MO97, PZ90a]. Nilpotent-by-Abelian [BT96]. Nilpotentes [MO97]. no [Sau90]. nodal [DR99]. Noetherian [II98]. Noise [FH94b]. Non [FP92, FM97b, Gut93, SZ90, Sta93, Til98b, Aya99, Bha99, BHM98, GB99, Gro98b, Mub99, NP96, Pat99a, Pot98, SH98, TZ99, wWS99]. non- [NP96]. non-collinear [Mub99]. non-decomposable [Gro98b]. Non-Euclidean [SZ90]. Non-everywhere [FP92]. Non-Hermitian [FM97b, Gut93, Til98b, Aya99, Bha99, TZ99]. non-linear [Pot98]. non-negative [Pat99a, wWS99]. Non-normal [Sta93]. non-primitive [SH98]. non-regular [BHM98]. non-zero [Gro98b]. NonAbelian [RS93]. Nonatomic [HR94]. Nonbarycentric [FH96b]. noncausal [Nie90]. Noncentral [Chi94]. Noncircular [MM94]. noncommutative [BK91a, SG99]. nonconstant [GdH99]. Nonconvexity [HHT93]. Noncore [Mit96]. Nondegenerate [ABBN94, FK93, FFK96, nCjH98b, FKK98]. Nonderogatory [DeA96, Koš94, Ple98]. nondistinct [Paa91]. Nonequivalence [GLW95b]. Nonexistence [Now96]. Nonextremal [BKS93]. Nongeneric [ME98]. Nonhomogeneous [GV92, Jod92, PV94, VS99]. Nonlinear [Bai97a, MM99, Mey97, RS97, Sak93, Guo99, KPY91, Oga99, Rum91]. nonminimal [Gał91]. Nonmodal [DR93]. Nonnegative [Alb96, AL95b, BT91a, BJS97, Bap98b, BM93a, Bor95b, BM97, BCJ96, CKN96, Chi91, CR93, CHR97b, DdlP96, ENN98, FOV97, FN98a, FN97, Gen96b, GT97b, Guo96, HRR92, Hei94, HHNS96, JS93, JV92, JLOvdD96, KK96, KH95b, KN95, Kir98, Kol93, Kra98, LRT97, Liu96, NW91, RR92a, RGS98, RS97, SM93, Son92, SH97b, Szy92, TK98, Val92, WC97, Wuw97, At91, BP91a, BM98c, CL98c, CP98b, – Dok90, EHP90, EH98, FJSv98, FN98b, HNR90, Her91, Her99, Hwa90a, HKM98, JB90, Laf98a, LM99, LHZ97, Mar99b, Mat90, Mer90, nM94, Nie98a, Smi90, Son91b, Yan98, ZT99, WC96]. Nonnegative-Matrix [KK96]. nonpolyhedral [BHL99]. Nonprincipal [Sil96, Sil97d, da 91]. nonregular [BR91]. Nonsingular [BY95b, BM92, Gib96a, Gib96b, GOV96, LLM95, Lim93a, LP95b, LMV96, LMOV96, MM93a, CP99b, HM90a, Loe90, Sha98a, Tho96, Wat90b]. Nonsingularity [Gil97, MOTV97, Roh90, SJ98, Szu92b, Szu92a, WW96, CZ90, Szu90b]. Nonsquare [Fau92, GS95, GS90]. Nonstationary [Cao98b, GT96a, MMPS96]. Nonstochastic [Fie95a]. Nonstrict [Sch95]. Nonsymmetric [BJ94, BO92, D’A97, HRW92, Jbi95, Kwa97, CC94, FLL97, GW99, PS98b, Ruh94, Sid99]. nontrivial [Hwa90b]. Nonuniformly [SS96]. Nonunique [MP96a, MP97d, Mit97]. Nonzero [BH95, Gib96b, FL99, LW98a, Lim91a, Sil91a]. Norm [ACN92, CLLT98, EF97, Gab97, Har96, HW93, Hor95, KH95a, Kwa97, LTZ97, LO93, Mah92, Mar94c, MBI95, MP98d, WZCZ93, Zha99, Zie93b, Abd95, Ant98, Auj99, CMN99, Fuh91, Fur94, Kri94, Mat90, SW98a, Kiw95b, MUV^{+}97, SO94a]. Normal [Adk91, BH94b, Bat94, BHK91, BR99, DS94c, Dru94a, EI97, EI98, FH93, GKS95, Hol92, HS96a, HS97a, Hol97, JS93, KH95b, KMS96, KH92, LZW97, MR98a, Neu96a, Rad96, Rus95, Sch96b, ST96b, Sun96a, SZ93b, Wat97a, Wim91, Yiu93, AZ90, CL98c, Chi91, Di 98, Dru98, Dru99a, HR90, Jin90, KKS90, LHZ97, New97, RGS99, Sid98, Sta93, SL97, Sto98, WZ98, Ral93]. Normal-Preserving [KH92]. Normality [JZ96, Bha97a]. Normalized [Nag95, FFK93a, FKK95, FKK97]. Normally [WL94, WM90]. norme [Abd95].

Normed [EGLL95, LN91]. Norms [Alb96, AG97, BP92a, CDS95, CFJ^{+}96, HHNS96, Hor95, Jen97, Lac95, Lav97, cL93, LM95, Mar94a, Mar96, Mar97, MD92, MP98d, Neu95, Rao96, Wat92b, Wat95b, Xie97a, Ziz98, AO91, AG94, Bou99, Daz98, Hla99, HJ90, JN91, Lav99a, Lav99b, LT91b, LT91a, Li94b, Li96, LWX99, MS91b, MS93b, Mer91a, Rao91]. Note [Abt94, AL95b, BP92a, BŠS96, Bha97b, CD92, CZ95, Cor93, DS92, Dan97, DD97a, Els93, Fie96, Fun96b, GL97, HWH95, Hua95, Jia92, Kir96a, Kir97b, Kir97a, Kit92, KT96a, Kop96a, LW96, Lim93b, Liv95, Mer94c, Mic97, Nab96, VL97, Xue97, Yan92, Zha96, ZG94, Ano96b, BP90a, CL98c, Dax90b, DH90, Dia93, Fri90b, Fur94, GKW91, Gro99c, HS90a, JM12, KH90, Kri99, Lav99a, Lev98, Liu90, Mar90, Mer98b, Nie99a, Oga99, RS98, SW91c, Tre90, Xie97b, YhG97]. Notes [Che90a, HY96, Lei99a, Mal92, NWZ98a, XX98, Xu93, MV93]. notion [BC95b, Téc96]. notions [VPV99]. novel [FLL97]. noyau [SC98]. NP [Kha96, Roh95]. NP-Hard [Roh95, Kha96]. nuclear [Bae90]. nuisance [NF95]. Null [BR91, HX92, Nyl98, BKGK94, Cho98, GKR90]. Null-pole [BR91, BKGK94]. Nullities [Bot96, Bot97]. Nullity [PR94b]. Nullstellensatz [BS91]. Number [BH92b, Cab95, CGK96, Cla92b, Ded97, DFF98, GOV96, JL97, KU93, MS97, MUV^{+}97, OŠ96a, Sch96b, Sil94, Soh97, WZCZ93, dSZ97, da 97, CL98d, ENV91, Fer99, FJK98, LS99, Lin98b]. Numbers [Cao98a, DF95, Ded97, EGLL95, Fle97, FN98a, GL97, Hig95a, KEK97, KH95b, LMR95, Mar92, MO96, Neu97b, She96c, dS95, BAG99, BG98b, CP90]. Numerical [CT94, Chi96c, CYYL98, Chk97, Dat94, Fie95b, GLW95a, HJRS97, HTP96, Hon92, HHT93, IKF^{+}93, KRS97, KM90, LS92b, Li94a, LT95, LSS98, MP97b, Mir98, NT91b, Roc98, Tho97, Tre91, Wat96c, BR99, Chi96b, CN99b, Daz94, Dru99a, Guo99, JAG98, LS90b, Li91a, MP98a, NWZ98b, Oml90, Ple98, QOQO98, TY99b]. Numerics [Bjö94].

O. [Str96]. obey [Par98a]. oblique [TY99a]. Observability [Fun96a, KSH94, Kra95, KL98a, LFMT94, Che91]. Observation [FM95, NP92c]. Observations [MCK95, Jin90]. Observer [DS91, Kra95]. obtain [GJMNV98]. obtained [CM95]. October [BdV92]. Odd [Tre93, Wan93b, Tre99a]. ODEs [SS95b]. Off [dS97, da 97, DHS94a]. Off-Diagonal [dS97, da 97, DHS94a]. Old [Tyr96]. Olga [Ano98b, Ano98d, Ano98e, Ano98f, Ano98h, Bau98, Hof98a, Joh98, Laf98b, Sch98b]. Oliveira [MV93, BKdP94, DC92, Kov99, MV93]. Oliveria [Kov94]. Olkin [Bon94, GPPS94, Mar98]. OLSE [LPW92, NP92c]. OMC [Dru99b]. omega [EV93]. Omitted [BW96]. On-line [DF98]. Once [Spi94]. One [BH96, Ghe92, LM98b, MH96, Scu93, TV92, WAS97, de 96a, BT91a, BCCR99, Ell94, HL91, OŠ93, Sau90, Scu91, TV90, VS94]. One-Connected [LM98b]. One-Step [Ghe92]. Ones [Wan95a, Wan96a, CL98d, Chi90, FM93c, WZ98]. Only [LS92a, LS94, Soh91]. onto [AGK^{+}98, Wei90a]. Opérateur [DC93, BC91b]. Opérateurs [El 98]. Operating [GW95]. Operation [Väl93, CF99, Miz91]. Operations [Ste97b, Tyr94]. Operator [ACS97, Auj97, BF92a, BJ94, Bha94a, Bha95, CHR97a, FN98a, FFK91, Fuj92, GKv98, GM90, GO93, HX92, Har93a, Har95, Hau94, Hou96, Jod92, JZ96, Kit94, Kru92, Lin98a, LO93, McE94, Mir92, MPŠV97, Put98, Raz92, Sak93, Wat95b, AO91, Auj99, BC91b, Bha97a, BC95b, CJY91, CX98b, CPR90, DC93, Gad90, GM98a, GMŠ98, HNN91, HNRR99, Hou91, JZ99, Jód91, KNW91, Str98, Ves98].

Operator-Norm [LO93]. Operators [AG97, ABBN94, BRW94, BS94a, BY95a, BY95b, BC95a, BLS97, BD95, BM98b, CAHJ98, Cai99b, Cam95, CZ96, CS96, Chi96c, CPR93, Ded97, DH97a, DR92, EGLL95, FW96, FK94, GW96, Gom95, HN93b, HW93, HS96a, HS97a, HHL92, ID92, KH95a, LMZ97, LPS98, Li94a, Mar95a, Met97, Nak97, Pis97, PR94b, Ptá96, Sil97a, Spa95b, Spa97, SC94, Tam92, WW94, Wan96b, Wat92a, Adu99, Ant98, App98, BL90, BP90d, BP91b, BHK91, Bog91, Bou99, Cai99a, CP99a, CP98b, Cow90, DH90, DS98, El 98, Eng95, Fri90a, Gu99a, Hou91, Kar99, Lam90, Lav98a, LR90, Lim91b, Mal98, MS91a, MLS90, MLS91, MLS95, MR98a, Mub99, Oml90, OŠ93, Ong91, Sal99, So99, Zal99, Ziz98]. Opposite [Aze95]. Optimal [AH97, BP95, Chu98, EV93, FR95a, GHN97, Gee94, HA94, Kan93, KMS96, Li93b, PS98b, ŽWB93, AG98, BDV98, CC94, Dan91b, HP99, LWX99, NWZ98a, TZ91, Tis91b]. optimal-basis [TZ91]. Optimization [Mey97, MCK95, OST93, Rol96, Shi95, HC98a]. Optimizing [DBRW91]. Orbit [Fle97, LM94b]. Orbits [BC95a, LT95, Gro98b]. Order [BK94b, BJS97, Bha94a, BCP93, BH95, CD92, DR93, DV95, Fur96, Gee96, Goe96, GGM96b, Hol96, IKF^{+}93, Jam91b, Jod92, LMZ97, LF93, MCK95, SS96, Tia94, WF92, Col96, DW98, Hei98b, KKY91, LVBL98, Lub98, MLS90, Mit91, Oua91b, Stu90, TB97, Wei99, Wer90a, WM90]. Order-reducing [Jam91b]. Ordered [Mar92, ER90]. Ordering [BK94b, BM93a, GR90, Gro99c, Kir99, Wu97, BNS90]. Orderings [GT97b, HM95a, Nie98b, BP90b, BH90, BP91a, BM91, BPS95, GHM99]. Orders [HL92, MH92, Ota97, Est91, Lav98a, Liu99c, Mat91]. Ordinary [MZ91, WAS97, ZM92]. orientation [GKS94]. Oriented [Li95a, Coo98]. Oriented-Graph [Li95a].

Origin [SS93]. originally [BS92a]. originating [fC91]. orthant [Lav99a, MS93b]. orthant-monotonic [Lav99a, MS93b]. Orthogonal [Bar96, DV95, ENN98, FB96, Fri93, Gen96a, GNP97, HF94, IOv98, KN97, KPGP97, Lim93a, MS93a, Pot93, Wan93b, Wan93a, Wat95a, DB99, GZ98a, Gro99d, HP90, HP99, HM99, HWZ90, Jam91a, Le 94, Téc96, YW90, Zha93, ZPJZ98]. orthogonale [Le 94, Téc96]. orthogonalité [BM95]. Orthogonality [BS94a, BŠ99a, DER97, FLM94, GZ98b, HOŠS96, Wat96a, BM95, GTT99]. Orthogonalization [Bjö94]. orthostochastic [AYC91]. Oscillatory [Eve96, Peñ97]. Osserman [Rak99]. Ostrowski [MG94]. Ostrowski-Reich [MG94]. Other [Kit93, Pat97b, Gad90, RR90a]. Outer [MH92, Mit99]. Output [KP94, Ter98, CMN99, LWYS99, Nie90]. Outputs [KSH94]. Overdetermined [Neu95]. Overestimations [Roh97]. Overgroups [– DL94]. Overlapping [HO98, Cao98b]. Overrelaxation [HK93a, Li97d].

P [SGR99]. P-matrix [SGR99]. Packings [BNP96]. Padé [ACF96, Gut93, XB90]. Pair [BCJ96, HR93a, LM98a, CH99b, EHL97, GdH99, Mit90, ÖA91, Sun90b]. Paired [HST97]. Pairs [AS94, DdO94, FGP97, FV97, Fri91a, Gom95, HJRS97, Ilt97, Len94, Wat95a, Wim93, WmL97, BT91b, BW91, BP90d, BA– DR91, DR91b, GZ94, Har91, Kar99, Mor97, Ruh94, Vel99, Yan98]. Pairwise [Chu98, HOŠS96]. Pang [DCPS96]. papers [BS92a, PS96b]. paracontractions [EKN90, Pot98]. Paradox [Had97]. Parallel [AB92, AP96a, Bai97b, BOC95, BP93, CHX92, FMP96, KKS90, Lau97, LY93c, LP97, Mal93, MMPS96, MP97d, MVV93, PPY93, QOQO98, WB97, Wan98b, Bai98, BKO99, CG98, Eij91, ENV91, KBŠ90, LC91b, Paa91, Wan91]. Parallelism [DK92, Coo98]. Parallelizable [Sha98b, de 91]. Parallelogram [Eğe95]. Parallelogram-Law-Type [Eğe95]. Parameter [Don92, Koš96, Mes95a, MV98c, vLS95, BGE91, Kri99]. Parameter-Dependent [MV98c]. Parameters [AB96b, Has96, WL94, Fio99, NF95, SW90]. Parametric [Hun90, Gem98, NF95]. Parametrization [AI96, Shm94, FP99]. Parametrizations [FFK93a, FKK95]. Parametrized [Don92]. Part [HK97b, Sil93b, gS96b, Tak96, Gro99b, Dav99a, Dav99b, Sch99a]. Partial [AL95a, BJ94, BJS97, CHR97a, Dan92, DS91, DER97, FH94a, Ghe92, GT97b, GHM99, GRS92, HL92, HM95a, HJ92, HHS97, Hou96, JTU97, Kru95, KS92, LM97, MH92, NW92, RS92, Sil97b, ST95, ZZ98, BP90b, BH90, BP91a, BM91, BPS95, CH99a, JTU98, KR96a, KR94, Liu99c, Mat91, Mit91, Stu90, VB97a, Wu97]. Partially [BLJW95, CD96, KO97a, Mar92, Sil94]. partite [BS95b]. Partition [Smi95, HOR98, ZPJZ98]. Partitioned [BS97]. Partitioning [PPY93, PPY95]. Partitions [God97, Row95, Sim95, Uhr96, Van97b, Woź97, Nar91]. Pascal [BT99, BP92c, Zha97b, ZL98, ZW98]. Pass [Mic97]. Path [BL91a, BL94b, KR91, Shi96, KM91, Pan99b, dRT91]. path-following [dRT91]. Path-positive [BL91a]. path-recursive [Pan99b]. Paths [Her94, WLL99]. Pattern [DR94, Esc93, EHL94, EHJL97, Kir93, Wat96a, EL99, LHE94, SEK99, Stu99, Thǎ99]. patterned [Nyl98]. Patterns [EJ93, IOv98, JM93, JMS95, LPS98, Lim93a, LMOV96, MOTV97, SJ98, EHL97, Hog98a, Hog98b, LHZ97, SJ95]. Paz [KNS98]. PBIB [Yan93, YW90]. Peak [Her91]. Peirce [BM95]. Penalty [Gee97]. Pencil [AL98, DER97, GES96, Li94c, Oar96, Pat93, BHM98, Del90, Erx90, FS99a, LMZZ98]. Pencils [BG96, DE95, FL98, Fie90b, GV95, GES96, Hua96c, Kar94, KL96, KSV95, LY93b, cL93, Li93a, Li94c, LSS98, NY93, RR97, Sch94, SS93, Ste94, Tho91, Tho95, Uhl94, Bol98, BGE91, CS91, GPS99, KM94, Li90b, LMX99, MOPvdD99, NY91, PC90, RR98, Wim91]. Penrose [Sil97a, BR92, BJP97, DS98, FM93a, GP91a, Gou94, HB90, Jen94, Mia91, MBI93, NL96, Pie96, PB92, Tia98, Wan97a, Xu92]. Pensacola [ORS94]. Pereyra [BKO99]. Pereyra-type [BKO99]. Perfect [BWS95, CCF97, SC94, An98, PZ90b]. Perfect-Reconstruction [SC94]. Performance [GTVV96, HL95]. perhermitian [Pre98]. period [Jia96, LHE94]. Periodic [BCHS97, CBSH94, Fd92, GS93, Kir93, RR92b, TK95, GL91, KSW99, Nie99a, VS94, Var96]. Periodicity [GV92, TV92]. périodique [Var96]. Permanent [BL94a, BC93, BH95, CCR97, Hu96, Hwa95, BCCR99, Chu97, HKM98, Lal98]. Permanental [Gib92, HHT93, KA95, Mal92, Sou94, Hwa90b]. Permanental-Dominance [Sou94]. Permanents [Bap93, BS94b, Dru99c, Fal97, Gol97c, Min94, Son95, SHJK97, dR98, Hwa90a, Son91a, Wan99]. Permutation [AYC91, LM95, Tei99, Wye97, SW91e]. permutations [CL90, SW94]. Permute [BP92b]. Perron [AK98, Bap98a, BCJ96, DR99, ERS95, Kir96b, Kir97a, Kir97c, Kir99, Kol93, LPS98, LSJ91, Lyu95, MOPvdD99, Rum97, Son92, TK98]. personal [Sch98b]. Perturbation [And96a, Bar94, BRL98, Bha94a, Bha95, BL93b, CWX96, DR94, Din93b, DH98, EH93, Els94, FB96, GZ94, HH92, Jia94, KNS98, KPGP97, KSP98, Kru92, Kuz97, Li93a, Li95c, Li97c, MV98b, McE93, Pan93, RS97, Sch92, Ste94, Sun91, Sun92a, Sun95, gS96c, gS97b, Wei90b, WP98, Wei98, Xue97, CX98b, DH94, HQ98, Jia98b, TS99, WW97, Loe92]. Perturbations [BH96, Bha94b, BO92, CAHJ98, DS94b, JLOvdD96, LMY94, ME98, SV95, TS97, VS93, JSOvdD94, Mal98, Rum98b, Wim99]. Perturbed [Els97, HRR92, Li93b, LS97, Roh97, CX98b, VS99]. Phenomenon [BL93b]. Physics [Nov97]. Picard [Che95, MW95]. Pick [CK94, nC98, nCjH98a, CL99, Pat97c, Wor97]. Piecewise [Die94, Kiw95b, ER90]. Pivot [Väl93]. Pivoting [Bom96, KO97a, ON96, KT91]. PL [Ral93]. PL-Normal [Ral93]. placement [CLR99, KR90]. Planar [CV93, CIM98, HS99]. plane [BKGK94, Gab97, Hur90]. Planes [ATST95, Yiu93, WM90]. Plus [BBB95, DD97a, Boz95, HB97, HR98b, HR98c, HB98, ORvE99, de 91]. plus-shaped [de 91]. Point [Chi96d, Gee97, GL95, Pat97c, VS93, Auj99, DF98, Din93b, DBRW91, Eij91, HKS96a, KPY91, KR90, LMS91, Meh91, TZ91, Wei90a]. Pointed [SH97a]. Points [AI96, GW95, HS95, ID92, Uhr92, Bor95a, CS98b, EFP98, NP96, Rei91]. pointwise [CHASG98]. Polar [BoG92, BvR^{+}97, GM97a, HM93, GHM99, Hig94, Mar99a, vdMRR99]. Polarized [Nov97]. Polarized-Light [Nov97]. Pole [BCU94, DER97, SW90, BR90, BR91, BKGK94, BK91a, CLR99, GKR90, Van90]. pole-shifting [BK91a]. pole-zero [Van90]. Poles [Sch94, Spa95b, WSCP91, CLR99]. Politecnica [BdV92]. Pollen [KT95b]. Pólya [Ric90a]. polydisk [Gu99a]. Polygons [van95c, – Dok90]. Polyhedra [Dah98]. Polyhedral [AGL98, LL91b, vvS99]. Polyhedron [BS92c]. Polymatroid [Nar95]. polynômiales [ABM93]. Polynomial [Bar96, BT97b, Bon97, BM92, BT95a, Die94, DdO94, Fle97, Gat91, HN93d, HWV93, HAP95, Hol95, Hol96, Ilt97, Jia99, JPR^{+}96, KO97b, KU93, Kar97, Kra94, LBC95, MH97, MP97a, MHR93, NP93, Sch93, Sem92, SV94, VP92, Win99, WM94, dS97, dJ95a, vH96, ABM93, BK90, Bee90, BEK90, DH90, ETS90, Fie90a, Fre91b, Fuh91, FS99b, Lin99, O’L91, Per98, Pie99, Rei91, Sau90, SL97, Väl91, Van90, YP91, ZX99]. polynomial-time [Väl91]. Polynomials [AGR92, ACR96, AC97, BP92b, BGM97, Bol95, Cab95, Chi92a, Chi92b, Chi94,

– Dok93, Dom96, DV95, EG93, EU93, Far95a, FV93, Fis95, Fle93, FV97, FN93, FN98a, FZ93, Gał96, Gen96a, Gen96b, GV96, GO93, HF94, KL98b, LM94a, LMY94, Lin98c, MS93a, MP97b, MS97, Nie93a, RZ97, Ric90b, Shi96, Sil94, SG96, Sim96, Soh97, Tho95, Yu95, dSZ97, da 97, BC98a, BK91b, Dav99a, Dav99b, – Dok90, Gem98, GS90, KK90, Lak99, Lin98b, MP98a, NWZ98b, Sch98a, Soh91, Tei99, VPV99, Zha93]. polynominals [dR98]. Polytope [BS94b, Hwa95, HS97b, LP95a, Soh97, SHJK97, Bor95a, CS98b, LSJ91, NP96]. Polytopes [Cuy95, DG95a, MW95, Dah99a, HOR98]. Poncelet [Mir98]. Popov [IW93]. Popov-Function [IW93]. Populations [Gut98]. Portugal [DLdO94]. Posed [San96b, Fle90, HN98]. Posets [BH92b, Hau96]. positifs [El 98]. Positions [TK95]. Positive [ABL98, AP94a, AH94a, AI96, AB96b, BJT93, BKL97, BM97, BCHS97, BHP97, CPR93, Dan92, FP97b, FMP95, FV96, FK93, GHJT99, GS93, GKT95, HR93a, HF96, Hu95, JOTV93, JT95, Kye95, LS93a, LP95a, Lau97, Lau98, Lei99b, LW97, LN97, LZ97b, LB96, MM94, MV98b, Nab96, Neu97a, OSiA96, Peñ95, SZ93a, SMB92, SS95a, Sim92, gS96b, Tam92, WG93, Wat93b, Wei97b, Woo96, Wu92a, XX98, van97a, vvS99, Abd95, BT91a, BL91a, Bar98, BHL99, Blo99, Buf91, BT95b, CL98a, CWL99, Cap98b, CGP99, Che99, Eng93, ERR93, Fri90a, Fri98b, GP91b, HNN91, Har99b, HH91, HM90b, IEs98, LR90, LJ91, MV98a, Mal98, Mat91, MR99, Osa91, Osa93, Pat99a, Per91, Pvv98, Pie91, Pin98, Ric90a, Sau90, SY98, Sto99, WXZ99, El 98]. Positivity [GP92, GP94a, Gla98, SZ93a, BL94b, FJ99]. Possible [DJ95b, Sil96, Sil97d]. Possibly [WY96a, Har99b]. Potapov [FFK93a, FKK95]. Potapov-Normalized [FFK93a, FKK95]. potent [SEK99, Stu99]. Potential [JMOV97, KY93, Met97, Ans91, Miz91, dRT91]. potential-reduction [dRT91]. Potentially [EL99, GW95, SJ95]. pour [DC93, EK99, RSDT97]. Power [BAG96, BO93, CL96b, nC98, ERS95, GKP93, HTP96, WG93, BT95b, CH97, CL99, EvdD99, Fri98b, Oua96]. power-associative [Oua96]. Powers [AH94a, ABCP97, Cec96, Chk97, GKK96, Lev97, MP94a, MP94b, Vis98, Chi91, KT93, KS90, nM94, RR99a, ZT99]. Practical [Coo98]. Practice [Far96]. Prasad [Mit97]. Pre [PT94b]. Pre-test [PT94b]. Precise [Soh97]. Precision [HK93b, CP99a]. Preconditioned [Ser97, Cap98a, Liu91]. preconditioner [Aya99, O’L91]. Preconditioners [CN93, CW95a, CNW96, HY96, Ng97a, Tyr95, TYZ97b, Wai97, BK99, HN98, PS98b]. Preconditioning [BW97, BJW98, Di 98, Fre91b, Rei91, San96a, Sha98b, Tis91b, Tis91a]. preconditionings [KKY91]. Predecessor [BCT93]. Predicting [Pot93]. Prediction [GNP97, NP92a, tKWS99]. Preface [Ano92, BS92b, Bru99b, HPUV97, van95d]. Preordering [NO95]. preorderings [GHM99]. preorders [MLS91, MLS95]. Prescribed [BLJW95, Bot96, Bot97, Cab95, HHS97, Mic93, Min94, Nyl96, Sil93b, Sil94, Sil96, Sil97d, dS97, dSZ97, da 97, BR90, Chu99, DHS94a, FS99b, GKR90, Lin98b, Sil91b, VSB90, da 91]. presentation [Mit99]. presented [BS92a, HHR93]. Preserve [BS94a, BY95b, Duf94, FMP95, IOv98, BP90d, BLL94, Hoc94, Loe90]. Preserver [– DL94, LT91b, Mol99a]. Preservers [BL93a, CL98e, Coe96, CD97, Fun96a, Lim97, LP94, LP95b, SH97b, Wat97b, de 96b, Dia93, Gur94, Mol99b, Sha99b]. Preserving [BY95a, BKL97, CZ96, Die94, HHL92, KH92, LPS98, Li94a, Lim93b, Mrč95, NO95, OŠ98, PŠ98a, SH97a, ABGR94a, BT97a, BP91b, Chr99, EH98, GMŠ98, JSOvdD94, Liu97, MS99a, Oml90, OŠ91, OŠ93, Pet99, So99, Zal99, tKWS99]. Press [Bon94]. primal [LMS91]. primal-dual [LMS91]. primary [Rob91]. Prime [Cho93a, Laz95]. Primeness [Fuh94b]. Primes [Pvv98]. Primitive [BK97b, GKP95, HN93d, Kir97b, Neu96b, SB97, She95, She96b, SN98a, jWySmJzZ95, Yan92, Art91, Li90a, LMWZ90, Liu90, Liu99a, SH98]. Primitivity [She96a, TZ96]. Principal [BH95, Cha96a, HAPSG95, JB92, JMOV97, LS98, dSZ97, BK91b, Far99, Liu97, Lon99, Nar91, ÖA91, TK99]. principal-ideal [BK91b]. Principle [LZ97b, TW95, Xin93a, HASG90, KN91, PSW98, Rak99]. Principles [BY95c, Thi95, BC91a]. priori [Gał91]. Priority [KY93]. Probability [Ton94]. Problem [AB94a, AL95a, ABL98, BJT93, BF97, Bol97, Bor95b, BCU94, CF96, CS97, CG94a, nCZpZ96, nC98, Ded97, Ebi95, EH93, FH94b, Goe96, GKv98, GTÖ96, Guo96, GT97c, HL96b, HKL96, IKF^{+}93, IAMC96, Ion96, JN93, JS96, Kan93, Kno92, KS91, LV97, Lau97, Li97b, LT97b, Mah96, MP98d, Næv97, Nar95, Pan93, PD97, Qin96, Que94, Rad96, Rob95, She96b, SN98b, Spa95b, SY95, Wan95b, Wei90b, Wu92a, Zam94, AG98, AH95, AM98, ACWY99, ABKW90, BKZ98, BF98, BK91a, CLR99, CF99, nCjH98a, CH99a, CH99b, Die90, Fri91b, Gam97, Gro99a, HRW99, HSW99, IC94, JM12, Joh98, JS99b, KM91, Kur99, Li95b, LC91b, MP90, Mit90, Nyl99, Roe90, ST93, Str98, Ten90, Tre91, VB97a, War90, WE91, Wei98, WC96, Xu99, DCPS96]. problème [CF99]. Problems [AP96b, BK96a, BY95c, BM98b, BSK94, Coh96, CG97, Dah98, DS91, Dat94, De 93, DRT93, DH98, – DL94, Fay92, Fie97, FK93, FFK96, Gem97a, GLW95b, GLW95a, HRW92, Hei92, JT95, JS98b, KO97a, KY93, Kiw95a, Kiw97, Koš96, Kwa97, Lau98, LR93, Mal93, Man92, Meh96, Nie93a, NU97, OSiA96, Ott93, San96b, Shm94, Sil97c, SG95, TT95, Zie96, Ano98d, AFK90, BŠ99a, BTT91, Bro91a, Bro91b, Che98, nCjH98b, CL99, CGV98, Dai99, Dax91, Di 97, Din93a, FLL97, Fle90, FKK98, Gem98, Gol98, HN98, Han90a, Jam91b, Jia98b, LR90, LT91b, Li91b, Li97a, Li98d, LWX99, Mol99a, NS99a, Oga99, PYH91, PI91, Ruh94, Rum91, San91, SP99, Thǎ99, YP91, ZGL99, vV90]. Procedure [HA94, HS98, Mai91]. Proceedings [Ano94c, BHR91, LS95b, ORS94]. Process [Car97, Jia97b, Jia98a, KS96, ZZ96, DF98, Dax90c, Paa91]. Processes [PP94, Dah90, Dax90b, GRO98a, See99]. Processing [HM96, BGM98, Che98, CG98]. processors [ENV91]. Producing [Pat97b]. Product [Abt94, BR93, BLD93, BZ94, BR95, BvR^{+}97, Che92, De 92, EHJL97, FM95, FM96a, FS94, GES96, HS96a, HS97a, Hol97, KT95b, LMZ97, McD94, MBI96, Pay95, RZ92, Wan96b, Woo93, Cro98, GP91a, GS99a, Gro99d, HH91, JB90, KN98, Kol99, Kom90, LC91a, Lie90, Lin98b, MR99, Neu98, Wan97a, WXZ99, vdMRR99]. Products [And95b, And96b, AVH95, BP92a, BM96, Bot97, Bot98, Bot99, BKN97, CZ95, DRVP94, DR96, Dru94a, EM91, Ell94, EF97, GK97, HO91, HK97a, HM92b, Hor95, Knü96, KL98b, Lei92, Lev93, Lev94, Mat93b, MP98d, NDS99, Pat97a, QS95, Rho97, SH97a, Sun96b, Thi95, Tia94, TV92, VW95, Vis98, WW91, WG93, Wan95c, WX97, Wei97a, YC99, Art91, BT95b, DL92, DL01, DW98, Dru99a, HR98a, Im97, KNW91, Kra91, Li98d, LV91b, Liu99d, Liu99b, Liu99c, Mat95, Neu90, NS99b, Par98b, Ri99, Rum98a, TV90, Wei99]. Program [Li93b]. Programming [AR95, AB96a, Abe97, Bal95, CHX92, KEK97, Kan93, Lam93, Ans91, Bea98, Bro91a, DBRW91, Egg90, Fre91a, HRZ91, HR90, KN91, KM91, KPY91, LVBL98, LMS91, Mit90, Miz91, Pan99a, TZ91, Tod91, dRT91]. Programs [Hua96c, Kiw95b, Wol96]. Projected [EG97]. Projection [FB96, Jbi95, Kiw95a, LT97b, Mes94, Saa92, ŽWB93, BK99, DB99, Din93b, MV98a, PW90, Téc96, Wei90a]. Projection-Minimization [Jbi95]. projection-space [PW90]. Projection-Type [LT97b]. projectionally [ST90]. Projections [AS94, DR92, HN93c, HN94, Hou96, LH97, O’L90, RK96, Spi93, Spi94, Ste89, SC94, Tur91, Wil93a, AGK^{+}98, CM95, CW90, DA90, Spi95]. Projective [Fri97b, D’A98, Pan99a, SZ90, Tod91]. projective-scaling [Tod91]. Projector [WY96a]. Projectors [GT97a, GS98a, Gro99d, TY99a]. Prolate [Var93]. Prone [Wai97]. Proof [BC98b, Che95, Els95, GV96, Gre92, HR94, Ikr94, Olk97, Ral93, Sch96b, She95, Zen93, Ikr95, Ikr99, Kol99, Ost99]. Proofs [Sea94, Har99a, LRT99]. proper [CP99b]. Properties [BM93a, BH97, CYYL98, Chi92b, Eva95, FH94a, FM97a, Gee93a, Gee96, GL97, HJRS97, IEs98, Jen94, Kar94, KL96, KS94a, KP94, Lav97, MP97c, Men97, PL95, PŠ98a, PR94a, Sch96a, SG96, Smi92, SK97, SG95, Thi95, Tis92, Wu94, XL95, ZL98, Zie93b, BPS95, BM98a, ER90, FL91, FN98b, Her94, JS90a, JZ99, JN91, LS90a, LV91a, NO90, Pre98, Sid98, SGR99, Til98a, XB90]. Property [BCT93, CV97, CR94b, Had94, LPS98, Wei97b, BTT90, Daz97, Pin98, Chr99]. Proportional [Col96, Rat97, Bap94a]. Proportionality [Neu96a]. proportionnelles [Col96]. propos [MV95]. propriété [Daz97]. proving [KT91]. Provo [BHR91]. Proxies [BW96]. Proximity [Bal95]. Pseudo [O’L90, Ste89, FG99, FHP90, KST99, TR92, Sim96]. pseudo- [FHP90]. pseudo-distance-regularity [FG99]. pseudo-eigenvalues [TR92]. Pseudo-Inverses [O’L90, Ste89]. Pseudo-Ritz [Sim96]. pseudo-Toeplitz [KST99]. Pseudoinverses [HH94b]. Pseudoirreducible [MS91a]. pseudomonotone [Gow90]. Pseudoorthogonal [HWZ92]. pseudoprimitive [MS91a]. Pseudoskeleton [GTZ97]. Pseudospectra [LW98b, GS98a, TG97]. pseudospectrum [GdH99]. pseudounitary [dRWZ90]. PTA [LS98]. PTA- [LS98]. Pták [Vav95]. Publications [Ano98e, Ano98f]. puissances [Oua96]. Punctured [Gol97b]. Purely [WmL97]. Pyramid [Har93c]. Pythagorean [Sch92].

QL [Jia92]. QMR [FM97b]. Quadrangles [GL95, Pay95]. Quadratic [BRL98, BMTZ98, Bsh95, Car96, CMLS95, DER97, DdlP96, Fri91b, Gee94, Gee97, GPRL94, GLW95a, HRW92, HK97b, Kur97, MP97a, NM96, Pis97, ST96a, Swe93, TK95, Wan95a, Wan95c, Wan96a, Wan96b, Wat95a, XL95, ZD92, ACWY99, BC98a, El 98, FHP90, MVKR99, PYH91, Téc97, Wat90b]. Quadratically [Fan93, HRZ91, Paa91]. Quadratics [VL97]. quadratique [Téc97]. quadratiques [El 98]. Quadrative [AG93]. Quadrature [Put98, SV94]. Quadrics [Kes92, Tab93]. quadruple [Zab99]. Quadruples [FGP96, GKv98, GPM99b, Gro98b]. Qualitative [HL95, OTV93].

Quantification [LA94]. quantique [CM96]. Quantitative [Tho92]. Quantized [HKH97]. quantum [CM96]. Quasi [FH93, Xu91, SLK93, OY94]. Quasi-Hankel [SLK93]. quasi-Hessenberg [Xu91]. quasi-Newton [OY94]. Quasi-Spectral [FH93]. Quasi-Toeplitz [SLK93]. Quasicomplete [BKZ98]. Quasidirect [FM93b]. quasilikelihood [SRT90]. Quasilinear [Sad93]. Quasipolyhedral [MPdS97]. Quasitriangular [Gor92]. Quaternion [Hua96a, Wil93b, MS99b]. Quaternionic [Cuy95, JW99, Kan95, Tho97, Bak99]. Quaternions [Ree93, Zha97a, Kar99]. Quelques [Le 94]. Question [BF92a, DG95b, Dru92a]. Questions [Gen96b, DL99]. Queueing [BM98b]. qui [Oua91c]. quiver [Ser98a]. Quivers [Gus95b]. Quotient [Li92a, Li94c]. Quotients [CHR97b, MQ95].

R [Bha97a, Li98a, Loe92, Uhl92, YM94]. R. [DCPS96]. Radhakrishna [PS96b]. Radial [Sun94]. Radical [Huy96]. Radii [Chk97, EHS93, FN98a, HRR92, Kir95, EHP90, LT91a]. Radius [BN93, CV93, CHR97b, DMB95, Els95, FOV97, FK94, Gri96, HHNS96, Hor97, JLOvdD96, LW95, Li94a, Mae96, OR97, Pol92, SW95, SWP97, Szy92, WW96, Wat96c, Bog91, Chi90, Fri90a, JB90, Lie90, RR90b, Rum98a, hSyLtP98, Zal99]. Radjavi [DR91a]. Random [Chi96d, FR95a, GM95b, Juh96, KS94b, NL96, WL94, Ede91, Pal90]. Random-Walk [GM95b]. Randomized [Alt94]. Range [BŠS96, BDB92, BH92a, CT94, Chi96c, CYYL98, Fie95b, Fuj92, HJRS97, HR94, HHT93, KRS97, MP97b, Pie91, Tho97, Chi96b, CN99b, Daz94, Dru99a, JAG98, LS90b, Li91a, MP98a, NWZ98b, NT91b, Oml90]. Ranges [LT95, LSS98, BL91b, BR99, Mec98, Mir98, Roc98, TY99b]. Rank [BR93, BM90, BŠS96, BH96, BKS95, BBD^{+}97, BCT93, CD96, DHM95, DZ95, EL92, FRLM97, FFK93a, FKK95, GKP95, GVS96, Har93c, HOŠS96, Hig95b, Hil92, HS96a, HZ98, HLY92, JM97, LY93a, LMY94, LR94b, LB96, Nyl96, Pee95, Rob96, Scu93, Sem92, Wat97b, Wes96, Woe94, WC97, Adu99, BT91a, BL90, Chi99, Cos91, FD99, FL99, GRO98a, Gee99, GWS99, Gro99c, Gro99e, Gu99a, GMŠ98, HL91, HLP97, KS99, Lam90, Lim91b, Mic91, Mol99a, OŠ93, QOQO98, Scu91, So99, TK99, VW90, Wat90b]. rank- [BL90]. Rank-One [BH96, Scu93, Scu91]. rank-subtractivity [Gro99c]. Ranks [BM93b, CP96, Fie93, SH97b, ZZ96, CR93, PT99]. Rao [Liu99d, PS96b]. Rapid [Jan97a]. Rate [BF92b, KS91, Li95c, RR94, Sou91, Str91]. Rates [Ach93, EN93, MR94, CMH90]. Ratio [BK94b, Tuc96, McL90]. Rational [AA90, AB96b, BK94a, BK96a, BSK94, nCZpZ96, nC98, Fau92, Fay94, Fri93, HF98, KO97a, LR95, LR93, Now96, RRR93, Rav97, Rea97, Ruh94, Van90, Var98, Adu99, AD90b, ABGR94a, ABKW90, BGR90, BR90, BR91, BKGK94, CH99a, Di 97, FN98b, GKR90, Gu99b, HSW99, HL99, JS90b, LR99, Lom99, RR90a, WSCP91]. Ratios [Car92]. Ray [LM97, MOTV97]. Rayleigh [CHR97b, LY93b, Li92a, Li94c, MQ95]. rays [Pat99a]. Re [Wu92a, WC96]. Re-nonnegative [WC96]. Re-Positive [Wu92a]. Reachability [BCHS97, GRO98a]. reachability/controllability [GRO98a]. Real [BD93, BJT93, BL93a, BRL98, Die96, Die98, ERS95, FN97, HW92, HF96, HS97a, HM93, IKF^{+}93, JT95, LY93a, LR95, Lau97, MS93a, MVKR99, RR94, Sch93, Sla98, Soh97, Tre93, Tur92, BM98c, CL90, Chu91, EHL97, Fie90b, LS99, Rum98a, Soh91, Str91, Tho91, Tre99a, Zal99]. Real-Coninvolutory [HM93]. Realizability [BCHS97]. Realization [AD96a, CC92b, CBSH94, HJ92, KL98b, KS92, MM93b, van97a, AR99, AP99b, CH99a, ETS90, Gał91, VB97a]. Realizations [Aze95, FH94a, HF94, HF96, LH97, PP94, AA90, App99, FN98b]. Reanalysis [Jan97a]. Rearrangement [Kom90]. Reciprocal [Len94, Swe93]. recognition [KvL93, NS99a]. Recognizing [HHR93]. recombination [CMH90]. Reconstructible [KO97a]. Reconstruction [SC94, DA90, EKN90, Fle90, PW90, SS90]. Rectangular [BCS94, Chi96d, Chi97, HM95a, Ilt97, Kni95, Ste94, WZCZ93, Gu99b, Mar99a]. Recurrence [DV95, YC92, ZW98]. Recurrences [Gut93]. recurring [SAZ91]. Recursions [RR92a, ZZ96, Sha98b]. Recursive [BM98a, GO93, HR95, Mes94, Ng97b, RV95, RV96, Sch96a, Bro90, JM99, Pan99b, Van90]. Redheffer [Hum97, Jar90]. Reduced [BS97, MV98c, Sch96a, BKGK94]. Reducible [Sha93, Stu99, Väl93, mJS91, Jia96, Li90a, Lon99]. Reducing [TT95, CF99, JS90a, Jam91b]. Reduction [Fay92, FP96, Fuh94a, KY93, NP93, Thi95, Wim93, YP96, BW91, BZ98, Chi99, Chu91, Miz91, ST91, dRT91]. reductions [Lu98]. réduisants [CF99]. Referees [Ano98g, Ano95c]. Refined [Bal95, Jia97b, Jia98a, Jia99]. Refinement [Jin98, JZ99]. Refinements [Cha96a]. Reflections [BNP93, RK96, Dru99b, WZ99]. Reflectors [SH93]. Reflexive [BPY92, DG95b, FNW96, DW98, FNW91, Lam90]. réflexivité [Del99]. Reflexivity [Din92]. Reformulation [MR98c]. Region [CC94, Ito97, ACWY99]. Region-dependent [CC94]. Regions [RSR94, PPG91]. Register [Lan96]. Regression [BW96, FC97, LS92b, Lis92, PT94b, WY96b, Chi99, HB90, Wie90]. Regressor [WAS97]. Regular [AGMP97, Bap98b, CLW90, Chi95, Cho93b, Cuy95, – Dok95, Eve96, GL95, GK95, GKR90, GKv93, GGM96a, Gou97, GV95, HT93, HWH95, Kir97a, KL98b, Li92b, LZ96, Pee95, PR96, RR97, Spe95, Woź97, Xin93b, de 96a, van95a, van95c, BHM98, Del90, FJK98, FS99a, HS90a, Li91c]. Regularity [CK97, GP94b, Hae96, Row95, BCG91, Cec90, FG99]. Regularization [San96b, CMN99, Fle90, Han90a, KM91]. regularized [SS90]. regularizing [BGBMN99]. Regulation [CS97]. Reich [MG94]. Related [AMPV97, BS92a, BS98, BT97b, BG94, CH92, CLLT98, FBO97, Fur96, Gen96b, Gon96, Gro97a, KO97b, Kir95, MW95, OSiA96, AK98, CK91, DL99, Far99, FM93c, GZ98a, WM90]. Relating [LT95]. Relation [Her93, Yip96, AH90, CHASG98, Hig94, KM91]. Relations [And95b, DV95, Han90a, MOPvdD99, Thi95, YZ95, Yu95, Zho96, Cro98, LN91, VW98]. Relationship [Kal96, Lis92, Hou99, Li91c]. Relationships [ER90, HC98b, YA94, TZ98]. Relative [CIR^{+}93, Drm96, Li97c, MV98b, Raz92, Sun90b, TS99, Wat92a, Zha93, D^{+}99]. Relatives [HKL96]. Relaxation [Kiw97, LP95a, Str97b, ACWY99, Bai98, Dax91]. Relaxed [MMPS96]. Reliable [O’N93]. remain [Cai99b]. Remark [Tyr97, Wer96, San96a, Thǎ99]. Remarks [AD97, GT97b, HF94, PR94b, PS96a, Szu95, KSV91]. Remarque [Che99]. reminiscences [Sch98b]. Renegar [dRT91]. Repeated [EJ93]. repetitive [GRO98a]. Report [BGH92a, BdV92, Da 94, HPUV97, Meh98, OPSS98]. Representability [BD91]. Representation [GK97, HPW94, KP94, Kit93, Ter98, Chi91, LS93b, MM98, WMP98]. Representations [AS97, Chi95, Fei94b, FH96a, HKH97, Hei95b, HR98c, HOR98, MM96a, MM96b, Ros93a, Wei94, Bou98, ETS90, HR98b, Ser98a, Vin90]. Representing [Spa95b]. Reproducing [AD93]. Reproduction [Car97]. Require [EJ93, LPS98]. Research [Dat94]. Residual [Drm96, Xie97a]. Residue [ABGR94b]. resistant [ZX99]. Resistive [SSS97]. resistor [CIM98]. Resolution [Mos96, Mos96]. Resolution- [Mos96]. Resolutions [ATZ95, HAP93]. Resolvent [FFK96, Phi91, Del90, FKK98]. Resolvents [MNS93]. Resolving [CD99]. resonance [MM99]. Respect [DeA96, Rao96, Cha96b, CP99b, PSW98, Rao91]. Response [FWWJL92]. Responses [YA94]. Restarted [BF97, Jia99]. restoration [KN98]. Restricted [DS94a, FH93, Hon92, Nie93b, RZ98, Van97b, BP90a]. Restrictions [BC92, BZ95, DB99, PT94b, Rum97, Chi99]. Result [Bd97, Gra96, HK97b, KH95b, Tam97, Tam98, Ikr99, Mar98, Ost99]. Resultant [KS96, BG98a, Wim90]. Results [BM96, Bal95, CLLT98, CF97b, DH98, FM96b, Har93b, HR93b, Sea94, SW95, SV96, Smi95, SW91a, Xu97a, CGV98, DR99, FI99, Guo91, HY91, Li90a, Li91a, Liu99d, McL90, MOPvdD99, SBKK98, Wer91, tKWS99]. Revealing [HLY92, HLP97]. Reversal [Kir95]. Reverse [AD96b, DW98, Ion96, MP95, Oar96, Tia94, Wei99]. Review [BB92, DCPS96, Gro95b, Kir98, Loe92, Nas92, Saa92, Sha92, Str96, Uhl92, Li98a]. Reviews [Ano94b]. Revisited [BNM92, CH96a, CC92b, GKT95, KSW99]. Revisiting [Fri91c]. Rhomboidal [Fie96]. Riccati [BA– DR91, CG92a, D’A97, D’A98, FP99, Fd92, FJAK96, FI99, HC98b, IW92, IW93, IO96, Ion96, Ion98, Jód91, Jon95b, JL94, KTK99, KSP98, KT96b, KM90, LRT97, Lew99, LY93c, LL93b, Oar96, RW94, Sch91, Sch95, Sto96, SS98, gS97a, gS98b, Wim91, Wim94b, XL95]. Rich [BBLP95]. Richard [Sha92]. Richardson [App99, Aze95, Rei91, SS91]. Riemann [Car92, Roe90]. Right [Bak99, FM97b, HH92, BM91, GW99]. Right-Hand [FM97b, HH92, GW99]. right-star [BM91]. Rigid [VW98]. rigidity [NS99a]. Ring [BR92, HL97, Jel92, MM96a, Rei93, BHH90, BK91b, ZJ91]. Rings [ATST95, AGMP97, AVH95, CM93, GRS92, Hau95, HAPSG96, IW96, Jøn95a, Kam94, LM92, Lan92, van96, Adk91, AP99a, BK91a, CX98a, Gur91, HASG90, NT91a, Ote90, Pra94, SG99, VW90]. Ritz [LY93b, Sim96, SV96]. RLS [BNP93]. Roads [Tho92]. Robinson [Ral93]. Robust [Fuh94a, Gom95, LWYS99, Wie90, GdOH98, LN98, dOGH99]. Robustness [Fei94a, NF95]. Rodman [Har90]. Roger [Uhl92]. Role [Ton97, Dor98, KT91]. Root [Kol93, Liu96, Shm94, Son92, TK98, Tho95, AK98, PPG91, Ple98, Sau90]. Roots [BGM97, BT95a, Far95a, Hav93, KU93, Kra94, Mar95a, Rea97, Sch92, BEK90,

– Dok90, FMMX99, Ote90, Pie91, Ri99, Sau90, Soh91, vdMRR99]. Rosenberg [Xin93b]. Rosenfeld [Kur99]. Rota [FN95b]. Rotation [Nov97]. Rotations [AP96b]. Roth [Wim94a]. Roth’s [HL97]. Rounding [Nar95, Sun92a, Roh90]. Rounding-Error [Sun92a]. Routh [Gen96a]. Row [BNP93, Fie96, FM97a, GKP92, Gut93, SM93, TV92, WW96, BH94a, Dax91, LMZZ98, TV90, WZ98, Zho99]. Row-Rhomboidal [Fie96]. Row-Stochastic [GKP92]. rows [FS99b]. Rule [Abt94, Zie93a]. Rules [Woo93, BS99b, Zie98]. Run [Chi96a]. Runge [CC94]. Ryser [CCHV97, Mic91, Sha92].

S [Ano95a, Nas92, Lav99b]. S-subinvariant [Lav99b]. S. [DCPS96, Mit97]. Saddle [Chi96d]. Saddle-Point [Chi96d]. Saddlepoints [Men97]. Sakamoto [Olk97]. salesman [War90]. Sample [Neu96a, Sch96a, Jin90, NW90]. Sampled [HFK94, YA94]. Sampled-Data [HFK94, YA94]. sampler [Har99b]. Samples [RAB93]. Sampling [DSW96, Grö93, Raz97, Ziz98]. Samuelson [Cam93, VP92]. Sandwich [Hua93, Hua96b, Gol98]. Satisfies [HAP95]. Satisfying [DdO94, RZ92, van96, MV95, PSW98]. Saturated [KSH94, MCK95]. Scalar [BR95, BvR^{+}97, BM92, nCZpZ96, HS96a, HS97a, Hol97, KL98b, Gay91, vdMRR99]. scale [BP98]. Scaled [Chu95, HN93c, O’L90, Ste89, Ans91, CH97, SW94]. Scaling [AP96b, BC98b, BCP93, KK96, Kal96, KEK97, Kha96, Neu97b, PN92, NR99, Tod91]. Scalings [RZ92]. Scattered [Sun94]. Scheduling [BK96c]. Scheme [PV95, San96a]. Schemes [GL95, Hig95b, Kiw95b, CC94, PP99, Vos99]. Schmidt [Bjö94, Gom95, Nie97b]. Schmidt-Mirsky [Nie97b]. Schneider [Bru99c]. Schoenberg [CF97b]. Schur [Ikr95, AMM99, AO91, AFK90, BD93, BT94, BGE91, Cha96a, CG92a, FGL^{+}98, FK93, FFK93a, FFK96, Gem97b, Gon97, Hla99, HM92b, Ikr94, KS94c, Lac95, Lei97, LRT99, LZ97b, Liu99b, LW99, Liu99c, Liv95, MS98, Nie98c, OS94, Par98a, Peñ97, Shi99a, Smi92, Soh90, WZZ99, WM94, de 96b]. Schur-Algorithm [CG92a]. Schur-like [AMM99]. Schur-Sequence [FFK93a]. Schwarz [BD95, HM90b, LN97, PPS96]. scores [tKWS99]. Search [GN93, Dax90a, PP99]. Seasons [Gol97a]. Second [Bha94a, DR93, DLdO94, FN97, GGM96b, Jod92, An98, Col96, LVBL98, Lub98, Oua91b, Da 94]. Second-Order [DR93, Jod92, Col96, LVBL98, Lub98, Oua91b]. Secular [And96a]. Segment [Li95c]. Segre [CU99]. Seidel [KKNU97]. selection [CLR99, WLN99]. Selections [Prz95, WY96a, LN91]. Self [AP96b, ABBN94, HM92a, Kar99, Pie96, RZ97, AIY90, vdMRR99]. Self-Adjoint [ABBN94, RZ97, Kar99, vdMRR99]. Self-Correcting [Pie96]. self-dual [AIY90]. Self-Equivalent [HM92a]. Self-Scaling [AP96b]. selfadjoint [MR98a]. Semencul [Huc98]. Semencul-Type [Huc98]. Semi [Gro97a, BG98b, BK91b, GTT99, Lei99b, MPdS97, Pat99a, Téc96, PV94]. semi-definite [BG98b, Pat99a]. semi-definiteness [Lei99b]. semi-Euclidean [Téc96]. semi-euclidien [Téc96]. semi-infinite [MPdS97]. semi-local [BK91b]. Semi-Markov [PV94]. semi-orthogonality [GTT99]. Semi-Star-Related [Gro97a]. Semidefinite [AH94a, Dan92, FMP95, JT95, LP95a, Lau97, Lau98, LN97, LZ97b, LB96, SMB92, Ste97c, WG93, Wim94b, Wol96, Woo96, BHL99, CWL99, Dru99c, GP91b, HH91, HM90b, Mat91, MR99, Pie91, RT94, Ri99, WXZ99]. Semigroup [Cho93b, Hua96b, VT99, YZ95, CX98a, GMŠ98, KB99, Zha91]. Semigroups [BC93, Gon97, Hua93, Okn97, ORR94, OR97, HS90a, RGS99]. Seminorms [AG93, GP95, HZ99]. Semipositivity [JMS95]. Semirings [BP92b, Scu93, BT97a, GB99, Scu91]. Semisimple [Chi95, FJK98]. Semistability [CHS92]. Sensitivity [Bal95, CP98a, OST93, gS98b, Cha98, Par98b, Rum91, Sun90a]. Sensor [HM96]. Separable [Bal95, Chi95, MM93b]. Separable-Denominator [MM93b]. separate [McL90]. Separating [Din92, Sti91]. Separation [Duc95, MP93, Mit94, US92, Xu97b]. Separators [GNP97, McC93]. Sequence [Cam95, Car92, FKK95, HP96, Kir96a, Mic93, RSR94, Szy92, TK98, AP99b, Le 94, Shi99b, FFK93a]. Sequences [AP94b, AI96, Aze95, BS95a, EN93, EI97, FK93, GO93, HHS97, JMOV97, Lan96, Rud98, App99, DS98, Die90, Lav98b, SAZ91, Til98a, ZW98]. Series [ERS95, Lau97, AG98, Osa93, PI90, Ric90a]. Series-Parallel [Lau97]. Sesquilinear [GM95a, AD90a]. Set [HN93c, Har93b, HM95a, ID92, Jan97b, Kan93, LW95, Mae96, SB97, Sil93a, Uhr92, WR94, Yan92, BM98c, Che98, Din93b, FPP98, FG99, Gel99, mJS91, Kni99, Kri99, Li90a, LMWZ90, MM99, Mub99, Rob91, Sch91]. Sets [DL92, EU93, Har93b, Har97, IS95, Jel92, Jeu95, JS98b, Kan95, MT94, Prz95, Uhr96, Wim94b, Yiu93, Hur90, MPdS97, SGR99, DL01]. Several [Dan90, Dru90, GV96, WmL97, CMH90, FKK98, Gel99, Pvv98, Zha93]. Sex [CMH90]. shadows [KV90]. Shanks [BNM92]. Shape [Cos95, HOR98]. Shaped [Uhr92, de 91]. Sharp [hCjXL96, HS99, Li93b]. Sherman [LV97]. Shift [Jia92, Lan96, Ptá96, Raz92, Wat96b, Zam94, KR91]. Shift-Invariant [Zam94]. Shifted [Bat94, Fre91a]. shifted-barrier-function [Fre91a]. shifting [BK91a, SW90]. Shifts [Wat96b]. shop [BKZ98]. Short [AK97, ZZ96, Ost99, dRT91]. short-step [dRT91]. Shortcut [Bom96]. Shorted [Met97, MP96a, Mit97]. Shortened [Gol97b]. Shortest [jWySmJzZ95]. shuffle [DPPR90]. SI [San91]. Sided [ABGR94b, AB95, CL96a, Kol95, Nab99]. Sides [FM97b, HH92, GW99]. Sign [AB94b, BY95b, DR94, Esc93, EJ93, EHL94, EHJL97, Eve96, GP94b, GB99, Gib96a, Gib96b, GOV96, JOv95, JMS95, LLM95, LPS98, LHZ97, Lim93a, LMV96, LMOV96, Rum97, SL96, SJ98, gS97b, Tho96, Wat96a, EHL97, EL99, Hig94, LHE94, Rum98a, Sha98a, Sha99a, Sti91, SEK99, Stu99, SJ95, Zal99]. Sign-Central [AB94b]. Sign-Invertibility [SL96]. Sign-non-singular [GB99]. Sign-Nonsingular [BY95b, Gib96a, Gib96b, LLM95, LMV96, LMOV96, Tho96]. Sign-Pattern [Esc93]. sign-real [Rum98a, Zal99]. Sign-Regular [Eve96]. Sign-Regularity [GP94b]. Signal [HM96, BGM98, Che98]. Signature [CG93a, – Dok93, FH93, Ghe92, RZ97, Her91]. Signed [GMJ95, GMJ90]. sim [Li95c]. Similar [BM97, Knü96, Lon95, Lon96, Roe97, CT91, MMP99]. Similaritäten [Gab97]. Similarities [EI97, EG90, Gab97]. Similarity [App98, BZ95, BT91b, BG94, CS92, FP92, FD99, Fri93, Gur91, HS96b, LR94a, LM94b, LT95, Lim93b, Sil96, Sil97d, Sol99, TY97, Thi97, WW94, Wat95a, Zab94, Cru91, DL99, GV98, Hen97, Lev94, Lin98b, Sha91b]. Similitudes [Har96]. Simple [BJT93, Chk97, GM97b, Gra96, Ikr94, Le 98, SB97, Sun96b, AH90, App99, Ikr95, Kol99, Nie90, TZ98, Yon99]. simples [EK99]. Simplest [GPS99]. simplex [Pan99a]. Simplices [ACE95, HKL96, Vel95, LZ98a]. Simplicity [TK99]. Simpson [Had97]. Simultaneous [BW91, BZ98, CD92, Kuc91, Mar99a, ST91, Shi99a, Wim93, BL91b, Chr99, Chu91, DL99, Mit90]. simultaneously [HO91]. Sine [CNW96, Vel95, Di 98, LZ98a]. Singh [Liu99d]. single [CLR99, Neu90, OTV93]. single-input [CLR99]. Singular [BP92a, BAG99, BOS93, BR95, CC96, Che92, DS94b, De 92, DD97b, Dei91, DG93, EF95, Gee93a, Gee93b, GKK96, HM92a, HK93b, HP92, Hon92, Jia94, JS98a, LS98, LZ96, Lin98c, LC92, LN97, LP94, MS93c, MST94, MT94, MT96, MZ91, Nie97b, PV95, Peñ97, PS97, PS96a, Que94, QS95, Ren96, SS93, Sea94, Sid92, Sil93a, Tam98, Til98b, TS97, Tyr94, WX97, WY96a, Yan93, Zha96, AH90, Bap91b, Chu99, CH99c, CL91, D^{+}99, GB99, HKS96a, Hua91, JN90, JL98b, Lev98, LT91b, Li98c, Liu99b, MS91b, RR98, RS98, SRT90, Sid99, ST93, Tar91, Van91b, YhG97]. Singular-Value [BR95, DD97b, HM92a, PV95]. Singularly [AI96, LS97]. Sinkhorn [Ach93, BC98b, Sou91]. SIRT [vV90]. SIRT- [vV90]. Situations [Neu96a]. Six [WAS97]. Sixth [BGH92b]. Size [RR97, RR98]. Skew [CP96, Die96, FK94, Fuh94b, Gib96a, Gib96b, GVS96, HR93a, Iwa98, Moo95, Mur95b, RRR93, HM99, LT91b, ST91, Tho91, FMMX99]. Skew-Centrosymmetric [CP96]. skew-Hamiltonian [FMMX99]. Skew-Symmetric [Die96, FK94, Gib96a, Gib96b, HR93a, Iwa98, Mur95b, RRR93, HM99]. slowly [Art91]. Smale [KM91]. Small [BC91c, Cla92b, Gol97c, Log98, Sch96a]. Smallest [HP92, Jia94, JS98a, WN95, Ede91, Van91b, YhG97]. Smith [GKS95, KMS96, New97, Rus95, SL97, Sto98]. Smith’s [HWS97]. Smoothed [ID92]. Smoothest [TK95]. Smoothing [GW96, HS98, LV97, PI90]. Society [BHR91, Da 94]. Softly [Sch99a]. Solution [Abe97, AP96a, AB96b, CWX96, Di 98, Fd92, Fra95, GK93, Gro99e, GLW95b, HR95, HTP96, HM96, Jan97b, JS98b, Kan93, Kit93, Kiw95b, Kri95, LY93c, Lin92, LP97, Oar96, PD97, PPY93, RV95, RV96, RS97, Sid92, SS95b, Str96, Woo96, XL95, Zam94, AM98, ABKW90, BTT90, BGE91, CW93, Eng93, ERR93, Fle90, Guo99, HKS96a, KSP98, KM90, LL93b, Meh91, ÖA91, PYH91, Sch91, Sid99, Tre91, VB97b, vV90]. Solutions [CN92, EU93, Eva95, FL96, GLW95a, Jod92, KT96a, KT96b, LRT97, Li93b, Lop94, MBI95, Pat97c, RR96a, Roh95, Roh97, Sak93, SZ93a, Sto96, Tis92, US92, Wim94b, Wol96, Wu92a, Ano98d, BK90, Bal99, BP98, D’A98, Dai90, DH94, Dob99, FP99, GM98a, Gro99a, IEs98, Lub98, MBI94, WC96, XWZ98, Xu99]. Solvability [Gee93b, JV92, Li91b, Li95b, Li97a]. Solvable [PZ90b, PZ90a, YP91]. solve [Bro91a, RG99]. Solver [CR92, KV98, VK98]. Solvers [AB92, Gut93, Sha98b, SP99]. Solving [Bea98, Cop93, FWWJL92, HP96, HK93b, HO93, Jód91, Mal93, Ple95, RW94, Shi95, Aya99, BKO99, BGS99, CC94, JZ91, Mai91, Tis91a]. Some [AD97, AB90, AB91, AV96, Auj97, BVB93, BPS95, BGM98, BM94c, BS95a, BS97, Bou99, BH95, CP98b, CLLT98, CG93a, Cox95, CRS99, Dat94, DO94, DeA96, – DL94, DA90, EHL97, Eva95, Far96, FM96a, FM96b, Fie97, Fie98, Ger98, GT97b, Guo91, HY91, HY96, HOŠS96, HR93b, HM95b, HKL96, Hwa90b, Hwa95, ID92, KY93, Kit94, KSV91, Laf98b, LS90a, Le 94, LR93, LR90, Li91a, Li91b, Li94b, Li96, Li98d, Lin98c, Liu99b, LW99, Liu99c, Mah92, Mal93, MP97c, MV93, Mol99a, Mol99b, MPŠV97, Nie93a, Pat97c, PPS96, PR94a, PS96a, Rez95, Sch98b, Sea94, She96c, Shm94, Smi92, Smi95, Szu95, Tis92, Ton94, TV92, Tyr96, WG93, WX97, WZZ99, WXZ99, Wat92b, WAS97, Wer91, Wu94, Xin93a, Xin93b, tKWS99, Ala99, BŠ99a, – DL95]. some [FI99, GM98a, Gay91, Gem98, KT91, KLS90, KPY91, LRT99, LS93b, LT91b, PT94a, PI90, PP99, RG99, SGR99, Ten90, TV90, Tyr91, VV92, Wan97a]. sont [Oua91c]. SOR [Dan91b, EV93, GHN97, GHN99, HY91, HN93a, HNN91, KPS91, Nie99a, PI91, SN91, de 91]. SOR- [PI91]. SOR-like [de 91]. sort [Bro91a]. Sorted [ZZ98]. Sound [Sei96]. Source [Mir92]. Sourour [EG99]. Sous [BC91b, CF99]. sous-espace [CF99]. Sous-espaces [BC91b, CF99]. Space [AB95, AS94, Beh94, BRL98, BN93, Cla92a, CPR93, HFK94, HN94, Lin98a, Ree93, RK96, Ren96, Vro93, Wan93b, Wan93a, van95c, AR99, ABGR94a, BH94a, BC91b, BDV98, CG99, DF98, Dru90, Gał91, GH91, GH95a, Loe97, MG99, Nyl98, Pap90, PW90, Téc96, Zho99]. Space-Time [Vro93]. Spaces [AD93, ABBN94, BRW94, BL93a, BvR^{+}97, CWX96, Chi97, DHM95, DH97a, EGLL95, FL99, HF98, HLMT98, HS96a, HS97a, Hol97, LY93a, LMZ97, LR94b, Lyu95, Mar92, Mes95b, OŠ96a, Pet96, Scu93, SH93, Wer96, Wes96, AD90b, AD90a, BAG99, Bog91, CX98b, CL90, Cow90, DH94, Dor98, El 98, Idr99, Lim91a, MOR91a, MR98a, SZ90, Scu91, Wim99, Win99, vdMRR99]. Spain [BdV92]. Spanning [BC92, SH97b]. Sparse [Alt94, GNP97, GLW95a, HL96a, Mas94, PPY93, Pot93, Bai98, BP99, CP98a, Gay91, Guo91, Har99b, JL98a, Mai91, Tis91a, Van91c, vV90]. Sparsity [IOv98]. Special [Ano94d, BMH94, BH96, BS94b, CD92, CL96a, Gro97b, GT97c, Jeu95, LW97, ORS94, BS92a, FJK98, GTT99, Nyl99, Sid98, Ten90]. Specific [BGM97, McE94]. Specified [CD96, GOV96, Kri94]. Spectra [ATST95, AH97, Boa97, Far92, GHN97, JOTV93, cL93, LW98b, LMOV96, Sim92, Tre93, TZ98, WLL99, BM98c, GHN99, KSV91, Li90b, So94b, Tre99a, Tre99b, de 91]. Spectral [BL93b, CV93, CP99a, CG92a, CALM97, CHR97b, CR94b, DR92, EHS93, Els95, FOV97, FL91, Far92, Fay92, FV96, FN98a, FN98b, FH93, FK94, Gri96, GPRL94, HRR92, HHNS96, Hol92, HOŠ96, Hor97, JZ99, JL97, JLOvdD96, Kat94, KP92a, KS94a, Kir95, LW95, Li98d, Mae96, Mal93, Mar96, MQ95, OR97, Ota97, Par98b, Pol92, RS97, SW95, SWP97, Shm94, Szy92, Til98a, Zie93b, Bog91, Chi90, CGV98, Del90, Dru98, EHP90, Fri90a, GS98a, HL91, Her94, JB90, Koš94, LL91b, LT91b, Lie90, Lim91b, Mat90, Phi91, RR90b, Rum98a, hSyLtP98, Zal99]. Spectral-Radius [Els95, Bog91]. spectroscopic [SS90]. Spectrum [BP90c, BH94b, Bro95, DS91, Els94, FF96, FK94, GS96a, Hae96, HN93d, KPRV97, Kir93, Kru92, Li95c, Mrč95, OŠ91, Pry92, Sun96a, Cig99, DST98, HNN91, Hoc94, Mal98, Pet99, Smi90]. Spectrum-preserving [OŠ91]. Spedicato [Saa92]. speed [Gay91]. Sphere [ST96a]. spheres [Pap90]. Spherical [Boy95, CF97b, Dav99a, Dav99b, Nak97, Rez95, Hab94, PST98]. Spinors [RK96]. Spline [TK95]. Splines [Jia97a, JS90b]. Splitting [Jin98, Van91c, Li91c]. Splittings [Li92b, LZ96, LP97, MNS93, McD94, Nab96, Woź97, Xin93b, CP98b, CP99b, Son91b].

Sporadic [GM97b]. Spread [JZ97].

Spreads [Tho95]. Spring [NU97, Nyl99]. Spring-Mass [NU97]. Springer [Li98a]. Sprinkling [JN93]. Square [CHR97a, CLLT98, Mar95a, Mit96, Sch92, SK97, FMMX99, Pie91, Ri99, vdMRR99]. Square-Zero [CHR97a]. Squares [AP96b, CWX96, De 93, DH98, FB96, Fra95, GT96b, Gro97b, HS95, Hu95, Jan97a, Jia97a, Ng97b, Pit97, SZ93a, Ton97, WAS97, Wei90b, WY96a, Woo96, BP90a, BM90, BP91a, BTT90, BTT91, Che98, DH94, Jam91b, Oga99, SRT90, WZ99, Wei98, Wer90b, vV90]. squaring [Fri91c]. SSOR [KP95, Buo90a, HN93a, HNT96, KKY91, KP92b, MTS96]. Ssytems [Rei94]. Stability [BK94a, BK93, BN93, Chu93, Dat99, FGP96, Gee94, GV95, Gur95, GH95b, Her92b, Hig99, JMOV97, LH97, Lin92, LS97, Mar97, MV98c, NS97, SWP97, SS96, WF92, WM94, YP96, vdMRR99, DS98, GPM97, Gem98, GdOH98, Her98, Kra91, LWYS99, Nie90, Shi99a, Soh90, dOGH99]. Stabilizability [Che91]. Stabilization [DR93, Fuh94a, MHR93]. stabilized [VK98]. Stabilizing [IW92, Oar96, Sto96]. Stable [AB96a, BOC95, BO92, CAHJ98, GW95, Gut93, KH95a, RR90a, Sun92b, Vel99, Buo90b, FGL^{+}98, Hol99, IHM99, PST98, SJ95, VW90]. stage [Cao98b, CWL99, CC94]. staircase [For90a]. Standard [LS92a, LS94, BH97, Han90a, RS90]. Star [Gro97a, HOŠS96, Row95, Uhr92, BP91a, BM91, Bel99, CRS99, JR99, Wer90a]. Star-Shaped [Uhr92]. starlike [CLW90]. stars [CA98, Lak99]. State [BN93, Cla92a, FV97, FH94b, Gee94, Gee97, HA94, HFK94, MM96a, MM96b, Nie97a, Rat97, AR99, ABGR94a, AA90, DF98, Gał91, RZ98]. state-feedback [RZ98]. State-Space [BN93, HFK94, Gał91]. state-variable [AA90]. Statement [Ito97]. static [BJW98, LWYS99]. Statics [HL95].

Stationary [Cao95, HK93b, PP94, Dah90, Dax90c, Dax90b]. statistic [McL90]. Statistical [BMH94, CV97, Dah98, Far96, HK97b, PPS96, Šem96, BP90c, Far99, HP90]. Statistics [BK94b, DRVP94, Rol96, Ton94, BS92a, WI99]. Steepest [KS91]. Stein [DJ95b, DeA96, Xin93b]. Step [Ghe92, HY96, LMMR95, Meh96, dRT91]. Stephen [Gro95b]. Stieltjes [ABGR94b]. Stiff [AP96b]. stiffness [San96a]. Stochastic [Ata92, BNP96, BGS92, BS94b, CF93, CHX92, CIR^{+}93, EF95, Fie95a, Gib92, GKP92, HM98, Hwa95, HS97b, Ito97, LS93a, Lei92, Min94, Per96, PS94, Rho97, Sen93, SMB92, SHJK97, WC97, – Dok90, GH91, GH95a, GP91b, HS90a, KS90, nM94, Nie90, NT91b, Sem91, Sin90, Son91a]. Stochastically [Lon95, Lon96]. Stokes [HX92]. Stone [DCPS96]. Straffin [Fis95]. Strategies [Alt94, Bom96, FR95a]. Strategy [ON96]. Stratification [GPM99a]. Strict [GP94b]. Strictly [Eve96, HLMT98, Hua91, Pin98]. Strong [ACWY99, Cec90, CK97, Fei94b, FP97c, Kra95, NS97, Yan92, KR91, Sha98a]. Stronger [WY98]. Strongly [HP90, LMY94, LPS98, Pee95, Sun92b, vH96, BP90d, HP99, Sch99b, WM90]. Structural [Coe93, FGP96, Juh96, Neu97b, RW97, van96, CP99a, GPM99b, GdOH98, HS90b, RSDT97, dOGH99]. Structure [Bap98b, BC93, BK97c, BZ96, BM93b, De 92, DIO97, Die94, Fie93, FV96, GNP97, Gut98, HM98, Hav93, HLWT96, HJ92, HH94a, HH94b, Her93, HZ98, KO97b, Kar94, Koš93, Koš97, Len94, LZW97, LV91b, Mas94, MO95, Mic93, Pot93, RR97, Ros93b, Sob96, Tam92, Tur92, WR94, BR90, Bol98, CK91, FPP98, GH91, GH95a, GS99a, Hei98b, Her99, Jin90, KLS90, Mar99b, Mic91, Nyl98, RR98, Tam99, TZ99, van91a]. Structure-Ranks [BM93b]. Structured [AD90b, AD90a, BSK94, De 93, GO94, Jen96, LS94, Rum98b, SDT97, gS99b, van94, CP98a, HQ98, Tre91, van91a]. structurelle [RSDT97]. Structures [AG93, ADH97, DIO97, GV92, Jan97a, Kru97, SK95, Sch94, Vro93, WZCZ93, CG99, MM98, SB90]. Student [Har99a]. Students [Ano98h]. Study [CRS93, ST90]. Sub [FJ99, BG98a]. Sub-direct [FJ99]. sub-resultant [BG98a]. Subalgebras [HWZ92, HWZ90, LL91a, TW99, dRWZ90]. Subclass [FV92, GC95, Kir99]. subcodes [VSB90]. Subdifferential [Wat92b]. Subdifferentials [Zie93c]. subdivision [Pra91]. Subdominant [HM98]. Subfield [ORR94]. Subgraphs [PPY93, PPY95]. Subgroup [CW95b, Zha91]. Subgroups [Mal96, Cha96b, RGS98]. subinvariant [Lav99b]. Subject [CD92, Gee94, Leo94]. Submatrices [BH95, CS92, Cab95, NS97, dSZ97, Lon99]. Submatrix [BK97b, Sil96, Sil97d]. submaximal [GKW91]. submodular [Nar91]. submodule [Lev91]. submodules [BS99c, II98]. Subnormal [FI92, FI95]. Subnormality [Ins97]. suboptimal [FFT94]. subpencil [FS99a]. Subpermanental [Mal92]. Subpermanents [Kop96a]. subset [Mom98]. subsets [Chi91]. Subshifts [Fri98a, Fri97a]. Subspace [HS98, HR92, McE93, PS97, Spa97, Uhr92, Wan93b, Wan93a, Wu92b, ZGL99, BK99, CF99, Cho98, LST98, Mol96, Ong91, Ple98, Sid98, Sid99, Van91b]. Subspaces [ABBN94, BZ95, Beh94, BC95a, CGK96, CF97a, DN97, FGP94, FPP96, Gee93a, GKv98, GV95, GT96b, JB92, MBI96, RRR93, RR94, Rie97, Sch94, Val92, BR91, BC91b, BA– DR91, BRS91, CF99, CG99, DBRW91, FPP98, GI99, Gro98b, Han90b, JS90a, LN91, MR98b, RR90a, RT94, Sun91, TS99, Vel99, WP98]. Substitution [Hau94, Hau95]. Substitutions [Lan96]. Subtracted [MT93]. Subtractivity [HOŠS96, Gro99c]. Successive [HA94, HK93a, AGK^{+}98]. Such [EU93]. Sufficient [CG92b, CR94b, DRT93, Li95b, Väl96a, Väl97, Xu93, Yip96, ERR93, Li91b, Soh91, Väl96b]. Sufficiently [NS97]. suite [Le 94]. Sujit [BH94a]. Sum [DHS94b, Dru92b, Dru94a, HHNS96, Huy96, Kop96a, MP97d, SLK93, Wan95b, van96, DST98, DH90, IHM99, Lev91, WZ98, WXZ99]. Summability [MR98c]. Sums [BMTZ98, Bsh95, FNW96, Gol93, HHS97, LMR94, Nie93b, NM96, QS95, Sch92, SM93, Spi93, Spi95, TV92, Wan95c, Wu90, vLS95, CH97, FJ99, HW91, NS99a, Neu90, TV90]. Sumsets [BWS95]. Super [CF97b]. Super- [CF97b]. Superalgebras [AE93]. Superfast [Gut93, VK98]. Superior [HN93a]. Supplement [MT96]. supported [BP98]. Surface [Car92, LV97, dS95]. Surfaces [Ste95, KV90]. Surrogate [Kiw95a, Kiw97]. Survey [CR94a, Mer94b, Sha91b]. SVD [Han90a, KN98]. Swapping [BD93]. Sylvester [BK90, BK97a, BM94c, BG94, BG96, DS91, HP96, HM96, HR92, KS96, LR93, Lew99]. Sylvester-Like [DS91]. Symbolic [LZW97, Rei94]. Symbols [KR96b, Adu99]. symétrique [BC95b]. symétriques [ABM93]. Symmetric [ATZ95, BPY92, BK96a, Bap93, BOC95, BL93a, BB96, Bot96, CL98a, CNW96, Chi92a, Chi97, Chu93, CD97, CHR97b, Dai96, DER97, DIO97, DHS94b, Die96, Don92, Dru92b, Duc95, ENN98, EG93, Fan93, FN95a, FF96, FV93, FV92, Fri93, FK94, FH94b, Gib96a, Gib96b, Gon96, GS93, HR93a, HW93, IKF^{+}93, IAMC96, Iwa98, Jia97a, JOTV93, KPRV97, KS91, KN97, KSV95, Lag92, LY93b, LR95, Lei97, Lim95, LR94b, Ma93, Mah96, Mar94a, Mar94b, Mar97, Mes95a, Mur95a, Mur95b, PT95, Pat97a, Que94, Rad96, RRR93, Rou93, Sch93, Sla98, Ste97b, Tre93, Uhl94, Vos99, Wan92, Woo96, ABM93, Bog91, BC95b, Buf91, BGS99, CW93, CL90, Dai90, Dai99, Erx90, Fie90a, Fie90b, Fie98, FL99, Fri98b, HNN91, HP91, Hig99, HM99, JS99b, LT91b, Lim91a, LMWZ90]. symmetric [LSJ91, LST98, Lu98, MV98a, Mel99, Pap90, RR90a, Rum98b, Sch97, SY98, SP99, Tei99, Tho91, TB97, Tre99a, Wat90a, van98]. Symmetries [BKN97]. symmetrised [GB99]. Symmetrization [HRW92]. Symmetrized [Dia96c, Fon99]. Symmetry [AIY90, MHR93, Thi95, Wim94a, LW98a, Lei99a, Pre98, TY99b]. Symplectic [BF97, FL98, LY93c, Pat93, RW94, SL96, Yan93, AM91, BF98, BFW99, Die98, Ell94, LMX99, Wim91]. System [BK97c, BN93, CG97, DR93, Hod93, KL96, Mus94, RS97, gS96c, YC92, GPM99b, Nie98a, Ple98, Str98, VS94, VS99, van91a]. systèmes [RSDT97]. Systems [AB92, AP96a, AGL98, BCHS97, Bsh95, CN93, CNW96, Cla92a, CR94b, CG97, DS92, DZ98, Fei94a, Fei94b, FV96, FZ93, FM97b, Fun96b, Gee93a, Gee93b, Gee94, Gee96, GV92, GT96a, GK93, GTÖ96, Gut93, HY96, Har92, HR95, HAPSG95, HAPSG96, HH92, HK93b, HPW94, HO93, IW96, Jan97b, Jbi95, JPR^{+}96, Kam94, Kat94, KP94, Kno92, KSH94, LBC95, Lin92, LS97, LFMT94, Log98, Lop94, LP97, MM93b, Neu95, Ng97a, Nie93b, Now96, NU97, O’N93, OST93, PV94, Rat97, RW97, RV95, RV96, Rud96, SK95, SO94a, Shi95, Sid92, SDT97, Ter98, TS97, US92, Wai97, WR94, Wei94, Wu94, YA94, Zam94, dS95, van97a, van94, Aya99, Bai98, Bea98, BRMM91, Bro91a, BGS99, BGBMN99, BJW98, CWL99, CHASG98, Che91, CC94, CH99c, CMN99, Dat91]. systems [ETS90, FM93c, Gał91, GPM99a, GL91, GW99, GJS91, HN90, HO98, HASG90, HPWZ98, II98, JM99, JZ91, KKY91, KTK99, KvL93, Lev91, LWYS99, LN98, Mai91, MPdS97, Mel99, MM98, Nyl99, Pot98, PS98b, RG99, Sch97, Sha99a, Sid99, SW90, gS99b, Tis91a, VB97b, VK98, Van91c, XK94, ZPJZ98, de 91, RSDT97]. Systems-Their [YA94]. syzygy [Lin99].

Szego [AGR92, ACR96, Ols98, Tre99b].

t [Eva95]. Table [Gut93]. Tableaux [Col96]. Tampere [BS92a]. Tangent [HF98]. tangential [BK90]. Tapered [KZ93]. Tasci [Ano96b]. task [Coo98]. task-oriented [Coo98]. Taussky [Ano98b, Ano98d, Ano98e, Ano98f, Ano98h, Bau98, Joh98, Laf98b, Sch98b]. Taussky-Todd [Bau98, Sch98b]. Taylor [HRR92]. teaching [Dor98, Har99a]. Technion [BGH92b]. Technique [Nar95, AS99, Din93a, TZ91, Tis91a]. Techniques [HB97, Sid92, Bea98, CGV98, DB99, HB98]. Technology [BGH92b]. Teng [IC94]. Tensor [BR93, BZ94, DRVP94, GES96, GK97, MP98d, Pat99b, Pay95]. Tensors [Dia96a, Dia96c, Gon96, GPRL94, Lei97, Lim95, Lim97, Fon99, LW98a]. Term [Gee97, MT93, Mic91]. Terms [Gee94, She96a, Ion98]. Tessellation [LC96]. Test [GP94b, PT94b, Väl91]. Testing [Dah98, Van97b, von90]. Tests [CV97]. th [Hav93, Kit93, Ote90, Rea97, Sha91a, Sha95]. Their [AEG93, AD96a, Aze95, Bha94b, Chi92a, Chi92b, Hig95a, Jen97, LM98b, Neu96a, SH97b, Wim93, YA94, YW97, AZ90, BP91a, Bha91, CG98, FRS99, Hun90, KV90, Mal98, RBRF91, Sid98, SW98b, Tia98, hXyY98]. Them [Mir92]. Theme [dJ95a]. Theorem [AB96b, BAG96, BC98b, hCjXL96, CCHV97, CKL96, Che95, CMLS95, Ded97, Dru94b, Els93, Els95, Far98, Fie95b, FN95b, Goe96, GV96, HL97, Ito97, Kal96, Kim95, Kra93, LM92, LS93a, Lim95, McD94, MG94, Olk97, Pry92, Ral93, Ree93, Šem96, Spa95a, Str97b, Szu95, Tab93, Vel95, ZG94, AK98, Auj99, BH90, Bap91a, Bap94b, Bap98a, BAG99, BC98a, BN99, Cap98a, Chu97, Cro98, Dan90, EG99, FY98, Han98, HDB99, JN90, Kes99, Kol99, KS94c, LZ98a, Li98b, Loe97, MO91, NT94, RR90b, Sal99, SW91d, Wei98, Xie97b]. théorème [EK99, MO91]. Theorems [EH93, Els94, HM92b, JL94, Kop96b, Mil95, Mit94, Nab96, Rum97, Tyr96, Wan98b, Wim94a, Woź97, Xin93b, Yan93, CS91, Cap98b, CP98b, CP99b, DR99, KT91, LRT99, WCM99]. Theoretic [FOV97, Rei94, Rud96, Ste97b, van94]. Theoretical [Fre91a, WY96a]. Theoretically [RR97]. Theory [And94, Ano96a, BS92a, BGH92b, BvR^{+}97, Dat94, DSW96, ERS95, Ebi95, Far96, Far97, FB96, Fuh94a, GKv93, GTZ97, Grö93, Gro97b, HH92, HL96b, IW93, IO96, JPR^{+}96, Jon95b, Kap95, KP92a, Len94, Li97c, Lyu95, Met97, Mir92, Rav97, RS97, Sha92, Ste94, Uhl98, WF92, Xue97, ABGR94a, BM90, Bon94, Dor98, FG99, FFT90, Gro98b, GKS94, HY91, Joh98, Koš94, LS93b, Li94b, Li96, Mit91, Rus94, Szu90a, TZ98, Wan97b, WW97, Wei98, BGH92a]. Theta [AK97]. Thiele [Gu99b]. Thiele-type [Gu99b]. Thin [TZ99]. Third [BS92a]. Thompson [AH94b, And96b, Bd97, DSW96, HP93, Tam97]. Thompson’s [Tam98]. Thorem [BB93, FFK93b]. Three [HL92, Kop96b, Pat97c, SSHS92, DBRW91, KB99, Mes97a, NDS99, SW90, TK99, Wat90b]. Three-Dimensional [SSHS92, DBRW91]. three-vectors [Mes97a]. three-way [TK99]. threefold [HP90, HP99]. tightest [FP99]. Tilings [Zho96]. Time [BK93, CBSH94, CR94b, DS92, Fei94a, Fei94b, FI99, Gee93b, Gee94, Gee96, Gee97, HC98b, IW92, IW93, IO96, Ion96, Ion98, Kam94, Kat94, KP94, Log98, Meh96, Oar96, RR96a, RR96b, Rud96, gS97a, Ter98, Vro93, Wim94b, XL95, Che91, Dru99c, ETS90, GL91, HPWZ98, Kat99, KTK99, LL93b, Nie90, gS98b, Väl91, VS94, Wim91, YP91, dOGH99]. Time-Dependent [RR96a, RR96b]. Time-Invariant [DS92, Gee93b, Kam94, VS94]. time-semidefinite [Dru99c]. Time-Varying [Fei94a, Fei94b, Rud96, Ter98, FI99, Ion98, ETS90]. times [Pal90]. Toda [ADH97, Chu95, Fay94]. Toda-like [ADH97, Chu95]. Todd [Ano98b, Ano98d, Ano98e, Ano98f, Ano98h, Bau98, Sch98b]. Toeplitz [Adu98, Adu99, ABL98, BM98a, BT91b, BM94a, BBB95, BP93, BM98b, Bol98, BG98b, Boz95, BJW98, Cap98b, CN93, CNW96, Chu93, CCR97, Dah90, Di 97, Di 98, DZ98, EL92, EGL96, Eng95, FZ93, FK93, GTVV96, Gem97b, GK93, Gor92, Gou93, Gov94, Gut93, GH95b, Haf95, HN98, Hei92, HJ92, Hei95a, HB97, HR98b, HR98c, HB98, Huc98, KO97a, Kar94, KH90, KST99, LS92a, LS94, LV91b, Lin92, LW98b, MV98a, MMP99, MP97a, Mel99, Ng97a, Ohs95, PD92, PS98b, PT99, Roc98, RS98, SLK93, Ser97, Ser98b, Sil93a, Sil97a, Ste97c, Til98a, Til98b, Tis91b, TR92, Tre90, Tre93, Tre99a, Tyr91, Tyr94, TZ98, TZ99, VB97b, Vos99, Woe94, Woo93, Xu92]. Toeplitz-Block [Ng97a]. Toeplitz-like [BP93, KO97a]. Toeplitz-plus-Hankel [HB97, HR98b, HR98c, HB98]. Tolerance [BMV95]. Tolerate [Mus94]. tomographic [EKN90]. tomography [KLS90, SS90]. tool [HN99]. Topics [Far98, Gon96, Uhl92]. Topological [AM95, FP96, Bak99]. Topology [DS94b, JAG98]. Total [De 93, FB96, GP92, GP94a, Gla98, Wei98]. Totally [BA– DR91, CP95, GS93, Hei94, Peñ95, Ric90a, Rie97, SS95a, Pin98]. Touching [Uhr92]. Toughness [Bro95]. Tournament [FR95a, FR95b, FK94, KS94a, Kir95, Kir96b, Kir97a, Kir97c, dGK^{+}92, Bor95a, Kir99, NP96]. Tournaments [BGS92]. Trace [BT95b, Dru98, Fur96, Hap97, HP93, JLOvdD96, KH95a, KSV95, LPW92, MV97, MT93, MT96, OS96b, OK96, Rei93, Sha99b, Tam97, Wan98a, Zie96, Abd95, BP90c, DR91a, JSOvdD94, Kit91, LM99, Lei99a, LMWZ90, da 91]. trace-preserving [JSOvdD94]. Traces [Bot97, LC92, Mar94b, MST94, ORR94]. Tracking [Har92, Spa95b]. Tracy [Liu99d].

Trades [KM95]. train [Cos91, Var96].

Trajectories [Gee94]. Transfer [ETS90, Nie93b, Wen96, Fli94]. Transfer-like [Nie93b]. transfert [Fli94]. Transform [AMM97, CNW96, FBO97, Di 98, Eng95, Fli94, GM98a, Ste90, BT97b]. Transformation [BNM92, DeA96, HB97, HB98, At91, BRMM91, Fir90, Fli94, Hab94, Hei97]. Transformations [Boa97, DHL97, Duf94, Fun96b, HB97, KH92, LMY94, Len94, NO95, Nov97, Ree93, SH97a, Wat97b, Wu92b, ŽWB93, Atk93, BLL94, Fer99, HR98c, Idr99, Ric90b, Vin90]. Transforms [KT95a, MP97c, MHR93, Ohs95, PST98]. Transistor [GW95]. transition [JZ99]. Transitive [BGS92, FRS99, PT95, PPY93, PPY95, Bor95a, NP96]. Transitively [AP94a]. translates [JS90b]. Transmission [O’N93, Wat96b, Ols98]. Transport [EG97, Jon95b]. transportation [Die90, LSJ91]. Transpose [GT93, Gow90, Sol99]. transvections [EM91, Ell94]. traveling [War90]. Treatment [Meh96, SB90]. Tree [JT97, MZ91, PPY95, An98, Nyl98, Rin98, SJ95]. tree-patterned [Nyl98]. Trees [CL97a, CL98b, Hof97, Sha95, Mer91b, Sha91a]. Triality [Hig95b]. Triangluar [Wim93]. triangulable [HO91]. Triangular [ACN92, CHS92, CH96b, CL98e, CM93, GKL96, Gol97c, GRS92, Her92a, Jøn95a, JTU97, Kop96b, Kru95, LZ96, PPY93, RR99a, RS92, SLK93, SS95a, Sil93b, Sil97b, gS96c, Tak96, Thi97, Wat95b, BT91b, BW91, BZ98, CU99, CX98a, Fri91a, Ger98, Hua91, JS91, KSW99, MS99a, NDS99, Sto98]. triangularizability [Hua99]. Triangularizable [Okn97]. Triangularization [GH92, Iwa98, BG98a, Chr99]. tribute [Bru99c]. Tridiagonal [AB92, BOC95, Chi96c, Gov94, IAMC96, Lop94, LP97, MP97a, New94, RR92b, Sch93, SJ96, Bap91a, CLW90, Hig99, KST99, Lu98, Nab99, Usm94]. Trigonometric [AL95a, Næv97, HR98c, PS98b]. Trigonometry [Gus95a]. Trilinear [Cox95]. triples [GPM97]. Triplets [Hon92]. Trudi [Gre92]. Truncated [San96b, nCjH98b, Fle90]. Truncation [ACN92, FP96, Wat95b]. trust [ACWY99]. trust-region [ACWY99]. Two [ABGR94b, AB95, Auj93, BFW98, CWL99, CS96, CL96a, CG92b, CF97b, DN97, Dru92b, Dru94a, EHJL97, Far92, FP97b, GES96, GTÖ96, Hof97, JPR^{+}96, Kol95, Koš96, Kri95, LS92a, LL91a, LMMR95, Mes95a, Moo95, Nab99, PV95, SW95, Spi94, gS96b, VW95, WW94, Wan95c, WLN99, Xu97b, Xu97a, Yan93, BT91a, BKZ98, BP98, Bot99, CT91, Cao98b, CG99, Daz98, DH90, GI99, KB99, LWX99, Mat91, Nie99a, Obe91, San91, Spe95, Wer90a]. Two-Block [GTÖ96]. Two-Dimensional [CS96, JPR^{+}96]. Two-generated [LL91a]. Two-Graphs [Moo95, Spe95]. Two-Level [Kri95]. Two-Parameter [Koš96, Mes95a]. two-periodic [Nie99a]. two-scale [BP98]. Two-Sided [ABGR94b, AB95, CL96a, Kol95, Nab99]. Two-stage [CWL99, Cao98b]. Two-Step [LMMR95]. Two-Way [PV95, WLN99]. Type [AH94b, And96b, BM94c, BKMS97, CW95b, Dru96, Eğe95, Gut98, HR93b, HLMT98, Huc98, LT97b, Nie98c, Tis92, Wan96b, Wan98b, ACWY99, Bai98, BKO99, BGS99, DR99, Ede91, Fri97a, GHN99, Gu99b, KP92b, KP95, LRT99, PI91, Rum97, Wan97b, WE91, vV90, CX98b]. types [FKK98]. typical [TK99].

Ultimate [BT95a]. Ultrametric [MNST95, NV95a, TV99]. Unambiguous [EHJL97]. unbalanced [GB99]. unbiased [BP90b, Sch99a]. Unbounded [DR96, Tyr95]. uncertain [LWYS99]. uncertainties [SNC98]. Unchanging [DMB95]. Unconditional [CS96]. Unconstrained [BSK94]. Uncontrollability [GN93, OTV93]. Underdetermined [HM96]. Underlie [RK96]. understanding [Har99a]. undirected [Li98c, Yon99]. Unfolding [BK97c]. unification [Joh98]. Unified [CS91, FH94b, Gro97b, Meh96, ZJ91, BM90, JS99a, Lew99, Mit91, Sid99, Vav91]. Uniform [JL98b, BRMM91, KN99, Li97d]. Uniformly [KZ93, LM98b]. Unifying [Tyr96]. Unimodular [CP95, Mer92, Mer94c, Wat94, Jam91a]. Unipotent [GT93, WW91]. unique [WLL99]. Uniqueness [FH94a, Gee94, Goe96, Lim95, Pee95, Xu99]. Unit [Gen96b, Mar94b, Mar94c, Mar96, Mar97, Str97a, Tof97, Mes97a]. Unitarily [Hor95, Mar94a, Mar94c, Mar97, Xie97a, Fur94]. Unitary [BR95, BKN97, Dob99, DS94c, Dru92a, Dru94b, EH93, EI97, Fle97, Fri96, Idr96, LM94b, LT95, Pis97, Ser98a, SC94, ABGR94a, AH95, BGE91, FJK98, Ols98, Sha91b, Wim99]. Units [Roe97]. Univariate [GV96]. Universally [FNW91]. Universidad [BdV92]. University [BHR91, ORS94]. unrestricted [BP90a]. Unstable [GTÖ96, OST93]. Unsymmetric [HK93a, Jia97b, Jia98a]. unzulässige [Pus90]. Update [Lam93, EI98]. Updates [Hil92]. Updating [Has96, MVV93, Di 97, WZ99]. Upper [ATST95, CZ97, CM93, GRS92, JL97, Jøn95a, JTU97, Kru97, LZ96, LZ97a, MUV^{+}97, RS92, SS95a, SJ96, Sil97b, Tak96, TK98, Thi97, Tho97, Tuc96, BT91b, CX98a, HKM98]. Use [Har99b, PC90, BP90b]. Using [ACE95, BK97a, BNP93, Che95, DZ95, Fri98a, GKP95, HOŠS96, HR94, Jin98, Kol98, LS94, LM94a, MST94, MV97, PŠ98a, Sim96, Sun94, Van97b, Atk93, DF98, Din93a, Gad90, HR98c, HPWZ98, Kra91, LN98, Meh91, Sti91, WZ99]. USSOR [HK93a, MTS96]. Utah [BHR91].

Valencia [BdV92]. Validity [Jen96, BKdP94]. valuation [Adk91]. Value [Bha94a, Bha95, BOS93, BR95, Che92, De 92, DD97b, GLW95b, HM92a, HP92, Hon92, Jia94, JS98a, Kir96b, LS98, MS93c, MT94, MZ91, PV95, Peñ97, Ren96, Zha96, D^{+}99, HJ90, Lev98, Wat90a, YhG97]. Valued [CK94, GM95a, SV94, AD90b, AD90a, BK91b, CL99, KM94]. Values [BP92a, BKL97, BRL98, DG93, Eie93, EF95, GKK96, Lin98c, LC92, MST94, MT96, Nie97b, Que94, QS95, Sil93a, Sim96, SV96, Sta93, Tam98, Til98b, Tyr94, WX97, Bap91b, Chu99, CL91, Dei91, Gam97, JN90, KK90, LT91b, Li98c, Liu99b, MS91b, RS98, Tar91, Van91b]. Vandermonde [CR92, CK91, EFP98, FH96a, FHR93, HR95, Hei98a, KO97b, MP98b, Ols98, RV95, RV96, RV98, Vav97]. Vanishing [Dom96, JM93]. Variable [HS98, Kit93, MH96, Rat97, AA90]. Variables [Auj93, BW96, dR98, Neu90]. Variance [Neu96a, ZD92, NW90]. variant [Kes99, dRT91]. variants [Cao97]. Variate [Hol96, MP97c]. Variation [Far92, Hol92, Kra94, Li97c, Spa95a, Sun96a, BEK90, Dru98, Li90b, Phi91]. Variational [BY95c, Man92, SY95]. Variations [CCHV97, cL93, Li94c, dJ95a, Li98d]. varieties [Coo97]. various [HO98]. Varying [Fei94a, Fei94b, Rud96, Ter98, Art91, ETS90, FI99, Ion98]. vecb [KNW91]. Vector [BRL98, BHP97, nCZpZ96, GM95a, GKP92, HR94, Hol96, Jen91, Kir97c, Kol98, Lin93, Sad93, SH93, Wan93b, Wan93a, WC97, AD90b, AD90a, Dor98, HR98b, HR98c, Kat99, Kir99, Le 94, Lom99, Sal98]. Vector-Valued [GM95a]. vectorielle [Le 94]. Vectors [AC97, Din92, GO94, Jel92, Kir97a, KS94b, LTZ97, Mes97a, Mub99]. Vegas [SL97]. Verdiere [vLS95]. vérifiant [MV95]. Vermutung [Pus90]. Version [hCjXL96, Had94, Šem96, BH90, Bap98a, WI99]. Versions [Lin98a, MPŠV97, DR99, WCM99]. Versus [van96, AIY90, BW96, Jam91b]. Vertex [BBD^{+}97, Meh91]. Vertical [SG95]. vertices [AH90, HOR98, Spe95]. via [AB96b, BK96c, BK99, BHP97, ETS90, GS98a, HM92a, Hwa99, Kat94, LBC95, Ols98, TZ98, Van90, WC97, vvS99]. Vibration [HL96b]. view [Par98a]. VII [Ano95a]. Virtanen [Dru92a]. vitae [Ano98b]. Vlastimil [Vav95]. Volume [Ano97, BB92, BI92, Car92]. Volumes [Ano95b, Ano95c, Vel95, Ano98g, Ano94a]. Vorobyev [Bre96]. Voting [FR95b].

W [Bon94, DCPS96]. Walk [GM95b, Pal90]. Walker [Tis92]. Wang [CA98]. Watkins [Nas92]. Watson [Car97]. Wavelet [FP96, KT95a]. Wavelets [KT95b, Zho96]. Way [HK97b, LS98, PV95, TK99, WLN99]. WBJ [FLM94]. Weak [Gee93b, Gee96, LZ96, Ma93, hXyY98, CP98b, CP99b, NS99b, TY99b, Vis95]. Weakly [Ata92, DdlP96, JJPT90]. Wedderburn [BRW94, LM92]. Wedderburn-Forney [BRW94]. weight [Fio99]. Weighted [AG94, AG97, BJP97, MR98c, Ng97b, Pol92, SRT90, Wan97a, AdlP98, Cow90, Har94, WZ99]. Weights [LL93a, MP94b, Shi99b]. Weitere [Pus90]. West [ORS94]. Where [BHM98]. Which [AEG93, BP92b, BS94a, BY95b, Fun96b, HN93d, Hua96b, Kye95, Val92, Wu92b, Bor95a, Cai99b, DL92, DL01, Fri91a, Loe90, SW98a]. Whittaker [MP97c, MP98c]. Whose [BH95, Kir95, LPS98, Peñ95, WN95, AYC91, TY99b]. Wielandt [BB93, Dru94b, Els93, FN98a, Gus99, HRW92, Mar92, WI99]. wild [Lak99]. Wilkinson [PD97, Pus90]. Winnipeg [OPSS98]. Wishart [Ede91, Per96]. Without [AM95, Gee94, GKS94, Rum97]. Woodbury [LV97].

Woodbury-Formula-Based [LV97]. work [BH94a, Laf98b, Vav95]. Workshop [BS92a, BMH94].

X [Ano98c, Eva95, LM97]. X-Ray [LM97].

Y-MP [Pap92]. Yakubovich [AB96b]. Yamamoto [BAG99, JN90]. year [Ano94a]. York [Bon94, Li98a]. Youla [SSS97]. Young [BHR91, Wei98, GHN99]. younger [Per98]. Yule [Tis92]. Yule-Walker [Tis92].

Z [KKNU97]. Z-Matrices [KKNU97]. Zero [AM95, BK97c, CHR97a, Fie94, Gem96, GM93, GOV96, JM93, Men97, Min94, Zie96, de 96a, BR90, Gro98b, Kit91, LM99, LMWZ90, Van90]. Zero-One [de 96a]. zero-pole [BR90]. Zero-Trace [Zie96, Kit91]. zeroes [BZ98, HL99]. Zeros [ACR96, CALM97, Gol97c, Hod93, IKF^{+}93, Lin98c, Soh97, CL98d, Chi90, WZ98, WSCP91]. Zhang [Bha97a]. ZME [Stu90]. ZME-matrix [Stu90].

References

Anderson:1990:RIS

[AA90] B. D. O. Anderson and A. C.

Antoulas. Rational interpolation and state-variable realizations. Linear Algebra and its Applications, 137/138:479– 509, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ashbaugh:1990:SEI

[AB90] Mark S. Ashbaugh and Rafael D. Benguria. Some eigenvalue inequalities for a class of Jacobi matrices. Linear Algebra and its Applications, 136:215–234, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ashbaugh:1991:ASE

[AB91] Mark S. Ashbaugh and Rafael D. Benguria. Addendum: “Some eigenvalue inequalities for a class of Jacobi matrices”. Linear Algebra and its Applications, 149: 277, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amodio:1992:PFP

[AB92] P. Amodio and L. Brugnano. Parallel factorization and parallel solvers for tridiagonal linear systems. Linear Algebra and its Applications, 172: 347–??, July 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1994:GMP

[AB94a] Daniel Alpay and Vladimir Bolotnikov. On a general moment problem and certain matrix equations. Linear Algebra and its Applications, 203/204: 3–43, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:1994:SCM

[AB94b] T. Ando and Richard A. Brualdi. Sign-central matrices. Linear Algebra and its Applications, 208/209:283– 295, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1995:TSI

[AB95] Daniel Alpay and Vladimir Bolotnikov. Two-sided interpolation for matrix functions with entries in the Hardy space. Linear Algebra and its Applications, 223/224(1– 3):31–56, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400074. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Abeledo:1996:SML

[AB96a] Hernán Abeledo and Yosef Blum. Stable matchings and linear programming. Linear Algebra and its Applications, 245(1–3):321–333, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9500052.

Arov:1996:SMI

[AB96b] D. Z. Arov and N. V. Bondarchuk. On the solution of matrix inequalities in the Kalman-Yakubovich theorem via hidden parameters of positive rational functions. Linear Algebra and its Applications, 246(1–3):1– 11, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400244.

Azizov:1994:NJS

[ABBN94] T. Ja. Azizov, P. A. Binding, J. Bognár, and B. Najman. Nondegenerate Jordan subspaces of self-adjoint operators in indefinite spaces. Linear Algebra and its Applications, 207:37–48, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Atkin:1997:PCB

[ABCP97] A. O. L. Atkin, Endre Boros, Katarína Cechlárová, and Uri N. Peled. Powers of circulants in bottleneck algebra. Linear Algebra and its Applications, 258(1–3):137– 148, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600198.

Anderson:1992:GFA

[ABD92] E. Anderson, Z. Bai, and J. Dongarra. Generalized QR factorization and its applications. Linear Algebra and its Applications, 162/164:243– 271, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Directions in matrix theory (Auburn, AL, 1990).

Abdelhak:1995:APC

[Abd95] Achouri Abdelhak. Approximation positive contractante en norme trace. (French) [approximation in the trace norm by a positive contraction]. Linear Algebra and its Applications, 225(1):163– 173, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300336.

Aberth:1997:SLI

[Abe97] Oliver Aberth. The solution of linear interval equations by a linear programming method. Linear Algebra and its Applications, 259(1–3): 271–279, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1997&volume=259&issue=1-
                                        3&aid=9600291.

Alpay:1994:UPA

[ABGR94a] Daniel Alpay, Joseph A. Ball, Israel Gohberg, and Leiba Rodman. J-unitary preserving automorphisms of rational matrix functions: state space theory, interpolation, and factorization. Linear Algebra and its Applications, 197/198: 531–566, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Alpay:1994:TSR

[ABGR94b] Daniel Alpay, Joseph A. Ball, Israel Gohberg, and Leiba Rodman. The two-sided residue interpolation in the Stieltjes class for matrix functions. Linear Algebra and its Applications, 208/209:485– 521, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:1990:SMR

[ABKW90] A. C. Antoulas, J. A. Ball, J. Kang, and J. C. Willems. On the solution of the minimal rational interpolation problem. Linear Algebra and its Applications, 137/138:511– 573, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1998:NPE

[ABL98] Daniel Alpay, Vladimir Bolotnikov, and Philippe Loubaton. On a new positive extension problem for block Toeplitz matrices. Linear Algebra and its Applications, 268(1–3):247–287, January 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700048.

Alcalde:1993:AIP

[ABM93] M. T. Alcalde, C. Burgueño, and C. Mallol. Les Pol(n, m)-algèbres: identités polynômiales symétriques dans des algèbres. (French) [Pol(n, m)- algebras: symmetric polynomial identities in algebras]. Linear Algebra and its Applications, 191:213–234, September 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abt:1994:NPC

[Abt94] Markus Abt. A note on the product correlation rule. Linear Algebra and its Applications, 199:171–177, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Augot:1997:CMP

[AC97] Daniel Augot and Paul Camion. On the computation of minimal polynomials, cyclic vectors, and Frobenius forms. Linear Algebra and its Applications, 260(1– 3):61–94, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600290.

Andersson:1995:CCM

[ACE95] Lars-Erik Andersson, Gengzhe Chang, and Tommy Elfving. Criteria for copositive matrices using simplices and barycentric coordinates. Linear Algebra and its Applications, 220(1–3):9– 30, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400161. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Arioli:1996:PMC

[ACF96] M. Arioli, B. Codenotti, and C. Fassino. The Padé method for computing the matrix exponential. Linear Algebra and its Applications, 240(1–3):

111–130, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400190.

Achilles:1993:ICR

[Ach93] Eva Achilles. Implications of convergence rates in Sinkhorn balancing. Linear Algebra and its Applications, 187:109–112, July 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Angelos:1992:TTF

[ACN92] James R. Angelos, Carl C.

Cowen, and Sivaram K. Narayan. Triangular truncation and finding the norm of a Hadamard multiplier. Linear Algebra and its Applications, 170:117–??, June 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ammar:1996:CMC

[ACR96] G. S. Ammar, D. Calvetti, and L. Reichel. Continuation methods for the computation of zeros of Szegő polynomials. Linear Algebra and its Applications, 249(1–3):125–155, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500324.

Andruchow:1997:GOI

[ACS97] E. Andruchow, G. Corach, and D. Stojanoff. Geometric operator inequalities. Linear Algebra and its Applications, 258(1–3):295– 310, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600201.

Anstreicher:1999:SDT

[ACWY99] Kurt Anstreicher, Xin Chen, Henry Wolkowicz, and Ya-Xiang Yuan. Strong duality for a trust-region type relaxation of the quadratic assignment problem. Linear Algebra and its Applications, 301(1–3):121–136, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/32/article.
pdf.

Alpay:1990:SISb

[AD90a] Daniel Alpay and Harry Dym. Structured invariant spaces of vector valued functions, sesquilinear forms, and a generalization of the Iohvidov laws. Linear Algebra and its Applications, 137/138:413– 451, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1990:SISa

[AD90b] Daniel Alpay and Harry Dym.

Structured invariant spaces of vector valued rational functions, Hermitian matrices, and a generalization of the Iohvidov laws. Linear Algebra and its Applications, 137/138:137– 181, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1993:NCR

[AD93] Daniel Alpay and Harry Dym.

On a new class of reproducing kernel spaces and a new generalization of the Iohvidov laws. Linear Algebra and its Applications, 178:109–183, January 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:1996:NCR

[AD96a] Daniel Alpay and Harry Dym.

On a new class of realization formulas and their application. Linear Algebra and its Applications, 241/243(1–3):3– 84, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500602. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Ambikkumar:1996:RHI

[AD96b] S. Ambikkumar and S. W. Drury. A reverse Hadamard inequality. Linear Algebra and its Applications, 247(1–3):83– 95, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9689196.

Ambikkumar:1997:SRC

[AD97] S. Ambikkumar and S. W. Drury. Some remarks on a conjecture of Boyle and Handelman. Linear Algebra and its Applications, 264(1–3):63–99, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9600402.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9600402.

Ashlock:1997:MSI

[ADH97] Daniel A. Ashlock, Kenneth R. Driessel, and Irvin R. Hentzel. On matrix structures invariant under Todalike isospectral flows. Linear Algebra and its Applications, 254(1–3):29–48, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600523. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Adkins:1991:NMH

[Adk91] William A. Adkins. Normal matrices over Hermitian discrete valuation rings. Linear Algebra and its Applications, 157:165–174, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:1998:CIW

[AdlP98] O. Araujo and J. A. de la Peña. The connectivity index of a weighted graph. Linear Algebra and its Applications, 283(1–3):171–177, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6215.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=283&issue=
                                        1-3&aid=6215.

Adukov:1998:GIB

[Adu98] Victor M. Adukov. Generalized inversion of block Toeplitz matrices. Linear Algebra and its Applications, 274(1–3):85–124, April 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=274&issue=1-
                                        3&aid=9700304.

Adukov:1999:GIF

[Adu99] Victor M. Adukov. Generalized inversion of finite rank Hankel and Toeplitz operators with rational matrix symbols. Linear Algebra and its Applications, 290(1–3):119–134, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        290/1-3/6322.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=290&issue=
                                        1-3&aid=6322.

Albuquerque:1993:GLS

[AE93] Helena Albuquerque and Alberto Elduque. On the generators of Lie superalgebras.

Linear Algebra and its Applications, 181:45–??, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adkins:1993:MDF

[AEG93] William A. Adkins, Jean-Claude Evard, and Robert M. Guralnick. Matrices over differential fields which commute with their derivative. Linear Algebra and its Applications, 190:253–261, September 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Apitzsch:1990:SAA

[AFK90] W. Apitzsch, B. Fritzsche, and B. Kirstein. A Schur analysis approach to minimum distance problems. Linear Algebra and its Applications, 141: 105–122, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arens:1993:QSJ

[AG93] Richard Arens and Moshe Goldberg. Quadrative seminorms and Jordan structures on algebras. Linear Algebra and its Applications, 181: 269–??, March 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arens:1994:WNM

[AG94] Richard Arens and Moshe Goldberg. Weighted ℓ_{∞} norms for matrices. Linear Algebra and its Applications, 201: 155–163, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arens:1997:MOW

[AG97] Richard Arens and Moshe Goldberg. Multilinear operators and weighted L^{1} norms. Linear Algebra and its Applications, 267(1–3):1–10, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9600641.

Abdel-Ghaffar:1998:LDL

[AG98] Khaled A. S. Abdel-Ghaffar. Long division for Laurent series matrices and the optimal assignment problem. Linear Algebra and its Applications, 280(2–3):189–197, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6142.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6142.

Angelos:1998:LCS

[AGK^{+}98] James Angelos, George Grossman, Edwin Kaufman, Terry Lenker, and Leela Rakesh. Limit cycles for successive projections onto hyperplanes in R^{n}. Linear Algebra and its Applications, 285(1–3): 201–228, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6233.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6233.

Anderson:1998:LPL

[AGL98] Edward J. Anderson, Miguel A.

Goberna, and Marco A. López. Locally polyhedral linear inequality systems. Linear Algebra and its Applications, 270(1–3):231– 253, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700241.

Ara:1997:DMR

[AGMP97] P. Ara, K. R. Goodearl, K. C. O. Meara, and E. Pardo. Diagonalization of matrices over regular rings. Linear Algebra and its Applications, 265(1–3):147–163, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600596.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600596.

Ammar:1992:DSP

[AGR92] G. S. Ammar, W. B. Gragg, and L. Reichel. Downdating of Szegő polynomials and data-fitting applications. Linear Algebra and its Applications, 172:315–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agha:1990:RBH

[AH90] Nader Agha and Daniel Hershkowitz. The relation between the height and level characteristics for a matrix with simple singular vertices. Linear Algebra and its Applications, 142:113–119, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:1994:IBP

[AH94a] T. Ando and F. Hiai. Inequality between powers of positive semidefinite matrices. Linear Algebra and its Applications, 208/209:65–71, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:1994:LMC

[AH94b] Tsuyoshi Ando and Fumio Hiai. Log majorization and complemetary Golden-Thompson type inequalities. Linear Algebra and its Applications, 197/198:113–131, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Ammar:1995:IEP

[AH95] Gregory S. Ammar and Chun Yang He. On an inverse eigenvalue problem for unitary Hessenberg matrices. Linear Algebra and its Applications, 218(1–3):263– 271, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300188.

Avdelas:1997:OCM

[AH97] George Avdelas and Apostolos Hadjidimos. Optimal 2-cyclic MSOR for “bowtie” spectra and the “continuous” Manteuffel algorithm. Linear Algebra and its Applications, 265(1–3):29–54, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600410.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600410.

Arasu:1991:MCD

[AHJP91] K. T. Arasu, Willem H. Haemers, Dieter Jungnickel, and Alexander Pott. Matrix constructions of divisible designs. Linear Algebra and its Applications, 153:123–133, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agler:1998:CHM

[AHS98] Jim Agler, J. William Helton, and Mark Stankus. Classification of hereditary matrices. Linear Algebra and its Applications, 274(1–3):125– 160, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=274&issue=1-
                                        3&aid=9700307.

Arimoto:1996:SPD

[AI96] Akio Arimoto and Takashi Ito. Singularly positive definite sequences and parametrization of extreme points. Linear Algebra and its Applications, 239(1–3):127–149, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400176.

Ajupov:1990:SVF

[AIY90] Sh. A. Ajupov, B. Iochum, and N. Dzh. Yadgorov. Symmetry versus facial homogeneity for self-dual cones. Linear Algebra and its Applications, 142:83–89, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adin:1997:SEM

[AK97] Ron M. Adin and Yaacov Kopeliovich. Short eigenvectors and multidimensional theta functions. Linear Algebra and its Applications, 257(1–3):49–63, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600037.

Alpin:1998:IPR

[AK98] Yu. A. Alpin and L. Yu.

Kolotilina. Inequalities for the Perron root related to Levinger’s theorem. Linear Algebra and its Applications, 283(1–3):99–113, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6203.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6203.

Alpay:1995:PTM

[AL95a] Daniel Alpay and Philippe Loubaton. The partial trigonometric moment problem on an interval: The matrix case. Linear Algebra and its Applications, 225(1):141– 161, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300330.

Artzrouni:1995:NCE

[AL95b] Marc Artzrouni and Xuefeng Li. A note on the coefficient of ergodicity of a column-allowable nonnegative matrix. Linear Algebra and its Applications, 214(1–3):93– 101, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1995&volume=214&issue=1-
3&aid=9300058.

Andrade:1996:CBA

[AL96] Raúl Andrade and Alicia Labra. On a class of Baric algebras. Linear Algebra and its Applications, 245(1–3):49– 53, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400207.

Ahues:1998:EBE

[AL98] Mario Ahues and Balmohan Limaye. On error bounds for eigenvalues of a matrix pencil. Linear Algebra and its Applications, 268(1–3):71– 89, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9789324.

Alaimia:1999:ASB

[Ala99]  M. R. Alaimia.  Auto-

morphisms of some Banach algebras of analytic functions. Linear Algebra and its Applications, 298 (1–3):87–97, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/105/17/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/23/article.
                                        pdf.

Albrecht:1996:MNN

[Alb96] J. Albrecht. Minimal norms of nonnegative irreducible matrices. Linear Algebra and its Applications, 249(1–3):255– 258, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500360.

Althofer:1994:SAR

[Alt94] Ingo Althöfer. On sparse approximations to randomized strategies and convex combinations. Linear Algebra and its Applications, 199:339–355, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ammar:1991:HSH

[AM91] Gregory Ammar and Volker Mehrmann. On Hamiltonian and symplectic Hessenberg forms. Linear Algebra and its Applications, 149:55– 72, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alefeld:1993:CMI

[AM93] G. Alefeld and G. Mayer. The Cholesky method for interval data. Linear Algebra and its Applications, 194:161–182, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arizmendi:1995:AGT

[AM95] H. Arizmendi and V. Müller.

On algebras without generalized topological divisors of zero. Linear Algebra and its Applications, 223/224(1– 3):65–71, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400355. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Andreani:1998:SEL

[AM98] Roberto Andreani and José Mario Martínez. On the solution of the extended linear complementarity problem. Linear Algebra and its Applications, 281(1–3):247–257, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6162.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6162.

Aharoni:1997:FTI

[AMM97] Ron Aharoni, Milton Maritz, and Roy Meshulam. Fourier transform imitations. Linear Algebra and its Applications, 257(1–3):201–209, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=257&issue=1-
                                        3&aid=9600153.

Ammar:1999:SLF

[AMM99] Gregory Ammar, Christian Mehl, and Volker Mehrmann. Schur-like forms for matrix Lie groups, Lie algebras and Jordan algebras. Linear Algebra and its Applications, 287(1–3):11–39, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6250.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6250.

Alic:1997:AGH

[AMPV97] M. Alić, B. Mond, J. Pečarić, and V. Volenec. The arithmetic-geometric-harmonic-mean and related matrix inequalities. Linear Algebra and its Applications, 264(1–3):55–62, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600471.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600471.

An:1998:BSL

[An98] Chang An. Bounds on the second largest eigenvalue of a tree with perfect matchings. Linear Algebra and its Applications, 283(1–3): 247–255, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6225.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6225.

Ando:1994:MIM

[And94] T. Ando. Majorizations and inequalities in matrix theory. Linear Algebra and its Applications, 199:17–67, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anderson:1995:LCN

[And95a] Charles A. Anderson. Loop and cyclic niche graphs. Linear Algebra and its Applications, 217(1–3):5–13, March 1, 1995. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=217&issue=1-
                                        3&aid=9400154. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Ando:1995:MRH

[And95b] T. Ando. Majorization relations for Hadamard products. Linear Algebra and its Applications, 223/224(1– 3):57–64, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400014. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Anderson:1996:SEE

[And96a] Joel Anderson. A secular equation for the eigenvalues of a diagonal matrix perturbation. Linear Algebra and its Applications, 246(1–3):49– 70, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400314.

Ando:1996:HPG

[And96b] T. Ando. Hadamard products and Golden-Thompson type inequalities. Linear Algebra and its Applications, 241/243(1–3):105–112, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500094. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Anonymous:1992:P

[Ano92] Anonymous. Preface. Linear Algebra and its Applications, 167:1–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1994:YIV

[Ano94a] Anonymous. 20 year index:

Volumes 101–200. Linear Algebra and its Applications, 200:83, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1994:BR

[Ano94b] Anonymous. Book reviews. Linear Algebra and its Applications, 200:5–??, March 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1994:CP

[Ano94c] Anonymous. Conference proceedings. Linear Algebra and its Applications, 200:3– ??, March 15, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1994:SI

[Ano94d] Anonymous. Special issues. Linear Algebra and its Applications, 200:1–??, March 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1995:AAH

[Ano95a] Anonymous. Announcement: Alston S. Householder Award VII. Linear Algebra and its Applications, 230(1–3):1– ??, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1995:AIV

[Ano95b] Anonymous. Author Index, volumes 201–220. Linear Algebra and its Applications, 220 (1–3):441–??, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1995:RV

[Ano95c] Anonymous. Referees, volumes 201–220. Linear Algebra and its Applications, 220 (1–3):435–??, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1996:CCM

[Ano96a] Anonymous. Call for challenges in matrix theory. Linear Algebra and its Applications, 233(1–3):1–??, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1996:ENC

[Ano96b] Anonymous. Editor’s note:

“On a conjecture by Goldberg and Newman” [Linear Algebra Appl. 215 (1995), 275– 277, MR 96i:15027] by Dursun Tasci. Linear Algebra and its Applications, 240(1–3):237, June ??, 1996. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1997:VAI

[Ano97] Anonymous. 20-volume author index. Linear Algebra and its Applications, 260(1–3): 329–??, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1998:A

[Ano98a] Anonymous. Acknowledgements. Linear Algebra and its Applications, 280(1):64, August 30, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:CVO

[Ano98b] Anonymous. [curriculum vitae of] Olga Taussky Todd. Linear Algebra and its Applications, 280(1):62–63, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:HAX

[Ano98c]  Anonymous.  Householder

Award X. Linear Algebra and its Applications, 283(1– 3):301–??, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1998:OTT

[Ano98d] Anonymous. Olga Taussky Todd problems and solutions. Linear Algebra and its Applications, 280(1):84–85, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:PAO

[Ano98e]  Anonymous.  Publications

about Olga Taussky Todd. Linear Algebra and its Applications, 280(1):59–61, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:POT

[Ano98f] Anonymous. Publications of Olga Taussky Todd. Linear Algebra and its Applications, 280(1):67–83, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:RV

[Ano98g] Anonymous. Referees for volumes 261–280. Linear Algebra and its Applications, 280 (2–3):337–340, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1998:SOT

[Ano98h] Anonymous. Students of Olga Taussky Todd. Linear Algebra and its Applications, 280(1): 65–66, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1999:Ia

[Ano99a] Anonymous. Index. Linear Algebra and its Applications, 297(1–3):203, August 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/30/
19/104/17/30/;   http://
www.elsevier.nl/gej-ng/
10/30/19/104/17/30/article.
pdf.

Anonymous:1999:Ib

[Ano99b] Anonymous. Index. Linear Algebra and its Applications, 298(1–3):209, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/30/
19/105/17/30/;   http://
www.elsevier.nl/gej-ng/
10/30/19/105/17/30/article.
pdf.

Anonymous:1999:Ic

[Ano99c] Anonymous. Index. Linear Algebra and its Applications, 299(1–3):201, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/32/article.
pdf.

Anonymous:1999:Id

[Ano99d] Anonymous. Index. Linear Algebra and its Applications, 300(1–3):9–20, October 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/119/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/119/25/26/article.
pdf.

Anonymous:1999:Ie

[Ano99e] Anonymous. Index. Linear Algebra and its Applications, 300(1–3):33–91, October 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/119/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/119/25/27/article.
pdf.

Anonymous:1999:If

[Ano99f] Anonymous. Index. Linear Algebra and its Applications, 301(1–3):203, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/39/article.
pdf.

Anonymous:1999:Ig

[Ano99g] Anonymous. Index. Linear Algebra and its Applications, 302-303(1–3):619–620, December 1, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/117/25/
                                        58/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/58/article.
                                        pdf.

Anstreicher:1991:MSP

[Ans91] Kurt M. Anstreicher. On monotonicity in the scaled potential algorithm for linear programming. Linear Algebra and its Applications, 152: 223–232, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Antoulas:1994:NAM

[Ant94] A. C. Antoulas. A new approach to modeling for control. Linear Algebra and its Applications, 203/204:45–65, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:1998:ALO

[Ant98] A. C. Antoulas. Approximation of linear operators in the 2-norm. Linear Algebra and its Applications, 278(1–3):309–316, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        278/1-3/6087.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?

year=1998&volume=278&issue=
1-3&aid=6087.

Ando:1991:INS

[AO91] T. Ando and K. Okubo. Induced norms of the Schur multiplier operator. Linear Algebra and its Applications, 147: 181–199, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andersson:1994:CMG

[AP94a] Steen A. Andersson and Michael D. Perlman. A characterization of matrix groups that act transitively on the cone of positive definite matrices. Linear Algebra and its Applications, 199:151–170, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arikati:1994:DSM

[AP94b] Srinivasa R. Arikati and Uri N. Peled. Degree sequences and majorization. Linear Algebra and its Applications, 199:179–211, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amodio:1996:PSA

[AP96a] P. Amodio and M. Paprzycki. Parallel solution of almost block diagonal systems on a hypercube. Linear Algebra and its Applications, 241/243(1–3):85–103, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500588. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Anda:1996:SSF

[AP96b] Andrew A. Anda and Haesun Park. Self-scaling fast rotations for stiff and equality-constrained linear least squares problems. Linear Algebra and its Applications, 234(1–3):137–161, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400092.

AparecidodeAndrade:1999:CDB

[AP99a] Antonio Aparecido de Andrade and Reginaldo Palazzo, Jr. Construction and decoding of BCH Codes over finite commutative rings. Linear Algebra and its Applications, 286(1–3):69–85, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6274.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6274.

Arikati:1999:RGD

[AP99b] Srinivasa R. Arikati and Uri N. Peled. The realization graph of a degree sequence with majorization gap 1 is Hamiltonian. Linear Algebra and its Applications, 290(1–3):213–235, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6334.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6334.

Appleby:1998:SCN

[App98] Glenn D. Appleby. Similarity classes for nilpotent operators over Dedekind domains. Linear Algebra and its Applications, 274(1–3):37– 59, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700244.

Appleby:1999:SAM

[App99] Glenn D. Appleby. A simple approach to matrix realizations for Littlewood– Richardson sequences. Linear Algebra and its Applications, 291(1–3):1–14, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6303.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6303.

Abeledo:1995:CLP

[AR95] Hernán G. Abeledo and Uriel G. Rothblum. Courtship and linear programming. Linear Algebra and its Applications, 216(1–3):111–124, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300122.

Allen:1999:MEA

[AR99] Brian M. Allen and Joachim Rosenthal. A matrix Euclidean algorithm induced by state space realization. Linear Algebra and its Applications, 288(1–3):105–121, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

tree/store/laa/sub/1999/
288/1-3/6296.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6296.

Artzrouni:1991:GIP

[Art91] Marc Artzrouni. On the growth of infinite products of slowly varying primitive matrices. Linear Algebra and its Applications, 145:33–57, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amrein:1994:PPH

[AS94] W. O. Amrein and Kalyan B.

Sinha. On pairs of projections in a Hilbert space. Linear Algebra and its Applications, 208/209:425–435, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anderson:1997:CDB

[AS97] T. W. Anderson and M. A.

Stephens. The continuous and discrete Brownian bridges: Representations and applications. Linear Algebra and its Applications, 264 (1–3):145–171, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700015.pdf;
http://www.elsevier.com/

cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700015.

Ashikhmin:1998:DI

[AS98] Alexey Ashikhmin and Juriaan Simonis. On the Delsarte inequalities. Linear Algebra and its Applications, 269(1):197–217, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700065.

Arnold:1999:MAT

[AS99] Bernhard F. Arnold and Peter Stahlecker. Minimax adjustment technique and fuzzy information. Linear Algebra and its Applications, 289 (1–3):25–39, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6005.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6005.

At:1991:DTM

[At91] Pham Van At. Diagonal transformation methods for computing the maximal eigenvalue and eigenvector of a nonnegative irreducible matrix. Linear Algebra and its Applications, 148:93–123, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Atalla:1992:CWM

[Ata92] Robert E. Atalla. Characterization of weakly monotone stochastic matrices. Linear Algebra and its Applications, 168:65–??, April 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Atkinson:1993:CUG

[Atk93] E. Neely Atkinson. Computing A^{T}A − B^{T}B = L^{T}DL using generalized hyperbolic transformations. Linear Algebra and its Applications, 194: 135–147, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Angel:1995:GSF

[ATST95] J. Angel, C. Trimble, B. Shook, and A. Terras. Graph spectra for finite upper half planes over rings. Linear Algebra and its Applications, 226/228 (1–3):423–457, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1995&volume=226-228&issue=
1-3&aid=9500173.

Aslaksen:1995:FRG

[ATZ95] Helmer Aslaksen, Eng-Chye Tan, and Chen-Bo Zhu. Free resolutions of generic symmetric matrices. Linear Algebra and its Applications, 230(1–3):21–34, November 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300355.

Ahmed:1999:IBE

[AU99] S. E. Ahmed and Bashir Ullah. Improved biased estimation in an ANOVA model. Linear Algebra and its Applications, 289(1–3):3–24, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6003.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6003.

Aujla:1993:MCF

[Auj93] Jaspal Singh Aujla. Matrix convexity of functions of two variables. Linear Algebra and its Applications, 194:149–160, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aujla:1997:SOE

[Auj97] Jaspal Singh Aujla. Some operator equalities involving connections and means. Linear Algebra and its Applications, 259(1–3):223–228, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600285.

Aujla:1999:FPT

[Auj99]  Jaspal Singh Aujla.  A

fixed point theorem and a norm inequality for operator means. Linear Algebra and its Applications, 290(1– 3):109–118, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6324.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6324.

Aujla:1996:SCM

[AV96] Jaspal Singh Aujla and H. L.

Vasudeva. Some convex and monotone matrix functions. Linear Algebra and its Applications, 248(1–3):47– 60, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500114.

Arlinghaus:1995:PIM

[AVH95] F. A. Arlinghaus, L. N. Vaserstein, and You Hong. Products of involutory matrices over rings. Linear Algebra and its Applications, 229(1–3):37– 47, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9300346.

Ayachour:1999:ESI

[Aya99] E. H. Ayachour. Expanded systems and the ILU preconditioner for solving non-Hermitian linear systems. Linear Algebra and its Applications, 293(1–3):243–256, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6398.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

year=1999&volume=293&issue=
1-3&aid=6398.

Au-Yeung:1991:PMW

[AYC91] Yik Hoi Au-Yeung and Che-Man Cheng. Permutation matrices whose convex combinations are orthostochastic. Linear Algebra and its Applications, 150:243–253, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Aupetit:1990:CNM

[AZ90] Bernard Aupetit and Jaroslav Zemánek. A characterization of normal matrices by their exponentials. Linear Algebra and its Applications, 132:119– 121, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Azenhas:1995:OLR

[Aze95]  Olga Azenhas.  Opposite

Littlewood–Richardson sequences and their matrix realizations. Linear Algebra and its Applications, 225(1):91– 116, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300324.

BaezaV:1990:NBA

[Bae90] Rodolfo Baeza V. On nuclear Bernstein algebras. Linear Algebra and its Applications, 142:19–22, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ben-Artzi:1996:MPM

[BAG96] A. Ben-Artzi and I. Gohberg. Monotone power method in indefinite metric and inertia theorem for matrices. Linear Algebra and its Applications, 241/243(1–3):153–166, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500603. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Ben-Artzi:1999:SNC

[BAG99] A. Ben-Artzi and I. Gohberg. Singular numbers of contractions in spaces with an indefinite metric and Yamamoto’s theorem. Linear Algebra and its Applications, 290(1–3):31–48, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/

290/1-3/6297.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6297.

Bai:1997:CIM

[Bai97a] Zhong-Zhi Bai. A class of iteration methods based on the Moser formula for nonlinear equations in Markov chains. Linear Algebra and its Applications, 266(1–3):219–241, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9786522.

Bai:1997:PHA

[Bai97b]  Zhong-Zhi Bai.  Parallel

hybrid algebraic multilevel iterative methods. Linear Algebra and its Applications, 267(1–3):281–315, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700024.

Bai:1998:CPD

[Bai98]  Zhong-Zhi Bai.  A class

of parallel decompositiontype relaxation methods for large sparse systems of linear equations. Linear Algebra and its Applications, 282 (1–3):1–24, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6212.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6212.

Baker:1999:REQ

[Bak99] Andrew Baker. Right eigenvalues for quaternionic matrices: A topological approach. Linear Algebra and its Applications, 286(1–3): 303–309, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6291.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6291.

Baldick:1995:RPS

[Bal95] Ross Baldick. Refined proximity and sensitivity results in linearly constrained convex separable integer programming. Linear Algebra and its Applications, 226/228 (1–3):389–407, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500172.

Balser:1999:CFF

[Bal99] Werner Balser. Computation of formal fundamental solutions. Linear Algebra and its Applications, 288(1– 3):293–312, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6336.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6336.

Bapat:1991:ITT

[Bap91a] R. B. Bapat. An interlacing theorem for tridiagonal matrices. Linear Algebra and its Applications, 150:331–340, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Bapat:1991:MSV

[Bap91b] R. B. Bapat. Majorization and singular values. III. Linear Algebra and its Applications, 145:59–70, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1993:SFM

[Bap93] R. B. Bapat. Symmetric function means and permanents. Linear Algebra and its Applications, 182:101–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1994:GIP

[Bap94a] R. B. Bapat. Generalized inverses with proportional minors. Linear Algebra and its Applications, 211:27–33, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Bapat:1994:KTB

[Bap94b] R. B. Bapat. König’s theorem and bimatroids. Linear Algebra and its Applications, 212/213:353–365, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Bapat:1998:MVP

[Bap98a] R. B. Bapat. A max version of the Perron–Frobenius theorem. Linear Algebra and its Applications, 275/276 (1–3):3–18, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6051.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6051. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Bapat:1998:SNR

[Bap98b] R. B. Bapat. Structure of a nonnegative regular matrix and its generalized inverses. Linear Algebra and its Applications, 268(1–3):31– 39, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700062.

Bapat:1999:LEM

[Bap99] R. B. Bapat. Linear estimation in models based on a graph. Linear Algebra and its Applications, 302-303(1–3):223–230, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/117/25/38/article.
                                        pdf.

Ben-Artzi:1991:TIS

[BA– DR91] Asher Ben-Artzi, Dragomir Ž. – Djokovič, and Leiba Rodman. Totally isotropic subspaces for pairs of Hermitian forms and applications to Riccati equations. Linear Algebra and its Applications, 159:121–128, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrlund:1994:PBG

[Bar94] Anders Barrlund. Perturbation bounds for the generalized QR factorization. Linear Algebra and its Applications, 207:251–271, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barnett:1996:LAO

[Bar96] Stephen Barnett. Leverrier’s algorithm for orthogonal polynomial bases. Linear Algebra and its Applications, 236(1–3): 245–263, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400158.

Barioli:1998:CPM

[Bar98]  Francesco Barioli.  Com-

pletely positive matrices with a book-graph. Linear Algebra and its Applications, 277 (1–3):11–31, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6060.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6060.

Batterson:1994:CFS

[Bat94] Steve Batterson. Convergence of the Francis shifted QR algorithm on normal matrices. Linear Algebra and its Applications, 207:181–195, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bauer:1992:DCF

[Bau92] Max Bauer. Dilatations and continued fractions. Linear Algebra and its Applications, 174:183–??, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bauer:1998:OTT

[Bau98] Friedrich L. Bauer. Olga Taussky-Todd 30.8.1906–7.10.1995. Linear Algebra and its Applications, 280(1):9–12, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Beasley:1992:RLA

[BB92] LeRoy B. Beasley and Debra A. S. Beasley. Review of Linear Algebra and Its Application, Volume 1: A First Course and Linear Algebra and Its Application, Volume 2: More Advanced, by D. H. Griffel. Linear Algebra and its Applications, 165:269–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1993:GHW

[BB93] Rajendra Bhatia and Tirthankar Bhattacharyya. A generalization of the Hoffman-Wielandt thorem. Linear Algebra and its Applications, 179:11–??, January 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bevilacqua:1996:ASL

[BB96] Roberto Bevilacqua and Enrico Bozzo. On algebras of symmetric Loewner matrices. Linear Algebra and its Applications, 248(1–3):241–251, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=248&issue=1-
                                        3&aid=9500202.

Bevilacqua:1995:ATP

[BBB95] Roberto Bevilacqua, Nazzareno Bonanni, and Enrico Bozzo. On algebras of Toeplitz plus Hankel matrices. Linear Algebra and its Applications, 223/224(1–3):99– 118, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400299. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Bevis:1997:RGA

[BBD^{+}97] Jean H. Bevis, Kevin K.

Blount, George J. Davis, Gayla S. Domke, and Valerie A. Miller. The rank of a graph after vertex addition. Linear Algebra and its Applications, 265 (1–3):55–69, November 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600513.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600513.

Barany:1995:RCA

[BBLP95] I. Barany, H. Bunting, D. G.

Larman, and J. Pach. Rich cells in an arrangement of hyperplanes. Linear Algebra and its Applications, 226/228 (1–3):567–575, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500244.

Bakonyi:1991:IPC

[BC91a] Mihai Bakonyi and Tiberiu Constantinescu. Inheritance principles for chordal graphs. Linear Algebra and its Applications, 148:125–143, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barraa:1991:SEI

[BC91b] Mohamed Barraa and Bernard Charles. Sous-espaces invariants d’un opérateur nilpotent sur un espace de Banach. (French) [invariant subspaces of a nilpotent operator on a Banach space]. Linear Algebra and its Applications, 153: 177–182, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1991:SML

[BC91c] R. A. Brualdi and J. Csima. Small matrices of large dimension. Linear Algebra and its Applications, 150:227–241, 1991. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Bapat:1992:EFS

[BC92] R. B. Bapat and G. Constantine. An enumerating function for spanning forests with color restrictions. Linear Algebra and its Applications, 173: 231–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1993:SGP

[BC93] LeRoy B. Beasley and Larry J.

Cummings. The structure of generalized permanent semigroups. Linear Algebra and its Applications, 184:235–247, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benabdallah:1995:OIS

[BC95a] Khalid Benabdallah and Bernard Charles. Orbits of invariant subspaces of algebraic linear operators. Linear Algebra and its Applications, 225(1):13–22, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300302.

Bouali:1995:ENO

[BC95b] S. Bouali and J. Charles. Extension de la notion d’opérateur D-symétrique. II. (French) [extension of the notion of Dsymmetric operator. II]. Linear Algebra and its Applications, 225(1):175–185, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9400003.

Bilir:1998:GIT

[BC98a] Bülent Bilir and Carmen Chicone. A generalization of the inertia theorem for quadratic matrix polynomials. Linear Algebra and its Applications, 280(2–3): 229–240, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6156.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6156.

Borobia:1998:MSG

[BC98b] Alberto Borobia and Rafael Cantó. Matrix scaling: A geometric proof of Sinkhorn’s theorem. Linear Algebra and its Applications, 268(1–3):1– 8, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700010.

Bernasconi:1999:HFC

[BCCR99] A. Bernasconi, B. Codenotti, V. Crespi, and G. Resta. How fast can one compute the permanent of circulant matrices? Linear Algebra and its Applications, 292(1–3):15–37, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6361.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6361.

Butkovic:1991:RMA

[BCG91] P. Butkovič and R. A.

Cuninghame-Green. On the regularity of matrices in min algebra. Linear Algebra and its Applications, 145:127–139, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:1997:GCR

[BCHS97] Rafael Bru, Carmen Coll, Vicente Hernández, and Elena Sánchez. Geometrical conditions for the reachability and realizability of positive periodic discrete systems. Linear Algebra and its Applications, 256(1–3):109– 124, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781113.

Bru:1996:PEB

[BCJ96] Rafael Bru, Joan-Josep Climent, and Charles R. Johnson. On the Perron eigenvalue of a block cocyclic pair of nonnegative matrices. Linear Algebra and its Applications, 236(1–3):231– 243, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400157.

Brown:1993:OIF

[BCP93] Jack B. Brown, Phillip J. Chase, and Arthur O. Pittenger. Order independence and factor convergence in iterative scaling. Linear Algebra and its Applications, 190: 1–38, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1994:RM

[BCS94] Richard A. Brualdi, Keith L.

Chavey, and Bryan L. Shader. Rectangular L-matrices. Linear Algebra and its Applications, 196:37–61, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:1993:PPF

[BCT93] Rafael Bru, Rafael Canto, and Bit-Shun Tam. Predecessor property, full combinatorial column rank, and the height characteristic of an M-matrix. Linear Algebra and its Applications, 183:1–??, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:1994:AMP

[BCU94] Rafael Brú, Juana Cerdán, and Ana M. Urbano. An algorithm for the multiinput pole assignment problem. Linear Algebra and its Applications, 199:427–444, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bouchet:1991:RMG

[BD91] A. Bouchet and A. Duchamp.

Representability of △-matroids over GF(2). Linear Algebra and its Applications, 146:67– 78, 1991. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:1993:SDB

[BD93] Zhaojun Bai and James W. Demmel. On swapping diagonal blocks in real Schur form. Linear Algebra and its Applications, 186:73–95, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1995:CSI

[BD95] Rajendra Bhatia and Chandler Davis. A Cauchy– Schwarz inequality for operators with applications. Linear Algebra and its Applications, 223/224(1–3):119– 129, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400344. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Bebiano:1997:DRM

[Bd97] Natália Bebiano and J. da Providência. On the dual of a result of Miranda and Thompson. Linear Algebra and its Applications, 262(1–3):119– 129, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600477.

Balasubramanian:1992:DRD

[BDB92] K. Balasubramanian, A. Dey, and P. Bhimasankaram. Diagonally range dominant matrices. Linear Algebra and its Applications, 176:45–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:1992:RMM

[BdV92] Rafael Bru, Graciano de Oliveira, and Jose Vitoria. Report: Meeting of Matrix Analysis and Applications, 3–5 October 1989, Universidad Politecnica de Valencia, Valencia, Spain. Linear Algebra and its Applications, 170:147–??, June 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bryant:1998:GOL

[BDV98] Darryn Bryant, Phil Diamond, and Igor Vladimirov. Generalized optimal lattice covering of finite-dimensional Euclidean space. Linear Algebra and its Applications, 282(1–3):311–324, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6200.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=282&issue=
                                        1-3&aid=6200.

Boyd:1993:MCM

[BE93] Stephen Boyd and Laurent El Ghaoui. Method of centers for minimizing generalized eigenvalues. Linear Algebra and its Applications, 188/189:63– 111, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beaumont:1998:SIL

[Bea98] O. Beaumont. Solving interval linear systems with linear programming techniques. Linear Algebra and its Applications, 281(1–3):293–309, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        281/1-3/6165.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6165.

Beezer:1990:DP

[Bee90] Robert A. Beezer. A disrespectful polynomial. Linear Algebra and its Applications, 128:139–146, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Behrends:1994:MCF

[Beh94]  Ehrhard Behrends.  On

matrix-closed families of subspaces of a Banach space. Linear Algebra and its Applications, 196:253–272, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1990:BVR

[BEK90] R. Bhatia, L. Elsner, and G. Krause. Bounds for the variation of the roots of a polynomial and the eigenvalues of a matrix. Linear Algebra and its Applications, 142: 195–209, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:1991:MIC

[Bel91] F. K. Bell. On the maximal index of connected graphs. Linear Algebra and its Applications, 144:135–151, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:1999:CLG

[Bel99] F. K. Bell. Characterizing line graphs by star complements. Linear Algebra and its Applications, 296(1–3):15–25, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/18/
abstract.html;   http://
www.elsevier.nl/gej-ng/

                                        10/30/19/47/17/18/article.
                                        pdf.

Beslin:1992:CM

[Bes92] Scott J. Beslin. Cofactor matrices. Linear Algebra and its Applications, 165:45–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bach:1992:CQO

[BF92a] Edward Bach and Takayuki Furuta. Counterexample to a question on the operator equation T (H^{1/n}T)^{n} = K. Linear Algebra and its Applications, 177:157–??, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunch:1992:CFA

[BF92b] James R. Bunch and Ricardo D. Fierro. A constantfalse-alarm-rate algorithm. Linear Algebra and its Applications, 172:231–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1995:CIM

[BF95] Richard A. Brualdi and Amélia Fonseca. Colorability of induced matroids. Linear Algebra and its Applications, 223/224(1–3):141– 153, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400167. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Benner:1997:IRS

[BF97] Peter Benner and Heike Faßbender. An implicitly restarted symplectic Lanczos method for the Hamiltonian eigenvalue problem. Linear Algebra and its Applications, 263(1–3):75–111, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600524.

Benner:1998:SEP

[BF98] Peter Benner and Heike Faßbender. The symplectic eigenvalue problem, the butterfly form, the SR algorithm, and the Lanczos method. Linear Algebra and its Applications, 275/276 (1–3):19–47, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6047.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-

                                        276&issue=1-3&aid=6047. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Benner:1998:TCB

[BFW98] Peter Benner, Heike Faßbender, and David S. Watkins. Two connections between the SR and HR eigenvalue algorithms. Linear Algebra and its Applications, 272(1–3):17– 32, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=272&issue=1-
                                        3&aid=9700279.

Benner:1999:ASB

[BFW99] Peter Benner, Heike Faßbender, and David S. Watkins. SR and SZ algorithms for the symplectic (butterfly) eigenproblem. Linear Algebra and its Applications, 287(1– 3):41–76, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6209.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6209.

Beitia:1994:LBS

[BG94] Ma. Asunción Beitia and Juan-M. Gracia. Local behavior of Sylvester matrix equations related to block similarity. Linear Algebra and its Applications, 199:253–279, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beitia:1996:SME

[BG96] Ma. Asunción Beitia and Juan-Miguel Gracia. Sylvester matrix equation for matrix pencils. Linear Algebra and its Applications, 232(1–3):155– 197, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:1998:FFF

[BG98a] Dario Andrea Bini and Luca Gemignani. Fast fraction-free triangularization of Bézoutians with applications to sub-resultant chain computation. Linear Algebra and its Applications, 284 (1–3):19–39, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6194.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6194.

Bottcher:1998:CNL

[BG98b] Albrecht Böttcher and Sergei M. Grudsky. On the condition numbers of large semi-definite Toeplitz matrices. Linear Algebra and its Applications, 279(1–3):285–301, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6113.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6113.

Bunse-Gerstner:1999:FDR

[BGBMN99] Angelika Bunse-Gerstner, Ralph Byers, Volker Mehrmann, and Nancy K. Nichols. Feedback design for regularizing descriptor systems. Linear Algebra and its Applications, 299(1–3):119–151, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/25/article.
pdf.

Bunse-Gerstner:1991:SPP

[BGE91] Angelika Bunse-Gerstner and Ludwig Elsner. Schur parameter pencils for the solution of the unitary eigenproblem. Linear Algebra and its Applications, 154/156:741– 778, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:1992:RHC

[BGH92a] Abraham Berman, Moshe Goldberg, and Daniel Hershkowitz. Report: Haifa 1990 Conference on Matrix Theory. Linear Algebra and its Applications, 167:205–??, April 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:1992:SHC

[BGH92b] Abraham Berman, Moshe Goldberg, and Daniel Hershkowitz, editors. Sixth Haifa Conference on Matrix Theory, Held on 11–14 June 1990 at the Technion, Israel Institute of Technology, Haifa, Israel, volume 167 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blake:1997:SIP

[BGM97] Ian F. Blake, Shuhong Gao, and Ronald C. Mullin. Specific irreducible polynomials with linearly independent roots over finite fields. Linear Algebra and its Applications, 253(1–3):227–249, March 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500744.

Barnabei:1998:SAA

[BGM98] M. Barnabei, C. Guerrini, and L. B. Montefusco. Some algebraic aspects of signal processing. Linear Algebra and its Applications, 284(1– 3):3–17, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6248.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6248.

Ball:1990:CMM

[BGR90] Joseph A. Ball, Israel Gohberg, and Leiba Rodman. Common minimal multiples and divisors for rational matrix functions. Linear Algebra and its Applications, 137/138: 621–662, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1992:GTT

[BGS92] Richard A. Brualdi and Hwang Geum-Sug. Generalized transitive tournaments and doubly stochastic matrices. Linear Algebra and its Applications, 172:151–??, July 15, 1992. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunse-Gerstner:1999:CGT

[BGS99] Angelika Bunse-Gerstner and Ronald Stöver. On a conjugate gradient-type method for solving complex symmetric linear systems. Linear Algebra and its Applications, 287(1–3):105–123, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6210.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6210.

Baksalary:1990:FAV

[BH90] Jerzy K. Baksalary and Jan Hauke. A further algebraic version of Cochran’s theorem and matrix partial orderings. Linear Algebra and its Applications, 127:157–169, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bevis:1991:IF

[BH91] Jean H. Bevis and Frank J.

Hall. Integer LU-factorizations. Linear Algebra and its Applications, 150:267–285, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Bevis:1992:NRC

[BH92a] Jean H. Bevis and Frank J. Hall. Nested range conditions for LU factorizations of integer matrices. Linear Algebra and its Applications, 172: 97–??, July 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1992:MMJ

[BH92b] Richard A. Brualdi and Chan Jung Hyung. Maximum and minimum jump number of posets from matrices. Linear Algebra and its Applications, 172:261–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1994:MRS

[BH94a] R. B. Bapat and Robert E. Hartwig. A master of the row space and the column space: the mathematical work of Sujit Kumar Mitra. Linear Algebra and its Applications, 211: 5–14, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Barnes:1994:BSN

[BH94b] E. R. Barnes and A. J. Hoffman. Bounds for the spectrum of normal matrices. Linear Algebra and its Applications, 201:79–90, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1995:MAW

[BH95] Richard A. Brualdi and Suk Geun Hwang. Matrices all of whose principal submatrices of some order have a nonzero permanent. Linear Algebra and its Applications, 226/228 (1–3):409–421, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500174.

Barnett:1996:SRO

[BH96] S. Barnett and R. E. Hartwig.

Special rank-one perturbations. Linear Algebra and its Applications, 235(1–3):171– 190, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400135.

Breen:1997:PSF

[BH97] Michael Breen and David Hume. Properties of a standard form for a Boolean matrix. Linear Algebra and its Applications, 254(1–3):49– 65, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9600368. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Bhat:1991:GCD

[Bha91] B. V. Rajarama Bhat. On greatest common divisor matrices and their applications. Linear Algebra and its Applications, 158:77–97, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1994:FSO

[Bha94a] Rajendra Bhatia. First and second order perturbation bounds for the operator absolute value. Linear Algebra and its Applications, 208/209: 367–376, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1994:MFT

[Bha94b] Rajendra Bhatia. Matrix factorizations and their perturbations. Linear Algebra and its Applications, 197/198:245– 276, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Bhatia:1995:PBO

[Bha95] Rajendra Bhatia. Perturbation bounds for the operator absolute value. Linear Algebra and its Applications, 226/228 (1–3):639–645, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500201.

Bhatia:1997:EOI

[Bha97a] R. Bhatia. Erratum: “An operator inequality and matrix normality” [Linear Algebra Appl. 240 (1996), 105– 110, MR 97b:47011] by C. R. Johnson and F. Zhang. Linear Algebra and its Applications, 260(1–3):323, July 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1997:NLE

[Bha97b] Rajendra Bhatia. A note on the Lyapunov equation. Linear Algebra and its Applications, 259(1–3):71–76, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600242.

Bhattacharyya:1999:FDC

[Bha99] Tirthankar Bhattacharyya. On finite-dimensional commutative non-Hermitian fusion algebras. Linear Algebra and its Applications, 287(1–3):87–103, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6252.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6252.

Bevis:1990:CCD

[BHH90] Jean H. Bevis, Frank J. Hall, and Robert E. Hartwig. Conpseudosimilarity and consemisimilarity over a division ring. Linear Algebra and its Applications, 136:181–188, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:1991:NAB

[BHK91] Rajendra Bhatia, Roger A. Horn, and Fuad Kittaneh. Normal approximants of binormal operators. Linear Algebra and its Applications, 147:169–179, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:1999:NCC

[BHL99] Wayne Barrett, H. Tracy Hall, and Raphael Loewy. A nonpolyhedral cone of class function inequalities for positive semidefinite matrices. Linear Algebra and its Applications, 302-303(1– 3):535–553, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
54/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/54/article.
pdf.

Byers:1998:WNN

[BHM98] Ralph Byers, Chunyang He, and Volker Mehrmann. Where is the nearest non-regular pencil? Linear Algebra and its Applications, 285(1–3): 81–105, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6239.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6239.

Brualdi:1997:VMP

[BHP97] Richard A. Brualdi, Suk-Geun Hwang, and Sung-Soo Pyo. Vector majorization via positive definite matrices. Linear Algebra and its Applications, 257(1–3):105– 120, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=257&issue=1-
                                        3&aid=9600042.

Barrett:1991:PFC

[BHR91] Wayne Barrett, Daniel. Hershkowitz, and Donald Robinson, editors. Proceedings of the First Conference of the International Linear Algebra Society: Brigham Young University, Provo, Utah, 12–15 August 1989, volume 150 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1991. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ben-Israel:1992:VAM

[BI92] Adi Ben-Israel. A volume associated with m × n matrices. Linear Algebra and its Applications, 167:87–??, April 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bakonyi:1994:ICP

[BJ94] Mihály Bakonyi and Charles R. Johnson. Invertible completions of partial operator matrices: The nonsymmetric case. Linear Algebra and its Applications, 208/209:327– 341, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bjorck:1994:NGS

[Bjö94]  Å. Björck.  Numerics of

Gram–Schmidt orthogonalization. Linear Algebra and its Applications, 197/198:297– 316, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Bapat:1997:WMP

[BJP97] R. B. Bapat, S. K. Jain, and S. Pati. Weighted Moore– Penrose inverse of a Boolean matrix. Linear Algebra and its Applications, 255(1–3):267– 279, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600777.

Bapat:1997:NIM

[BJS97] R. B. Bapat, S. K. Jain, and L. E. Snyder. Nonnegative idempotent matrices and the minus partial order. Linear Algebra and its Applications, 261(1–3):143– 154, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600364.

Barrett:1993:RPD

[BJT93] Wayne Barrett, Charles R. Johnson, and Pablo Tarazaga. The real positive definite completion problem for a simple cycle. Linear Algebra and its Applications, 192:3–31, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Burrage:1998:BTP

[BJW98] K. Burrage, Z. Jackiewicz, and B. Welfert. Block-Toeplitz preconditioning for static and dynamic linear systems. Linear Algebra and its Applications, 279(1–3):51–74, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6105.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6105.

Ball:1990:MPS

[BK90] Joseph A. Ball and Jeongook Kang. Matrix polynomial solutions of tangential Lagrange–Sylvester interpolation conditions of low McMillan degree. Linear Algebra and its Applications, 137/138: 699–746, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brewer:1991:PSP

[BK91a] James Brewer and Lee Klingler. On the pole-shifting problem for noncommutative rings. Linear Algebra and its Applications, 157:147–159, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brewer:1991:RIV

[BK91b] James Brewer and Lee Klingler. The ring of integervalued polynomials of a semilocal principal-ideal domain. Linear Algebra and its Applications, 157:141–145, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhaya:1993:DTD

[BK93] Amit Bhaya and Eugenius Kaszkurewicz. On discretetime diagonal and D-stability. Linear Algebra and its Applications, 187:87–104, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ball:1994:SMD

[BK94a] Joseph A. Ball and Jeongook Kim. Stability and McMillan degree for rational matrix interpolants. Linear Algebra and its Applications, 196:207– 232, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1994:LRO

[BK94b] R. B. Bapat and Subhash C. Kochar. On likelihood-ratio ordering of order statistics. Linear Algebra and its Applications, 199:281–291, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ball:1996:BIP

[BK96a] Joseph A. Ball and Jeongook Kim. Bitangential interpolation problems for symmetric rational matrix functions. Linear Algebra and its Applications, 241/243(1–3):113– 152, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500562. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Bart:1996:CBM

[BK96b] H. Bart and L. G. Kroon.

Companion based matrix functions: Description and minimal factorization. Linear Algebra and its Applications, 248(1–3):1–46, November 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9400045.

Bart:1996:FJS

[BK96c] Harm Bart and Leo G. Kroon.

Factorization and job scheduling: A connection via companion based matrix functions. Linear Algebra and its Applications, 248(1–3):111–136, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500163.

Basilio:1997:ACM

[BK97a] J. C. Basilio and B. Kouvaritakis. An algorithm for coprime matrix fraction description using Sylvester matrices. Linear Algebra and its Applications, 266(1–3):107–125, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=266&issue=1-
                                        3&aid=9600636.

Beasley:1997:EPM

[BK97b] LeRoy B. Beasley and Steve Kirkland. On the exponent of a primitive matrix containing a primitive submatrix. Linear Algebra and its Applications, 261(1–3):195– 205, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=261&issue=1-
                                        3&aid=9600401.

Berg:1997:UZS

[BK97c] Jordan M. Berg and Harry G. Kwatny. Unfolding the zero structure of a linear control system. Linear Algebra and its Applications, 258(1–3):19– 39, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=258&issue=1-
                                        3&aid=9600156.

Boman:1999:CPS

[BK99] E. Boman and I. Koltracht. Computing preconditioners via subspace projection. Linear Algebra and its Applications, 302-303(1–3):347–353, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
45/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/45/article.
pdf.

Bebiano:1994:VMO

[BKdP94] Natália Bebiano, Alexander Kovačec, and João da Providência. The validity of the Marcus-de Oliveira conjecture for essentially Hermitian matrices. Linear Algebra and its Applications, 197/198: 411–427, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Ball:1994:CRR

[BKGK94] J. A. Ball, M. A. Kaashoek, G. Groenewald, and J. Kim. Column reduced rational matrix functions with given nullpole data in the complex plane. Linear Algebra and its Applications, 203/204:67– 110, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1997:PLM

[BKL97] LeRoy B. Beasley, An-Hyun Kim, and Woo Young Lee. On a positive linear map preserving absolute values. Linear Algebra and its Applications, 260(1–3):311–318, July 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=260&issue=1-
                                        3&aid=9600312.

Bhaya:1991:ABI

[BKM91] Amit Bhaya, Eugenius Kaszkurewicz, and Francisco Mota. Asynchronous block-iterative methods for almost linear equations. Linear Algebra and its Applications, 154/156:487– 508, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bussemaker:1997:DMC

[BKMS97] Frans Bussemaker, Irving Kaplansky, Brendan McKay, and Jacob Seidel. Determinants of matrices of the conference type. Linear Algebra and its Applications, 261(1–3):275– 292, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1997&volume=261&issue=1-
3&aid=9600412.

Bunger:1997:PSU

[BKN97] Florian Bünger, Frieder Knüppel, and Klaus Nielsen. Products of symmetries in unitary groups. Linear Algebra and its Applications, 260(1– 3):9–42, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600237.

Boros:1999:FPB

[BKO99] Tibor Boros, Thomas Kailath, and Vadim Olshevsky. A fast parallel Björck–Pereyra-type algorithm for solving Cauchy linear equations. Linear Algebra and its Applications, 302-303(1–3):265–293, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/41/article.
pdf.

Bland:1993:NCB

[BKS93] R. G. Bland, C. W. Ko, and B. Sturmfels. A nonextremal Camion basis. Linear Algebra and its Applications, 187: 195–199, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1995:RC

[BKS95] LeRoy B. Beasley, Stephen J. Kirkland, and Bryan L. Shader. Rank comparisons. Linear Algebra and its Applications, 221(1–3):171–188, May ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300242.

Bart:1998:QFT

[BKZ98] Harm Bart, Leo Kroon, and Rob Zuidwijk. Quasicomplete factorization and the two machine flow shop problem. Linear Algebra and its Applications, 278(1–3):195–219, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        278/1-3/6082.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=278&issue=
                                        1-3&aid=6082.

Beasley:1990:LOM

[BL90] LeRoy B. Beasley and Thomas J. Laffey. Linear operators on matrices: the invariance of rank-k matrices. Linear Algebra and its Applications, 133:

175–184, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1991:PPG

[BL91a] R. B. Bapat and A. K. Lal.

Path-positive graphs. Linear Algebra and its Applications, 149:125–149, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Binding:1991:JRH

[BL91b] Paul Binding and Chi-Kwong Li. Joint ranges of Hermitian matrices and simultaneous diagonalization. Linear Algebra and its Applications, 151: 157–167, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourque:1992:GLM

[BL92] Keith Bourque and Steve Ligh. On GCD and LCM matrices. Linear Algebra and its Applications, 174:65– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baruch:1993:LPS

[BL93a] Ehud Moshe Baruch and Raphael Loewy. Linear preservers on spaces of Hermitian or real symmetric matrices. Linear Algebra and its Applications, 183:89–??, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bohl:1993:PSI

[BL93b] Erich Bohl and Peter Lancaster. Perturbation of spectral inverses applied to a boundary layer phenomenon arising in chemical networks. Linear Algebra and its Applications, 180:35–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1994:IP

[BL94a] R. B. Bapat and A. K. Lal. Inequalities for the qpermanent. Linear Algebra and its Applications, 197/198: 397–409, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Bapat:1994:PPI

[BL94b] R. B. Bapat and A. K. Lal. Path positivity and infinite Coxeter groups. Linear Algebra and its Applications, 196: 19–35, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourque:1995:MAM

[BL95] Keith Bourque and Steve Ligh. Matrices associated with multiplicative functions. Linear Algebra and its Applications, 216(1–3):267–275, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300154.

Bebiano:1993:PDE

[BLD93] Natalia Bebiano, Chi-Kwong Li, and Joao Da Providencia. Product of diagonal elements of matrices. Linear Algebra and its Applications, 178:185– ??, January 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:1995:CBD

[BLJW95] Wayne W. Barrett, Michael E.

Lundquist, Charles R. Johnson, and Hugo J. Woerdeman. Completing a block diagonal matrix with a partially prescribed inverse. Linear Algebra and its Applications, 223/224(1–3):73– 87, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400108. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Boley:1992:LAH

[BLL92] Daniel L. Boley, Tong J. Lee, and Franklin T. Luk. The Lanczos algorithm and Hankel matrix factorization. Linear Algebra and its Applications, 172:109–135, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1994:LTP

[BLL94] LeRoy B. Beasley, Gwang-Yeon Lee, and Sang-Gu Lee. Linear transformations that preserve the assignment. Linear Algebra and its Applications, 212/213:387–395, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Blower:1999:AFP

[Blo99] Gordon Blower. On analytic factorization of positive hermitian matrix functions over the bidisc. Linear Algebra and its Applications, 295(1–3):149–158, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6456.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6456.

Bercovici:1997:CLA

[BLS97] Hari Bercovici, Wing Suet Li, and Thomas Smotzer. Classical linear algebra inequalities for the Jordan models of C_{0} operators. Linear Algebra and its Applications, 251(1–3):341–350, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500609.

Boley:1998:FMD

[BLV98] Daniel L. Boley, Franklin T.

Luk, and David Vandevoorde. A fast method to diagonalize a Hankel matrix. Linear Algebra and its Applications, 284(1–3):41–52, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6220.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6220.

Baksalary:1990:RIC

[BM90] Jerzy K. Baksalary and Thomas Mathew. Rank invariance criterion and its application to the unified theory of least squares. Linear Algebra and its Applications, 127: 393–401, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1991:LSR

[BM91] Jerzy K. Baksalary and Sujit Kumar Mitra. Left-star and right-star partial orderings. Linear Algebra and its Applications, 149:73–89, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brawley:1992:SPF

[BM92] Joel V. Brawley and Gary L. Mullen. Scalar polynomial functions on the nonsingular matrices over a finite field. Linear Algebra and its Applications, 174:1–??, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhimasankaram:1993:OPG

[BM93a] P. Bhimasankaram and Thomas Mathew. On ordering properties of generalized inverses of nonnegative definite matrices. Linear Algebra and its Applications, 183:131–??, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1993:MSR

[BM93b] Richard A. Brualdi and Jennifer J. Q. Massey. More on structure-ranks of matrices. Linear Algebra and its Applications, 183:193–199, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Basor:1994:FHC

[BM94a] Estelle L. Basor and Kent E.

Morrison. The Fisher– Hartwig conjecture and Toeplitz eigenvalues. Linear Algebra and its Applications, 202:129– 142, April 15, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:1994:RDI

[BM94b] N. Bebiano and M. E. Miranda. On a recent determinantal inequality. Linear Algebra and its Applications, 201: 99–102, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beckermann:1994:GDI

[BM94c] B. Beckermann and G. Mühlbach.

A general determinantal identity of Sylvester type and some applications. Linear Algebra and its Applications, 197/198: 93–112, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Burgueno:1995:MPO

[BM95] C. Burgueño and C. Mallol.

Morphismes de Peirce et orthogonalité dans les algèbres de Bernstein. (French) [Peirce morphisms and orthogonality in Bernstein algebras].

Linear Algebra and its Applications, 219(1–3):179–186, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300209.

Baksalary:1996:FRI

[BM96] Jerzy K. Baksalary and Augustyn Markiewicz. Further results on invariance of the eigenvalues of matrix products involving generalized inverses. Linear Algebra and its Applications, 237/238(1–3):115– 121, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400001. Special

issue honoring Calyampudi Radhakrishna Rao.

Borobia:1997:NMS

[BM97] Alberto Borobia and Julio Moro. On nonnegative matrices similar to positive matrices. Linear Algebra and its Applications, 266(1–3):365– 379, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9700362.

Barnabei:1998:RPT

[BM98a] M. Barnabei and L. B. Montefusco. Recursive properties of Toeplitz and Hurwitz matrices. Linear Algebra and its Applications, 274(1–3):367– 388, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700370.

Bini:1998:IBT

[BM98b] Dario Andrea Bini and Beatrice Meini. Inverting block Toeplitz matrices in block Hessenberg form by means of displacement operators: Application to queueing problems. Linear Algebra and its Applications, 272(1–3):1– 16, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700273.

Borobia:1998:BSR

[BM98c] Alberto Borobia and Julio Moro. On the boundary of the set of real spectra of nonnegative matrices. Linear Algebra and its Applications, 278(1–3):287–293, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6084.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6084.

Bapat:1994:SIG

[BMH94] Ravi Bapat, Sujit Kumar Mitra, and Robert Hartwig, editors. Special issue: generalized inverses: Workshop held 11–16 December 1993 at the Indian Statistical Institute, volume 211 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1994. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolla:1998:ESH

[BMTZ98] Marianna Bolla, György Michaletzky, Gábor Tusnády, and Margit Ziermann. Extrema of sums of heterogeneous quadratic forms. Linear Algebra and its Applications, 269(1):331–365, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1998&volume=269&issue=1&
aid=9700230.

Brigham:1995:TCG

[BMV95] R. C. Brigham, F. R. Mc-Morris, and R. P. Vitray. Tolerance competition graphs. Linear Algebra and its Applications, 217(1–3):41– 52, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400059. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Byers:1993:SRG

[BN93] Ralph Byers and N. K.

Nichols. On the stability radius of a generalized statespace system. Linear Algebra and its Applications, 188/189: 113–134, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunger:1999:MDT

[BN99] Florian Bünger and Klaus Nielsen. A matrix-decomposition theorem for GL_{n}(K). Linear Algebra and its Applications, 298(1–3):39–50, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/

                                        20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/20/article.
                                        pdf.

Beckermann:1992:STR

[BNM92] B. Beckermann, A. Neuber, and G. Muhlbach. Shanks’s transformation revisited. Linear Algebra and its Applications, 173:191–??, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bojanczyk:1993:BRU

[BNP93] Adam W. Bojanczyk, James G. Nagy, and Robert J. Plemmons. Block RLS using row Householder reflections. Linear Algebra and its Applications, 188/189:31–61, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:1996:DSM

[BNP96] Alberto Borobia, Zeev Nutov, and Michal Penn. Doubly stochastic matrices and dicycle covers and packings in Eulerian digraphs. Linear Algebra and its Applications, 246(1–3):361–371, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=246&issue=1-
                                        3&aid=9500056.

Baksalary:1990:LOA

[BNS90] Jerzy K. Baksalary, Kenneth Nordström, and George P. H. Styan. Löwner-ordering antitonicity of generalized inverses of Hermitian matrices. Linear Algebra and its Applications, 127:171–182, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burke:1992:SPN

[BO92] James V. Burke and Michael L.

Overton. Stable perturbations of nonsymmetric matrices. Linear Algebra and its Applications, 171:249–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Braker:1993:PAM

[BO93] J. G. Braker and G. J. Olsder. The power algorithm in max algebra. Linear Algebra and its Applications, 182: 67–??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boasso:1997:JSN

[Boa97] Enrico Boasso. Joint spectra and nilpotent Lie algebras of linear transformations. Linear Algebra and its Applications, 263(1–3):49–62, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600481.

Bar-On:1995:FSP

[BOC95] Ilan Bar-On and Bruno Codenotti. A fast and stable parallel QR algorithm for symmetric tridiagonal matrices. Linear Algebra and its Applications, 220(1–3):63– 95, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9300360. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Bognar:1991:SRF

[Bog91] János Bognár. A spectralradius formula for symmetric operators in finite-dimensional Kreĭn spaces. Linear Algebra and its Applications, 147: 3–16, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bar-on:1992:GPD

[BoG92] J. R. Bar-on and C. W. Gray. A generalized polar decomposition. Linear Algebra and its Applications, 170:75–??, June 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boldt:1995:MDC

[Bol95] Axel Boldt. Methods to determine Coxeter polynomials. Linear Algebra and its Applications, 230(1–3):151–164, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9400004.

Bolotnikov:1997:GMP

[Bol97] Vladimir Bolotnikov. On a general moment problem on the half axis. Linear Algebra and its Applications, 255(1–3): 57–112, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9500786.

Boley:1998:ASP

[Bol98] Daniel L. Boley. The algebraic structure of pencils and block Toeplitz matrices. Linear Algebra and its Applications, 279(1–3):255–279, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

                                        tree/store/laa/sub/1998/
                                        279/1-3/6120.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6120.

Bomze:1996:BPS

[Bom96] Immanuel M. Bomze. Block pivoting and shortcut strategies for detecting copositivity. Linear Algebra and its Applications, 248(1–3):161–184, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=248&issue=1-
                                        3&aid=9500165.

Bondar:1994:CCI

[Bon94] James V. Bondar. Comments on and complements to: Inequalities: theory of majorization and its applications [Academic Press, New York, 1979, MR 81b:00002] by A. W. Marshall and I. Olkin. Linear Algebra and its Applications, 199:115–130, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bondari:1997:CPI

[Bon97] Siamack Bondari. Constructing the polynomial identities and central identities of degree < 9 of 3 × 3 matrices. Linear Algebra and its Applications, 258(1–3):233– 249, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600208.

Borobia:1995:MWE

[Bor95a] Alberto Borobia. (0, \frac{1}{2}, 1)

matrices which are extreme points of the generalized transitive tournament polytope. Linear Algebra and its Applications, 220(1–3):97– 110, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9300312. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Borobia:1995:NEP

[Bor95b] Alberto Borobia. On the nonnegative eigenvalue problem. Linear Algebra and its Applications, 223/224(1–3):131– 140, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400343. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Bojanczyk:1993:EHS

[BOS93] Adam W. Bojanczyk, Ruth Onn, and Allan O. Steinhardt. Existence of the hyperbolic singular value decomposition. Linear Algebra and its Applications, 185:21–30, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:1996:FMS

[Bot96] J. D. Botha. Factorizations of a matrix into symmetric matrices with prescribed nullities. Linear Algebra and its Applications, 240(1–3):79– 86, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=240&issue=1-
                                        3&aid=9400187.

Botha:1997:PMP

[Bot97]  J. D. Botha.  Products

of matrices with prescribed nullities and traces. Linear Algebra and its Applications, 252(1–3):173–198, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500680.

Botha:1998:PDM

[Bot98]  J. D. Botha.  Products

of diagonalizable matrices. Linear Algebra and its Applications, 273(1–3):65–82, April 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700344.

Botha:1999:PDM

[Bot99] J. D. Botha. Products of diagonalizable matrices over a field of characteristic two. Linear Algebra and its Applications, 286(1–3):37–44, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6229.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6229.

Bouchet:1998:MIC

[Bou98] André Bouchet. Multimatroids. IV. Chain-group representations. Linear Algebra and its Applications, 277(1– 3):271–289, June 15, 1998.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6041.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6041.

Bourin:1999:SIN

[Bou99] Jean-Christophe Bourin. Some inequalities for norms on matrices and operators. Linear Algebra and its Applications, 292(1–3):139–154, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6373.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6373.

Boyvalenkov:1995:CDD

[Boy95] Peter Boyvalenkov. Computing distance distributions of spherical designs. Linear Algebra and its Applications, 226/228(1–3):277–286, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1995&volume=226-228&issue=
1-3&aid=9500153.

Bozzo:1995:AHD

[Boz95] Enrico Bozzo. Algebras of higher dimension for displacement decompositions and computations with Toeplitz plus Hankel matrices. Linear Algebra and its Applications, 230(1–3):127–150, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300370.

Baksalary:1990:NCU

[BP90a] Jerzy K. Baksalary and Paweł R. Pordzik. A note on comparing the unrestricted and restricted least-squares estimators. Linear Algebra and its Applications, 127:371–378, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1990:CBL

[BP90b] Jerzy K. Baksalary and Simo Puntanen. Characterizations of the best linear unbiased estimator in the general Gauss– Markov model with the use of matrix partial orderings. Linear Algebra and its Applications, 127:363–370, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1990:STI

[BP90c] Jerzy K. Baksalary and Simo Puntanen. Spectrum and trace invariance criterion and its statistical applications. Linear Algebra and its Applications, 142:121–128, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1990:LOS

[BP90d] LeRoy B. Beasley and Norman J. Pullman. Linear operators that strongly preserve commuting pairs of Boolean matrices. Linear Algebra and its Applications, 132:137–143, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1991:LMS

[BP91a] Jerzy K. Baksalary and Friedrich Pukelsheim. On the Löwner, minus, and star partial orderings of nonnegative definite matrices and their squares. Linear Algebra and its Applications, 151:135–141, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1991:LOP

[BP91b] LeRoy B. Beasley and Norman J. Pullman. Linear operators preserving idempotent matrices over fields. Linear Algebra and its Applications, 146:7–20, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1992:NIE

[BP92a] Jerzy K. Baksalary and Tarmo Pukkila. A note on invariance of the eigenvalues, singular values, and norms of matrix products involving generalized inverses. Linear Algebra and its Applications, 165: 125–??, March 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1992:PWP

[BP92b] LeRoy B. Beasley and Norman J. Pullman. Polynomials which permute matrices over commutative antinegative semirings. Linear Algebra and its Applications, 165: 167–??, March 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brawer:1992:LAP

[BP92c] Robert Brawer and Magnus Pirovino. The linear algebra of the Pascal matrix. Linear Algebra and its Applications, 174:13–??, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:1993:IPC

[BP93] Dario Bini and Victor Pan.

Improved parallel computations with Toeplitz-like and Hankel-like matrices. Linear Algebra and its Applications, 188/189:3–29, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1995:PMA

[BP95] R. B. Bapat and Permanents. Permanents, max algebra and optimal assignment. Linear Algebra and its Applications, 226/228(1–3):73–86, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500304.

Berg:1998:CSS

[BP98] Lothar Berg and Gerlind Plonka. Compactly supported solutions of two-scale difference equations. Linear Algebra and its Applications, 275/276(1–3):49–75, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6021.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6021. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Barry:1999:MCE

[BP99] Ronald Paul Barry and R. Kelley Pace. Monte Carlo estimates of the log determinant of large sparse matrices. Linear Algebra and its Applications, 289(1– 3):41–54, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6009.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6009.

Baksalary:1995:ESP

[BPS95] Jerzy K. Baksalary, Friedrich Pukelsheim, and George P. H. Styan. Erratum: “Some properties of matrix partial orderings” [Linear Algebra Appl. 119 (1989), 57–85, MR 90h:15022]. Linear Algebra and its Applications, 220(1–3): 3, April 1995. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1992:CCA

[BPY92] Jerzy K. Baksalary, Simo Puntanen, and Haruo Yanai. Canonical correlations associated with symmetric reflexive generalized inverses of the dispersion matrix. Linear Algebra and its Applications, 176: 61–74, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ball:1990:MMD

[BR90] Joseph A. Ball and Marek Rakowski. Minimal McMillan degree rational matrix functions with prescribed local zero-pole structure. Linear Algebra and its Applications, 137/138:325–349, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ball:1991:NPS

[BR91] Joseph A. Ball and Marek Rakowski. Null-pole subspaces of nonregular rational matrix functions. Linear Algebra and its Applications, 159: 81–120, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:1992:MPI

[BR92] R. B. Bapat and Donald W. Robinson. The Moore– Penrose inverse over a commutative ring. Linear Algebra and its Applications, 177: 89–??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bailey:1993:MRE

[BR93] R. A. Bailey and C. A. Rowley. Maximal rank of an element of a tensor product. Linear Algebra and its Applications, 182:1–??, March 15, 1993. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolshakov:1995:UEI

[BR95] Yuri Bolshakov and Boris Reichstein. Unitary equivalence in an indefinite scalar: Product an analogue of singularvalue decomposition. Linear Algebra and its Applications, 222(1–3):155–226, June ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300295.

Bolotnikov:1999:NFJ

[BR99] V. Bolotnikov and L. Rodman. Normal forms and joint numerical ranges of doubly commuting matrices. Linear Algebra and its Applications, 301(1–3):187–194, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/37/article.
pdf.

Brezinski:1996:MVL

[Bre96] C. Brezinski. The methods of Vorobyev and Lanczos. Linear Algebra and its Applications, 234(1–3):21–41, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400081.

Brezinski:1999:MDM

[Bre99] C. Brezinski. Multiparameter descent methods. Linear Algebra and its Applications, 296(1–3):113–141, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/23/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/23/article.
pdf.

Bernhardsson:1998:RPV

[BRL98] Bo Bernhardsson, Anders Rantzer, and Qiu Li. Real perturbation values and real quadratic forms in a complex vector space. Linear Algebra and its Applications, 270(1–3): 131–154, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700232.

Boros:1991:UAT

[BRMM91] T. Boros, P. Rózsa, L. Mironovskij, and N. Mihajlov. A uniform algorithm for the transformation of multivariable systems into canonical forms. Linear Algebra and its Applications, 147:441–467, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Broyden:1990:LMR

[Bro90] C. G. Broyden. Lemke’s method—a recursive approach. Linear Algebra and its Applications, 136:257–272, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brockett:1991:DSS

[Bro91a] R. W. Brockett. Dynamical systems that sort lists, diagonalize matrices, and solve linear programming problems. Linear Algebra and its Applications, 146:79–91, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Broyden:1991:DLC

[Bro91b] C. G. Broyden. On degeneracy in linear complementarity problems. Linear Algebra and its Applications, 143: 99–110, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brockett:1994:DEM

[Bro94] Roger Brockett. Differential equations and matrix inequalities on isospectral families. Linear Algebra and its Applications, 203/204:189– 207, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brouwer:1995:TSG

[Bro95]  A. E. Brouwer.  Tough-

ness and spectrum of a graph. Linear Algebra and its Applications, 226/228 (1–3):267–271, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500154.

Bapat:1990:GII

[BRP90] R. B. Bapat, K. P. S. Bhaskara Rao, and K. Manjunatha Prasad. Generalized inverses over integral domains. Linear Algebra and its Applications, 140:181–196, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:1991:EJB

[BRS91] Rafael Brú, Leiba Rodman, and Hans Schneider. Extensions of Jordan bases for invariant subspaces of a matrix. Linear Algebra and its Applications, 150:209–225, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Brualdi:1993:EC

[Bru93] Richard Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 180: 1–??, February 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1995:EC

[Bru95] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 220 (1–3):1–??, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1996:EC

[Bru96] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 240(1–3):235–238, June ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9600229.

Brualdi:1997:EC

[Bru97] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 260 (1–3):323–??, July 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1998:EC

[Bru98] R. A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 280 (2–3):335–336, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1999:EC

[Bru99a] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 300(1–3):1–2, October 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/119/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/119/25/25/article.
pdf.

Brualdi:1999:P

[Bru99b] Richard A. Brualdi. Preface. Linear Algebra and its Applications, 302-303(1– 3):1–2, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/25/article.
pdf.

Brualdi:1999:THS

[Bru99c] Richard A. Brualdi. A tribute to Hans Schneider. Linear Algebra and its Applications, 302-303(1–3):3–16, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/26/article.
pdf.

Ball:1994:COW

[BRW94] Joseph A. Ball, Marek Rakowski, and Bostwick F. Wyman. Coupling operators, wedderburnforney spaces, and generalized inverses. Linear Algebra and its Applications, 203/204:111– 138, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berenstein:1991:RIC

[BS91] Carlos A. Berenstein and Daniele C. Struppa. Recent improvements in the complexity of the effective Nullstellensatz. Linear Algebra and its Applications, 157:203–215, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1992:TSI

[BS92a] J. K. (Jerzy K.) Baksalary and George P. H. Styan, editors. Third special issue on linear algebra and statistics: including papers originally presented at the International Workshop on Linear Models, Experimental Designs, and Related Matrix Theory held in Tampere, Finland, 6-8 August 1990, volume 176 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1992:P

[BS92b] Jerzy K. Baksalary and George P. H. Styan. Preface. Linear Algebra and its Applications, 176:1–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bergthaller:1992:DP

[BS92c] C. Bergthaller and Ivan Singer. The distance to a polyhedron. Linear Algebra and its Applications, 169:111–??, May 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:1994:LOW

[BS94a] LeRoy B. Beasley and Daniel J. Scully. Linear operators which preserve combinatorial orthogonality. Linear Algebra and its Applications, 201: 171–180, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:1994:MPS

[BS94b] Richard A. Brualdi and Bryan L. Shader. Minimum permanents on special faces of the polytope of doubly stochastic matrices. Linear Algebra and its Applications, 201:103–111, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bernstein:1995:SCS

[BS95a] M. Bernstein and N. J. A.

Sloane. Some canonical sequences of integers. Linear Algebra and its Applications, 226/228(1–3):57–72, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400245.

Boyer:1995:BDC

[BS95b] Elizabeth D. Boyer and Bryan L. Shader. On biclique decompositions of complete tpartite graphs. Linear Algebra and its Applications, 217(1–3): 31–40, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400140. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Bhimasankaram:1997:PLM

[BS97] P. Bhimasankaram and Rita SahaRay. On a partitioned linear model and some associated reduced models. Linear Algebra and its Applications, 264(1–3):329–339, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9700011.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9700011.

Ballalij:1998:NIR

[BS98] M. Ballalij and H. Sadok. New interpretation of related Huang’s methods. Linear Algebra and its Applications, 269(1):183–195, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9700064.

Bhatia:1999:OMS

[BŠ99a] Rajendra Bhatia and Peter Šemrl. Orthogonality of matrices and some distance problems. Linear Algebra and its Applications, 287(1–3):77–85, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6251.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6251.

Bhatia:1999:DDC

[BS99b] Rajendra Bhatia and Kalyan B.

Sinha. Derivations, derivatives and chain rules. Linear Algebra and its Applications, 302-303(1–3):231–244, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/39/article.
pdf.

Brewer:1999:CS

[BS99c] Jim Brewer and Wiland Schmale. (A, B)-cyclic submodules. Linear Algebra and its Applications, 301 (1–3):63–80, November 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/115/25/28/article.
                                        pdf.

Bshouty:1995:MCD

[Bsh95] Nader H. Bshouty. Multiplicative complexity of direct sums of quadratic systems. Linear Algebra and its Applications, 215(1–3):183– 223, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=215&issue=1-
                                        3&aid=9300088.

Boros:1994:SMU

[BSK94] Tibor Boros, Ali H. Sayed, and Thomas Kailath. Structured matrices and unconstrained rational interpolation problems. Linear Algebra and its Applications, 203/204:155– 188, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:1996:NRA

[BŠS96] Jerzy K. Baksalary, Peter Šemrl, and George P. H. Styan. A note on rank additivity and range additivity. Linear Algebra and its Applications, 237/238(1–3):489– 498, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400118. Special

issue honoring Calyampudi Radhakrishna Rao.

Baksalary:1991:NPD

[BT91a] Jerzy K. Baksalary and Götz Trenkler. Nonnegative and positive definiteness of matrices modified by two matrices of rank one. Linear Algebra and its Applications, 151: 169–184, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:1991:SIP

[BT91b] H. Bart and G. Ph. A. Thijsse. Similarity invariants for pairs of upper triangular Toeplitz matrices. Linear Algebra and its Applications, 147:17–44, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:1994:CSC

[BT94] H. Bart and V. È. Tsekanovskiĭ.

Complementary Schur complements. Linear Algebra and its Applications, 197/198: 651–658, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon,

Brugnano:1995:PRU

[BT95a] L. Brugnano and D. Trigiante. Polynomial roots: The ultimate answer? Linear Algebra and its Applications, 225(1):207–219, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300341.

Bushell:1995:ETI

[BT95b] P. J. Bushell and G. B. Trustrum. Erratum: “Trace inequalities for positive definite matrix power products” [Linear Algebra Appl. 132 (1990), 173–178, MR 91d:15034]. Linear Algebra and its Applications, 220(1–3): 2, April 1995. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bowman:1996:ANA

[BT96] Kevin Bowman and David A. Towers. On almost nilpotentby-Abelian Lie algebras. Linear Algebra and its Applications, 247(1–3):159–167, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=247&issue=1-
3&aid=9500097.

Barker:1997:BSB

[BT97a] George Phillip Barker and Bit-Shun Tam. Baer semirings and Baer ^{∗}-semirings of cone-preserving maps. Linear Algebra and its Applications, 256(1–3):165–183, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781116.

Baszenski:1997:FPM

[BT97b] Günter Baszenski and Manfred Tasche. Fast polynomial multiplication and convolutions related to the Discrete Cosine Transform. Linear Algebra and its Applications, 252(1–3):1–25, February ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500696.

Bekker:1997:GGI

[Bt97c] Paul A. Bekker and Jos M. F. ten Berge. Generic global identification in factor analysis. Linear Algebra and its Applications, 264(1–3):255–263, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9600363.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9600363.

Bayat:1999:LAG

[BT99] M. Bayat and H. Teimoori. The linear algebra of the generalized Pascal functional matrix. Linear Algebra and its Applications, 295(1–3):81– 89, July 01, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6409.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6409.

Ben-Tal:1990:GPL

[BTT90] Aharon Ben-Tal and Marc Teboulle. A geometric property of the least squares solution of linear equations. Linear Algebra and its Applications, 139:165–170, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bierlaire:1991:IAL

[BTT91] M. Bierlaire, Ph. L. Toint, and D. Tuyttens. On iterative algorithms for linear least squares problems with bound constraints. Linear Algebra and its Applications, 143:111– 143, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buffoni:1991:IMP

[Buf91] Giuseppe Buffoni. On the inverse of a matrix with positive definite symmetric component. Linear Algebra and its Applications, 159:1–10, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buoni:1990:ES

[Buo90a] John J. Buoni. Extensions of SSOR. Linear Algebra and its Applications, 139:67–87, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buoni:1990:SMI

[Buo90b] John J. Buoni. A stable method for the incomplete factorization of H-matrices. Linear Algebra and its Applications, 129:143–154, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baeza-Vega:1993:ABF

[BVB93] R. Baeza-Vega and R. Benavides. Associative bilinear forms in some Baric algebras. Linear Algebra and its Applications, 180:219–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolshakov:1997:PDF

[BvR^{+}97] Yuri Bolshakov, Cornelis V. M. van der Mee, André C. M. Ran, Boris Reichstein, and Leiba Rodman. Polar decompositions in finite dimensional indefinite scalar product spaces: General theory. Linear Algebra and its Applications, 261(1–3):91– 141, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600317.

Bart:1991:SRT

[BW91] H. Bart and H. K. Wimmer. Simultaneous reduction to triangular and companion forms of pairs of matrices: the case rank(I − AZ) = 1. Linear Algebra and its Applications, 150:443–461, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Bekker:1996:PVO

[BW96] Paul A. Bekker and Tom J.

Wansbeek. Proxies versus omitted variables in regression analysis. Linear Algebra and its Applications, 237/238(1–3):301– 312, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400591. Special

issue honoring Calyampudi Radhakrishna Rao.

Bai:1997:CNH

[BW97] Zhongzhi Bai and Deren Wang. A class of new hybrid algebraic multilevel preconditioning methods. Linear Algebra and its Applications, 260(1–3):223–255, July 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600307.

Blokhuis:1995:PSF

[BWS95] Aart Blokhuis, Henny A.

Wilbrink, and Attila Sali. Perfect sumsets in finite Abelian groups. Linear Algebra and its Applications, 226/228(1–3):47–56, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500232.

Beasley:1995:LOP

[BY95a] LeRoy B. Beasley and Shumin Ye. Linear operators preserving L-matrices. Linear Algebra and its Applications, 223/224(1–3):89– 98, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400186. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Beasley:1995:LOW

[BY95b] LeRoy B. Beasley and Shumin Ye. Linear operators which preserve sign-nonsingular matrices. Linear Algebra and its Applications, 217(1–3):15– 29, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=217&issue=1-
                                        3&aid=9400172. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Binding:1995:VPI

[BY95c] Paul Binding and Qiang Ye. Variational principles for indefinite eigenvalue problems. Linear Algebra and its Applications, 218(1–3):251– 262, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300187.

Beitia:1994:IBT

[BZ94] Ma. Asunción Beitia and Ion Zaballa. Invariants of the block tensor product. Linear Algebra and its Applications, 197/198:589–622, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Baragana:1995:BSI

[BZ95] I. Baragaña and I. Zaballa.

Block similarity invariants of restrictions to (A, B)- invariant subspaces. Linear Algebra and its Applications, 220(1–3):31–62, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1995&volume=220&issue=1-
3&aid=9400127. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Bevilacqua:1996:SAC

[BZ96] Roberto Bevilacqua and Paolo Zellini. Structure of algebras of commutative matrices. Linear Algebra and its Applications, 233(1–3):5– 41, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400017.

Bart:1998:SRT

[BZ98] H. Bart and R. A. Zuidwijk. Simultaneous reduction to triangular forms after extension with zeroes. Linear Algebra and its Applications, 281 (1–3):105–135, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6157.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6157.

Chetty:1998:FWC

[CA98] C. S. Karuppan Chetty and S. Maria Arulraj. Falsity of Wang’s conjecture on stars. Linear Algebra and its Applications, 277(1–3):49–56, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6038.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6038.

Cabral:1995:MPS

[Cab95] Isabel Cabral. Matrices with prescribed submatrices and number of invariant polynomials. Linear Algebra and its Applications, 219(1–3):207– 224, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300215.

Cain:1998:MPS

[CAHJ98] Bryan Cain, Luz M. De Alba, Leslie Hogben, and Charles R. Johnson. Multiplicative perturbations of stable and convergent operators. Linear Algebra and its Applications, 268(1–3):151– 169, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=268&issue=1-
                                        3&aid=9700038.

Cain:1999:CMC

[Cai99a]  Bryan E. Cain.  Conver-

gent multiples of convergent operators. Linear Algebra and its Applications, 299(1–3):171–173, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/112/21/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/112/21/28/article.
                                        pdf.

Cain:1999:OWR

[Cai99b]  Bryan E. Cain.  Opera-

tors which remain convergent when multiplied by certain Hermitian operators. Linear Algebra and its Applications, 297(1–3):57–61, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/20/article.
                                        pdf.

Clements:1997:SFI

[CALM97] D. J. Clements, B. D. O. Anderson, A. J. Laub, and J. B. Matson. Spectral factorization with imaginary-axis zeros. Linear Algebra and its Applications, 250(1–3):225– 252, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500525.

Campbell:1993:DSM

[Cam93] L. Andrew Campbell. Decomposing Samuelson maps. Linear Algebra and its Applications, 187:227–238, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cameron:1995:SOG

[Cam95] Peter J. Cameron. Sequence operators from groups. Linear Algebra and its Applications, 226/228(1–3):109–113, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500352.

Cao:1995:CNS

[Cao95] Zhi Hao Cao. On convergence of nested stationary iterative methods. Linear Algebra and its Applications, 221(1–3):

159–170, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300239.

Cao:1997:ABV

[Cao97] Zhi-Hao Cao. Avoiding breakdown in variants of the BI-CGSTAB algorithm. Linear Algebra and its Applications, 263(1–3):113–132, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=263&issue=1-
                                        3&aid=9600525.

Cao:1998:BEC

[Cao98a] Dasong Cao. Bounds on eigenvalues and chromatic numbers. Linear Algebra and its Applications, 270(1–3):1– 13, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=270&issue=1-
                                        3&aid=9600199.

Cao:1998:NTS

[Cao98b] Zhi-Hao Cao. Nonstationary two-stage multisplitting methods with overlapping blocks. Linear Algebra and its Applications, 285(1–3): 153–163, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6235.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6235.

Capizzano:1998:ETC

[Cap98a] Stefano Serra Capizzano. An ergodic theorem for classes of preconditioned matrices. Linear Algebra and its Applications, 282(1–3):161–183, October 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6176.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6176.

Capizzano:1998:KTL

[Cap98b] Stefano Serra Capizzano. Korovkin theorems and linear positive Gram matrix algebra approximations of Toeplitz matrices. Linear Algebra and its Applications, 284(1–3): 307–334, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        284/1-3/6191.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=284&issue=
                                        1-3&aid=6191.

Carstensen:1991:LCC

[Car91] C. Carstensen. Linear construction of companion matrices. Linear Algebra and its Applications, 149:191–214, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carey:1992:RHS

[Car92] John Corning Carey. The Riemann hypothesis as a sequence of surface to volume ratios. Linear Algebra and its Applications, 165:131–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carvalho:1996:QLF

[Car96]  L. A. V. Carvalho.  On

quadratic Liapunov functionals for linear difference equations. Linear Algebra and its Applications, 240(1–3):41– 64, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1996&volume=240&issue=1-
3&aid=9400183.

Carvalho:1997:JEE

[Car97] Maria Lucília Carvalho. A joint estimator for the eigenvalues of the reproduction mean matrix of a multitype Galton–Watson process. Linear Algebra and its Applications, 264(1–3):189–203, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9782946.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9782946.

Coll:1994:DTL

[CBSH94] Carmen Coll, Rafael Brú, Elena Sánchez, and Vicente Hernández. Discrete-time linear periodic realization in the frequency domain. Linear Algebra and its Applications, 203/204:301–326, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chao:1992:AMM

[CC92a] Kung-Mei Chao and Song Wong Chi. Applications of Mmatrices to majorization. Linear Algebra and its Applications, 169:31–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chernyak:1992:JRM

[CC92b] A. A. Chernyak and Z. A. Chernyak. Joint realization of (0, 1) matrices revisited. Linear Algebra and its Applications, 174:25–35, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1994:RDO

[CC94] Mei-Qin Chen and Chichia Chiu. Region-dependent optimal m-stage Runge–Kutta schemes for solving a class of nonsymmetric linear systems. Linear Algebra and its Applications, 212/213:523– 546, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Canto:1996:SGE

[CC96] Rafael Cantó and Joan-Josep Climent. Singular graph and extension of Jordan chains of an M-matrix. Linear Algebra and its Applications, 241/243(1–3):167–189, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500584. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Conforti:1997:PM

[CCF97] Michele Conforti, Gérard Cornuéjols, and Carla De Francesco. Perfect 0, ±1 matrices. Linear Algebra and its Applications, 253(1–3):299– 309, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9600773.

Cao:1997:VTR

[CCHV97] Dasong Cao, V. Chvátal, A. J. Hoffman, and A. Vince. Variations on a theorem of Ryser. Linear Algebra and its Applications, 260(1–3):215– 222, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600306.

Chang:1992:CIR

[CCK92] Stanley S. Chang and Li Chi-Kwong. Certain isometries on R^{n}. Linear Algebra and its Applications, 165:251–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Codenotti:1997:PCT

[CCR97] Bruno Codenotti, Valentino Crespi, and Giovanni Resta. On the permanent of certain (0, 1) Toeplitz matrices. Linear Algebra and its Applications, 267(1–3):65–100, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9600644.

Calvin:1992:NMS

[CD92] James A. Calvin and Richard L. Dykstra. A note on maximizing a special concave function subject to simultaneous Loewner order constraints. Linear Algebra and its Applications, 176:37–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:1996:MRH

[CD96] Nir Cohen and Jerome Dancis. Maximal rank Hermitian completions of partially specified Hermitian matrices. Linear Algebra and its Applications, 244(1–3):265–276, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=244&issue=1-
3&aid=9400235.

Coelho:1997:LPI

[CD97] M. Purificação Coelho and M. Antónia Duffner. Linear preservers of immanants on symmetric matrices. Linear Algebra and its Applications, 255(1–3):315–334, April 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600160.

Cechlarova:1999:RIE

[CD99] Katarína Cechlárová and Pavel Diko. Resolving infeasibility in extremal algebras. Linear Algebra and its Applications, 290(1–3): 267–273, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6351.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6351.

Cowen:1995:GNH

[CDS95] Carl C. Cowen, Kelly E.

Debro, and Peter D. Sepanski. Geometry and the norms of Hadamard multipliers. Linear Algebra and its Applications, 218(1–3):239– 249, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=218&issue=1-
                                        3&aid=9300186.

Cechlarova:1990:SRM

[Cec90] Katarína Cechlárová. Strong regularity of matrices in a discrete bottleneck algebra. Linear Algebra and its Applications, 128:35–50, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cechlarova:1992:EBA

[Cec92] Katarina Cechlarova. Eigenvectors in bottleneck algebra. Linear Algebra and its Applications, 175:63–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cechlarova:1996:PMB

[Cec96] Katarína Cechlárová. On the powers of matrices in bottleneck/fuzzy algebra. Linear Algebra and its Applications, 246(1–3):97–111, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400338.

Christensen:1993:MSM

[CF93] J. P. R. Christensen and P. Fischer. Multidimensional stochastic matrices and errorcorrecting codes. Linear Algebra and its Applications, 183: 255–276, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:1996:MLC

[CF96] Menglin Cao and Michael C.

Ferris. P_{c}-matrices and the linear complementarity problem. Linear Algebra and its Applications, 246(1–3):299– 312, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400362.

Compta:1997:ISH

[CF97a] A. Compta and J. Ferrer. On (A, B)t-invariant subspaces having extendible Brunovsky bases. Linear Algebra and its Applications, 255(1–3):185– 201, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1997&volume=255&issue=1-
3&aid=9600001.

Critchley:1997:SSD

[CF97b] Frank Critchley and Bernard Fichet. On (super-) spherical distance matrices and two results from Schoenberg. Linear Algebra and its Applications, 251(1–3):145–165, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500558.

Charles:1999:FJE

[CF99] Bernard Charles and Abdelkhalek Faouzi. Forme de Jordan des extensions d’opérateurs linéaires (problème de Carlson) et sous-espaces réduisants minimaux contenant un sous-espace donné [french: Jordan form of linear operation extensions (Carlson’s problem) and minimal reducing subspaces containing a given subspace]. Linear Algebra and its Applications, 288(1–3):45–51, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6283.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

year=1999&volume=288&issue=
1-3&aid=6283.

Cowen:1996:FNH

[CFJ^{+}96] C. C. Cowen, P. A. Ferguson, D. K. Jackman, E. A. Sexauer, C. Vogt, and H. J. Woolf. Finding norms of Hadamard multipliers. Linear Algebra and its Applications, 247(1–3):217–235, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500108.

Chang:1992:SAB

[CG92a] Chin Chang and Tryphon T.

Georgiou. On a Schur-Algorithm based approach to spectral factorization: Connection with the Riccati equation. Linear Algebra and its Applications, 171:233–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cottle:1992:TCS

[CG92b] Richard W. Cottle and Sy-Ming Guu. Two characterizations of sufficient matrices. Linear Algebra and its Applications, 170:65–??, June 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Constantinescu:1993:NSS

[CG93a] T. Constantinescu and A. Gheondea. The negative signature of some Hermitian matrices. Linear Algebra and its Applications, 178:17–??, January 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:1993:IBA

[CG93b] Robert Costa and Henrique Guzzo, Jr. Indecomposable Baric algebras. Linear Algebra and its Applications, 183: 223–??, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:1994:GAC

[CG94a] Chin Chang and Tryphon T. Georgiou. Geometric aspects of the Carathéodory extension problem. Linear Algebra and its Applications, 203/204:209– 251, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:1994:IBA

[CG94b] Roberto Costa and Henrique Guzzo, Jr. Indecomposable Baric algebras. II. Linear Algebra and its Applications, 196:233–242, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Curtain:1997:ASP

[CG97] Ruth Curtain and Michael Green. Analytic system problems and J-lossless coprime factorization for infinitedimensional linear systems. Linear Algebra and its Applications, 257(1–3):121–161, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600175.

Chu:1998:FAT

[CG98] Eleanor Chu and Alan George.

FFT algorithms and their adaptation to parallel processing. Linear Algebra and its Applications, 284(1–3): 95–124, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6205.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6205.

Coulton:1999:IST

[CG99] P. Coulton and H. Gauchman. Invariant subspaces of two Hermitian structures on a Euclidean space. Linear Algebra and its Applications, 301(1–3):113–119, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/115/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/115/25/31/article.
                                        pdf.

Cuninghame-Green:1995:EEB

[CGB95] R. A. Cuninghame-Green and P. Butkovič. Extremal eigenproblem for bivalent matrices. Linear Algebra and its Applications, 222(1–3):77– 89, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=222&issue=1-
                                        3&aid=9300285.

Carpraux:1996:CNK

[CGK96] J. F. Carpraux, S. K. Godunov, and S. V. Kuznetsov. Condition number of the Krylov bases and subspaces. Linear Algebra and its Applications, 248(1–3):137–160, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=248&issue=1-
                                        3&aid=9500164.

Carnicer:1999:LCP

[CGP99] J. M. Carnicer, T. N. T. Goodman, and J. M. Peña. Linear conditions for positive determinants. Linear Algebra and its Applications, 292 (1–3):39–59, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6360.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6360.

Codenotti:1998:HRS

[CGV98] Bruno Codenotti, Ivan Gerace, and Sebastiano Vigna. Hardness results and spectral techniques for combinatorial problems on circulant graphs. Linear Algebra and its Applications, 285 (1–3):123–142, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6243.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6243.

Cheon:1992:MMF

[CH92] Gi-Sang Cheon and Suk-Geun Hwang. Maximization of a matrix function related to the Dittert conjecture. Linear Algebra and its Applications, 165:153–??, March 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1996:HIR

[CH96a] Xu-Zhou Chen and Robert E. Hartwig. The hyperpower iteration revisited. Linear Algebra and its Applications, 233(1–3):207–229, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400076.

Chen:1996:GIT

[CH96b] Xuzhou Chen and Robert E. Hartwig. The group inverse of a triangular matrix. Linear Algebra and its Applications, 237/238(1–3):97– 108, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500561. Special

issue honoring Calyampudi Radhakrishna Rao.

Chen:1997:CPS

[CH97] Xuzhou Chen and Robert E. Hartwig. On the convergence of power scaled Cesáro sums. Linear Algebra and its Applications, 267(1–3):335–358, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700026.

Chen:1999:IGB

[CH99a] Gong-Ning Chen and Yong-Jian Hu. Inversion of a generalized block Loewner matrix, the minimal partial realization, and matrix rational interpolation problem. Linear Algebra and its Applications, 286(1–3):223–259, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6268.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6268.

Cheng:1999:NDP

[CH99b] Sheung Hun Cheng and Nicholas J. Higham. The nearest definite pair for the Hermitian generalized eigenvalue problem. Linear Algebra and its Applications, 302-303(1–3):63–76, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/

                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/30/article.
                                        pdf.

Chu:1999:IEA

[CH99c] Delin Chu and D. W. C. Ho. Infinite eigenvalue assignment for singular systems. Linear Algebra and its Applications, 298(1–3):21–37, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/105/17/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/19/article.
                                        pdf.

Chan:1996:RSI

[Cha96a] Chi Fai Chan. Refinements of the Schur inequality for principal characters. Linear Algebra and its Applications, 239(1–3): 15–27, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=239&issue=1-
                                        3&aid=9400159.

Chang:1996:DCD

[Cha96b] Shu-Chu Chang. Double coset decomposition for SL(r + s, Z) with respect to congruence subgroups. Linear Algebra and its Applications, 244(1–3):1– 11, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400201.

Chang:1998:SD

[Cha98] Xiao-Wen Chang. On the sensitivity of the SR decomposition. Linear Algebra and its Applications, 282(1– 3):297–310, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6196.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6196.

Carriegos:1998:PFR

[CHASG98] M. Carriegos, J. A. Hermida-Alonso, and T. Sánchez-Giralda. The pointwise feedback relation for linear dynamical systems. Linear Algebra and its Applications, 279(1–3):119–134, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6118.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1998&volume=279&issue=
                                        1-3&aid=6118.

Chen:1990:NM

[Che90a] Ying Chen. Notes on F_{0}- matrices. Linear Algebra and its Applications, 142:167–172, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1990:GBD

[Che90b] Yong Lin Chen. The generalized Bott–Duffin inverse and its applications. Linear Algebra and its Applications, 134: 71–91, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1991:SCO

[Che91] Wan Yi Chen. Stabilizability, controllability and observability of linear continuous-time systems defined over a commutative Banach algebra. Linear Algebra and its Applications, 144:1–10, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:1992:CES

[Che92] Che-Man Cheng. Cases of equality for a singular value inequality for the Hadamard product. Linear Algebra and its Applications, 177:209– ??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1995:PGP

[Che95] Wayne W. Chen. A proof of the generalized Picard’s little theorem using matrices. Linear Algebra and its Applications, 214(1–3):187– 192, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300065.

Chen:1998:DSB

[Che98] Mei-Qin Chen. A direction set based algorithm for least squares problems in adaptive signal processing. Linear Algebra and its Applications, 284(1–3):73–94, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6247.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6247.

Cherki:1999:RAP

[Che99] S. Cherki. Remarque sur l’approximation positive contractante. Linear Algebra and its Applications, 288(1– 3):243–247, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        288/1-3/6323.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=288&issue=
                                        1-3&aid=6323.

Ching:1990:BSR

[Chi90] Li Ching. A bound on the spectral radius of matrices of zeros and ones. Linear Algebra and its Applications, 132: 179–183, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chillag:1991:NMB

[Chi91] David Chillag. Nonnegative matrices, Brauer characters, normal subsets, and powers of representation modules. Linear Algebra and its Applications, 157:69–99, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chikuse:1992:GHL

[Chi92a] Yasuko Chikuse. Generalized Hermite and Laguerre polynomials in multiple symmetric matrix arguments and their applications. Linear Algebra and its Applications, 176:261– ??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chikuse:1992:PHL

[Chi92b] Yasuko Chikuse. Properties of Hermite and Laguerre polynomials in matrix argument and their applications. Linear Algebra and its Applications, 176:237–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ching:1993:MDL

[Chi93] Li Ching. The maximum determinant of an n × n lower Hessenberg (0, 1) matrix. Linear Algebra and its Applications, 183:147–??, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chikuse:1994:GNH

[Chi94] Yasuko Chikuse. Generalized noncentral Hermite and Laguerre polynomials in multiple matrices. Linear Algebra and its Applications, 210: 209–226, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chillag:1995:RRS

[Chi95] David Chillag. Regular representations of semisimple algebras, separable field extensions, group characters, generalized circulants, and generalized cyclic codes. Linear Algebra and its Applications, 218(1–3):147–183, March 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300167.

Chi:1996:LRB

[Chi96a] Mei-Hsiu Chi. The long-run behavior of Markov chains. Linear Algebra and its Applications, 244(1–3):111–121, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400216.

Chien:1996:EBN

[Chi96b]  M. T. Chien.  Erratum:

“Boundedness of the numerical range” [Linear Algebra Appl. 134 (1990), 25–30, MR 91d:15052]. Linear Algebra and its Applications, 240(1– 3):237, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:1996:NRT

[Chi96c] Mao-Ting Chien. On the numerical range of tridiagonal operators. Linear Algebra and its Applications, 246(1–3):203–214, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400353.

Chikuse:1996:ESP

[Chi96d]  Yasuko Chikuse.  Edge-

worth and saddle-point approximations for random rectangular matrices. Linear Algebra and its Applications, 237/238(1–3):133– 153, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500118. Special

issue honoring Calyampudi Radhakrishna Rao.

Chikuse:1997:DES

[Chi97]  Yasuko Chikuse.  Density

estimation on the spaces of symmetric and rectangular matrices. Linear Algebra and its Applications, 264(1–3):265–290, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700027.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700027.

Chipman:1999:LRR

[Chi99] John S. Chipman. Linear restrictions, rank reduction, and biased estimation in linear regression. Linear Algebra and its Applications, 289 (1–3):55–74, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        289/1-3/6255.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=289&issue=
                                        1-3&aid=6255.

Chkliar:1997:NRS

[Chk97] Vitali Chkliar. Numerical radii of simple powers. Linear Algebra and its Applications, 265(1–3):119–121, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600594.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600594.

Cain:1995:IMA

[CHL95] Bryan E. Cain, Roger A. Horn, and Luoluo Li. Inequalities for monotonic arrangements of eigenvalues. Linear Algebra and its Applications, 222(1–3):1–8, June ??, 1995. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300261.

Cho:1993:PBM

[Cho93a] Han H. Cho. Prime Boolean matrices and factorizations. Linear Algebra and its Applications, 190:87–98, September 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:1993:RMS

[Cho93b] Han H. Cho. Regular matrices in the semigroup of Hall matrices. Linear Algebra and its Applications, 191:151–163, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:1998:MND

[Cho98] Soojin Cho. Minimal null designs of subspace lattice over finite fields. Linear Algebra and its Applications, 282(1–3):199–220, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6183.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6183.

Choi:1997:COP

[CHR97a] Man-Duen Choi, Jinchuan Hou, and Peter Rosenthal. Completion of operator partial matrices to square-zero contractions. Linear Algebra and its Applications, 256(1– 3):1–30, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600774.

Coppersmith:1997:IRQ

[CHR97b] Don Coppersmith, Alan J. Hoffman, and Uriel G. Rothblum. Inequalities of Rayleigh quotients and bounds on the spectral radius of nonnegative symmetric matrices. Linear Algebra and its Applications, 263(1–3):201–220, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600534.

Christensen:1999:IPL

[Chr99] Erik Christensen. On invertibility preserving linear mappings, simultaneous triangularization and Property L. Linear Algebra and its Applications, 301 (1–3):153–170, November 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/35/article.
pdf.

Carlson:1992:BDS

[CHS92] David Carlson, Daniel Hershkowitz, and Dafna Shasha. Block diagonal semistability factors and Lyapunov semistability of block triangular matrices. Linear Algebra and its Applications, 172:1–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:1991:CJL

[Chu91] Moody T. Chu. A continuous Jacobi-like approach to the simultaneous reduction of real matrices. Linear Algebra and its Applications, 147:75– 96, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:1993:SGS

[Chu93] Moody T. Chu. The stability group of symmetric Toeplitz matrices. Linear Algebra and its Applications, 185:119–123, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:1995:STL

[Chu95]  Moody T. Chu.  Scaled

Toda-like flows. Linear Algebra and its Applications, 215(1–3):261–273, January 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300091.

Chu:1997:DPM

[Chu97]  Wenchang Chu.  Deter-

minant, permanent, and MacMahon’s master theorem. Linear Algebra and its Applications, 255(1–3):171– 183, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9500774.

Chu:1998:OCA

[Chu98] Moody T. Chu. On the optimal consistent approximation to pairwise comparison matrices. Linear Algebra and its Applications, 272(1–3):155– 168, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700329.

Chu:1999:CMP

[Chu99] Moody T. Chu. On constructing matrices with prescribed singular values and diagonal elements. Linear Algebra and its Applications, 288(1–3):11–22, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6241.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6241.

Chung:1992:PSD

[CHX92] S. L. Chung, F. B. Hanson, and H. H. Xu. Parallel stochastic dynamic programming: Finite element methods. Linear Algebra and its Applications, 172:197–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cigler:1999:MGF

[Cig99]  Grega Cigler.  On ma-

trix groups with finite spectrum. Linear Algebra and its Applications, 286(1–3): 287–295, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        286/1-3/6286.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=286&issue=
                                        1-3&aid=6286.

Curtis:1998:CPG

[CIM98] E. B. Curtis, D. Ingerman, and J. A. Morrow. Circular planar graphs and resistor networks. Linear Algebra and its Applications, 283 (1–3):115–150, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        283/1-3/6206.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=283&issue=
                                        1-3&aid=6206.

Cohen:1993:REU

[CIR^{+}93] Joel E. Cohen, Yoh Iwasa, Gh. Rautu, Mary Beth Ruskai, and Eugene Seneta. Relative entropy under mappings by stochastic matrices. Linear Algebra and its Applications, 179:211–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:1991:COB

[CJY91] Raymond H. Chan, Xiao Qing Jin, and Man-Chung Yeung.

The circulant operator in the Banach algebra of matrices. Linear Algebra and its Applications, 149:41–53, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chun:1991:DSH

[CK91] J. Chun and T. Kailath. Displacement structure for Hankel, Vandermonde, and related (derived) matrices. Linear Algebra and its Applications, 151:199–227, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1994:CMV

[CK94] Guan Rong Chen and Çetin Kaya Koç. Computing matrixvalued Nevanlinna–Pick interpolation. Linear Algebra and its Applications, 203/204:253– 263, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cechlarova:1997:SRM

[CK97] Katarína Cechlárová and Karol Kolesár. Strong regularity of matrices in a discrete bounded bottleneck algebra. Linear Algebra and its Applications, 256(1–3):141– 152, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1997&volume=256&issue=1-
3&aid=9781115.

Chandrasekaran:1996:ETF

[CKL96] R. Chandrasekaran, S. N. Kabadi, and S. Lakshminarayanan. An extension of a theorem of Fulkerson and Gross. Linear Algebra and its Applications, 246(1–3):23– 29, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400308.

Chen:1996:NAC

[CKN96] Yonghong Chen, Stephen J. Kirkland, and Michael Neumann. Nonnegative alternating circulants leading to Mmatrix group inverses. Linear Algebra and its Applications, 233(1–3):81–97, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400055.

Cheng:1990:CIS

[CL90] Ching-Shui Cheng and Ker-Chau Li. Characterization of invariant spaces under a symmetric real matrix and its permutations. Linear Algebra and its Applications, 127: 503–516, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:1991:BSV

[CL91] Nir Cohen and Izchak Lewkowicz. Bounds for singular values based on the eigenvectors. Linear Algebra and its Applications, 149:227–248, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1993:NCM

[cL93] Ren cang Li. Norms of certain matrices with applications to variations of the spectra of matrices and matrix pencils. Linear Algebra and its Applications, 182:199–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:1996:TSE

[CL96a] Shu-Chu Chang and Cheng Lee. Two-sided equivalence on the special linear group. Linear Algebra and its Applications, 246(1–3):191–202, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400352.

Chen:1996:CPC

[CL96b] William Y. C. Chen and James D. Louck. The combinatorial power of the companion matrix. Linear Algebra and its Applications, 232(1– 3):261–278, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:1997:HII

[CL97a] Onn Chan and T. K. Lam. Hook Immanantal inequalities for Laplacians of trees. Linear Algebra and its Applications, 261(1–3):23–47, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600305.

Cohen:1997:CIC

[CL97b] Nir Cohen and Izchak Lewkowicz. Convex invertible cones and the Lyapunov equation. Linear Algebra and its Applications, 250(1–3):105–131, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500424.

Cao:1998:SMS

[CL98a] Zhi-Hao Cao and Zhong-Yun Liu. Symmetric multisplitting of a symmetric positive definite matrix. Linear Algebra and its Applications, 285 (1–3):309–319, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6262.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6262.

Chan:1998:HII

[CL98b] Onn Chan and T. K. Lam.

Hook Immanantal inequalities for trees explained. Linear Algebra and its Applications, 273(1–3):119–131, April 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700369.

Chen:1998:NNN

[CL98c] Zhourong Chen and Wen Li.

A note on nonnegative normal matrices. Linear Algebra and its Applications, 279(1–3):281–283, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6132.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6132.

Cheng:1998:MZO

[CL98d] Bo Cheng and Bolian Liu. Matrices of zeros and ones with the maximum jump number. Linear Algebra and its Applications, 277(1–3):83–95, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6039.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6039.

Chooi:1998:LPT

[CL98e] W. L. Chooi and M. H. Lim. Linear preservers on triangular matrices. Linear Algebra and its Applications, 269(1):241–255, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700069.

Chen:1999:NPI

[CL99] Gong-Ning Chen and Xiao-Qing Li. The Nevanlinna– Pick interpolation problems and power moment problems for matrix-valued functions. Linear Algebra and its Applications, 288(1–3): 123–148, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6298.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6298.

Clark:1992:DHD

[Cla92a] Kenneth D. Clark. Decomposition of Hessenberg DAE systems to state space form. Linear Algebra and its Applications, 172:33–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Clavijo:1992:LAS

[Cla92b] M. Pilar Benito Clavijo. Lie alegebras with a small number of ideals. Linear Algebra and its Applications, 177: 233–??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coleman:1998:IIN

[CLLT98] Todd Coleman, Chi-Kwong Li, Michael Lundquist, and Thomas Travison. Isometries for the induced c-norm on square matrices and some related results. Linear Algebra and its Applications, 271(1–3): 235–256, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700270.

Calvetti:1999:SPS

[CLR99] D. Calvetti, B. Lewis, and L. Reichel. On the selection of poles in the singleinput pole placement problem. Linear Algebra and its Applications, 302-303(1– 3):331–345, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/44/article.
pdf.

Cheng:1990:RSD

[CLW90] Sui-Sun Cheng, Tzon-Tzer Lu, and Shu Hei Wu. Regular starlike domains of tridiagonal matrices. Linear Algebra and its Applications, 129:29– 54, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:1993:DUT

[CM93] Sonia P. Coelho and C. Polcino Milies. Derivations of upper triangular matrix rings. Linear Algebra and its Applications, 187:263–??, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calderbank:1995:LCO

[CM95] A. R. Calderbank and Gary McGuire. Z_{4}-linear codes obtained as projections of Kerdock and Delsarte-Goethals codes. Linear Algebra and its Applications, 226/228 (1–3):647–665, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500239.

Corpet:1996:AGM

[CM96] Florence Corpet and Artibano Micali. Algèbres genétiques et mutation quantique. (French) [genetic algebras and quantum mutation]. Linear Algebra and its Applications, 239 (1–3):1–14, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=239&issue=1-
3&aid=9400075.

Campos:1990:SDG

[CMH90] Tânia M. M. Campos, Silvia D. A. Machado, and Philip Holgate. Sex dependent genetic recombination rates for several loci. Linear Algebra and its Applications, 136:165– 172, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Crouzeix:1995:ATQ

[CMLS95] J.-P. Crouzeix, J.-E. Martínez-Legaz, and A. Seeger. An alternative theorem for quadratic forms and extensions. Linear Algebra and its Applications, 215(1–3):121–134, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300082.

Chu:1999:MNR

[CMN99] Delin Chu, Volker Mehrmann, and Nancy K. Nichols. Minimum norm regularization of descriptor systems by mixed output feedback. Linear Algebra and its Applications, 296(1–3):39–77, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/20/

abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/20/article.
pdf.

Cain:1992:SXA

[CN92] B. E. Cain and S. A. Nelson.

Solutions to x − axb = w. Linear Algebra and its Applications, 165:229–232, March 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:1993:TPH

[CN93] Raymond H. Chan and Kwok-Po Ng. Toeplitz preconditioners for Hermitian Toeplitz systems. Linear Algebra and its Applications, 190:181–208, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1997:MGI

[CN97] Yonghong Chen and M. Neumann. M-matrix generalized inverses of M-matrices. Linear Algebra and its Applications, 256(1–3):263–285, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781117.

Chan:1999:HII

[CN99a] Onn Chan and Bee-San Ng. Hook Immanantal inequalities for Hadamard’s function. Linear Algebra and its Applications, 299(1– 3):175–190, September 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/29/article.
pdf.

Chien:1999:BGC

[CN99b] Mao-Ting Chien and Hiroshi Nakazato. Boundary generating curves of the cnumerical range. Linear Algebra and its Applications, 294 (1–3):67–84, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6402.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6402.

Chan:1996:STB

[CNW96] Raymond H. Chan, Michael K. Ng, and C. K. Wong. Sine transform based preconditioners for symmetric Toeplitz systems. Linear Algebra and its Applications, 232(1–3):237– 259, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:1993:AGS

[Coe93] Sônia P. Coelho. The automorphism group of a structural matrix algebra. Linear Algebra and its Applications, 195:35–58, December 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:1996:LPI

[Coe96] M. Purificação Coelho. On linear preservers of immanants. Linear Algebra and its Applications, 247(1–3):265–271, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500115.

Cohen:1996:AAC

[Coh96] A. M. Cohen. An algebraic approach to certain differential eigenvalue problems. Linear Algebra and its Applications, 240(1–3):183– 198, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400198.

Collombier:1996:TFM

[Col96] D. Collombier. Tableaux à fréquences marginales d’ordre 2 proportionnelles. (French) [proportional second-order marginal frequency arrays]. Linear Algebra and its Applications, 237/238(1–3):509– 537, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400178. Special

issue honoring Calyampudi Radhakrishna Rao.

Cooperstein:1997:EFV

[Coo97]  B. N. Cooperstein.  Ex-

ternal flats to varieties in PG(M_{n,n}(GF(q))). Linear Algebra and its Applications, 267(1–3):175–186, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700009.

Cooperman:1998:PTO

[Coo98] Gene Cooperman. Practical task-oriented parallelism for Gaussian elimination in distributed memory. Linear Algebra and its Applications, 275/276(1–3):107–120, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6033.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6033. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Copper:1992:GDM

[Cop92] Mark Copper. On the Gauss decomposition of a matrix. Linear Algebra and its Applications, 177:175–??, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coppersmith:1993:SLE

[Cop93] Don Coppersmith. Solving linear equations over GF(2): block Lanczos algorithm. Linear Algebra and its Applications, 192:33–60, October 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Cortes:1993:NLD

[Cor93] Teresa Cortes. A note on the latice definability of Bernstein algebras. Linear Algebra and its Applications, 179: 203–??, January 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:1991:TAR

[Cos91] R. Costa. On train algebras of rank 3. Linear Algebra and its Applications, 148:1– 12, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:1995:SIG

[Cos95] Roberto Costa. Shape identities in genetic algebras. Linear Algebra and its Applications, 214(1–3):119–131, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=214&issue=1-
                                        3&aid=9300060.

Cowen:1990:AHM

[Cow90] Carl C. Cowen. An application of Hadamard multiplication to operators on weighted Hardy spaces. Linear Algebra and its Applications, 133:21– 32, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coxeter:1995:SAT

[Cox95] H. S. M. Coxeter. Some applications of trilinear coordinates. Linear Algebra and its Applications, 226/228 (1–3):375–388, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500169.

Chou:1990:CNA

[CP90] S. H. Chou and T. A.

Porsching. Contraction numbers for additive correction methods. Linear Algebra and its Applications, 139:207–215, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coullard:1995:TUL

[CP95] Collette R. Coullard and H. Ng Peh. Totally unimodular Leontief directed hypergraphs. Linear Algebra and its Applications, 230(1–3):101– 125, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300369.

Culler:1996:RSC

[CP96] Kristen W. Culler and Geoffrey L. Price. On the ranks of skew-centrosymmetric matrices over finite fields. Linear Algebra and its Applications, 248(1–3):317–325, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500249.

Chang:1998:SAF

[CP98a] Xiao-Wen Chang and Christopher C. Paige. Sensitivity analyses for factorizations of sparse or structured matrices. Linear Algebra and its Applications, 284(1–3): 53–71, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6219.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6219.

Climent:1998:SCT

[CP98b] Joan-Josep Climent and Carmen Perea. Some comparison theorems for weak nonnegative splittings of bounded operators. Linear Algebra and its Applications, 275/276 (1–3):77–106, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6058.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/

laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6058. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Capizzano:1999:SSA

[CP99a] Stefano Serra Capizzano and Cristina Tablino Possio. Spectral and structural analysis of high precision finite difference matrices for elliptic operators. Linear Algebra and its Applications, 293(1–3):85–131, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6371.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6371.

Climent:1999:CTW

[CP99b] Joan-Josep Climent and Carmen Perea. Comparison theorems for weak splittings in respect to a proper cone of nonsingular matrices. Linear Algebra and its Applications, 302-303(1– 3):355–366, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
46/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/117/25/46/article.
                                        pdf.

Corach:1990:OI

[CPR90] G. Corach, H. Porta, and L. Recht. An operator inequality. Linear Algebra and its Applications, 142:153–158, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Corach:1993:JFS

[CPR93] G. Corach, H. Porta, and L. Recht. Jacobi fields on space of positive operators. Linear Algebra and its Applications, 179:271–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calvetti:1992:CVS

[CR92] D. Calvetti and L. Reichel. A Chebychev-Vandermonde solver. Linear Algebra and its Applications, 172:219–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:1993:NRD

[CR93] Joel E. Cohen and Uriel G. Rothblum. Nonnegative ranks, decompositions, and factorizations of nonnegative matrices. Linear Algebra and its Applications, 190:149–168, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:1994:SCD

[CR94a] Man Duen Choi and Peter Rosenthal. A survey of Chandler Davis. Linear Algebra and its Applications, 208/209: 3–18, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Curtain:1994:NSC

[CR94b] Ruth F. Curtain and Alejandro Rodríguez. Necessary and sufficient conditions for Jspectral factorizations with a J-lossless property for infinitedimensional systems in continuous and discrete time. Linear Algebra and its Applications, 203/204:327–358, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cramer:1999:EMC

[Cra99] Erhard Cramer. Estimation of the mean and the covariance matrix under a marginal independence assumption — an application of matrix differential calculus. Linear Algebra and its Applications, 288(1–3):219–228, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6320.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6320.

Cross:1998:ITP

[Cro98] R. W. Cross. An index theorem for the product of linear relations. Linear Algebra and its Applications, 277(1–3):127–134, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        277/1-3/6065.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6065.

Cvetkovic:1993:SEG

[CRS93] Dragoš Cvetković, Peter Rowlinson, and Slobodan Simić. A study of eigenspaces of graphs. Linear Algebra and its Applications, 182:45– 66, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:1994:ECE

[CRS94] Man Duen Choi, Mary Beth Ruskai, and Eugene Seneta. Equivalence of certain entropy contraction coefficients. Linear Algebra and its Applications, 208/209:29–36, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:1999:SCG

[CRS99] Dragoš Cvetković, Peter Rowlinson, and Slobodan K.

Simić. Some characterizations of graphs by star complements. Linear Algebra and its Applications, 301(1–3):81–97, November 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/29/article.
pdf.

Cruz:1991:LLS

[Cru91] Ricardo N. Cruz. Linear and Lipschitz similarity. Linear Algebra and its Applications, 151:17–25, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cabral:1991:UTC

[CS91] Isabel Cabral and Fernando C.

Silva. Unified theorems on completions of matrix pencils. Linear Algebra and its Applications, 159:43–54, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cabral:1992:SIC

[CS92] Isabel Cabral and Fernando C.

Silva. Similarity invariants of completions of submatrices. Linear Algebra and its Applications, 169:151–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chalmers:1996:ECU

[CS96] B. L. Chalmers and B. Shekhtman. Extension constants of unconditional two-dimensional operators. Linear Algebra and its Applications, 240(1–3): 173–182, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400196.

Cevik:1997:RIP

[CS97] M. K. K. Cevik and J. M. Schumacher. Regulation as an interpolation problem. Linear Algebra and its Applications, 253(1–3):311– 340, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500739.

Chalmers:1998:AC

[CS98a] B. Chalmers and B. Shekhtman. Actions that characterize l_{∞}^{(n)}. Linear Algebra and its Applications, 270(1–3): 155–169, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700233.

Cheon:1998:EPM

[CS98b] Gi-Sang Cheon and Seok-Zun Song. On the extreme points on the majorization polytope Ω_{3}(y ≺ x). Linear Algebra and its Applications, 269(1): 47–52, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700055.

Campbell:1991:WTE

[CT91] James T. Campbell and Elizabeth C. Trouy. When are two elements of GL(2, Z) similar. Linear Algebra and its Applications, 157:175–184, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:1994:CNR

[CT94] Mao Ting Chien and Bit Shun Tam. Circularity of the numerical range. Linear Algebra and its Applications, 201: 113–133, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Canto:1999:SCB

[CU99] Rafael Cantó and Ana M. Urbano. On the Segre characteristic of a block triangular matrix. Linear Algebra and its Applications, 302-303(1–3):245–263, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/40/article.
pdf.

Cuypers:1995:RQP

[Cuy95]  Hans Cuypers.  Regular

quaternionic polytopes. Linear Algebra and its Applications, 226/228(1–3):311–329, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500149.

Cao:1993:SRP

[CV93] Da Song Cao and Andrew Vince. The spectral radius of a planar graph. Linear Algebra and its Applications, 187: 251–257, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chaganty:1997:IPC

[CV97] N. Rao Chaganty and A. K. Vaish. An invariance property of common statistical tests. Linear Algebra and its Applications, 264 (1–3):421–437, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700032.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700032.

Choi:1990:CCP

[CW90] Man Duen Choi and Pei Yuan Wu. Convex combinations of projections. Linear Algebra and its Applications, 136:25– 42, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:1993:SSM

[CW93] Xiao Wen Chang and Jia Song Wang. The symmetric solution of the matrix equations AX + Y A = C, AXA^{T} + BY B^{T} = C, and (A^{T}XA, B^{T}XB) = (C, D). Linear Algebra and its Applications, 179:171–189, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:1995:BCC

[CW95a] Raymond H. Chan and C. K.

Wong. Best-conditioned circulant preconditioners. Linear Algebra and its Applications, 218(1–3):205–211, March 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300171.

Cohen:1995:MSL

[CW95b] Arjeh M. Cohen and David B. Wales. SL(3, 3) is not a maximal subgroup of the Lie group of type F_{4}. Linear Algebra and its Applications, 226/228 (1–3):287–296, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500151.

Cao:1999:TSI

[CWL99] Zhi-Hao Cao, He-Bing Wu, and Zhong-Yun Liu. Twostage iterative methods for consistent Hermitian positive semidefinite systems. Linear Algebra and its Applications, 294(1–3):217–238, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6422.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

year=1999&volume=294&issue=
1-3&aid=6422.

Chen:1996:PAL

[CWX96] Guoliang Chen, Musheng Wei, and Yifeng Xue. Perturbation analysis of the least squares solution in Hilbert spaces. Linear Algebra and its Applications, 244(1–3):69– 80, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400210.

Cao:1998:MSA

[CX98a] Chongguang Cao and Zhang Xian. Multiplicative semigroup automorphisms of upper triangular matrices over rings. Linear Algebra and its Applications, 278(1–3):85– 90, July 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6035.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6035.

Chen:1998:EGI

[CX98b] Guoliang Chen and Yifeng Xue. The expression of the generalized inverse of the perturbed operator under Type I perturbation in Hilbert spaces. Linear Algebra and its Applications, 285(1–3):1–6, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6187.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6187.

Cao:1995:DEG

[CY95] Da Song Cao and Hong Yuan. The distribution of eigenvalues of graphs. Linear Algebra and its Applications, 216(1–3):211–224, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300135.

Cheng:1997:MEF

[CY97] Hon-Wing Cheng and Stephen S.-T. Yau. More explicit formulas for the matrix exponential. Linear Algebra and its Applications, 262(1–3):131– 163, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600478.

Chien:1998:GPN

[CYYL98] Mao-Ting Chien, Lina Yeh, Yean-Tzuoh Yeh, and Fun-Zin Lin. On geometric properties of the numerical range. Linear Algebra and its Applications, 274(1–3):389– 410, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700373.

Chen:1990:INB

[CZ90] Yong Lin Chen and Bing Jun Zhou. On g-inverses and the nonsingularity of a bordered

matrix ({AB \atop C0}). Linear Algebra and its Applications, 133: 133–151, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Censor:1991:ICM

[CZ91] Yair Censor and Stavros A.

Zenios. Interval-constrained matrix balancing. Linear Algebra and its Applications, 150:393–421, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Chen:1994:MLM

[CZ94] Gong Ning Chen and Hui Pin Zhang. More on Loewner matrices. Linear Algebra and its Applications, 203/204:265– 300, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1995:NPB

[CZ95] Gong Ning Chen and Hui Pin Zhang. Note on products of Bezoutians and Hankel matrices. Linear Algebra and its Applications, 225(1):23– 35, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300305.

Cao:1996:AOP

[CZ96] Chongguang Cao and Xian Zhang. Additive operators preserving idempotent matrices over fields and applications. Linear Algebra and its Applications, 248(1–3):327– 338, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9689195.

Chen:1997:UBG

[CZ97] Xi Chen and Ke Min Zhang.

The upper bounds of the generalized maximum density index of irreducible Boolean matrices. Linear Algebra and its Applications, 256(1–3):75– 93, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600169.

Demmel:1999:CSV

[D^{+}99] James Demmel et al. Computing the singular value decomposition with high relative accuracy. Linear Algebra and its Applications, 299(1–3):21–80, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/22/article.
pdf.

daGracaMarques:1991:MPN

[da 91] Maria da Graça Marques. Matrices with prescribed nonprincipal blocks and the trace function. Linear Algebra and its Applications, 146:1– 6, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DaSilva:1994:RSC

[Da 94] J. A. Da Silva. Report: Second Conference of the International Linear Algebra Society (Lisbon, 1992). Linear Algebra and its Applications, 197/198: 793–??, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daGracaMarques:1997:NIPb

[da 97] Maria da Graça Marques. The number of invariant polynomials of a matrix with prescribed off-diagonal blocks. Linear Algebra and its Applications, 258(1–3):149–158, June ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=258&issue=1-
                                        3&aid=9600202.

daFonseca:1998:ICH

[da 98] C. M. da Fonseca. The inertia of certain Hermitian block matrices. Linear Algebra and its Applications, 274(1–3):193– 210, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=274&issue=1-
                                        3&aid=9700338.

Dusaussoy:1990:SNM

[DA90] Nicolas J. Dusaussoy and Ikram E. Abdou. Some new multiplicative algorithms for image reconstruction from projections. Linear Algebra and its Applications, 130:111– 132, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

DAlessandro:1997:GAR

[D’A97] Domenico D’Alessandro. Geometric aspects of the Riccati difference equation in the nonsymmetric case. Linear Algebra and its Applications, 255 (1–3):1–18, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600738.

DAlessandro:1998:IMP

[D’A98] Domenico D’Alessandro. Invariant manifolds and projective combinations of solutions of the Riccati differential equation. Linear Algebra and its Applications, 279(1–3):181–193, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6111.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6111.

Dahlhaus:1990:AIT

[Dah90] R. Dahlhaus. Approximations for the inverse of Toeplitz matrices with applications to stationary processes. Linear Algebra and its Applications, 127:27–40, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:1998:MPS

[Dah98] Geir Dahl. Majorization, polyhedra, and statistical testing problems. Linear Algebra and its Applications, 272(1–3): 205–225, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=272&issue=1-
                                        3&aid=9700336.

Dahl:1999:MP

[Dah99a]  Geir Dahl.  Majorization

polytopes. Linear Algebra and its Applications, 297(1–3):157–175, August 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/

10/30/19/104/17/26/article.
pdf.

Dahl:1999:MM

[Dah99b]  Geir Dahl.  Matrix ma-

jorization. Linear Algebra and its Applications, 288(1– 3):53–73, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6285.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6285.

Dai:1990:SSL

[Dai90] Hua Dai. On the symmetric solutions of linear matrix equations. Linear Algebra and its Applications, 131:1– 7, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dai:1996:CSB

[Dai96] Hua Dai. Completing a symmetric 2 × 2 block matrix and its inverse. Linear Algebra and its Applications, 235(1–3): 235–245, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400144.

Dai:1999:ASG

[Dai99] Hua Dai. An algorithm for symmetric generalized inverse eigenvalue problems. Linear Algebra and its Applications, 296(1–3):79–98, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/47/17/21/
                                        abstract.html;   http://
                                        www.elsevier.nl/gej-ng/
                                        10/30/19/47/17/21/article.
                                        pdf.

Dancis:1990:SCI

[Dan90] Jerome Dancis. Several consequences of an inertia theorem. Linear Algebra and its Applications, 136:43–61, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dancis:1991:ICB

[Dan91a] Jerome Dancis. Invertible completions of band matrices. Linear Algebra and its Applications, 150:125–138, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Dancis:1991:OBS

[Dan91b] Jerome Dancis. The optimal ω is not best for the SOR iteration method. Linear Algebra and its Applications, 154/156:

819–845, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dancis:1992:PSC

[Dan92]  Jerome Dancis.  Positive

semidefinite completions of partial Hermitian matrices. Linear Algebra and its Applications, 175:97–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Danao:1997:NM

[Dan97] R. A. Danao. A note on E′-matrices. Linear Algebra and its Applications, 259(1–3): 299–305, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600294.

Datta:1991:ILS

[Dat91] K. B. Datta. Invertibility of linear systems described by differential-difference equations. Linear Algebra and its Applications, 151:57–83, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:1994:LNL

[Dat94] Biswa Nath Datta. Linear and numerical linear algebra in control theory: Some research problems. Linear Algebra and its Applications, 197/198:755– 790, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Datta:1999:SI

[Dat99] Biswa Nath Datta. Stability and inertia. Linear Algebra and its Applications, 302-303(1–3):563–600, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
56/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/56/article.
pdf.

Davis:1999:SFGa

[Dav99a] A. W. Davis. Spherical functions on the Grassmann manifold and generalized Jacobi polynomials — Part 1. Linear Algebra and its Applications, 289(1–3):75–94, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6288.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6288.

Davis:1999:SFGb

[Dav99b] A. W. Davis. Spherical functions on the Grassmann manifold and generalized Jacobi polynomials — Part 2. Linear Algebra and its Applications, 289(1–3):95–119, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6289.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6289.

Dax:1990:LSA

[Dax90a] Achiya Dax. Line search acceleration of iterative methods. Linear Algebra and its Applications, 130:43–63, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Dax:1990:NCL

[Dax90b] Achiya Dax. A note on the convergence of linear stationary iterative processes. Linear Algebra and its Applications, 129:131–142, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dax:1990:HAL

[Dax90c] Achiya Dax. On hybrid acceleration of a linear stationary iterative process. Linear Algebra and its Applications, 130: 99–110, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Dax:1991:RRM

[Dax91] Achiya Dax. A row relaxation method for large ℓ_{1} problems. Linear Algebra and its Applications, 154/156:793– 818, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dazord:1994:NRM

[Daz94] Jean Dazord. On the Cnumerical range of a matrix. Linear Algebra and its Applications, 212/213:21– 29, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Dazord:1997:PED

[Daz97]  Jean Dazord.  Une pro-

priété extrémale de la diagonale d’une matrice. (French) [An extremal property of the diagonal of a matrix]. Linear Algebra and its Applications, 254(1–3):67–77, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1997&volume=254&issue=1-
3&aid=9600516. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Dazord:1998:CTN

[Daz98] Jean Dazord. Comparison of two norms of matrices. Linear Algebra and its Applications, 280(2–3):173–187, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6137.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6137.

DeSamblanx:1999:RIF

[DB99] G. De Samblanx and A. Bultheel.

Restrictions on implicit filtering techniques for orthogonal projection methods. Linear Algebra and its Applications, 286(1–3):45–68, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6264.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6264.

Domich:1991:OTD

[DBRW91] Paul D. Domich, Paul T.

Boggs, Janet E. Rogers, and Christoph Witzgall. Optimizing over three-dimensional subspaces in an interior-point method for linear programming. Linear Algebra and its Applications, 152:315–342, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Drury:1992:DCM

[DC92] S. W. Drury and B. Cload. On the determinantal conjecture of Marcus and de Oliveira. Linear Algebra and its Applications, 177:105–??, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delai:1993:DAP

[DC93] M. Benlarbi Delai and B. Charles. Description de AlgLat A pour un opérateur A algébrique. (French) [Description of AlgLat A for an algebraic operator A]. Linear Algebra and its Applications, 187:105–??, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeLeone:1996:RLC

[DCPS96] Renato De Leone, R. W. Cottle, J-S Pang, and R. E. Stone. Review of The Linear Complementarity Problem by R. W. Cottle, J.-S. Pang and R. E. Stone. Linear Algebra and its Applications, 235(1–3):

275–276, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400202.

Datta:1990:ECF

[DD90] B. N. Datta and Karabi Datta.

On eigenvalue and canonical form assignments. Linear Algebra and its Applications, 131:161–182, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeSchutter:1997:NCE

[DD97a] Bart De Schutter and Bart De Moor. A note on the characteristic equation in the max-plus algebra. Linear Algebra and its Applications, 261(1–3):237–250, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600407.

DeSchutter:1997:SVD

[DD97b] Bart De Schutter and Bart De Moor. The singularvalue decomposition in the extended max algebra. Linear Algebra and its Applications, 250(1–3):143–176, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500455.

Dean:1996:AWN

[DdlP96] A. Dean and J. A. de la Peña. Algorithms for weakly nonnegative quadratic forms. Linear Algebra and its Applications, 235(1–3):35–46, March 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400109.

Duffner:1994:PMS

[DdO94] Maria Antónia Duffner and G. N. de Oliveira. Pairs of matrices satisfying certain polynomial identities. Linear Algebra and its Applications, 197/198:177–188, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

DiasdaSilva:1990:CM

[DdPC90] J. A. Dias da Silva and Ma. da Purificação Coelho. (λ, G)- critical matrices. Linear Algebra and its Applications, 140: 1–11, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

dePillis:1991:PSL

[de 91] J. de Pillis. A parallelizable SOR-like method: systems with plus-shaped and linear spectra. Linear Algebra and its Applications, 154/156: 551–582, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeMoor:1992:SGP

[De 92] Bart L. R. De Moor. On the structure and geometry of the product singular value decomposition. Linear Algebra and its Applications, 168: 95–??, April 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeMoor:1993:STL

[De 93] Bart De Moor. Structured total least squares and L_{2} approximation problems. Linear Algebra and its Applications, 188/189:163–??, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deCaen:1996:DIC

[de 96a] D. de Caen. Degree of indecomposability of certain highly regular zero-one matrices. Linear Algebra and its Applications, 247(1–3):151– 158, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500111.

deOliveira:1996:LPS

[de 96b] G. N. de Oliveira. Linear preservers of Schur functions. Linear Algebra and its Applications, 241/243(1–3):203– 210, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500586. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Demmel:1995:DMM

[DE95] James W. Demmel and Alan Edelman. The dimension of matrices (matrix pencils) with given Jordan (Kronecker) canonical forms. Linear Algebra and its Applications, 230(1–3):61–87, November 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1995&volume=230&issue=1-
3&aid=9300362.

DeAlba:1996:IST

[DeA96]  Luz M. DeAlba.  Inertia

of the Stein transformation with respect to some nonderogatory matrices. Linear Algebra and its Applications, 241/243(1–3):191–201, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500546. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Dedieu:1997:COC

[Ded97] Jean-Pierre Dedieu. Condition operators, condition numbers, and condition number theorem for the generalized eigenvalue problem. Linear Algebra and its Applications, 263(1–3):1–24, September 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600366.

Deif:1991:SVI

[Dei91] A. S. Deif. Singular values of an interval matrix. Linear Algebra and its Applications, 151:125–133, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delebecque:1990:RAS

[Del90] F. Delebecque. On the resolvent approach to the spectral decomposition of a regular matrix pencil. Linear Algebra and its Applications, 129:63– 75, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delai:1999:EOA

[Del99] M’hammed Benlarbi Delai. Extensions d’opérateurs auto-adjoints et défaut de réflexivité. Linear Algebra and its Applications, 297(1–3):81–85, August 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/22/article.
                                        pdf.

Datta:1997:OPP

[DER97] Biswa N. Datta, Sylvan Elhay, and Yitshak M. Ram. Orthogonality and partial pole assignment for the symmetric definite quadratic pencil. Linear Algebra and its Applications, 257(1–3):29– 48, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600036.

Dargen:1995:CCN

[DF95] Karen Casey Dargen and Kathryn Fraughnaugh. Conditional chromatic numbers with forbidden cycles. Linear Algebra and its Applications, 217(1–3):53–66, March 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400139. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Daumer:1998:LCP

[DF98] Martin Daumer and Markus Falk. On-line change-point detection (for state space models) using multi-process Kalman filters. Linear Algebra and its Applications, 284 (1–3):125–135, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6249.pdf;   http:
//www.elsevier.com/cgi-

                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=284&issue=
                                        1-3&aid=6249.

DiasdaSilva:1998:CNE

[DFF98] J. A. Dias da Silva, Rosário Fernandes, and Amélia Fonseca. The covering number of the elements of a matroid and generalized matrix functions. Linear Algebra and its Applications, 271(1–3):191– 219, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=271&issue=1-
                                        3&aid=9700268.

Demmel:1993:CAS

[DG93] James W. Demmel and William Gragg. On computing accurate singular values and eigenvalues of matrices with acyclic graphs. Linear Algebra and its Applications, 185:203– 217, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deza:1995:DPC

[DG95a] Michel Deza and Viatcheslav Grishukhin. Delaunay polytopes of cut lattices. Linear Algebra and its Applications, 226/228 (1–3):667–685, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500240.

Ding:1995:QGA

[DG95b] Songqing Ding and William H.

Gustafson. A question of Guralnick about reflexive algebras. Linear Algebra and its Applications, 214(1–3):11– 16, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300050.

deCaen:1992:AME

[dGK^{+}92] D. de Caen, D. A. Gregory, S. J. Kirkland, N. J. Pullman, and J. S. Maybee. Algebraic multiplicity of the eigenvalues of a tournament matrix. Linear Algebra and its Applications, 169:179–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiasdaSilva:1990:NMP

[DH90] J. A. Dias da Silva and Yahya Ould Hamidoune. A note on the minimal polynomial of the Kronecker sum of two linear operators. Linear Algebra and its Applications, 141:283–287, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:1994:PLS

[DH94] J. Ding and L. J. Huang. On the perturbation of the least squares solutions in Hilbert spaces. Linear Algebra and its Applications, 212/213:487– 500, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Ding:1997:CGI

[DH97a] J. Ding and L. J. Huang. On the continuity of generalized inverses of linear operators in Hilbert spaces. Linear Algebra and its Applications, 262(1–3):229–242, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=262&issue=1-
                                        3&aid=9600509.

Hou:1997:CCM

[dH97b] Xiang dong Hou. On the centralizer of the centralizer of a matrix. Linear Algebra and its Applications, 256(1–3): 251–261, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/

browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600035.

Ding:1998:NPR

[DH98] Jiu Ding and Weihua Hang.

New perturbation results for equality-constrained least squares problems. Linear Algebra and its Applications, 272(1–3):181–192, March 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700332.

Datta:1997:ALT

[DHL97] Karabi Datta, YooPyo Hong, and Ran Baik Lee. Applications of linear transformations to matrix equations. Linear Algebra and its Applications, 267(1–3):221–240, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700014.

Davidson:1995:RDI

[DHM95] K. R. Davidson, K. J. Harrison, and U. A. Mueller. Rank decomposability in incident spaces. Linear Algebra and its Applications, 230(1–3):3– 19, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=230&issue=1-
                                        3&aid=9300351.

DiasdaSilva:1994:EMP

[DHS94a] J. A. Dias da Silva, Daniel Hershkowitz, and Hans Schneider. Existence of matrices with prescribed off-diagonal block element maxima. Linear Algebra and its Applications, 212/213:367–385, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

DiasdaSilva:1994:SDS

[DHS94b] J. A. Dias da Silva, Daniel Hershkowitz, and Hans Schneider. Sum decompositions of symmetric matrices. Linear Algebra and its Applications, 208/209:523–537, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiBenedetto:1997:GUP

[Di 97] Fabio Di Benedetto. Generalized updating problems and computation of the eigenvalues of rational Toeplitz matrices. Linear Algebra and its Applications, 267(1–3):187– 219, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700012.

DiBenedetto:1998:STN

[Di 98] Fabio Di Benedetto. Solution of Toeplitz normal equations by sine transform based preconditioning. Linear Algebra and its Applications, 285 (1–3):229–255, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6232.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6232.

DiasdaSilva:1993:NPD

[Dia93] J. A. Dias da Silva. A note on preservers of decomposability. Linear Algebra and its Applications, 186:215–225, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiasdaSilva:1996:FET

[Dia96a] J. A. Dias da Silva. Flags and equality of tensors. Linear Algebra and its Applications, 232 (1–3):55–75, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiasdaSilva:1996:MAR

[Dia96b] J. A. Dias da Silva. Multilinear algebra: Recent applications. Linear Algebra and its Applications, 241/243(1–3):211–223, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500684. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

DiasdaSilva:1996:NCE

[Dia96c] J. A. Dias da Silva. New conditions for equality of decomposable symmetrized tensors. Linear Algebra and its Applications, 245(1–3):353–372, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=245&issue=1-
                                        3&aid=9400347.

Dietrich:1990:MSA

[Die90] Brenda L. Dietrich. Monge sequences, antimatroids, and the transportation problem with forbidden arcs. Linear Algebra and its Applications, 139:133–145, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dieci:1994:SPP

[Die94] Luca Dieci. Structure preserving piecewise polynomial interpolation for definite matrices. Linear Algebra and its Applications, 202:25–32, April 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dieci:1996:CCR

[Die96]  Luca Dieci.  Considera-

tions on computing real logarithms of matrices, Hamiltonian logarithms, and skewsymmetric logarithms. Linear Algebra and its Applications, 244(1–3):35–54, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400206.

Dieci:1998:RHL

[Die98] Luca Dieci. Real Hamiltonian logarithm of a symplectic matrix. Linear Algebra and its Applications, 281 (1–3):227–246, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        281/1-3/6164.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6164.

Ding:1992:SVR

[Din92] Lifeng Ding. Separating vectors and reflexivity. Linear Algebra and its Applications, 174:37–??, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:1993:CAC

[Din93a] Jiu Ding. A continuation algorithm for a class of linear complementarity problems using an extrapolation technique. Linear Algebra and its Applications, 186:199–214, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:1993:PAP

[Din93b] Jiu Ding. Perturbation analysis for the projection of a point to an affine set. Linear Algebra and its Applications, 191:199– 212, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dekimpe:1997:FDC

[DIO97] Karel Dekimpe, Paul Igodt, and Veerle Ongenae. The five-dimensional complete leftsymmetric algebra structures compatible with an Abelian Lie algebra structure. Linear Algebra and its Applications, 263(1–3):349–375, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600591.

delaHarpe:1995:CIF

[dJ95a] Pierre de la Harpe and François Jaeger. Chromatic invariants for finite graphs: Theme and polynomial variations. Linear Algebra and its Applications, 226/228 (1–3):687–722, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500248.

DeAlba:1995:PIC

[DJ95b] Luz M. DeAlba and Charles R.

Johnson. Possible inertia combinations in the Stein and Lyapunov equations. Linear Algebra and its Applications, 222(1–3):227–240, June ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=222&issue=1-
                                        3&aid=9300296.

DuPre:1992:DPB

[DK92] Arthur M. DuPre and Seymour Kass. Distance and parallelism between flats in R^{n}. Linear Algebra and its Applications, 171:99–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Daubechies:1992:SMA

[DL92] Ingrid Daubechies and Jeffrey C. Lagarias. Sets of matrices all infinite products of which converge. Linear Algebra and its Applications, 161(??):227–263, ???? 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [DL01].

DiasdaSilva:1999:SSM

[DL99] J. A. Dias da Silva and T. J. Laffey. On simultaneous similarity of matrices and related questions. Linear Algebra and its Applications, 291(1–3):167–184, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        291/1-3/6350.pdf;   http:
                                        //www.elsevier.com/cgi-

bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6350.

Daubechies:2001:CAS

[DL01] Ingrid Daubechies and Jeffrey C. Lagarias. Corrigendum/addendum to: Sets of matrices all infinite products of which converge. Linear Algebra and its Applications, 327(1–3):69–83, April 15, 2001. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej%2Dng/10/30/19/152/25/
31/abstract.html;   http:
//www.elsevier.nl/gej%2Dng/
10/30/19/152/25/31/article.
pdf.

DiasdaSilva:1994:SCI

[DLdO94] Jose Dias de Silva, Chi-Kwong Li, and Graciano de Oliveira, editors. Second Conference of the ILAS, Lisbon, Portugal, August 3–7, 1992, volume 197/198 of Linear Algebra and its Applications. Elsevier Science Publishers, Amsterdam, The Netherlands, 1994. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dutton:1995:CUR

[DMB95] Ronald D. Dutton, Sirisha R.

Medidi, and Robert C. Brigham. Changing and unchanging of the radius of a graph. Linear Algebra and its Applications, 217(1–3):67– 82, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=217&issue=1-
                                        3&aid=9400153. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Dauxois:1997:CAT

[DN97] J. Dauxois and G. M. Nkiet. Canonical analysis of two Euclidean subspaces and its applications. Linear Algebra and its Applications, 264 (1–3):355–388, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9600244.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9600244.

Davidson:1994:SED

[DO94] Kenneth R. Davidson and Marc S. Ordower. Some exact distance constants. Linear Algebra and its Applications, 208/209:37–55, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dobovisek:1999:USE

[Dob99]  Mirko Dobovišek.  Uni-

tary solutions of the equation cu + u^{∗}c = 2d. Linear Algebra and its Applications, 296(1–3):213–225, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/29/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/29/article.
pdf.

deOliveira:1999:LCS

[dOGH99] M. C. de Oliveira, J. C.

Geromel, and Liu Hsu. LMI characterization of structural and robust stability: the discrete-time case. Linear Algebra and its Applications, 296(1–3):27–38, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/19/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/19/article.
pdf.

Domokos:1996:CVE

[Dom96] M. Domokos. Criteria for vanishing of Eulerian polynomials on n × n matrices. Linear Algebra and its Applications, 234(1–3):181–195, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=234&issue=1-
                                        3&aid=9400101.

Dongyi:1992:WGE

[Don92] Ye Dongyi. When is the greatest eigenvalue of a parametrized symmetric matrix a convex function of the parameter? Linear Algebra and its Applications, 169:103– ??, May 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dorier:1998:RFT

[Dor98] Jean-Luc Dorier. The role of formalism in the teaching of the theory of vector spaces. Linear Algebra and its Applications, 275/276(1– 3):141–160, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6053.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6053. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Dixon:1990:SAJ

[DPPR90] J. D. Dixon, J. C. Poland, I. S. Pressman, and L. Ribes. The shuffle algorithm and Jordan blocks. Linear Algebra and its Applications, 142:159– 165, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dehghan:1991:MAR

[DR91a] M. A. Dehghan and M. Radjabalipour. Matrix algebras and Radjavi’s trace condition. Linear Algebra and its Applications, 148:19–25, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dlab:1991:CFP

[DR91b] Vlastimil Dlab and Claus Michael Ringel. Canonical forms of pairs of complex matrices. Linear Algebra and its Applications, 147:387–410, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Doust:1992:SPH

[DR92] Ian Doust and Werner J.

Ricker. Spectral projections for Hermitian operators. Linear Algebra and its Applications, 175:75–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:1993:FSS

[DR93] Biswa Nath Datta and Fernando Rincón. Feedback stabilization of a second-order system: A nonmodal approach. Linear Algebra and its Applications, 188/189:135– 161, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deif:1994:ISP

[DR94] A. Deif and J. Rohn. On the invariance of the sign pattern of matrix eigenvectors under perturbation. Linear Algebra and its Applications, 196: 63–70, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dehghan:1996:PUC

[DR96] M. A. Dehghan and M. Radjabalipour. On products of unbounded collections of matrices. Linear Algebra and its Applications, 233(1–3):43– 49, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400038.

Arias-de-Reyna:1998:GVP

[dR98] J. Arias de Reyna. Gaussian variables, polynominals and permanents. Linear Algebra and its Applications, 285 (1–3):107–114, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6242.pdf;   http:
//www.elsevier.com/cgi-

bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6242.

Duval:1999:PFT

[DR99] Art M. Duval and Victor Reiner. Perron–Frobenius type results and discrete versions of nodal domain theorems. Linear Algebra and its Applications, 294(1–3): 259–268, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6436.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6436.

Drazin:1992:EDG

[Dra92] Michael P. Drazin. Extremal definitions of generalized inverses. Linear Algebra and its Applications, 165:185–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drmac:1996:RRB

[Drm96] Zlatko Drmač. On relative residual bounds for the eigenvalues of a Hermitian matrix. Linear Algebra and its Applications, 244(1–3):155–163, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=244&issue=1-
                                        3&aid=9400221.

denHertog:1991:PRV

[dRT91] D. den Hertog, C. Roos, and T. Terlaky. A potentialreduction variant of Renegar’s short-step path-following method for linear programming. Linear Algebra and its Applications, 152:43–68, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

DenHertog:1993:LCP

[DRT93] D. Den Hertog, C. Roos, and T. Terlaky. The linear complementarity problems, sufficient matrices, and the crisscross method. Linear Algebra and its Applications, 187:1–??, July 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1990:SGC

[Dru90] S. W. Drury. Several graded contractions on Hilbert space. Linear Algebra and its Applications, 142:23–41, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1992:CQM

[Dru92a] S. W. Drury. A counterexample to a question of Merikoski and Virtanen on the compounds of unitary matrices. Linear Algebra and its Applications, 168:251–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1992:SFE

[Dru92b] S. W. Drury. On symmetric functions of the eigenvalues of the sum of two Hermitian matrices. Linear Algebra and its Applications, 176: 211–??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1994:BDC

[Dru94a]  S. W. Drury.  A bound

for the determinant of certain Hadamard products and for the determinant of the sum of two normal matrices. Linear Algebra and its Applications, 199:329–338, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1994:TWC

[Dru94b] S. W. Drury. On a theorem of Wielandt and the compounds of unitary matrices. Linear Algebra and its Applications, 199:391–412, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:1996:DCH

[Dru96]  S. W. Drury.  The de-

terminantal conjecture and Hadamard type inequalities. Linear Algebra and its Applications, 246(1–3):279–297, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400361.

Drury:1998:TCM

[Dru98] S. W. Drury. Trace class multipliers and spectral variation of normal matrices. Linear Algebra and its Applications, 280(2–3):217–227, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6134.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6134.

Drury:1999:NRP

[Dru99a] S. W. Drury. The numerical range of products of normal matrices. Linear Algebra and its Applications, 288 (1–3):283–292, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6330.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6330.

Drury:1999:OHR

[Dru99b] S. W. Drury. OMC for Householder reflections. Linear Algebra and its Applications, 298(1–3):159–169, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/27/article.
pdf.

Drury:1999:PLT

[Dru99c] S. W. Drury. Permanents and Lorentzian time-semidefinite matrices. Linear Algebra and its Applications, 294(1– 3):155–168, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6405.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6405.

Dauxois:1994:TPS

[DRVP94] J. Dauxois, Y. Romain, and S. Viguier-Pla. Tensor products and statistics. Linear Algebra and its Applications, 210:59–88, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

delOlmo:1990:MAS

[dRWZ90] M. A. del Olmo, M. A. Rodríguez, P. Winternitz, and H. Zassenhaus. Maximal Abelian subalgebras of pseudounitary Lie algebras. Linear Algebra and its Applications, 135:79–151, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:1991:AML

[DS91] Biswa Nath Datta and Youcef Saad. Arnoldi methods for large Sylvester-like observer problems, and an associated algorithm for partial spectrum assignment. Linear Algebra and its Applications, 154/156: 225–244, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dale:1992:NET

[DS92] Wilbur N. Dale and Malcolm C. Smith. A note on eventually time-invariant systems. Linear Algebra and its Applications, 171:225–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:1994:RCC

[DS94a] Shubhabrata Das and Pranab Kumar Sen. Restricted canonical correlations. Linear Algebra and its Applications, 210:29– 47, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeDoes:1994:CSP

[DS94b] J. De Does and J. M. Schumacher. Continuity of singular perturbations in the graph topology. Linear Algebra and its Applications, 205/206: 1121–1143, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dorfler:1994:CUN

[DS94c] P. Dörfler and G. Schmeisser.

Construction of unitary and normal companion matrices. Linear Algebra and its Applications, 202:193–220, April 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deGraaf:1995:CHS

[dS95] Maurits de Graaf and Alexander Schrijver. Characterizing homotopy of systems of curves on a compact surface by crossing numbers. Linear Algebra and its Applications, 226/228 (1–3):519–528, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500161.

daGracaMarques:1997:CPM

[dS97] M. da Graça Marques and Fernando C. Silva. The characteristic polynomial of a matrix with prescribed off-diagonal blocks. Linear Algebra and its Applications, 250(1–3):21– 29, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500393.

Didenko:1998:EAM

[DS98] V. D. Didenko and B. Silbermann. Extension of C^{∗}- algebras and Moore–Penrose stability of sequences of additive operators. Linear Algebra and its Applications, 275/276(1–3):121–140, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6023.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6023. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Dong:1998:MCM

[DSG98] Xue-Dong Dong, Cheong Boon Soh, and Erry Gunawan. Matrix characterization of MDS linear codes over modules. Linear Algebra and its Applications, 277(1–3):57–61, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6063.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6063.

Day:1998:SHM

[DST98] Jane Day, Wasin So, and Robert C. Thompson. The spectrum of a Hermitian matrix sum. Linear Algebra and its Applications, 280(2–3): 289–332, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6140.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6140.

Dol:1996:MAS

[DSW96] Wietse Dol, Ton Steerneman, and Tom Wansbeek. Matrix algebra and sampling theory: The case of the Horvitz–Thompson estimator. Linear Algebra and its Applications, 237/238(1–3):225– 238, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500781. Special

issue honoring Calyampudi Radhakrishna Rao.

daGracaMarques:1997:NIPa

[dSZ97] M. da Graça Marques, Fernando C. Silva, and Yu Lin Zhang. The number of invariant polynomials of a matrix with prescribed complementary principal submatrices. Linear Algebra and its Applications, 251(1–3):167– 179, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500688.

Duchamp:1995:CSS

[Duc95] Alain Duchamp. Circuit separation for symmetric matroids. Linear Algebra and its Applications, 231(1–3):87– 103, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400011.

Duffner:1994:LTP

[Duf94] M. Antónia Duffner. Linear transformations that preserve immanants. Linear Algebra and its Applications, 197/198: 567–588, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Duran:1995:OMP

[DV95] A. J. Durán and W. Van Assche. Orthogonal matrix polynomials and higherorder recurrence relations. Linear Algebra and its Applications, 219(1–3):261–280, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300218.

DePierro:1998:ROL

[DW98] Alvaro R. De Pierro and Musheng Wei. Reverse order law for reflexive generalized inverses of products of matrices. Linear Algebra and its Applications, 277(1– 3):299–311, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6072.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6072.

DiFiore:1995:MDU

[DZ95] Carmine Di Fiore and Paolo Zellini. Matrix decompositions using displacement rank and classes of commutative matrix algebras. Linear Algebra and its Applications, 229(1–3):49–99, October 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9300347.

DiFiore:1998:MDD

[DZ98] Carmine Di Fiore and Paolo Zellini. Matrix displacement decompositions and applications to Toeplitz linear systems. Linear Algebra and its Applications, 268(1–3):197– 225, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700044.

Ebiefung:1995:ETM

[Ebi95] Aniekan A. Ebiefung. Existence theory and Q-matrix characterization for the generalized linear complementarity problem. Linear Algebra and its Applications, 223/224(1–3):155–169, July ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500091. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Edelman:1991:DMS

[Ede91] Alan Edelman. The distribution and moments of the smallest eigenvalue of a random matrix of Wishart type. Linear Algebra and its Applications, 159:55–80, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1995:SVD

[EF95] L. Elsner and S. Friedland. Singular values, doubly stochastic matrices, and applications. Linear Algebra and its Applications, 220(1–3): 161–169, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9500111. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Elsner:1997:NCC

[EF97] L. Elsner and S. Friedland. Norm conditions for convergence of infinite products. Linear Algebra and its Applications, 250(1–3):133– 142, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500423.

Eisinberg:1998:VMC

[EFP98] A. Eisinberg, G. Franzé, and P. Pugliese. Vandermonde matrices on Chebyshev points. Linear Algebra and its Applications, 283 (1–3):205–219, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6211.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6211.

Evard:1990:SCA

[EG90] Jean-Claude Evard and Juan-Miguel Gracia. On similarities of class C^{p} and applications to matrix differential equations. Linear Algebra and its Applications, 137/138:363– 386, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estes:1993:MPI

[EG93] Dennis R. Estes and Robert M.

Guralnick. Minimal polynomials of integral symmetric matrices. Linear Algebra and its Applications, 192:83–99, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Ern:1997:PIA

[EG97] Alexandre Ern and Vincent Giovangigli. Projected iterative algorithms with application to multicomponent transport. Linear Algebra and its Applications, 250(1–3):289– 315, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500502.

Ellers:1999:GST

[EG99] Erich W. Ellers and Nikolai Gordeev. A generalization of Sourour’s theorem. Linear Algebra and its Applications, 286(1–3):187–196, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        286/1-3/6278.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=286&issue=
                                        1-3&aid=6278.

Egecioglu:1995:PLT

[Eğe95] Ömer Eğecioğlu. Parallelogramlaw-type identities. Linear Algebra and its Applications, 225(1):1–12, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=225&issue=1&
                                        aid=9300301.

Eggermont:1990:MIA

[Egg90] P. P. B. Eggermont. Multiplicative iterative algorithms for convex programming. Linear Algebra and its Applications, 130:25–42, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Ellis:1996:CBT

[EGL96] Robert L. Ellis, Israel Gohberg, and David C. Lay. On a class of block Toeplitz matrices. Linear Algebra and its Applications, 241/243(1–3):225–245, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500260. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Enflo:1995:ENL

[EGLL95] P. Enflo, V. I. Gurariĭ, V. Lomonosov, and Yu. I. Lyubich. Exponential numbers of linear operators in normed spaces. Linear Algebra and its Applications, 219(1–3): 225–260, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300217.

Elsner:1993:PIT

[EH93] Ludwig Elsner and Chunyang He. Perturbation and interlace theorems for the unitary eigenvalue problem. Linear Algebra and its Applications, 188/189:207–229, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1998:HFP

[EH98] Ludwig Elsner and Daniel Hershkowitz. Hadamard functions preserving nonnegative H-matrices. Linear Algebra and its Applications, 279(1– 3):13–19, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6094.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6094.

Eschenbach:1997:GUE

[EHJL97] Carolyn A. Eschenbach, Frank J. Hall, Charles R. Johnson, and Zhongshan Li. The graphs of the unambiguous entries in the product of two (+, −) sign pattern matrices. Linear Algebra and its Applications, 260(1–3):95– 118, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1997&volume=260&issue=1-
3&aid=9600293.

Eschenbach:1994:SPM

[EHL94] Carolyn A. Eschenbach, Frank J.

Hall, and Zhongshan Li. Sign pattern matrices and generalized inverses. Linear Algebra and its Applications, 211:53– 66, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Eschenbach:1997:SSP

[EHL97] Carolyn A. Eschenbach, Frank J.

Hall, and Zhongshan Li. Some sign patterns that allow a real inverse pair B and B^{−1}. Linear Algebra and its Applications, 252(1–3):299–321, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9600672.

Elsner:1990:FIS

[EHP90] Ludwig Elsner, Daniel Hershkowitz, and Allan Pinkus. Functional inequalities for spectral radii of nonnegative matrices. Linear Algebra and its Applications, 129:103–130, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1993:SRC

[EHS93] Ludwig Elsner, Daniel Hershkowitz, and Hans Schneider. Spectral radii of certain iteration matrices and cycle means of digraphs. Linear Algebra and its Applications, 192:61–81, October 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Elsner:1997:CFN

[EI97] L. Elsner and Kh. D. Ikramov. On a condensed form for normal matrices under finite sequences of elementary unitary similarities. Linear Algebra and its Applications, 254(1–3): 79–98, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600526. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Elsner:1998:NMU

[EI98] L. Elsner and Kh. D. D. Ikramov. Normal matrices: An update. Linear Algebra and its Applications, 285 (1–3):291–303, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6272.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6272.

Eiermann:1993:FVI

[Eie93] Michael Eiermann. Fields of values and iterative methods. Linear Algebra and its Applications, 180:167–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eijkhout:1991:API

[Eij91] Victor Eijkhout. Analysis of parallel incomplete point factorizations. Linear Algebra and its Applications, 154/156: 723–740, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eschenbach:1993:SPR

[EJ93] Carolyn A. Eschenbach and Charles R. Johnson. Sign patterns that require repeated eigenvalues. Linear Algebra and its Applications, 190:169– 179, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ebiefung:1993:GM

[EK93] A. A. Ebiefung and M. M.

Kostreva. Generalized P_{0}- and Z-matrices. Linear Algebra and its Applications, 195:165– 179, December 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Essannouni:1999:THP

[EK99] H. Essannouni and A. Kaidi. Le théorème de Hua pour les algèbres artiniennes simples. Linear Algebra and its Applications, 297(1–3):9–22, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/104/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/104/17/18/article.
pdf.

Elsner:1990:CAP

[EKN90] L. Elsner, I. Koltracht, and M. Neumann. On the convergence of asynchronous paracontractions with application to tomographic reconstruction from incomplete data. Linear Algebra and its Applications, 130:65–82, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

ElKadiri:1998:OPC

[El 98] Mohamed El Kadiri. Opérateurs positifs sur certains espaces de formes quadratiques. (French) [Positive operators on certain quadratic form spaces]. Linear Algebra and its Applications, 283(1–3):273–288, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6222.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6222.

Ellis:1992:FFR

[EL92] Robert L. Ellis and David C.

Lay. Factorization of finite rank Hankel and Toeplitz matrices. Linear Algebra and its Applications, 173:19–??, August 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eschenbach:1997:HMN

[EL97] Carolyn A. Eschenbach and Zhongshan Li. How many negative entries can A^{2} have? Linear Algebra and its Applications, 254(1–3):99–117, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600510. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Eschenbach:1999:PNS

[EL99] Carolyn A. Eschenbach and Zhongshan Li. Potentially nilpotent sign pattern matrices. Linear Algebra and its Applications, 299(1–3):81–99, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/112/21/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/112/21/23/article.
                                        pdf.

Ellers:1994:PTO

[Ell94] Erich W. Ellers. Products of transvections in one conjugacy class in the symplectic group over GF(3). Linear Algebra and its Applications, 202:1–23, April 15, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1993:NHW

[Els93] L. Elsner. A note on the Hoffman-Wielandt theorem. Linear Algebra and its Applications, 182:235–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1994:PTJ

[Els94] Ludwig Elsner. Perturbation theorems for the joint spectrum of commuting matrices: A conservative approach. Linear Algebra and its Applications, 208/209:83–95, 1994.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1995:GSR

[Els95]  L. Elsner.  The gener-

alized spectral-radius theorem: An analytic-geometric proof. Linear Algebra and its Applications, 220(1–3):151– 159, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9300320. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Elsner:1997:BDP

[Els97] L. Elsner. Bounds for determinants of perturbed Mmatrices. Linear Algebra and its Applications, 257(1–3): 283–288, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600174.

Ellers:1991:PTS

[EM91] Erich W. Ellers and Jerry Malzan. Products of λtransvections in Sp(2n, q).

Linear Algebra and its Applications, 146:121–132, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:1993:MSR

[EN93] L. Elsner and M. Neumann. Monotonic sequences and rates of convergence of asynchronized iterative methods. Linear Algebra and its Applications, 180:17–33, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Engwerda:1993:EPD

[Eng93] Jacob C. Engwerda. On the existence of a positive definite solution of the matrix equation X + A^{T}X^{−1}A = I. Linear Algebra and its Applications, 194:91–108, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Englis:1995:TOB

[Eng95] Miroslav Engliš. Toeplitz operators and the Berezin transform on H^{2}. Linear Algebra and its Applications, 223/224(1–3):171– 204, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400056. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Elsner:1998:OBL

[ENN98] L. Elsner, R. Nabben, and M. Neumann. Orthogonal bases that lead to symmetric nonnegative matrices. Linear Algebra and its Applications, 271(1–3):323– 343, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700302.

Elsner:1991:ENP

[ENV91] L. Elsner, M. Neumann, and B. Vemmer. The effect of the number of processors on the convergence of the parallel block Jacobi method. Linear Algebra and its Applications, 154/156:311–330, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eaves:1990:RPP

[ER90] B. Curtis Eaves and Uriel G.

Rothblum. Relationships of properties of piecewise affine maps over ordered fields. Linear Algebra and its Applications, 132:1–63, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Engwerda:1993:NSC

[ERR93] Jacob C. Engwerda, André C. M. Ran, and Arie L. Rijkeboer. Necessary and sufficient conditions for the existence of a positive definite solution of the matrix equation X + A^{∗}X^{−1}A = Q. Linear Algebra and its Applications, 186:255–275, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eaves:1995:PFT

[ERS95] B. Curtis Eaves, Uriel G. Rothblum, and Hans Schneider. Perron–Frobenius theory over real closed fields and fractional power series expansions. Linear Algebra and its Applications, 220(1–3):123– 150, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400053. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Erxiong:1990:AFG

[Erx90] Jiang Erxiong. An algorithm for finding generalized eigenpairs of a symmetric definite matrix pencil. Linear Algebra and its Applications, 132:65– 91, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eschenbach:1993:ISP

[Esc93] Carolyn Eschenbach. Idempotence for sign-pattern matrices. Linear Algebra and its Applications, 180:153–??, February 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estes:1991:FHO

[Est91] Dennis R. Estes. Factorization in hereditary orders. Linear Algebra and its Applications, 157:161–164, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Emre:1990:TMR

[ETS90] Erol Emre, Heng Ming Tai, and Jin H. Seo. Transfer matrices, realization, and control of continuous-time linear time-varying systems via polynomial fractional representations. Linear Algebra and its Applications, 141:79–104, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Evard:1993:CSP

[EU93] Jean-Claude Evard and Frank Uhlig. Comparison of the sets of polynomials p and q such that X = p(A) and Y = q(B) are solutions of the matrix equations f(X) = A and f(Y) = B. Linear Algebra and its Applications, 185:41– 47, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eiermann:1993:OOB

[EV93] M. Eiermann and R. S. Varga. Is the optimal omega best for the SOR iteration method? Linear Algebra and its Applications, 182:257–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Evard:1995:ICP

[Eva95] Jean-Claude Evard. Invariance and commutativity properties of some classes of solutions of the matrix differential equation X(t)X’(t) = X’(t)X(t). Linear Algebra and its Applications, 218(1–3):89– 102, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300160.

Elsner:1999:PMM

[EvdD99] Ludwig Elsner and P. van den Driessche. On the power method in max algebra. Linear Algebra and its Applications, 302-303(1–3):17–32, December 1, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/27/article.
pdf.

Eveson:1996:EDO

[Eve96] S. P. Eveson. The eigenvalue distribution of oscillatory and strictly sign-regular matrices. Linear Algebra and its Applications, 246(1–3):17– 21, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400303.

Falikman:1997:IPH

[Fal97] Dmitry Falikman. Inequalities for permanents of Hermitian matrices. Linear Algebra and its Applications, 263(1–3):63– 74, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600508.

Fan:1993:QCL

[Fan93]  Michael K. H. Fan.  A

quadratically convergent local algorithm on minimizing the largest eigenvalue of a symmetric matrix. Linear Algebra and its Applications, 188/189: 231–??, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farid:1992:SVT

[Far92] F. O. Farid. The spectral variation for two matrices with spectra on two intersecting lines. Linear Algebra and its Applications, 177:251–??, December 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Faria:1995:MIR

[Far95a] Isabel Faria. Multiplicity of integer roots of polynomials of graphs. Linear Algebra and its Applications, 229(1–3):15– 35, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=229&issue=1-
                                        3&aid=9300337.

Farid:1995:CID

[Far95b] F. O. Farid. Criteria for invertibility of diagonally dominant matrices. Linear Algebra and its Applications, 215(1–3): 63–93, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=215&issue=1-
                                        3&aid=9300072.

Farebrother:1996:SES

[Far96] Richard William Farebrother.

Some early statistical contributions to the theory and practice of linear algebra. Linear Algebra and its Applications, 237/238(1–3):205– 224, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500600. Special

issue honoring Calyampudi Radhakrishna Rao.

Farebrother:1997:CAC

[Far97] Richard William Farebrother.

A. C. Aitken and the consolidation of matrix theory. Linear Algebra and its Applications, 264(1–3):3–12, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600398.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600398.

Farid:1998:TGG

[Far98] F. O. Farid. Topics on a generalization of Gershgorin’s theorem. Linear Algebra and its Applications, 268(1–3):91– 116, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700030.

Farebrother:1999:CSE

[Far99] Richard William Farebrother. A class of statistical estimators related to principal components. Linear Algebra and its Applications, 289(1– 3):121–126, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6006.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6006.

Faulkenberry:1992:CKN

[Fau92] Richard E. Faulkenberry. The continuous kernel of a nonsquare rational matrix function. Linear Algebra and its Applications, 172:85–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Faybusovich:1992:RGH

[Fay92] Leonid Faybusovich. Reduction to generalized Hessenberg form and inverse spectral problems. Linear Algebra and its Applications, 165:

71–??, March 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Faybusovich:1994:RFT

[Fay94] Leonid Faybusovich. Rational functions, Toda flows, and LR-like algorithms. Linear Algebra and its Applications, 203/204:359–383, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fierro:1996:PTO

[FB96] Ricardo D. Fierro and James R.

Bunch. Perturbation theory for orthogonal projection methods with applications to least squares and total least squares. Linear Algebra and its Applications, 234(1–3):71– 96, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400209.

Feig:1997:ARD

[FBO97] Ephraim Feig and Michael Ben-Or. On algebras related to the discrete cosine transform. Linear Algebra and its Applications, 266(1–3):81– 106, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=266&issue=1-
                                        3&aid=9600634.

Chuan:1991:MOC

[fC91] Wai fong Chuan. Matrices originating from certain difference equations. Linear Algebra and its Applications, 148: 27–42, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fortiana:1997:FMD

[FC97] J. Fortiana and C. M. Cuadras. A family of matrices, the discretized Brownian bridge, and distancebased regression. Linear Algebra and its Applications, 264(1–3):173–188, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9700051.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9700051.

Fragoso:1992:MSC

[Fd92] Marcelo D. Fragoso and Carlos E. de Souza. Maximal solution of a certain class of periodic Riccati differential equations. Linear Algebra and its Applications, 169:61–??, May 1992. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Flick-DOrnano:1999:SMC

[FD99] J.-C. Flick-D’Ornano. Similarity and matrices of constant rank. Linear Algebra and its Applications, 295(1– 3):145–148, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6437.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6437.

Feintuch:1994:GMR

[Fei94a] Avraham Feintuch. Gap metric robustness for linear timevarying systems. Linear Algebra and its Applications, 208/209:209–221, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feintuch:1994:SGR

[Fei94b] Avraham Feintuch. Strong graph representations for linear time-varying systems. Linear Algebra and its Applications, 203/204:385–399, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:1999:CNE

[Fer99] Rosário Fernandes. The covering number of elements of a matroid and associated transformations. Linear Algebra and its Applications, 298(1–3):51–71, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/21/article.
pdf.

Feinsilver:1996:SSK

[FF96] Philip Feinsilver and Robert Fitzgerald. The spectrum of symmetric Krawtchouk matrices. Linear Algebra and its Applications, 235(1–3):121– 139, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400123.

Fujii:1991:OFA

[FFK91] Masatoshi Fujii, Takayuki Furuta, and Eizaburo Kamei. Operator functions associated with Furuta’s inequality. Linear Algebra and its Applications, 149:91–96, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:1993:SSP

[FFK93a] Bernd Fritzsche, Stefan Fuchs, and Bernd Kirstein. Schur-Sequence parametrizations of Potapov-Normalized full-rank j_{pq}-elementary factors. Linear Algebra and its Applications, 191:107–149, September 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:1993:FIA

[FFK93b] Masatoshi Fujii, Takayuki Furuta, and Eizaburo Kamei. Furuta’s inequality and its application to Ando’s thorem. Linear Algebra and its Applications, 179:161–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:1996:RMN

[FFK96] Bernd Fritzsche, Stefan H.

Fuchs, and Bernd Kirstein. On resolvent matrices for nondegenerate matricial Schur problems. Linear Algebra and its Applications, 235(1–3):191– 220, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400138.

Foias:1990:CME

[FFT90] Ciprian Foias, Arthur Frazho, and Allen Tannenbaum. On certain minimal entropy extensions appearing in dilation theory. I. Linear Algebra and its Applications, 137/138:213– 238, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Foias:1994:CSI

[FFT94] Ciprian Foias, Arthur Frazho, and Allen Tannenbaum. On combined H^{∞}-H^{2} suboptimal interpolants. Linear Algebra and its Applications, 203/204: 443–469, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiol:1999:ATP

[FG99] M. A. Fiol and E. Garriga. On the algebraic theory of pseudo-distance-regularity around a set. Linear Algebra and its Applications, 298(1–3):115–141, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/105/17/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/25/article.
                                        pdf.

Fleming:1998:SSM

[FGL^{+}98] R. Fleming, G. Grossman, T. Lenker, S. Narayan, and S.-C. Ong. On Schur Dstable matrices. Linear Algebra and its Applications, 279(1–3):39–50, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/

tree/store/laa/sub/1998/
279/1-3/6179.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6179.

Ferrer:1994:DFS

[FGP94] J. Ferrer, Ma. I. García, and F. Puerta. Differentiable families of subspaces. Linear Algebra and its Applications, 199: 229–252, March 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferrer:1996:SSQ

[FGP96] J. Ferrer and Ma. I. García-Planas. Structural stability of quadruples of matrices. Linear Algebra and its Applications, 241/243(1–3):279– 290, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500523. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Ferrer:1997:BLF

[FGP97] J. Ferrer, Ma. I. García, and F. Puerta. Brunowsky local form of a holomorphic family of pairs of matrices. Linear Algebra and its Applications, 253(1–3):175– 198, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=253&issue=1-
                                        3&aid=9500742.

Freund:1993:RSN

[FH93] Roland W. Freund and Thomas Huckle. A restricted signature normal form for Hermitian matrices, quasispectral decompositions, and applications. Linear Algebra and its Applications, 184: 165–??, April 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feldmann:1994:UPM

[FH94a] Sven Feldmann and George Heinig. Uniqueness properties of minimal partial realizations. Linear Algebra and its Applications, 203/204:401– 427, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujioka:1994:SCA

[FH94b] Hisaya Fujioka and Shinji Hara. State covariance assignment problem with measurement noise: A unified approach based on a symmetric matrix equation. Linear Algebra and its Applications, 203/204:579–605, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feldmann:1996:VFC

[FH96a] Sven Feldmann and Georg Heinig. Vandermonde factorization and canonical representations of block Hankel matrices. Linear Algebra and its Applications, 241/243(1–3):247–278, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500501. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Fischer:1996:CNC

[FH96b] Ismor Fischer and Suk-Geun Hwang. Certain nonbarycentric cohesive matrices. Linear Algebra and its Applications, 239(1–3):185–200, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400191.

First:1990:MCP

[FHP90] Zvi First, Steven T. Hackman, and Ury Passy. Matrix criteria for the pseudo-P -convexity of a quadratic form. Linear Algebra and its Applications, 136: 235–255, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Finck:1993:IFF

[FHR93] Tilo Finck, Georg Heinig, and Karla Rost. An inversion formula and fast algorithms for Cauchy–Vandermonde matrices. Linear Algebra and its Applications, 183:179–??, April 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedberg:1992:HMS

[FI92] Stephen H. Friedberg and Arnold J. Insel. Hyponormal 2 × 2 matrices are subnormal. Linear Algebra and its Applications, 175:31–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedberg:1995:CSM

[FI95] Stephen H. Friedberg and Arnold J. Insel. Characterizations of subnormal matrices. Linear Algebra and its Applications, 231(1–3):1–13, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1995&volume=231&issue=1-
3&aid=9400006.

Freiling:1999:TVD

[FI99] Gerhard Freiling and Vlad Ionescu. Time-varying discrete Riccati equation: some monotonicity results. Linear Algebra and its Applications, 286(1–3):135–148, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6277.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6277.

Fiedler:1990:EPC

[Fie90a] Miroslav Fiedler. Expressing a polynomial as the characteristic polynomial of a symmetric matrix. Linear Algebra and its Applications, 141: 265–270, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1990:PRS

[Fie90b] Miroslav Fiedler. Pencils of real symmetric matrices and real algebraic curves. Linear Algebra and its Applications, 141:53–60, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1993:SRM

[Fie93] Miroslav Fiedler. Structure ranks of matrices. Linear Algebra and its Applications, 179:119–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1994:EMZ

[Fie94] Miroslav Fiedler. Elliptic matrices with zero diagonal. Linear Algebra and its Applications, 197/198:337–347, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Fiedler:1995:ENE

[Fie95a] Miroslav Fiedler. An estimate for the nonstochastic eigenvalues of doubly stochastic matrices. Linear Algebra and its Applications, 214(1–3):133– 143, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=214&issue=1-
                                        3&aid=9300061.

Fiedler:1995:NRM

[Fie95b] Miroslav Fiedler. Numerical range of matrices and Levinger’s theorem. Linear Algebra and its Applications, 220(1–3):171–180, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400073. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Fiedler:1996:NRR

[Fie96] Miroslav Fiedler. A note on the row-rhomboidal form of a matrix. Linear Algebra and its Applications, 232(1– 3):149–154, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1997:SIP

[Fie97] Miroslav Fiedler. Some inverse problems for acyclic matrices. Linear Algebra and its Applications, 253(1–3):113– 123, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500692.

Fiedler:1998:SCS

[Fie98] Miroslav Fiedler. Some characterizations of symmetric inverse M-matrices. Linear Algebra and its Applications, 275/276(1–3):179–187, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6027.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6027. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Fiol:1999:EIW

[Fio99] M. A. Fiol. Eigenvalue interlacing and weight parameters of graphs. Linear Algebra and its Applications, 290(1–3):275–301, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6341.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6341.

Firoozi:1990:TIC

[Fir90] Fathali K. Firoozi. A transformation of the inequalityconstrained linear model. Linear Algebra and its Applications, 133:153–163, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fisher:1995:CPS

[Fis95] David C. Fisher. Characteristic polynomials of Straffin digraphs. Linear Algebra and its Applications, 218(1–3): 103–111, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300164.

Fallat:1999:SDS

[FJ99] Shaun M. Fallat and Charles R.

Johnson. Sub-direct sums and positivity classes of matrices. Linear Algebra and its Applications, 288(1–3): 149–173, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6304.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6304.

Freiling:1996:GRD

[FJAK96] G. Freiling, G. Jank, and H. Abou-Kandil. Generalized Riccati difference and differential equations. Linear Algebra and its Applications, 241/243(1–3):291–303, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500587. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Fleischmann:1998:NRS

[FJK98] P. Fleischmann, I. Janiszczak, and R. Knörr. The number of regular semisimple classes of special linear and unitary groups. Linear Algebra and its Applications, 274(1–3):17– 26, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700224.

Fallat:1998:ELN

[FJSv98] Shaun M. Fallat, Charles R. Johnson, Ronald L. Smith, and P. van den Driessche. Eigenvalue location for nonnegative and Z-matrices. Linear Algebra and its Applications, 277(1–3):187–198, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/

277/1-3/6081.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6081.

Fritzsche:1993:IPP

[FK93] Bernd Fritzsche and Bernd Kirstein. Inverse problems for positive Hermitian block Toeplitz matrices and nondegenerate Schur sequences. Linear Algebra and its Applications, 179:237–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1994:MSR

[FK94] S. Friedland and M. Katz. On the maximal spectral radius of even tournament matrices and the spectrum of almost skew-symmetric compact operators. Linear Algebra and its Applications, 208/209:455– 469, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:1995:CSP

[FKK95] Bernd Fritzsche, Bernd Kirstein, and Vera Krug. Carathéodory sequence parametrizations of Potapov-Normalized full-rank J_{q}-elementary factors. Linear Algebra and its Applications, 214(1–3):145–185, January 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300062.

Fritzsche:1997:ANE

[FKK97] Bernd Fritzsche, Bernd Kirstein, and Vera Krug. On Arovnormalized j_{qq}−J_{q}-elementary factors. Linear Algebra and its Applications, 267(1–3):11– 40, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9600632.

Fritzsche:1998:STR

[FKK98] Bernd Fritzsche, Bernd Kirstein, and Vera Krug. On several types of resolvent matrices of nondegenerate matricial Carathéodory problems. Linear Algebra and its Applications, 281(1–3): 137–170, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6155.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6155.

Farid:1991:SPD

[FL91] F. O. Farid and P. Lancaster.

Spectral properties of diagonally dominant infinite matrices. II. Linear Algebra and its Applications, 143:7–17, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferrante:1996:HSE

[FL96] Augusto Ferrante and Bernard C.

Levy. Hermitian solutions of the equation X = Q + NX^{−1}N. Linear Algebra and its Applications, 247(1–3):359–373, November 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500121.

Ferrante:1998:CFS

[FL98] Augusto Ferrante and Bernard C.

Levy. Canonical form of symplectic matrix pencils. Linear Algebra and its Applications, 274(1–3):259–300, April 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700347.

Friedland:1999:SSM

[FL99] S. Friedland and R. Loewy. Spaces of symmetric matrices containing a nonzero matrix of bounded rank. Linear Algebra and its Applications, 287(1–3):161–170, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6295.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6295.

Fleming:1990:ERT

[Fle90] Henry E. Fleming. Equivalence of regularization and truncated iteration in the solution of ill-posed image reconstruction problems. Linear Algebra and its Applications, 130:133–150, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Fleischmann:1993:CBA

[Fle93] Peter Fleischmann. Connections between the algorithms of Berlekamp and Niederreiter for factoring polynomials over F_{q}. Linear Algebra and its Applications, 192:101–108, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Fleischmann:1997:PIO

[Fle97] Peter Fleischmann. Polynomial identities for orbit numbers of general linear and unitary groups over finite fields. Linear Algebra and its Applications, 253(1–3):341– 362, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9600740.

Fliess:1994:IAT

[Fli94]  Michel Fliess.  Une in-

terprétation algébrique de la transformation de Laplace et des matrices de transfert. (French) [an algebraic interpretation of the Laplace transform and transfer matrices]. Linear Algebra and its Applications, 203/204:429–442, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferng:1997:NNL

[FLL97] William R. Ferng, Kun-Yi Lin, and Wen-Wei Lin. A novel nonsymmetric K -lanczos algorithm for the generalized nonsymmetric K -eigenvalue problems. Linear Algebra and its Applications, 252(1–3):81– 105, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=252&issue=1-
                                        3&aid=9600670.

Fuenzalida:1994:OWA

[FLM94] Ana Fuenzalida, Alicia Labra, and Cristian Mallol. Orthogonality in WBJ algebras. Linear Algebra and its Applications, 196:125–138, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Flournoy:1990:DMM

[Flo90] Nancy Flournoy. Dependency in multivariate Markov chains. Linear Algebra and its Applications, 127:85–106, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1992:CMC

[FM92] Miroslav Fiedler and Thomas L. Markham. A classification of matrices of class Z. Linear Algebra and its Applications, 173:115–124, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1993:CMP

[FM93a] Miroslav Fiedler and Thomas L. Markham. A characterization of the Moore–Penrose inverse.

Linear Algebra and its Applications, 179:129–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1993:QAM

[FM93b] Miroslav Fiedler and Thomas L.

Markham. Quasidirect addition of matrices and generalized inverses. Linear Algebra and its Applications, 191:165– 182, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Frommer:1993:LSD

[FM93c] A. Frommer and G. Mayer.

Linear systems with Ωdiagonally dominant matrices and related ones. Linear Algebra and its Applications, 186: 165–181, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fu:1993:EDE

[FM93d] Enzhong Fu and Thomas L.

Markham. On the eigenvalues and diagonal entries of a Hermitian matrix. Linear Algebra and its Applications, 179: 7–??, January 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1995:OHP

[FM95] Miroslav Fiedler and Thomas L.

Markham. An observation on the Hadamard product of Hermitian matrices.

Linear Algebra and its Applications, 215(1–3):179–182, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300087.

Fiedler:1996:SIH

[FM96a] Miroslav Fiedler and Thomas L. Markham. Some inequalities for the Hadamard product of matrices. Linear Algebra and its Applications, 246(1–3):13– 16, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400297.

Fiedler:1996:SRB

[FM96b] Miroslav Fiedler and Thomas L. Markham. Some results on the Bergstrom and Minkowski inequalities. Linear Algebra and its Applications, 232(1– 3):199–211, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1997:CCR

[FM97a] Miroslav Fiedler and Thomas L. Markham. Consecutivecolumn and -row properties of matrices and the Loewner– Neville factorization. Linear Algebra and its Applications, 266(1–3):243–259, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9786523.

Freund:1997:BQA

[FM97b] Roland W. Freund and Manish Malhotra. A block QMR algorithm for non-Hermitian linear systems with multiple right-hand sides. Linear Algebra and its Applications, 254(1–3):119–157, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600529. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Fassbender:1999:HSR

[FMMX99] Heike Faßbender, D. Steven Mackey, Niloufer Mackey, and Hongguo Xu. Hamiltonian square roots of skew-Hamiltonian matrices. Linear Algebra and its Applications, 287(1–3):125–159, January 15, 1999. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6254.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6254.

FitzGerald:1995:FPF

[FMP95] Carl H. FitzGerald, Charles A. Micchelli, and Allan Pinkus. Functions that preserve families of positive semidefinite matrices. Linear Algebra and its Applications, 221(1–3):83– 102, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300232.

Fuster:1996:PCE

[FMP96] R. Fuster, V. Migallón, and J. Penadés. Parallel chaotic extrapolated Jacobilike methods. Linear Algebra and its Applications, 247(1–3):237–250, November 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500109.

Fischer:1998:MDM

[FMS98] Klaus G. Fischer, Walter Morris, and Jay Shapiro. Mixed dominating matrices. Linear Algebra and its Applications, 270(1–3):191–214, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700239.

Forster:1993:DMP

[FN93] K.-H. Förster and B. Nagy.

Decomposable matrix polynomials. Linear Algebra and its Applications, 187:41–58, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:1995:MLE

[FN95a] Michael K. H. Fan and Batool Nekooie. On minimizing the largest eigenvalue of a symmetric matrix. Linear Algebra and its Applications, 214(1–3):225– 246, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300068.

Fujii:1995:RTH

[FN95b] Masatoshi Fujii and Ritsuo Nakamoto. Rota’s theorem and Heinz inequalities. Linear Algebra and its Applications, 214(1–3):271– 275, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300080.

Friedland:1997:SRE

[FN97] Shmuel Friedland and Reinhard Nabben. On the second real eigenvalue of nonnegative and Z-matrices. Linear Algebra and its Applications, 255(1–3):303–313, April 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600033.

Forster:1998:LSR

[FN98a] K.-H. Förster and B. Nagy. Local spectral radii and Collatz–Wielandt numbers of monic operator polynomials with nonnegative coefficients. Linear Algebra and its Applications, 268(1–3):41– 57, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9789323.

Forster:1998:SPR

[FN98b] K.-H. Förster and B. Nagy.

Spectral properties of rational matrix functions with nonnegative realizations. Linear Algebra and its Applications, 275/276(1–3):189–200, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6032.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6032. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Fuller:1991:URA

[FNW91] K. R. Fuller, W. K. Nicholson, and J. F. Watters. Universally reflexive algebras. Linear Algebra and its Applications, 157:195–201, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuller:1996:DSR

[FNW96] K. R. Fuller, W. K. Nicholson, and J. F. Watters. Direct sums of reflexive modules. Linear Algebra and its Applications, 239(1–3):201– 214, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=239&issue=1-
                                        3&aid=9400194.

Fonseca:1999:EFD

[Fon99] Amélia Fonseca. On the equality of families of decomposable symmetrized tensors. Linear Algebra and its Applications, 293(1–3):1–14, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        293/1-3/6356.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=293&issue=
                                        1-3&aid=6356.

Foregger:1990:END

[For90a] Thomas H. Foregger. Embedding nearly decomposable matrices into certain staircase matrices. Linear Algebra and its Applications, 135:153–165, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fornasini:1990:MC

[For90b] Ettore Fornasini. 2D Markov chains. Linear Algebra and its Applications, 140:101–127, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:1997:GTA

[FOV97] Shaun Fallat, D. D. Olesky, and P. Van Den Driessche. Graph theoretic aspects of maximizing the spectral radius of nonnegative matrices. Linear Algebra and its Applications, 253(1–3):61– 77, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500691.

Ferrer:1992:SNE

[FP92] Josep Ferrer and Ferran Puerta. Similarity of noneverywhere defined linear maps. Linear Algebra and its Applications, 168:27–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Freedman:1996:TWM

[FP96] Michael H. Freedman and William H. Press. Truncation of wavelet matrices: Edge effects and the reduction of topological control. Linear Algebra and its Applications, 234(1–3):1– 19, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400039.

Fiedler:1997:BAC

[FP97a] Miroslav Fiedler and Vlastimil Pták. Block analogies of comparison matrices. Linear Algebra and its Applications, 256(1–3):95– 108, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600004.

Fiedler:1997:NPD

[FP97b] Miroslav Fiedler and Vlastimil Pták. A new positive definite geometric mean of two positive definite matrices. Linear Algebra and its Applications, 251(1–3):1– 20, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500540.

Fiedler:1997:SMH

[FP97c] Miroslav Fiedler and Vlastimil Pták. Strong majorization for Hermitian matrices. Linear Algebra and its Applications, 255(1–3):281– 301, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600029.

Ferrante:1999:TLF

[FP99] Augusto Ferrante and Michele Pavon. The algebraic Riccati inequality: parametrization of solutions, tightest local frames and generalized feedback matrices. Linear Algebra and its Applications, 292(1–3):187–206, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6378.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6378.

Ferrer:1996:GCC

[FPP96] J. Ferrer, F. Puerta, and X. Puerta. Geometric characterization and classification of marked subspaces. Linear Algebra and its Applications, 235(1–3):15–34, March 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400107.

Ferrer:1998:DSS

[FPP98] J. Ferrer, F. Puerta, and X. Puerta. Differentiable structure of the set of controllable (A, B)^{t}-invariant subspaces. Linear Algebra and its Applications, 275/276(1– 3):161–177, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6057.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6057. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Fisher:1995:OSR

[FR95a] David C. Fisher and Richard B. Reeves. Optimal strategies for random tournament games. Linear Algebra and its Applications, 217(1–3):83– 85, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1995&volume=217&issue=1-
3&aid=9400212. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Fisher:1995:TGC

[FR95b] David C. Fisher and Jennifer Ryan. Tournament games and Condorcet voting. Linear Algebra and its Applications, 217(1–3):87– 100, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400066. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Franklin:1995:LSS

[Fra95] Joel Franklin. Least-squares solution of equations of motion under inconsistent constraints. Linear Algebra and its Applications, 222(1–3):9– 13, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300267.

Freund:1991:TES

[Fre91a] Robert M. Freund. Theoretical efficiency of a shiftedbarrier-function algorithm for linear programming. Linear Algebra and its Applications, 152:19–41, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Freund:1991:PPA

[Fre91b] Roland Freund. On polynomial preconditioning and asymptotic convergence factors for indefinite Hermitian matrices. Linear Algebra and its Applications, 154/156:259– 288, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1990:CSR

[Fri90a] Shmuel Friedland. Characterizations of the spectral radius of positive operators. Linear Algebra and its Applications, 134:93–105, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1990:NDI

[Fri90b] Shmuel Friedland. A note on a determinantal inequality. Linear Algebra and its Applications, 141:221–222, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1991:PMW

[Fri91a] Shmuel Friedland. Pairs of matrices which do not admit a complementary triangular form. Linear Algebra and its Applications, 150:119– 123, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Friedland:1991:QFG

[Fri91b] Shmuel Friedland. Quadratic forms and the graph isomorphism problem. Linear Algebra and its Applications, 150: 423–442, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Friedland:1991:RMS

[Fri91c] Shmuel Friedland. Revisiting matrix squaring. Linear Algebra and its Applications, 154/156:59–63, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1992:LBF

[Fri92]  Shmuel Friedland.  Lower

bounds for the first eigenvalue of certain M-matrices associated with graphs. Linear Algebra and its Applications, 172: 71–??, July 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:1993:ROS

[Fri93] Shmuel Friedland. Rational orthogonal similarity of rational symmetric matrices. Linear Algebra and its Applications, 192:109–114, October 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Friedland:1996:DGU

[Fri96] Shmuel Friedland. Discrete groups of unitary isometries and balls in hyperbolic manifolds. Linear Algebra and its Applications, 241/243(1–3):305–341, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500259. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Friedland:1997:ESF

[Fri97a] Shmuel Friedland. On the entropy of Z^{d} subshifts of finite type. Linear Algebra and its Applications, 252(1–3):199– 220, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500676.

Fripertinger:1997:CIL

[Fri97b]  Harald Fripertinger.  Cy-

cle indices of linear, affine, and projective groups. Linear Algebra and its Applications, 263(1–3):133–156, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600530.

Friedland:1998:CHD

[Fri98a]  Shmuel Friedland.  Com-

puting the Hausdorff dimension of subshifts using matrices. Linear Algebra and its Applications, 273(1–3):133– 167, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700371.

Friedman:1998:EBP

[Fri98b] Joel Friedman. Error bounds on the power method for determining the largest eigenvalue of a symmetric, positive definite matrix. Linear Algebra and its Applications, 280(2–3):199–216, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6141.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6141.

FernandezLopez:1997:JCF

[FRLM97] A. Fernández López, E. García Rus, M. Gómez Lozano, and M. Siles Molina. Jordan canonical form for finite rank elements in Jordan algebras. Linear Algebra and its Applications, 260(1–3): 151–167, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600298.

Farkas:1999:TMT

[FRS99] András Farkas, Pál Rózsa, and Etelka Stubnya. Transitive matrices and their applications. Linear Algebra and its Applications, 302-303(1–3):423–433, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/

50/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/50/article.
pdf.

Friedland:1994:PME

[FS94] Shmuel Friedland and Wasin So. On the product of matrix exponentials. Linear Algebra and its Applications, 196:193– 205, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:1999:ERS

[FS99a] Susana Furtado and Fernando C. Silva. Embedding a regular subpencil into a general linear pencil. Linear Algebra and its Applications, 295(1–3):61–72, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6397.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6397.

Furtado:1999:CPM

[FS99b] Susana Furtado and Fernando C. Silva. On the characteristic polynomial of matrices with prescribed rows. Linear Algebra and its Applications, 293(1–3):51–72, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        293/1-3/6369.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=293&issue=
                                        1-3&aid=6369.

Fuhrmann:1991:PAH

[Fuh91] P. A. Fuhrmann. A polynomial approach to Hankel norm and balanced approximations. Linear Algebra and its Applications, 146:133–220, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:1994:DTR

[Fuh94a] P. A. Fuhrmann. A duality theory for robust stabilization and model reduction. Linear Algebra and its Applications, 203/204:471–578, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:1994:SPI

[Fuh94b] P. A. Fuhrmann. On skew primeness of inner functions. Linear Algebra and its Applications, 208/209:539–551, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:1992:OMR

[Fuj92] Jun Ichi Fujii. Operator means and range inclusion. Linear Algebra and its Applications, 170:137–??, June 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fung:1996:LPC

[Fun96a]  Hon-Kwok Fung.  Linear

preservers of controllability and/or observability. Linear Algebra and its Applications, 246(1–3):335–360, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400364.

Fung:1996:NLT

[Fun96b] Hon-Kwok Fung. A note on linear transformations which leave controllable multiinput descriptor systems controllable. Linear Algebra and its Applications, 247(1–3):327– 336, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500119.

Furuta:1994:NAG

[Fur94] Takayuki Furuta. A note on the arithmetic-geometric mean inequality for every unitarily invariant matrix norm. Linear Algebra and its Applications, 208/209:223–228, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:1995:EFI

[Fur95] Takayuki Furuta. Extension of the Furuta inequality and Ando-Hiai log-majorization. Linear Algebra and its Applications, 219(1–3):139–155, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=219&issue=1-
                                        3&aid=9300203.

Furuta:1996:GKT

[Fur96] Takayuki Furuta. Generalizations of Kosaki trace inequalities and related trace inequalities on chaotic order. Linear Algebra and its Applications, 235(1–3):153– 161, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400132.

Fielder:1992:SSL

[FV92] Miroslav Fielder and Zdenek Vavrin. A subclass of symmetric Loewner matrices. Linear Algebra and its Applications, 170:47–??, June 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:1993:PCS

[FV93] Miroslav Fiedler and Zdeněk Vavřín. Polynomials compatible with a symmetric Loewner matrix. Linear Algebra and its Applications, 190:235–251, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fornasini:1996:SCS

[FV96] Ettore Fornasini and Maria Elena Valcher. On the spectral and combinatorial structure of 2D positive systems. Linear Algebra and its Applications, 245(1–3):223–258, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400242.

Fornasini:1997:DGS

[FV97] Ettore Fornasini and Maria Elena Valcher. Directed graphs, 2D state models, and characteristic polynomials of irreducible matrix pairs. Linear Algebra and its Applications, 263(1–3):275–310, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600540.

Fong:1996:BDO

[FW96] Che Kao Fong and Pei Yuan Wu. Band-diagonal operators. Linear Algebra and its Applications, 248(1–3):185–204, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500168.

Flaschka:1992:KAS

[FWWJL92] Ulrike Flaschka, Lin Wen-Wei, and Wu Jy-Liang. A KQZ algorithm for solving linearresponse eigenvalue equations. Linear Algebra and its Applications, 165:93–??, March 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujishige:1998:LAT

[FY98] Satoru Fujishige and Zaifu Yang. A lexicographic algebraic theorem and its applications. Linear Algebra and its Applications, 279(1– 3):75–91, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/

279/1-3/6100.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6100.

Freund:1993:FBP

[FZ93] Roland W. Freund and Hong Yuan Zha. Formally biorthogonal polynomials and a look-ahead Levinson algorithm for general Toeplitz systems. Linear Algebra and its Applications, 188/189:255– 303, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gabriel:1997:MEN

[Gab97]  Richard Gabriel.  Min-

imierung der Euklidischen Norm einer komplexen Matrix durch ebene Similaritäten. (German) [Minimizing the Euclidean norm of a complex matrix by plane similarities]. Linear Algebra and its Applications, 252(1–3):261–297, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500378.

Gader:1990:DOB

[Gad90] Paul D. Gader. Displacement operator based decompositions of matrices using circulants or other group matrices. Linear Algebra and its Applications, 139:111–131, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Galkowski:1991:PNS

[Gał91] Krzysztof Gałkowski. An a priori nonminimal state-space realization of n-D systems. Linear Algebra and its Applications, 151:185–198, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Galkowski:1996:MDM

[Gał96] Krzysztof Gałkowski. Matrix description of multivariable polynomials. Linear Algebra and its Applications, 234(1–3):209–226, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=234&issue=1-
                                        3&aid=9500104.

Gamage:1997:GVM

[Gam97] Jinadasa K. Gamage. Generalized p-values and the multivariate Behrens–Fisher problem. Linear Algebra and its Applications, 253(1–3):369– 377, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9600166.

Gattazzo:1991:PMG

[Gat91] Remo Gattazzo. Polynomial matrices with given determinant. Linear Algebra and its Applications, 144:107–120, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gay:1991:MMB

[Gay91] David M. Gay. Massive memory buys little speed for complete, in-core sparse Cholesky factorizations on some scalar computers. Linear Algebra and its Applications, 152:291– 314, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Gaubert:1999:SNS

[GB99] S. Gaubert and P. Butkovic.

Sign-non-singular matrices and matrices with unbalanced determinant in symmetrised semirings. Linear Algebra and its Applications, 301(1–3):195–201, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/38/article.
pdf.

Guu:1995:S

[GC95] Sy-Ming Guu and Richard W. Cottle. On a subclass of P_{0}. Linear Algebra and its Applications, 223/224(1–3):325– 335, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9300271. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Gracia:1999:NPM

[GdH99] Juan-Miguel Gracia and Inmaculada de Hoyos. Nearest pair with more nonconstant invariant factors and pseudospectrum. Linear Algebra and its Applications, 298(1–3):143–158, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/105/17/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/26/article.
                                        pdf.

Geromel:1998:LCS

[GdOH98] José C. Geromel, Maurício C. de Oliveira, and Liu Hsu. LMI characterization of structural and robust stability. Linear Algebra and its Applications, 285(1–3):69–80, December 1, 1998. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6240.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6240.

Geerrts:1993:ISI

[Gee93a] Ton Geerrts. Invariant subspaces and invertibility properties for singular systems: The general case. Linear Algebra and its Applications, 183: 61–??, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geerts:1993:SCC

[Gee93b]  Ton Geerts.  Solvability

conditions, consistency, and weak consistency for linear differential-algebraic equations and time-invariant singular systems: The general case. Linear Algebra and its Applications, 181:111–130, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geerts:1994:LQC

[Gee94] Ton Geerts. Linear-quadratic control with and without stability subject to general implicit continuous-time systems: Coordinate-free interpretations of the optimal costs in terms of dissipation inequality and linear matrix inequality: Existence and uniqueness of optimal controls and state trajectories. Linear Algebra and its Applications, 203/204:607–658, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geerts:1996:HOC

[Gee96] Tom Geerts. Higher-order continuous-time implicit systems: Consistency and weak consistency, impulse controllability, geometric concepts, and invertibility properties. Linear Algebra and its Applications, 244(1–3):203–253, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400224.

Geerts:1997:IHL

[Gee97] Ton Geerts. Infinite-horizon linear-quadratic control with end-point state penalty term: The discrete-time case. Linear Algebra and its Applications, 259(1–3):251–269, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600289.

Geelen:1999:MRM

[Gee99]  James F. Geelen.  Maxi-

mum rank matrix completion. Linear Algebra and its Applications, 288(1–3): 211–217, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6316.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6316.

Gelonch:1999:CES

[Gel99] J. Gelonch. The contragredient equivalence for several matrices: a set of invariants. Linear Algebra and its Applications, 291(1–3):37–49, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6318.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6318.

Gemignani:1996:CEA

[Gem96] Luca Gemignani. Computationally efficient applications of the Euclidean algorithm to zero location. Linear Algebra and its Applications, 249(1–3):79–91, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500267.

Gemignani:1997:FAG

[Gem97a] Luca Gemignani. A fast algorithm for generalized Hankel matrices arising in finitemoment problems. Linear Algebra and its Applications, 267(1–3):41–52, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9600633.

Gemignani:1997:SCB

[Gem97b] Luca Gemignani. Schur complements of Bezoutians and the inversion of block Hankel and block Toeplitz matrices. Linear Algebra and its Applications, 253(1–3):39– 59, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500690.

Gemignani:1998:HAC

[Gem98] Luca Gemignani. A hybrid approach to the computation of the inertia of a parametric family of Bézoutians with application to some stability problems for bivariate polynomials. Linear Algebra and its Applications, 283 (1–3):221–238, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6207.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6207.

Genin:1996:EAO

[Gen96a] Yves V. Genin. Euclid algorithm, orthogonal polynomials, and generalized Routh– Hurwitz algorithm. Linear Algebra and its Applications, 246(1–3):131–158, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400342.

Genin:1996:PNU

[Gen96b] Yves V. Genin. On polynomials nonnegative on the unit circle and related questions. Linear Algebra and its Applications, 247(1–3):317– 325, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500112.

Gerth:1998:SAT

[Ger98]  Frank Gerth, III.  Some

asymptotics for triangular matrices over finite fields. Linear Algebra and its Applications, 282(1–3):249–261, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6188.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6188.

Gracia:1996:MPG

[GES96] Juan-Miguel Gracia, Lourdes Ortiz De Elguea, and Maria-Jose Sodupe. Matrix pencil generated by a tensor product from two matrix pencils. Linear Algebra and its Applications, 249(1–3):289–310, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500364.

Gonzalez:1996:RBA

[GGM96a] S. González, J. C. Gutiérrez, and C. Martínez. On regular Bernstein algebras. Linear Algebra and its Applications, 241/243(1–3):389–400, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500589. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Gonzalez:1996:SOB

[GGM96b] S. González, J. C. Gutiérrez, and C. Martínez. Second order Bernstein algebras of dimension 4. Linear Algebra and its Applications, 233(1–3):243– 273, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400085.

Gonzalez:1991:SSI

[GH91] Raul Ernesto Gonzalez and D. J. Hartfiel. On the structure of the stochastic idempotent matrix space. Linear Algebra and its Applications, 145:141–158, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gingold:1992:GAT

[GH92] Harry Gingold and Po-Fang Hsieh. Globally analytic triangularization of a matrix function. Linear Algebra and its Applications, 169:75–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:1995:ESS

[GH95a] R. E. Gonzalez and D. J. Hartfiel. Erratum: “On the structure of the stochastic idempotent matrix space” [Linear Algebra Appl. 145 (1991), 141– 158, MR 92d:15017]. Linear Algebra and its Applications, 220(1–3):2, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutknecht:1995:SIF

[GH95b] Martin H. Gutknecht and Marlis Hochbruck. The stability of inversion formulas for Toeplitz matrices. Linear Algebra and its Applications, 223/224(1–3):307– 324, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/

browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400218. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Gheondea:1992:OSC

[Ghe92] Aurelian Gheondea. Onestep completions of Hermitian partial matrices with minimal negative signature. Linear Algebra and its Applications, 173:99–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Glunt:1999:PDC

[GHJT99] W. Glunt, T. L. Hayden, Charles R. Johnson, and P. Tarazaga. Positive definite completions and determinant maximization. Linear Algebra and its Applications, 288(1–3):1–10, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6317.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6317.

Gross:1999:POPb

[GHM99] Jürgen Groß, Jan Hauke, and Augustyn Markiewicz. Partial orderings, preorderings, and the polar decomposition of matrices. Linear Algebra and its Applications, 289(1– 3):161–168, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6226.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6226.

Galanis:1997:OCS

[GHN97] S. Galanis, A. Hadjidimos, and D. Noutsos. Optimal pcyclic SOR for complex spectra. Linear Algebra and its Applications, 263(1–3):233– 260, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600538.

Galanis:1999:YET

[GHN99] S. Galanis, A. Hadjidimos, and D. Noutsos. A Young– Eidson’s type algorithm for complex p-cyclic SOR spectra. Linear Algebra and its Applications, 286(1–3): 87–106, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6276.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6276.

Gao:1996:CGD

[GhW96] Yi-Ming Gao and Xiao hui Wang. Criteria for generalized diagonally dominant matrices and M-matrices. II. Linear Algebra and its Applications, 248(1–3):339–353, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500251.

George:1999:CIS

[GI99] Alan George and Khakim D.

Ikramov. Common invariant subspaces of two matrices. Linear Algebra and its Applications, 287(1–3): 171–179, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6261.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6261.

Gibson:1992:PMD

[Gib92] Peter M. Gibson. Permanental mates of doubly stochastic matrices. Linear Algebra and its Applications, 177: 197–??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gibson:1996:SNSa

[Gib96a]  Peter M. Gibson.  Sign-

nonsingular skew-symmetric matrices. Linear Algebra and its Applications, 240(1–3): 207–229, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=240&issue=1-
                                        3&aid=9400199.

Gibson:1996:SNSb

[Gib96b]  Peter M. Gibson.  Sign-

nonsingular skew-symmetric matrices with the most nonzero entries. Linear Algebra and its Applications, 244(1–3):97–109, September 1, 1996. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=244&issue=1-
                                        3&aid=9400215.

Gil:1993:EEM

[Gil93] Michael I. Gil. On equalities for eigenvalues of matrices. Linear Algebra and its Applications, 184:201–??, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gil:1995:EIE

[Gil95] Michael I. Gil’. Erratum:

“On inequalities for eigenvalues of matrices” [Linear Algebra Appl. 184 (1993), 201– 206, MR 94a:15027]. Linear Algebra and its Applications, 220(1–3):1–2, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gil:1997:NCM

[Gil97] M. I. Gil’. A nonsingularity criterion for matrices. Linear Algebra and its Applications, 253(1–3):79– 87, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500704.

Gomez:1998:AOL

[GJMNV98] J. R. Gómez, A. Jiménez-Merchán, and J. Núñez-Valdés. An algorithm to obtain laws of families of filiform Lie algebras. Linear Algebra and its Applications, 279(1–3):1–12, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        279/1-3/6077.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6077.

Gunawardena:1991:MIM

[GJS91] Ananda D. Gunawardena, S. K. Jain, and Larry Snyder. Modified iterative methods for consistent linear systems. Linear Algebra and its Applications, 154/156:123–143, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Glentis:1993:EAS

[GK93] George-Othon Glentis and Nicholas Kalouptsidis. Efficient algorithms for the solution of block linear systems with Toeplitz entries. Linear Algebra and its Applications, 179:85–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godsil:1995:MED

[GK95] C. D. Godsil and J. H. Koolen. On the multiplicity of eigenvalues of distanceregular graphs. Linear Algebra and its Applications, 226/228 (1–3):273–275, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500152.

Grunenfelder:1997:RCM

[GK97] Luzius Grunenfelder and Tomaž Košir. On a representation of commuting maps by tensor products. Linear Algebra and its Applications, 251(1–3):215–222, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9600555.

Gohberg:1996:ABS

[GKK96] I. Gohberg, M. A. Kaashoek, and J. Kos. The asymptotic behavior of the singular values of matrix powers and applications. Linear Algebra and its Applications, 245(1–3):55– 76, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9682454.

Gohberg:1996:FBL

[GKL96] I. Gohberg, M. A. Kaashoek, and L. Lerer. Factorization of banded lower triangular infinite matrices. Linear Algebra and its Applications, 247(1–3):347–357, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500122.

Gregory:1992:RSM

[GKP92] D. A. Gregory, S. J. Kirkland, and J. J. Pullman. Rowstochastic matrices with a common left fixed vector. Linear Algebra and its Applications, 169:131–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gregory:1993:PCB

[GKP93] D. A. Gregory, S. Kirkland, and N. J. Pullman. Power convergent Boolean matrices. Linear Algebra and its Applications, 179:105–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gregory:1995:BEP

[GKP95] D. A. Gregory, S. J. Kirkland, and N. J. Pullman. A bound on the exponent of a primitive matrix using Boolean rank. Linear Algebra and its Applications, 217(1–3):101– 116, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9200003. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Gohberg:1990:RRM

[GKR90] I. Gohberg, M. A. Kaashoek, and A. C. M. Ran. Regular rational matrix functions with prescribed null and pole data except at infinity. Linear Algebra and its Applications, 137/138:387–412, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grossman:1994:AGT

[GKS94] Jerrold W. Grossman, Devadatta M. Kulkarni, and Irwin E. Schochetman. Algebraic graph theory without orientation. Linear Algebra and its Applications, 212/213: 289–307, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Grossman:1995:MIM

[GKS95] Jerrold W. Grossman, Devadatta M. Kulkarni, and Irwin E. Schochetman. On the minors of an incidence matrix and its Smith Normal Form. Linear Algebra and its Applications, 218(1–3): 213–224, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300173.

Gritzmann:1995:CPM

[GKT95] Peter Gritzmann, Victor Klee, and Bit Shun Tam. Crosspositive matrices revisited. Linear Algebra and its Applications, 223/224(1–3):285– 305, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9300364. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Gohberg:1993:LTR

[GKv93] I. Gohberg, M. A. Kaashock, and F. van Schagen. On the local theory of regular analytic matrix functions. Linear Algebra and its Applications, 182: 9–??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:1998:OBQ

[GKv98] I. Gohberg, M. A. Kaashoek, and F. van Schagen. Operator blocks and quadruples of subspaces: Classification and the eigenvalue completion problem. Linear Algebra and its Applications, 269(1): 65–89, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9781513.

Gohberg:1991:NEB

[GKW91] I. Gohberg, M. A. Kaashoek, and H. J. Woerdeman. A note on extensions of band matrices with maximal and submaximal invertible blocks. Linear Algebra and its Applications, 150:157–166, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Grasselli:1991:GAL

[GL91] Osvaldo Maria Grasselli and Sauro Longhi. The geometric approach for linear periodic discrete-time systems. Linear Algebra and its Applications, 158:27–60, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghinelli:1995:GQR

[GL95] Dina Ghinelli and Stefan Löwe. Generalized quadrangles with a regular point and association schemes. Linear Algebra and its Applications, 226/228(1–3):87–104, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400305.

Gonzaga:1997:NPC

[GL97] Clovis C. Gonzaga and Hugo J. Lara. A note on properties of condition numbers. Linear Algebra and its Applications, 261(1–3):269– 273, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600409.

Gladwell:1998:TPA

[Gla98] G. M. L. Gladwell. Total positivity and the QR algorithm. Linear Algebra and its Applications, 271(1–3):257– 272, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700275.

Guo:1995:NSL

[GLW95a] Jong-Sheng Guo, Wen-Wei Lin, and Chern-Shuh Wang. Numerical solutions for large sparse quadratic eigenvalue problems. Linear Algebra and its Applications, 225(1):57– 89, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300318.

Guo:1995:NDS

[GLW95b] Jong-Shenq Guo, Wen-Wei Lin, and Chern-Shuh Wang. Nonequivalence deflation for the solution of matrix latent value problems. Linear Algebra and its Applications, 231(1–3):15–45, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400008.

Green:1990:OMM

[GM90] William L. Green and T. D.

Morley. Operator means and matrix functions. Linear Algebra and its Applications, 137/138:453–465, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:1993:BAZ

[GM93] S. González and C. Martínez. Bernstein algebras with zero derivation algebra. Linear Algebra and its Applications, 191:235–244, September 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garrity:1995:IVV

[GM95a] Thomas Garrity and Robert Mizner. Invariants of vectorvalued bilinear and sesquilinear forms. Linear Algebra and its Applications, 218(1–3): 225–237, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300174.

Goodman:1995:RWI

[GM95b] Jonathan Goodman and Neal Madras. Random-walk interpretations of classical iteration methods. Linear Algebra and its Applications, 216(1–3):61– 79, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300101.

Getz:1997:DMP

[GM97a] Neil H. Getz and Jerrold E.

Marsden. Dynamical methods for polar decomposition and inversion of matrices. Linear Algebra and its Applications, 258(1–3):311– 343, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600235.

Gollan:1997:CDS

[GM97b] Holger W. Gollan and Gerhard O. Michler. Construction of a 45, 694-dimensional simple module of Lyons’ sporadic group over GF(2). Linear Algebra and its Applications, 256(1–3):185–197, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600002.

Gavrilyuk:1998:EAS

[GM98a] Ivan P. Gavrilyuk and Vladimir L.

Makarov. Exact and approximate solutions of some operator equations based on the Cayley transform. Linear Algebra and its Applications, 282(1–3):97–121, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        282/1-3/6171.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=282&issue=
                                        1-3&aid=6171.

Gondran:1998:EEF

[GM98b] M. Gondran and M. Minoux. Eigenvalues and eigenfunctionals of diagonally dominant endomorphisms in Min-Max analysis. Linear Algebra and its Applications, 282(1–3):47–61, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        282/1-3/6160.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=282&issue=
                                        1-3&aid=6160.

Graves-Morris:1990:CHC

[GMJ90] P. R. Graves-Morris and C. R. Johnson. Confluent Hermitian Cauchy matrices are diagonally signed. Linear Algebra and its Applications, 142: 129–140, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Graves-Morris:1995:DID

[GMJ95] P. R. Graves-Morris and C. R. Johnson. Determinantal inequalities for diagonally signed matrices and an application to Gram-Cauchy matrices. Linear Algebra and its Applications, 225(1):125– 140, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9590928.

Gyory:1998:LRC

[GMŠ98] Máté Györy, Lajos Molnár, and Peter Šemrl. Linear rank and corank preserving maps on B(H) and an application to ^{∗}-semigroup isomorphisms of operator ideals. Linear Algebra and its Applications, 280(2–3):253–266, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6166.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6166.

Gao:1993:GMS

[GN93] Mei Gao and Michael Neumann. A global minimum search algorithm for estimating the distance to uncontrollability. Linear Algebra and its Applications, 188/189:305– 350, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grady:1997:GIM

[GN97] Michael Grady and Morris Newman. The geometry of an interchange: Minimal matrices and circulants. Linear Algebra and its Applications, 262(1–3):11–25, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=262&issue=1-
                                        3&aid=9600465.

Gilbert:1997:SSP

[GNP97] John R. Gilbert, Esmond G. Ng, and Barry W. Peyton. Separators and structure prediction in sparse orthogonal factorization. Linear Algebra and its Applications, 262(1–3):83–97, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=262&issue=1-
                                        3&aid=9600473.

Grunenfelder:1993:LRS

[GO93] L. Grunenfelder and M. Omladic. Linearly recursive sequences and operator polynomials. Linear Algebra and its Applications, 182:127– ??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:1994:CMV

[GO94] I. Gohberg and V. Olshevsky.

Complexity of multiplication with vectors for structured matrices. Linear Algebra and its Applications, 202:163–192, April 15, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godsil:1997:CGE

[God97]  C. D. Godsil.  Compact

graphs and equitable partitions. Linear Algebra and its Applications, 255(1–3):259– 266, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9783595.

Goeleven:1996:UTG

[Goe96] D. Goeleven. A uniqueness theorem for the generalizedorder linear complementary problem associated with Mmatrices. Linear Algebra and its Applications, 235(1–3): 221–227, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400141.

Gollan:1993:CCS

[Gol93] Holger W. Gollan. Conjugacy class sums for induced modules: Construction and applications. Linear Algebra and its Applications, 192:115–135, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Goldberg:1997:MND

[Gol97a]  Moshe Goldberg.  Mor-

ris Newman — A discrete mathematician for all seasons. Linear Algebra and its Applications, 254(1–3):7– 18, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9500779. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Goldwasser:1997:SPC

[Gol97b] John L. Goldwasser. Shortened and punctured codes and the MacWilliams identities. Linear Algebra and its Applications, 253(1–3):1– 13, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500671.

Goldwasser:1997:TBZ

[Gol97c] John L. Goldwasser. Triangular blocks of zeros in (0, 1) matrices with small permanents. Linear Algebra and its Applications, 252(1–3):367– 374, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500782.

Golumbic:1998:MSP

[Gol98] Martin Charles Golumbic.

Matrix sandwich problems. Linear Algebra and its Applications, 277(1–3):239–251, June 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6075.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6075.

Gombani:1995:SPM

[Gom95] Andrea Gombani. On the Schmidt pairs of multivariable Hankel operators and robust control. Linear Algebra and its Applications, 223/224(1–3): 243–265, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9500049. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Gong:1996:GST

[Gon96] Ming-Peng Gong. Generalized symmetric tensors and related topics. Linear Algebra and its Applications, 236(1–3): 113–129, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400136.

Gong:1997:SDG

[Gon97] Ming-Peng Gong. Semigroups defined by generalized Schur functions. Linear Algebra and its Applications, 257(1– 3):65–75, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600039.

Gorodetsky:1992:IQB

[Gor92] Michael Gorodetsky. Inversion of quasitriangular block Toeplitz matrices. Linear Algebra and its Applications, 167:119–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gouveia:1993:GIH

[Gou93] M. C. Gouveia. Generalized invertibility of Hankel and Toeplitz matrices. Linear Algebra and its Applications, 193:95–106, November 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gouveia:1994:GMP

[Gou94] M. C. Gouveia. Group and Moore–Penrose invertibility of Bezoutians. Linear Algebra and its Applications, 197/198: 495–509, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Gouveia:1997:RHM

[Gou97] M. C. Gouveia. Regular Hankel matrices over integral domains. Linear Algebra and its Applications, 255(1–3):335– 347, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600003.

Gover:1994:ETT

[Gov94] M. J. C. Gover. The eigenproblem of a tridiagonal 2-Toeplitz matrix. Linear Algebra and its Applications, 197/198:63–78, January 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Green:1996:CSN

[GOV96] Bryan C. J. Green, D. D. Olesky, and P. Van Den Driessche. Classes of sign nonsingular matrices with a specified number of zero entries. Linear Algebra and its Applications, 248(1–3):253–275, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500198.

Govaerts:1997:DFM

[Gov97] W. Govaerts. Defining functions for manifolds of matrices. Linear Algebra and its Applications, 266(1–3):49–68, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9700630.

Gowda:1990:TPM

[Gow90] M. Seetharama Gowda. On the transpose of a pseudomonotone matrix and the LCP. Linear Algebra and its Applications, 140:129–137, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1990:EBM

[GP90] Robert Grone and Stephen Pierce. Extremal bipartite matrices. Linear Algebra and its Applications, 131:39–50, 1990.

Gouveia:1991:AGI

[GP91a] M. C. Gouveia and R. Puystjens. About the group inverse and Moore–Penrose inverse of a product. Linear Algebra and its Applications, 150:361– 369, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Grone:1991:EPS

[GP91b] Bob Grone and Steve Pierce. Extremal positive semidefinite doubly stochastic matrices. Linear Algebra and its Applications, 150:107–117, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Gasca:1992:TPN

[GP92] M. Gasca and J. M. Pena. Total positivity and Neville elimination. Linear Algebra and its Applications, 165:25– ??, March 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gasca:1994:MDN

[GP94a] M. Gasca and J. M. Peña. A matricial description of Neville elimination with applications to total positivity. Linear Algebra and its Applications, 202:33–53, April 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gasca:1994:TSS

[GP94b] M. Gasca and J. M. Peña. A test for strict sign-regularity. Linear Algebra and its Applications, 197/198:133–142, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Grigorieff:1995:MES

[GP95] R. D. Grigorieff and R. Plato.

On a minimax equality for seminorms. Linear Algebra and its Applications, 221(1–3): 227–243, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300258.

Garcia-Planas:1997:DST

[GPM97] Ma. I. García-Planas and Ma. D. Magret. Deformation and stability of triples of matrices. Linear Algebra and its Applications, 254(1–3): 159–192, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600404. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Garcia-Planas:1999:SLS

[GPM99a] M. I. García-Planas and M. D.

Magret. Stratification of linear systems. Bifurcation diagrams for families of linear systems.

Linear Algebra and its Applications, 297(1–3):23–56, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/104/17/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/104/17/19/article.
pdf.

Garcia-Planas:1999:ASS

[GPM99b] M. Isabel García-Planas and M. Dolors Magret. An alternative system of structural invariants of quadruples of matrices. Linear Algebra and its Applications, 291(1–3):83–102, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6332.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6332.

Gleser:1994:BBA

[GPPS94] Leon Jay Gleser, Michael D. Perlman, S. James Press, and Allan R. Sampson. A brief biography and appreciation of Ingram Olkin. Linear Algebra and its Applications, 199:1–15, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1994:SBD

[GPRL94] Robert Grone, Stephen Pierce, James Ross, and Chi-Kwong Li. Spectral bounds derived from quadratic forms on decomposable tensors. Linear Algebra and its Applications, 201:181–198, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Planas:1999:SMD

[GPS99] M. Isabel García-Planas and Vladimir V. Sergeichuk. Simplest miniversal deformations of matrices, matrix pencils, and contragredient matrix pencils. Linear Algebra and its Applications, 302-303(1–3):45–61, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/29/article.
pdf.

Grone:1990:ECM

[GPW90] Robert Grone, Stephen Pierce, and William Watkins. Extremal correlation matrices. Linear Algebra and its Applications, 134:63–70, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Giovagnoli:1990:GMO

[GR90] Alessandra Giovagnoli and Mario Romanazzi. A group majorization ordering for correlation matrices. Linear Algebra and its Applications, 127: 139–155, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gradl:1996:REF

[Gra96] Hans Gradl. A result on exponents of finite-dimensional simple Lie algebras and its application to Kac–Moody algebras. Linear Algebra and its Applications, 233(1–3):189– 206, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400069.

Greene:1992:PCI

[Gre92] Curtis Greene. Proof of a conjecture on immanants of the Jacobi-Trudi matrix. Linear Algebra and its Applications, 171:65–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gripenberg:1996:CJS

[Gri96] Gustaf Gripenberg. Computing the joint spectral radius. Linear Algebra and its Applications, 234(1–3):43– 60, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400082.

Grone:1991:GLG

[Gro91] Robert Grone. On the geometry and Laplacian of a graph. Linear Algebra and its Applications, 150:167–178, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Grochenig:1993:DTI

[Grö93] Karlheinz Gröchenig. A discrete theory of irregular sampling. Linear Algebra and its Applications, 193:129–150, November 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1994:BMM

[Gro94] Robert Grone. A biography of Marvin Marcus. Linear Algebra and its Applications, 201: 1–20, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1995:EBM

[Gro95a] Robert Grone. Erratum: “A biography of Marvin Marcus”

[Linear Algebra Appl. 201 (1994), 1–20]. Linear Algebra and its Applications, 220(1–3): 6, April 1995. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:1995:RMM

[Gro95b] Robert Grone. Review of Matrices: Methods and Applications by Stephen Barnett. Linear Algebra and its Applications, 214(1–3):277– 279, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300241.

Gross:1997:SSR

[Gro97a]  Jürgen Groß.  On semi-

star-related matrices. Linear Algebra and its Applications, 264(1–3):495–503, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700005.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700005.

Gross:1997:SGI

[Gro97b] Jürgen Groß. Special generalized inverse matrices connected with the theory of unified least squares. Linear Algebra and its Applications, 264(1–3):325–327, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9507778.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9507778.

Galkowski:1998:MRB

[GRO98a] K. Galkowski, E. Rogers, and D. H. Owens. Matrix rank based conditions for reachability/controllability of discrete linear repetitive processes. Linear Algebra and its Applications, 275/276(1– 3):201–224, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6056.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6056. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Grosshans:1998:ITO

[Gro98b] Frank D. Grosshans. Invariant theory, orbits and non-decomposable quadruples of subspaces having nonzero defect. Linear Algebra and its Applications, 282(1– 3):63–81, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6174.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6174.

Gross:1999:ESM

[Gro99a] Jürgen Gross. Explicit solutions to the matrix inverse problem AX=B. Linear Algebra and its Applications, 289(1–3):131–134, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6008.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6008.

Gross:1999:IHP

[Gro99b] Jürgen Groß. Idempotency of the Hermitian part of a complex matrix. Linear Algebra and its Applications, 289(1–3):135–139, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

tree/store/laa/sub/1999/
289/1-3/6045.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6045.

Gross:1999:NRS

[Gro99c] Jürgen Groß. A note on the rank-subtractivity ordering. Linear Algebra and its Applications, 289(1–3): 151–160, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6339.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6339.

Gross:1999:POPa

[Gro99d] Jürgen Groß. On the product of orthogonal projectors. Linear Algebra and its Applications, 289(1–3): 141–150, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6185.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6185.

Gross:1999:SRE

[Gro99e] Jürgen Groß. Solution to a rank equation. Linear Algebra and its Applications, 289(1–3):127–130, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        289/1-3/6001.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=289&issue=
                                        1-3&aid=6001.

Gurvits:1992:CCP

[GRS92] Leonid Gurvits, Leiba Rodman, and Tamir Shalom. Controllability and completion of partial upper triangular matrices over rings. Linear Algebra and its Applications, 172: 135–??, July 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:1990:BNM

[GS90] I. Gohberg and T. Shalom. On Bezoutians of nonsquare matrix polynomials and inversion of matrices with nonsquare blocks. Linear Algebra and its Applications, 137/138: 249–323, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goodman:1993:FTP

[GS93] T. N. T. Goodman and A. Sharma. Factorization of totally positive, symmetric, periodic, banded matrices with applications. Linear Algebra and its Applications, 178:85–??, January 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gekhtman:1995:ING

[GS95] M. I. Gekhtman and M. Shmoĭsh.

On invertibility of nonsquare generalized Bezoutians. Linear Algebra and its Applications, 223/224(1–3):205– 241, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400310. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Godunov:1996:EDM

[GS96a] S. K. Godunov and M. Sadkane. Elliptic dichotomy of a matrix spectrum. Linear Algebra and its Applications, 248(1–3):205–232, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500195.

Sun:1996:TFM

[gS96b]  Ji guang Sun.  On two

functions of a matrix with positive definite Hermitian part. Linear Algebra and its Applications, 244(1–3):55– 68, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=244&issue=1-
                                        3&aid=9400208.

Sun:1996:PAS

[gS96c]  Ji guang Sun.  Pertur-

bation analysis of system Hessenberg and Hessenbergtriangular forms. Linear Algebra and its Applications, 241/243(1–3):811–849, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500107. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Sun:1997:BED

[gS97a] Ji guang Sun. Backward error for the discrete-time algebraic Riccati equation. Linear Algebra and its Applications, 259(1–3):183–208, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600283.

Sun:1997:PAM

[gS97b] Ji guang Sun. Perturbation analysis of the matrix sign function. Linear Algebra and its Applications, 250(1–3):177–206, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500522.

Godunov:1998:CPS

[GS98a] S. K. Godunov and M. Sadkane. Computation of pseudospectra via spectral projectors. Linear Algebra and its Applications, 279(1–3): 163–175, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6109.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6109.

Sun:1998:SAD

[gS98b] Ji guang Sun. Sensitivity analysis of the discrete-time algebraic Riccati equation. Linear Algebra and its Applications, 275/276(1–3):595–615, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6024.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6024. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Govaerts:1999:MMJ

[GS99a] W. Govaerts and B. Sijnave. Matrix manifolds and the Jordan structure of the bialternate matrix product. Linear Algebra and its Applications, 292(1–3):245–266, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        292/1-3/6387.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=292&issue=
                                        1-3&aid=6387.

Sun:1999:SBE

[gS99b] Ji guang Sun. Structured backward errors for KKT systems. Linear Algebra and its Applications, 288(1–3): 75–88, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6294.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6294.

Gow:1993:GJU

[GT93] R. Gow and M. C. Tamburini.

Generation of SL(n, Z) by a Jordan unipotent matrix and its transpose. Linear Algebra and its Applications, 181: 63–??, March 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Georgiou:1996:NCB

[GT96a] A. C. Georgiou and N. Tsantas. Nonstationary cyclic behavior in Markov systems. Linear Algebra and its Applications, 237/238(1–3):549– 578, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400302. Special

issue honoring Calyampudi Radhakrishna Rao.

Gross:1996:LSD

[GT96b] Jürgen Gross and Götz Trenkler. On the least squares distance between affine subspaces. Linear Algebra and its Applications, 237/238(1–3): 269–276, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500648. Special

issue honoring Calyampudi Radhakrishna Rao.

Gross:1997:GHP

[GT97a] Jürgen Groß and Götz Trenkler. Generalized and hypergeneralized projectors. Linear Algebra and its Applications, 264(1–3):463–474, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9600541.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9600541.

Gross:1997:SRP

[GT97b] Jürgen Groß and Sven-Oliver Troschke. Some remarks on partial orderings of nonnegative definite matrices. Linear Algebra and its Applications, 264(1–3):457–461, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600470.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600470.

Guseinov:1997:EPM

[GT97c] Gusein Guseinov and Mehmet Terziler. The eigenvalue problem for a matrix of a special form and its applications. Linear Algebra and its Applications, 259(1–3):329– 346, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600296.

Gu:1996:TBP

[GTÖ96] Caixing Gu, Onur Toker, and Hitay Özbay. On the two-block H(∞) problem for a class of unstable distributed systems. Linear Algebra and its Applications, 234(1–3):227–244, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1996&volume=234&issue=1-
                                        3&aid=9400112.

Gross:1999:SOS

[GTT99] Jürgen Groß, Götz Trenkler, and Sven-Oliver Troschke. On semi-orthogonality and a special class of matrices. Linear Algebra and its Applications, 289(1–3):169–182, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        289/1-3/6002.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=289&issue=
                                        1-3&aid=6002.

Gallivan:1996:HPA

[GTVV96] K. A. Gallivan, S. Thirumalai, P. Van Dooren, and V. Vermaut. High performance algorithms for Toeplitz and block Toeplitz matrices. Linear Algebra and its Applications, 241/243(1–3):343– 388, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500649. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Goreinov:1997:TPA

[GTZ97] S. A. Goreinov, E. E. Tyrtyshnikov, and N. L. Zamarashkin. A theory of pseudoskeleton approximations. Linear Algebra and its Applications, 261(1–3):1– 21, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600301.

Gu:1999:FRH

[Gu99a] Caixing Gu. Finite rank Hankel operators on the polydisk. Linear Algebra and its Applications, 288(1–3): 269–281, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6329.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6329.

Gu:1999:TTL

[Gu99b] Chuanqing Gu. Thiele-type and Lagrange-type generalized inverse rational interpolation for rectangular complex matrices. Linear Algebra and its Applications, 295 (1–3):7–30, July 01, 1999.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6384.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6384.

Guo:1991:SRS

[Guo91] Chun Hua Guo. Some results on sparse block factorization iterative methods. Linear Algebra and its Applications, 145:187–199, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:1996:IEP

[Guo96] Wuwen Guo. An inverse eigenvalue problem for nonnegative matrices. Linear Algebra and its Applications, 249(1–3):67– 78, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500268.

Guo:1999:NSN

[Guo99] Chun-Hua Guo. On the numerical solution of a nonlinear matrix equation in Markov chains. Linear Algebra and its Applications, 288(1–3): 175–186, February 1, 1999.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6300.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6300.

Guralnick:1991:SMC

[Gur91] Robert M. Guralnick. Similarity of matrices over commutative rings. Linear Algebra and its Applications, 157:55– 68, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guralnick:1994:IPA

[Gur94] Robert M. Guralnick. Invertible preservers and algebraic groups. Linear Algebra and its Applications, 212/213:249– 257, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Gurvits:1995:SDL

[Gur95]  Leonid Gurvits.  Stabil-

ity of discrete linear inclusion. Linear Algebra and its Applications, 231(1–3):47–85, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1995&volume=231&issue=1-
                                        3&aid=9400010.

Gustafson:1991:MM

[Gus91] William H. Gustafson. Modules and matrices. Linear Algebra and its Applications, 157:3–19, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gustafson:1994:A

[Gus94] Karl Gustafson. Antieigenvalues. Linear Algebra and its Applications, 208/209:437– 454, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gustafson:1995:MT

[Gus95a]  Karl Gustafson.  Matrix

trigonometry. Linear Algebra and its Applications, 217(1–3): 117–140, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=217&issue=1-
                                        3&aid=9400213. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Gustafson:1995:QME

[Gus95b] William H. Gustafson. Quivers and matrix equations. Linear Algebra and its Applications, 231(1–3):159–174, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400022.

Gustafson:1999:GMK

[Gus99] Karl Gustafson. The geometrical meaning of the Kantorovich-Wielandt inequalities. Linear Algebra and its Applications, 296(1–3):143– 151, July 15, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/24/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/24/article.
pdf.

Gutknecht:1993:SRR

[Gut93] Martin H. Gutknecht. Stable row recurrences for the Padé table and generically superfast lookahead solvers for non-Hermitian Toeplitz systems. Linear Algebra and its Applications, 188/189:351– 421, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

GutierrezFernandez:1998:SBP

[Gut98] J. Carlos Gutiérrez Fernández.

Structure of Bernstein populations of type (3, n − 3).

Linear Algebra and its Applications, 269(1):17–32, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9781512.

Georgiou:1992:PAA

[GV92] A. C. Georgiou and P.- C. G. Vassiliou. Periodicity of asymptotically attainable structures in nonhomogeneous Markov systems. Linear Algebra and its Applications, 176: 137–174, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gracia:1995:SIS

[GV95] Juan M. Gracia and Francisco E. Velasco. Stability of invariant subspaces of regular matrix pencils. Linear Algebra and its Applications, 221(1–3): 219–226, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300256.

Gonzalez-Vega:1996:EPB

[GV96] Laureano González-Vega. An elementary proof of Barnett’s theorem about the greatest common divisor of several univariate polynomials. Linear Algebra and its Applications, 247(1–3):185–202, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500099.

Gracia:1998:GAB

[GV98] Juan M. Gracia and Francisco E. Velasco. Global analytic block similarity to a Brunovsky form. Linear Algebra and its Applications, 282(1–3):233–247, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6186.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6186.

Gregory:1996:SRD

[GVS96] David A. Gregory, Kevin N.

Vander Meulen, and Bryan L. Shader. Skew rank decompositions. Linear Algebra and its Applications, 244(1–3):123– 154, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=244&issue=1-
                                        3&aid=9400217.

Green:1995:TCP

[GW95] Michael M. Green and Alan N. Willson, Jr. Transistor circuits and potentially stable operating points. Linear Algebra and its Applications, 223/224(1–3):267– 283, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400156. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Giovagnoli:1996:CMS

[GW96] Alessandra Giovagnoli and Henry P. Wynn. Cyclic majorization and smoothing operators. Linear Algebra and its Applications, 239(1–3):215– 225, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=239&issue=1-
                                        3&aid=9400359.

Gu:1999:BAN

[GW99] Gui-Ding Gu and He-Bing Wu. A block EN algorithm for nonsymmetric linear systems with multiple right-hand sides. Linear Algebra and its Applications, 299(1–3):1–20, September 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/21/article.
pdf.

Gregory:1999:BDH

[GWS99] David A. Gregory, Valerie L.

Watts, and Bryan L. Shader. Biclique decompositions and Hermitian rank. Linear Algebra and its Applications, 292(1–3):267–280, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6390.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6390.

Golub:1994:PAC

[GZ94] Gene H. Golub and Hong Yuan Zha. Perturbation analysis of the canonical correlations of matrix pairs. Linear Algebra and its Applications, 210:3– 28, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gibson:1998:COM

[GZ98a] Peter M. Gibson and Guo-Hui Zhang. Combinatorially orthogonal matrices and related graphs. Linear Algebra and its Applications, 282(1–3):83–95, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        282/1-3/6172.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=282&issue=
                                        1-3&aid=6172.

Gower:1998:OAA

[GZ98b] J. C. Gower and B. Zeilman. Orthogonality and its approximation in the analysis of asymmetry. Linear Algebra and its Applications, 278(1–3):183–193, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        278/1-3/6090.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=278&issue=
                                        1-3&aid=6090.

Hagiwara:1994:SOC

[HA94] Tomomichi Hagiwara and Mituhiko Araki. A successive optimal construction procedure for state feedback gains. Linear Algebra and its Applications, 203/204:659–673, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Habibi:1994:SGG

[Hab94] Javad Faghih Habibi. The spherical gap of the graph of a linear transformation. Linear Algebra and its Applications, 212/213:501–503, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Hadwin:1994:AVM

[Had94] Don Hadwin. An algebraic version of the multiplication property of the Fredholm index. Linear Algebra and its Applications, 208/209:229– 230, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjicostas:1997:CMS

[Had97] Petros Hadjicostas. Copositive matrices and Simpson’s paradox. Linear Algebra and its Applications, 264(1–3):475–488, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600536.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.

cgi?year=1997&volume=264&
issue=1-3&aid=9600536.

Haemers:1995:IEG

[Hae95] Willem H. Haemers. Interlacing eigenvalues and graphs. Linear Algebra and its Applications, 226/228 (1–3):593–616, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500199.

Haemers:1996:DRS

[Hae96] Willem H. Haemers. Distanceregularity and the spectrum of graphs. Linear Algebra and its Applications, 236(1–3): 265–278, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400166.

Hafner:1995:EAF

[Haf95]  James Lee Hafner.  Ex-

plicit and asymptotic formulas for LDM^{t} factorization of banded Toeplitz matrices. Linear Algebra and its Applications, 222(1–3):97– 126, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9594777.

Hansen:1990:RBS

[Han90a] Per Christian Hansen. Relations between SVD and GSVD of discrete regularization problems in standard and general form. Linear Algebra and its Applications, 141: 165–176, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:1990:ALS

[Han90b] Sönke Hansen. An algorithm for Lagrangian subspaces. Linear Algebra and its Applications, 129:155–176, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanlon:1998:LMT

[Han98] Phil Hanlon. To the Latimer– Macduffee theorem and beyond! Linear Algebra and its Applications, 280(1):21– 37, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Hermida-Alonso:1993:IFE

[HAP93] José Á. Hermida-Alonso and Miriam Pisonero. Invariant factors of an endomorphism and finite free resolutions. Linear Algebra and its Applications, 187:201–226, July 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hermida-Alonso:1995:WPS

[HAP95] José Á. Hermida-Alonso and Miriam Pisonero. What polynomial satisfies a given endomorphism? Linear Algebra and its Applications, 216(1–3):125–138, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300125.

Happel:1997:TCM

[Hap97] Dieter Happel. The trace of the Coxeter matrix and Hochschild cohomology. Linear Algebra and its Applications, 258(1–3):169–177, June ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600195.

Hermida-Alonso:1995:FIL

[HAPSG95] José Á. Hermida-Alonso, M. Pilar Perez, and Tomás Sánchez-Giralda. Feedback invariants for linear dynamical systems over a principal ideal domain. Linear Algebra and its Applications, 218(1–3):29– 45, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300153.

Hermida-Alonso:1996:BCF

[HAPSG96] Jose A. Hermida-Alonso, M. Pilar Perez, and Tomas Sanchez-Giralda. Brunovsky’s canonical form for linear dynamical systems over commutative rings. Linear Algebra and its Applications, 233(1–3):131–147, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400062.

Hartman:1990:CGR

[Har90] Jim Hartman. On a conjecture of Gohberg and Rodman. Linear Algebra and its Applications, 140:267–278, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hari:1991:PAD

[Har91] Vjeran Hari. On pairs of almost diagonal matrices. Linear Algebra and its Applications, 148:193–223, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartfiel:1992:TMS

[Har92] D. J. Hartfiel. Tracking in matrix systems. Linear Algebra and its Applications, 165: 233–??, March 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Harris:1993:CFC

[Har93a] Lawrence A. Harris. Computation of functions of certain operator matrices. Linear Algebra and its Applications, 194:31–34, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartfiel:1993:RLS

[Har93b] D. J. Hartfiel. Results on limiting sets of Markov set chains. Linear Algebra and its Applications, 195:155–163, December 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:1993:PDR

[Har93c] Robert E. Hartwig. The pyramid decomposition and rank minimization. Linear Algebra and its Applications, 191:53– 76, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:1994:WCN

[Har94] Robert E. Hartwig. The weighted ∗-core-nilpotent decomposition. Linear Algebra and its Applications, 211:101– 111, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Harris:1995:FOM

[Har95] Lawrence A. Harris. Factorizations of operator matrices. Linear Algebra and its Applications, 225(1):37– 41, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300308.

Hartfiel:1996:SN

[Har96] D. J. Hartfiel. Similitudes and the l_{1}-norm. Linear Algebra and its Applications, 245(1–3):305–319, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9500053.

Hartfiel:1997:ESI

[Har97] D. J. Hartfiel. Eigenvector sets for intervals of matrices. Linear Algebra and its Applications, 262(1–3):1– 10, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600446.

Harel:1999:SUP

[Har99a] Guershon Harel. Students’ understanding of proofs: a historical analysis and implications for the teaching of geometry and linear algebra. Linear Algebra and its Applications, 302-303(1–3):601–613, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
57/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/57/article.
pdf.

Harville:1999:UGS

[Har99b] David A. Harville. Use of the Gibbs sampler to invert large, possibly sparse, positive definite matrices. Linear Algebra and its Applications, 289(1–3):203–224, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6293.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6293.

Haslett:1996:ULM

[Has96] Stephen Haslett. Updating linear models with dependent errors to include additional data and/or parameters. Linear Algebra and its Applications, 237/238(1–3): 329–349, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500255. Special

issue honoring Calyampudi Radhakrishna Rao.

Hermida-Alonso:1990:DPL

[HASG90] J. A. Hermida-Alonso and T. Sánchez-Giralda. On the duality principle for linear dynamical systems over commutative rings. Linear Algebra and its Applications, 139:175– 180, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haukkanen:1992:HDG

[Hau92] Pentti Haukkanen. Higherdimensional GCD matrices. Linear Algebra and its Applications, 170:53–??, June 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hautus:1994:OS

[Hau94] M. L. J. Hautus. Operator substitution. Linear Algebra and its Applications, 205/206: 713–739, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hautus:1995:SMR

[Hau95]  M. L. J. Hautus.  Sub-

stitution of matrices over rings. Linear Algebra and its Applications, 226/228 (1–3):353–370, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500155.

Haukkanen:1996:MMP

[Hau96] Pentti Haukkanen. On meet matrices on posets. Linear Algebra and its Applications, 249(1–3):111–123, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500349.

Have:1993:SRM

[Hav93] Gabrielle Ten Have. Structure of the n th roots of a matrix. Linear Algebra and its Applications, 187:59–??, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hawkins:1990:AMP

[HB90] Douglas M. Hawkins and Dan Bradu. Application of the Moore–Penrose inverse of a data matrix in multiple regression. Linear Algebra and its Applications, 127:403–425, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1997:TTT

[HB97] Georg Heinig and Adam Bojanczyk. Transformation techniques for Toeplitz and Toeplitz-plus-Hankel matrices. I. transformations. Linear Algebra and its Applications, 254(1–3):193–226, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600527. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Heinig:1998:TTT

[HB98] Georg Heinig and Adam Bojanczyk. Transformation techniques for Toeplitz and Toeplitz-plus-Hankel matrices II. algorithms. Linear Algebra and its Applications, 278 (1–3):11–36, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6043.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6043.

Higham:1998:MIM

[HC98a] Nicholas J. Higham and Sheung Hun Cheng. Modifying the inertia of matrices arising in optimization. Linear Algebra and its Applications, 275/276(1–3):261–279, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6011.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6011. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Hung:1998:RBD

[HC98b] Y. S. Hung and D. L. Chu. Relationships between discretetime and continuous-time algebraic Riccati inequalities. Linear Algebra and its Applications, 270(1–3):287–313,

????  ??, 1998.  CO-

DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700246.

Hung:1998:ELF

[HC98c] Y. S. Hung and Delin Chu. On extended (J.J′)-lossless factorization. Linear Algebra and its Applications, 271(1–3): 117–138, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700257.

Cao:1996:SVK

[hCjXL96] Zhi hao Cao, Jin jun Xie, and Ren-Cang Li. A sharp version of Kahan’s theorem on clustered eigenvalues. Linear Algebra and its Applications, 245(1–3):147–155, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400226.

Ho:1999:GHB

[HDB99] Ming-Tzu Ho, Aniruddha Datta, and S. P. Bhattacharyya. Generalizations of the Hermite-Biehler theorem. Linear Algebra and its Applications, 302-303(1– 3):135–153, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/33/article.
pdf.

Haffaf:1999:GMC

[HDTR99] Hafid Haffaf, Geneviève Dauphin-Tanguy, and Mustapha Kamel Rahmouni. Graphical matroid for causality assignment in bond graphs. Linear Algebra and its Applications, 293(1–3):257–272, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6399.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6399.

He:1994:NCM

[He94] Bing Sheng He. New contraction methods for linear inequalities. Linear Algebra and its Applications, 207:115–133, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1992:IPH

[Hei92] Georg Heinig. Inverse problems for Hankel and Toeplitz matrices. Linear Algebra and its Applications, 165:1–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heiligers:1994:TNM

[Hei94] Berthold Heiligers. Totally nonnegative moment matrices. Linear Algebra and its Applications, 199:213–227, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1995:GIH

[Hei95a]  Georg Heinig.  General-

ized inverses of Hankel and Toeplitz mosaic matrices. Linear Algebra and its Applications, 216(1–3):43–59, February ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300097.

Heinig:1995:MRB

[Hei95b] Georg Heinig. Matrix representations of Bezoutians. Linear Algebra and its Applications, 223/224(1–3):337– 354, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500048. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Heinig:1997:GIT

[Hei97] Georg Heinig. The group inverse of the transformation S(X) = AX − XB. Linear Algebra and its Applications, 257(1–3):321–342, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600165.

Heinig:1998:GCV

[Hei98a] Georg Heinig. Generalized Cauchy–Vandermonde matrices. Linear Algebra and its Applications, 270(1–3):45– 77, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700034.

Heinig:1998:MHO

[Hei98b]  Georg Heinig.  Matrices

with higher order displacement structure. Linear Algebra and its Applications, 278(1–3):295–301, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6085.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6085.

Henniger:1997:FS

[Hen97] J. P. Henniger. Factorization and similarity in GL(2, Z). Linear Algebra and its Applications, 251(1–3):223–237, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9600557.

Hershkowitz:1991:PCN

[Her91] Daniel Hershkowitz. Peak characteristic and nonnegative signature. Linear Algebra and its Applications, 147:55– 73, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1992:HCB

[Her92a]  Daniel Hershkowitz.  The

height characteristic of block triangular matrices. Linear Algebra and its Applications, 167:3–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1992:RDM

[Her92b] Daniel Hershkowitz. Recent directions in matrix stability. Linear Algebra and its Applications, 171:161–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1993:RBJ

[Her93] Daniel Hershkowitz. The relation between the Jordan structure of a matrix and its graph. Linear Algebra and its Applications, 184:55–69, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1994:PDG

[Her94] Daniel Hershkowitz. Paths in directed graphs and spectral properties of matrices. Linear Algebra and its Applications, 212/213:309–337, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Hershkowitz:1998:CS

[Her98] Daniel Hershkowitz. On cones and stability. Linear Algebra and its Applications, 275/276(1–3):249–259, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/

tree/store/laa/sub/1998/
275-276/1-3/6015.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6015. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Hershkowitz:1999:CSG

[Her99] Daniel Hershkowitz. The combinatorial structure of generalized eigenspaces — from nonnegative matrices to general matrices. Linear Algebra and its Applications, 302-303(1–3):173–191, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/36/article.
pdf.

Hoffmann:1994:ROP

[HF94] J. Hoffmann and P. A.

Fuhrmann. Remarks on orthogonal polynomials and balanced realizations. Linear Algebra and its Applications, 205/206:741–776, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffmann:1996:BRB

[HF96] J. W. Hoffmann and P. A.

Fuhrmann. On balanced realizations of bounded real and positive real functions. Linear Algebra and its Applications, 245(1–3):107–146, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400225.

Helmke:1998:TSR

[HF98] U. Helmke and P. A. Fuhrmann. Tangent spaces of rational matrix functions. Linear Algebra and its Applications, 271(1– 3):1–40, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700252.

Hara:1994:HSS

[HFK94] Shinji Hara, Hisaya Fujioka, and Pierre T. Kabamba. A hybrid state-space approach to sampled-data feedback control. Linear Algebra and its Applications, 205/206:675– 712, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:1991:JCF

[HH91] Yoo Pyo Hong and Roger A. Horn. The Jordan canonical form of a product of a Hermitian and a positive semidefinite matrix. Linear Algebra and its Applications, 147:373– 386, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:1992:CPT

[HH92] Desmond J. Higham and Nicholas J. Higman. Componentwise perturbation theory for linear systems with multiple right-hand sides. Linear Algebra and its Applications, 174:111–130, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1994:DSG

[HH94a] Georg Heinig and Frank Hellinger. Displacement structure of generalized inverse matrices. Linear Algebra and its Applications, 211:67–83, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Heinig:1994:DSP

[HH94b] Georg Heinig and Frank Hellinger. Displacement structure of pseudoinverses. Linear Algebra and its Applications, 197/198:623–649, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Hong:1992:LOP

[HHL92] Yoo-Pyo Hong, Roger A. Horn, and Chi-Kwong Li. Linear operators preserving tcongruence on matrices. Linear Algebra and its Applications, 175:191–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hershkowitz:1996:MNS

[HHNS96] Daniel Hershkowitz, Wenchao Huang, Michael Neumann, and Hans Schneider. Minimization of norms and the spectral radius of a sum of nonnegative matrices under diagonal equivalence. Linear Algebra and its Applications, 241/243(1–3):431–453, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500521. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Havas:1993:RBP

[HHR93] George Havas, Derek F. Holt, and Sarah Rees. Recognizing badly presented Z-modules. Linear Algebra and its Applications, 192:137–163, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Hershkowitz:1997:ESM

[HHS97] Daniel Hershkowitz, Alan J.

Hoffman, and Hans Schneider. On the existence of sequences and matrices with prescribed partial sums of elements. Linear Algebra and its Applications, 265(1–3):71–92, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600535.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600535.

Hu:1993:NPN

[HHT93] Shu-An Hu, James F. Hurley, and Tin-Yau Tam. Nonconvexity of the permanental numerical range. Linear Algebra and its Applications, 180: 121–??, February 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:1994:MSD

[Hig94] Nicholas J. Higham. The matrix sign decomposition and its relation to the polar decomposition. Linear Algebra and its Applications, 212/213: 3–20, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Higham:1995:CNT

[Hig95a] Desmond J. Higham. Condition numbers and their condition numbers. Linear Algebra and its Applications, 214(1–3):193–213, January 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300066.

Higman:1995:RAS

[Hig95b] D. G. Higman. Rank 5 association schemes and triality. Linear Algebra and its Applications, 226/228 (1–3):197–222, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500102.

Higham:1999:SBL

[Hig99] Nicholas J. Higham. Stability of block LDL^{T} factorization of a symmetric tridiagonal matrix. Linear Algebra and its Applications, 287 (1–3):181–189, January 15, 1999. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6193.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6193.

Hill:1992:CRU

[Hil92] R. O. Hill, Jr. Certain rank 2K updates act like rank K. Linear Algebra and its Applications, 177:173–174, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hirschbuhl:1990:CCL

[Hir90] Ralf Hirschbühl. Commutators in classical Lie algebras. Linear Algebra and its Applications, 142:91–111, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Howe:1990:EVN

[HJ90] Eric C. Howe and Charles R.

Johnson. Expected-value norms on matrices. Linear Algebra and its Applications, 139:21–29, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1992:KSB

[HJ92] Georg Heinig and Peter Jankowski. Kernel structure of block Hankel and Toeplitz matrices and partial realization. Linear Algebra and its Applications, 175:1–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hill:1997:EII

[HJK97] Richard O. Hill, Charles R. Johnson, and Brenda K. Kroschel. Extended interlacing intervals. Linear Algebra and its Applications, 254(1–3): 227–239, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600304. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Hillman:1997:DPN

[HJRS97] J. A. Hillman, B. R. F. Jefferies, W. J. Ricker, and B. Straub. Differential properties of the numerical range map of pairs of matrices. Linear Algebra and its Applications, 267(1–3):317–334, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700025.

Herceg:1993:CUS

[HK93a] Dragoslav Herceg and Nataša Krejić. On the convergence of the unsymmetric successive overrelaxation (USSOR) method. Linear Algebra and its Applications, 185:49–60, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:1993:FPB

[HK93b] Nicholas J. Higham and Philip A. Knight. Finite precision behavior of stationary iteration for solving singular systems. Linear Algebra and its Applications, 192:165–186, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Hartwig:1997:PEM

[HK97a] Robert E. Hartwig and Irving J. Katz. On products of EP matrices. Linear Algebra and its Applications, 252(1–3):339–345, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500693.

Harville:1997:AWE

[HK97b] David A. Harville and Oscar Kempthorne. An alternative way to establish the necessity part of the classical result on the statistical independence of quadratic forms. Linear Algebra and its Applications, 264(1–3):205–215, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600590.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600590.

Hanna:1997:RLA

[HKH97] L. A-M. Hanna, M. E. Khalifa, and S. S. Hassan. On representations of Lie algebras for quantized Hamiltonians. Linear Algebra and its Applications, 266(1–3):69– 79, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600631.

Hudelson:1996:LSC

[HKL96] Matthew Hudelson, Victor Klee, and David Larman. Largest j-simplices in dcubes: Some relatives of the Hadamard maximum determinant problem. Linear Algebra and its Applications, 241/243(1–3):519–598, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500541. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Hwang:1998:UBP

[HKM98] Suk-Geun Hwang, Arnold R.

Kräuter, and T. S. Michael. An upper bound for the permanent of a nonnegative matrix. Linear Algebra and its Applications, 281(1–3):259– 263, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6161.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6161.

Hsieh:1996:CFM

[HKS96a] Po-Fang Hsieh, Mitsuhiko Kohno, and Yasutaka Sibuya. Construction of a fundamental matrix solution at a singular point of the first kind by means of the SN decomposition of matrices. Linear Algebra and its Applications, 239(1–3):29– 76, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=239&issue=1-
                                        3&aid=9400165.

Hwang:1996:CCM

[HKS96b] Suk-Geun Hwang, Si-Ju Kim, and Seok-Zun Song. On convertible complex matrices. Linear Algebra and its Applications, 233(1–3):167– 173, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400065.

Han:1991:SCR

[HL91] De Guang Han and Balmohan V. Limaye. On a spectral characterization of rank one matrices. Linear Algebra and its Applications, 143:1– 6, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:1992:MEU

[HL92] Robert E. Hartwig and Raphael Loewy. Maximal elements under the three partial orders. Linear Algebra and its Applications, 175:39–??, October 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hale:1995:QCS

[HL95] Douglas Hale and George Lady. Qualitative comparative statics and audits of model performance. Linear Algebra and its Applications, 217(1–3): 141–154, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400125.

Hofmeister:1996:CSA

[HL96a] Thomas Hofmeister and Hanno Lefmann. Computing sparse approximations deterministically. Linear Algebra and its Applications, 240(1–3):9– 19, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400175.

Hua:1996:LME

[HL96b] Dai Hua and Peter Lancaster.

Linear matrix equations from an inverse problem of vibration theory. Linear Algebra and its Applications, 246(1–3):

31–47, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400311.

Huang:1997:ERT

[HL97] Liping Huang and Jianzhou Liu. The extension of Roth’s theorem for matrix equations over a ring. Linear Algebra and its Applications, 259(1–3): 229–235, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600286.

Horn:1999:MGM

[HL99] Roger A. Horn and Alexey K. Lopatin. The moment and Gram matrices, distinct eigenvalues and zeroes, and rational criteria for diagonalizability. Linear Algebra and its Applications, 299(1– 3):153–163, September 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/26/article.
pdf.

Hladnik:1999:SNB

[Hla99] Milan Hladnik. Schur norms of bicirculant matrices. Linear Algebra and its Applications, 286(1–3):261–272, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6282.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6282.

Horn:1996:DBG

[HLM96] Roger A. Horn, Chi-Kwong Li, and Dennis I. Merino. Distances between the graphs of matrices. Linear Algebra and its Applications, 240(1– 3):65–77, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400185.

Hjorth:1998:FMS

[HLMT98] Poul Hjorth, Petr Lisonĕk, Steen Markvorsen, and Carsten Thomassen. Finite metric spaces of strictly negative type. Linear Algebra and its Applications, 270(1–3):255– 273, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=270&issue=1-
                                        3&aid=9700242.

Hwang:1997:IBR

[HLP97] Tsung-Min Hwang, Wen-Wei Lin, and Daniel Pierce. Improved bound for rank revealing LU factorizations. Linear Algebra and its Applications, 261(1–3):173–186, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=261&issue=1-
                                        3&aid=9600397.

Hayden:1996:BMM

[HLWT96] T. L. Hayden, Jon Lee, Jim Wells, and P. Tarazaga. Block matrices and multispherical structure of distance matrices. Linear Algebra and its Applications, 247(1–3):203– 216, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500100.

Hwang:1992:RRF

[HLY92] Tsung-Min Hwang, Wen-Wei Lin, and Eugene K. Yang.

Rank revealing LU factorizations. Linear Algebra and its Applications, 175:115–??, October 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heuvers:1990:BCF

[HM90a] Konrad J. Heuvers and Daniel S. Moak. The Binet– Cauchy functional equation and nonsingular multiindexed matrices. Linear Algebra and its Applications, 140:197–215, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:1990:CSI

[HM90b] Roger A. Horn and Roy Mathias. Cauchy–Schwarz inequalities associated with positive semidefinite matrices. Linear Algebra and its Applications, 142:63–82, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Helmke:1992:SVD

[HM92a] U. Helmke and J. B. Moore.

Singular-value decomposition via gradient and selfequivalent flows. Linear Algebra and its Applications, 169:223–??, May 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:1992:BMG

[HM92b] Roger A. Horn and Roy Mathias. Block-matrix generalizations of Schur’s basic theorems on Hadamard products. Linear Algebra and its Applications, 172:337–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:1993:RCA

[HM93] Roger A. Horn and Dennis I. Merino. A real-coninvolutory analog of the polar decomposition. Linear Algebra and its Applications, 190:209–227, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hauke:1995:POS

[HM95a] Jan Hauke and Augustyn Markiewicz. On partial orderings on the set of rectangular matrices. Linear Algebra and its Applications, 219(1–3): 187–193, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300212.

Horn:1995:CEC

[HM95b] Roger A. Horn and Dennis I. Merino. Contragredient equivalence: A canonical form and some applications. Linear Algebra and its Applications, 214(1–3):43– 92, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300056.

Hodel:1996:SUS

[HM96] A. Scottedward Hodel and Pradeep Misra. Solution of underdetermined Sylvester equations in sensor array signal processing. Linear Algebra and its Applications, 249(1–3): 1–14, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500236.

Hartfiel:1998:SSM

[HM98] D. J. Hartfiel and Carl D.

Meyer. On the structure of stochastic matrices with a subdominant eigenvalue near 1. Linear Algebra and its Applications, 272(1–3):193– 203, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700333.

Horn:1999:JCF

[HM99] Roger A. Horn and Dennis I. Merino. The Jordan Canonical Forms of complex orthogonal and skew-symmetric matrices. Linear Algebra and its Applications, 302-303(1–3):411–421, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        49/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/49/article.
                                        pdf.

Hanke:1990:AKM

[HN90] Martin Hanke and Wilhelm Niethammer. On the acceleration of Kaczmarz’s method for inconsistent linear systems. Linear Algebra and its Applications, 130:83–98, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Hadjidimos:1993:DSC

[HN93a] A. Hadjidimos and Michael Neumann. On domains of superior convergence of the SSOR method to that of the SOR method. Linear Algebra and its Applications, 187: 67–85, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:1993:MIC

[HN93b] Apostolos Hadjidimos and Dimitrios Noutsos. On a matrix identity connecting iteration operators associated with a p-cyclic matrix. Linear Algebra and its Applications, 182: 157–??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanke:1993:GSS

[HN93c] Martin Hanke and Michael Neumann. The geometry of the set of scaled projections. Linear Algebra and its Applications, 190:137–148, September 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:1993:BEP

[HN93d] Robert E. Hartwig and Michael Neumann. Bounds on the exponent of primitive which depend on the spectrum and the minimal polynomial. Linear Algebra and its Applications, 184:103–??, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hassi:1994:PSI

[HN94] Seppo Hassi and Kenneth Nordström. On projections in a space with an indefinite metric. Linear Algebra and its Applications, 208/209:401– 417, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanke:1998:ITP

[HN98] Martin Hanke and James Nagy. Inverse Toeplitz preconditioners for ill-posed problems. Linear Algebra and its Applications, 284(1–3): 137–156, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6167.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6167.

Hadi:1999:FDT

[HN99] Ali S. Hadi and Hans Nyquist. Frechet distance as a tool for diagnosing multivariate data. Linear Algebra and its Applications, 289(1–3): 183–201, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6340.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6340.

Hanke:1991:SSO

[HNN91] Martin Hanke, Michael Neumann, and Wilhelm Niethammer. On the spectrum of the SOR operator for symmetric positive definite matrices. Linear Algebra and its Applications, 154/156:457–472, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:1990:AAA

[HNR90] Robert E. Hartwig, Michael Neumann, and Nicholas J. Rose. An algebraic-analytic approach to nonnegative bases. Linear Algebra and its Applications, 133:77–88, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holbrook:1999:OE

[HNRR99] John Holbrook, Eric Nordgren, Heydar Radjavi, and Peter Rosenthal. On the operator equation AX = XAX. Linear Algebra and its Applications, 295(1–3): 113–116, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6438.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6438.

Hadjidimos:1996:ECS

[HNT96] A. Hadjidimos, D. Noutsos, and M. Tzoumas. On the exact p-cyclic SSOR convergence domains. Linear Algebra and its Applications, 232(1– 3):213–236, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hannah:1991:PST

[HO91] John Hannah and K. C. O’Meara. Products of simultaneously triangulable idempotent matrices. Linear Algebra and its Applications, 149: 185–190, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:1993:KMA

[HO93] Chiou-Ming Huang and Dianne P. O’Leary. A Krylov multisplitting algorithm for solving linear systems of equations. Linear Algebra and its Applications, 194:9–29, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanzon:1998:OBB

[HO98] Bernard Hanzon and Raimund J. Ober. Overlapping blockbalanced canonical forms for various classes of linear systems. Linear Algebra and its Applications, 281(1–3): 171–225, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        281/1-3/6177.pdf;   http:
                                        //www.elsevier.com/cgi-

bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6177.

Hobart:1995:KCC

[Hob95] S. A. Hobart. Krein conditions for coherent configurations. Linear Algebra and its Applications, 226/228 (1–3):499–508, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500159.

Hochwald:1994:MMM

[Hoc94] Scott H. Hochwald. Multiplicative maps on matrices that preserve the spectrum. Linear Algebra and its Applications, 212/213:339– 351, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Hodel:1993:CSZ

[Hod93] A. Scottedward Hodel. Computation of system zeros with balancing. Linear Algebra and its Applications, 188/189:423– 436, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hofmeister:1997:TLE

[Hof97]  M. Hofmeister.  On the

two largest eigenvalues of trees. Linear Algebra and its Applications, 260(1–3):43– 59, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600249.

Hoffman:1998:WOD

[Hof98a] Alan Hoffman. What Olga did for me. Linear Algebra and its Applications, 280(1): 13, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Hoffmann:1998:GHA

[Hof98b] Walter Hoffmann. The Gauss– Huard algorithm and LU factorization. Linear Algebra and its Applications, 275/276 (1–3):281–286, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6026.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6026. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Hogben:1998:CIM

[Hog98a]  Leslie Hogben.  Comple-

tions of inverse M-matrix patterns. Linear Algebra and its Applications, 282(1–3): 145–160, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6175.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6175.

Hogben:1998:CMP

[Hog98b] Leslie Hogben. Completions of M-matrix patterns. Linear Algebra and its Applications, 285(1–3):143–152, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6237.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6237.

Holbrook:1992:SVN

[Hol92] John A. Holbrook. Spectral variation of normal matrices. Linear Algebra and its Applications, 174:131–??, September 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holbrook:1995:IEP

[Hol95] John A. Holbrook. Interpenetration of ellipsoids and the polynomial bound of a matrix. Linear Algebra and its Applications, 229(1–3):151– 166, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=229&issue=1-
                                        3&aid=9300350.

Holmquist:1996:VVH

[Hol96] Björn Holmquist. The dvariate vector Hermite polynomial of order k. Linear Algebra and its Applications, 237/238(1–3):155– 190, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500595. Special

issue honoring Calyampudi Radhakrishna Rao.

Holtz:1997:INM

[Hol97] O. V. Holtz. On indecomposable normal matrices in spaces with indefinite scalar product. Linear Algebra and its Applications, 259(1–3):155– 168, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600247.

Holtz:1999:AGM

[Hol99] Olga Holtz. Not all GKK τ-matrices are stable. Linear Algebra and its Applications, 291(1–3):235–244, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6353.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6353.

Hong:1990:HCF

[Hon90] Yoopyo Hong. A Hermitian canonical form for complex matrices under consimilarity. Linear Algebra and its Applications, 133:1–19, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:1991:CFU

[Hon91] Yoo Pyo Hong. A canonical form under ϕ-equivalence. Linear Algebra and its Applications, 147:501–549, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hongyuan:1992:NAC

[Hon92] Zha Hongyuan. A numerical algorithm for computing the restricted singular value decomposition of matrix triplets. Linear Algebra and its Applications, 168:1–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:1998:BDM

[Hon98] Shaofang Hong. Bounds for determinants of matrices associated with classes of arithmetical functions. Linear Algebra and its Applications, 281 (1–3):311–322, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        281/1-3/6173.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6173.

Horn:1995:NBH

[Hor95]  Roger A. Horn.  Norm

bounds for Hadamard products and an arithmetic– geometric mean inequality for unitarily invariant norms. Linear Algebra and its Applications, 223/224(1–3):355– 361, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400034. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Horne:1997:LBS

[Hor97] Bill G. Horne. Lower bounds for the spectral radius of a matrix. Linear Algebra and its Applications, 263(1–3):261– 273, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600539.

Hwang:1998:RCV

[HOR98] Frank K. Hwang, Shmuel Onn, and Uriel G. Rothblum. Representations and characterizations of vertices of bounded-shape partition polytopes. Linear Algebra and its Applications, 278(1– 3):263–284, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6092.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1998&volume=278&issue=
                                        1-3&aid=6092.

Holbrook:1996:MSD

[HOŠ96] John Holbrook, Matjaž Omladič, and Peter Šemrl. Maximal spectral distance. Linear Algebra and its Applications, 249(1–3):197–205, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500353.

Hartwig:1996:SCP

[HOŠS96] Robert E. Hartwig, Matjaž Omladič, Peter Šemrl, and George P. H. Styan. On some characterizations of pairwise star orthogonality using rank and dagger additivity and subtractivity. Linear Algebra and its Applications, 237/238(1–3):499– 507, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9400133. Special

issue honoring Calyampudi Radhakrishna Rao.

Hou:1991:OIL

[Hou91] Jin Chuan Hou. On operator inequalities and linear combinations of operators. Linear Algebra and its Applications, 153:35–51, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:1996:COP

[Hou96] Jinchuan Hou. Completion of operator partial matrices to projections. Linear Algebra and its Applications, 246(1–3): 71–82, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400336.

Hou:1999:RBC

[Hou99]  Su-Mei Hou.  The rela-

tionship between the class A_{2}(R, S) of (0, 1, 2)-matrices and the collection of constellation matrices. Linear Algebra and its Applications, 290(1–3):49–59, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6305.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6305.

Hedayat:1990:STO

[HP90] A. S. Hedayat and H. Pesotan.

Strongly threefold orthogonal matrices with statistical applications. Linear Algebra and its Applications, 136:1–23, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hentzel:1991:BAG

[HP91] Irvin Roy Hentzel and Luiz Antonio Peresi. Bernstein algebras given by symmetric bilinear forms. Linear Algebra and its Applications, 145:213– 219, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:1992:LBS

[HP92] Y. P. Hong and C.-T Pan. A lower bound for the smallest singular value. Linear Algebra and its Applications, 172: 27–??, July 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hiai:1993:GTT

[HP93] Fumio Hiai and Denes Petz. The Golden-Thompson trace inequality is complemented. Linear Algebra and its Applications, 181:153–??, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanzon:1996:FSM

[HP96] Bernard Hanzon and Ralf L. M. Peeters. A Faddeev sequence method for solving Lyapunov and Sylvester equations. Linear Algebra and its Applications, 241/243(1–3):401–430, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500683. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Hedayat:1999:COS

[HP99] A. S. Hedayat and H. Pesotan.

Column optimal strongly threefold orthogonal matrices in a class index eight. Linear Algebra and its Applications, 298(1–3):171–191, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/28/article.
pdf.

Havlicek:1998:LGI

[HPP98] Miloslav Havlícek, Jiří Patera, and Edita Pelantova. On Lie gradings. II. Linear Algebra and its Applications, 277(1–3):97–125, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        277/1-3/6040.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6040.

Hall:1997:PCR

[HPUV97] Frank Hall, Alex Pothen, Frank Uhlig, and Paul Van Dooren. Preface and conference report. Linear Algebra and its Applications, 254(1– 3):1–5, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffmann:1994:DMB

[HPW94] J. Hoffmann and D. Prätzel-Wolters. Dipolynomial minimal bases and linear systems in AR representation. Linear Algebra and its Applications, 205/206:777–803, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffmann:1998:BCF

[HPWZ98] Jörg Hoffmann, Dieter Prätzel-Wolters, and Eva Zerz. A balanced canonical form for discrete-time minimal systems using characteristic maps. Linear Algebra and its Applications, 277(1–3):63–81, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/

tree/store/laa/sub/1998/
277/1-3/6064.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6064.

Hu:1998:SPH

[HQ98] Tingshu Hu and Li Qiu. On structured perturbation of Hermitian matrices. Linear Algebra and its Applications, 275/276(1–3):287–314, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6059.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6059. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Hung:1990:AHN

[HR90] Ming S. Hung and Walter O.

Rom. An application of the Hermite normal form in integer programming. Linear Algebra and its Applications, 140:163–179, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:1992:KSM

[HR92] D. Y. Hu and L. Reichel.

Krylov-subspace methods for the Sylvester equation. Linear Algebra and its Applications, 172:283–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hari:1993:MPA

[HR93a] Vjeran Hari and Noah H. Rhee. A matrix pair of an almost diagonal skewsymmetric matrix and a symmetric positive definite matrix. Linear Algebra and its Applications, 195:117– 144, December 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasselbarth:1993:ELC

[HR93b] W. Hasselbarth and E. Ruch. On the extension of linear contractions: Some results of the Hahn-Banach type. Linear Algebra and its Applications, 184:145–??, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffman:1994:PCR

[HR94] Alan Hoffman and Uriel G. Rothblum. A proof of the convexity of the range of a nonatomic vector measure using linear inequalities. Linear Algebra and its Applications, 199:373–379, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:1995:RSC

[HR95] Georg Heinig and Karla Rost.

Recursive solution of Cauchy– Vandermonde systems of equations. Linear Algebra and its Applications, 218(1–3):59– 72, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300157.

Hartfiel:1998:CIP

[HR98a] D. J. Hartfiel and Uriel G.

Rothblum. Convergence of inhomogeneous products of matrices and coefficients of ergodicity. Linear Algebra and its Applications, 277 (1–3):1–9, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6097.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6097.

Heinig:1998:DRT

[HR98b] Georg Heinig and Karla Rost. DFT representations of Toeplitz-plus-Hankel Bézoutians with application to fast matrix-vector multiplication. Linear Algebra and its Applications, 284(1–3): 157–175, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6195.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6195.

Heinig:1998:RTP

[HR98c] Georg Heinig and Karla Rost. Representations of Toeplitz-plus-Hankel matrices using trigonometric transformations with application to fast matrix-vector multiplication. Linear Algebra and its Applications, 275/276(1– 3):225–248, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6029.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6029. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Haviv:1992:TEE

[HRR92] Moshe Haviv, Ya’acov Ritov, and Uriel G. Rothblum. Taylor expansions of eigenvalues of perturbed matrices with applications to spectral radii of nonnegative matrices. Linear Algebra and its Applications, 168:159–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:1998:EIE

[HRS98] Roger A. Horn, Noah H.

Rhee, and Wasin So. Eigenvalue inequalities and equalities. Linear Algebra and its Applications, 270(1–3):29– 44, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700031.

Hadley:1992:SNQ

[HRW92] S. W. Hadley, F. Rendl, and H. Wolkowicz. Symmetrization of nonsymmetric quadratic assignment problems and the Hoffman-Wielandt inequality. Linear Algebra and its Applications, 167:53–64, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hayden:1999:MCD

[HRW99] Thomas L. Hayden, Robert Reams, and James Wells. Methods for constructing distance matrices and the inverse eigenvalue problem. Linear Algebra and its Applications, 295(1–3):97–112, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6412.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6412.  See note

[JM12].

Herzel:1991:QCM

[HRZ91] Stefano Herzel, Maria Cristina Recchioni, and Francesco Zirilli. A quadratically convergent method for linear programming. Linear Algebra and its Applications, 152:255– 289, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Hartfiel:1990:NSR

[HS90a] D. J. Hartfiel and E. Seneta. A note on semigroups of regular stochastic matrices. Linear Algebra and its Applications, 141:47–51, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Henderson:1990:GDM

[HS90b] Harold V. Henderson and Shayle R. Searle. Generalized dispersion matrices for covariance structural analysis. Linear Algebra and its Applications, 127:543–565, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Helmke:1995:CPM

[HS95] Uwe Helmke and Mark A.

Shayman. Critical points of matrix least squares distance functions. Linear Algebra and its Applications, 215(1–3):1– 19, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300070.

Holtz:1996:CNO

[HS96a] O. V. Holtz and V. A. Strauss.

Classification of normal operators in spaces with indefinite scalar product of rank 2. Linear Algebra and its Applications, 241/243(1–3):455– 517, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500605. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Hu:1996:SCM

[HS96b] Shu-An Hu and Eugene Spiegel. The similarity class of a matrix. Linear Algebra and its Applications, 233(1–3):231–241, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400083.

Holtz:1997:CNO

[HS97a] O. V. Holtz and V. A. Strauss. On classification of normal operators in real spaces with indefinite scalar product. Linear Algebra and its Applications, 255(1–3): 113–155, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=255&issue=1-
                                        3&aid=9500770.

Hwang:1997:FPD

[HS97b] Suk-Geun Hwang and Sun-Jeong Shin. A face of the polytope of doubly stochastic matrices associated with certain matrix expansions. Linear Algebra and its Applications, 253(1–3):125–140, March 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500780.

Heyouni:1998:VSP

[HS98] M. Heyouni and H. Sadok.

On a variable smoothing procedure for Krylov subspace methods. Linear Algebra and its Applications, 268(1–3):131–149, January 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700037.

Hong:1999:SLB

[HS99] Yuan Hong and Jin-Long Shu. Sharp lower bounds of the least eigenvalue of planar graphs. Linear Algebra and its Applications, 296(1–3):227–232, July 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/30/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/30/article.
pdf.

Heinig:1997:PCM

[HST97] G. Heinig, L. A. Sakhnovich, and I. F. Tidniuk. Paired Cauchy matrices. Linear Algebra and its Applications, 251(1–3):189–214, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500607.

Heemels:1999:RCP

[HSW99] W. P. M. H. Heemels, J. M. Schumacher, and S. Weiland. The rational complementarity problem. Linear Algebra and its Applications, 294 (1–3):93–135, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6407.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6407.

Shih:1998:ISR

[hSyLtP98] Mau hsiang Shih, Yung yih Lur, and Chin tzong Pang. An inequality for the spectral radius of an interval matrix. Linear Algebra and its Applications, 274(1–3):27– 36, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700238.

Hayden:1993:DMR

[HT93] T. L. Hayden and P. Tarazaga.

Distance matrices and regular figures. Linear Algebra and its Applications, 195:9–16, December 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hodel:1996:NSL

[HTP96] A. Scottedward Hodel, Bruce Tenison, and Kameshwar R. Poolla. Numerical solution of the Lyapunov equation by approximate power iteration. Linear Algebra and its Applications, 236(1–3):205– 230, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400155.

Hu:1995:PDC

[Hu95]  H. Hu.  Positive definite

constrained least-squares estimation of matrices. Linear Algebra and its Applications, 229(1–3):167–174, October 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=229&issue=1-
                                        3&aid=9400024.

Hu:1996:MPH

[Hu96] Shu-An Hu. Maximum permanent and Hermitian matrices. Linear Algebra and its Applications, 235(1–3):93– 105, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400121.

Huang:1990:CBM

[Hua90] Dao De Huang. On circulant Boolean matrices. Linear Algebra and its Applications, 136:107–117, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:1991:SGJ

[Hua91] Wen Chao Huang. On the singular graph and Jordan diagram of strictly lower triangular matrices and M-matrices. Linear Algebra and its Applications, 143:145–169, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:1993:SSC

[Hua93] Wenchao Huang. On the sandwich semigroups of circulant Boolean matrices. Linear Algebra and its Applications, 179:135–??, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:1995:NGD

[Hua95]  Tin-Zhu Huang.  A note

on generalized diagonally dominant matrices. Linear Algebra and its Applications, 225(1):237–242, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300368.

Huang:1996:MEQ

[Hua96a] Liping Huang. The matrix equation AXB − GXD = E over the quaternion field. Linear Algebra and its Applications, 234(1–3):197–208, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400103.

Huang:1996:MWB

[Hua96b] Wenchao Huang. Matrices which belong to an idempotent in a sandwich semigroup of circulant Boolean matrices.

Linear Algebra and its Applications, 249(1–3):157–167, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500325.

Huang:1996:CFP

[Hua96c] Ying Huang. A canonical form for pencils of matrices with applications to asymptotic linear programs. Linear Algebra and its Applications, 234(1–3):97– 123, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9500090.

Huang:1999:CTL

[Hua99] Wenxue Huang. A characterization of triangularizability of a linear associative algebra. Linear Algebra and its Applications, 299(1–3):165–169, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/27/article.
pdf.

Huckle:1998:CGS

[Huc98]  Thomas Huckle.  Com-

putations with Gohberg– Semencul-type formulas for Toeplitz matrices. Linear Algebra and its Applications, 273(1–3):169–198, April 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700372.

Humphries:1997:CGM

[Hum97] Stephen P. Humphries. Cogrowth of groups and a matrix of Redheffer. Linear Algebra and its Applications, 265(1–3):101–117, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600593.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600593.

Hunter:1990:PFG

[Hun90] Jeffrey J. Hunter. Parametric forms for generalized inverses of Markovian kernels and their applications. Linear Algebra and its Applications, 127:71– 84, 1990. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Hurkens:1990:BCS

[Hur90] C. A. J. Hurkens. Blowing up convex sets in the plane. Linear Algebra and its Applications, 134:121–128, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huylebrouck:1996:GIS

[Huy96] D. Huylebrouck. The generalized inverse of a sum with radical element: Applications. Linear Algebra and its Applications, 246(1–3):159– 175, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=246&issue=1-
                                        3&aid=9400350.

Heitmann:1991:DSI

[HW91] Raymond Heitmann and Roger Wiegand. Direct sums of ideals. Linear Algebra and its Applications, 157:21–36, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hill:1992:KRK

[HW92] Richard D. Hill and Steven R. Waters. On kappa-real and kappa-Hermitian matrices. Linear Algebra and its Applications, 169:17–??, May 1992.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Helton:1993:SHO

[HW93] J. William Helton and Hugo J.

Woerdeman. Symmetric Hankel operators: Minimal norm extensions and eigenstructures. Linear Algebra and its Applications, 185:1–19, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:1990:MPC

[Hwa90a] Suk Geun Hwang. Maximum permanents on certain classes of nonnegative matrices. Linear Algebra and its Applications, 132:103–113, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:1990:SNP

[Hwa90b] Suk Geun Hwang. Some nontrivial permanental mates. Linear Algebra and its Applications, 140:89–100, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:1995:MPS

[Hwa95] Suk Geun Hwang. Minimizing the permanent over some faces of the polytope of doubly stochastic matrices. Linear Algebra and its Applications, 219(1–3):195– 205, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300214.

Hwang:1999:MGH

[Hwa99] Suk-Geun Hwang. Majorization via generalized Hessenberg matrices. Linear Algebra and its Applications, 292(1–3):127–138, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6372.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6372.

Hoffman:1995:NAR

[HWH95] A. J. Hoffman, P. Wolfe, and M. Hofmeister. A note on almost regular matrices. Linear Algebra and its Applications, 226/228 (1–3):105–108, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400307.

Hayden:1991:CDM

[HWLT91] T. L. Hayden, Jim Wells, Wei Min Liu, and P. Tarazaga. The cone of distance matrices. Linear Algebra and its Applications, 144:153–169, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haukkanen:1997:SD

[HWS97] Pentti Haukkanen, Jun Wang, and Juha Sillanpää. On Smith’s determinant. Linear Algebra and its Applications, 258(1–3):251–269, June ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600192.

Helmberg:1993:FLM

[HWV93] Gilbert Helmberg, Peter Wagner, and Gerhard Veltkamp. On Faddeev-Leverrier’s method for the computation of the characteristic polynomial of a matrix and of eigenvectors. Linear Algebra and its Applications, 185:219–233, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hussin:1990:MAS

[HWZ90] V. Hussin, P. Winternitz, and H. Zassenhaus. Maximal Abelian subalgebras of complex orthogonal Lie algebras. Linear Algebra and its Applications, 141:183–220, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hussin:1992:MAS

[HWZ92] V. Hussin, P. Winternitz, and H. Zassenhaus. Maximal Abelian subalgebras of pseudoorthogonal Lie algebras. Linear Algebra and its Applications, 173:125–??, August 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hall:1992:CNB

[HX92] Charles A. Hall and Ye Xiu. Construction of null bases for the divergence operator associated with incompressible Navier–Stokes equations. Linear Algebra and its Applications, 171:9–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xiang:1998:WBD

[hXyY98] Shu huang Xiang and Zhao yong You. Weak block diagonally dominant matrices, weak block H-matrix and their applications. Linear Algebra and its Applications, 282(1–3):263–274, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

tree/store/laa/sub/1998/
282/1-3/6189.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6189.

Hadjidimos:1991:SRR

[HY91] A. Hadjidimos and A. K.

Yeyios. Some recent results on the modified SOR theory. Linear Algebra and its Applications, 154/156:5–21, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:1996:SNM

[HY96] A. Hadjidimos and A. K.

Yeyios. Some notes on multisplitting methods and m-step preconditioners for linear systems. Linear Algebra and its Applications, 248(1–3):277– 301, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500247.

Hong:1998:SRC

[HZ98] Shao-Fang Hong and Mou-Cheng Zhang. On the structure rank of a class of structure matrices. Linear Algebra and its Applications, 268(1–3):355–364, January 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700053.

Horn:1999:ISL

[HZ99] Roger A. Horn and Xingzhi Zhan. Inequalities for C-S seminorms and Lieb functions. Linear Algebra and its Applications, 291(1–3): 103–113, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6348.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6348.

Ikebe:1996:EPI

[IAMC96] Yasuhiko Ikebe, Nobuyoshi Asai, Yoshinori Miyazaki, and DongSheng Cai. The eigenvalue problem for infinite complex symmetric tridiagonal matrices with application. Linear Algebra and its Applications, 241/243(1–3):599– 618, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=241-243&issue=
1-3&aid=9500699. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Ikramov:1994:TIE

[IC94] Khakim D. Ikramov and Vadim N. Chugunov. On the Teng inverse eigenvalue problem. Linear Algebra and its Applications, 208/209:397– 399, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Iusem:1992:CSF

[ID92] Alfredo N. Iusem and Alvaro R. De Pierro. A characterization of the set of fixed points of some smoothed operators. Linear Algebra and its Applications, 177:1–??, December 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Idris:1996:UGM

[Idr96] Ismail M. Idris. On unitary groups modulo infinitesimals. Linear Algebra and its Applications, 235(1–3):63– 76, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400115.

Idris:1999:TES

[Idr99] Ismail M. Idris. On transformations of elliptic spaces. Linear Algebra and its Applications, 301(1–3):51–62, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/27/article.
pdf.

Ivanov:1998:PPD

[IEs98] Ivan G. Ivanov and Salah M. El-sayed. Properties of positive definite solutions of the equation X + A ∗ X − 2A = I. Linear Algebra and its Applications, 279(1– 3):303–316, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6121.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6121.

Ito:1999:DMS

[IHM99] Yoshimichi Ito, Satoshi Hattori, and Hajime Maeda. On the decomposition of a matrix into the sum of stable matrices. Linear Algebra and its Applications, 297(1–3):177–182, August 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/104/17/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/104/17/27/article.
pdf.

Ito:1998:DFI

[II98] Naoharu Ito and Hiroshi Inaba. Dynamic feedback (A, B)-invariant submodules for linear systems over commutative Noetherian domains. Linear Algebra and its Applications, 282(1–3):123–129, October 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6169.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6169.

Ikebe:1993:EPI

[IKF^{+}93] Yasuhiko Ikebe, Yasushi Kikuchi, Issei Fujishiro, Nobuyoshi Asai, Kouichi Takanashi, and Minoru Harada. The eigenvalue problem for infinite compact complex symmetric matrices with application to the numerical computation of complex zeros of J_{0}(z)−iJ_{1}(z) and of Bessel functions J_{m}(z) of any real order m. Linear Algebra and its Applications, 194:35–70, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:1994:SPG

[Ikr94] Khakim D. Ikramov. A simple proof of the generalized Schur inequality. Linear Algebra and its Applications, 199:143–149, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:1995:ESP

[Ikr95] K. D. Ikramov. Erratum: “A simple proof of the generalized Schur inequality” [Linear Algebra Appl. 199 (1994), 143– 149, MR 95i:15024]. Linear Algebra and its Applications, 220(1–3):4, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:1999:APR

[Ikr99] Khakim D. Ikramov. An algebraic proof of a result by Gonzaga and Lara. Linear Algebra and its Applications, 299(1–3):191–194, September 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/112/21/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/112/21/30/article.
                                        pdf.

Iltyakov:1997:PIJ

[Ilt97] A. V. Iltyakov. On polynomial identities of Jordan pairs of rectangular matrices. Linear Algebra and its Applications, 260(1–3):257– 271, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600309.

Im:1997:NEH

[Im97] Eric Iksoon Im. Narrower eigenbounds for Hadamard products. Linear Algebra and its Applications, 264 (1–3):141–144, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700017.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700017.

Insel:1997:LS

[Ins97] Arnold J. Insel. Levels of subnormality. Linear Algebra and its Applications, 262(1–3):27– 53, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600466.

Ionescu:1996:GCT

[IO96] Vlad Ionescu and Cristian Oară. Generalized continuoustime Riccati theory. Linear Algebra and its Applications, 232(1–3):111–130, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ionescu:1996:RDT

[Ion96]  Vlad Ionescu.  Reverse

discrete-time Riccati equation and extended Nehari’s problem. Linear Algebra and its Applications, 236(1–3):59– 94, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400129.

Ionescu:1998:TVD

[Ion98] Vlad Ionescu. Time-varying discrete Riccati equation in terms of Ben Artzi-Gohberg dichotomy. Linear Algebra and its Applications, 277(1– 3):313–336, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6042.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6042.

Iwata:1998:PPS

[IOv98] S. Iwata, D. D. Olesky, and P. van den Driessche. Patterns that preserve sparsity in orthogonal factorization. Linear Algebra and its Applications, 268(1–3):345– 354, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700052.

Ionin:1995:EFS

[IS95] Yury J. Ionin and Mohan S.

Shrikhande. Equidistant families of sets. Linear Algebra and its Applications, 226/228 (1–3):223–235, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500103.

Ishibashi:1995:IEE

[Ish95]  Hiroyuki Ishibashi.  Invo-

lutary expressions for elements in GL_{n}(Z) and SL_{n}(Z). Linear Algebra and its Applications, 219(1–3):165–177, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=219&issue=1-
                                        3&aid=9300206.

Ito:1997:NSA

[Ito97] Hisashi Ito. A new statement about the theorem determining the region of eigenvalues of stochastic matrices. Linear Algebra and its Applications, 267(1–3):241–246, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=267&issue=1-
                                        3&aid=9700018.

Ionescu:1992:CSD

[IW92] Vlad Ionescu and Martin Weiss. On computing the stabilizing of the discrete-time Riccati equation. Linear Algebra and its Applications, 174: 229–??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ionescu:1993:CDT

[IW93] Vlad Ionescu and Martin Weiss. Continuous and discrete-time Riccati theory: A Popov-Function approach.

Linear Algebra and its Applications, 193:173–209, November 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Inaba:1996:BDL

[IW96] Hiroshi Inaba and Weiben Wang. Block decoupling for linear systems over rings. Linear Algebra and its Applications, 241/243(1–3):619– 634, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500685. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Iwata:1998:BTS

[Iwa98] Satoru Iwata. Block triangularization of skew-symmetric matrices. Linear Algebra and its Applications, 273(1–3): 215–226, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700375.

Jonckheere:1998:DTN

[JAG98] Edmond A. Jonckheere, Farooq Ahmad, and Eugene Gutkin. Differential topology of numerical range. Linear Algebra and its Applications, 279(1–3):227–254, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        279/1-3/6119.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6119.

James:1991:GOG

[Jam91a] Donald G. James. Generators for orthogonal groups of unimodular lattices. Linear Algebra and its Applications, 157: 101–111, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

James:1991:ORC

[Jam91b]  Douglas James.  Order-

reducing conjugate gradients versus block AOR for constrained least-squares problems. Linear Algebra and its Applications, 154/156:23–43, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

James:1994:HAI

[Jam94] Gordon James. Hecke algebras and immanants. Linear Algebra and its Applications, 197/198:659–670, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Jang:1997:CLS

[Jan97a]  Ho-Jong Jang.  A con-

strained least-squares approach to the rapid reanalysis of structures. Linear Algebra and its Applications, 265(1–3):185–202, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600602.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600602.

Jansson:1997:CEB

[Jan97b] C. Jansson. Calculation of exact bounds for the solution set of linear interval systems. Linear Algebra and its Applications, 251(1–3):321– 340, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9600681.

Jarvis:1990:DNE

[Jar90] Tyler J. Jarvis. A dominant negative eigenvalue of a matrix of Redheffer. Linear Algebra and its Applications, 142: 141–152, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1990:SRP

[JB90] Charles R. Johnson and Rafael Brú. The spectral radius of a product of nonnegative matrices. Linear Algebra and its Applications, 141: 227–240, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jianming:1992:PAB

[JB92] Miao Jianming and Adi Ben-Israel. On principal angles between subspaces in R^{n}. Linear Algebra and its Applications, 171:81–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jbilou:1995:PMM

[Jbi95] Khalide Jbilou. Projectionminimization methods for nonsymmetric linear systems. Linear Algebra and its Applications, 229(1–3):101–125, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=229&issue=1-
                                        3&aid=9300348.

Jelonek:1992:MSL

[Jel92] Zbigniew Jelonek. Maximal sets of linearly independent vectors in a free module over a commutative ring. Linear Algebra and its Applications, 168:137–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jensen:1991:VEM

[Jen91] D. R. Jensen. Vector efficiency in multiparameter estimation. Linear Algebra and its Applications, 151:143–155, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jensen:1994:MPM

[Jen94] D. R. Jensen. Minimal properties of Moore–Penrose inverses. Linear Algebra and its Applications, 196:175– 182, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jensen:1996:SDV

[Jen96]  D. R. Jensen.  Struc-

tured dispersion and validity in linear inference. Linear Algebra and its Applications, 249(1–3):189–196, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500354.

Jensen:1997:BMN

[Jen97] D. R. Jensen. Bounds on Mahalanobis norms and their applications. Linear Algebra and its Applications, 264 (1–3):127–139, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9700235.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9700235.

Jeurissen:1995:SSL

[Jeu95] R. H. Jeurissen. Special sets of lines in PG(3, 2). Linear Algebra and its Applications, 226/228(1–3):617–638, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500200.

Jiang:1992:NDS

[Jia92] Erxiong Jiang. A note on the double-shift QL algorithm. Linear Algebra and its Applications, 171:121–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:1994:BSS

[Jia94] Er Xiong Jiang. Bounds for the smallest singular value of a Jordan block with an application to eigenvalue perturbation. Linear Algebra and its Applications, 197/198:691– 707, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Jiang:1996:ICR

[Jia96] Zhi-Ming Jiang. On the indices of convergence of reducible Boolean matrices with the generalized period f_{0} ≥ 2. Linear Algebra and its Applications, 248(1–3):355–374, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500252.

Jia:1997:SMS

[Jia97a] Rong-Qing Jia. Symmetric magic squares and multivariate splines. Linear Algebra and its Applications, 250(1–3): 69–103, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500451.

Jia:1997:RIA

[Jia97b]  Zhongxiao Jia.  Refined

iterative algorithms based on Arnoldi’s process for large unsymmetric eigenproblems. Linear Algebra and its Applications, 259(1–3): 1–23, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600238.

Jia:1998:RIA

[Jia98a] Zhongxiao Jia. A refined iterative algorithm based on the block Arnoldi process for large unsymmetric eigenproblems. Linear Algebra and its Applications, 270(1–3):171– 189, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=270&issue=1-
                                        3&aid=9700023.

Jiang:1998:CEP

[Jia98b]  Erxiong Jiang.  Challeng-

ing eigenvalue perturbation problems. Linear Algebra and its Applications, 278(1– 3):303–307, July 15, 1998.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6086.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6086.

Jia:1999:PCA

[Jia99] Zhongxiao Jia. Polynomial characterizations of the approximate eigenvectors by the refined Arnoldi method and an implicitly restarted refined Arnoldi algorithm. Linear Algebra and its Applications, 287(1–3):191–214, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6307.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6307.

Jinadasa:1990:IMS

[Jin90] K. G. Jinadasa. The information matrix of a sample of observations with missing data from a multivariate normal distribution with a covariance structure. Linear Algebra and its Applications, 131: 9–15, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jin:1998:IRU

[Jin98] Yun Yuan Jin. Iterative refinement using splitting methods. Linear Algebra and its Applications, 273(1–3):199– 214, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=273&issue=1-
                                        3&aid=9700374.

Jurg:1990:WCM

[JJPT90] A. P. Jurg, M. J. M. Jansen, T. Parthasarathy, and S. H. Tijs. On weakly completely mixed bimatrix games. Linear Algebra and its Applications, 141:61–74, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Juang:1994:CTM

[JL94] Jonq Juang and Mei Tzu Lee. Comparison theorems for the matrix Riccati equation. Linear Algebra and its Applications, 196:183–191, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:1997:UBS

[JL97] Erxiong Jiang and Peter C. B. Lam. An upper bound for the spectral condition number of a diagonalizable matrix. Linear Algebra and its Applications, 262(1–3):165– 178, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600000.

Johnson:1998:LIM

[JL98a] Charles R. Johnson and Michael Lundquist. Local inversion of matrices with sparse inverses. Linear Algebra and its Applications, 277 (1–3):33–39, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6061.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6061.

Junghanns:1998:UCF

[JL98b] Peter Junghanns and Uwe Luther. Uniform convergence of a fast algorithm for Cauchy singular integral equations. Linear Algebra and its Applications, 275/276(1– 3):327–347, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6016.pdf;

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6016. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Johnson:1996:MSR

[JLOvdD96] Charles R. Johnson, Raphael Loewy, D. D. Olesky, and P. van den Driessche. Maximizing the spectral radius of fixed trace diagonal perturbations of nonnegative matrices. Linear Algebra and its Applications, 241/243(1–3):635– 654, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500258. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Johnson:1993:VMC

[JM93] Charles R. Johnson and John S. Maybee. Vanishing minor conditions for inverse zero patterns. Linear Algebra and its Applications, 178: 1–??, January 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1997:RDU

[JM97] Charles R. Johnson and Jeremy Miller. Rank decomposition under combinatorial constraints. Linear Algebra and its Applications, 251(1–3):97–104, January 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500548.

Jbilou:1999:MRI

[JM99] K. Jbilou and A. Messaoudi.

Matrix recursive interpolation algorithm for block linear systems. Linear Algebra and its Applications, 294 (1–3):137–154, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6403.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6403.

Jaklic:2012:NSC

[JM12] Gasper Jaklic and Jolanda Modic. A note on “Methods for constructing distance matrices and the inverse eigenvalue problem”. Linear Algebra and its Applications, 437(11):2781–2792, December 1, 2012. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379512005162.  See

[HRW99].

Johnson:1997:NSP

[JMOV97] Charles R. Johnson, John S. Maybee, D. D. Olesky, and P. Van Den Driessche. Nested sequences of principal minors and potential stability. Linear Algebra and its Applications, 262(1–3):243–257, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600511.

Johnson:1995:SPA

[JMS95] Charles R. Johnson, William D. McCuaig, and David P. Stanford. Sign patterns that allow minimal semipositivity. Linear Algebra and its Applications, 223/224(1–3):363– 373, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400181. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Johnson:1990:YTG

[JN90] Charles R. Johnson and Peter Nylen. Yamamoto’s theorem for generalized singular values. Linear Algebra and its Applications, 128:147–158, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1991:MPN

[JN91] Charles R. Johnson and Peter Nylen. Monotonicity properties of norms. Linear Algebra and its Applications, 148:43– 58, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1993:SP

[JN93] Charles R. Johnson and Peter M. Nylen. The sprinkling problem. Linear Algebra and its Applications, 181:131–152, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jodar:1991:SAD

[Jód91] Lucas Jódar. Solving algebraic and differential Riccati operator equations. Linear Algebra and its Applications, 144:71– 83, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jodar:1992:NSO

[Jod92] Lucas Jodar. On the nonhomogeneous second-order Euler operator differential equation: Explicit solutions. Linear Algebra and its Applications, 177:145–??, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1998:OMT

[Joh98] Charles R. Johnson. Olga, matrix theory and the Taussky unification problem. Linear Algebra and its Applications, 280(1):39–49, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). With the assistance of Shaun Fallat. Special issue in memory of Olga Taussky Todd.

Jondrup:1995:ADU

[Jøn95a] S. Jøndrup. Automorphisms and derivations of upper triangular matrix rings. Linear Algebra and its Applications, 221(1–3):205–218, May ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300255.

Jong:1995:EAM

[Jon95b] Juang Jong. Existence of algebraic matrix Riccati equations arising in transport theory. Linear Algebra and its Applications, 230(1–3):89–100, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300366.

Johnson:1993:SPE

[JOTV93] Charles R. Johnson, D. D.

Olesky, Michael Tsatsomeros, and P. Van Den Driessche. Spectra with positive elementary symmetric functions. Linear Algebra and its Applications, 180:247–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1995:SDF

[JOv95] Charles R. Johnson, D. Dale Olesky, and P. van den Driessche. Sign determinancy in LU factorization of P - matrices. Linear Algebra and its Applications, 217(1–3): 155–166, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400061. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Johnson:1999:EBF

[JOvdD99] Charles R. Johnson, D. D.

Olesky, and P. van den Driessche. Elementary bidiagonal factorizations. Linear Algebra and its Applications, 292 (1–3):233–244, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6385.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6385.

Johnson:1996:PMT

[JPR^{+}96] D. S. Johnson, A. C. Pugh, E. Rogers, G. E. Hayton, and D. H. Owens. A polynomial matrix theory for a certain class of twodimensional linear systems. Linear Algebra and its Applications, 241/243(1–3):669– 703, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500430. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Jackson:1999:GCB

[JR99] P. S. Jackson and P. Rowlinson. On graphs with complete bipartite star complements. Linear Algebra and its Applications, 298 (1–3):9–20, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/18/article.
pdf.

Jafarian:1990:CPH

[JS90a] A. A. Jafarian and A. R.

Sourour. Continuity properties of Hyperlat and reducing subspaces. Linear Algebra and its Applications, 141: 253–264, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jia:1990:LII

[JS90b] R. Q. Jia and N. Sivakumar. On the linear independence of integer translates of box splines with rational directions. Linear Algebra and its Applications, 135:19– 31, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1991:EJF

[JS91] Charles R. Johnson and Erik A. Schreiner. Explicit Jordan form for certain block triangular matrices. Linear Algebra and its Applications, 150:297–314, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Jain:1993:NNM

[JS93] S. K. Jain and L. E. Snyder. Nonnegative normal matrices. Linear Algebra and its Applications, 182:147–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1996:CPM

[JS96] Charles R. Johnson and Ronald L. Smith. The completion problem for M-matrices and inverse M-matrices. Linear Algebra and its Applications, 241/243(1–3):655–667, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500429. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Johnson:1998:FLB

[JS98a] Charles R. Johnson and Tomasz Szulc. Further lower bounds for the smallest singular value. Linear Algebra and its Applications, 272(1–3): 169–179, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700330.

Jones:1998:CSS

[JS98b] Cristen Jones and M. Seetharama Gowda. On the connectedness of solution sets in linear complementarity problems. Linear Algebra and its Applications, 272(1–3):33– 44, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700282.

Jammalamadaka:1999:CGL

[JS99a] S. R. Jammalamadaka and D. Sengupta. Changes in the general linear model: A unified approach. Linear Algebra and its Applications, 289(1–3):225–242, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6046.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6046.

Johnson:1999:SIM

[JS99b] Charles R. Johnson and Ronald L. Smith. The symmetric inverse M-matrix completion problem. Linear Algebra and its Applications, 290(1–3):193–212, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        290/1-3/6335.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=290&issue=
                                        1-3&aid=6335.

Johnson:1994:DEU

[JSOvdD94] Charles R. Johnson, David P. Stanford, D. Dale Olesky, and P. van den Driessche. Dominant eigenvalues under tracepreserving diagonal perturbations. Linear Algebra and its Applications, 212/213:415– 435, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Johnson:1995:CBR

[JT95] Charles R. Johnson and Pablo Tarazaga. Connections between the real positive semidefinite and distance matrix completion problems. Linear Algebra and its Applications, 223/224(1–3):375– 391, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500096. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Jacobs:1997:DTN

[JT97] David P. Jacobs and Vilmar Trevisan. The determinant of a tree’s neighborhood matrix. Linear Algebra and its Applications, 256(1–3):235– 249, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600038.

Jordan:1997:JFC

[JTU97] C. Jordán, J. R. Torregrosa, and A. Urbano. On the Jordan form of completions of partial upper triangular matrices. Linear Algebra and its Applications, 254(1–3):241– 250, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600472. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Jordan:1998:ANC

[JTU98] C. Jordán, J. R. Torregrosa, and A. Urbano. An algorithm for nilpotent completions of partial Jordan matrices. Linear Algebra and its Applications, 275/276(1– 3):315–325, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6025.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6025. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Juhasz:1996:SEB

[Juh96] Ferenc Juhász. On the structural eigenvalues of block random matrices. Linear Algebra and its Applications, 246(1–3):225–231, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=246&issue=1-
                                        3&aid=9400356.

Jang:1992:NSL

[JV92] Ruey-Jen Jang and Harold Dean Victory. On nonnegative solvability of linear integral equations. Linear Algebra and its Applications, 165:197–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1999:QMG

[JW99] Norman W. Johnson and Asia Ivić Weiss. Quaternionic modular groups. Linear Algebra and its Applications, 295(1–3):159–189, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6452.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6452.

Wu:1995:EPM

[jWySmJzZ95] Xiao jun Wu, Jia yu Shao, Zhi ming Jiang, and Xi zhao Zhou. On the exponents of primitive, ministrong digraphs with shortest elementary circuit length s. Linear Algebra and its Applications, 222(1–3): 41–61, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1995&volume=222&issue=1-
                                        3&aid=9300274.

Jiang:1991:EBA

[JZ91] Feng Jiang and Rong Fu Zhou. An error bound on the AOR method for solving linear systems. Linear Algebra and its Applications, 158:61–75, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:1996:OIM

[JZ96] Charles R. Johnson and Fuzhen Zhang. An operator inequality and matrix normality. Linear Algebra and its Applications, 240(1–3):105– 110, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=240&issue=1-
                                        3&aid=9400189.

Jiang:1997:LBS

[JZ97] Erxiong Jiang and Xingzhi Zhan. Lower bounds for the spread of a Hermitian matrix. Linear Algebra and its Applications, 256(1–3):153– 163, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=256&issue=1-
                                        3&aid=9500785.

Jia:1999:SPT

[JZ99] Rong-Qing Jia and Shurong Zhang. Spectral properties of the transition operator associated to a multivariate refinement equation. Linear Algebra and its Applications, 292(1–3):155–178, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6375.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6375.

Karuppanchetty:1995:PMH

[KA95] C. S. Karuppanchetty and S. Maria Arulraj. Permanental mates and Hwang’s conjecture. Linear Algebra and its Applications, 216(1–3):225– 237, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300139.

Kalantari:1996:TAM

[Kal96] Bahman Kalantari. A theorem of the alternative for multihomogeneous functions and its relationship to diagonal scaling of matrices. Linear Algebra and its Applications, 236(1–3):1– 24, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400162.

Kamen:1994:BFC

[Kam94] E. W. Kamen. Block-form control of linear time-invariant discrete-time systems defined over commutative rings. Linear Algebra and its Applications, 205/206:805–829, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kantor:1993:DOS

[Kan93] Isaiah L. Kantor. Description of the optimal solution set of the linear programming problem and the dimension formula. Linear Algebra and its Applications, 179:19–32, January 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kantor:1995:QLS

[Kan95] William M. Kantor. Quaternionic line-sets and quaternionic Kerdock codes. Linear Algebra and its Applications, 226/228(1–3):749–779, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500549.

Kaplansky:1995:CNT

[Kap95] Irving Kaplansky. A contribution to von Neumann’s theory of games. II. Linear Algebra and its Applications, 226/228 (1–3):371–373, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500167.

Karcanias:1994:MBM

[Kar94] Nicos Karcanias. Minimal bases of matrix pencils: Algebraic Toeplitz structure and geometric properties. Linear Algebra and its Applications, 205/206:831–868, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karampetakis:1997:CGI

[Kar97] N. P. Karampetakis. Computation of the generalized inverse of a polynomial matrix and applications. Linear Algebra and its Applications, 252(1–3):35–60, February ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=252&issue=1-
                                        3&aid=9500695.

Karow:1999:SAO

[Kar99] Michael Karow. Self-adjoint operators and pairs of Hermitian forms over the quaternions. Linear Algebra and its Applications, 299(1–3): 101–117, September 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/112/21/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/112/21/24/article.
                                        pdf.

Katayama:1994:SFA

[Kat94] Tohru Katayama. A spectral factorization algorithm for discrete-time descriptor systems via generalized eigenproblems. Linear Algebra and its Applications, 205/206:869– 891, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Katirtzoglou:1999:CTV

[Kat99] Eleni Katirtzoglou. The cycle time vector of D–A–D functions. Linear Algebra and its Applications, 293(1– 3):133–144, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6380.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6380.

Kokol-Bukovsek:1999:MSH

[KB99] Damjana Kokol-Bukovšek.

Matrix semigroup homomorphisms from dimension two to three. Linear Algebra and its Applications, 296(1–3):99–112, July 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/22/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/22/article.
pdf.

Kastner:1998:EMA

[KBDZ98] J. Kästner, A. Borobia, A. Dress, and Ch. Zylka. Extremal majorizing and antimajorizing matrices. Linear Algebra and its Applications, 278(1–3):133–145, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6036.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6036.

Kaszkurewicz:1990:CPA

[KBŠ90] E. Kaszkurewicz, A. Bhaya, and D. D. Šiljak. On the convergence of parallel asynchronous block-iterative computations. Linear Algebra and its Applications, 131:139–160, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalantari:1997:LPM

[KEK97] B. Kalantari and M. R. Emamy-K. On linear programming and matrix scaling over the algebraic numbers. Linear Algebra and its Applications, 262(1–3):283–306, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=262&issue=1-
                                        3&aid=9600519.

Kestenband:1992:QHF

[Kes92] Barbu C. Kestenband. Quadrics as hyperplanes in finite affine geometries. Linear Algebra and its Applications, 174:165– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Keska:1999:VHT

[Kes99] Sergiusz Keska. A variant of the Hausdorff theorem for multiindex matrices I. Linear Algebra and its Applications, 295(1–3):201–221, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6448.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6448.

Keliba:1990:NCT

[KH90] Ntumba Tshimvunda Keliba and Dirk Huylebrouck. A note on conjugate Toeplitz matrices. Linear Algebra and its Applications, 139:103–109, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kunicki:1992:NPL

[KH92] Catherine M. Kunicki and Richard D. Hill. Normalpreserving linear transformations. Linear Algebra and its Applications, 170:107–??, June 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kenney:1995:TNB

[KH95a] Charles Kenney and Gary Hewer. Trace norm bounds for stable Lyapunov operators. Linear Algebra and its Applications, 221(1–3):1– 18, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300219.

Kirkland:1995:CRC

[KH95b] Steve Kirkland and Robert E. Hartwig. A convexity result for complex numbers with applications to nonnegative normal matrices. Linear Algebra and its Applications, 221(1–3): 195–204, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300254.

Khachiyan:1996:DMS

[Kha96] Leonid Khachiyan. Diagonal matrix scaling is NP-hard. Linear Algebra and its Applications, 234(1–3):173–179, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400099.

Kimelfeld:1995:GMT

[Kim95] Boris Kimelfeld. A generalization of Muirhead’s theorem. Linear Algebra and its Applications, 216(1–3):205– 209, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300133.

Kirkland:1993:SLM

[Kir93] Steve Kirkland. On the spectrum of a Leslie matrix with a near-periodic fecundity pattern. Linear Algebra and its Applications, 178:261–??, January 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:1995:SRT

[Kir95]  Steve Kirkland.  Spectral

radii of tournament matrices whose graphs are related by an arc reversal. Linear Algebra and its Applications, 217(1–3): 179–202, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400160. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Kirkland:1996:NSB

[Kir96a] Steve Kirkland. A note on the sequence of Brualdi–Li matrices. Linear Algebra and its Applications, 248(1–3):233– 240, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500196.

Kirkland:1996:MPV

[Kir96b] Steve Kirkland. On the minimum Perron value for an irreducible tournament matrix. Linear Algebra and its Applications, 244(1–3):277–304, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400236.

Kirkland:1997:NPV

[Kir97a]  Steve Kirkland.  A note

on Perron vectors for almost regular tournament matrices. Linear Algebra and its Applications, 266(1–3):43– 47, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600628.

Kirkland:1997:NEP

[Kir97b] Steve Kirkland. A note on the eigenvalues of a primitive matrix with large exponent. Linear Algebra and its Applications, 253(1–3):103– 112, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500701.

Kirkland:1997:PVB

[Kir97c] Steve Kirkland. Perron vector bounds for a tournament matrix with applications to a conjecture of Brualdi and Li. Linear Algebra and its Applications, 262(1–3):209–227, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600483.

Kirkland:1998:RNM

[Kir98] S. Kirkland. Review of Nonnegative Matrices and Applications. Linear Algebra and its Applications, 277(1–3): 357–??, June 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:1999:PVO

[Kir99] Steve Kirkland. Perron vector ordering for a subclass of tournament matrices. Linear Algebra and its Applications, 291(1–3):151–166, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        291/1-3/6352.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=291&issue=
                                        1-3&aid=6352.

Kittaneh:1991:ZTM

[Kit91] Fuad Kittaneh. On zero-trace matrices. Linear Algebra and its Applications, 151:119–124, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittaneh:1992:NAG

[Kit92] Fuad Kittaneh. A note on the arithmetic-geometric-mean inequality for matrices. Linear Algebra and its Applications, 171:1–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittappa:1993:RSO

[Kit93] R. Kit Kittappa. A representation of the solution of the n th other linear difference equation with variable coefficients. Linear Algebra and its Applications, 193:211–222, November 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittaneh:1994:SOI

[Kit94] Fuad Kittaneh. On some operator inequalities. Linear Algebra and its Applications, 208/209:19–28, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kiwiel:1995:BIS

[Kiw95a] Krzysztof C. Kiwiel. Blockiterative surrogate projection methods for convex feasibility problems. Linear Algebra and its Applications, 215(1–3):225–259, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300089.

Kiwiel:1995:ISL

[Kiw95b] Krzysztof C. Kiwiel. Iterative schemes for the least 2-Norm solution of piecewise linear programs. Linear Algebra and its Applications, 229(1–3): 1–7, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9300314.

Kiwiel:1997:MGM

[Kiw97] Krzysztof C. Kiwiel. Monotone Gram matrices and deepest surrogate inequalities in accelerated relaxation methods for convex feasibility problems. Linear Algebra and its Applications, 252(1–3):27– 33, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500608.

Knopfmacher:1990:DVP

[KK90] Arnold Knopfmacher and John Knopfmacher. The distribution of values of polynomials over a finite field. Linear Algebra and its Applications, 134:145–151, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalantari:1996:CNM

[KK96] Bahman Kalantari and Leonid Khachiyan. On the complexity of nonnegative-matrix scaling. Linear Algebra and its Applications, 240(1–3):87– 103, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=240&issue=1-
3&aid=9400188.

Kohno:1997:IMG

[KKNU97] Toshiyuki Kohno, Hisashi Kotakemori, Hiroshi Niki, and Masataka Usui. Improving the modified Gauss– Seidel method for Z-matrices. Linear Algebra and its Applications, 267(1–3):113–123, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700063.

Kaltofen:1990:PAM

[KKS90] Erich Kaltofen, M. S. Krishnamoorthy, and B. David Saunders. Parallel algorithms for matrix normal forms. Linear Algebra and its Applications, 136:189–208, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaporin:1991:BSP

[KKY91] I. E. Kaporin, L. Yu.

Kolotilina, and A. Yu. Yeremin. Block SSOR preconditionings for high-order 3D FE systems. II. incomplete BSSOR preconditionings. Linear Algebra and its Applications, 154/156:647– 674, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karcanias:1996:GIM

[KL96] Nicos Karcanias and John Leventides. Grassmann invariants, matrix pencils, and linear system properties. Linear Algebra and its Applications, 241/243(1–3):705–731, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500590. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Kratz:1998:LCO

[KL98a] Werner Kratz and Dirk Liebscher. A local characterization of observability. Linear Algebra and its Applications, 269(1):115–137, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700061.

Krupnik:1998:LRS

[KL98b] Ilya Krupnik and Peter Lancaster. Linearizations, realization, and scalar products for regular matrix polynomials. Linear Algebra and its Applications, 272(1–3):45– 57, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700301.

Kreinovich:1996:ALA

[KLN96] V. Kreinovich, A. V. Lakeyev, and S. I. Noskov. Approximate linear algebra is intractable. Linear Algebra and its Applications, 232(1–3):45– 54, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koltracht:1990:SSM

[KLS90] Israel Koltracht, Peter Lancaster, and Digby Smith. The structure of some matrices arising in tomography. Linear Algebra and its Applications, 130:193–218, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Kunkel:1990:NSD

[KM90] P. Kunkel and V. Mehrmann.

Numerical solution of differential algebraic Riccati equations. Linear Algebra and its Applications, 137/138:39– 66, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kojima:1991:RBP

[KM91] Masakazu Kojima and Nimrod Megiddo. The relation between the path of centers and Smale’s regularization of the linear programming problem. Linear Algebra and its Applications, 152:135–139, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Kunkel:1994:NLP

[KM94] Peter Kunkel and Volker Mehrmann. A new look at pencils of matrix valued functions. Linear Algebra and its Applications, 212/213:215– 248, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Khosrovshahi:1995:BT

[KM95] G. B. Khosrovshahi and Ch. Maysoori. On the bases for trades. Linear Algebra and its Applications, 226/228 (1–3):731–748, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500456.

Koukouvinos:1996:SNF

[KMS96] C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. On the Smith normal form of D-optimal designs. Linear Algebra and its Applications, 247(1–3):277–295, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500117.

Kim:1991:DPA

[KN91] K. Kim and J. L. Nazareth.

The decomposition principle and algorithms for linear programming. Linear Algebra and its Applications, 152:119– 133, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Kirkland:1995:GIM

[KN95] Stephen J. Kirkland and Michael Neumann. Group inverses of M-matrices associated with nonnegative matrices having few eigenvalues. Linear Algebra and its Applications, 220(1–3):181– 213, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400301. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Kollo:1997:DOM

[KN97] Tônu Kollo and Heinz Neudecker. The derivative of an orthogonal matrix of eigenvectors of a symmetric matrix. Linear Algebra and its Applications, 264(1–3):489–493, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700016.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700016.

Kamm:1998:KPS

[KN98] Julie Kamm and James G. Nagy. Kronecker product and SVD approximations in image restoration. Linear Algebra and its Applications, 284 (1–3):177–192, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6145.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

year=1998&volume=284&issue=
1-3&aid=6145.

Kirkland:1999:GIM

[KN99] Stephen J. Kirkland and Michael Neumann. On group inverses of M-matrices with uniform diagonal entries. Linear Algebra and its Applications, 296(1–3):153–170, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/25/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/25/article.
pdf.

Knight:1995:FRM

[Kni95]  Philip A. Knight.  Fast

rectangular matrix multiplication and QR decomposition. Linear Algebra and its Applications, 221(1–3):69– 81, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300230.

Knizhnerman:1999:EBA

[Kni99] Leonid Knizhnerman. Error bounds for the Arnoldi method: a set of extreme eigenpairs. Linear Algebra and its Applications, 296(1–3):191–211, July 15, 1999. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/47/17/28/
                                        abstract.html;   http://
                                        www.elsevier.nl/gej-ng/
                                        10/30/19/47/17/28/article.
                                        pdf.

Knowles:1992:MPA

[Kno92] James K. Knowles. On a minimization problem associated with linear dynamical systems. Linear Algebra and its Applications, 165:173–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:1998:API

[KNS98] Stephen J. Kirkland, Michael Neumann, and Bryan L. Shader. Applications of Paz’s inequality to perturbation bounds for Markov chains. Linear Algebra and its Applications, 268(1–3):183– 196, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=268&issue=1-
                                        3&aid=9700042.

Knuppel:1996:PSM

[Knü96] Frieder Knüppel. Products of similar matrices. Linear Algebra and its Applications, 233(1–3):99–108, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400057.

Koning:1991:BKP

[KNW91] Ruud H. Koning, Heinz Neudecker, and Tom Wansbeek. Block Kronecker products and the vecb operator. Linear Algebra and its Applications, 149:165–184, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kailath:1997:DPP

[KO97a] Thomas Kailath and Vadim Olshevsky. Diagonal pivoting for partially reconstructible Cauchy-like matrices, with applications to Toeplitz-like linear equations and to boundary rational matrix interpolation problems. Linear Algebra and its Applications, 254(1–3): 251–302, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600288. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Kailath:1997:DSA

[KO97b] Thomas Kailath and Vadim Olshevsky. Displacementstructure approach to polynomial Vandermonde and related matrices. Linear Algebra and its Applications, 261(1–3): 49–90, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=261&issue=1-
                                        3&aid=9600315.

Kohnen:1997:G

[Koh97] Winfried Kohnen. On the generators of Sp_{n}(Z). Linear Algebra and its Applications, 253(1–3):363–367, March 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=253&issue=1-
                                        3&aid=9500708.

Kolotilina:1993:LBP

[Kol93] L. Yu. Kolotilina. Lower bounds for the Perron root of a nonnegative matrix. Linear Algebra and its Applications, 180:133–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:1995:TSB

[Kol95] L. Yu. Kolotilina. Two-sided bounds for the inverse of an H-matrix. Linear Algebra and its Applications, 225(1):117– 123, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300325.

Kolotilina:1998:EBI

[Kol98] L. Yu. Kolotilina. Eigenvalue bounds and inequalities using vector aggregation of matrices. Linear Algebra and its Applications, 271(1–3): 139–167, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700260.

Koliha:1999:SPP

[Kol99] J. J. Koliha. A simple proof of the product theorem for EP matrices. Linear Algebra and its Applications, 294(1–3):213–215, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6413.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=294&issue=
                                        1-3&aid=6413.

Komaroff:1990:RMP

[Kom90] N. Komaroff. Rearrangement and matrix product inequalities. Linear Algebra and its Applications, 140:155–161, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kopotun:1996:NCS

[Kop96a] Kirill A. Kopotun. A note on the convexity of the sum of subpermanents. Linear Algebra and its Applications, 245(1–3):157–169, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=245&issue=1-
                                        3&aid=9400228.

Koppinen:1996:TAT

[Kop96b] M. Koppinen. Three automorphism theorems for triangular matrix algebras. Linear Algebra and its Applications, 245(1–3):295–304, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1996&volume=245&issue=1-
3&aid=9400309.

Kosir:1993:SCM

[Koš93] Tomaž Košir. On the structure of commutative matrices. Linear Algebra and its Applications, 187:163–182, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kosir:1994:FDM

[Koš94] Tomaž Košir. Finite-dimensional multiparameter spectral theory: the nonderogatory case. Linear Algebra and its Applications, 212/213:45–70, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Kosir:1996:KBL

[Koš96]  Tomaž Košir.  Kronecker

bases for linear matrix equations, with application to twoparameter eigenvalue problems. Linear Algebra and its Applications, 249(1–3):259– 288, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500361.

Kosir:1997:SCM

[Koš97] Tomaž Košir. On the structure of commutative matrices.

II. Linear Algebra and its Applications, 261(1–3):293– 305, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600413.

Kovacec:1994:CMO

[Kov94] Alexander Kovačec. On a conjecture of Marcus and de Oliveria. Linear Algebra and its Applications, 201:91–97, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kovacec:1999:MOC

[Kov99]  Alexander Kovačec.  The

Marcus-de Oliveira conjecture, bilinear forms, and cones. Linear Algebra and its Applications, 289(1–3): 243–259, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6004.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6004.

Kirkland:1992:BST

[KP92a] Stephen Kirkland and Norman J. Pullman. Boolean spectral theory. Linear Algebra and its Applications, 175: 177–??, October 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:1992:IBF

[KP92b] Lily Kolotilina and Ben Polman. On incomplete block factorization methods of generalized SSOR type for Hmatrices. Linear Algebra and its Applications, 177:111– ??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kishore:1994:KRP

[KP94] A. P. Kishore and J. B. Pearson, Jr. Kernel representation and properties of discrete-time input-output systems. Linear Algebra and its Applications, 205/206:893–908, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:1995:EIB

[KP95] L. Kolotilina and B. Polman. Erratum: “On incomplete block factorization methods of generalized SSOR type for H-matrices” [Linear Algebra Appl. 177 (1992), 111– 136, MR 93k:65028]. Linear Algebra and its Applications, 220(1–3):6, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Konstantinov:1997:PAO

[KPGP97] M. M. Konstantinov, P. Hr. Petkov, D.-W. Gu, and I. Postlethwaite. Perturbation analysis of orthogonal canonical forms. Linear Algebra and its Applications, 251(1–3):267–291, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500647.

Kandic:1997:ECH

[KPRV97] Dragan B. Kandić, Beresford Parlett, Branimir D. Reljin, and Petar M. Vasić. Explicit construction of hyperdominant symmetric matrices with assigned spectrum. Linear Algebra and its Applications, 258(1–3):41– 51, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600161.

Kontovasilis:1991:BCS

[KPS91] Kimon Kontovasilis, Robert J. Plemmons, and William J. Stewart. Block cyclic SOR for Markov chains with p-cyclic infinitesimal generator. Linear Algebra and its Applications, 154/156:145–223, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kortanek:1991:SEI

[KPY91] K. O. Kortanek, Florian Potra, and Yinyu Ye. On some efficient interior point methods for nonlinear convex programming. Linear Algebra and its Applications, 152:169– 189, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Kouvaritakis:1990:BPP

[KR90] B. Kouvaritakis and J. A.

Rossiter. Branch-point placement. Linear Algebra and its Applications, 140:217–249, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:1991:PCM

[KR91] K. H. Kim and F. W. Roush.

Path components of matrices and strong shift equivalence over Q^{+}. Linear Algebra and its Applications, 145:177–186, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:1992:ALM

[KR92] K. H. Kim and F. W. Roush.

Automorphisms of lambdamatrices. Linear Algebra and its Applications, 174:75– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krupnik:1994:CPJ

[KR94] Mark Krupnik and Leiba Rodman. Completions of partial Jordan and Hessenberg matrices. Linear Algebra and its Applications, 212/213:267– 287, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Krupnick:1996:ECP

[KR96a] M. Krupnick and L. Rodman. Erratum: “Completions of partial Jordan and Hessenberg matrices” [Linear Algebra Appl. 212/213 (1994), 267–287, MR 95k:15018]. Linear Algebra and its Applications, 240(1–3):235–236, June ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krupnik:1996:ISB

[KR96b] Naum Krupnik and Steffen Roch. On invertibility symbols in Banach algebras. Linear Algebra and its Applications, 245(1–3):77–81, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400214.

Kraaijevanger:1991:CLD

[Kra91] J. F. B. M. Kraaijevanger.

A characterization of Lyapunov diagonal stability using Hadamard products. Linear Algebra and its Applications, 151:245–254, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kratz:1993:LTM

[Kra93] Werner Kratz. A limit theorem for monotone matrix functions. Linear Algebra and its Applications, 194:205–222, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krause:1994:BVM

[Kra94] Gerd M. Krause. Bounds for the variation of matrix eigenvalues and polynomial roots. Linear Algebra and its Applications, 208/209:73–82, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kratz:1995:CSO

[Kra95] Werner Kratz. Characterization of strong observability and construction of an observer. Linear Algebra and its Applications, 221(1–3):31– 40, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300221.

Krafft:1998:AHM

[Kra98] Olaf Krafft. An arithmeticharmonic-mean inequality for nonnegative definite matrices. Linear Algebra and its Applications, 268(1–3):243– 246, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700046.

Krieg:1994:CMM

[Kri94] Aloys Krieg. Counting modular matrices with specified maximum norm. Linear Algebra and its Applications, 196: 273–278, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krieger:1995:TLM

[Kri95]  Udo R. Krieger.  On a

two-level multigrid solution method for finite Markov chains. Linear Algebra and its Applications, 223/224(1–3): 415–438, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500166. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Krijnen:1999:NPS

[Kri99] Wim P. Krijnen. A note on the parameter set for factor analysis models. Linear Algebra and its Applications, 289(1–3):261–266, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6044.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6044.

Keeler:1997:NRM

[KRS97] Dennis S. Keeler, Leiba Rodman, and Ilya M. Spitkovsky. The numerical range of 3 × 3 matrices. Linear Algebra and its Applications, 252(1–3):115–139, February ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500674.

Krupnik:1992:CSO

[Kru92] Mark Krupnik. Changing the spectrum of an operator by perturbation. Linear Algebra and its Applications, 167: 113–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krupnik:1995:GMC

[Kru95] Mark Krupnik. Geometric multiplicities of completions of partial triangular matrices. Linear Algebra and its Applications, 220(1–3):215– 227, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=220&issue=1-
                                        3&aid=9400042. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Krupnik:1997:JSU

[Kru97] Mark Krupnik. Jordan structures of upper equivalent matrices. Linear Algebra and its Applications, 261(1–3):167– 172, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1997&volume=261&issue=1-
3&aid=9600367.

Krafft:1990:CPC

[KS90] Olaf Krafft and Martin Schaefer. Convergence of the powers of a circulant stochastic matrix. Linear Algebra and its Applications, 127:59–69, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knyazev:1991:EEC

[KS91] A. V. Knyazev and A. L.

Skorokhodov. On exact estimates of the convergence rate of the steepest ascent method in the symmetric eigenvalue problem. Linear Algebra and its Applications, 154/156:245– 257, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuijper:1992:RPF

[KS92] M. Kuijper and J. M. Schumacher. Realization and partial fractions. Linear Algebra and its Applications, 169:195– ??, May 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:1994:TME

[KS94a] Stephen J. Kirkland and Bryan L. Shader. Tournament matrices with extremal spectral properties. Linear Algebra and its Applications, 196: 1–17, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knott:1994:GCM

[KS94b] M. Knott and C. S. Smith. On a generalization of cyclic monotonicity and distances among random vectors. Linear Algebra and its Applications, 199:363–371, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kovacec:1994:CTC

[KS94c] Alexander Kovačec and Ana Paula Santana. A combinatorial theorem on the Cartan invariants of the Schur algebra S(B^{+}). Linear Algebra and its Applications, 197/198:79–91, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Klein:1996:FIM

[KS96] André Klein and Peter Spreij. On Fisher’s information matrix of an ARMAX process and Sylvester’s resultant matrices. Linear Algebra and its Applications, 237/238(1–3): 579–590, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500552. Special

issue honoring Calyampudi Radhakrishna Rao.

Kolodziejczak:1999:CCM

[KS99] Barbara Kolodziejczak and Tomasz Szulc. Convex combinations of matrices — full rank characterization. Linear Algebra and its Applications, 287(1–3):215–222, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6253.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6253.

Koplon:1994:OLS

[KSH94] Renée Koplon, Eduardo D.

Sontag, and M. L. J. Hautus. Observability of linear systems with saturated outputs. Linear Algebra and its Applications, 205/206:909–936, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Konstantinov:1998:PBC

[KSP98] M. M. Konstantinov, M. O.

Stanislavova, and P. H. Petkov. Perturbation bounds and characterisation of the solution of the associated algebraic Riccati equation. Linear Algebra and its Applications, 285(1–3):7–31, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

                                        tree/store/laa/sub/1998/
                                        285/1-3/6213.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=285&issue=
                                        1-3&aid=6213.

Kulkarni:1999:ETP

[KST99] Devadatta Kulkarni, Darrell Schmidt, and Sze-Kai Tsui. Eigenvalues of tridiagonal pseudo-Toeplitz matrices. Linear Algebra and its Applications, 297(1–3):63–80, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/21/article.
                                        pdf.

Kovac-Striko:1991:SRS

[KSV91] J. Kovač-Striko and K. Veselić. Some remarks on the spectra of Hermitian matrices. Linear Algebra and its Applications, 145:221–229, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kovac-Striko:1995:TMD

[KSV95] J. Kovač-Striko and K. Veselić. Trace minimization and definiteness of symmetric pencils. Linear Algebra and its Applications, 216(1–3):139– 158, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300126.

Karlovich:1999:APF

[KSW99] Yuri I. Karlovich, Ilya M.

Spitkovsky, and Ronald A. Walker. Almost periodic factorization of block triangular matrix functions revisited. Linear Algebra and its Applications, 293(1– 3):199–232, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6389.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6389.

Klafszky:1991:RPP

[KT91] Emil Klafszky and Tamás Terlaky. The role of pivoting in proving some fundamental theorems of linear algebra. Linear Algebra and its Applications, 151:97–118, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kantor:1993:FPF

[KT93] Isaiah L. Kantor and Ivan M.

Trishin. Formulas for powers and functions of matrices.

Linear Algebra and its Applications, 186:1–13, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kautsky:1995:DBW

[KT95a] Jaroslav Kautsky and Radka Turcajová. Discrete biorthogonal wavelet transforms as block circulant matrices. Linear Algebra and its Applications, 223/224(1–3):393– 413, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400298. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Kautsky:1995:PPF

[KT95b] Jaroslav Kautsky and Radka Turcajová. Pollen product factorization and construction of higher multiplicity wavelets. Linear Algebra and its Applications, 222(1–3): 241–260, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=222&issue=1-
                                        3&aid=9300298.

Klatte:1996:NLC

[KT96a] Diethard Klatte and Gisbert Thiere. A note of Lipschitz constants for solutions of linear inequalities and equations. Linear Algebra and its Applications, 244(1–3):365–374, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400349.

Kouikoglou:1996:EBS

[KT96b] Vassilis S. Kouikoglou and Nikos C. Tsourveloudis. Eigenvalue bounds on the solutions of coupled Lyapunov and Riccati equations. Linear Algebra and its Applications, 235(1–3): 247–259, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400145.

Kawamoto:1999:GAR

[KTK99] A. Kawamoto, K. Takaba, and T. Katayama. On the generalized algebraic Riccati equation for continuous-time descriptor systems. Linear Algebra and its Applications, 296(1–3):1–14, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/47/17/17/
                                        abstract.html;   http://
                                        www.elsevier.nl/gej-ng/
                                        10/30/19/47/17/17/article.
                                        pdf.

Kalinina:1993:DNR

[KU93] Elizabeth A. Kalinina and Alexei Yu. Uteshev. Determination of the number of roots of a polynomial lying in a given algebraic domain. Linear Algebra and its Applications, 185:61–81, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kucharz:1991:SDA

[Kuc91] Wojciech Kucharz. Simultaneous diagonalization of an algebraic family of matrices. Linear Algebra and its Applications, 157:49–53, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuroda:1997:QFA

[Kur97] Mitsuru Kuroda. A quadratic form tr(AX^{2}) and its application. Linear Algebra and its Applications, 266(1–3):181– 198, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=266&issue=1-
                                        3&aid=9700002.

Kuroda:1999:RP

[Kur99]  M. Kuroda.  On Rosen-

feld’s problem. Linear Algebra and its Applications, 295(1–3):241–247, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6466.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6466.

Kuznetsov:1997:PBK

[Kuz97] S. V. Kuznetsov. Perturbation bounds of the Krylov bases and associated Hessenberg forms. Linear Algebra and its Applications, 265(1–3):1–28, November 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600299.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600299.

Karl:1990:CST

[KV90] W. C. Karl and G. C. Verghese. Curvatures of surfaces and their shadows. Linear Algebra and its Applications, 130:231–255, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Kravanja:1998:FHS

[KV98] Peter Kravanja and Marc Van Barel. A fast Hankel solver based on an inversion formula for Loewner matrices. Linear Algebra and its Applications, 282(1–3):275–295, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        282/1-3/6198.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=282&issue=
                                        1-3&aid=6198.

Klee:1993:RS

[KvL93] Victor Klee, Balder von Hohenbalken, and Ted Lewis. On the recognition of S-systems. Linear Algebra and its Applications, 192:187–204, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Kos:1995:HCB

[KW95] Johan Kos and Hugo J. Woerdeman. Hermitian completions of band matrices and applications. Linear Algebra and its Applications, 215(1–3):21– 62, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300071.

Kwak:1997:NEA

[Kwa97] Do Y. Kwak. A norm estimate for the ADI method for nonsymmetric problems. Linear Algebra and its Applications, 266(1–3):127–141, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600638.

Kaliski:1993:EPR

[KY93] John A. Kaliski and Yinyu Ye. An extension of the potential reduction algorithm for linear complementarity problems with some priority goals. Linear Algebra and its Applications, 193:35–50, November 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kye:1995:PLM

[Kye95] Seung-Hyeok Kye. Positive linear maps between matrix algebras which fix diagonals. Linear Algebra and its Applications, 216(1–3):239–256, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=216&issue=1-
                                        3&aid=9300140.

Kastner:1993:GUT

[KZ93] J. Kastner and Ch Zylka. On generalized uniformly tapered matrices. Linear Algebra and its Applications, 181:221–??, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:1994:OSC

[– DL94] Dragomir Ž. – Doković and Chi-Kwong Li. Overgroups of some classical linear groups with applications to linear preserver problems. Linear Algebra and its Applications, 197/198:31–61, January 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Dokovic:1995:CHM

[– DL95] Dragomir Ž. – Doković and Leonardo Legorreta. Congruence of Hermitian matrices over some localizations of C[z, z^{−1}]. Linear Algebra and its Applications, 219(1– 3):35–48, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300189.

Latham:1994:AQE

[LA94] Geoff A. Latham and Robert S.

Anderssen. Assessing quantification for the EMS algorithm. Linear Algebra and its Applications, 210:89–122, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:1968:LAA

[LAA68] Linear algebra and its applications, 1968. ISSN 0024-3795 (print), 1873-1856 (electronic). North-Holland Publishing Co., Amsterdam, The Netherlands.

Labahn:1992:ICB

[Lab92] George Labahn. Inversion components of block Hankellike matrices. Linear Algebra and its Applications, 177: 7–??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lacruz:1995:NSM

[Lac95] Miguel Lacruz. Norms of Schur multipliers. Linear Algebra and its Applications, 219(1–3):157–163, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300205.

Laffey:1998:ENM

[Laf98a] Thomas J. Laffey. Extreme nonnegative matrices. Linear Algebra and its Applications, 275/276(1–3):349–357, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6052.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6052. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Laffey:1998:SAO

[Laf98b] Thomas J. Laffey. Some aspects of Olga Taussky’s work in algebra. Linear Algebra and its Applications, 280(1): 51–57, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Lagarias:1992:LBD

[Lag92] J. C. Lagarias. A Lie bracket decomposition and its application to flows on symmetric matrices. Linear Algebra and its Applications, 175:239–??, October 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lakatos:1999:CPW

[Lak99] Piroska Lakatos. On the Coxeter polynomials of wild stars. Linear Algebra and its Applications, 293(1–3): 159–170, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6381.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6381.

Lal:1998:IPI

[Lal98] A. K. Lal. Inequalities for the q-permanent. II. Linear Algebra and its Applications, 274(1–3):1–16, April 15, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700083.

Lambrou:1990:ROR

[Lam90] M. S. Lambrou. On the rank of operators in reflexive algebras. Linear Algebra and its Applications, 142:211–235, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lamond:1993:EBU

[Lam93] Bernard F. Lamond. An efficient basis update of asymptotic linear programming. Linear Algebra and its Applications, 184:83–??, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lanski:1992:IMR

[Lan92] Charles Lanski. An identity for matrix rings with involution. Linear Algebra and its Applications, 174:91– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lange:1996:SLS

[Lan96] Ehler Lange. Substitutions for linear shift register sequences and the factorization algorithms of Berlekamp and Niederreiter. Linear Algebra and its Applications, 249(1–3):217–228, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500358.

Laurent:1997:RPS

[Lau97] Monique Laurent. The real positive semidefinite completion problem for series-parallel graphs. Linear Algebra and its Applications, 252(1–3):347– 366, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500741.

Laurent:1998:CBP

[Lau98] Monique Laurent. A connection between positive semidefinite and Euclidean distance matrix completion problems. Linear Algebra and its Applications, 273(1–3):9– 22, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9783714.

Lavric:1997:MPC

[Lav97] Boris Lavrič. Monotonicity properties of certain classes of norms. Linear Algebra and its Applications, 259(1–3): 237–250, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600287.

Lavric:1998:CLO

[Lav98a] Boris Lavrič. Compatible lattice orders and linear operators on R^{n}. Linear Algebra and its Applications, 285 (1–3):189–200, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        285/1-3/6234.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=285&issue=
                                        1-3&aid=6234.

Lavric:1998:CCB

[Lav98b] Boris Lavrič. Convergence of certain bounded sequences. Linear Algebra and its Applications, 278(1–3):1–10, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        278/1-3/6073.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=278&issue=
                                        1-3&aid=6073.

Lavric:1999:NOM

[Lav99a] Boris Lavrič. A note on ∗ orthant-monotonic norms. Linear Algebra and its Applications, 299(1–3):195–200, September 15, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/112/21/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/112/21/31/article.
pdf.

Lavric:1999:SN

[Lav99b] Boris Lavrič. S-subinvariant norms. Linear Algebra and its Applications, 288(1–3): 199–210, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6312.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6312.

Lazarus:1995:ECM

[Laz95] Andrew J. Lazarus. Eigenvectors of circulant matrices of prime dimension. Linear Algebra and its Applications, 221(1–3):111–116, May ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300234.

Lundquist:1996:RIP

[LB96] Michael Lundquist and Wayne Barrett. Rank inequalities for positive semidefinite matrices. Linear Algebra and its Applications, 248(1–3):91– 100, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=248&issue=1-
                                        3&aid=9500170.

Labahn:1995:IMH

[LBC95] George Labahn, Bernhard Beckermann, and Stan Cabay. Inversion of mosaic Hankel matrices via matrix polynomial systems. Linear Algebra and its Applications, 221(1–3): 253–279, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300262.

Li:1991:BHP

[LC91a] Ching Li and Ji Cheng Chen. On a bound for the Hadamard product of an M-matrix and its inverse. Linear Algebra and its Applications, 144:171– 178, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:1991:AMC

[LC91b] Wen Wei Lin and C. W.

Chen. An acceleration method for computing the generalized eigenvalue problem on a parallel computer. Linear Algebra and its Applications, 146: 49–65, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ling:1992:ISV

[LC92] Chen Ling and Wong Chi Song. Inequalities for singular values and traces. Linear Algebra and its Applications, 171:109–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lakshminarayanan:1996:TTC

[LC96] Sambhavi Lakshminarayanan and R. Chandrasekaran. Tessellation and g-tessellation of circulants, Q_{6}, and Q_{6}^{t}. Linear Algebra and its Applications, 245(1–3):191–222, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400239.

LeFerrand:1994:QID

[Le 94] Hervé Le Ferrand. Quelques inégalités sur les déterminants de Gram d’une suite vectorielle de Krylov associée à une matrice orthogonale de R^{p}. (French) [some inequalities on the Gram determinants of a Krylov vector sequence associated with an orthogonal matrix in R^{p}]. Linear Algebra and its Applications, 196:243– 252, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

LeBruyn:1998:GGC

[Le 98] Lieven Le Bruyn. Generating graded central simple algebras. Linear Algebra and its Applications, 268(1–3):323– 344, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700050.

Leizarowitz:1992:IPS

[Lei92] Arie Leizarowitz. On infinite products of stochastic matrices. Linear Algebra and its Applications, 168:189–??, April 15, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lei:1997:GSF

[Lei97] Tian-Gang Lei. Generalized Schur functions and generalized decomposable symmetric tensors. Linear Algebra and its Applications, 263(1–3):311– 332, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600542.

Lei:1999:NCS

[Lei99a] Tian-Gang Lei. Notes on Cartesian symmetry classes and generalized trace functions. Linear Algebra and its Applications, 292(1–3): 281–288, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6395.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6395.

Lei:1999:PSD

[Lei99b]  Tian-Gang Lei.  Positive

semi-definiteness in a group algebra. Linear Algebra and its Applications, 292(1– 3):289–296, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6396.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1999&volume=292&issue=
                                        1-3&aid=6396.

Lenard:1994:STR

[Len94] Andrew Lenard. Structure theory of reciprocal pairs of linear transformations. Linear Algebra and its Applications, 207:21–35, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leon:1994:MBF

[Leo94] Steven J. Leon. Maximizing bilinear forms subject to linear constraints. Linear Algebra and its Applications, 210: 49–58, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Levy:1991:DSC

[Lev91] Lawrence S. Levy. Directsum cancellation of submodule systems. Linear Algebra and its Applications, 157:37– 48, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lev:1993:PCC

[Lev93]  Arieh Lev.  Products of

cyclic conjugacy classes in the groups PSL(n, F). Linear Algebra and its Applications, 179:59–83, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lev:1994:PCS

[Lev94] Arieh Lev. Products of cyclic similarity classes in the groups GL_{n}(F). Linear Algebra and its Applications, 202:235–266, April 15, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lev:1997:PCC

[Lev97] Arieh Lev. Powers of 1cyclic conjugacy classes in the groups gl_{n}(F) and sl_{n}(F). Linear Algebra and its Applications, 257(1–3):243–281, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600206.

Levy:1998:NHS

[Lev98] Bernard C. Levy. A note on the hyperbolic singular value decomposition. Linear Algebra and its Applications, 277(1–3):135–142, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6066.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6066.

Lewis:1998:ECF

[Lew98] A. S. Lewis. Eigenvalueconstrained faces. Linear Algebra and its Applications, 269(1):159–181, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9781515.

Lewkowicz:1999:CIC

[Lew99] Izchak Lewkowicz. Convex invertible cones of matrices — a unified framework for the equations of Sylvester, Lyapunov and Riccati. Linear Algebra and its Applications, 286(1–3):107–133, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        286/1-3/6275.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=286&issue=
                                        1-3&aid=6275.

Li:1993:CCM

[LF93] Ping Li and Yu Yu Feng. Criteria for copositive matrices of order four. Linear Algebra and its Applications, 194:109– 124, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loeliger:1994:MOG

[LFMT94] Hans-Andrea Loeliger, G. David Forney, Jr., Thomas Mittelholzer, and Mitchell D. Trott. Minimality and observability of group systems. Linear Algebra and its Applications, 205/206:937–963, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lam:1997:SPB

[LH97] James Lam and Y. S. Hung.

On the stability of projections of balanced realizations. Linear Algebra and its Applications, 257(1–3):163– 182, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600148.

Li:1994:PBS

[LHE94] Zhongshan Li, Frank Hall, and Carolyn Eschenbach. On the period and base of a sign pattern matrix. Linear Algebra and its Applications, 212/213: 101–120, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Li:1997:SPN

[LHZ97] Zhongshan Li, Frank Hall, and Fuzhen Zhang. Sign patterns of nonnegative normal matrices. Linear Algebra and its Applications, 254(1–3):335– 354, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9600469. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Li:1990:NRE

[Li90a] Bing Li. New results on the exponent set of primitive nearly reducible matrices. Linear Algebra and its Applications, 139:1–19, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1990:VSM

[Li90b] Ren Cang Li. On the variation of the spectra of matrix pencils. Linear Algebra and its Applications, 139:147– 164, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1991:SRG

[Li91a] Chi-Kwong Li. Some results on the generalized k-numerical range. Linear Algebra and its Applications, 146:21–29, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1991:SSC

[Li91b] Luo Luo Li. Some sufficient conditions for the solvability of inverse eigenvalue problems. Linear Algebra and its Applications, 148:225–236, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1991:RRS

[Li91c] Wen Li. The relationship of a regular splitting to a graph compatible splitting. Linear Algebra and its Applications, 144:101–105, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1992:ERQ

[Li92a] Ren-Cang Li. On eigenvalues of a Rayleigh quotient matrix. Linear Algebra and its Applications, 169:249–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1992:RSG

[Li92b] Wen Li. On regular splittings and graph compatible splittings of an M-matrix. Linear Algebra and its Applications, 177:163–172, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1993:PBD

[Li93a] Ren Cang Li. A perturbation bound for definite pencils. Linear Algebra and its Applications, 179:191–202, January 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1993:SLC

[Li93b] Wu Li. The sharp Lipschitz constants for feasible and optimal solutions of a perturbed linear program. Linear Algebra and its Applications, 187: 15–40, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1994:LOP

[Li94a] Chi-Kwong Li. Linear operators preserving the (p, q) numerical radius. Linear Algebra and its Applications, 201: 21–42, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1994:SAT

[Li94b] Chi-Kwong Li. Some aspects of the theory of norms. Linear Algebra and its Applications, 212/213:71–100, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Li:1994:EVR

[Li94c] Ren Cang Li. On eigenvalue variations of Rayleigh quotient matrix pencils of a definite pencil. Linear Algebra and its Applications, 208/209: 471–483, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1995:EOG

[Li95a] Jiong-Sheng Li. Eigenvalues of oriented-graph matrices. Linear Algebra and its Applications, 221(1–3):103– 110, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300233.

Li:1995:SCS

[Li95b] Luoluo Li. Sufficient conditions for the solvability of an algebraic inverse eigenvalue problem. Linear Algebra and its Applications, 221(1–3): 117–129, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300225.

Li:1995:CRC

[Li95c] Xie Zhang Li. The convergence rate of the Chebyshev sim under a perturbation of a complex line-segment spectrum. Linear Algebra and its Applications, 230(1–3):47– 60, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300361.

Li:1996:ESA

[Li96] C. K. Li. Erratum: “Some aspects of the theory of norms” [Linear Algebra Appl. 212/213 (1994), 71–100, MR 95h:15043]. Linear Algebra and its Applications, 240(1–3): 237–238, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1997:SIE

[Li97a] Luoluo Li. On solvability of inverse eigenvalue problems with Hermitian matrices. Linear Algebra and its Applications, 253(1–3):89– 101, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500783.

Li:1997:MIE

[Li97b]  Nian Li.  A matrix in-

verse eigenvalue problem and its application. Linear Algebra and its Applications, 266(1–3):143–152, November 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600639.

Li:1997:RPT

[Li97c] Ren-Cang Li. Relative perturbation theory. III. more bounds on eigenvalue variation. Linear Algebra and its Applications, 266(1–3):337– 345, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9700029.

Li:1997:UEB

[Li97d] Xiezhang Li. A uniform error bound for the overrelaxation methods. Linear Algebra and its Applications, 254(1–3): 315–333, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9600313. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Li:1998:ERM

[Li98a] Chi-Kwong Li. Essay review of Matrix analysis [Springer, New York, 1997; 1 477 662] by R. Bhatia. Linear Algebra and its Applications, 273(1–3): 227–229, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=273&issue=1-
                                        3&aid=9700376.

Li:1998:IKF

[Li98b] Luoluo Li. An improvement on Ky Fan’s theorem of matrix eigenvalues. Linear Algebra and its Applications, 279(1–3):111–117, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        279/1-3/6108.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6108.

Li:1998:UGE

[Li98c] Luoluo Li. The undirected graph and estimates of matrix singular values. Linear Algebra and its Applications, 285(1–3):181–188, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6245.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6245.

Li:1998:SVH

[Li98d] Ren-Cang Li. Spectral variations and Hadamard products: Some problems. Linear Algebra and its Applications, 278(1–3):317–326, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6088.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6088.

Li:1998:NM

[Li98e] Wen Li. On Nekrasov matrices. Linear Algebra and its Applications, 281(1–3): 87–96, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6152.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6152.

Lieb:1990:SRP

[Lie90] Elliott H. Lieb. On the spectral radius of the product of matrix exponentials. Linear Algebra and its Applications, 141:271–273, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:1991:LMS

[Lim91a] M. H. Lim. Linear maps on symmetric spaces: the invariance of nonzero decomposable elements. Linear Algebra and its Applications, 148:89– 92, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Limaye:1991:SCO

[Lim91b] Balmohan V. Limaye. A spectral characterization of operators having rank k. Linear Algebra and its Applications, 143:57–66, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:1993:NSPa

[Lim93a] Chjan C. Lim. Nonsingular sign patterns and the orthogonal group. Linear Algebra and its Applications, 184:1–??, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:1993:NSPb

[Lim93b] M. H. Lim. A note on similarity preserving linear maps on matrices. Linear Algebra and its Applications, 190:229– 233, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:1995:UTS

[Lim95] M. H. Lim. A uniqueness theorem on symmetric tensors. Linear Algebra and its Applications, 231(1–3):153–157, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400021.

Lim:1997:IPC

[Lim97] M. H. Lim. Invertible preservers of certain decomposable tensors. Linear Algebra and its Applications, 266(1–3):153–159, November 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600642.

Linzer:1992:SSM

[Lin92] Elliot Linzer. On the stability of solution methods for band Toeplitz systems. Linear Algebra and its Applications, 170:1–??, June 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Linton:1993:VE

[Lin93] S. A. Linton. On vector enumeration. Linear Algebra and its Applications, 192:235–248, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Lin:1998:OVI

[Lin98a] C.-S. Lin. Operator versions of inequalities and equalities on a Hilbert space. Linear Algebra and its Applications, 268(1–3):365–374, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700054.

Lin:1998:NIP

[Lin98b] Zhang Yu Lin. On the number of invariant polynomials of the product of matrices with prescribed similarity classes. Linear Algebra and its Applications, 277(1–3):253–269, June 15, 1998. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6070.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6070.

Linden:1998:BZP

[Lin98c] Hansjörg Linden. Bounds for the zeros of polynomials from eigenvalues and singular values of some companion matrices. Linear Algebra and its Applications, 271(1–3): 41–82, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700254.

Lin:1999:SMP

[Lin99]  Zhiping Lin.  On syzygy

modules for polynomial matrices. Linear Algebra and its Applications, 298(1–3): 73–86, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/22/article.
pdf.

Liski:1992:RBM

[Lis92] Erkki P. Liski. On a relationship between minimax-linear and admissible estimators in linear regression. Linear Algebra and its Applications, 176: 109–??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:1990:NEP

[Liu90] Bo Lian Liu. A note on the exponents of primitive (0, 1) matrices. Linear Algebra and its Applications, 140:45–51, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:1991:CPI

[Liu91]  Xing Ping Liu.  Conver-

gence of preconditioned iterative methods. Linear Algebra and its Applications, 146: 93–110, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:1996:BGC

[Liu96] Shu-Lin Liu. Bounds for the greatest characteristic root of a nonnegative matrix. Linear Algebra and its Applications, 239(1–3):151– 160, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1996&volume=239&issue=1-
3&aid=9400177.

Liu:1997:LMP

[Liu97] Shaowu Liu. Linear maps preserving idempotence on matrix modules over principal ideal domains. Linear Algebra and its Applications, 258(1–3): 219–231, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600203.

Liu:1999:EIP

[Liu99a] Bolian Liu. On exponent of indecomposability for primitive Boolean matrices. Linear Algebra and its Applications, 298(1–3):1–8, September 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
17/00006429.pdf;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/17/abstract.
html.

Liu:1999:SISa

[Liu99b]  Jianzhou Liu.  Some in-

equalities for singular values and eigenvalues of generalized Schur complements of products of matrices. Linear Algebra and its Applications, 286(1–3):209–221, January 1, 1999. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6267.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6267.

Liu:1999:SLP

[Liu99c] Jianzhou Liu. Some Löwner partial orders of Schur complements and Kronecker products of matrices. Linear Algebra and its Applications, 291(1–3):143–149, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6344.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6344.

Liu:1999:MRK

[Liu99d] Shuangzhe Liu. Matrix results on the Khatri–Rao and Tracy–Singh products. Linear Algebra and its Applications, 289(1–3):267–277, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6315.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6315.

Livshits:1995:NSM

[Liv95] Leo Livshits. A note on 0-1 Schur multipliers. Linear Algebra and its Applications, 222(1–3):15–22, June ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300268.

Lundquist:1991:LCP

[LJ91] Michael E. Lundquist and Charles R. Johnson. Linearly constrained positive definite completions. Linear Algebra and its Applications, 150: 195–207, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Laffey:1991:TGC

[LL91a] Thomas J. Laffey and Susan Lazarus. Two-generated commutative matrix subalgebras. Linear Algebra and its Applications, 147:249–273, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:1991:SMC

[LL91b] Jon Lee and Yew Sing Lee. A spectral method for concordant polyhedral faces. Linear Algebra and its Applications, 159:11–41, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1993:MIW

[LL93a] Hongxiang Li and Gangsong Leng. Matrix inequality with weights and its applications. Linear Algebra and its Applications, 185:273–??, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:1993:IAS

[LL93b] Lin Zhang Lu and Wen Wei Lin. An iterative algorithm for the solution of the discretetime algebraic Riccati equation. Linear Algebra and its Applications, 188/189:465– 488, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1998:ICM

[LLH^{+}98] Bishan Li, Lei Li, Masunori Harada, Hiroshi Niki, and Michael J. Tsatsomeros. An iterative criterion for Hmatrices. Linear Algebra and its Applications, 271(1–3): 179–190, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700267.

Lady:1995:NSN

[LLM95] George M. Lady, Thomas J.

Lundy, and John Maybee. Nearly sign-nonsingular matrices. Linear Algebra and its Applications, 220(1–3):229– 248, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400152. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Lee:1997:GFM

[LLS97] G.-Y. Lee, S.-G. Lee, and H.-G. Shin. On the Kgeneralized Fibonacci matrix Q_{k}^{∗}. Linear Algebra and its Applications, 251(1–3):73– 88, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500553.

Laffey:1992:EFT

[LM92] Thomas J. Laffey and Eleanor Meehan. An extension of a factorization theorem of Wedderburn to matrix rings. Linear Algebra and its Applications, 172:243–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:1994:FPU

[LM94a] Thomas J. Laffey and Eleanor Meehan. Factorization of polynomials using commuting matrices. Linear Algebra and its Applications, 196:85– 103, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1994:AMU

[LM94b] Chi-Kwong Li and Maria Emília Miranda. Antidiagonals of matrices in a unitary similarity orbit. Linear Algebra and its Applications, 197/198: 317–336, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Li:1995:PIN

[LM95] Chi-Kwong Li and Paras P. Mehta. Permutation invariant norms. Linear Algebra and its Applications, 219(1–3): 93–110, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300198.

Li:1997:MIP

[LM97] Chi-Kwong Li and Roy Mathias. Matrix inequalities and partial isometries that arise in X-ray crystallography. Linear Algebra and its Applications, 254(1–3):303– 314, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600172. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Li:1998:GED

[LM98a] Chi-Kwong Li and Roy Mathias. Generalized eigenvalues of a definite Hermitian matrix pair. Linear Algebra and its Applications, 271(1–3): 309–321, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700281.

Lundy:1998:UOC

[LM98b] Thomas J. Lundy and John Maybee. Uniformly oneconnected matrices and their inverses. Linear Algebra and its Applications, 268(1–3):227–242, January 1, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=268&issue=1-
                                        3&aid=9700045.

Laffey:1999:CTZ

[LM99] Thomas J. Laffey and Eleanor Meehan. A characterization of trace zero nonnegative 5 × 5 matrices. Linear Algebra and its Applications, 302-303(1–3):295–302, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        42/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/42/article.
                                        pdf.

Lundgren:1995:CGI

[LMMR95] J. Richard Lundgren, Sarah K. Merz, John S. Maybee, and Craig W. Rasmussen. A characterization of graphs with interval two-step graphs. Linear Algebra and its Applications, 217(1–3):203–223, March 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400173. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Lundy:1996:SIS

[LMOV96] Thomas Lundy, John S. Maybee, D. D. Olesky, and P. Van Den Driessche. Spectra and inverse sign patterns of nearly sign-nonsingular matrices. Linear Algebra and its Applications, 249(1–3):325– 339, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500363.

Laurie:1994:SI

[LMR94] C. Laurie, B. Mathes, and H. Radjavi. Sums of idempotents. Linear Algebra and its Applications, 208/209:175– 197, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lundgren:1995:CNC

[LMR95] J. Richard Lundgren, Sarah K.

Merz, and Craig W. Rasmussen. Chromatic numbers of competition graphs.

Linear Algebra and its Applications, 217(1–3):225–239, March 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400227. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Lustig:1991:CEP

[LMS91] Irvin J. Lustig, Roy E. Marsten, and David F. Shanno. Computational experience with a primal-dual interior point method for linear programming. Linear Algebra and its Applications, 152: 191–222, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Lundy:1995:IM

[LMS95] Thomas Lundy, John S. Maybee, and Asok K. Sen. S-inverse matrices. Linear Algebra and its Applications, 217(1–3):241–253, March 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-

3&aid=9400233. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Lundy:1996:MSN

[LMV96] Thomas J. Lundy, John Maybee, and James Van Buskirk. On maximal sign-nonsingular matrices. Linear Algebra and its Applications, 247(1–3):55– 81, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500054.

Liu:1990:ESS

[LMWZ90] Bo Lian Liu, Brendan D.

McKay, Nicholas C. Wormald, and Ke Min Zhang. The exponent set of symmetric primitive (0, 1) matrices with zero trace. Linear Algebra and its Applications, 133:121–131, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:1999:CFH

[LMX99] Wen-Wei Lin, Volker Mehrmann, and Hongguo Xu. Canonical forms for Hamiltonian and symplectic matrices and pencils. Linear Algebra and its Applications, 302-303(1–3):469–533, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        53/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/53/article.
                                        pdf.

Lancaster:1994:LRP

[LMY94] Peter Lancaster, Alexander S. Markus, and Qiang Ye. Low rank perturbations of strongly definitizable transformations and matrix polynomials. Linear Algebra and its Applications, 197/198:3–29, January 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Lancaster:1997:ONL

[LMZ97] P. Lancaster, A. S. Markus, and P. Zizler. The order of neutrality for linear operators on inner product spaces. Linear Algebra and its Applications, 259(1–3): 25–29, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600239.

Loiseau:1998:AKI

[LMZZ98] Jean Jacques Loiseau, Sabine Mondié, Ion Zaballa, and Petr Zagalak. Assigning the Kronecker invariants of a matrix pencil by row or column completions. Linear Algebra and its Applications, 278(1–3):327–336, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6089.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6089.

Lee:1991:NLR

[LN91] S. J. Lee and M. Z. Nashed.

Normed linear relations: domain decomposability, adjoint subspaces, and selections. Linear Algebra and its Applications, 153:135–159, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:1997:KCS

[LN97] Shuangzhe Liu and Heinz Neudecker. Kantorovich and Cauchy–Schwarz inequalities involving positive semidefinite matrices, and efficiency comparisons for a singular linear model. Linear Algebra and its Applications, 259(1–3): 209–221, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600284.

Littleboy:1998:MCL

[LN98] Darren M. Littleboy and N. K. Nichols. Modal coupling in linear control systems using robust eigenstructure assignment. Linear Algebra and its Applications, 275/276(1– 3):359–379, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6049.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6049. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Livshits:1993:IMO

[LO93] Leonya Livshits and Sing-Cheong Ong. On the invertibility of the map T → ST S^{−1} + S^{−1}T S and operator-norm inequalities. Linear Algebra and its Applications, 183:117–129, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:1990:LMW

[Loe90] Raphael Loewy. Linear maps which preserve a balanced nonsingular inertia class. Linear Algebra and its Applications, 134:165–179, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:1992:RPB

[Loe92] Raphael Loewy. Review of Perturbation Bounds for Matrix Eigenvalues by R. Bhatia. Linear Algebra and its Applications, 173:265–??, August 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:1997:ITL

[Loe97] R. Loewy. An inertia theorem for Lyapunov’s equation and the dimension of a controllability space. Linear Algebra and its Applications, 260 (1–3):1–7, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600197.

Logemann:1998:DES

[Log98] Hartmut Logemann. Destabilizing effects of small time delays on feedback-controlled descriptor systems. Linear Algebra and its Applications, 272(1–3):131–153, March 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700328.

Lomadze:1999:AVB

[Lom99] Vakhtang Lomadze. Applications of vector bundles to factorization of rational matrices. Linear Algebra and its Applications, 288(1–3): 249–258, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6326.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6326.

London:1995:DMS

[Lon95] David London. Diagonals of matrices stochastically similar to a given matrix. Linear Algebra and its Applications, 214(1–3):1–10, January 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300048.

London:1996:DMD

[Lon96] David London. On diagonals of matrices doubly stochastically similar to a given matrix. Linear Algebra and its Applications, 244(1–3):305– 340, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9688855.

London:1999:IMR

[Lon99] David London. Irreducible matrices with reducible principal submatrices. Linear Algebra and its Applications, 290(1–3):257–266, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6345.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6345.

Lopez:1994:BSC

[Lop94] L. Lopez. Bounds for the solutions of a class of tridiagonal linear systems. Linear Algebra and its Applications, 202: 221–233, April 15, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:1994:LPB

[LP94] Raphael Loewy and Stephen Pierce. Linear preservers of balanced singular inertia classes. Linear Algebra and its Applications, 201:61–77, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laurent:1995:PSR

[LP95a] Monique Laurent and Svatopluk Poljak. On a positive semidefinite relaxation of the cut polytope. Linear Algebra and its Applications, 223/224(1–3):439– 461, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500271. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Loewy:1995:LPB

[LP95b] Raphael Loewy and Stephen Pierce. Linear preservers of balanced nonsingular inertia classes. Linear Algebra and its Applications, 223/224(1–3): 483–499, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9300084. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Lopez:1997:TSC

[LP97] L. Lopez and T. Politi. Tridiagonal splittings in the conditioning and parallel solution of banded linear systems. Linear Algebra and its Applications, 251(1–3):249– 265, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500694.

Lee:1998:LOS

[LPS98] Sang-Gu Lee, Se-Won Park, and Han-Guk Seol. Linear operators strongly preserving matrices whose sign patterns require the Perron property. Linear Algebra and its Applications, 269(1):53– 63, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700056.

Liski:1992:BTD

[LPW92] Erkki P. Liski, Simo Puntanen, and Song Gui Wang. Bounds for the trace of the difference of the covariance matrices of the OLSE and BLUE. Linear Algebra and its Applications, 176:121–130, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1990:SEP

[LR90] Chi-Kwong Li and Leiba Rodman. Some extremal problems for positive definite matrices and operators. Linear Algebra and its Applications, 140: 139–154, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lerer:1993:SLE

[LR93] Leonid Lerer and Leiba Rodman. Sylvester and Lyapunov equations and some interpolation problems for rational matrix functions. Linear Algebra and its Applications, 185:83– 117, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:1994:ISC

[LR94a] Thomas J. Laffey and Robert Reams. Integral similarity and commutators of integral matrices. Linear Algebra and its Applications, 197/198:671– 689, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Loewy:1994:SSM

[LR94b] Raphael Loewy and Nizar Radwan. Spaces of symmetric matrices of bounded rank. Linear Algebra and its Applications, 197/198:189– 215, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Lancaster:1995:MSF

[LR95] P. Lancaster and L. Rodman.

Minimal symmetric factorizations of symmetric real and complex rational matrix functions. Linear Algebra and its Applications, 220(1–3):249– 282, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400151. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Lerer:1999:BRM

[LR99] Leonid Lerer and Leiba Rodman. Bezoutians of rational matrix functions, matrix equations and factorizations. Linear Algebra and its Applications, 302-303(1–3):105–133, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/32/article.
                                        pdf.

Langer:1997:NSA

[LRT97] H. Langer, A. C. M. Ran, and D. Temme. Nonnegative solutions of algebraic Riccati equations. Linear Algebra and its Applications, 261(1–3):317–352, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=261&issue=1-
                                        3&aid=9600444.

Leite:1999:GPS

[LRT99] R. S. Leite, T. R. W. Richa, and C. Tomei. Geometric proofs of some theorems of Schur–Horn type. Linear Algebra and its Applications, 286(1–3):149–173, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        286/1-3/6280.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=286&issue=
                                        1-3&aid=6280.

Lauterbach:1990:SPA

[LS90a] J. Lauterbach and P. Stahlecker.

Some properties of [tr(Q^{2p})]^{1/2p} with application to linear minimax estimation. Linear Algebra and its Applications, 127: 301–325, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lenferink:1990:GNR

[LS90b] H. W. J. Lenferink and M. N.

Spijker. A generalization of the numerical range of a matrix. Linear Algebra and its Applications, 140:251–266, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Labahn:1992:ITM

[LS92a] George Labahn and Tamir Shalom. Inversion of Toeplitz matrices with only two standard equations. Linear Algebra and its Applications, 175: 143–158, October 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lauterbach:1992:NMA

[LS92b] Jorg Lauterbach and Peter Stahlecker. A numerical method for an approximate minimax estimator in linear regression. Linear Algebra and its Applications, 176:91– ??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Landau:1993:BTD

[LS93a] L. J. Landau and R. F. Streater. On Birkhoff’s theorem for doubly stochastic completely positive maps of matrix algebras. Linear Algebra and its Applications, 193: 107–127, November 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lempken:1993:CSR

[LS93b] W. Lempken and R. Staszewski. A construction of ˆ3McL and some representation theory in characteristic 5. Linear Algebra and its Applications, 192: 205–234, October 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Labahn:1994:ITS

[LS94] George Labahn and Tamir Shalom. Inversion of Toeplitz structured matrices using only standard equations. Linear Algebra and its Applications, 207:49–70, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lovasz:1995:CLB

[LS95a] László Lovász and Ákos Seress. The cocycle lattice of binary matroids. II. Linear Algebra and its Applications, 226/228 (1–3):553–565, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500243.

Lundgren:1995:PCG

[LS95b] J. Richard Lundgren and Hans Schneider, editors. Proceedings of a Conference on Graphs and Matrices in honor of John Maybee, Held in Boulder, Colorado on 7–8 May 1993, volume 217 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1995. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:1997:NCS

[LS97] W. Q. Liu and V. Sreeram.

A new characterization on stability bounds for singularly perturbed systems. Linear Algebra and its Applications, 263(1–3):377–388, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600597.

Leibovici:1998:SVD

[LS98] Didier Leibovici and Robert Sabatier. A singular value decomposition of a k-way array for a principal component analysis of multiway data, PTA-k. Linear Algebra and its Applications, 269(1):307– 329, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9781516.

Lewis:1999:DMC

[LS99] Debra Lewis and Mike Shub. The distribution of the maximum condition number on great circles through a fixed 2 × 2 real matrix. Linear Algebra and its Applications, 297(1–3):193–202, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/29/article.
                                        pdf.

Loewy:1991:PES

[LSJ91] R. Loewy, D. R. Shier, and C. R. Johnson. Perron eigenvectors and the symmetric transportation polytope. Linear Algebra and its Applications, 150:139–155, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Li:1998:EHN

[LSS98] Chi-Kwong Li, Ilya Spitkovsky, and Sudheer Shukla. Equality of higher numerical ranges of matrices and a conjecture of Kippenhahn on Hermitian pencils. Linear Algebra and its Applications, 270(1–3):323– 349, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700251.

Losche:1998:MBN

[LST98] Ralf Lösche, Hubert Schwetlick, and Gisela Timmermann. A modified block Newton iteration for approximating an invariant subspace of a symmetric matrix. Linear Algebra and its Applications, 275/276(1–3):381–400, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6013.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-

                                        276&issue=1-3&aid=6013. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Li:1991:INR

[LT91a] Chi-Kwong Li and Nam-Kiu Tsing. G-invariant norms and G(c)-radii. Linear Algebra and its Applications, 150:179–194, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Li:1991:DBS

[LT91b] Chi-Kwong Li and Nam-Kiu Tsing. Duality between some linear preserver problems. III. c-spectral norms and (skew)- symmetric matrices with fixed singular values. Linear Algebra and its Applications, 143: 67–97, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:1995:LMR

[LT95] Chi-Kwong Li and Nam-Kiu Tsing. Linear maps relating different unitary similarity orbits or different generalized numerical ranges. Linear Algebra and its Applications, 223/224(1–3):463– 481, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500105. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Li:1997:DDD

[LT97a] Bishan Li and M. J. Tsatsomeros. Doubly diagonally dominant matrices. Linear Algebra and its Applications, 261(1–3):221–235, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600406.

Luo:1997:GPT

[LT97b] Xiao-Dong Luo and Paul Tseng. On a global projectiontype error bound for the linear complementarity problem. Linear Algebra and its Applications, 253(1–3):251– 278, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500707.

Li:1997:NHV

[LTZ97] Chi-Kwong Li, Nam-Kiu Tsing, and Fuzhen Zhang. Norm hull of vectors and matrices. Linear Algebra and its Applications, 257(1–3):1– 27, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600099.

Lu:1998:ESM

[Lu98] Ya Yan Lu. Exponentials of symmetric matrices through tridiagonal reductions. Linear Algebra and its Applications, 279(1–3):317–324, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6115.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6115.

Lubary:1998:MSS

[Lub98] José A. Lubary. Multiplicity of solutions of second-order linear differential equations on networks. Linear Algebra and its Applications, 274(1–3): 301–315, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1998&volume=274&issue=1-
3&aid=9700348.

Lundgren:1995:BHJ

[Lun95] J. Richard Lundgren. A brief history of John Maybee’s career. Linear Algebra and its Applications, 217(1– 3):1–4, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400315. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

LeBruyn:1991:APL

[LV91a] L. Le Bruyn and M. Van den Bergh. Algebraic properties of linear cellular automata. Linear Algebra and its Applications, 157:217–234, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Linzer:1991:SAC

[LV91b] Elliot Linzer and Martin Vetterli. Structure and arithmetic complexity of products and inverses of Toeplitz matrices. Linear Algebra and its Applications, 145:159–175, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lai:1997:SMW

[LV97] Shang-Hong Lai and B. C. Vemuri. Sherman–Morrison– Woodbury-formula-based algorithms for the surface smoothing problem. Linear Algebra and its Applications, 265(1–3):203–229, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9780366.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9780366.

Lobo:1998:ASO

[LVBL98] Miguel Sousa Lobo, Lieven Vandenberghe, Stephen Boyd, and Herve Lebret. Applications of second-order cone programming. Linear Algebra and its Applications, 284 (1–3):193–228, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6153.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6153.

Lagarias:1995:FCG

[LW95] Jeffrey C. Lagarias and Yang Wang. The finiteness conjecture for the generalized spectral radius of a set of matrices. Linear Algebra and its Applications, 214(1–3):17– 42, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300052.

Lee:1996:NCF

[LW96] Jeffrey M. Lee and David A.

Weinberg. A note on canonical forms for matrix congruence. Linear Algebra and its Applications, 249(1–3):207–215, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500357.

Li:1997:SCP

[LW97] Chi-Kwong Li and Hugo J.

Woerdeman. Special classes of positive and completely positive maps. Linear Algebra and its Applications, 255(1–3): 247–258, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1997&volume=255&issue=1-
                                        3&aid=9600776.

Lei:1998:IND

[LW98a] Tian-Gang Lei and Bo-Ying Wang. Indices and nonzero decomposable elements of a symmetry class of tensors. Linear Algebra and its Applications, 279(1–3):21–30, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        279/1-3/6103.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6103.

Lumsdaine:1998:SPB

[LW98b] Andrew Lumsdaine and Deyun Wu. Spectra and pseudospectra of block Toeplitz matrices. Linear Algebra and its Applications, 272(1–3):103– 130, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=272&issue=1-
                                        3&aid=9700311.

Liu:1999:SISb

[LW99] Jianzhou Liu and Jian Wang. Some inequalities for Schur complements. Linear Algebra and its Applications, 293 (1–3):233–241, May 15, 1999.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6394.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6394.

Lin:1999:CON

[LWX99] Wen-Wei Lin, Chern-Shuh Wang, and Quan-Fu Xu. On the computation of the optimal H_{∞} norms for two feedback control problems. Linear Algebra and its Applications, 287(1–3):223–255, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6309.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6309.

Lin:1999:RCR

[LWYS99] Chong Lin, Jian Liang Wang, Guang-Hong Yang, and C. B. Soh. Robust controllability and robust closed-loop stability with static output feedback for a class of uncertain descriptor systems. Linear Algebra and its Applications, 297(1–3):133–155, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/25/article.
                                        pdf.

Lam:1993:LSR

[LY93a] Kee Yuen Lam and Paul Yiu. Linear spaces of real matrices of constant rank. Linear Algebra and its Applications, 195: 69–79, December 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lancaster:1993:RRL

[LY93b] Peter Lancaster and Qiang Ye. Rayleigh–Ritz and Lanczos methods for symmetric matrix pencils. Linear Algebra and its Applications, 185:173– 201, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:1993:SAM

[LY93c] W. W. Lin and S. S. You. A symplectic acceleration method for the solution of the algebraic Riccati equation on a parallel computer. Linear Algebra and its Applications, 188/189:437–463, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lyubich:1995:PFT

[Lyu95]  Yu. Lyubich.  Perron–

Frobenius theory for finitedimensional spaces with a hyperbolic cone. Linear Algebra and its Applications, 220(1–3): 283–309, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400126. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Li:1996:UTB

[LZ96] Wen Li and Mou-Cheng Zhang. On upper triangular block weak regular splittings of a singular M-matrix. Linear Algebra and its Applications, 233(1–3):175–187, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400068.

Li:1997:NUB

[LZ97a] Jiong-Sheng Li and Xiao-Dong Zhang. A new upper bound for eigenvalues of the Laplacian matrix of a graph. Linear Algebra and its Applications, 265 (1–3):93–100, November 1, 1997. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600592.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600592.

Liu:1997:MPE

[LZ97b] Jianzhou Liu and Li Zhu. A minimum principle and estimates of the eigenvalues for Schur complements of positive semidefinite Hermitian matrices. Linear Algebra and its Applications, 265(1–3):123–145, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600595.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600595.

Leng:1998:GST

[LZ98a] Gangsong Leng and Yao Zhang. The generalized sine theorem and inequalities for simplices. Linear Algebra and its Applications, 278(1– 3):237–247, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/

278/1-3/6096.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6096.

Li:1998:LEG

[LZ98b] Jiong-Sheng Li and Xiao-Dong Zhang. On the Laplacian eigenvalues of a graph. Linear Algebra and its Applications, 285(1–3):305–307, December 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6260.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6260.

Lim:1999:MCG

[LZ99] Arthur Lim and Chen-Bo Zhu. Multiresolution on compact groups. Linear Algebra and its Applications, 293 (1–3):15–38, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6367.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6367.

Li:1997:DSJ

[LZW97] T. Y. Li, Zhinan Zhang, and Tianjun Wang. Determining the structure of the Jordan normal form of a matrix by symbolic computation. Linear Algebra and its Applications, 252(1–3):221– 259, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=252&issue=1-
                                        3&aid=9500669.

Ma:1993:WIS

[Ma93] Wenxin Ma. The weak identity T_{2}n for n × n symmetric matrices. Linear Algebra and its Applications, 180:239–??, February 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maesumi:1996:ELB

[Mae96] Mohsen Maesumi. An efficient lower bound for the generalized spectral radius of a set of matrices. Linear Algebra and its Applications, 240(1– 3):1–7, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=240&issue=1-
                                        3&aid=9400171.

Maher:1992:SNI

[Mah92] P. J. Maher. Some norm inequalities concerning generalized inverses. Linear Algebra and its Applications, 174: 99–??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mahony:1996:CNM

[Mah96] R. E. Mahony. The constrained Newton method on a Lie group and the symmetric eigenvalue problem. Linear Algebra and its Applications, 248(1–3):67–89, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500171.

Mai:1991:CAP

[Mai91] Tsun-Zee Mai. Composite adaptive procedure for solving large sparse linear systems. Linear Algebra and its Applications, 154/156:509– 534, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Malek:1992:NPS

[Mal92] Massoud Malek. Notes on permanental and subpermanental inequalities. Linear Algebra and its Applications, 174:

53–??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Malyshev:1993:PAS

[Mal93] Alexander N. Malyshev. Parallel algorithm for solving some spectral problems of linear algebra. Linear Algebra and its Applications, 188/189: 489–520, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maliakas:1996:CFI

[Mal96]  Mihalis Maliakas.  On a

character formula involving Borel subgroups. Linear Algebra and its Applications, 247(1–3):273–275, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500116.

Malamud:1998:CMP

[Mal98] S. M. Malamud. On a class of multiparameter perturbations of positive definite operators with fixed bounds on their spectrum. Linear Algebra and its Applications, 274(1–3): 239–257, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700345.

Mangasarian:1992:GEB

[Man92] O. L. Mangasarian. Global error bounds for monotone affine variational inequality problems. Linear Algebra and its Applications, 174:153– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martins:1990:NCM

[Mar90] M. Madalena Martins. A note on the convergence of the MSOR method. Linear Algebra and its Applications, 141: 223–226, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcus:1991:MML

[Mar91] Marvin Marcus. Multilinear methods in linear algebra. Linear Algebra and its Applications, 150:41–59, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Marek:1992:CWN

[Mar92] Ivo Marek. Collatz–Wielandt numbers in general partially ordered spaces. Linear Algebra and its Applications, 173:

165–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MarquesdeSa:1994:EFD

[Mar94a] Eduardo Marques de Sá. Exposed faces and duality for symmetric and unitarily invariant norms. Linear Algebra and its Applications, 197/198: 429–450, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

MarquesdeSa:1994:FTU

[Mar94b] Eduardo Marques de Sá. Faces and traces of the unit ball of a symmetric gauge function. Linear Algebra and its Applications, 197/198:349– 395, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

MarquesdeSa:1994:FUB

[Mar94c] Eduardo Marques de Sá. Faces of the unit ball of a unitarily invariant norm. Linear Algebra and its Applications, 197/198:451–493, January 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Marek:1995:SRO

[Mar95a]  Ivo Marek.  On square

roots of M-operators. Linear Algebra and its Applications, 223/224(1–3):501– 520, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400358. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Markham:1995:EED

[Mar95b] T. Markham. Erratum: “On the eigenvalues and diagonal entries of a Hermitian matrix” [Linear Algebra Appl. 179 (1993), 7–10, MR 93k:15012] by E. Fu and Markham. Linear Algebra and its Applications, 220(1–3):6, April 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MarquesdeSa:1996:FUB

[Mar96] Eduardo Marques de Sá. The faces of the unit balls of cnorms and c-spectral norms. Linear Algebra and its Applications, 246(1–3):177–189, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400351.

MarquesdeSa:1997:SUB

[Mar97] Eduardo Marques de Sá. The stability of the unit balls of symmetric and unitarily invariant norms. Linear Algebra and its Applications, 250(1–3):207–218, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500597.

Marcus:1998:DRO

[Mar98] Marvin Marcus. On a determinant result of I. Olkin. Linear Algebra and its Applications, 277(1–3):237–238, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6074.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6074.

Markiewicz:1999:SPD

[Mar99a] Augustyn Markiewicz. Simultaneous polar decomposition of rectangular complex matrices. Linear Algebra and its Applications, 289(1–3):279–284, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6223.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6223.

Marwaha:1999:DSN

[Mar99b] Alka Marwaha. Decomposability and structure of nonnegative bands in M_{n}(R). Linear Algebra and its Applications, 291(1–3):63–82, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6338.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6338.

Mascarenhas:1994:SES

[Mas94] Walter F. Mascarenhas. The structure of the eigenvectors of sparse matrices. Linear Algebra and its Applications, 207: 1–20, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1990:SNN

[Mat90] Roy Mathias. The spectral norm of a nonnegative matrix. Linear Algebra and its Applications, 139:269–284, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1991:ETP

[Mat91] Roy Mathias. The equivalence of two partial orders on a convex cone of positive semidefinite matrices. Linear Algebra and its Applications, 151:27– 55, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathai:1993:AHF

[Mat93a] A. M. Mathai. Appell’s and Humbert’s functions of matrix arguments. Linear Algebra and its Applications, 183: 201–221, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1993:AGH

[Mat93b] Roy Mathias. An arithmeticgeometric-harmonic mean inequality involving Hadamard products. Linear Algebra and its Applications, 184:71–??, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1995:EAG

[Mat95]  Roy Mathias.  Erratum:

“An arithmetic-geometric-harmonic mean inequality involving Hadamard products” [Linear Algebra Appl. 184 (1993), 71–78, MR 94b:15019]. Linear Algebra and its Applications, 220(1–3):4, April 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miao:1993:MMP

[MBI93] Jian Ming Miao and Adi Ben-Israel. Minors of the Moore– Penrose inverse. Linear Algebra and its Applications, 195: 191–207, December 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miao:1994:ASL

[MBI94] Jian Ming Miao and Adi Ben-Israel. On ℓ_{p}-approximate solutions of linear equations. Linear Algebra and its Applications, 199:305–327, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miao:1995:GBA

[MBI95] Jian Ming Miao and Adi Ben-Israel. The geometry of basic, approximate, and minimumnorm solutions of linear equations. Linear Algebra and its Applications, 216(1–3):25– 41, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300093.

Miao:1996:PCA

[MBI96] Jianming Miao and Adi Ben-Israel. Product cosines of angles between subspaces. Linear Algebra and its Applications, 237/238(1–3):71– 81, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500425. Special

issue honoring Calyampudi Radhakrishna Rao.

McCullough:1993:MSC

[McC93]  Scott McCullough.  Mini-

mal separators of 2-Chordal graphs. Linear Algebra and its Applications, 184:187–199, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:1994:PIT

[McD94] J. J. McDonald. A product index theorem with applications to splittings of Mmatrices. Linear Algebra and its Applications, 197/198:511– 530, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

McEachin:1993:CGS

[McE93] R. McEachin. Closing the gap in a subspace perturbation bound. Linear Algebra and its Applications, 180:7–16, February 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McEachin:1994:ASC

[McE94] R. McEachin. Analyzing specific cases of an operator inequality. Linear Algebra and its Applications, 208/209:343– 365, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moyssiadis:1995:OEF

[MCK95] C. Moyssiadis, S. Chadjiconstantinidis, and S. Kounias. A-optimization of exact firstorder saturated designs for N ≡ 1 mod 4 observations. Linear Algebra and its Applications, 216(1–3):159–176, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300130.

McLaren:1990:ARD

[McL90] Mara Lee McLaren. Asymptotic results for decomposing a likelihood-ratio statistic into separate components. Linear Algebra and its Applications, 127:601–616, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merikoski:1992:MNM

[MD92] Jorma Kaarlo Merikoski and Graciano De Oliveira. On k-major norms and k-minor antinorms. Linear Algebra and its Applications, 176:197– ??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mongeau:1993:CDM

[MDD93] Jean-Pierre Mongeau, Gilles Deslauriers, and Serge Dubuc. Continuous and differentiable multidimensional iterative interpolation. Linear Algebra and its Applications, 180:95– 120, February 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ma:1998:NEP

[ME98] Yanyuan Ma and Alan Edelman. Nongeneric eigenvalue perturbations of Jordan blocks. Linear Algebra and its Applications, 273(1– 3):45–63, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=273&issue=1-
                                        3&aid=9700342.

Mecheri:1998:DR

[Mec98] Salah Mecheri. Derivation ranges. Linear Algebra and its Applications, 279(1–3):

31–38, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6116.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6116.

Mehrotra:1991:FVS

[Meh91] Sanjay Mehrotra. On finding a vertex solution using interior point methods. Linear Algebra and its Applications, 152: 233–253, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Mehrmann:1996:STU

[Meh96] Volker Mehrmann. A step toward a unified treatment of continuous and discrete time control problems. Linear Algebra and its Applications, 241/243(1–3):749–779, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500257. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Mehrmann:1998:CCR

[Meh98] V. Mehrmann. Chemnitz 1996 Conference report. Linear Algebra and its Applications, 275/276(1–3):627–??, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Melman:1999:SAT

[Mel99]  A. Melman.  A symmet-

ric algorithm for Toeplitz systems. Linear Algebra and its Applications, 301 (1–3):145–152, November 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/115/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/115/25/34/article.
                                        pdf.

Mendlovitz:1997:ESZ

[Men97] Mark A. Mendlovitz. Eigenvector saddlepoints and zero properties of the eigenpolynomial factors. Linear Algebra and its Applications, 259(1–3): 281–298, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600292.

Merikoski:1990:ANM

[Mer90] Jorma Kaarlo Merikoski. On ℓ_{p_{1},p_{2}} antinorms of nonnegative matrices. Linear Algebra and its Applications, 140:31– 44, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merikoski:1991:NAC

[Mer91a] Jorma Kaarlo Merikoski. On c-norms and c-antinorms on cones. Linear Algebra and its Applications, 150:315–329, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Merris:1991:AAT

[Mer91b] Russell Merris. Almost all trees are co-immanantal. Linear Algebra and its Applications, 150:61–66, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Merris:1992:UEG

[Mer92] Russell Merris. Unimodular equivalence of graphs. Linear Algebra and its Applications, 173:181–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merris:1994:DMG

[Mer94a] Russell Merris. Degree maximal graphs are Laplacian integral. Linear Algebra and its Applications, 199:381–389, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merris:1994:LMG

[Mer94b] Russell Merris. Laplacian matrices of graphs: A survey. Linear Algebra and its Applications, 197/198:143– 176, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Merris:1994:NUC

[Mer94c]  Russell Merris.  A note

on unimodular congruence of graphs. Linear Algebra and its Applications, 201:57–60, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merris:1998:LGE

[Mer98a] Russell Merris. Laplacian graph eigenvectors. Linear Algebra and its Applications, 278(1–3):221–236, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6080.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6080.

Merris:1998:NLG

[Mer98b]  Russell Merris.  A note

on Laplacian graph eigenvalues. Linear Algebra and its Applications, 285(1–3): 33–35, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6259.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6259.

Messaoudi:1994:MRP

[Mes94] A. Messaoudi. Matrix recursive projection and interpolation algorithms. Linear Algebra and its Applications, 202: 71–89, April 15, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meshulam:1995:TPF

[Mes95a] Roy Meshulam. On twoparameter families of symmetric matrices. Linear Algebra and its Applications, 216(1–3): 93–96, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=216&issue=1-
                                        3&aid=9300108.

Meshulam:1995:SHM

[Mes95b] Roy Meshulam. Spaces of Hankel matrices over finite fields. Linear Algebra and its Applications, 218(1–3):73– 76, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=218&issue=1-
                                        3&aid=9300158.

Messaoudene:1997:UMB

[Mes97a] Mohamed Messaoudene. The unit mass ball of threevectors in R^{6}. Linear Algebra and its Applications, 265(1–3):247–298, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/265/1-3/9600629.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=265&
                                        issue=1-3&aid=9600629.

Messaoudi:1997:MEA

[Mes97b] A. Messaoudi. Matrix extrapolation algorithms. Linear Algebra and its Applications, 256(1–3):49–73, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781112.

Metz:1997:SOA

[Met97] Volker Metz. Shorted operators: An application in potential theory. Linear Algebra and its Applications, 264(1–3):439–455, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600303.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600303.

Meyer:1997:NEA

[Mey97] Renate Meyer. Nonlinear eigenvector algorithms for local optimization in multivariate data analysis. Linear Algebra and its Applications, 264(1–3):225–246, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600635.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.

                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9600635.

Missirlis:1994:EOR

[MG94] N. M. Missirlis and N. G. Gaitanos. Extensions of the Ostrowski-Reich theorem. Linear Algebra and its Applications, 207:147–157, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moreland:1999:IHS

[MG99] T. Moreland and S. Gudder. Infima of Hilbert space effects. Linear Algebra and its Applications, 286 (1–3):1–17, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        286/1-3/6236.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=286&issue=
                                        1-3&aid=6236.

Gao:1992:CGD

[mGhW92] Yi ming Gao and Xiao hui Wang. Criteria for generalized diagonally dominant matrices and M-matrices. Linear Algebra and its Applications, 169:257–??, May 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitra:1992:POB

[MH92] Sujit Kumar Mitra and Robert E. Hartwig. Partial orders based on outer inverses. Linear Algebra and its Applications, 176:3–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mani:1996:BBI

[MH96] Vaidyanath Mani and Robert E.

Hartwig. Black box interpolation. II. the one variable derogatory case. Linear Algebra and its Applications, 249(1–3):229–253, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500359.

Mani:1997:GPB

[MH97] Vaidyanath Mani and Robert E.

Hartwig. Generalized polynomial bases and the Bezoutian. Linear Algebra and its Applications, 251(1–3):293–320, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500585.

Moore:1993:SSF

[MHR93] Sean S. B. Moore, Dennis M. Healy, Jr., and Daniel N. Rockmore. Symmetry stabilization for fast discrete monomial transforms and polynomial evaluation. Linear Algebra and its Applications, 192: 249–299, October 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Miao:1991:GEM

[Mia91] Jian Ming Miao. General expressions for the Moore– Penrose inverse of a 2 × 2 block matrix. Linear Algebra and its Applications, 151: 1–15, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Michael:1991:SMG

[Mic91] T. S. Michael. The structure matrix and a generalization of Ryser’s maximum term rank formula. Linear Algebra and its Applications, 145:21– 31, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Michael:1993:SMC

[Mic93] T. S. Michael. The structure matrix of the class of rmultigraphs with a prescribed degree sequence. Linear Algebra and its Applications, 183: 155–??, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Michaletzky:1997:NHA

[Mic97]  György Michaletzky.  A

note on Hankel approximation of all-pass functions. Linear Algebra and its Applications, 255(1–3):19–55, April 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9500775.

Milaszewicz:1995:CTM

[Mil95] J. P. Milaszewicz. Comparison theorems for monotone Newton-Fourier iterations and applications in functional elimination. Linear Algebra and its Applications, 220(1–3): 343–357, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400313. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Minc:1994:MPD

[Min94] Henryk Minc. Minimum permanents of doubly stochastic matrices with prescribed zero entries on the main diagonal. Linear Algebra and its Applications, 201:135–154, April 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mirman:1992:SCO

[Mir92] Boris Mirman. A source of counterexamples in operator theory and how to construct them. Linear Algebra and its Applications, 169:49–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miranker:1997:CAM

[Mir97]  Willard L. Miranker.  A

cache for an associative memory. Linear Algebra and its Applications, 251(1–3):181– 188, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=251&issue=1-
                                        3&aid=9500606.

Mirman:1998:NRP

[Mir98]  Boris Mirman.  Numer-

ical ranges and Poncelet Curves. Linear Algebra and its Applications, 281(1–3): 59–85, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6158.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6158.

Mitra:1990:PSL

[Mit90] Sujit Kumar Mitra. A pair of simultaneous linear matrix equations A_{1}XB_{1} = C_{1}, A_{2}XB_{2} = C_{2} and a matrix programming problem. Linear Algebra and its Applications, 131:107–123, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitra:1991:MPO

[Mit91] Sujit Kumar Mitra. Matrix partial order through generalized inverses: unified theory. Linear Algebra and its Applications, 148:237–263, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitra:1994:ST

[Mit94] Sujit Kumar Mitra. Separation theorems. Linear Algebra and its Applications, 208/209: 239–256, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitra:1996:NSMb

[Mit96] Sujit Kumar Mitra. Noncore square matrices miscellany. Linear Algebra and its Applications, 249(1–3):47–66, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500256.

Mitra:1997:ENS

[Mit97]  S. K. Mitra.  Erratum:

“The nonunique shorted matrix” [Linear Algebra Appl. 237/238 (1996), 41–70, MR 97a:15012] by S. K. Mitra and K. M. Prasad. Linear Algebra and its Applications, 260 (1–3):323, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitra:1999:DPI

[Mit99] Sujit Kumar Mitra. Diagrammatic presentation of inner and outer inverses: Sdiagrams. Linear Algebra and its Applications, 287(1– 3):271–288, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6302.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6302.

Mizuno:1991:IOP

[Miz91]  Shinji Mizuno.  O(n^{ρ}L)-

iteration and O(n^{3}L)-operation potential reduction algorithms for linear programming. Linear Algebra and its Applications, 152:155–168, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Jiang:1991:SIC

[mJS91] Zhi min Jiang and Jia Yu Shao. On the set of indices of convergence for reducible matrices. Linear Algebra and its Applications, 148:265–278, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meenakshi:1998:EM

[MK98] A. R. Meenakshi and S. Krishnamoorthy. On k-EP matrices. Linear Algebra and its Applications, 269(1):219– 232, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700066.

Martinez-Legaz:1990:LOL

[MLS90] Juan-Enrique Martínez-Legaz and Ivan Singer. Lexicographical order, lexicographical index, and linear operators. Linear Algebra and its Applications, 128:65–95, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martinez-Legaz:1991:CPL

[MLS91] Juan-Enrique Martínez-Legaz and Ivan Singer. Compatible preorders and linear operators on R^{n}. Linear Algebra and its Applications, 153:53– 66, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martinez-Legaz:1995:ECP

[MLS95] J.-E. Martínez-Legaz and I. Singer. Erratum: “Compatible preorders and linear operators on R^{n}” [Linear Algebra Appl. 153 (1991), 53–66, MR 92e:06018]. Linear Algebra and its Applications, 220 (1–3):5–6, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martinez-Legaz:1998:CCF

[MLS98] Juan-Enrique Martínez-Legaz and Ivan Singer. On Φconvexity of convex functions. Linear Algebra and its Applications, 278(1–3): 163–181, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6101.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6101.

Milaszewicz:1993:NM

[MM93a] J. P. Milaszewicz and L. P.

Moledo. On nonsingular Mmatrices. Linear Algebra and its Applications, 195:1–8, December 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Misra:1993:BRS

[MM93b] Pradeep Misra and Thulasinath Manickam. Balanced realization of separabledenominator multidimensional systems. Linear Algebra and its Applications, 188/189:521– 547, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Monga-Made:1993:ABB

[MM93c] Magolu Monga-Made. Analytical bounds for block approximate factorization methods. Linear Algebra and its Applications, 179:33–??, January 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Markus:1994:FFP

[MM94] Alexander Markus and Vladimir Matsaev. The failure of factorization of positive matrix functions on noncircular contours. Linear Algebra and its Applications, 208/209:231– 237, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moeller:1996:ARFa

[MM96a] Thomas L. Moeller and Jaime Milstein. Algebraic representations for finite-state machines. I. monoid-ring formulation. Linear Algebra and its Applications, 239(1–3):109– 126, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400174.

Moeller:1996:ARFb

[MM96b] Thomas L. Moeller and Jaime Milstein. Algebraic representations for finite-state machines. II module formulation. Linear Algebra and its Applications, 247(1–3):133– 150, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500110.

Moeller:1998:GAS

[MM98] Thomas L. Moeller and Jaime Milstein. Generalized algebraic structures for the representation of discrete systems. Linear Algebra and its Applications, 274(1–3):161– 191, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700327.

Margulies:1999:NRS

[MM99] Caryl Margulies and William Margulies. Nonlinear resonance set for nonlinear matrix equations. Linear Algebra and its Applications, 293(1–3):187–197, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6388.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6388.

Mallol:1991:ABI

[MMO91] Cristian Mallol, Artibano Micali, and Moussa Ouattara. Sur les algèbres de Bernstein. IV. (French) [on Bernstein algebras. IV]. Linear Algebra and its Applications, 158:1– 26, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mackey:1999:EMS

[MMP99] D. Steven Mackey, Niloufer Mackey, and Srdjan Petrovic. Is every matrix similar to a Toeplitz matrix? Linear Algebra and its Applications, 297(1–3):87–105, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/104/17/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/104/17/23/article.
pdf.

Mas:1996:NPR

[MMPS96] Jose Mas, Violeta Migallon, Jose Penades, and Daniel B. Szyld. Nonstationary parallel relaxed multisplitting methods. Linear Algebra and its Applications, 241/243(1–3):733–747, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500583. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Magolu:1991:CAB

[MN91] Monga-Made Magolu and Yvan Notay. On the conditioning analysis of block approximate factorization methods. Linear Algebra and its Applications, 154/156:583– 599, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:1993:RMM

[MNS93] J. J. McDonald, M. Neumann, and H. Schneider. Resolvents of minus M-matrices and splittings of M-matrices. Linear Algebra and its Applications, 195:17–33, December 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:1995:IMI

[MNST95] J. J. McDonald, M. Neumann, H. Schneider, and M. J. Tsatsomeros. Inverse Mmatrix inequalities and generalized ultrametric matrices. Linear Algebra and its Applications, 220(1–3):321– 341, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400077. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Micali:1991:DAT

[MO91] Artibano Micali and Moussa Ouattara. Dupliquée d’une algèbre et le théorème d’Etherington. (French) [duplication of an algebra and Etherington’s theorem]. Linear Algebra and its Applications, 153:193–207, 1991. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Micali:1995:SAB

[MO95] Artibano Micali and Moussa Ouattara. Structure des algèbres de Bernstein. (French) [structure of Bernstein algebras]. Linear Algebra and its Applications, 218(1–3):77– 88, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=218&issue=1-
                                        3&aid=9300159.

Matsaev:1996:CDK

[MO96] Vladimir Matsaev and Vadim Olshevsky. Cyclic dimensions, kernel multiplicities, and Gohberg-Kaashoek numbers. Linear Algebra and its Applications, 239(1–3):161– 174, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=239&issue=1-
                                        3&aid=9400179.

Micali:1997:AGD

[MO97] Artibano Micali and Moussa Ouattara. Algèbres Génétiques dimensionnellement nilpotentes. (French) [Dimensionallynilpotent genetic matrices].

Linear Algebra and its Applications, 266(1–3):271–290, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9786524.

Molnar:1996:CSI

[Mol96] Lajos Molnár. A condition for a subspace of B(H) to be an ideal. Linear Algebra and its Applications, 235(1–3): 229–234, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400143.

Molnar:1999:SLP

[Mol99a] Lajos Molnár. Some linear preserver problems on B ( H ) concerning rank and corank. Linear Algebra and its Applications, 286(1– 3):311–321, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6299.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1999&volume=286&issue=
                                        1-3&aid=6299.

Molnar:1999:SMP

[Mol99b] Lajos Molnár. Some multiplicative preservers on B(H). Linear Algebra and its Applications, 301(1–3):1–13, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/115/25/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/115/25/25/article.
                                        pdf.

Mom:1998:EDM

[Mom98] Alain Mom. Eigenstructure of distance matrices with an equal distance subset. Linear Algebra and its Applications, 280(2–3):245–251, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        280/2-3/6148.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=280&issue=
                                        2-3&aid=6148.

Moorhouse:1995:TGS

[Moo95] G. Eric Moorhouse. Twographs and skew two-graphs in finite geometries. Linear Algebra and its Applications, 226/228(1–3):529–551, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500242.

Mehrmann:1999:RBP

[MOPvdD99] V. Mehrmann, D. D. Olesky, T. X. T. Phan, and P. van den Driessche. Relations between Perron–Frobenius results for matrix pencils. Linear Algebra and its Applications, 287 (1–3):257–269, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6204.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6204.

Mathes:1991:LSN

[MOR91a] Ben Mathes, Matjaž Omladič, and Heydar Radjavi. Linear spaces of nilpotent matrices. Linear Algebra and its Applications, 149:215–225, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Morgan:1991:CIE

[Mor91b] Ronald B. Morgan. Computing interior eigenvalues of large matrices. Linear Algebra and its Applications, 154/156:

289–309, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Morris:1997:EP

[Mor97] Walter D. Morris, Jr. Extended P -pairs. Linear Algebra and its Applications, 257(1–3):211–223, May ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=257&issue=1-
                                        3&aid=9600167.

Mosler:1994:MED

[Mos94] Karl Mosler. Majorization in economic disparity measures. Linear Algebra and its Applications, 199:91–114, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mosler:1995:EME

[Mos95] Karl Mosler. Erratum: “Majorization in economic disparity measures” [linear algebra appl. 199 (1994), 91–114]. Linear Algebra and its Applications, 220(1–3):3–4, April 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mossadeq:1996:FRG

[Mos96] A. El Mossadeq. Fractions de résolution w et groupes abéliens finis. (French) [resolutionw fractions and finite Abelian groups]. Linear Algebra and its Applications, 237/238(1– 3):539–547, April ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.elsevier.com/cgi-bin/
cas/tree/store/laa/cas_
sub/browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500689. Special

issue honoring Calyampudi Radhakrishna Rao.

Maybee:1992:CEM

[MOTV92] John S. Maybee, D. D. Olesky, Michael Tsatsomeros, and P. Van Den Driessche. Combinatorial eigenvalues of matrices. Linear Algebra and its Applications, 175:213–??, October 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:1997:RPM

[MOTV97] J. J. McDonald, D. D. Olesky, M. J. Tsatsomeros, and P. Van Den Driessche. Ray patterns of matrices and nonsingularity. Linear Algebra and its Applications, 267(1–3):359–373, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700041.

Mathias:1990:EBL

[MP90] Roy Mathias and Jong-Shi Pang. Error bounds for the linear complementarity problem with a P -matrix. Linear Algebra and its Applications, 132:123–136, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miller:1993:ESG

[MP93] Zevi Miller and Dan Pritikin. Eigenvalues and separation in graphs. Linear Algebra and its Applications, 181:187–??, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mond:1994:IIP

[MP94a] B. Mond and J. E. Pečarić. Inequalities involving powers of generalized inverses. Linear Algebra and its Applications, 199:293–303, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mond:1994:IWP

[MP94b] B. Mond and J. E. Pečarić. Inequalities with weights for powers of generalized inverses II. Linear Algebra and its Applications, 210:265–272, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mond:1995:RFC

[MP95] B. Mond and J. E. Pečarić. Reverse forms of a convex matrix inequality. Linear Algebra and its Applications, 220(1–3):

359–364, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400300. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Mitra:1996:NSMa

[MP96a] Sujit Kumar Mitra and K. Manjunatha Prasad. The nonunique shorted matrix. Linear Algebra and its Applications, 237/238(1–3):41– 70, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500599. Special

issue honoring Calyampudi Radhakrishna Rao.

Mond:1996:MAM

[MP96b] B. Mond and J. E. Pečarić.

A mixed arithmetic-mean-harmonic-mean matrix inequality. Linear Algebra and its Applications, 237/238(1–3):449–454, April ??, 1996. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500269. Special

issue honoring Calyampudi Radhakrishna Rao.

Marcellan:1997:ETT

[MP97a] F. Marcellán and J. Petronilho. Eigenproblems for tridiagonal 2-Toeplitz matrices and quadratic polynomial mappings. Linear Algebra and its Applications, 260(1–3): 169–208, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=260&issue=1-
                                        3&aid=9600300.

Maroulas:1997:BNR

[MP97b] J. Maroulas and P. Psarrakos. The boundary of the numerical range of matrix polynomials. Linear Algebra and its Applications, 267(1–3):101–111, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=267&issue=1-
                                        3&aid=9700004.

Mathai:1997:SPM

[MP97c] A. M. Mathai and Giorgio Pederzoli. Some properties of matrix-variate Laplace transforms and matrix-variate Whittaker functions. Linear Algebra and its Applications, 253(1–3):209–226, March 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500705.

Mitra:1997:NPS

[MP97d] Sujit Kumar Mitra and K. Manjunatha Prasad. The nonunique parallel sum. Linear Algebra and its Applications, 259(1–3):77–99, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600243.

Maroulas:1998:CNR

[MP98a] J. Maroulas and P. Psarrakos. On the connectedness of numerical range of matrix polynomials. Linear Algebra and its Applications, 280(2–3):97–108, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6110.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=280&issue=
                                        2-3&aid=6110.

Martinez:1998:FCV

[MP98b] J. J. Martínez and J. M. Peña. Factorizations of Cauchy–Vandermonde matrices. Linear Algebra and its Applications, 284(1–3):229– 237, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        284/1-3/6192.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=284&issue=
                                        1-3&aid=6192.

Mathai:1998:WFM

[MP98c] A. M. Mathai and Giorgio Pederzoli. A Whittaker function of matrix argument. Linear Algebra and its Applications, 269(1):91– 103, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9700059.

MorenoGalindo:1998:ANT

[MP98d] A. Moreno Galindo and A. Rodríguez Palacios. Algebra norms on tensor products of algebras, and the norm extension problem. Linear Algebra and its Applications, 269(1):257–305, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700225.

Marchi:1997:QSL

[MPdS97] Ezio Marchi, Rubén Puente, and Virginia N. Vera de Serio. Quasipolyhedral sets in linear semi-infinite inequality systems. Linear Algebra and its Applications, 255(1–3): 157–169, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600164.

Murthy:1997:CCL

[MPS97] G. S. R. Murthy, T. Parthasarathy, and B. Sriparna. Constructive characterization of Lipschitzian Q_{0}-matrices. Linear Algebra and its Applications, 252(1–3):323–337, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=252&issue=1-
                                        3&aid=9600158.

Mond:1997:OVS

[MPŠV97] B. Mond, J. Pečarić, J. Šunde, and S. Varošanec. Operator versions of some classical inequalities. Linear Algebra and its Applications, 264(1–3):117–126, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9700234.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1997&volume=264&
                                        issue=1-3&aid=9700234.

Martins:1995:SIG

[MQ95] Rosa Amélia Martins and João Filipe Queiró. Spectral inequalities for generalized Rayleigh quotients. Linear Algebra and its Applications, 220(1–3):311–320, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=220&issue=1-
                                        3&aid=9400110. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Meng:1994:GCR

[MR94] Xiao-Li Meng and Donald B.

Rubin. On the global and componentwise rates of convergence of the EM algorithm. Linear Algebra and its Applications, 199:413–425, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McCullough:1998:NGS

[MR98a] Scott A. McCullough and Leiba Rodman. Normal generalized selfadjoint operators in Krein spaces. Linear Algebra and its Applications, 283 (1–3):239–245, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6208.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6208.

Meshulam:1998:LSN

[MR98b] Roy Meshulam and Nizar Radwan. On linear subspaces of nilpotent elements in a Lie algebra. Linear Algebra and its Applications, 279(1–3):195–199, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6131.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6131.

Moricz:1998:ERS

[MR98c] Ferenc Móricz and Billy E. Rhoades. An equivalent reformulation of summability by weighted mean methods. Linear Algebra and its Applications, 268(1–3):171– 181, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700039.

Meenakshi:1999:PPS

[MR99] A. R. Meenakshi and C. Rajian. On a product of positive semidefinite matrices. Linear Algebra and its Applications, 295(1–3):3–6, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6363.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6363.

Mrcun:1995:LSP

[Mrč95]  Janez Mrčun.  Lipschitz

spectrum preserving mappings on algebras of matrices. Linear Algebra and its Applications, 215(1–3):113– 120, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300078.

Marek:1991:PPB

[MS91a] Ivo Marek and Daniel B.

Szyld. Pseudoirreducible and pseudoprimitive bounded operators. Linear Algebra and its Applications, 154/156:779– 791, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MarquesdeSa:1991:GSV

[MS91b] Eduardo Marques de Sá and María-José Sodupe. Generalized singular values, interlacing inequalities, and monotonic norms. Linear Algebra and its Applications, 148:179– 192, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcellan:1993:CMO

[MS93a] Francisco Marcellan and Gabriela Sansigre. On a class of matrix orthogonal polynomials on the real line. Linear Algebra and its Applications, 181:97–??, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MarquesdeSa:1993:CMN

[MS93b] Eduardo Marques de Sá and María-José Sodupe. Characterizations of ^{∗} orthantmonotonic norms. Linear Algebra and its Applications, 193:1–9, November 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathias:1993:BAS

[MS93c] R. Mathias and G. W. Stewart. A block QR algorithm and the singular value decomposition. Linear Algebra and its Applications, 182:91– 100, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martins:1997:NIP

[MS97] Enide Andrade Martins and Fernando C. Silva. On the number of invariant polynomials of matrix commutators. Linear Algebra and its Applications, 262(1–3):179–188, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600480.

Markham:1998:SCI

[MS98] Thomas L. Markham and Ronald L. Smith. A Schur complement inequality for certain P -matrices. Linear Algebra and its Applications, 281 (1–3):33–41, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6144.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6144.

Marcoux:1999:CPL

[MS99a] Laurent W. Marcoux and Ahmed R. Sourour. Commutativity preserving linear maps and Lie automorphisms of triangular matrix algebras. Linear Algebra and its Applications, 288(1–3): 89–104, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6292.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6292.

Merino:1999:LAQ

[MS99b] Dennis I. Merino and Vladimir V.

Sergeichuk. Littlewood’s algorithm and quaternion matrices. Linear Algebra and its Applications, 298 (1–3):193–208, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/105/17/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/105/17/29/article.
                                        pdf.

Merikoski:1994:BSV

[MST94] Jorma Kaarlo Merikoski, Humberto Sarria, and Pablo Tarazaga. Bounds for singular values using traces. Linear Algebra and its Applications, 210:227–254, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miranda:1993:TIS

[MT93] H. Miranda and Robert C. Thompson. A trace inequality with a subtracted term. Linear Algebra and its Applications, 185:165–??, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miranda:1994:GMC

[MT94] Hector F. Miranda and Robert C. Thompson. Group majorization, the convex hulls of sets of matrices and the diagonal element–singular value inequalities. Linear Algebra and its Applications, 199:131– 141, March 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miranda:1996:SNT

[MT96] Hector Miranda and Robert C.

Thompson. A supplement to the von Neumann trace inequality for singular values. Linear Algebra and its Applications, 248(1–3):61–66, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500157.

Martins:1996:EBS

[MTS96] M. Madalena Martins, M. Estela Trigo, and M. Madalena Santos. An error bound for the SSOR and USSOR methods. Linear Algebra and its Applications, 232(1–3):131–147, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mubarakzjanov:1999:FSN

[Mub99] R. Mubarakzjanov. Finiteness of a set of non-collinear vectors generated by a family of linear operators. Linear Algebra and its Applications, 294(1–3):9–33, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6379.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6379.

Muhlbach:1990:EDI

[Müh90] G. Mühlbach. On extending determinantal identities. Linear Algebra and its Applications, 132:145–162, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Murota:1995:IBM

[Mur95a] Kazuo Murota. An identity for bipartite matching and symmetric determinant. Linear Algebra and its Applications, 222(1–3):261– 274, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300300.

Murota:1995:IMS

[Mur95b] Kazuo Murota. An identity for matching and skew-symmetric determinant. Linear Algebra and its Applications, 218(1–3): 1–27, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300149.

Mustafa:1994:HMI

[Mus94] Denis Mustafa. How much integral action can a control system tolerate? Linear Algebra and its Applications, 205/206: 965–970, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merikoski:1997:BUB

[MUV^{+}97] Jorma Kaarlo Merikoski, Uoti Urpala, Ari Virtanen, Tin-Yau Tam, and Frank Uhlig. A best upper bound for the 2-Norm condition number of a matrix. Linear Algebra and its Applications, 254(1–3): 355–365, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600474. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Merikoski:1993:SFN

[MV93] Jorma Kaarlo Merikoski and Ari Virtanen. Some further notes on the Marcus-de Oliveira determinantal conjecture. erratum to: “Some notes on de Oliveira’s determinantal conjecture” [Linear Algebraic Appl. 121 (1989), 345–352, MR 90g:15014]. Linear Algebra and its Applications, 187: 259–262, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mallol:1995:PAV

[MV95] C. Mallol and R. Varro. À propos des algèbres vérifiant x^{[3]} = ω(x)^{3}x. (French) [on algebras satisfying x^{[3]} = ω(x)^{3}x]. Linear Algebra and its Applications, 225(1):187– 194, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300339.

Merikoski:1997:BEU

[MV97] Jorma Kaarlo Merikoski and Ari Virtanen. Bounds for eigenvalues using the trace and determinant. Linear Algebra and its Applications, 264(1–3):101–108, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700067.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.

cgi?year=1997&volume=264&
issue=1-3&aid=9700067.

Mackens:1998:PMC

[MV98a] Wolfgang Mackens and Heinrich Voss. A projection method for computing the minimum eigenvalue of a symmetric positive definite Toeplitz matrix. Linear Algebra and its Applications, 275/276(1–3):401–415, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6017.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6017. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Mathias:1998:RPB

[MV98b] Roy Mathias and Krešimir Veselić. A relative perturbation bound for positive definite matrices. Linear Algebra and its Applications, 270(1–3): 315–321, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700248.

Moro:1998:RSP

[MV98c] Julio Moro and José Manuel Vegas. Reduced stability of parameter-dependent matrices. Linear Algebra and its Applications, 268(1–3):289– 321, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=268&issue=1-
                                        3&aid=9700049.

Mira:1999:RQF

[MVKR99] J. A. Cuenca Mira, R. De Los Santos Villodres, A. Kaidi, and A. Rochdi. Real quadratic flexible division algebras. Linear Algebra and its Applications, 290(1–3):1–22, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        290/1-3/6230.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=290&issue=
                                        1-3&aid=6230.

Moonen:1993:AUD

[MVV93] Marc Moonen, Paul Van Dooren, and Filiep Vanpoucke. On the QR algorithm and updating the SV D and the URV decomposition in parallel. Linear Algebra and its Applications, 188/189:

549–568, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Monson:1995:PRP

[MW95] B. Monson and Asia Ivić Weiss. Polytopes related to the Picard group. Linear Algebra and its Applications, 218(1–3):185–204, March 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300170.

Moor:1991:TGO

[MZ91] Bart De Moor and Hong Yuan Zha. A tree of generalizations of the ordinary singular value decomposition. Linear Algebra and its Applications, 147: 469–500, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nabben:1996:NCT

[Nab96] Reinhard Nabben. A note on comparison theorems for splittings and multisplittings of Hermitian positive definite matrices. Linear Algebra and its Applications, 233(1–3):67– 80, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400050.

Nabben:1997:MIM

[Nab97] Reinhard Nabben. Z-matrices and inverse Z-matrices. Linear Algebra and its Applications, 256(1–3):31–48, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781111.

Nabben:1999:TSB

[Nab99] Reinhard Nabben. Two-sided bounds on the inverses of diagonally dominant tridiagonal matrices. Linear Algebra and its Applications, 287 (1–3):289–305, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6257.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6257.

Naevdal:1997:GTM

[Næv97] Geir Nævdal. On a generalization of the trigonometric moment problem. Linear Algebra and its Applications, 258(1– 3):1–18, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600151.

Nagyova:1995:CNM

[Nag95] Zuzana Nagyová. On classes of normalized matrices. Linear Algebra and its Applications, 219(1–3):111–137, April 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300202.

Nakamoto:1997:SGO

[Nak97] Ritsuo Nakamoto. The spherical gap of operators. Linear Algebra and its Applications, 251(1–3):89–95, January 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500697.

Narayanan:1991:PLP

[Nar91] H. Narayanan. The principal lattice of partitions of a submodular function. Linear Algebra and its Applications, 144:179–216, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Narayanan:1995:RTP

[Nar95] H. Narayanan. A rounding technique for the polymatroid membership problem. Linear Algebra and its Applications, 221(1–3):41– 57, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300222.

Nash:1992:RFM

[Nas92] Stephen G. Nash. Review of Fundamentals of Matrix Computations by David S. Watkins. Linear Algebra and its Applications, 171:275–??, July 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:1998:GRI

[nC98] Gong ning Chen. The general rational interpolation problem and its connection with the Nevanlinna–Pick interpolation and power moment problem. Linear Algebra and its Applications, 273(1–3):83– 117, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700346.

Chen:1998:MNP

[nCjH98a] Gong ning Chen and Yong jian Hu. On the multiple Nevanlinna–Pick matrix interpolation in the class C_{p} and the Carathéodory matrix coefficient problem. Linear Algebra and its Applications, 283 (1–3):179–203, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6214.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6214.

Chen:1998:THM

[nCjH98b] Gong ning Chen and Yong jian Hu. The truncated Hamburger matrix moment problems in the nondegenerate and degenerate cases, and matrix continued fractions. Linear Algebra and its Applications, 277(1–3):199–236, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6076.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6076.

Chen:1996:GRI

[nCZpZ96] Gong ning Chen, Bin Zhao, and Hui pin Zhang. The general rational interpolation problem in the scalar case and its Hankel vector. Linear Algebra and its Applications, 244(1–3):165–201, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=244&issue=1-
                                        3&aid=9400223.

Nagarajan:1999:PTT

[NDS99] K. R. Nagarajan, M. Paul Devasahayam, and T. Soundararajan. Products of three triangular matrices. Linear Algebra and its Applications, 292 (1–3):61–71, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        292/1-3/6354.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=292&issue=
                                        1-3&aid=6354.

Neuman:1990:FSP

[Neu90] František Neuman. Finite sums of products of functions in single variables. Linear Algebra and its Applications, 134:153–164, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neumaier:1995:HNB

[Neu95] Arnold Neumaier. Hybrid norms and bounds for overdetermined linear systems. Linear Algebra and its Applications, 216(1–3):257–265, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300152.

Neudecker:1996:AVM

[Neu96a] Heinz Neudecker. The asymptotic variance matrices of the sample correlation matrix in elliptical and normal situations and their proportionality. Linear Algebra and its Applications, 237/238(1–3):127– 132, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500351. Special

issue honoring Calyampudi Radhakrishna Rao.

Neufeld:1996:DBE

[Neu96b] Stewart W. Neufeld. A diameter bound on the exponent of a primitive directed graph. Linear Algebra and its Applications, 245(1–3):27– 47, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=245&issue=1-
                                        3&aid=9400203.

Neubauer:1997:IPD

[Neu97a] Michael G. Neubauer. An inequality for positive definite matrices with applications to combinatorial matrices. Linear Algebra and its Applications, 267(1–3):163– 174, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=267&issue=1-
                                        3&aid=9700006.

Neumaier:1997:SSC

[Neu97b] Arnold Neumaier. Scaling and structural condition numbers. Linear Algebra and its Applications, 263(1–3):157–165, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600531.

Neumann:1998:CCH

[Neu98] M. Neumann. A conjecture concerning the Hadamard product of inverses of Mmatrices. Linear Algebra and its Applications, 285(1–3): 277–290, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6266.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6266.

Newman:1994:TM

[New94] Morris Newman. Tridiagonal matrices. Linear Algebra and its Applications, 201: 51–55, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Newman:1997:SNF

[New97] Morris Newman. The Smith normal form. Linear Algebra and its Applications, 254(1–3): 367–381, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1997&volume=254&issue=1-
                                        3&aid=9600163. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Nordstrom:1995:ECD

[NF95] K. Nordström and J. Fellman. Erratum: “Characterizations and dispersion-matrix robustness of efficiently estimable parametric functionals in linear models with nuisance parameters” [Linear Algebra Appl. 127 (1990), 341– 361, MR 91d:62098]. Linear Algebra and its Applications, 220(1–3):4–5, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ng:1997:BPB

[Ng97a]  Michael K. Ng.  Band

preconditioners for block-Toeplitz–Toeplitz-block systems. Linear Algebra and its Applications, 259(1–3):307– 327, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600295.

Ng:1997:FBE

[Ng97b] Michael K. Ng. FFT-Based exponentially weighted recursive least squares computations. Linear Algebra and its Applications, 263(1–3):167– 191, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600532.

Nieuwenhuis:1990:SCN

[Nie90] J. W. Nieuwenhuis. From simple to complicated: noncausal bounded input, bounded output stability in linear discrete time models in the deterministic and stochastic cases. Linear Algebra and its Applications, 141:153–164, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niederreiter:1993:FPS

[Nie93a] Harald Niederreiter. Factorization of polynomials and some linear-algebra problems over finite fields. Linear Algebra and its Applications, 192: 301–328, October 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Nieuwenhuis:1993:RSL

[Nie93b] J. W. Nieuwenhuis. Restricted sums of linear systems and transfer-like sums. Linear Algebra and its Applications, 178:237–??, January 1, 1993.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nieuwenhuis:1997:ALM

[Nie97a] J. W. Nieuwenhuis. About limit matrices of finite-state Markov chains. Linear Algebra and its Applications, 250(1–3):219–223, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500524.

Nievergelt:1997:SMM

[Nie97b] Yves Nievergelt. Schmidt-Mirsky matrix approximation with linearly constrained singular values. Linear Algebra and its Applications, 261(1–3):207–219, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600403.

Nieuwenhuis:1998:WCL

[Nie98a] J. W. Nieuwenhuis. When to call a linear system nonnegative. Linear Algebra and its Applications, 281(1– 3):43–58, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6154.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6154.

Niezgoda:1998:ACE

[Nie98b] M. Niezgoda. An analytical characterization of effective and of irreducible groups inducing cone orderings. Linear Algebra and its Applications, 269(1):105– 114, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700060.

Niezgoda:1998:GMS

[Nie98c] M. Niezgoda. Group majorization and Schur type inequalities. Linear Algebra and its Applications, 268(1–3):9– 30, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9789322.

Niethammer:1999:NEC

[Nie99a]  Wilhelm Niethammer.  A

note on the extended convergence of SOR for two-periodic Markov chains. Linear Algebra and its Applications, 287 (1–3):315–322, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6221.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6221.

Niezgoda:1999:MIL

[Nie99b] M. Niezgoda. G-majorization inequalities for linear maps. Linear Algebra and its Applications, 292(1–3):207–231, May 01, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1999/292/1-3/6382.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1999&volume=292&
                                        issue=1-3&aid=6382.

Neudecker:1996:DMP

[NL96] Heinz Neudecker and Shuangzhe Liu. The density of the Moore–Penrose inverse of a random matrix. Linear Algebra and its Applications, 237/238(1–3):123– 126, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9400240. Special

issue honoring Calyampudi Radhakrishna Rao.

Muoneke:1994:DSP

[nM94] n’Ekwunife Muoneke. Doubly stochastic powers of nonnegative irreducible matrices. Linear Algebra and its Applications, 212/213:259–266, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Nordstrom:1996:ISM

[NM96] Kenneth Nordström and Thomas Mathew. Inequalities for sums of matrix quadratic forms. Linear Algebra and its Applications, 237/238(1–3): 429–447, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500356. Special

issue honoring Calyampudi Radhakrishna Rao.

Nanda:1990:MPI

[NO90] H. Nanda and I. Olkin. Matrix properties of an interbattery factor analytic model. Linear Algebra and its Applications, 127:617–630, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:1995:DTP

[NO95] M. Niezgoda and Z. Otachel. Differentiable transformations preserving a cone preordering. Linear Algebra and its Applications, 221(1–3):59– 67, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=221&issue=1-
                                        3&aid=9300229.

Nakazi:1998:CM

[NO98] Takahiko Nakazi and Kazuyoshi Okubo. ρ-contraction and 2 × 2 matrix. Linear Algebra and its Applications, 283 (1–3):165–169, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        283/1-3/6216.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=283&issue=
                                        1-3&aid=6216.

Norman:1997:JFF

[Nor97] C. W. Norman. On the Jordan form of a family of linear mappings. Linear Algebra and its Applications, 257(1–3): 225–241, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600155.

Notay:1991:CAM

[Not91] Yvan Notay. Conditioning analysis of modified block incomplete factorizations. Linear Algebra and its Applications, 154/156:711–722, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

November:1997:RLT

[Nov97] Laurence J. November. Rotation and Lorentz transformations in 2 × 2 and 4 × 4 complex matrices and in polarizedlight physics. Linear Algebra and its Applications, 254(1–3): 383–408, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600515. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Nowicki:1996:NRF

[Now96] Andrzej Nowicki. On the nonexistence of rational first integrals for systems of linear differential equations. Linear Algebra and its Applications, 235(1–3):107–120, March 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400122.

Nagy:1992:IFA

[NP92a] James Nagy and Robert J. Plemmons. An inverse factorization algorithm for linear prediction. Linear Algebra and its Applications, 172:169–??, July 15, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neal:1992:GAG

[NP92b] Larry Neal and George Poole. A geometric analysis of Gaussian elimination. II. Linear Algebra and its Applications, 173:239–264, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nurhonen:1992:EDO

[NP92c] Markku Nurhonen and Simo Puntanen. Effect of deleting an observation on the equality of the OLSE and BLUE.

Linear Algebra and its Applications, 176:131–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neven:1993:CRP

[NP93] W. H. L. Neven and C. Praagman. Column reduction of polynomial matrices. Linear Algebra and its Applications, 188/189:569–589, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nutov:1996:NEP

[NP96] Zeev Nutov and Michal Penn.

On non-\{0, \frac{1}{2}, 1\} extreme points of the generalized transitive tournament polytope. Linear Algebra and its Applications, 233(1–3):149–159, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400063.

Neubauer:1997:MDM

[NR97] M. G. Neubauer and A. J.

Radcliffe. The maximum determinant of ±1 matrices. Linear Algebra and its Applications, 257(1–3):289– 306, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600147.

Nemirovski:1999:CMS

[NR99] Arkadi Nemirovski and Uriel Rothblum. On complexity of matrix scaling. Linear Algebra and its Applications, 302-303(1–3):435–460, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
51/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/51/article.
pdf.

Nieuwenhuis:1997:SMS

[NS97] Herman J. Nieuwenhuis and Lambert Schoonbeek. Stability of matrices with sufficiently strong negativedominant-diagonal submatrices. Linear Algebra and its Applications, 258(1–3):195– 217, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600193.

Neto:1999:RRP

[NS99a] O. Neto and F. C. Silva. On the recognition and rigidity problems for sums of matrices.

Linear Algebra and its Applications, 293(1–3):73–84, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6368.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6368.

Neumann:1999:CGP

[NS99b] Michael Neumann and Hans Schneider. The convergence of general products of matrices and the weak ergodicity of Markov chains. Linear Algebra and its Applications, 287(1–3):307–314, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6306.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6306.

Newman:1991:MRA

[NT91a] Morris Newman and Robert C.

Thompson. Matrices over rings of algebraic integers. Linear Algebra and its Applications, 145:1–20, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nylen:1991:NRD

[NT91b] Peter Nylen and Tin Yau Tam. Numerical range of a doubly stochastic matrix. Linear Algebra and its Applications, 153:161–176, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Newman:1994:CCG

[NT94] Morris Newman and Robert C. Thompson. A counterexample connected with Geršgorin’s theorem. Linear Algebra and its Applications, 201:165–169, April 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nylen:1997:IEP

[NU97] Peter Nylen and Frank Uhlig. Inverse eigenvalue problems associated with springmass systems. Linear Algebra and its Applications, 254(1–3): 409–425, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9600316. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Nabben:1995:GUM

[NV95a] Reinhard Nabben and Richard S. Varga. Generalized ultrametric matrices — A class of inverse M-matrices. Linear Algebra and its Applications, 220(1–3):365–390, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9400086. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Nabben:1995:CIM

[NV95b] Reinhard Nabben and Richard S.

Varga. On classes of inverse Z-matrices. Linear Algebra and its Applications, 223/224(1–3):521– 552, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9592718. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Neudecker:1990:AVM

[NW90] H. Neudecker and A. M. Wesselman. The asymptotic variance matrix of the sample correlation matrix. Linear Algebra and its Applications, 127: 589–599, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neumann:1991:NGI

[NW91] M. Neumann and H. J. Werner. Nonnegative group inverses. Linear Algebra and its Applications, 151:85–96, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Naevdal:1992:PMC

[NW92] Geir Naevdal and Hugo J. Woerdeman. Partial matrix contractions and intersections of matrix balls. Linear Algebra and its Applications, 175: 225–??, October 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neubauer:1998:NOD

[NWZ98a] Michael G. Neubauer, William Watkins, and Joel Zeitlin. Notes on D-optimal designs. Linear Algebra and its Applications, 280(2–3):109–127, September 4, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        280/2-3/6136.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=280&issue=
                                        2-3&aid=6136.

Neubauer:1998:CNR

[NWZ98b] Michael G. Neubauer, William Watkins, and Joel Zeitlin. On the connectedness of numerical range of matrix polynomials. Linear Algebra and its Applications, 280(2–3):97– ??, September 4, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Najman:1991:MCE

[NY91] B. Najman and Q. Ye. A minimax characterization for eigenvalues of Hermitian pencils. Linear Algebra and its Applications, 144:217–230, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Najman:1993:MCE

[NY93] B. Najman and Q. Ye. A minimax characterization for eigenvalues of Hermitian pencils. II. Linear Algebra and its Applications, 191:183–197, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nylen:1996:MRM

[Nyl96] Peter M. Nylen. Minimumrank matrices with prescribed graph. Linear Algebra and its Applications, 248(1–3):303– 316, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=248&issue=1-
3&aid=9500238.

Nylen:1998:NSS

[Nyl98] Peter Nylen. Null space structure of tree-patterned matrices. Linear Algebra and its Applications, 279(1–3): 153–161, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6107.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6107.

Nylen:1999:IEP

[Nyl99] Peter Nylen. Inverse eigenvalue problem: existence of special mass-damper-spring systems. Linear Algebra and its Applications, 297 (1–3):107–132, August 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/104/17/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/104/17/24/article.
pdf.

Ozguler:1991:CSP

[ÖA91] A. Bülent Özgüler and Nail Akar. A common solution to a pair of linear matrix equations over a principal ideal domain. Linear Algebra and its Applications, 144:85–99, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oara:1996:SSR

[Oar96] Cristian Oară. Stabilizing solution to the reverse discretetime Riccati equation: A matrix-pencil-based approach. Linear Algebra and its Applications, 246(1–3):113–130, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400341.

Obert:1991:ABT

[Obe91] David G. Obert. The angle between two cones. Linear Algebra and its Applications, 144: 63–70, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ogasawara:1999:NEC

[Oga99] Hideho Ogasawara. A note on the equivalence of a class of factorized Broyden families for nonlinear least squares problems. Linear Algebra and its Applications, 297(1–3):183–191, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/28/article.
                                        pdf.

Olson:1996:DM

[OH96] Curtis J. Olson and Richard D. Hill. On dissipative matrices. Linear Algebra and its Applications, 249(1–3):93– 109, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=249&issue=1-
                                        3&aid=9500348.

Ohsmann:1995:FTT

[Ohs95] Martin Ohsmann. Fast transforms of Toeplitz matrices. Linear Algebra and its Applications, 231(1–3):181–192, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=231&issue=1-
                                        3&aid=9400043.

Dokovic:1990:CPR

[– Dok90] Dragomir Ž. – Doković. Cyclic polygons, roots of polynomials with decreasing nonnegative coefficients, and eigenvalues of stochastic matrices. Linear Algebra and its Applications, 142:173–193, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:1993:FHM

[– Dok93] Dragomir Ž. – Doković. Factorization of Hermitian matrix polynomials with constant signature. Linear Algebra and its Applications, 194: 85–90, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:1995:CCG

[– Dok95] Dragomir Ž. – Doković. Cosets of centralizers in the general linear group contain regular elements. Linear Algebra and its Applications, 220(1–3): 111–121, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9300311. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Ortner:1996:LBD

[OK96] Balder Ortner and Arnold R.

Kräuter. Lower bounds for the determinant and the trace of a class of Hermitian matrices. Linear Algebra and its Applications, 236(1–3):147– 180, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400148.

Okninski:1997:TSM

[Okn97] Jan Okniński. Triangularizable semigroups of matrices. Linear Algebra and its Applications, 262(1–3):111–118, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600476.

OLeary:1990:BSP

[O’L90]  Dianne P. O’Leary.  On

bounds for scaled projections and pseudo-inverses. Linear Algebra and its Applications, 132:115–117, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [Ste89].

OLeary:1991:YAP

[O’L91]  Dianne P. O’Leary.  Yet

another polynomial preconditioner for the conjugate gradient algorithm. Linear Algebra and its Applications, 154/156:

377–388, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olkin:1997:DPC

[Olk97] Ingram Olkin. A determinantal proof of the Craig-Sakamoto theorem. Linear Algebra and its Applications, 264(1–3):217–223, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600484.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600484.

Olshevsky:1998:ECA

[Ols98]  Vadim Olshevsky.  Eigen-

vector computation for almost unitary Hessenberg matrices and inversion of Szegő– Vandermonde matrices via discrete transmission lines. Linear Algebra and its Applications, 285(1–3):37–67, December 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6218.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6218.

Omladic:1990:OPN

[Oml90] Matjaž Omladič. On operators preserving the numerical range. Linear Algebra and its Applications, 134:31– 51, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

ONeill:1993:MBR

[O’N93]  William D. O’Neill.  M-

matrices and bounds for reliable transmission of information in communication systems and economic markets. Linear Algebra and its Applications, 195:145–154, December 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olschowka:1996:NPS

[ON96] Markus Olschowka and Arnold Neumaier. A new pivoting strategy for Gaussian elimination. Linear Algebra and its Applications, 240(1–3):131– 151, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=240&issue=1-
                                        3&aid=9400192.

Ong:1991:EIS

[Ong91] Sing-Cheong Ong. On equality of invariant subspace lattices of operators. Linear Algebra and its Applications, 144:23– 27, 1991. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Olech:1991:AML

[OPR91] C. Olech, T. Parthasarathy, and G. Ravindran. Almost N-matrices and linear complementarity. Linear Algebra and its Applications, 145:107– 125, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

OLeary:1998:WCR

[OPSS98] D. O’Leary, R. Plemmons, H. Schneider, and P. N. Shivakumar. Winnipeg 1997 Conference report. Linear Algebra and its Applications, 284(1–3): 357–??, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omladic:1997:ISM

[OR97] Matjaž Omladič and Heydar Radjavi. Irreducible semigroups with multiplicative spectral radius. Linear Algebra and its Applications, 251(1–3):59–72, January 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500544.

Omladic:1994:SMT

[ORR94] M. Omladič, M. Radjabalipour, and H. Radjavi. On semigroups of matrices with traces in a subfield. Linear Algebra and its Applications, 208/209:419–424, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

OLeary:1994:SIP

[ORS94] Dianne O. O’Leary, Leiba Rodman, and Helene Shapiro, editors. Special issue: proceedings of the 3rd ILAS Conference at the University of West Florida, March 17–20, 1993, Pensacola, Florida, volume 212/213 of Linear Algebra and its Applications. North-Holland Publishing Co., Amsterdam, The Netherlands, 1994. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olsder:1999:EAC

[ORvE99] Geert-Jan Olsder, Kees Roos, and Robert-Jan van Egmond. An efficient algorithm for critical circuits and finite eigenvectors in the max-plus algebra. Linear Algebra and its Applications, 295(1– 3):231–240, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6465.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6465.

Omladic:1991:SPA

[OŠ91] Matjaž Omladič and Peter Šemrl. Spectrum-preserving additive maps. Linear Algebra and its Applications, 153: 67–72, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omladic:1993:AMP

[OŠ93] Matjaž Omladič and Peter Šemrl. Additive mappings preserving operators of rank one. Linear Algebra and its Applications, 182:239– 256, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olevskii:1994:ESM

[OS94] V. Olevskii and M. Solomyak.

An estimate for Schur multipliers in S_{p}-classes. Linear Algebra and its Applications, 208/209:57–64, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omladic:1996:MSB

[OŠ96a] Matjaž Omladič and Peter Šemrl. Matrix spaces with bounded number of eigenvalues. Linear Algebra and its Applications, 249(1–3):29–46, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500253.

ORyan:1996:TFH

[OS96b] Manuel O’Ryan and Daniel B. Shapiro. On trace forms of higher degree. Linear Algebra and its Applications, 246(1–3):313–333, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400363.

Omladic:1998:PD

[OŠ98] Matjaž Omladič and Peter Šemrl. Preserving diagonalisability. Linear Algebra and its Applications, 285 (1–3):165–179, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6228.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6228.

Osaka:1991:IPM

[Osa91]  H. Osaka.  Indecompos-

able positive maps in lowdimensional matrix algebras. Linear Algebra and its Applications, 153:73–83, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Osaka:1993:SAI

[Osa93] Hiroyuki Osaka. A series of absolutely indecomposable positive maps in matrix algebras. Linear Algebra and its Applications, 186:45–53, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ohara:1996:DDG

[OSiA96] Atsumi Ohara, Nobuhide Suda, and Shun ichi Amari. Dualistic differential geometry of positive definite matrices and its applications to related problems. Linear Algebra and its Applications, 247(1–3):31– 53, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9400348.

Ozbay:1993:MSO

[OST93] Hitay Ozbay, Malcolm C.

Smith, and Allen Tannenbaum. Mixed-sensitivity optimization for a class of unstable infinite-dimensional systems. Linear Algebra and its Applications, 178:43–??, January 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ostrovskii:1999:SPR

[Ost99] M. I. Ostrovskii. A short proof of the result on actions that characterize l_{∞}^{n}. Linear Algebra and its Applications, 294(1–3):193–195, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        294/1-3/6425.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=294&issue=
                                        1-3&aid=6425.

Otachel:1997:SOI

[Ota97]  Z. Otachel.  Spectral or-

ders and isotone functionals. Linear Algebra and its Applications, 252(1–3):159–172, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=252&issue=1-
                                        3&aid=9500596.

Otero:1990:ERM

[Ote90] Daniel E. Otero. Extraction of m th roots in matrix rings over fields. Linear Algebra and its Applications, 128:1– 26, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Otto:1993:MCC

[Ott93] James S. Otto. Multigrid convergence for convectiondiffusion problems on composite grids. Linear Algebra and its Applications, 190:39– 70, September 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olesky:1993:QCU

[OTV93] D. D. Olesky, Michael Tsatsomeros, and P. Van Den Driessche. Qualitative controllability and uncontrollability by a single entry. Linear Algebra and its Applications, 187: 183–194, July 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ouattara:1991:AJF

[Oua91a] Moussa Ouattara. Sur les T - algèbres de Jordan. (French) [On Jordan T -algebras]. Linear Algebra and its Applications, 144:11–21, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ouattara:1991:ABO

[Oua91b] Moussa Ouattara. Sur les algèbres de Bernstein d’ordre 2. (French) [on second-order Bernstein algebras]. Linear Algebra and its Applications, 144:29–38, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ouattara:1991:ABQ

[Oua91c] Moussa Ouattara. Sur les algèbres de Bernstein qui sont des T -algèbres. (French) [Bernstein algebras that are T -algebras]. Linear Algebra and its Applications, 148:171– 178, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ouattara:1996:CAP

[Oua96] Moussa Ouattara. Sur une classe d’algèbres à puissances associatives. (French) [on a class of power-associative algebras]. Linear Algebra and its Applications, 235(1–3):47– 62, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400113.

OLeary:1994:LAB

[OY94] Dianne P. O’Leary and A. Yeremin. The linear algebra of block quasi-Newton algorithms. Linear Algebra and its Applications, 212/213: 153–168, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Paardekooper:1991:QCP

[Paa91] M. H. C. Paardekooper. A quadratically convergent parallel Jacobi process for diagonally dominant matrices with nondistinct eigenvalues. Linear Algebra and its Applications, 145:71–87, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Palacios:1990:BEH

[Pal90] José Luis Palacios. Bounds on expected hitting times for a random walk on a connected graph. Linear Algebra and its Applications, 141:241–252, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:1993:PAP

[Pan93] C.-T. Pan. A perturbation analysis of the problem of downdating a Cholesky factorization. Linear Algebra and its Applications, 183:103–??, April 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:1999:PSM

[Pan99a] Ping-Qi Pan. A projective simplex method for linear programming. Linear Algebra and its Applications, 292 (1–3):99–125, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6359.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=292&issue=
                                        1-3&aid=6359.

Panttaja:1999:CPR

[Pan99b] Melody K. Panttaja. Classification of path-recursive graphs. Linear Algebra and its Applications, 295(1–3):91– 95, July 01, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6408.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6408.

Papantoniou:1990:CSS

[Pap90] B. J. Papantoniou. A characterization of the symmetric space SU(n)/SO(n) by geodesic spheres. Linear Algebra and its Applications, 136: 133–164, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Paprzycki:1992:CGE

[Pap92] Marcin Paprzycki. Comparisons of Gaussian elimination algorithms on a Cray Y-MP. Linear Algebra and its Applications, 172:57–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parring:1992:ACM

[Par92] A.-M Parring. About the concept of the matrix derivative. Linear Algebra and its Applications, 176:223–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Park:1996:GBM

[Par96] Hong Goo Park. p-groups in the Betti-Mathieu group. Linear Algebra and its Applications, 234(1–3):125–135, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400091.

Parker:1998:SCO

[Par98a] D. Stott Parker. Schur complements obey Lambek’s categorial grammar: Another view of Gaussian elimination and LU decomposition. Linear Algebra and its Applications, 278(1–3):63–84, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6037.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6037.

Parlett:1998:SSP

[Par98b] Beresford N. Parlett. Spectral sensitivity of products of bidiagonals. Linear Algebra and its Applications, 275/276(1–3):417–431, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        275-276/1-3/6055.pdf;
                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        laa/cas_sub/browse/browse.
                                        cgi?year=1998&volume=275-
                                        276&issue=1-3&aid=6055. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Patel:1993:CES

[Pat93] R. V. Patel. On computing the eigenvalues of a symplectic pencil. Linear Algebra and its Applications, 188/189:591– 611, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:1994:III

[Pat94] Thomas H. Pate. Inequalities involving immanants. Linear Algebra and its Applications, 212/213:31–44, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Pate:1997:EPE

[Pat97a] Thomas H. Pate. Exterior products, elementary symmetric functions, and the Fischer determinant inequality. Linear Algebra and its Applications, 255(1–3):203– 242, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600044.

Pate:1997:MPI

[Pat97b] Thomas H. Pate. A machine for producing inequalities involving immanants and other generalized matrix functions. Linear Algebra and its Applications, 254(1–3):427– 466, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600512. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Patton:1997:SSM

[Pat97c] Linda J. Patton. Some solutions to the matrix equation for three point Nevanlinna– Pick interpolation on the bidisc. Linear Algebra and its Applications, 254(1–3):467– 484, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600442. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Pate:1999:ERC

[Pat99a] Thomas H. Pate. Extreme rays for cones of Hermitian matrix functions that are nonnegative on the positive semidefinite matrices. Linear Algebra and its Applications, 291(1–3):201–225, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6343.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6343.

Pate:1999:TIF

[Pat99b]  Thomas H. Pate.  Ten-

sor inequalities, ξ-functions, and inequalities involving immanants. Linear Algebra and its Applications, 295 (1–3):31–59, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6383.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6383.

Pavani:1999:NDC

[Pav99] Raffaella Pavani. Nearly defective complex eigenvalues in Bessel matrices. Linear Algebra and its Applications, 295(1–3):117–132, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6435.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6435.

Payne:1995:TPA

[Pay95] S. E. Payne. A tensor product action on q-clan generalized quadrangles with q = 2e. Linear Algebra and its Applications, 226/228 (1–3):115–137, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400345.

Prasad:1992:GMP

[PB92] K. Manjunatha Prasad and R. B. Bapat. The generalized Moore–Penrose inverse. Linear Algebra and its Applications, 165:59–??, March 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:1990:UIP

[PC90] B. N. Parlett and H. C. Chen. Use of indefinite pencils for computing damped natural modes. Linear Algebra and its Applications, 140:53– 88, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pasupathy:1992:GTM

[PD92] J. Pasupathy and R. A. Damodar. The Gaussian Toeplitz matrix. Linear Algebra and its Applications, 171: 133–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:1997:FSW

[PD97] Beresford N. Parlett and Inderjit S. Dhillon. Fernando’s solution to Wilkinson’s problem: An application of double factorization. Linear Algebra and its Applications, 267(1–3):247–279, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700022.

Peeters:1995:USR

[Pee95] René Peeters. Uniqueness of strongly regular graphs having minimal p-rank. Linear Algebra and its Applications, 226/228(1–3):9–31, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500184.

Pena:1995:MWI

[Peñ95] J. M. Peña. M-matrices whose inverses are totally positive. Linear Algebra and its Applications, 221(1–3):189– 193, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300244.

Pena:1997:SSV

[Peñ97] J. M. Peña. On the Schur and singular value decompositions of oscillatory matrices. Linear Algebra and its Applications, 261(1–3):307– 315, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=261&issue=1-
                                        3&aid=9600443.

Peris:1991:NCI

[Per91] Josep E. Peris. A new characterization of inverse-positive matrices. Linear Algebra and its Applications, 154/156:45– 58, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Perlman:1996:SME

[Per96] Michael D. Perlman. On stochastic majorization of the eigenvalues of a Wishart matrix. Linear Algebra and its Applications, 237/238(1–3): 405–428, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500594. Special

issue honoring Calyampudi Radhakrishna Rao.

Peretz:1998:LCP

[Per98] Ronen Peretz. Linear conditions for a polynomial P (X, Y) to have younger mates. Linear Algebra and its Applications, 285(1–3): 115–121, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
285/1-3/6238.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=285&issue=
1-3&aid=6238.

Petz:1996:MMM

[Pet96] Dénes Petz. Monotone metrics on matrix spaces. Linear Algebra and its Applications, 244(1–3):81–96, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400211.

Petek:1999:MPS

[Pet99] Tatjana Petek. Mappings preserving spectrum and commutativity on Hermitian matrices. Linear Algebra and its Applications, 290(1–3): 167–191, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6331.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6331.

Phillips:1991:RBS

[Phi91] Dennis Phillips. Resolvent bounds and spectral variation. Linear Algebra and its Applications, 149:35–40, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pierro:1990:ABS

[PI90] Alvaro R. De Pierro and Alfredo N. Iusem. On the asymptotic behavior of some alternate smoothing series expansion iterative methods. Linear Algebra and its Applications, 130:3–24, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Pierro:1991:CSJ

[PI91] Alvaro R. De Pierro and Alfredo N. Iusem. On the convergence of SOR- and JORtype methods for convex linear complementarity problems. Linear Algebra and its Applications, 154/156:601– 614, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pierce:1991:RDG

[Pie91] Donald Pierce. Range decompositions and generalized square roots of positive semidefinite matrices. Linear Algebra and its Applications, 145:201–212, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pierce:1996:SCM

[Pie96]  William H. Pierce.  A

self-correcting matrix iteration for the Moore–Penrose generalized inverse. Linear Algebra and its Applications, 244(1–3):357–363, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400306.

Pierce:1999:CPM

[Pie99] Stephen Pierce. Congruence of polynomial matrices. Linear Algebra and its Applications, 294(1–3):1–8, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6377.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6377.

Pinkus:1998:IPE

[Pin98]  Allan Pinkus.  An inter-

lacing property of eigenvalues of strictly totally positive matrices. Linear Algebra and its Applications, 279(1–3):201–206, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6114.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6114.

Pisier:1997:QFU

[Pis97]  Gilles Pisier.  Quadratic

forms in unitary operators. Linear Algebra and its Applications, 267(1–3):125–137, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9700003.

Pittenger:1997:MLS

[Pit97] Arthur O. Pittenger. Mappings of Latin squares. Linear Algebra and its Applications, 261(1–3):251–268, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600408.

Paz:1995:FPL

[PL95] Azaria Paz and Mody Lempel. Factorization properties of lattices over the integers. Linear Algebra and its Applications, 214(1–3):247– 270, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300069.

Plesken:1995:SI

[Ple95] W. Plesken. Solving XX^{tr} = A over the integers. Linear Algebra and its Applications, 226/228(1–3):331–344, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500156.

Plestenjak:1998:NAC

[Ple98] Bor Plestenjak. A numerical algorithm for computing a basis for the root subspace at a nonderogatory eigenvalue of a multiparameter system. Linear Algebra and its Applications, 285 (1–3):257–276, December 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        285/1-3/6265.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=285&issue=
                                        1-3&aid=6265.

Poole:1991:GAG

[PN91] George Poole and Larry Neal. A geometric analysis of Gaussian elimination. I. Linear Algebra and its Applications, 149:249–272, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Poole:1992:GEW

[PN92] G. Poole and L. Neal. Gaussian elimination: When is scaling beneficial? Linear Algebra and its Applications, 162/164:309–324, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Poljak:1992:MSR

[Pol92] Svatopluk Poljak. Minimum spectral radius of a weighted graph. Linear Algebra and its Applications, 171:53–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pothen:1993:PSS

[Pot93] Alex Pothen. Predicting the structure of sparse orthogonal factors. Linear Algebra and its Applications, 194:183– 203, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pott:1998:CAI

[Pot98] Matthias Pott. On the convergence of asynchronous iteration methods for non-linear paracontractions and consistent linear systems. Linear Algebra and its Applications, 283(1–3):1–33, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6181.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6181.

Picci:1994:AMB

[PP94] Giorgio Picci and Stefano Pinzoni. Acausal models and balanced realizations of stationary processes. Linear Algebra and its Applications, 205/206:997–1043, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pryde:1999:MES

[PP99] A. J. Pryde and R. M. Phatarfod. Multiplicities of eigenvalues of some linear search schemes. Linear Algebra and its Applications, 291(1– 3):115–124, April 15, 1999.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6349.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6349.

Paldi:1991:IRR

[PPG91] Eytan Paldi, Hedi Taub Paldi, and Shaul Gutman. The intersection of regions and root clustering. Linear Algebra and its Applications, 149:151–164, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pecaric:1996:SFM

[PPS96] Josip E. Pečarić, Simo Puntanen, and George P. H. Styan. Some further matrix extensions of the Cauchy–Schwarz and Kantorovich inequalities, with some statistical applications. Linear Algebra and its Applications, 237/238(1–3): 455–476, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500679. Special

issue honoring Calyampudi Radhakrishna Rao.

Peyton:1993:PCG

[PPY93] Barry W. Peyton, Alex Pothen, and Xiaoqing Yuan. Partitioning a chordal graph into transitive subgraphs for parallel sparse triangular solution. Linear Algebra and its Applications, 192:329–353, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Peyton:1995:CTA

[PPY95] Barry W. Peyton, Alex Pothen, and Xiaoqing Yuan. A clique-tree algorithm for partitioning a chordal graph into transitive subgraphs. Linear Algebra and its Applications, 223/224(1–3):553– 588, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9500092. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Parthasarathy:1990:M

[PR90] T. Parthasarathy and G. Ravindran. N-matrices. Linear Algebra and its Applications, 139:89–102, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pierce:1994:LPS

[PR94a] Stephen Pierce and Leiba Rodman. Lipschitz properties of some actions of matrix groups. Linear Algebra and its Applications, 208/209: 135–174, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Przeworska-Rolewicz:1994:RCO

[PR94b] D. Przeworska-Rolewicz. Remarks on commutators with operators having either finite nullity or finite deficiency. Linear Algebra and its Applications, 208/209:199– 207, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Prasad:1996:BRM

[PR96] K. Manjunatha Prasad and K. P. S. Bhaskara Rao. On bordering of regular matrices. Linear Algebra and its Applications, 234(1–3):245– 259, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=234&issue=1-
                                        3&aid=9400116.

Prautzsch:1991:LS

[Pra91] Hartmut Prautzsch. Linear subdivision. Linear Algebra and its Applications, 143:223– 230, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Prasad:1994:GIM

[Pra94] K. Manjunatha Prasad. Generalized inverses of matrices over commutative rings. Linear Algebra and its Applications, 211:35–52, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Prasad:1991:GII

[PRB91] K. Manjunatha Prasad, K. P.

S. Bhaskara Rao, and R. B. Bapat. Generalized inverses over integral domains. II. group inverses and Drazin inverses. Linear Algebra and its Applications, 146:31–47, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pressman:1998:MMS

[Pre98] Irwin S. Pressman. Matrices with multiple symmetry properties: Applications of centrohermitian and perhermitian matrices. Linear Algebra and its Applications, 284 (1–3):239–258, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6256.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1998&volume=284&issue=
                                        1-3&aid=6256.

Pryde:1992:BFT

[Pry92] A. J. Pryde. A Bauer–Fike theorem for the joint spectrum of commuting matrices. Linear Algebra and its Applications, 173:221–??, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Przeslawski:1995:FCS

[Prz95] Krzysztof Przesławski. Faces of convex sets and Minkowski additive selections. Linear Algebra and its Applications, 214(1–3):215–224, January 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=214&issue=1-
                                        3&aid=9300067.

Peters:1994:MCS

[PS94] M. A. Peters and A. A. Stoorvogel. Mixed H_{2}|H_{∞} control in a stochastic framework. Linear Algebra and its Applications, 205/206:971– 996, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puntanen:1996:SFR

[PS96a] Simo Puntanen and Alastair J. Scott. Some further remarks on the singular linear model. Linear Algebra and its Applications, 237/238(1–3): 313–327, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500560. Special

issue honoring Calyampudi Radhakrishna Rao.

Puntanen:1996:BBA

[PS96b] Simo Puntanen and George P. H. Styan. A brief biography and appreciation of Calyampudi Radhakrishna Rao [b. 1920], with a bibliography of his books and papers. Linear Algebra and its Applications, 237/238(1–3):1– 40, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500027. Special

issue honoring Calyampudi Radhakrishna Rao.

Philippe:1997:CFS

[PS97] B. Philippe and M. Sadkane.

Computation of the fundamental singular subspace of a large matrix. Linear Algebra and its Applications, 257(1–3): 77–104, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=257&issue=1-
                                        3&aid=9600040.

Petek:1998:CJH

[PŠ98a] Tatjana Petek and Peter Šemrl. Characterization of Jordan homomorphisms on M_{n} using preserving properties. Linear Algebra and its Applications, 269(1):33– 46, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=269&issue=1&
                                        aid=9700021.

Potts:1998:OTP

[PS98b] Daniel Potts and Gabriele Steidl. Optimal trigonometric preconditioners for nonsymmetric Toeplitz systems. Linear Algebra and its Applications, 281(1–3): 265–292, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        281/1-3/6163.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=281&issue=
                                        1-3&aid=6163.

Parthasarathy:1997:LIM

[PSS97] T. Parthasarathy, R. Sridhar, and B. Sriparna. On Lipschitzian Q_{0} and INS matrices. Linear Algebra and its Applications, 263(1–3):193–199, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600533.

Potts:1998:FSA

[PST98] Daniel Potts, Gabriele Steidl, and Manfred Tasche. Fast and stable algorithms for discrete spherical Fourier transforms. Linear Algebra and its Applications, 275/276(1– 3):433–450, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6022.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6022. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Puhl:1998:MSM

[PSW98] H. Pühl, W. Schirotzek, and M. R. Weber. On matrices satisfying a maximum principle with respect to a cone. Linear Algebra and its Applications, 277(1–3):337–356, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        277/1-3/6093.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6093.

Pesterev:1994:ISM

[PT94a] A. V. Pesterev and G. A. Tavrizov. On inversion of some meromorphic matrices. Linear Algebra and its Applications, 212/213:505–521, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Pordzik:1994:PTE

[PT94b] Pawełr R. Pordzik and Götz Trenkler. Pre-test estimation in the linear regression model with competing restrictions. Linear Algebra and its Applications, 210:123–137, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parker:1995:LCD

[PT95] Christopher Parker and Vladimir D. Tonchev. Linear codes and doubly transitive symmetric designs. Linear Algebra and its Applications, 226/228 (1–3):237–246, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500104.

Price:1999:RTM

[PT99] Geoffrey L. Price and Glenn H.

Truitt. On the ranks of Toeplitz matrices over finite fields. Linear Algebra and its Applications, 294 (1–3):49–66, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6400.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6400.

Ptak:1992:ICM

[Pta92] V. Ptak. Infinite companion matrices. Linear Algebra and its Applications, 166:65– 95, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ptak:1995:DIC

[Ptá95] Vlastimil Pták. Decomposition of the infinite companion and interpolation. Linear Algebra and its Applications, 215(1–3):161–178, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=215&issue=1-
                                        3&aid=9300085.

Ptak:1996:MFS

[Ptá96] Vlastimil Pták. Möbius functions of the shift and extremal operators. Linear Algebra and its Applications, 236(1–3):43– 51, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400119.

Puschmann:1990:WUA

[Pus90] Heinrich Puschmann. Weitere unzulässige Abwandlungen der Vermutung von Wilkinson. (German) [Further invalid modifications of Wilkinson’s conjecture]. Linear Algebra and its Applications, 141: 75–78, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Putinar:1998:MAF

[Put98] Mihai Putinar. Matrix analysis and the Friedrichs operator of a quadrature domain. Linear Algebra and its Applications, 270(1–3):215– 229, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700240.

Ptak:1993:BHL

[PV93] Vlastimil Ptak and Zdenek Vavrin. Bezout, Hankel, and Loewner matrices. Linear Algebra and its Applications, 184:13–??, April 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Papadopoulou:1994:ABN

[PV94] A. A. Papadopoulou and P.- C. G. Vassiliou. Asymptotic behavior of nonhomogeneous Semi-Markov systems. Linear Algebra and its Applications, 210:153–198, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Park:1995:TWB

[PV95] Haesun Park and Sabine Van Huffel. Two-way bidiagonalization scheme for downdating the singular-value decomposition. Linear Algebra and its Applications, 222(1– 3):23–39, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=222&issue=1-
                                        3&aid=9300273.

Picci:1998:PSC

[Pvv98] G. Picci, J. M. van den Hof, and J. H. van Schuppen. Primes in several classes of the positive matrices. Linear Algebra and its Applications, 277(1–3):149–185, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        277/1-3/6068.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6068.

Prince:1990:GPS

[PW90] Jerry L. Prince and Alan S. Willsky. A geometric projectionspace reconstruction algorithm. Linear Algebra and its Applications, 130:151–191, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Paige:1994:HGD

[PW94] C. C. Paige and M. Wei. History and generality of the CS decomposition. Linear Algebra and its Applications, 208/209: 303–326, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pardalos:1991:ASQ

[PYH91] Panos M. Pardalos, Yinyu Ye, and Chi-Geun Han. Algorithms for the solution of quadratic knapsack problems. Linear Algebra and its Applications, 152:69–91, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Patera:1990:CSL

[PZ90a] J. Patera and H. Zassenhaus.

The construction of solvable Lie algebras from equidimensional nilpotent algebras. Linear Algebra and its Applications, 133:89–120, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Patera:1990:SLA

[PZ90b] J. Patera and H. Zassenhaus.

Solvable Lie algebras of dimension ≤ 4 over perfect fields. Linear Algebra and its Applications, 142:1–17, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qin:1996:FPG

[Qin96] Hourong Qin. Fermat’s problem and Goldbach’s problem over M_{n}Z. Linear Algebra and its Applications, 236(1–3): 131–135, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400137.

Quintana-Orti:1998:PCC

[QOQO98] Gregorio Quintana-Ortí and Enrique S. Quintana-Ortí. Parallel codes for computing the numerical rank. Linear Algebra and its Applications, 275/276(1–3):451–470, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6031.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6031. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Queiro:1995:SVI

[QS95] João F. Queiró and Eduardo M. Sá. Singular values and invariant factors of matrix sums and products. Linear Algebra and its Applications, 225(1):43– 56, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9300317.

Queiro:1994:IPS

[Que94] João Filipe Queiró. An inverse problem for singular values and the Jacobian of the elementary symmetric functions. Linear Algebra and its Applications, 197/198:277– 282, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Richardson:1993:MFD

[RAB93] D. W. Richardson, B. D. O.

Anderson, and N. K. Bose. Matrix-fraction description from frequency samples. Linear Algebra and its Applications, 181:251–??, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Radwan:1996:IEP

[Rad96] Nizar Radwan. An inverse eigenvalue problem for symmetric and normal matrices. Linear Algebra and its Applications, 248(1–3):101–109, November 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1996&volume=248&issue=1-
                                        3&aid=9500162.

Rakic:1999:DPO

[Rak99] Zoran Rakić. On duality principle in Osserman manifolds. Linear Algebra and its Applications, 296(1–3):183–189, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/47/17/27/
                                        abstract.html;   http://
                                        www.elsevier.nl/gej-ng/
                                        10/30/19/47/17/27/article.
                                        pdf.

Ralph:1993:NPR

[Ral93] Daniel Ralph. A new proof of Robinson’s homeomorphism theorem for PL-Normal maps. Linear Algebra and its Applications, 178:249–??, January 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rao:1991:GIR

[Rao91] K. P. S. Bhaskara Rao. Generalized inverses with respect to general norms. II. Linear Algebra and its Applications, 148:13–18, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rao:1996:GIR

[Rao96] K. P. S. Bhaskara Rao. Generalized inverses with respect to general norms. III. Linear Algebra and its Applications, 237/238(1–3):109– 113, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500237. Special

issue honoring Calyampudi Radhakrishna Rao.

Rath:1997:DPS

[Rat97]  Werner Rath.  Derivative

and proportional state feedback for linear descriptor systems with variable coefficients. Linear Algebra and its Applications, 260(1–3):273– 310, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600311.

Ravi:1997:GMR

[Rav97] M. S. Ravi. Geometric methods in rational interpolation theory. Linear Algebra and its Applications, 258(1–3):159– 168, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600209.

Razpet:1992:RSO

[Raz92] Marko Razpet. A relative of the shift operator. Linear Algebra and its Applications, 169:1–??, May 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Razafinjatovo:1997:DIS

[Raz97] Haja N. Razafinjatovo. Discrete irregular sampling with larger gaps. Linear Algebra and its Applications, 251(1–3):351–372, January 15, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=251&issue=1-
                                        3&aid=9600610.

Rozsa:1991:BMT

[RBRF91] Pál Rózsa, Roberto Bevilacqua, Francesco Romani, and Paola Favati. On band matrices and their inverses. Linear Algebra and its Applications, 150:287–295, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Reams:1997:GAR

[Rea97] Robert Reams. A Galois approach to m th roots of matrices with rational entries. Linear Algebra and its Applications, 258(1–3):187– 194, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9700191.

Reed:1993:GMT

[Ree93] Irving S. Reed. Generalized de Moivre’s theorem, quaternions, and Lorentz transformations on a Minkowski space. Linear Algebra and its Applications, 191:15–40, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reichel:1991:ALP

[Rei91] Lothar Reichel. The application of Leja points to Richardson iteration and polynomial preconditioning. Linear Algebra and its Applications, 154/156:389–414, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reichstein:1993:AMI

[Rei93] Zinovy Reichstein. On automorphisms of matrix invariants induced from the trace ring. Linear Algebra and its Applications, 193:51–74, November 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reinschke:1994:GTA

[Rei94] Kurt J. Reinschke. Graphtheoretic approach to symbolic analysis of linear descriptor ssytems. Linear Algebra and its Applications, 197/198: 217–244, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Renardy:1996:SVD

[Ren96]  Michael Renardy.  Singu-

lar value decomposition in Minkowski space. Linear Algebra and its Applications, 236(1–3):53–58, March 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400124.

Reznick:1995:SCS

[Rez95] Bruce Reznick. Some constructions of spherical 5-Designs. Linear Algebra and its Applications, 226/228 (1–3):163–196, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500101.

Rubio:1999:CEA

[RG99] P. Rubió and J. Gelonch.

The contragredient equivalence: Application to solve some matrix systems. Linear Algebra and its Applications, 290(1–3):145–166, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6337.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6337.

Rosales:1998:NES

[RGS98] J. C. Rosales and Pedro A.

García-Sánchez. Nonnegative elements of subgroups of Z^{n}. Linear Algebra and its Applications, 270(1–3):351– 357, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700309.

Rosales:1999:NAS

[RGS99] J. C. Rosales and P. A. García-Sánchez. On normal affine semigroups. Linear Algebra and its Applications, 286(1– 3):175–186, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6263.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6263.

Rhodius:1993:EFE

[Rho93] Adolf Rhodius. On explicit forms for ergodicity coefficients. Linear Algebra and its Applications, 194:71–83, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rhodius:1997:MEC

[Rho97] Adolf Rhodius. On the maximum of ergodicity coefficients, the Dobrushin ergodicity coefficients, and products of stochastic matrices. Linear Algebra and its Applications, 253(1–3):141– 154, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500706.

Reams:1999:SRP

[Ri99] Robert Reams and Hadamard inverses. square roots and products of almost semidefinite matrices. Linear Algebra and its Applications, 288(1–3):35–43, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6273.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6273.

Richards:1990:TPK

[Ric90a] Donald St. P. Richards. Totally positive kernels, Pólya frequency functions, and generalized hypergeometric series. Linear Algebra and its Applications, 137/138:467– 478, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Richman:1990:PLT

[Ric90b] Fred Richman. Polynomials and linear transformations. Linear Algebra and its Applications, 131:131–137, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rieck:1997:TIS

[Rie97] M. Q. Rieck. Totally isotropic subspaces, complementary subspaces, and generalized inverses. Linear Algebra and its Applications, 251(1–3):239– 248, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9600554.

Ringel:1998:EMT

[Rin98] Claus Michael Ringel. Exceptional modules are tree modules. Linear Algebra and its Applications, 275/276(1– 3):471–493, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6054.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6054. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Rjasanow:1994:EAC

[Rja94] Sergej Rjasanow. Effective algorithms with circulant-block matrices. Linear Algebra and its Applications, 202:55–69, April 15, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reed:1996:RSP

[RK96] Irving S. Reed and Wolfgang F. Kraske. Reflections, spinors, and projections on a Minkowski space underlie Dirac’s equation. Linear Algebra and its Applications, 239(1–3):227–262, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9500197.

Kim:1995:CG

[rKMMR95] Suh ryung Kim, Terry A.

McKee, F. R. McMorris, and Fred S. Roberts. pcompetition graphs. Linear Algebra and its Applications, 217(1–3):167–178, March 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=217&issue=1-
3&aid=9400060. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Robinson:1991:PDS

[Rob91] Donald W. Robinson. The primary decomposition of a set of matrices. Linear Algebra and its Applications, 147:45– 53, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robinson:1994:MIM

[Rob94] Donald W. Robinson. The Moore idempotents of a matrix. Linear Algebra and its Applications, 211:15–26, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Roberts:1995:PCM

[Rob95] Fred S. Roberts. On the problem of consistent marking of a graph. Linear Algebra and its Applications, 217(1–3): 255–263, March 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=217&issue=1-
                                        3&aid=9400193. Proceedings

of a Conference on Graphs and Matrices in honor of John Maybee (Boulder, CO, 1993).

Robinson:1996:DRI

[Rob96] Donald W. Robinson. The determinantal-rank idempotents of a matrix. Linear Algebra and its Applications, 237/238(1–3):83– 96, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500450. Special

issue honoring Calyampudi Radhakrishna Rao.

Robinson:1998:IAM

[Rob98] Donald W. Robinson. The image of the adjoint mapping. Linear Algebra and its Applications, 277(1–3): 143–148, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6067.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6067.

Roch:1998:NRL

[Roc98]  Steffen Roch.  Numerical

ranges of large Toeplitz matrices. Linear Algebra and its Applications, 282(1–3): 185–198, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6178.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6178.

Roesler:1990:RHE

[Roe90] Friedrich Roesler. Riemann’s hypothesis as an eigenvalue problem. III. Linear Algebra and its Applications, 141:1– 46, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roesler:1997:BES

[Roe97]  Friedrich Roesler.  Bate-

man’s equation and similar units. Linear Algebra and its Applications, 250(1–3):253– 273, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500528.

Rohn:1990:NUD

[Roh90] Jiří Rohn. Nonsingularity under data rounding. Linear Algebra and its Applications, 139:171–174, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rohn:1995:CBS

[Roh95] Jiří Rohn. Checking bounds on solutions of linear interval equations is NP-Hard. Linear Algebra and its Applications, 223/224(1–3):589– 596, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400219. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Rohn:1997:OBS

[Roh97]  Jiří Rohn.  Overestima-

tions in bounding solutions of perturbed linear equations. Linear Algebra and its Applications, 262(1–3):55– 65, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600467.

Rolle:1996:OFM

[Rol96] Jean-Daniel Rolle. Optimization of functions of matrices with an application in statistics. Linear Algebra and its Applications, 234(1–3):261– 275, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400087.

Rost:1993:GCM

[Ros93a] Karla Rost. Generalized companion matrices and matrix representations for generalized Bezoutians. Linear Algebra and its Applications, 193:151– 172, November 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rost:1993:GLE

[Ros93b]  Karla Rost.  Generalized

Lyapunov equations, matrices with displacement structure, and generalized Bezoutians. Linear Algebra and its Applications, 193:75–93, November 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rousseau:1993:ECS

[Rou93] G. Rousseau. An expansion for certain symmetric determinants. Linear Algebra and its Applications, 181:233–??, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rowlinson:1995:SPR

[Row95] Peter Rowlinson. Star partitions and regularity in graphs. Linear Algebra and its Applications, 226/228 (1–3):247–265, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9593232.

Rowlinson:1998:GME

[Row98] Peter Rowlinson. On graphs with multiple eigenvalues. Linear Algebra and its Applications, 283(1–3):75–85, November 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6201.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6201.

Ran:1990:SIL

[RR90a] André C. M. Ran and Leiba Rodman. Stable invariant Lagrangian subspaces: factorization of symmetric rational matrix functions and other applications. Linear Algebra and its Applications, 137/138:575– 620, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Redheffer:1990:SRL

[RR90b] Ray Redheffer and Reinhard Redlinger. The spectral radius and Liapunov’s theorem. Linear Algebra and its Applications, 128:169–180, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roitman:1992:LRN

[RR92a] Moshe Roitman and Zalman Rubinstein. On linear recursions with nonnegative coefficients. Linear Algebra and its Applications, 167:151–??, April 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rozsa:1992:PBT

[RR92b] Pal Rozsa and Francesco Romani. On periodic blocktridiagonal matrices. Linear Algebra and its Applications, 167:35–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ran:1994:RCR

[RR94] André C. M. Ran and Leiba Rodman. The rate of convergence of real invariant subspaces. Linear Algebra and its Applications, 207:197–224, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rabier:1996:CGS

[RR96a] Patrick J. Rabier and Werner C. Rheinboldt. Classical and generalized solutions of timedependent linear differentialalgebraic equations. Linear Algebra and its Applications, 245(1–3):259–293, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=245&issue=1-
3&aid=9400243.

Rabier:1996:TDL

[RR96b] Patrick J. Rabier and Werner C.

Rheinboldt. Time-dependent linear DAEs with discontinuous inputs. Linear Algebra and its Applications, 247(1–3): 1–29, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9400312.

Robenack:1997:GTD

[RR97] Klaus Röbenack and Kurt J.

Reinschke. Graph-theoretically determined Jordan-block-size structure of regular matrix pencils. Linear Algebra and its Applications, 263(1–3):333– 348, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600589.

Robenack:1998:DBD

[RR98] Klaus Röbenack and Kurt J.

Reinschke. Digraph based determination of Jordan block size structure of singular matrix pencils. Linear Algebra and its Applications, 275/276 (1–3):495–507, May 15, 1998.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6028.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6028. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Ratliff:1999:TPI

[RR99a] L. J. Ratliff, Jr. and D. E. Rush. Triangular powers of integers from determinants of binomial coefficient matrices. Linear Algebra and its Applications, 291(1– 3):125–142, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6347.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6347.

Robinson:1999:BAE

[RR99b] Allen C. Robinson and Donald W. Robinson. A best approximation evaluation of a finite element calculation. Linear Algebra and its Applications, 302-303(1–3):367–375, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
47/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/47/article.
pdf.

Ran:1993:DCI

[RRR93] André C. M. Ran, Leiba Rodman, and Jonathan E. Rubin. Direct complements of invariant Lagrangian subspaces and minimal factorizations of skew-symmetric rational matrix functions. Linear Algebra and its Applications, 180: 61–94, February 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roch:1990:ASI

[RS90] S. Roch and B. Silbermann.

On algebras with standard identities. Linear Algebra and its Applications, 137/138:239– 247, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:1992:JFC

[RS92] Lieba Rodman and Tamir Shalom. Jordan forms of completions of partial upper triangular matrices. Linear Algebra and its Applications, 168: 221–??, April 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roggenkamp:1993:ANC

[RS93] K. W. Roggenkamp and L. L. Scott. Automorphisms and NonAbelian cohomology: An algorithm. Linear Algebra and its Applications, 192:355–382, October 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Computational linear algebra in algebraic and related problems (Essen, 1992).

Rothblum:1997:SNS

[RS97] Uriel G. Rothblum and Hans Schneider. A solution of a nonlinear system arising in spectral perturbation theory of nonnegative matrices. Linear Algebra and its Applications, 258(1–3):53– 80, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600154.

Roch:1998:NSV

[RS98] Steffen Roch and Bernd Silbermann. A note on singular values of Cauchy–Toeplitz matrices. Linear Algebra and its Applications, 275/276(1– 3):531–536, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/

275-276/1-3/6012.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6012. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Rahmani:1997:ABG

[RSDT97] A. Rahmani, C. Sueur, and G. Dauphin-Tanguy. Approche des bond graphs pour l’analyse structurelle des systèmes linéaires. (French) [Bond graph approach for the structural analysis of linear systems]. Linear Algebra and its Applications, 259(1–3): 101–131, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600245.

Rojo:1994:DSE

[RSR94] Óscar Rojo, Ricardo L. Soto, and Héctor Rojo. A decreasing sequence of eigenvalue localization regions. Linear Algebra and its Applications, 196: 71–84, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ran:1994:DMI

[RT94] A. C. M. Ran and D. Temme. Dissipative matrices and invariant maximal semidefinite subspaces. Linear Algebra and its Applications, 212/213:169– 214, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Rudolph:1996:DTV

[Rud96] J. Rudolph. Duality in timevarying linear systems: A module theoretic approach. Linear Algebra and its Applications, 245(1–3):83–106, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=245&issue=1-
                                        3&aid=9400222.

Rudolf:1998:MSD

[Rud98] Rüdiger Rudolf. On Monge sequences in d-dimensional arrays. Linear Algebra and its Applications, 268(1–3):59– 70, January 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=268&issue=1-
                                        3&aid=9700007.

Ruhe:1994:RKA

[Ruh94] Axel Ruhe. Rational Krylov algorithms for nonsymmetric eigenvalue problems. II. matrix pairs. Linear Algebra and its Applications, 197/198: 283–295, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon, 1992).

Rump:1991:ESL

[Rum91] S. M. Rump. Estimation of the sensitivity of linear and nonlinear algebraic problems. Linear Algebra and its Applications, 153:1–34, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rump:1997:TPF

[Rum97]  S. M. Rump.  Theorems

of Perron–Frobenius type for matrices without sign restrictions. Linear Algebra and its Applications, 266(1–3):1– 42, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600522.

Rump:1998:SRS

[Rum98a] Siegfried M. Rump. The sign-real spectral radius and cycle products. Linear Algebra and its Applications, 279(1–3):177–180, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6112.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6112.

Rump:1998:SPS

[Rum98b] Siegfried M. Rump. Structured perturbations and symmetric matrices. Linear Algebra and its Applications, 278(1–3):121–132, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6078.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6078.

Russakovskii:1994:TBA

[Rus94] Eugene M. Russakovskii. The theory of V -bezoutians and its applications. Linear Algebra and its Applications, 212/213: 437–460, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Rushanan:1995:ESN

[Rus95] Joseph J. Rushanan. Eigenvalues and the Smith normal form. Linear Algebra and its Applications, 216(1–3):177– 184, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300131.

Rost:1995:RSL

[RV95] Karla Rost and Zdeněk Vavřín. Recursive solution of Löwner-Vandermonde systems of equations II. Linear Algebra and its Applications, 223/224(1–3):597– 617, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400220. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Rost:1996:RSL

[RV96] Karla Rost and Zdeněk Vavřín. Recursive solution of Löwner-Vandermonde systems of equations I. Linear Algebra and its Applications, 233(1–3):51–65, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=233&issue=1-
3&aid=9400040.

Rost:1998:IFF

[RV98] Karla Rost and Zdeněk Vavřín. Inversion formulas and fast algorithms for Löwner-Vandermonde matrices. Linear Algebra and its Applications, 275/276(1– 3):537–549, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6020.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6020. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Raines:1994:CHS

[RW94] Allen C. Raines, III and David S. Watkins. A class of Hamiltonian-symplectic methods for solving the algebraic Riccati equation. Linear Algebra and its Applications, 205/206:1045–1060, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reinschke:1997:DCS

[RW97] Kurt J. Reinschke and Gunter Wiedemann. Digraph characterization of structural controllability for linear descriptor systems. Linear Algebra and its Applications, 266(1–3):199–217, November 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9786521.

Ryan:1990:CGI

[Rya90] Jennifer Ryan. The constraints of the group of an integral monoid. Linear Algebra and its Applications, 139: 285–292, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ryan:1991:DFG

[Rya91] Jennifer Ryan. Decomposing finitely generated integral monoids by elimination. Linear Algebra and its Applications, 153:209–217, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rothblum:1992:SMS

[RZ92] Uriel G. Rothblum and Stavros A. Zenios. Scalings of matrices satisfying lineproduct constraints and generalizations. Linear Algebra and its Applications, 175: 159–??, October 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ran:1997:SAM

[RZ97] André Ran and Peter Zizler. On self-adjoint matrix polynomials with constant signature. Linear Algebra and its Applications, 259(1–3):133– 153, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600246.

Roca:1998:RIF

[RZ98] A. Roca and I. Zaballa. Restricted invariant factor assignment under state-feedback. Linear Algebra and its Applications, 275/276(1–3):509– 529, May 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6030.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6030. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Saad:1992:RAP

[Saa92] Youcef Saad. Review of ABS projection algorithms by J. Abaffy and E. Spedicato. Linear Algebra and its Applications, 165:267–??, March 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sadok:1993:QVE

[Sad93] Hassane Sadok. Quasilinear vector extrapolation methods. Linear Algebra and its Applications, 190:71–85, September 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sakhnovich:1993:ESN

[Sak93] Alexander Sakhnovich. Exact solutions of nonlinear equations and the method of operator identities. Linear Algebra and its Applications, 182: 109–126, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salam:1998:WVH

[Sal98] A. Salam. What is a vector Hankel determinant. Linear Algebra and its Applications, 278(1–3):147–161, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6091.pdf;   http:
//www.elsevier.com/cgi-

bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6091.

Salas:1999:GTM

[Sal99]  Héctor N. Salas.  Gersh-

gorin’s theorem for matrices of operators. Linear Algebra and its Applications, 291 (1–3):15–36, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6325.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6325.

Santos:1991:AJJ

[San91] Newton R. Santos. Applying the JII, JII-SI, and CG methods to two classes of linear problems. Linear Algebra and its Applications, 158: 99–127, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sang:1996:RED

[San96a] Dong Kim Sang. A remark on the eigenvalue distribution of the preconditioning collocation scheme with P_{3} finite element stiffness matrix. Linear Algebra and its Applications, 236(1–3):35– 42, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400117.

Santos:1996:ERT

[San96b] Reginaldo J. Santos. Equivalence of regularization and truncated iteration for general ill-posed problems. Linear Algebra and its Applications, 236(1–3):25–33, March 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400114.

Sauber:1990:WPE

[Sau90] Beatriz I. Sauber. When a polynomial has exactly one positive root and no roots in (−1, 0). Linear Algebra and its Applications, 128:107–115, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singh:1991:CFR

[SAZ91] Surjeet Singh and Hassan Al-Zaid. On a canonical form for recurring sequences. Linear Algebra and its Applications, 157:185–194, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shapiro:1990:TCS

[SB90] A. Shapiro and M. W. Browne. On the treatment of correlation structures as covariance structures. Linear Algebra and its Applications, 127:567–587, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:1997:SGE

[SB97] Jia-Yu Shao and Li Bin. The set of generalized exponents of primitive simple graphs. Linear Algebra and its Applications, 258(1–3):95– 127, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600196.

Su:1998:FRC

[SBKK98] Yangfeng Su, Amit Bhaya, Eugenius Kaszkurewicz, and Victor S. Kozyakin. Further results on convergence of asynchronous linear iterations. Linear Algebra and its Applications, 281(1–3): 11–24, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6151.pdf;   http:
//www.elsevier.com/cgi-

bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6151.

Stewart:1994:PRF

[SC94] Michael Stewart and George Cybenko. Perfect-reconstruction filtering with unitary operators and projections. Linear Algebra and its Applications, 208/209:97–133, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seddik:1998:IND

[SC98] Ameur Seddik and Josette Charles. Sur l’image et le noyau d’une dérivation généralisée. (French) [On the image and kernel of a generalized derivation]. Linear Algebra and its Applications, 274(1–3):77–83, April 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700265.

Scherer:1991:SSA

[Sch91] Carsten Scherer. The solution set of the algebraic Riccati equation and the algebraic Riccati inequality. Linear Algebra and its Applications, 153:99–122, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmitt:1992:PBM

[Sch92] Bernhard A. Schmitt. Perturbation bounds for matrix square roots and Pythagorean sums. Linear Algebra and its Applications, 174:215– 228, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmeisser:1993:RST

[Sch93] Gerhard Schmeisser. A real symmetric tridiagonal matrix with a given characteristic polynomial. Linear Algebra and its Applications, 193:11– 18, November 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schrader:1994:DSP

[Sch94] Cheryl B. Schrader. Dynamical structures on pencils, poles, and fundamental subspaces. Linear Algebra and its Applications, 205/206:1061– 1079, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Scherer:1995:GNA

[Sch95] Carsten W. Scherer. The general nonstrict algebraic Riccati inequality. Linear Algebra and its Applications, 219 (1–3):1–33, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300175.

Schneeweiss:1996:RFR

[Sch96a] Hans Schneeweiss. The reduced form of recursive models: Small sample properties. Linear Algebra and its Applications, 237/238(1–3):277– 300, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500527. Special

issue honoring Calyampudi Radhakrishna Rao.

Scholtes:1996:PBN

[Sch96b] Stefan Scholtes. A proof of the branching number bound for normal manifolds. Linear Algebra and its Applications, 246(1–3):83–95, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400337.

Schmale:1997:SAC

[Sch97] Wiland Schmale. A symmetric approach to cyclization for systems over C[y]. Linear Algebra and its Applications, 263(1–3):221–232, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600537.

Schmale:1998:MCC

[Sch98a] Wiland Schmale. Matrix cyclization over complex polynomials. Linear Algebra and its Applications, 275/276(1– 3):551–562, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6018.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6018. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Schneider:1998:SPR

[Sch98b] Hans Schneider. Some personal reminiscences of Olga Taussky-Todd. Linear Algebra and its Applications, 280(1): 15–19, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Schaffrin:1999:SUE

[Sch99a] Burkhard Schaffrin. Softly unbiased estimation Part 1: The Gauss–Markov model. Linear Algebra and its Applications, 289(1–3):285–296, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6314.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6314.

Schwarz:1999:CCS

[Sch99b] Binyamin Schwarz. A conjecture concerning strongly connected graphs. Linear Algebra and its Applications, 286(1–3):197–208, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6269.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6269.

Scully:1991:MRO

[Scu91] Daniel J. Scully. Maximal rank-one spaces of matrices over chain semirings. I. uspaces. Linear Algebra and its Applications, 150:67–80, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Scully:1993:MRO

[Scu93] Daniel J. Scully. Maximal rank-one spaces of matrices over chain semirings. II (u, i)- spaces. Linear Algebra and its Applications, 181:29–??, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sueur:1997:CIS

[SDT97] C. Sueur and G. Dauphin-Tanguy. Controllability indices for structured systems. Linear Algebra and its Applications, 250(1–3):275–287, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500598.

Searle:1994:ESR

[Sea94] S. R. Searle. Extending some results and proofs for the singular linear model. Linear Algebra and its Applications, 210:139–151, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Searle:1997:MHB

[Sea97] Shayle R. Searle. The matrix handling of BLUE and BLUP in the mixed linear model. Linear Algebra and its Applications, 264 (1–3):291–311, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600400.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600400.

Seeger:1999:EAE

[See99] Alberto Seeger. Eigenvalue analysis of equilibrium processes defined by linear complementarity conditions. Linear Algebra and its Applications, 292(1–3):1–14, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6355.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6355.

Seidel:1996:SEL

[Sei96] J. J. Seidel. A sound equilibrium for linear algebra. Linear Algebra and its Applications, 241/243(1–3):781– 785, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500428. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Stuart:1999:ISP

[SEK99] Jeffrey Stuart, Carolyn Eschenbach, and Steve Kirkland. Irreducible sign k-potent sign pattern matrices. Linear Algebra and its Applications, 294(1–3):85–92, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6423.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6423.

Semal:1991:IAL

[Sem91] Pierre Semal. Iterative algorithms for large stochastic matrices. Linear Algebra and its Applications, 154/156:65– 103, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semrl:1992:RAP

[Sem92] Peter Semrl. On rank additivity and the polynomial equation p(A) = 0. Linear Algebra and its Applications, 176: 21–??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semal:1995:MIM

[Sem95] Pierre Semal. Monotone iterative methods for Markov chains. Linear Algebra and its Applications, 230(1–3):35– 46, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9300357.

Semrl:1996:MVC

[Šem96] Peter Šemrl. On a matrix version of Cochran’s statistical theorem. Linear Algebra and its Applications, 237/238(1–3):477– 487, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500395. Special

issue honoring Calyampudi Radhakrishna Rao.

Seneta:1993:EFE

[Sen93] Eugene Seneta. Explicit forms for ergodicity coefficients of stochastic matrices. Linear Algebra and its Applications, 191:245–252, September 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra:1997:ECG

[Ser97] Stefano Serra. The extension of the concept of the generating function to a class of preconditioned Toeplitz matrices. Linear Algebra and its Applications, 267(1–3):139–161, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=267&issue=1-
                                        3&aid=9782204.

Sergeichuk:1998:UER

[Ser98a] Vladimir V. Sergeichuk. Unitary and Euclidean representations of a quiver. Linear Algebra and its Applications, 278(1–3):37–62, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        278/1-3/6104.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=278&issue=
                                        1-3&aid=6104.

Serra:1998:EEH

[Ser98b] Stefano Serra. On the extreme eigenvalues of Hermitian (block) Toeplitz matrices. Linear Algebra and its Applications, 270(1–3):109– 129, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9700231.

Sznajder:1995:GPE

[SG95] Roman Sznajder and M. Seetharama Gowda. Generalizations of P_{0}- and P-properties, extended vertical and horizontal linear complementarity problems. Linear Algebra and its Applications, 223/224(1–3): 695–715, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9300184. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Simoncini:1996:CPB

[SG96] V. Simoncini and E. Gallopoulos. Convergence properties of block GMRES and matrix polynomials. Linear Algebra and its Applications, 247(1–3):97–119, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500093.

Song:1999:DIM

[SG99] Guangtian Song and Xuejun Guo. Diagonability of idempotent matrices over noncommutative rings. Linear Algebra and its Applications, 297(1–3):1–7, August 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/104/17/
                                        17/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/104/17/17/article.
                                        pdf.

Shen:1999:EI

[SGN99] Jian Shen, David Gregory, and Stewart Neufeld. Exponents of indecomposability. Linear Algebra and its Applications, 288(1–3):229–241, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        288/1-3/6319.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=288&issue=
                                        1-3&aid=6319.

Song:1999:SPP

[SGR99] Yoon Song, M. Seetharama Gowda, and G. Ravindran. On some properties of Pmatrix sets. Linear Algebra and its Applications, 290(1– 3):237–246, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6333.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6333.

Siler:1993:RVS

[SH93] Joseph R. Siler and Richard D.

Hill. Reflectors on vector spaces. Linear Algebra and its Applications, 179:1–6, January 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Siler:1997:MIP

[SH97a] Joseph R. Siler and Richard D.

Hill. Matricial inner products and pointed cones of Hermitian-preserving linear transformations. Linear Algebra and its Applications, 257(1–3):183–191, May ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600205.

Song:1997:SCR

[SH97b] Seok-Zun Song and Suk-Geun Hwang. Spanning column ranks and their preservers of nonnegative matrices. Linear Algebra and its Applications, 254(1–3):485– 495, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=254&issue=1-
                                        3&aid=9600514. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Shao:1998:GEN

[SH98] Jia-Yu Shao and Suk-Geun Hwang. Generalized exponents of non-primitive graphs. Linear Algebra and its Applications, 279(1–3):207–225, August 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        279/1-3/6135.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6135.

Shao:1991:BET

[Sha91a] Jia Yu Shao. Bounds on the k th eigenvalues of trees and forests. Linear Algebra and its Applications, 149:19– 34, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shapiro:1991:SCF

[Sha91b] Helene Shapiro. A survey of canonical forms and invariants for unitary similarity. Linear Algebra and its Applications, 147:101–167, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shader:1992:RCM

[Sha92] Bryan L. Shader. Review of Combinatorial Matrix Theory by Richard A. Brualdi and Herbert J. Ryser. Linear Algebra and its Applications, 173: 273–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shader:1993:CND

[Sha93]  Bryan L. Shader.  Con-

vertible, nearly decomposable, and nearly reducible matrices. Linear Algebra and its Applications, 184:37–53, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:1995:LET

[Sha95]  Jia Yu Shao.  On the

largest k th eigenvalues of trees. Linear Algebra and its Applications, 221(1–3):131– 157, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300236.

Shao:1998:DFC

[Sha98a] Jia-Yu Shao. On digraphs and forbidden configurations of strong sign nonsingular matrices. Linear Algebra and its Applications, 282(1– 3):221–232, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6184.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6184.

Shapira:1998:PAS

[Sha98b] Yair Shapira. Parallelizable approximate solvers for recursions arising in preconditioning. Linear Algebra and its Applications, 274(1–3):211– 237, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700339.

Shapiro:1998:I

[Sha98c] Helene Shapiro. Introduction. Linear Algebra and its Applications, 280(1):5–8, August 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0024379598100022.

Shao:1999:SIL

[Sha99a] Jia-Yu Shao. On sign inconsistent linear systems. Linear Algebra and its Applications, 296(1–3):245–257, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/32/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/32/article.
pdf.

Shapiro:1999:TFP

[Sha99b] Daniel B. Shapiro. Trace form preservers. Linear Algebra and its Applications, 288(1–3):23–33, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6271.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6271.

Shen:1995:PCA

[She95]  Jian Shen.  Proof of a

conjecture about the exponent of primitive matrices. Linear Algebra and its Applications, 216(1–3):185–203, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300132.

Shen:1996:BEP

[She96a] Jian Shen. A bound on the exponent of primitivity in terms of diameter. Linear Algebra and its Applications, 244(1–3):21–33, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400205.

Shen:1996:PEP

[She96b] Jian Shen. A problem on the exponent of primitive digraphs. Linear Algebra and its Applications, 244(1–3):255– 264, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400229.

Shen:1996:SEF

[She96c] Jian Shen. Some estimated formulas for the Frobenius numbers. Linear Algebra and its Applications, 244(1–3):13– 20, September 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=244&issue=1-
3&aid=9400204.

Shi:1995:SLS

[Shi95]  Yixun Shi.  Solving lin-

ear systems involved in constrained optimization. Linear Algebra and its Applications, 229(1–3):175–189, October 1, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9400007.

Shi:1996:PPG

[Shi96] Ronghua Shi. Path polynomials of a graph. Linear Algebra and its Applications, 236(1–3): 181–187, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400149.

Shih:1999:SSS

[Shi99a] Mau-Hsiang Shih. Simultaneous Schur stability. Linear Algebra and its Applications, 287(1–3):323–336, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6190.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=287&issue=
                                        1-3&aid=6190.

Shiromoto:1999:BES

[Shi99b] Keisuke Shiromoto. A basic exact sequence for the Lee and Euclidean weights of linear codes over Z_{l}. Linear Algebra and its Applications, 295(1–3):191–200, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6470.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6470.

Song:1997:MPF

[SHJK97] Seok-Zun Song, Sung Min Hong, Young-Bae Jun, and Seon-Jeong Kim. Minimum permanents on a face of the polytope of doubly stochastic matrices. Linear Algebra and its Applications, 259(1–3): 169–182, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600248.

Shmoish:1994:GSF

[Shm94] Michael Shmoish. On generalized spectral functions, the parametrization of block Hankel and block Jacobi matrices, and some root location problems. Linear Algebra and its Applications, 202:91–128, April 15, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sidi:1992:DIT

[Sid92] Avram Sidi. Development of iterative techniques and extrapolation methods for Drazin inverse solution of consistent or inconsistent singular linear systems. Linear Algebra and its Applications, 167: 171–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sidi:1998:KSM

[Sid98] Avram Sidi. Krylov subspace methods for eigenvalues with special properties and their analysis for normal matrices. Linear Algebra and its Applications, 280 (2–3):129–162, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6150.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6150.

Sidi:1999:UAK

[Sid99] Avram Sidi. A unified approach to Krylov subspace methods for the Drazininverse solution of singular nonsymmetric linear systems. Linear Algebra and its Applications, 298(1–3): 99–113, September 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/105/17/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/105/17/24/article.
pdf.

Silva:1991:MNE

[Sil91a] Fernando C. Silva. Matrices with nonzero entries. Linear Algebra and its Applications, 146:111–119, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:1991:MPE

[Sil91b] Fernando C. Silva. Matrices with prescribed eigenvalues and blocks. Linear Algebra and its Applications, 148: 59–73, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silbermann:1993:LSS

[Sil93a] Bernd Silbermann. On the limiting set of singular values of Toeplitz matrices. Linear Algebra and its Applications, 182:35–??, March 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:1993:MPL

[Sil93b] Fernando C. Silva. Matrices with prescribed lower triangular part. Linear Algebra and its Applications, 182: 27–??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:1994:NIP

[Sil94] Fernando C. Silva. On the number of invariant polynomials of partially prescribed matrices. Linear Algebra and its Applications, 197/198:709– 754, January 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Second Conference of the International Linear Algebra Society (ILAS) (Lisbon,

Silva:1996:PBS

[Sil96] Fernando C. Silva. Possible block similarity classes of a matrix with a prescribed nonprincipal submatrix. Linear Algebra and its Applications, 249(1–3):359– 373, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500394.

Silbermann:1997:AMP

[Sil97a] Bernd Silbermann. Asymptotic Moore-penrose inversion of Toeplitz operators. Linear Algebra and its Applications, 256(1–3):219–234, April 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600043.

Silva:1997:CAJ

[Sil97b] Fernando C. Silva. On a conjecture about the Jordan form of completions of partial upper triangular matrices. Linear Algebra and its Applications, 260(1–3):319– 321, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600314.

Silva:1997:FEC

[Sil97c] Fernando C. Silva. On feedback equivalence and completion problems. Linear Algebra and its Applications, 265(1–3):231–245, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600627.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600627.

Silva:1997:PBS

[Sil97d] Fernando C. Silva. Possible block similarity classes of a matrix with a prescribed nonprincipal submatrix. II. Linear Algebra and its Applications, 262(1–3):259–282, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600518.

Simonic:1992:MGP

[Sim92] Aleksander Simonic. Matrix groups with positive spectra. Linear Algebra and its Applications, 173:57–??, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Simonis:1995:MIC

[Sim95] Juriaan Simonis. MacWilliams identities and coordinate partitions. Linear Algebra and its Applications, 216(1–3):81– 91, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300106.

Simoncini:1996:RPR

[Sim96]  V. Simoncini.  Ritz and

Pseudo-Ritz values using matrix polynomials. Linear Algebra and its Applications, 241/243(1–3):787–801, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500682. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Sinkhorn:1990:CHS

[Sin90] Richard Sinkhorn. Concerning hypernormal stochastic matrices. Linear Algebra and its Applications, 141:275–281, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Summers:1995:EPS

[SJ95] T. Summers and C. R. Johnson. Erratum: “The potentially stable tree sign patterns for dimensions less than five” [Linear Algebra Appl. 126 (1989), 1–13, MR 91f:15043]. Linear Algebra and its Applications, 220(1–3):2–3, April 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shivakumar:1996:ULB

[SJ96] P. N. Shivakumar and Chuanxiang Ji. Upper and lower bounds for inverse elements of finite and infinite tridiagonal matrices. Linear Algebra and its Applications, 247(1–3):297–316, November 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500113.

Shivakumar:1998:NMC

[SJ98] P. N. Shivakumar and Chuanxiang Ji. On the nonsingularity of matrices with certain sign patterns. Linear Algebra and its Applications, 273(1– 3):29–43, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700341.

Sayed:1995:FAG

[SK95] Ali H. Sayed and Thomas Kailath. Fast algorithms for generalized displacement structures and lossless systems. Linear Algebra and its Applications, 219(1–3):49– 78, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300193.

Smith:1997:CPS

[SK97] C. S. Smith and M. Knott. Cycle properties of square matrices. Linear Algebra and its Applications, 258(1–3):345– 350, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600236.

Skarabot:1990:ECM

[Ška90] Jure Škarabot. The extremal case of a matrix inequality. Linear Algebra and its Applications, 134:129–136, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Skula:1998:IMG

[Sku98] Ladislav Skula. Involutions for matrices and generalized inverses. Linear Algebra and its Applications, 271(1–3):283– 308, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700280.

Schmidt:1996:FSI

[SL96] David A. Schmidt and Chjan C. Lim. Full signinvertibility and symplectic matrices. Linear Algebra and its Applications, 232(1–3):97– 110, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Storjohann:1997:FVA

[SL97] Arne Storjohann and George Labahn. A fast Las Vegas algorithm for computing the Smith normal form of a polynomial matrix. Linear Algebra and its Applications, 253(1–3):

155–173, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500743.

Slapnicar:1998:CAD

[Sla98] Ivan Slapničar. Componentwise analysis of direct factorization of real symmetric and Hermitian matrices. Linear Algebra and its Applications, 272(1–3):227– 275, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700334.

Sayed:1993:FTF

[SLK93] Ali H. Sayed, Hanoch Lev-Ari, and Thomas Kailath. Fast triangular factorization of the sum of Quasi-Toeplitz and Quasi-Hankel matrices. Linear Algebra and its Applications, 191:77–106, September 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shaked-Monderer:1993:ECD

[SM93] Naomi Shaked-Monderer. Extreme chordal doubly nonnegative matrices with given row sums. Linear Algebra and its Applications, 183:23–??, April 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shaked-Monderer:1992:MEP

[SMB92] Naomi Shaked-Monderer and Abraham Berman. More on extremal positive semidefinite doubly stochastic matrices. Linear Algebra and its Applications, 167:17–??, April 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smith:1990:BSN

[Smi90] Ronald L. Smith. Bounds on the spectrum of nonnegative matrices and certain Zmatrices. Linear Algebra and its Applications, 129:13–28, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smith:1992:SIP

[Smi92] Ronald L. Smith. Some interlacing properties of the Schur complement of a Hermitian matrix. Linear Algebra and its Applications, 177:137–??, December 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smith:1995:SRP

[Smi95] Ronald L. Smith. Some results on a partition of Z-matrices. Linear Algebra and its Applications, 223/224(1–3):619– 629, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400168. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Starke:1991:SXC

[SN91] Gerhard Starke and Wilhelm Niethammer. SOR for AX − XB = C. Linear Algebra and its Applications, 154/156: 355–375, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shen:1998:LEP

[SN98a] Jian Shen and Stewart Neufeld. Local exponents of primitive digraphs. Linear Algebra and its Applications, 268(1–3):117–129, January 1, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=268&issue=1-
3&aid=9700036.

Shen:1998:PL

[SN98b] Jiang Shen and Stewart Neufeld. On a problem of Lewin. Linear Algebra and its Applications, 274(1–3):411– 426, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700389.

Sayed:1998:ECB

[SNC98] A. H. Sayed, V. H. Nascimento, and S. Chandrasekaran. Estimation and control with bounded data uncertainties. Linear Algebra and its Applications, 284(1–3):259–306, November 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6246.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6246.

Sefton:1994:HNA

[SO94a] J. A. Sefton and R. J. Ober.

Hankel-Norm approximation and control systems. Linear Algebra and its Applications, 205/206:1081–1120, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

So:1994:CSH

[So94b] Wasin So. Commutativity and spectra of Hermitian matrices. Linear Algebra and its Applications, 212/213:121– 129, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

So:1999:LOP

[So99]  Wasin So.  Linear oper-

ators preserving the minimal rank. Linear Algebra and its Applications, 302-303(1–3):461–468, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        52/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/52/article.
                                        pdf.

Sobczyk:1996:SFA

[Sob96] Garret Sobczyk. Structure of factor algebras and Clifford algebra. Linear Algebra and its Applications, 241/243(1–3):803–810, July/ August 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=241-243&issue=
                                        1-3&aid=9500604. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Soh:1990:SSC

[Soh90] C. B. Soh. Schur stability of convex combination of matrices. Linear Algebra and its Applications, 128:159–168, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soh:1991:NSC

[Soh91] C. B. Soh. Necessary and sufficient conditions for interval polynomials to have only real distinct roots. Linear Algebra and its Applications, 144: 121–133, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soh:1997:LPP

[Soh97] C. B. Soh. On the largest polytope of polynomials with a precise number of distinct real zeros. Linear Algebra and its Applications, 265(1–3):173–183, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600599.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600599.

Solomon:1999:SCM

[Sol99] Louis Solomon. Similarity of the companion matrix and its transpose. Linear Algebra and its Applications, 302-303(1–3):555–561, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/117/25/
                                        55/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/55/article.
                                        pdf.

Song:1991:MPC

[Son91a] Seok-Zun Song. Minimum permanents on certain doubly stochastic matrices. Linear Algebra and its Applications, 143:49–56, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:1991:CNS

[Son91b] Yong Zhong Song. Comparisons of nonnegative splittings of matrices. Linear Algebra and its Applications, 154/156: 433–455, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:1992:LBP

[Son92]  Yongzhong Song.  Lower

bounds for the Perron root of a nonnegative matrix. Linear Algebra and its Applications, 169:269–??, May 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:1995:CP

[Son95] Seok-Zun Song. A conjecture on permanents. Linear Algebra and its Applications, 222(1–3):91–95, June ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300286.

Song:1997:CGA

[Son97] Yongzhong Song. On the convergence of the generalized AOR method. Linear Algebra and its Applications, 256(1–3): 199–218, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600028.

Soules:1991:RCS

[Sou91] George W. Soules. The rate of convergence of Sinkhorn balancing. Linear Algebra and its Applications, 150:3–40, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Soules:1994:APD

[Sou94] George W. Soules. An approach to the permanentaldominance conjecture. Linear Algebra and its Applications, 201:211–229, April 1, 1994. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Smit:1999:EIS

[SP99] P. Smit and M. H. C. Paardekooper. The effects of inexact solvers in algorithms for symmetric eigenvalue problems. Linear Algebra and its Applications, 287 (1–3):337–357, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6308.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6308.

Spain:1995:DVK

[Spa95a] P. G. Spain. A discrete variation on Kronecker’s theorem. Linear Algebra and its Applications, 223/224(1–3):631– 636, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400340. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Spain:1995:TPR

[Spa95b] P. G. Spain. Tracking poles, representing Hankel operators, and the Nehari problem.

Linear Algebra and its Applications, 223/224(1–3):637– 694, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400106. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Spanoudakis:1997:OFD

[Spa97] N. K. Spanoudakis. Operators in finite distributive subspace lattices. III. Linear Algebra and its Applications, 262(1–3):189–207, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600482.

Spence:1995:RTG

[Spe95]  Edward Spence.  Regu-

lar two-graphs on 36 vertices. Linear Algebra and its Applications, 226/228 (1–3):459–497, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1995&volume=226-228&issue=
                                        1-3&aid=9500158.

Spiegel:1993:SP

[Spi93] Eugene Spiegel. Sums of projections. Linear Algebra and its Applications, 187:239–249, July 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Spitkovsky:1994:OMA

[Spi94] Ilya Spitkovsky. Once more on algebras generated by two projections. Linear Algebra and its Applications, 208/209: 377–395, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Spiegel:1995:ESP

[Spi95] E. Spiegel. Erratum: “Sums of projections” [Linear Algebra Appl. 187 (1993), 239– 249, MR 94m:15001]. Linear Algebra and its Applications, 220(1–3):6, April 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Scott:1990:WLS

[SRT90] A. J. Scott, J. N. K. Rao, and D. Roland Thomas. Weighted least-squares and quasilikelihood estimation for categorical data under singular models. Linear Algebra and its Applications, 127:427–447, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salakhov:1990:RAL

[SS90] M. Kh. Salakhov and N. K.

Shcherbakova. A regularized algorithm for local emission reconstruction in spectroscopic tomography. Linear Algebra and its Applications, 130:219–229, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Saylor:1991:IAA

[SS91] Paul E. Saylor and Dennis C.

Smolarski. Implementation of an adaptive algorithm for Richardson’s method. Linear Algebra and its Applications, 154/156:615–646, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schweitzer:1993:LEP

[SS93] Paul J. Schweitzer and G. W.

Stewart. The Laurent expansion of pencils that are singular at the origin. Linear Algebra and its Applications, 183: 237–254, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shapiro:1995:BTP

[SS95a] B. Z. Shapiro and M. Z.

Shapiro. On the boundary of totally positive upper triangular matrices. Linear Algebra and its Applications, 231(1–3):105–109, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400013.

Sloboda:1995:ISL

[SS95b] Fridrich Sloboda and Fiorella Sgallari. On iterative solution of linear equations arising in BVPs of ODEs. Linear Algebra and its Applications, 225(1):195–205, August ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9500340.

Spijker:1996:SEF

[SS96] M. N. Spijker and F. A. J. Straetemans. Stability estimates for families of matrices of nonuniformly bounded order. Linear Algebra and its Applications, 239(1–3):77– 102, May ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400169.

Stoorvogel:1998:DAR

[SS98] Anton A. Stoorvogel and Ali Saberi. The discrete algebraic Riccati equation and linear matrix inequality. Linear Algebra and its Applications, 274(1–3):317– 365, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=274&issue=1-
3&aid=9700365.

Sung:1992:GCM

[SSHS92] Je Cho Sung, Kye Seung-Hyeok, and Ge Lee Sa. Generalized Choi maps in threedimensional matrix algebra. Linear Algebra and its Applications, 171:213–??, July 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Styan:1997:IEA

[SSS97] George P. H. Styan and Gerald E. Subak-Sharpe. Inequalities and equalities associated with the Campbell-Youla generalized inverse of the indefinite admittance matrix of resistive networks. Linear Algebra and its Applications, 250(1–3):349–370, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500601.

Sung:1990:SPE

[ST90] Chen Han Sung and Bit-Shun Tam. A study of projectionally exposed cones. Linear Algebra and its Applications, 139:225–252, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Santi:1991:SRC

[ST91] Barbara A. Li Santi and Robert C. Thompson. Simultaneous reduction of a complex skew matrix and a Hermitian matrix. Linear Algebra and its Applications, 147: 275–322, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stoorvogel:1993:FHS

[ST93] A. A. Stoorvogel and H. L. Trentelman. The finitehorizon singular H_{∞} control problem with dynamic measurement feedback. Linear Algebra and its Applications, 187:113–161, July 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Storme:1995:APF

[ST95] L. Storme and J. A. Thas. karcs and partial flocks. Linear Algebra and its Applications, 226/228(1–3):33–45, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400231.

Saldanha:1996:ADQ

[ST96a] Nicolau C. Saldanha and Carlos Tomei. The accumulated distribution of quadratic forms on the sphere. Linear Algebra and its Applications, 245(1–3):335–351, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400339.

Sultan:1996:MCM

[ST96b] Shagufta A. Sultan and Derrick S. Tracy. Moments of the complex multivariate normal distribution. Linear Algebra and its Applications, 237/238(1–3):191– 204, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500559. Special

issue honoring Calyampudi Radhakrishna Rao.

Starke:1993:FVA

[Sta93] Gerhard Starke. Fields of values and the ADI method for non-normal matrices. Linear Algebra and its Applications, 180:199–??, February 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:1989:SPP

[Ste89] G. W. Stewart. On scaled projections and pseudo-inverses. Linear Algebra and its Applications, 112(??):189–193, January 1989. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0024379589905946. See an-

swer [O’L90] to a question raised in this paper.

Steidl:1990:GAD

[Ste90] Gabriele Steidl. Generalization of the algebraic discrete Fourier transform with application to fast convolutions. Linear Algebra and its Applications, 139:181–206, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:1993:IFD

[Ste93] David E. Stewart. An index formula for degenerate LCPs. Linear Algebra and its Applications, 191:41–51, September 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:1994:PTR

[Ste94] G. W. Stewart. Perturbation theory for rectangular matrix pencils. Linear Algebra and its Applications, 208/209:297– 301, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sterk:1995:LSD

[Ste95] Hans Sterk. Lattices and K3 surfaces of degree 6. Linear Algebra and its Applications, 226/228(1–3):297–309, September/October 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500150.

Stepniak:1997:MLC

[Ste97a] Czesław Stepniak. Matrix loss in comparison of linear experiments. Linear Algebra and its Applications, 264(1–3):341–348, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700019.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.

cgi?year=1997&volume=264&
issue=1-3&aid=9700019.

Stewart:1997:GTM

[Ste97b] D. E. Stewart. A graphtheoretic model of symmetric Givens operations and its implications. Linear Algebra and its Applications, 257(1–3): 311–320, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600157.

Stewart:1997:CFS

[Ste97c] Michael Stewart. Cholesky factorization of semidefinite Toeplitz matrices. Linear Algebra and its Applications, 254(1–3):497–525, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600517. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Stewart:1998:AM

[Ste98] G. W. Stewart. On the adjugate matrix. Linear Algebra and its Applications, 283 (1–3):151–164, November 1, 1998. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6217.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6217.

Stevanovic:1999:WNG

[Ste99] Dragan Stevanović. When is NEPS of graphs connected? Linear Algebra and its Applications, 301 (1–3):137–144, November 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/33/article.
pdf.

Stickel:1991:SEU

[Sti91] Eberhard U. Stickel. Separating eigenvalues using the matrix sign function. Linear Algebra and its Applications, 148:75–88, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stoorvogel:1996:SSA

[Sto96] Anton A. Stoorvogel. Stabilizing solutions of the H_{∞} algebraic Riccati equation. Linear Algebra and its Applications, 240(1–3):153–172, June ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400195.

Storjohann:1998:CHS

[Sto98] Arne Storjohann. Computing Hermite and Smith normal forms of triangular integer matrices. Linear Algebra and its Applications, 282(1–3):25–45, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6133.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6133.

Stojanoff:1999:IHM

[Sto99] Demetrio Stojanoff. Index of Hadamard multiplication by positive matrices. Linear Algebra and its Applications, 290(1–3):95–108, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6321.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

year=1999&volume=290&issue=
1-3&aid=6321.

Strakos:1991:RCR

[Str91] Z. Strakoš. On the real convergence rate of the conjugate gradient method. Linear Algebra and its Applications, 154/156:535–549, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strakos:1996:RIS

[Str96] Zdenek Strakos. Review of Iterative Solution Methods by O. Axelsson. Linear Algebra and its Applications, 240(1–3): 231–233, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400230.

Strang:1997:EUM

[Str97a] Gilbert Strang. Every unit matrix is a LULU. Linear Algebra and its Applications, 265(1–3):165–172, November 1, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/265/1-3/9600598.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=265&
issue=1-3&aid=9600598.

Strikwerda:1997:CTC

[Str97b] John C. Strikwerda. A convergence theorem for chaotic asynchronous relaxation. Linear Algebra and its Applications, 253(1–3):15–24, March 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=253&issue=1-
                                        3&aid=9500698.

Strauss:1998:EEI

[Str98] V. A. Strauss. On an estimate of the eigenvalues for an infinite-dimensional matrix and its application to the problem of the completeness of an eigenvector system of a completely continuous operator. Linear Algebra and its Applications, 274(1–3): 61–76, April 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=274&issue=1-
                                        3&aid=9700247.

Stuart:1990:POG

[Stu90] Jeffrey L. Stuart. The partial order graph for a ZMEmatrix. Linear Algebra and its Applications, 141:123–152, 1990. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Stuart:1999:RSP

[Stu99]  Jeffrey Stuart.  Reducible

sign k-potent sign pattern matrices. Linear Algebra and its Applications, 294(1– 3):197–211, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6414.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6414.

Su:1996:GMA

[Su96] Yangfeng Su. Generalized multisplitting asynchronous iteration. Linear Algebra and its Applications, 235(1–3):77– 92, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400120.

Sun:1990:MES

[Sun90a] Ji Guang Sun. Multiple eigenvalue sensitivity analysis. Linear Algebra and its Applications, 137/138:183–211, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1990:RED

[Sun90b] Ji Guang Sun. Relative eigenvalues of a definite matrix pair. Linear Algebra and its Applications, 139:253–267, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1991:PEI

[Sun91] Ji Guang Sun. Perturbation expansions for invariant subspaces. Linear Algebra and its Applications, 153:85–97, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1992:REP

[Sun92a] Ji-Guang Sun. Roundingerror and perturbation bounds for the Cholesky and LDL^{T} factorizations. Linear Algebra and its Applications, 173: 77–98, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1992:ICS

[Sun92b] W. Sun. The interior and closure of strongly stable matrices. Linear Algebra and its Applications, 165:53–??, March 1, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1994:SHI

[Sun94] Xingping Sun. Scattered Hermite interpolation using radial basis functions. Linear Algebra and its Applications, 207: 135–146, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:1995:PBF

[Sun95] Ji Guang Sun. On perturbation bounds for the QR factorization. Linear Algebra and its Applications, 215(1–3):95– 111, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300077.

Sun:1996:VSN

[Sun96a]  Ji-Guang Sun.  On the

variation of the spectrum of a normal matrix. Linear Algebra and its Applications, 246(1–3):215–223, October ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400354.

Sunwoo:1996:SAA

[Sun96b] Hasik Sunwoo. Simple algorithms about Kronecker products in the linear model. Linear Algebra and its Applications, 237/238(1–3):351– 358, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500675. Special

issue honoring Calyampudi Radhakrishna Rao.

Sinap:1994:PIG

[SV94] Ann Sinap and Walter Van Assche. Polynomial interpolation and Gaussian quadrature for matrix-valued functions. Linear Algebra and its Applications, 207:71–114, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Slapnicar:1995:PEF

[SV95] Ivan Slapničar and Krešimir Veselić. Perturbations of the eigenprojections of a factorized Hermitian matrix. Linear Algebra and its Applications, 218(1–3):273–280, March 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=218&issue=1-
                                        3&aid=9300190.

Sleijpen:1996:FRC

[SV96] G. L. G. Sleijpen and A. Van Der Sluis. Further results on the convergence behavior of conjugate-gradients and Ritz values. Linear Algebra and its Applications, 246(1–3):233– 278, October ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=246&issue=1-
3&aid=9400360.

Say:1998:MEB

[SV98] Song Goh Say and Bing Yap Von. Matrix extension and biorthogonal multiwavelet construction. Linear Algebra and its Applications, 269(1): 139–157, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.elsevier.com/cgi-bin/
cas/tree/store/laa/cas_
sub/browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9781514.

Slapnicar:1999:BCH

[SV99] Ivan Slapničar and Krešimir Veselic´ . A bound for the condition of a hyperbolic eigenvector matrix. Linear Algebra and its Applications, 290(1–3):247–255, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6346.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?

                                        year=1999&volume=290&issue=
                                        1-3&aid=6346.

Sontag:1990:PSF

[SW90] Eduardo D. Sontag and Yuan Wang. Pole shifting for families of linear systems depending on at most three parameters. Linear Algebra and its Applications, 137/138:3– 38, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stern:1991:RIC

[SW91a] R. J. Stern and H. Wolkowicz. Results on invariant cones. Linear Algebra and its Applications, 166:1–26, 1991. Proceedings from the Haifa Matrix Theory Conference, June 1990.

Stern:1991:IEC

[SW91b] Ronald J. Stern and Henry Wolkowicz. Invariant ellipsoidal cones. Linear Algebra and its Applications, 150:81– 106, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Stern:1991:NGI

[SW91c] Ronald J. Stern and Henry Wolkowicz. A note on generalized invariant cones and the Kronecker canonical form. Linear Algebra and its Applications, 147:97–100, 1991.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strikwerda:1991:CMK

[SW91d] John C. Strikwerda and Bruce A. Wade. Cesàro means and the Kreiss matrix theorem. Linear Algebra and its Applications, 145:89–106, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stuart:1991:MCP

[SW91e] Jeffrey L. Stuart and James R.

Weaver. Matrices that commute with a permutation matrix. Linear Algebra and its Applications, 150:255–265, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Stuart:1994:DSP

[SW94] Jeffrey L. Stuart and James R.

Weaver. Diagonally scaled permutations and circulant matrices. Linear Algebra and its Applications, 212/213:397– 411, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Shih:1995:TRS

[SW95] Mau Hsiang Shih and Jinn Wen Wu. Two results on the spectral radius of a complex matrix, with application to instability. Linear Algebra and its Applications, 222(1–3): 147–153, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300293.

Schneider:1998:MWN

[SW98a] Hans Schneider and Hans F. Weinberger. On matrices for which norm bounds are attained. Linear Algebra and its Applications, 275/276(1– 3):563–577, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6019.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6019. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Stuart:1998:FMT

[SW98b] Jeffrey L. Stuart and James R. Weaver. Fiedler matrices and their factorization. Linear Algebra and its Applications, 275/276(1–3):579–594, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6048.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6048. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Sweet:1993:QCR

[Swe93] Daniel Sweet. Quadratic characterizations for reciprocal linear difference equations. Linear Algebra and its Applications, 194:1–7, November 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shih:1997:ASG

[SWP97] Mau-Hsiang Shih, Jinn-Wen Wu, and Chin-Tzong Pang. Asymptotic stability and generalized Gelfand spectral radius formula. Linear Algebra and its Applications, 252(1–3): 61–70, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500592.

Stern:1995:VAE

[SY95] Ronald J. Stern and J. J.

Ye. Variational analysis of an extended eigenvalue problem. Linear Algebra and its Applications, 220(1–3):391– 417, April 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=220&issue=1-
3&aid=9200016. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Simons:1998:AIS

[SY98] Gordon Simons and Yi-Ching Yao. Approximating the inverse of a symmetric positive definite matrix. Linear Algebra and its Applications, 281 (1–3):97–103, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6159.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6159.

Schwarz:1990:NEM

[SZ90] Binyamin Schwarz and Abraham Zaks. Non-Euclidean motions in projective matrix spaces. Linear Algebra and its Applications, 137/138:351– 361, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salce:1993:CPM

[SZ93a] Luigi Salce and Paolo Zanardo. Completely positive matrices and positivity of least squares solutions. Linear Algebra and its Applications, 178:201–??, January 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Syrmos:1993:CNE

[SZ93b] Vassilis Syrmos and Petr Zagalak. Computing normal external descriptions and feedback design. Linear Algebra and its Applications, 188/189: 613–639, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szulc:1990:CTM

[Szu90a] Tomasz Szulc. A contribution to the theory of P -matrices. Linear Algebra and its Applications, 139:217–224, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szulc:1990:CNM

[Szu90b] Tomasz Szulc. A criterion for the nonsingularity of matrices. Linear Algebra and its Applications, 136:209–213, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szulc:1992:NCC

[Szu92a] Tomasz Szulc. A nonsingularity criterion for complex matrices. Linear Algebra and its Applications, 168:145–??, April 15, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szulc:1992:CNC

[Szu92b] Tomasz Szulc. On a criterion for the nonsingularity of complex matrices. Linear Algebra and its Applications, 173: 39–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szulc:1995:SRT

[Szu95]  Tomasz Szulc.  Some re-

marks on a theorem of Gudkov. Linear Algebra and its Applications, 225(1):221– 235, August ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=225&issue=1&
aid=9500343.

Szyld:1992:SLB

[Szy92]  Daniel B. Szyld.  A se-

quence of lower bounds for the spectral radius of nonnegative matrices. Linear Algebra and its Applications, 174: 239–??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taboada:1993:TLA

[Tab93] Mario Taboada. A theorem in linear algebra with applications to the geometry of quadrics. Linear Algebra and its Applications, 185:31– ??, May 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takahashi:1996:EMG

[Tak96] Katsutoshi Takahashi. Eigenvalues of matrices with given block upper triangular part. Linear Algebra and its Applications, 239(1–3):175–184, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400182.

Tam:1992:SCP

[Tam92] Bit-Shun Tam. On the structure of the cone of positive operators. Linear Algebra and its Applications, 167:65–??, April 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tam:1997:MTT

[Tam97] Tin-Yau Tam. Miranda and Thompson’s trace inequality and a log convexity result. Linear Algebra and its Applications, 262(1–3):307–325, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600521.

Tam:1998:LMT

[Tam98] Tin-Yau Tam. On Lei, Miranda, and Thompson’s result on singular values and diagonal elements. Linear Algebra and its Applications, 272(1–3): 91–101, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700310.

Tam:1999:MCS

[Tam99]  Bit-Shun Tam.  On ma-

trices with cyclic structure. Linear Algebra and its Applications, 302-303(1– 3):377–410, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
48/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/48/article.
pdf.

Tan:1998:EEM

[Tan98] Yi-Jia Tan. Eigenvalues and eigenvectors for matrices over distributive lattices. Linear Algebra and its Applications, 283(1–3):257–272, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6224.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6224.

Tarazaga:1991:MEE

[Tar91] Pablo Tarazaga. More estimates for eigenvalues and singular values. Linear Algebra and its Applications, 149:97– 110, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tasci:1995:CGN

[Taş95] Dursun Taşci. On a conjecture by Goldberg and Newman. Linear Algebra and its Applications, 215(1–3):275– 277, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300099.

Towers:1997:HOB

[TB97] D. A. Towers and K. Bowman.

Higher-order Bernstein algebras given by symmetric bilinear forms. Linear Algebra and its Applications, 252(1–3):71– 79, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500673.

Techene:1996:NPO

[Téc96] Jean-Jacques Téchené. Sur la notion de projection orthogonale dans un espace semieuclidien. (French) [on the notion of orthogonal projection in a semi-euclidean space]. Linear Algebra and its Applications, 237/238(1–3):239– 268, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=237-238&issue=
1-3&aid=9500526. Special

issue honoring Calyampudi Radhakrishna Rao.

Techene:1997:AFA

[Téc97] Jean-Jacques Téchené. Les aspects fondamentaux de l’admissibilité en approximation quadratique d’applications linéaires. (French) [Fundamental aspects of admissibility in the quadratic approximation of linear mappings]. Linear Algebra and its Applications, 264(1–3):389–419, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600405.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600405.

Teitloff:1999:PPS

[Tei99] Timothy C. Teitloff. Permutation polynomials on symmetric matrices. Linear Algebra and its Applications, 296(1–3):233–243, July 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/31/
abstract.html;   http://
www.elsevier.nl/gej-ng/
10/30/19/47/17/31/article.
pdf.

Teng:1990:IEP

[Ten90] De Gui Teng. Inverse eigenvalue problem for some special matrices. Linear Algebra and its Applications, 139:63– 65, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Terrell:1998:IOR

[Ter98] William J. Terrell. An inputoutput representation for implicit linear time-varying systems. Linear Algebra and its Applications, 271(1–3):221– 234, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=271&issue=1-
3&aid=9700269.

Tsing:1994:AFI

[TFV94] Nam-Kiu Tsing, Michael K. H. Fan, and Erik I. Verriest. On analyticity of functions involving eigenvalues. Linear Algebra and its Applications, 207:159–180, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trefethen:1997:DPM

[TG97] Lloyd N. Trefethen and A. Greenbaum. Do the pseudospectra of a matrix determine its behavior? Linear Algebra and its Applications, ??: ??, 1997. To appear. The first author reports that the paper has been accepted, subject to minor revisions.

Thang:1999:RPP

[Thǎ99] Nguyêñ Quôc´ Thǎn´ g. A remark on pattern problems for matrix groups. Linear Algebra and its Applications, 292(1–3):179–185, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6365.pdf;   http:

//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6365.

Thijsse:1995:SPR

[Thi95] G. Philip A. Thijsse. Symmetry properties of and reduction principles for divisibility relations between the invariant factors of products of holomorphic matrix functions. Linear Algebra and its Applications, 218(1–3):113– 145, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300165.

Thijsse:1997:UTS

[Thi97] Philip Thijsse. Upper triangular similarity of upper triangular matrices. Linear Algebra and its Applications, 260(1–3): 119–149, July 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600297.

Thompson:1991:PCR

[Tho91] Robert C. Thompson. Pencils of complex and real symmetric and skew matrices. Linear Algebra and its Applications, 147:323–371, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thompson:1992:HLQ

[Tho92] R. C. Thompson. High, low, and quantitative roads in linear algebra. Linear Algebra and its Applications, 162/164: 23–64, 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thompson:1995:RSP

[Tho95] Robert C. Thompson. Root spreads for polynomials and Hermitian matrix pencils. Linear Algebra and its Applications, 220(1–3):419–433, April 15, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=220&issue=1-
                                        3&aid=9400051. Proceedings

of the Workshop “Nonnegative Matrices, Applications and Generalizations” and the Eighth Haifa Matrix Theory Conference (Haifa, 1993).

Thomassen:1996:ESN

[Tho96]  C. Thomassen.  Erratum:

“Sign-nonsingular matrices and even cycles in directed graphs” [Linear Algebra Appl. 75 (1986), 27–41, MR 87k:05120]. Linear Algebra and its Applications, 240 (1–3):238, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thompson:1997:UNR

[Tho97] Robert C. Thompson. The upper numerical range of a quaternionic matrix is not a complex numerical range. Linear Algebra and its Applications, 254(1–3):19–28, March 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9781634. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

Tarazaga:1996:CED

[THW96] P. Tarazaga, T. L. Hayden, and Jim Wells. Circum-Euclidean distance matrices and faces. Linear Algebra and its Applications, 232(1–3):77– 96, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:1994:ROL

[Tia94] Yong Ge Tian. Reverse order laws for the generalized inverses of multiple matrix products. Linear Algebra and its Applications, 211:85–100, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Generalized inverses (1993).

Tian:1998:MPI

[Tia98] Yongge Tian. The Moore– Penrose inverses of m × n block matrices and their applications. Linear Algebra and its Applications, 283(1– 3):35–60, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6170.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6170.

Tilli:1998:LTS

[Til98a] P. Tilli. Locally Toeplitz sequences: Spectral properties and applications. Linear Algebra and its Applications, 278(1–3):91–120, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6079.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6079.

Tilli:1998:SVE

[Til98b] Paolo Tilli. Singular values and eigenvalues of non-Hermitian block Toeplitz matrices. Linear Algebra and its Applications, 272(1–3):59– 89, March 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=272&issue=1-
3&aid=9700308.

Tismenetsky:1991:NPT

[Tis91a]  M. Tismenetsky.  A new

preconditioning technique for solving large sparse linear systems. Linear Algebra and its Applications, 154/156:331– 353, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tismenetsky:1991:DTM

[Tis91b] Miron Tismenetsky. A decomposition of Toeplitz matrices and optimal circulant preconditioning. Linear Algebra and its Applications, 154/156: 105–121, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tismenetsky:1992:SPS

[Tis92] Miron Tismenetsky. Some properties of solutions of Yule-Walker type equations. Linear Algebra and its Applications, 173:1–??, August 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Toraichi:1995:KPS

[TK95] Kazuo Toraichi and Masaru Kamada. Knot positions for the smoothest periodic quadratic spline interpolation of equispaced data. Linear Algebra and its Applications, 221(1–3):245–251, May ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300259.

Tasci:1998:SUB

[TK98] Dursun Taşçi and Steve Kirkland. A sequence of upper bounds for the Perron root of a nonnegative matrix. Linear Algebra and its Applications, 273(1–3):23– 28, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9700337.

TenBerge:1999:SCA

[TK99] Jos M. F. Ten Berge and Henk A. L. Kiers. Simplicity of core arrays in three-way principal component analysis and the typical rank of p × q × 2 arrays. Linear Algebra and its Applications, 294(1–3):169–179, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6404.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6404.

tenBerge:1999:SNR

[tKWS99] Jos M. F. ten Berge, Wim P.

Krijnen, Tom Wansbeek, and Alexander Shapiro. Some new results on correlationpreserving factor scores prediction methods. Linear Algebra and its Applications, 289(1–3):311–318, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6007.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6007.

Todd:1991:ASD

[Tod91] Michael J. Todd. The affinescaling direction for linear programming is a limit of projective-scaling directions. Linear Algebra and its Applications, 152:93–105, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Toffoli:1997:AEU

[Tof97] Tommaso Toffoli. Almost every unit matrix is a ULU. Linear Algebra and its Applications, 259(1–3):31–38, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=259&issue=1-
                                        3&aid=9600240.

Tong:1994:SRD

[Ton94] Y. L. Tong. Some recent developments on majorization inequalities in probability and statistics. Linear Algebra and its Applications, 199:69–90, March 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tong:1997:RCM

[Ton97] Y. L. Tong. The role of the covariance matrix in the least-squares estimation for a common mean. Linear Algebra and its Applications, 264(1–3):313–323, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/laa/sub/
                                        1997/264/1-3/9600411.pdf;
                                         http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/

laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600411.

Trefethen:1992:EPE

[TR92] Lloyd N. Trefethen and L. Reichel. Eigenvalues and pseudoeigenvalues of Toeplitz matrices. Linear Algebra and its Applications, 162/164:153– 185, 1992.

Trench:1990:NTI

[Tre90] William F. Trench. A note on a Toeplitz inversion formula. Linear Algebra and its Applications, 129:55–61, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:1991:NSE

[Tre91] William F. Trench. Numerical solution of the eigenvalue problem for efficiently structured Hermitian matrices. Linear Algebra and its Applications, 154/156:415– 432, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:1993:IEO

[Tre93] William F. Trench. Interlacement of the even and odd spectra of real symmetric Toeplitz matrices. Linear Algebra and its Applications, 195:59–68, December 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:1999:ADE

[Tre99a] William F. Trench. Asymptotic distribution of the even and odd spectra of real symmetric Toeplitz matrices. Linear Algebra and its Applications, 302-303(1– 3):155–162, December 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/34/article.
pdf.

Trench:1999:ADS

[Tre99b] William F. Trench. Asymptotic distribution of the spectra of a class of generalized Kac–Murdock–Szegő matrices. Linear Algebra and its Applications, 294(1– 3):181–192, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6427.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6427.

Tolsa:1997:CSP

[TS97] Javier Tolsa and Miquel Salichs. Convergence of singular perturbations in singular linear systems. Linear Algebra and its Applications, 251(1–3):105–143, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500556.

Truhar:1999:RPB

[TS99] Ninoslav Truhar and Ivan Slapničar. Relative perturbation bound for invariant subspaces of graded indefinite Hermitian matrices. Linear Algebra and its Applications, 301(1–3):171–185, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/115/25/36/article.
pdf.

Taussky:1991:ALM

[TT91] Olga Taussky and John Todd.

Another look at a matrix of Mark Kac. Linear Algebra and its Applications, 150:341– 360, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Tutuncu:1995:RHL

[TT95] Reha H. Tütüncü and Michael J. Todd. Reducing horizontal linear complementarity problems. Linear Algebra and its Applications, 223/224(1–3): 717–729, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=223-224&issue=
                                        1-3&aid=9400147. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Tuck:1996:UBG

[Tuc96] Sang Leong Tuck. An upper bound on the growth ratio of Gaussian elimination. Linear Algebra and its Applications, 235(1–3):141– 152, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=235&issue=1-
                                        3&aid=9400128.

Turner:1991:CPK

[Tur91] Kathryn Turner. Computing projections for the Karmarkar algorithm. Linear Algebra and its Applications, 152:141– 154, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Turkowski:1992:SRL

[Tur92] P. Turkowski. Structure of real Lie algebras. Linear Algebra and its Applications, 171: 197–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tsaklidis:1990:IPM

[TV90] G. Tsaklidis and P.-C. G. Vassiliou. Infinite products of matrices with some negative elements and row sums equal to one. Linear Algebra and its Applications, 127:41–58, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tsaklidis:1992:PIP

[TV92] G. Tsaklidis and P.-C. G. Vassiliou. Periodicity of infinite products of matrices with some negative elements and row sums equal to one. Linear Algebra and its Applications, 176:175–??, November 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Todd:1999:HIU

[TV99] John Todd and Richard S.

Varga. Hardy’s inequality and ultrametric matrices. Linear Algebra and its Applications, 302-303(1–3):33–43, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/28/article.
pdf.

Turke:1995:MPI

[TW95] C. Türke and M. Weber. On a maximum principle for inverse monotone matrices. Linear Algebra and its Applications, 218(1–3):47– 57, March 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=218&issue=1-
3&aid=9300156.

Thomova:1999:MAS

[TW99] Z. Thomova and P. Winternitz. Maximal Abelian subalgebras of e(p, q) algebras. Linear Algebra and its Applications, 291(1–3):245–274, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6357.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6357.

Tartar:1997:SCA

[TY97] Luc Tartar and Arieh Yakir.

Similarity classes of affine maps. Linear Algebra and its Applications, 261(1–3):155– 165, August ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600365.

Takane:1999:OP

[TY99a] Yoshio Takane and Haruo Yanai. On oblique projectors. Linear Algebra and its Applications, 289(1–3): 297–310, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
289/1-3/6290.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=289&issue=
1-3&aid=6290.

Tam:1999:MWN

[TY99b] Bit-Shun Tam and Shangjun Yang. On matrices whose numerical ranges have circular or weak circular symmetry. Linear Algebra and its Applications, 302-303(1–3):193–221, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/

                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/117/25/37/article.
                                        pdf.

Tyrtyshnikov:1991:CTM

[Tyr91] Evgenij E. Tyrtyshnikov. Cauchy–Toeplitz matrices and some applications. Linear Algebra and its Applications, 149:1–18, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tyrtyshnikov:1994:IMO

[Tyr94] Evgenij E. Tyrtyshnikov. Influence of matrix operations on the distribution of eigenvalues and singular values of Toeplitz matrices. Linear Algebra and its Applications, 207:225–249, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tyrtyshnikov:1995:CPU

[Tyr95] E. E. Tyrtyshnikov. Circulant preconditioners with unbounded inverses. Linear Algebra and its Applications, 216(1–3):1–23, February ??, 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=216&issue=1-
                                        3&aid=9300092.

Tyrtyshnikov:1996:UAS

[Tyr96] Evgenij E. Tyrtyshnikov. A unifying approach to some old and new theorems on distribution and clustering. Linear Algebra and its Applications, 232(1–3):1–43, January 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tyrtyshnikov:1997:MBD

[Tyr97]  E. Tyrtyshnikov.  Matrix

Bruhat decompositions with a remark on the QR (GR) algorithm. Linear Algebra and its Applications, 250(1–3):61– 68, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500453.

Tam:1997:IIM

[TYZ97a] Bit-Shun Tam, Shangjun Yang, and Xiaodong Zhang. Invertibility of irreducible matrices. Linear Algebra and its Applications, 259(1–3): 39–70, July 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600241.

Tyrtyshnikov:1997:CPC

[TYZ97b] E. E. Tyrtyshnikov, A. Yu. Yeremin, and N. L. Zamarashkin. Clusters, preconditioners, convergence. Linear Algebra and its Applications, 263(1–3):25–48, September 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=263&issue=1-
3&aid=9600445.

Tapia:1991:OBI

[TZ91] R. A. Tapia and Yin Zhang. An optimal-basis identification technique for interiorpoint linear programming algorithms. Linear Algebra and its Applications, 152:343–363, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Tan:1996:PGC

[TZ96] Yi-Jia Tan and Mou-Cheng Zhang. Primitivity of generalized circulant Boolean matrices. Linear Algebra and its Applications, 234(1–3):61– 69, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=234&issue=1-
3&aid=9400084.

Tyrtyshnikov:1998:SMT

[TZ98] E. E. Tyrtyshnikov and N. L. Zamarashkin. Spectra of multilevel Toeplitz matrices: advanced theory via simple matrix relationships. Linear Algebra and its Applications, 270(1–3):15– 27, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=270&issue=1-
3&aid=9780001.

Tyrtyshnikov:1999:TSE

[TZ99] E. E. Tyrtyshnikov and N. L.

Zamarashkin. Thin structure of eigenvalue clusters for non-Hermitian Toeplitz matrices. Linear Algebra and its Applications, 292(1– 3):297–310, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6392.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6392.

Uhlig:1992:RTM

[Uhl92] Frank Uhlig. Review of Topics in Matrix Analysis by Roger A. Horn and Charles R. Johnson. Linear Algebra and its Applications, 174:243– ??, September 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhlig:1994:CIS

[Uhl94] Frank Uhlig. Computing the inertias in symmetric matrix pencils. Linear Algebra and its Applications, 201:199–209, April 1, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhlig:1998:CMT

[Uhl98] F. Uhlig. Challenges in Matrix Theory 1998. Linear Algebra and its Applications, 278(1–3): 285–??, July 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhrin:1992:TPS

[Uhr92] B. Uhrin. Touching points of a star-shaped set with an affine subspace. Linear Algebra and its Applications, 171: 149–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhrin:1996:IAP

[Uhr96] B. Uhrin. Inner aperiodicities and partitions of sets. Linear Algebra and its Applications, 241/243(1–3):851– 876, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500194. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Uteshev:1992:HMS

[US92] Alexei Yu. Uteshev and Sergei G. Shulyak. Hermite’s method of separation of solutions of systems of algebraic equations and its applications. Linear Algebra and its Applications, 177:49–??, December 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Usmani:1994:ITJ

[Usm94] Riaz A. Usmani. Inversion of a tridiagonal Jacobi matrix. Linear Algebra and its Applications, 212/213:413– 414, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Valiaho:1991:PTT

[Väl91] H. Väliaho. A polynomialtime test for M-matrices. Linear Algebra and its Applications, 153:183–192, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Valiaho:1992:DSW

[Val92] H. Valiaho. Determining subspaces on which a matrix is nonnegative definite. Linear Algebra and its Applications, 168:71–??, April 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Valiaho:1993:BPO

[Väl93] H. Väliaho. The Boolean pivot operation, M-matrices, and reducible matrices. Linear Algebra and its Applications, 185:143–163, May 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Valiaho:1996:CSM

[Väl96a]  H. Väliaho.  Criteria for

sufficient matrices. Linear Algebra and its Applications, 233(1–3):109–129, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400058.

Valiaho:1996:MJS

[Väl96b] Hannu Väliaho. P_{∗}-matrices are just sufficient. Linear Algebra and its Applications, 239(1–3):103–108, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=239&issue=1-
3&aid=9400170.

Valiaho:1997:DHS

[Väl97] H. Väliaho. Determining the handicap of a sufficient matrix. Linear Algebra and its Applications, 253(1–3):279– 298, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500703.

VanDooren:1990:RPM

[Van90] P. Van Dooren. Rational and polynomial matrix factorizations via recursive pole-zero cancellation. Linear Algebra and its Applications, 137/138: 663–697, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderWoude:1991:SIS

[van91a] J. W. van der Woude. On the structure at infinity of a structured system. Linear Algebra and its Applications, 148: 145–169, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanHuffel:1991:IAC

[Van91b] Sabine Van Huffel. Iterative algorithms for computing the singular subspace of a matrix associated with its smallest singular values. Linear Algebra and its Applications, 154/156:675–709, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vanderbei:1991:SDC

[Van91c] Robert J. Vanderbei. Splitting dense columns in sparse linear systems. Linear Algebra and its Applications, 152: 107–117, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

vanderWoude:1994:GTM

[van94] J. W. van der Woude. Graph theoretic methods for the computation of disturbance decoupling feedback matrices for structured systems. Linear Algebra and its Applications, 196:139–162, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:1995:RGF

[van95a] Edwin R. van Dam. Regular graphs with four eigenvalues. Linear Algebra and its Applications, 226/228 (1–3):139–162, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9400346.

vandenHeuvel:1995:HCE

[van95b] J. van den Heuvel. Hamilton cycles and eigenvalues of graphs. Linear Algebra and its Applications, 226/228 (1–3):723–730, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500254.

vanderBlij:1995:RPE

[van95c]  F. van der Blij.  Regu-

lar polygons in Euclidean space. Linear Algebra and its Applications, 226/228 (1–3):345–352, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500148.

vanLint:1995:P

[van95d] J. H. van Lint. Preface.

Linear Algebra and its Applications, 226/228(1–3):1– 8, September/October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanWyk:1996:MRS

[van96] Leon van Wyk. Matrix rings satisfying column sum conditions versus structural matrix rings. Linear Algebra and its Applications, 249(1–3):15– 28, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500241.

vandenHof:1997:RPL

[van97a] J. M. van den Hof. Realization of positive linear systems. Linear Algebra and its Applications, 256(1–3):287– 308, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9600030.

VanDerGenugten:1997:TRL

[Van97b] B. B. Van Der Genugten. Testing in the restricted linear model using canonical partitions. Linear Algebra and its Applications, 264 (1–3):349–353, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9600150.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9600150.

vanLeijenhorst:1998:HSC

[van98] D. C. van Leijenhorst. How symmetric can a function be? Linear Algebra and its Applications, 281(1–3): 1–10, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6149.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6149.

Varah:1993:PM

[Var93] J. M. Varah. The prolate matrix. Linear Algebra and its Applications, 187:269–278, July 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Varro:1996:DTA

[Var96] Richard Varro. Dupliquée d’une train algèbre et d’une algèbre de Bernstein périodique. (French) [duplicate of a train algebra and of a periodic Bernstein algebra]. Linear Algebra and its Applications, 233(1–3):161–165, January 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=233&issue=1-
                                        3&aid=9400064.

Varga:1998:CCF

[Var98] A. Varga. Computation of coprime factorizations of rational matrices. Linear Algebra and its Applications, 271(1–3): 83–115, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=271&issue=1-
                                        3&aid=9700256.

Vavrin:1991:UAL

[Vav91] Zdeněk Vavřín. A unified approach to Loewner and Hankel matrices. Linear Algebra and its Applications, 143:171– 222, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vavrin:1995:MFV

[Vav95]  Zdeněk Vavřín.  Miroslav

Fiedler and Vlastimil Pták: life and work. Linear Algebra and its Applications, 223/224 (1–3):3–29, July 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue honoring Miroslav Fiedler and Vlastimil Pták.

Vavrin:1997:CCC

[Vav97]  Zdeněk Vavřín.  Conflu-

ent Cauchy and Cauchy– Vandermonde matrices. Linear Algebra and its Applications, 258(1–3):271–293, June ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600200.

VanBarel:1997:GMP

[VB97a] Marc Van Barel and Adhemar Bultheel. A generalized minimal partial realization problem. Linear Algebra and its Applications, 254(1–3): 527–551, March 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=254&issue=1-
3&aid=9600528. Proceedings

of the Fifth Conference of the International Linear Algebra Society (Atlanta, GA, 1995).

VanBarel:1997:LAS

[VB97b] Marc Van Barel and Adhemar Bultheel. A lookahead algorithm for the solution of block Toeplitz systems.

Linear Algebra and its Applications, 266(1–3):291–335, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9700013.

vanderVeen:1994:LCA

[vD94] Alle-Jan van der Veen and Patrick Dewilde. On lowcomplexity approximation of matrices. Linear Algebra and its Applications, 205/206: 1145–1201, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:1998:CKM

[vD98]  Edwin R. van Dam.  A

Cauchy–Khinchin matrix inequality. Linear Algebra and its Applications, 280(2–3): 163–172, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6117.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6117.

vanderMee:1999:SSA

[vdMRR99] Cornelis V. M. van der Mee, André C. M. Ran, and Leiba Rodman. Stability of selfadjoint square roots and polar decompositions in indefinite scalar product spaces. Linear Algebra and its Applications, 302-303(1–3):77–104, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/31/article.
pdf.

Veljan:1995:STI

[Vel95] Darko Veljan. The sine theorem and inequalities for volumes of simplices and determinants. Linear Algebra and its Applications, 219(1– 3):79–91, April 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=219&issue=1-
3&aid=9300196.

Velasco:1999:SSM

[Vel99] Francisco E. Velasco. Stable subspaces of matrix pairs. Linear Algebra and its Applications, 301(1–3):15–49, November 1, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
26/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/115/25/26/article.
                                        pdf.

Veselic:1998:EOE

[Ves98] K. Veselić. Estimating the operator exponential. Linear Algebra and its Applications, 280(2–3):241–244, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        280/2-3/6147.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=280&issue=
                                        2-3&aid=6147.

vandenEssen:1996:PMS

[vH96] Arno van den Essen and Engelbert Hubbers. Polynomial maps with strongly nilpotent Jacobian matrix and the Jacobian conjecture. Linear Algebra and its Applications, 247(1–3):121–132, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500095.

Villarreal:1998:EDM

[Vil98] César E. Villarreal. Ergodic decomposition of Markov chains. Linear Algebra and its Applications, 283(1–3):

61–73, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6199.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6199.

Vinnikov:1990:ETD

[Vin90] Victor Vinnikov. Elementary transformations of determinantal representations of algebraic curves. Linear Algebra and its Applications, 135: 1–18, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Visick:1995:WMI

[Vis95] George Visick. A weak majorization involving the matrices A ∘ B and AB. Linear Algebra and its Applications, 223/224(1–3):731– 744, July ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=223-224&issue=
1-3&aid=9400009. Special is-

sue honoring Miroslav Fiedler and Vlastimil Pták.

Visick:1998:MHP

[Vis98]  George Visick.  Majoriza-

tions of Hadamard products of matrix powers. Linear Algebra and its Applications, 269(1):233–240, January 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700068.

VanBarel:1998:SSS

[VK98] Marc Van Barel and Peter Kravanja. A stabilized superfast solver for indefinite Hankel systems. Linear Algebra and its Applications, 284 (1–3):335–355, November 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
284/1-3/6197.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=284&issue=
1-3&aid=6197.

Volaufova:1997:NIQ

[VL97] Júlia Volaufová and Lynn Roy LaMotte. A note on invariant quadratics. Linear Algebra and its Applications, 264(1–3):247–253, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700043.pdf;

 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700043.

vanderHolst:1995:ICV

[vLS95] Hein van der Holst, László Lovász, and Alexander Schrijver. On the invariance of Colin de Verdiere’s graph parameter under clique sums. Linear Algebra and its Applications, 226/228(1–3):509– 517, September/October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.elsevier.com/cgi-bin/
cas/tree/store/laa/cas_
sub/browse/browse.cgi?year=
1995&volume=226-228&issue=
1-3&aid=9500160.

Volkmer:1999:CLE

[Vol99] Hans Volkmer. Convexity of the Lyapunov exponent. Linear Algebra and its Applications, 294(1–3):35–48, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6391.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6391.

vonRosen:1990:MFT

[von90] Dietrich von Rosen. A matrix formula for testing linear hypotheses in linear models. Linear Algebra and its Applications, 127:457–461, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vonzurGathen:1991:MBC

[von91] Joachim von zur Gathen. Maximal bilinear complexity and codes. Linear Algebra and its Applications, 144:49– 61, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VonRosen:1993:HME

[Von93] Dietrich Von Rosen. Homogeneous matrix equations and multivariate linear models. Linear Algebra and its Applications, 193:19–33, November 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Voss:1999:SSC

[Vos99] Heinrich Voss. Symmetric schemes for computing the minimum eigenvalue of a symmetric Toeplitz matrix. Linear Algebra and its Applications, 287(1–3):359–371, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        287/1-3/6258.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?

year=1999&volume=287&issue=
1-3&aid=6258.

VanDenEssen:1992:PMC

[VP92] Arno Van Den Essen and T. Parthasarathy. Polynomial maps and a conjecture of Samuelson. Linear Algebra and its Applications, 177: 191–??, December 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanBarel:1999:ENC

[VPV99] Marc Van Barel, Vlastimil Pták, and Zdeněk Vavřín. Extending the notions of companion and infinite companion to matrix polynomials. Linear Algebra and its Applications, 290(1–3):61–94, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6311.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6311.

Vroegindeweij:1993:ASS

[Vro93] P. G. Vroegindeweij. Algebraic structures on spacetime algebra. Linear Algebra and its Applications, 178: 217–??, January 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Veselic:1993:FPP

[VS93] Krešimir Veselić and Ivan Slapničar. Floating-point perturbations of Hermitian matrices. Linear Algebra and its Applications, 195:81–116, December 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanDooren:1994:WPD

[VS94] Paul Van Dooren and J. Sreedhar. When is a periodic discrete-time system equivalent to a time-invariant one? Linear Algebra and its Applications, 212/213:131–151, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Vassiliou:1999:PNM

[VS99] P.-C. G. Vassiliou and M. A. Symeonaki. The perturbed nonhomogeneous Markov system. Linear Algebra and its Applications, 289(1–3): 319–332, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        289/1-3/6146.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=289&issue=
                                        1-3&aid=6146.

Vardy:1990:BDC

[VSB90] Alexander Vardy, Jakov Snyders, and Yair Be’ery. Bounds on the dimension of codes and subcodes with prescribed contraction index. Linear Algebra and its Applications, 142: 237–261, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vigneron-Tenorio:1999:SIL

[VT99] A. Vigneron-Tenorio. Semigroup ideals and linear diophantine equations. Linear Algebra and its Applications, 295(1–3):133–144, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6433.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6433.

vanderSluis:1990:SCT

[vV90] A. van der Sluis and H. A.

Van der Vorst. SIRT- and CG-type methods for the iterative solution of sparse linear least-squares problems. Linear Algebra and its Applications, 130:257–303, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Vuik:1992:CSG

[VV92] C. Vuik and H. A. Van der Vorst. A comparison of some GMRES-like methods. Linear Algebra and its Applications, 160:131–162, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vandenHof:1999:PMF

[vvS99] Jacqueline M. van den Hof and Jan H. van Schuppen. Positive matrix factorization via extremal polyhedral cones. Linear Algebra and its Applications, 293(1–3):171–186, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        293/1-3/6386.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=293&issue=
                                        1-3&aid=6386.

Vaserstein:1990:CCM

[VW90] L. N. Vaserstein and E. Wheland. Commutators and companion matrices over rings of stable rank 1. Linear Algebra and its Applications, 142: 263–277, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vaserstein:1995:PCC

[VW95] L. Vaserstein and E. Wheland. Products of conjugacy classes of two by two matrices.

Linear Algebra and its Applications, 230(1–3):165–188, November 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=230&issue=1-
3&aid=9400005.

Vaserstein:1998:RR

[VW98] L. Vaserstein and E. Wheland. Rigid relations in GL_{2}F. Linear Algebra and its Applications, 281(1–3): 25–31, September 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
281/1-3/6143.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=281&issue=
1-3&aid=6143.

Wai:1997:CPF

[Wai97] Ki Ching Wai. Circulant preconditioners for failure prone manufacturing systems. Linear Algebra and its Applications, 266(1–3):161–180, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

                                        1997&volume=266&issue=1-
                                        3&aid=9700001.

Wang:1991:CPM

[Wan91] De Ren Wang. On the convergence of the parallel multisplitting AOR algorithm. Linear Algebra and its Applications, 154/156:473–486, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1992:CEM

[Wan92] Shougen Wang. Computing the eigenvalues of modified symmetric matrices. Linear Algebra and its Applications, 168:57–??, April 15, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wan:1993:OISb

[Wan93a] Zhe-Xian Wan. On the orthogonal invariants of a subspace of a vector space over a finite field of even characteristics. Linear Algebra and its Applications, 184:135–??, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wan:1993:OISa

[Wan93b] Zhe-Xian Wan. On the orthogonal invariants of a subspace of a vector space over a finite field of odd characteristic. Linear Algebra and its Applications, 184:123–??, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1994:CGF

[Wan94] Xin Min Wang. Convergence for a general form of the GAOR method and its application to the MSOR method. Linear Algebra and its Applications, 196:105–123, January 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1995:FMQ

[Wan95a]  Jin Hsien Wang.  Fac-

torization of matrices into quadratic ones. II. Linear Algebra and its Applications, 231(1–3):111–152, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400016.

Wang:1995:LPS

[Wan95b] Jin Hsien Wang. The length problem for a sum of idempotents. Linear Algebra and its Applications, 215(1–3):135– 159, January 15, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=215&issue=1-
3&aid=9300083.

Wang:1995:SPT

[Wan95c] Jin Hsien Wang. Sums and products of two quadratic matrices. Linear Algebra and its Applications, 229(1–3):127– 149, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9300349.

Wang:1996:FMQ

[Wan96a] Jin-Hsien Wang. Factorization of matrices into quadratic ones. III. Linear Algebra and its Applications, 240(1– 3):21–39, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400180.

Wang:1996:PIO

[Wan96b]  Jin-Hsien Wang.  Prod-

uct of invertible operators of quadratic type. Linear Algebra and its Applications, 245(1–3):1–26, September 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=245&issue=1-
3&aid=9400200.

Wang:1997:WMP

[Wan97a] Guorong Wang. Weighted Moore–Penrose, Drazin, and group inverses of the Kronecker product A ⊗ B, and some applications. Linear Algebra and its Applications, 250(1–3):39–50, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500427.

Wang:1997:CTG

[Wan97b]  Xinmin Wang.  Conver-

gence theory for the general GAOR type iterative method and the MSOR iterative method applied to Hmatrices. Linear Algebra and its Applications, 250(1–3):1– 19, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500391.

Wang:1998:TI

[Wan98a] Derming Wang. On a trace inequality. Linear Algebra and its Applications, 273(1– 3):1–8, April 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=273&issue=1-
3&aid=9600520.

Wang:1998:CTC

[Wan98b] Xinmin Wang. Comparison theorems for a class of parallel multisplitting AOR type iterative methods. Linear Algebra and its Applications, 269(1): 1–16, January 15, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1998&volume=269&issue=1&
aid=9700008.

Wanless:1999:HDC

[Wan99] I. M. Wanless. The Holens- – Doković Conjecture on permanents fails! Linear Algebra and its Applications, 286(1–3):273–285, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6287.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6287.

Warren:1990:CMT

[War90] Richard H. Warren. Classes of matrices for the traveling salesman problem. Linear Algebra and its Applications, 139:53–62, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watson:1997:SCS

[WAS97] Geoffrey S. Watson, Gulhan Alpargu, and George P. H. Styan. Some comments on six inequalities associated with the inefficiency of ordinary least squares with one regressor. Linear Algebra and its Applications, 264(1–3):13–54, October ??, 1997. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700228.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700228.

Waterhouse:1990:AVE

[Wat90a] William C. Waterhouse. The absolute-value estimate for symmetric multilinear forms. Linear Algebra and its Applications, 128:97–105, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:1990:CNQ

[Wat90b] William C. Waterhouse. Classifying nonsingular quadratic forms of rank three. Linear Algebra and its Applications, 142:55–61, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:1992:DOR

[Wat92a] William C. Waterhouse. Displacement operators relative to group matrices. Linear Algebra and its Applications, 169:41–??, May 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watson:1992:CSS

[Wat92b] G. A. Watson. Characterization of the subdifferential of some matrix norms. Linear Algebra and its Applications, 170:33–??, June 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:1993:DCE

[Wat93a] William C. Waterhouse. The degrees of the cyclotomic extension fields. Linear Algebra and its Applications, 195: 181–189, December 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watkins:1993:CPG

[Wat93b] William Watkins. The cone of positive generalized matrix functions. Linear Algebra and its Applications, 181:1–??, March 1, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watkins:1994:UCL

[Wat94] William Watkins. Unimodular congruence of the Laplacian matrix of a graph. Linear Algebra and its Applications, 201:43–49, April 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:1995:OSP

[Wat95a] William C. Waterhouse. Orthogonal similarity and pairs of quadratic forms. Linear Algebra and its Applications, 231(1–3):175–179, December ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=231&issue=1-
3&aid=9400023.

Watson:1995:EHO

[Wat95b] G. A. Watson. Estimating Hadamard operator norms, with application to triangular truncation. Linear Algebra and its Applications, 216(1–3):97–110, February ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=216&issue=1-
3&aid=9300118.

Waters:1996:SPM

[Wat96a] Charles Waters. Sign pattern matrices that allow orthogonality. Linear Algebra and its Applications, 235(1– 3):1–13, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400098.

Watkins:1996:TSS

[Wat96b] David S. Watkins. The transmission of shifts and shift blurring in the QR algorithm. Linear Algebra and its Applications, 241/243(1–3):877– 896, July/August 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=241-243&issue=
1-3&aid=9500545. Proceed-

ings of the Fourth Conference of the International Linear Algebra Society (Rotterdam, 1994).

Watson:1996:CNR

[Wat96c] G. A. Watson. Computing the numerical radius.

Linear Algebra and its Applications, 234(1–3):163–172, February ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=234&issue=1-
3&aid=9400097.

Waterhouse:1997:ENM

[Wat97a] William C. Waterhouse. Eigenmatrices of a normal matrix. Linear Algebra and its Applications, 255(1–3):243– 246, April 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=255&issue=1-
3&aid=9600168.

Watkins:1997:BTM

[Wat97b]  William Watkins.  Bilin-

ear transformations on matrices: Rank preservers. Linear Algebra and its Applications, 250(1–3):31–38, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=250&issue=1-
3&aid=9500426.

Wang:1997:PMI

[WB97] De-Ren Wang and Zhong-Zhi Bai. Parallel multilevel iterative methods. Linear Algebra and its Applications, 250(1–3):317–347, January 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=250&issue=1-
                                        3&aid=9500547.

Wu:1996:RND

[WC96] Lei Wu and Bryan Cain. The Re-nonnegative definite solutions to the matrix inverse problem AX = B. Linear Algebra and its Applications, 236(1–3):137–146, March 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400142.

Wong:1997:VMN

[WC97] Chi Song Wong and Hua Cheng. Vector majorization via nonnegative definite doubly stochastic matrices of maximum rank. Linear Algebra and its Applications, 261(1–3):187–194, August ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=261&issue=1-
3&aid=9600399.

Wong:1999:MVC

[WCM99] Chi Song Wong, Hua Cheng, and Joe Masaro. Multivariate versions of Cochran theorems. Linear Algebra and its Applications, 291(1– 3):227–234, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6342.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6342.

Watkins:1991:CAD

[WE91] D. S. Watkins and L. Elsner. Convergence of algorithms of decomposition type for the eigenvalue problem. Linear Algebra and its Applications, 143:19–47, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:1990:EEP

[Wei90a] Mu Sheng Wei. On the error estimate for the projection of a point onto a linear manifold. Linear Algebra and its Applications, 133:53–75, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:1990:PLS

[Wei90b] Mu Sheng Wei. Perturbation of the least squares problem. Linear Algebra and its Applications, 141:177–182, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Weiland:1994:BAB

[Wei94] Siep Weiland. A behavioral approach to balanced representations of dynamical systems. Linear Algebra and its Applications, 205/206:1227– 1252, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:1997:ECG

[Wei97a] Musheng Wei. Equivalent conditions for generalized inverses of products. Linear Algebra and its Applications, 266(1–3):347–363, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=266&issue=1-
                                        3&aid=9700035.

Weissenhofer:1997:PMP

[Wei97b] S. Weissenhofer. The positive minorant property on matrices. Linear Algebra and its Applications, 258(1–3):81– 94, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600152.

Wei:1998:PTE

[Wei98]  Musheng Wei.  Perturba-

tion theory for the Eckart– Young–Mirsky theorem and the constrained total least squares problem. Linear Algebra and its Applications, 280 (2–3):267–287, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
280/2-3/6139.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=280&issue=
2-3&aid=6139.

Wei:1999:ROL

[Wei99] Musheng Wei. Reverse order laws for generalized inverses of multiple matrix products. Linear Algebra and its Applications, 293(1– 3):273–288, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6401.pdf;   http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=293&issue=
                                        1-3&aid=6401.

Wen:1996:TMF

[Wen96]  Zhi-Xiong Wen.  On the

transfer matrices of fractals. Linear Algebra and its Applications, 236(1–3):189– 203, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=236&issue=1-
                                        3&aid=9400150.

Werner:1990:CFF

[Wer90a] Hans Joachim Werner. A closed form formula for the intersection of two complex matrices under the star order. Linear Algebra and its Applications, 140:13–30, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Werner:1990:ICG

[Wer90b] Hans Joachim Werner. On inequality constrained generalized least-squares estimation. Linear Algebra and its Applications, 127:379–392, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Werner:1991:SFR

[Wer91] Hans Joachim Werner. Some further results on matrix monotonicity. Linear Algebra and its Applications, 150: 371–392, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Werner:1994:WGI

[Wer94] Hans Joachim Werner. When is B^{−}A^{−} a generalized inverse of AB? Linear Algebra and its Applications, 210: 255–263, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wermuth:1996:REH

[Wer96] Edgar M. E. Wermuth. A remark on equidistance in Hilbert spaces. Linear Algebra and its Applications, 236(1–3): 105–111, March 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400131.

Westwick:1996:SMF

[Wes96]  R. Westwick.  Spaces of

matrices of fixed rank. II. Linear Algebra and its Applications, 235(1–3):163–169, March 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400134.

Willems:1992:STH

[WF92] J. C. Willems and P. A. Fuhrmann. Stability theory for high order equations. Linear Algebra and its Applications, 167:131–??, April 1, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1993:SEI

[WG93] Bo-Ying Wang and Ming-Peng Gong. Some eigenvalue inequalities for positive semidefinite matrix power products. Linear Algebra and its Applications, 184:249–??, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1999:MVW

[WI99] Song-Gui Wang and Wai-Cheung Ip. A matrix version of the Wielandt inequality and its applications to statistics. Linear Algebra and its Applications, 296(1–3):171–181, July 15, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/47/17/26/
abstract.html;   http://

www.elsevier.nl/gej-ng/
10/30/19/47/17/26/article.
pdf.

Wiens:1990:RMD

[Wie90] Douglas P. Wiens. Robust minimax designs for multiple linear regression. Linear Algebra and its Applications, 127: 327–340, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wiener:1997:IME

[Wie97] Zvi Wiener. An interesting matrix exponent formula. Linear Algebra and its Applications, 257(1–3):307– 310, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=257&issue=1-
3&aid=9600194.

Wilde:1993:NCP

[Wil93a] Alan C. Wilde. Nearly commuting projections. Linear Algebra and its Applications, 181:73–??, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wilker:1993:QFM

[Wil93b] J. B. Wilker. The quaternion formalism for mobius groups in four or fewer dimensions. Linear Algebra and its Applications, 190:99–??, September 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1990:HBR

[Wim90] Harald K. Wimmer. On the history of the Bezoutian and the resultant matrix. Linear Algebra and its Applications, 128:27–34, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1991:NFS

[Wim91] Harald K. Wimmer. Normal forms of symplectic pencils and the discrete-time algebraic Riccati equation. Linear Algebra and its Applications, 147:411–440, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1993:PCM

[Wim93] Harald K. Wimmer. Pairs of companion matrices and their simultaneous reduction to complementary triangluar forms. Linear Algebra and its Applications, 182:179– ??, March 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1994:RTM

[Wim94a] H. K. Wimmer. Roth’s theorems for matrix equations with symmetry constraints. Linear Algebra and its Applications, 199:357–362, March 1, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1994:GCB

[Wim94b] Harald K. Wimmer. A Galois correspondence between sets of semidefinite solutions of continuous-time algebraic Riccati equations. Linear Algebra and its Applications, 205/206:1253–1270, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:1999:CAU

[Wim99] Harald K. Wimmer. Canonical angles of unitary spaces and perturbations of direct complements. Linear Algebra and its Applications, 287 (1–3):373–379, January 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
287/1-3/6138.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=287&issue=
1-3&aid=6138.

Winterhof:1999:PSF

[Win99] Arne Winterhof. Polynomial spaces over finite fields. Linear Algebra and its Applications, 295(1–3):223–229, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        295/1-3/6447.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=295&issue=
                                        1-3&aid=6447.

Wang:1993:NEL

[WL93] Guorong Wang and Yuhua Lin. A new extension of Leverrier’s algorithm. Linear Algebra and its Applications, 180: 227–??, February 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1994:EPN

[WL94] D. Q. Wang and C. R. O. Lawoko. Estimation of parameters for normally distributed random matrices. Linear Algebra and its Applications, 210: 199–208, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1998:ME

[WL98] Yaokun Wu and Qiao Li. On the matrix equation A^{l} + A^{l+k} = J_{n}. Linear Algebra and its Applications, 277 (1–3):41–48, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        277/1-3/6062.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/

cas_sub/browse/browse.cgi?
year=1998&volume=277&issue=
1-3&aid=6062.

Wu:1999:ME

[WLH99] Yaokun Wu, Qiao Li, and Xuyan Huang. On the matrix equation A^{k} = J − I. Linear Algebra and its Applications, 295(1–3):249–260, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6370.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6370.

Wu:1999:SEC

[WLL99] Yaokun Wu, Qiao Li, and Jingjie Li. Spectra and elementary cycles of the digraphs with unique paths of fixed length. Linear Algebra and its Applications, 293 (1–3):145–158, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6376.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6376.

Wang:1999:TWS

[WLN99] Song-Gui Wang, Erkki P.

Liski, and Tapio Nummi. Two-way selection of covariables in multivariate growth curve models. Linear Algebra and its Applications, 289(1–3):333–342, March 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        289/1-3/6270.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=289&issue=
                                        1-3&aid=6270.

Wong:1990:NRP

[WM90] Yung Chow Wong and Kam Ping Mok. Normally related nplanes and isoclinic n-planes in R^{2n} and strongly linearly independent matrices of order n. Linear Algebra and its Applications, 139:31–52, 1990. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1994:SHS

[WM94] Q. H. Wu and M. Mansour. On the stability hyperellipsoid of a Schur polynomial. Linear Algebra and its Applications, 205/206:1271–1288, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wong:1995:ECG

[WMD95] Chi Song Wong, Joe Masaro, and Wei Cai Deng. Estimating covariance in a growth curve model. Linear Algebra and its Applications, 214(1–3):103– 118, January 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=214&issue=1-
3&aid=9300057.

Wu:1997:CHC

[WmL97] Zhi-Ming Wu and Wei min Liu. Criterion of highcodimensional bifurcations with several pairs of purely imaginary eigenvalues. Linear Algebra and its Applications, 267(1–3):53–63, December ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=267&issue=1-
3&aid=9600643.

Wei:1998:CRG

[WMP98] Yimin Wei, J. Maroulas, and P. Psarrakos. A characterization and representation of the generalized inverse A_{T,S}^{(2)} and its applications. Linear Algebra and its Applications, 280(2–3):87–96, September 4, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/

                                        280/2-3/6106.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=280&issue=
                                        2-3&aid=6106.

Woo:1995:GWS

[WN95] Renee Woo and Arnold Neumaier. On graphs whose smallest eigenvalue is at least −1 − \sqrt{2}. Linear Algebra and its Applications, 226/228 (1–3):577–591, September/ October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1995&volume=226-228&issue=
                                        1-3&aid=9500245.

Woerdeman:1994:TMR

[Woe94] Hugo J. Woerdeman. Toeplitz minimal rank completions. Linear Algebra and its Applications, 202:267–278, April 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wolkowicz:1996:ESI

[Wol96] Henry Wolkowicz. Explicit solutions for interval semidefinite linear programs. Linear Algebra and its Applications, 236(1–3):95–104, March 15, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=236&issue=1-
3&aid=9400130.

Wong:1990:ECM

[Won90] Chi Song Wong. Extensions of tr[(A^{−1}−B^{−1})(A−B)] ≤ 0 for covariance matrices A, B. Linear Algebra and its Applications, 127:517–529, 1990. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wood:1993:PRD

[Woo93] David H. Wood. Product rules for the displacement of Near-Toeplitz matrices. Linear Algebra and its Applications, 188/189:641–663, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Woodgate:1996:LSS

[Woo96] Keith G. Woodgate. Leastsquares solution of F = P G over positive semidefinite symmetric P. Linear Algebra and its Applications, 245(1–3):171–190, September 15, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=245&issue=1-
3&aid=9400238.

Woracek:1997:NPI

[Wor97] Harald Woracek. Nevanlinna– Pick interpolation: The degenerated case. Linear Algebra and its Applications, 252(1–3):141–158, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500700.

Woznicki:1997:CTR

[Woź97] Zbigniew I. Woźnicki. Comparison theorems for regular splittings on block partitions. Linear Algebra and its Applications, 253(1–3):199– 207, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500702.

Wei:1998:PAC

[WP98] Musheng Wei and Alvaro R. De Pierro. Perturbation analysis of the canonical subspaces. Linear Algebra and its Applications, 279(1–3):135–151, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/

tree/store/laa/sub/1998/
279/1-3/6102.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=279&issue=
1-3&aid=6102.

Wang:1994:CSS

[WR94] Xiaochang Wang and Joachim Rosenthal. A cell structure for the set of autoregressive systems. Linear Algebra and its Applications, 205/206:1203– 1225, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wyman:1991:PZM

[WSCP91] Bostwick F. Wyman, Michael K.

Sain, Giuseppe Conte, and Anna-Maria Perdon. Poles and zeros of matrices of rational functions. Linear Algebra and its Applications, 157: 113–139, 1991. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1990:SIM

[Wu90] Pei Yuan Wu. Sums of idempotent matrices. Linear Algebra and its Applications, 142: 43–54, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1992:RPD

[Wu92a] Lei Wu. The re-positive definite solutions to the matrix inverse problem AX = B. Linear Algebra and its Applications, 174:145–??, September 1992. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1992:WLT

[Wu92b] Pei Yuan Wu. Which linear transformations have isomorphic hyperinvariant subspace lattices? Linear Algebra and its Applications, 169:163–??, May 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1994:SPL

[Wu94] Bao Wei Wu. Some properties of linear systems defined over a commutative Banach algebra. Linear Algebra and its Applications, 203/204:139– 154, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:1997:POD

[Wu97]  Jun Wu.  A partial or-

dering of CS_{m} defined by the generalized matrix functions. Linear Algebra and its Applications, 257(1–3):193– 200, May ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1997&volume=257&issue=1-
                                        3&aid=9600204.

Wuwen:1997:ENM

[Wuw97] Guo Wuwen. Eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 266(1–3):261–270, November 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=266&issue=1-
3&aid=9600007.

Wang:1991:PUM

[WW91] Jin Hsien Wang and Pei Yuan Wu. Products of unipotent matrices of index 2. Linear Algebra and its Applications, 149:111–123, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1994:DSM

[WW94] Jin Hsien Wang and Pei Yuan Wu. Difference and similarity models of two idempotent operators. Linear Algebra and its Applications, 208/209:257– 282, 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wainberg:1996:MRL

[WW96] Bella I. Wainberg and Hugo J.

Woerdeman. The maximum row length nonsingularity radius. Linear Algebra and its Applications, 247(1–3):251– 263, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=247&issue=1-
3&aid=9500106.

Wei:1997:PTD

[WW97] Yimin Wei and Guorong Wang. The perturbation theory for the Drazin inverse and its applications. Linear Algebra and its Applications, 258(1–3):179–186, June ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600159.

Wang:1999:INN

[wWS99] Ming wei Wang and Jeffrey Shallit. An inequality for non-negative matrices. Linear Algebra and its Applications, 290(1–3):135–144, March 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6313.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6313.

Wang:1997:SIS

[WX97] Bo-Ying Wang and Bo-Yan Xi. Some inequalities for singular values of matrix products. Linear Algebra and its Applications, 264 (1–3):109–115, October ??, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1997/264/1-3/9700020.pdf;
 http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1997&volume=264&
issue=1-3&aid=9700020.

Wang:1999:SIS

[WXZ99] Bo-Ying Wang, Bo-Yan Xi, and Fuzhen Zhang. Some inequalities for sum and product of positive semidefinite matrices. Linear Algebra and its Applications, 293 (1–3):39–49, May 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
293/1-3/6362.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=293&issue=
1-3&aid=6362.

Werner:1996:ICG

[WY96a] H. J. Werner and C. Yapar. On inequality constrained generalized least squares selections in the general possibly singular Gauss–Markov model: A projector theoretical approach. Linear Algebra and its Applications, 237/238(1–3):359– 393, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9400357. Special

issue honoring Calyampudi Radhakrishna Rao.

Werner:1996:BDG

[WY96b] Hans Joachim Werner and Cemil Yapar. A BLUE decomposition in the general linear regression model. Linear Algebra and its Applications, 237/238(1–3):395– 404, April ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=237-238&issue=
                                        1-3&aid=9500542. Special

issue honoring Calyampudi Radhakrishna Rao.

Wang:1998:SEH

[WY98] Lie Heng Wang and Ya Xiang Yuan. A stronger extension of the Hardy inequality. Linear Algebra and its Applications, 270(1–3):275– 286, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=

1998&volume=270&issue=1-
3&aid=9700245.

Wyels:1997:PEH

[Wye97] Cynthia J. Wyels. Permutation equivalence and the Hermite invariant. Linear Algebra and its Applications, 256(1–3): 125–140, April 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=256&issue=1-
3&aid=9781114.

Wang:1998:NMZ

[WZ98] Bo-Ying Wang and Fuzhen Zhang. On normal matrices of zeros and ones with fixed row sum. Linear Algebra and its Applications, 275/276(1–3):617–626, May 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
275-276/1-3/6014.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1998&volume=275-
276&issue=1-3&aid=6014. Pro-

ceedings of the Sixth Conference of the International Linear Algebra Society (Chemnitz, 1996).

Wang:1999:IUD

[WZ99] Weiguo Wang and Jinxi Zhao. Inverse updating and downdating for weighted linear least squares using Minvariant reflections. Linear Algebra and its Applications, 291(1–3):185–199, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6358.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6358.

Wang:1993:SIM

[WZCZ93] Cheng Wang, Dao Sheng Zheng, Guo-Liang Chen, and Shu Qin Zhao. Structures of p-isometric matrices and rectangular matrices with minimum p-norm condition number. Linear Algebra and its Applications, 184:261–278, April 15, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:1999:SIG

[WZZ99] Bo-Ying Wang, Xiuping Zhang, and Fuzhen Zhang. Some inequalities on generalized Schur complements. Linear Algebra and its Applications, 302-303(1–3):163–172, December 1, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
35/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/30/19/117/25/35/article.
pdf.

Xu:1990:MPA

[XB90] Guo Liang Xu and Adhemar Bultheel. Matrix Padé approximation: definitions and properties. Linear Algebra and its Applications, 137/138:67– 136, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xie:1997:RBU

[Xie97a]  Jian-Jun Xie.  Residual

bounds for unitarily invariant norms on clustered eigenvalues. Linear Algebra and its Applications, 252(1–3):107– 113, February ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=252&issue=1-
3&aid=9500677.

Xie:1997:NDK

[Xie97b] Jianjun Xie. A note on the Davis-Kahan sin Θ theorem. Linear Algebra and its Applications, 258(1–3):129– 135, June ??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=258&issue=1-
3&aid=9600207.

Xinmin:1993:GEP

[Xin93a] Wang Xinmin. Generalized extrapolation principle and convergence of some generalized iterative methods. Linear Algebra and its Applications, 185:235–??, May 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xinmin:1993:GSR

[Xin93b] Wang Xinmin. Generalized Stein–Rosenberg theorems for the regular splittings and convergence of some generalized iterative methods. Linear Algebra and its Applications, 184:207–??, April 15, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xin:1994:LFD

[XK94] Xin Xin and Hidenori Kimura. (J, J′)-lossless factorization for descriptor systems. Linear Algebra and its Applications, 205/206:1289–1318, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:1995:PQM

[XL95] Hong Guo Xu and Lin Zhang Lu. Properties of a quadratic matrix equation and the solution of the continuoustime algebraic Riccati equation. Linear Algebra and its Applications, 222(1–3):127– 145, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300292.

Xu:1991:AIQ

[Xu91] Zhong Xu. An algorithm for the inversion of quasihessenberg matrices. Linear Algebra and its Applications, 144:39–47, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:1992:MPI

[Xu92] Zhong Xu. On Moore–Penrose inverses of Toeplitz matrices. Linear Algebra and its Applications, 169:9–??, May 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:1993:NSM

[Xu93] Song Xu. Notes on sufficient matrices. Linear Algebra and its Applications, 191:1– 13, September 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:1995:BI

[Xu95] Chang Qin Xu. Bellman’s inequality. Linear Algebra and its Applications, 229(1–3):9– 14, October 1, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=229&issue=1-
3&aid=9300331.

Xu:1997:TRA

[Xu97a] Hongguo Xu. Two results about the matrix exponential. Linear Algebra and its Applications, 262(1–3):99– 109, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600475.

Xu:1997:LBE

[Xu97b] Shu-Fang Xu. Lower bound estimation for the separation of two matrices. Linear Algebra and its Applications, 262(1–3):67–82, September 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=262&issue=1-
3&aid=9600468.

Xu:1999:LUS

[Xu99]  Song Xu.  On local w-

uniqueness of solutions to linear complementarity problem. Linear Algebra and its Applications, 290(1–3):23– 29, March 15, 1999. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
290/1-3/6231.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=290&issue=
1-3&aid=6231.

Xue:1997:NEP

[Xue97] Jungong Xue. A note on entrywise perturbation theory for Markov chains. Linear Algebra and its Applications, 260(1–3):209–213, July 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=260&issue=1-
3&aid=9600302.

Xu:1998:SME

[XWZ98] Guiping Xu, Musheng Wei, and Daosheng Zheng. On solutions of matrix equation AXB + CY D = F. Linear Algebra and its Applications, 279(1–3):93–109, August 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
279/1-3/6099.pdf;   http:
//www.elsevier.com/cgi-

                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=279&issue=
                                        1-3&aid=6099.

Xiang:1998:NCP

[XX98] Shuhuang Xiang and Shuwen Xiang. Notes on completely positive matrices. Linear Algebra and its Applications, 271(1–3):273–282, March 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=271&issue=1-
                                        3&aid=9700278.

Yamamoto:1994:FRS

[YA94] Yutaka Yamamoto and Mituhiko Araki. Frequency responses for sampled-data systems-their equivalence and relationships. Linear Algebra and its Applications, 205/206: 1319–1339, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:1992:NES

[Yan92] Shangjun Yang. A note on the exponent set of primitive minimally strong digraphs. Linear Algebra and its Applications, 171:187–??, July 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yan:1993:TET

[Yan93] Wang Yan. Two enumeration theorems in singular symplectic geometry and a class of PBIB designs. Linear Algebra and its Applications, 181: 85–??, March 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:1998:BCP

[Yan98] Shangjun Yang. On block cocyclic pairs of nonnegative matrices. Linear Algebra and its Applications, 282(1–3):131–143, October 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
282/1-3/6168.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=282&issue=
1-3&aid=6168.

Yastrebov:1996:CCA

[Yas96] Alexander Yastrebov. On a construction of the Cliffordean algebra. Linear Algebra and its Applications, 244(1–3):341–356, September 1, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1996&volume=244&issue=1-
3&aid=9400241.

Yaacoby:1992:DDS

[YC92] Yoav Yaacoby and Peter R. Cappello. Decoupling the dimensions of a system of affine recurrence equations. Linear Algebra and its Applications, 167:157–??, April 1, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:1999:PCD

[YC99] Hong You and Sheng Chen. Products of commutators of dilatations. Linear Algebra and its Applications, 291 (1–3):51–61, April 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
291/1-3/6328.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=291&issue=
1-3&aid=6328.

Yu:1997:NLB

[YhG97] Yi-Sheng Yu and Dun he Gu. A note on a lower bound for the smallest singular value. Linear Algebra and its Applications, 253(1–3):25– 38, March 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/

browse/browse.cgi?year=
1997&volume=253&issue=1-
3&aid=9500784.

Yip:1996:NSC

[Yip96] E. L. Yip. A necessary and sufficient condition for M-matrices and its relation to block LU factorization. Linear Algebra and its Applications, 235(1–3):261– 274, March 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=235&issue=1-
3&aid=9400164.

Yiu:1993:MNS

[Yiu93] Paul Yiu. Maximal normal sets of n-planes in R^{2n}. Linear Algebra and its Applications, 181:241–249, March 1, 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yanai:1994:LAL

[YM94] H. Yanai and S.-I Mayekawa.

Linear Algebra and Linear Models, by R. B. Bapat. Linear Algebra and its Applications, 207:273–??, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yong:1999:DES

[Yon99] Xuerong Yong. On the distribution of eigenvalues of a simple undirected graph. Linear Algebra and its Applications, 295(1–3):73–80, July 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
295/1-3/6424.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=295&issue=
1-3&aid=6424.

Ye:1991:CLC

[YP91] Yinyu Ye and Panos M. Pardalos. A class of linear complementarity problems solvable in polynomial time. Linear Algebra and its Applications, 152:3–17, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Interior point methods for linear programming.

Yalamov:1996:SBC

[YP96] Plamen Yalamov and Velisar Pavlov. Stability of the block cyclic reduction. Linear Algebra and its Applications, 249(1–3):341–358, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500392.

Yanai:1992:CCA

[YT92] Haruo Yanai and Yoshio Takane. Canonical correlation analysis with linear constraints. Linear Algebra and its Applications, 176:75– ??, November 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:1995:RBJ

[Yu95] Jie Tai Yu. On relations between Jacobians and minimal polynomials. Linear Algebra and its Applications, 221(1– 3):19–29, May ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=221&issue=1-
3&aid=9300220.

Yang:1990:FOG

[YW90] Ben Fu Yang and Wan Di Wei.

Finite orthogonal geometries with characteristic ≠ 2 and PBIB designs. I. Linear Algebra and its Applications, 134: 1–23, 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yong:1997:EMT

[YW97] Xuerong Yong and Zheng Wang. Elliptic matrices and their eigenpolynomials. Linear Algebra and its Applications, 259(1–3):347–356, July 1, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=259&issue=1-
3&aid=9600310.

Yong:1999:CFM

[YW99] Xuerong Yong and Zheng Wang. On a conjecture of Fiedler and Markham. Linear Algebra and its Applications, 288(1–3):259–267, February 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
288/1-3/6327.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=288&issue=
1-3&aid=6327.

Yang:1995:GRM

[YZ95] Shangjun Yang and Ronghua Zhang. Green’s relations in the matrix semigroup M_{n}(S). Linear Algebra and its Applications, 222(1–3):63– 76, June ??, 1995. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1995&volume=222&issue=1-
3&aid=9300276.

Zaballa:1994:SBS

[Zab94] Ion Zaballa. Similarity and block similarity. Linear Algebra and its Applications, 212/213:461–485, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the 3rd ILAS Conference (Pensacola, FL, 1993).

Zaballa:1999:FIM

[Zab99]  I. Zaballa.  Feedback in-

variants of matrix quadruple completions. Linear Algebra and its Applications, 292 (1–3):73–97, May 01, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
292/1-3/6366.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=292&issue=
1-3&aid=6366.

Zalar:1999:LOP

[Zal99] Bojana Zalar. Linear operators preserving the signreal spectral radius. Linear Algebra and its Applications, 301(1–3):99–112, November 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/115/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/115/25/30/article.
                                        pdf.

Zampieri:1994:SCP

[Zam94] Sandro Zampieri. A solution of the Cauchy problem for multidimensional discrete linear shift-invariant systems. Linear Algebra and its Applications, 202:143–162, April 15, 1994. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zymslony:1992:JAB

[ZD92] Roman Zymslony and Hilmar Drygas. Jordan algebras and Bayesian quadratic estimation of variance components. Linear Algebra and its Applications, 168:259–??, April 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zeng:1993:BPM

[Zen93] Jiang Zeng. A bijective proof of Muir’s identity and the Cauchy-Binet formula. Linear Algebra and its Applications, 184:79–??, April 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:1994:NBT

[ZG94] Xian Zhang and Dun He Gu. A note on A. Brauer’s theorem. Linear Algebra and its Applications, 196:163– 174, January 1, 1994. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:1999:SIM

[ZGL99] T. Zhang, G. H. Golub, and K. H. Law. Subspace iterative methods for eigenvalue problems. Linear Algebra and its Applications, 294(1–3):239–258, June 15, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
294/1-3/6421.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=294&issue=
1-3&aid=6421.

Zhang:1991:MSS

[Zha91]  Mou Cheng Zhang.  On

the maximal subgroup of the semigroup of generalized circulant Boolean matrices. Linear Algebra and its Applications, 151:229–243, 1991. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:1993:RGL

[Zha93] John Zhang. Relative growth of linear iterations and orthogonal polynomials on several intervals. Linear Algebra and its Applications, 186: 97–115, 1993. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zha:1996:NEH

[Zha96] Hongyuan Zha. A note on the existence of the hyperbolic singular value decomposition. Linear Algebra and its Applications, 240(1–3):199– 205, June ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=240&issue=1-
3&aid=9400197.

Zhang:1997:QMQ

[Zha97a]  Fuzhen Zhang.  Quater-

nions and matrices of quaternions. Linear Algebra and its Applications, 251(1–3):21– 57, January 15, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1997&volume=251&issue=1-
3&aid=9500543.

Zhang:1997:LAG

[Zha97b] Zhizheng Zhang. The linear algebra of the generalized Pascal matrix. Linear Algebra and its Applications, 250(1–3): 51–60, January 1, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=

1997&volume=250&issue=1-
3&aid=9500452.

Zhan:1999:NIC

[Zha99] Xingzhi Zhan. Norm inequalities for Cartesian decompositions. Linear Algebra and its Applications, 286(1–3):297–301, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6284.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6284.

Zhou:1996:LDR

[Zho96]  Ding-Xuan Zhou.  Lin-

ear dependence relations in wavelets and tilings. Linear Algebra and its Applications, 249(1–3):311–323, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9500365.

Zhong:1999:CCR

[Zho99] Lipin Zhong. On the contribution of cardinalities of row space of Boolean matrices. Linear Algebra and its Applications, 288(1–3): 187–198, February 1, 1999.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1999/
                                        288/1-3/6301.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1999&volume=288&issue=
                                        1-3&aid=6301.

Ziebur:1993:CRF

[Zie93a] Allen D. Ziebur. The chain rule for functions with a matrix argument. Linear Algebra and its Applications, 194:125– 133, November 15, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zietak:1993:PLA

[Zie93b] K. Zietak. Properties of linear approximations of matrices in the spectral norm. Linear Algebra and its Applications, 183:41–??, April 1, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zietak:1993:SFD

[Zie93c] K. Zietak. Subdifferentials, faces, and dual matrices. Linear Algebra and its Applications, 185:125–141, May 1993. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zietak:1996:APZ

[Zie96] Krystyna Zietak. On approximation problems with zero-trace matrices. Linear Algebra and its Applications, 247(1–3):169–183, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=247&issue=1-
3&aid=9500098.

Ziebur:1998:CRF

[Zie98]  Allen D. Ziebur.  Chain

rules for functions of matrices. Linear Algebra and its Applications, 283(1–3): 87–97, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
283/1-3/6202.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=283&issue=
1-3&aid=6202.

Zizler:1998:NSO

[Ziz98] P. Zizler. Norms of sampling operators. Linear Algebra and its Applications, 277(1–3):291–298, June 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
277/1-3/6071.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/

                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=277&issue=
                                        1-3&aid=6071.

Zhu:1991:UCF

[ZJ91] J. Zhu and C. D. Johnson. Unified canonical forms for matrices over a differential ring. Linear Algebra and its Applications, 147:201–248, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:1998:EGP

[ZL98] Zhizheng Zhang and Maixue Liu. An extension of the generalized Pascal matrix and its algebraic properties. Linear Algebra and its Applications, 271(1–3):169– 177, March 1, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=271&issue=1-
                                        3&aid=9700266.

Zhu:1992:LOD

[ZM92] J. Zhu and C. H. Morales. On linear ordinary differential equations with functionally commutative coefficient matrices. Linear Algebra and its Applications, 170:81–??, June 1992. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:1998:GPS

[ZPJZ98] Yingshan Zhang, Shanqi Pang, Zhengming Jiao, and Wenzheng Zhao. Group partition and systems of orthogonal idempotents. Linear Algebra and its Applications, 278(1–3):249–262, July 15, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1998/
278/1-3/6095.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1998&volume=278&issue=
1-3&aid=6095.

Zaslavsky:1999:JFI

[ZT99] Boris G. Zaslavsky and Bit-Shun Tam. On the Jordan form of an irreducible matrix with eventually nonnegative powers. Linear Algebra and its Applications, 302-303(1–3):303–330, December 1, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/117/25/
43/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/117/25/43/article.
pdf.

Zhang:1998:GPM

[ZW98] Zhizheng Zhang and Tianming Wang. Generalized Pascal matrix and recurrence sequences. Linear Algebra and its Applications, 283(1–3): 289–299, November 1, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/laa/sub/1998/
                                        283/1-3/6227.pdf;   http:
                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/laa/
                                        cas_sub/browse/browse.cgi?
                                        year=1998&volume=283&issue=
                                        1-3&aid=6227.

Zigic:1993:CTA

[ŽWB93] Dragan Žigić, Layne T. Watson, and Christopher Beattie. Contragredient transformations applied to the optimal projection equations. Linear Algebra and its Applications, 188/189:665–676, 1993. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhan:1996:ME

[ZX96] Xingzhi Zhan and Jianjun Xie. On the matrix equation X + A^{T}X^{−1}A = I. Linear Algebra and its Applications, 247(1–3):337–345, November 1, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1996&volume=247&issue=1-
                                        3&aid=9500120.

Zhou:1999:ARP

[ZX99] Guangcai Zhou and Xiang-Gen Xia. Ambiguity resistant polynomial matrices. Linear Algebra and its Applications, 286(1–3):19–35, January 1, 1999. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/laa/sub/1999/
286/1-3/6244.pdf;   http:
//www.elsevier.com/cgi-
bin/cas/tree/store/laa/
cas_sub/browse/browse.cgi?
year=1999&volume=286&issue=
1-3&aid=6244.

Zha:1996:APS

[ZZ96] Hongyuan Zha and Zhenyue Zhang. The Arnoldi process, short recursions, and displacement ranks. Linear Algebra and its Applications, 249(1–3):169–188, December ??, 1996. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/laa/cas_sub/
browse/browse.cgi?year=
1996&volume=249&issue=1-
3&aid=9782725.

Zha:1998:SPJ

[ZZ98] Hongyuan Zha and Zhenyue Zhang. A sorted partial Jacobi method and its convergence analysis. Linear Algebra and its Applications, 270(1–3): 79–108, ???? ??, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/laa/cas_sub/
                                        browse/browse.cgi?year=
                                        1998&volume=270&issue=1-
                                        3&aid=9700227.